A Complete Bibliography of Publications in Future Generation Computer Systems: 2020–2029

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

24 May 2024 Version 3.20

Title word cross-reference

2 [LGW^{+}21, NQH^{+}20, SCK^{+}22, TJG^{+}20]. 3 [BÖ20a, CIJM20, GLF^{+}22, HIdAR^{+}20, HYRZ20, JZZD21, JL21, JPJO22, NMRK21, QJZ^{+}20, VHP^{+}22]. 5 [KSSR20]. 7μ [RSFB23]. ^{2} [KAK^{+}23, KMK^{+}23]. _{2.5} [ZG23]. A^{⋆} [HYC^{+}21]. B [WCWC19, WCWC20]. B^{+} [WX24]. c [MGW23]. Δ [GLM21]. ℓ_{0} [CCL^{+}22]. H^{2} [LDCZ20]. K [BJ22, WGW^{+}20, WGL^{+}24, CQA^{+}24, EGD24, HPY20, LRML21, MBM^{+}20, ŠHDT21, WFL^{+}20, WGS24]. λ [PTZ^{+}20]. m [NQB^{+}23]. N_{1} + N_{2} [BJ22]. π [TLC^{+}20]. Q [SPL24]. w_{ρ}^{3} [See20]. X [Ben23b].

* [BFM23]. *-chain [BFM23].

-architecture [PTZ^{+}20]. -based [RKG20, ETH20]. -causal [GLM21]. -D [LGW^{+}21]. -Hub [TLC^{+}20]. -learning [SPL24]. -mer [EGD24]. -norm [CCL^{+}22]. -opt-based [SCK^{+}22]. -pixel [Ben23b]. -simplex [NQB^{+}23]. -sphere [CIJM20]. -stage [BÖ20a]. -standard [WGL^{+}24]. -tree [WX24]. -vertex [ŠHDT21].

/M [BJ22].

1/Pascal [GDS^{+}20]. 10-year [BPLFRL20]. 100 [SME^{+}21]. 105 [LYYG20b]. 108 [DP21a]. 114 [FGB21a]. 115 [LCLW24]. 129 [ARA^{+}23]. 137C [KF23]. 150 [MTHA24a]. 15th [FIABC^{+}20]. 19 [BCT^{+}21, MMC^{+}23, VCM^{+}21].

2 [PK22]. 2-phase [AOF21]. 2017 [DPG20]. 2020 [Ano20m, Ano20w, Ano20r, Ano20n, Ano20o, Ano20s, Ano20t, Ano20v, Ano20x, Ano20p]. 2021 [Ano21m, Ano21x, Ano21n, Ano21t, Ano21q, Ano21p]. 2022 [Ano22u, Ano22s, Ano22r, Ano22w, Ano22p, Ano22m, Ano22n, Ano22x, Ano22o, Ano22q, Ano22v, Ano22t]. 2023 [Ano23x, Ano23u, Ano23w, Ano23n, Ano23o, Ano23s, Ano23p, Ano23r, Ano23t, Ano23v, Ano23q, Ano23m]. 2024 [Ano24q, Ano24r, Ano24j, Ano24m, Ano24n, Ano24k, Ano24o, Ano24l, Ano24p].

3 [VDSB22]. 3000 [LGZ^{+}24]. 3DM [LMO^{+}22].

4.0 [AEN^{+}23, ASH^{+}23, MDDZ21].

5G [BBTC20, URN^{+}20, AZA23, AT20, LM20, LZCGMVV20, MBJ^{+}20, RNA^{+}22, SCGVP20, SNMWC21, SHB22, TPD^{+}24, URN^{+}20, WLL24, YGE21]. 5G-based [BBTC20]. 5G-oriented [AT20]. 5GTopoNet [SNMWC21].

6G [Dao23, HYC^{+}23, MYL^{+}23, SGDG23, TRB^{+}23a, WYX^{+}23a, WZS^{+}23, WLX^{+}24, XLG^{+}23, ZWC^{+}22]. 6G-Envisioned [ZWC^{+}22].

800 [OMPSPL20]. 800-90B [OMPSPL20]. 85 [ABMESM22]. 86 [ABMM22, LBJ^{+}24]. 88 [GHEB^{+}23]. 89 [ABMMC22].

90 [AB19]. 90B [OMPSPL20]. 91 [Bo20b, DP20c, DP21a, DP21b, ZMZ^{+}20]. 98 [HZX^{+}20, JLC^{+}20, WWP20, WCWC20, YWG^{+}20b, YTQ20a, YTQ20b, wZcZN^{+}20].

ABAC [ASSG22]. ABATA [ELS20]. ABCDM [BNA^{+}21]. ABE [XRHS21, ZZQ21, ZPK^{+}23]. ABI [PJL^{+}24]. ABI-free [PJL^{+}24]. abilities [EKK23]. ability [XLCB20, XZ20]. Abnormal [BSOK^{+}20, CCC^{+}23, HYWY22, KIJ^{+}24, LTB^{+}22]. Abstract [RPF21, ACD^{+}20]. Abstracting [MBGC20, CCML20]. academic [KLW^{+}21]. accelerate [QRS^{+}21]. Accelerated [SLH^{+}20, SGBC^{+}20, JCP^{+}20, JCK24, RKI^{+}23, WYX^{+}23a]. Accelerating [GVCUGF20, MNFQ24, SCA22, SXF22, LJW^{+}20, WGW^{+}20, LZL^{+}23]. Acceleration [GDEBC20, LWJ^{+}23, LDX^{+}23, LGS^{+}23, TBB^{+}23, XWR24]. accelerations [LP21b]. accelerator [HMSA^{+}23, KCJ23, LGW^{+}21, SHF23, TWY^{+}23, TZW^{+}22, WXC^{+}24, YSL^{+}24, YTW^{+}20, LZL^{+}23]. Access [ASSG22, AMR^{+}20, KCR20, KRW^{+}20, MR23b, YYKK20, ASA^{+}20, ATK^{+}22, AAG^{+}20, BBTC20, BKHD20, BBB22, CF21, FLTQ20, FFAW20, HXL^{+}23, KAA^{+}21, KHES21, LYYG20a, LYYG20b, LWNH22, LHW^{+}23, LGCY22, LLY^{+}20, MMZI22, MWK^{+}21, NAT20, OWK^{+}23, QRS^{+}21, SAD24, SCE23, SP24, SXF22, SXW^{+}22, SKX^{+}20, TLMP20, WCXW22, WDSK21, WXZ23, YC22, YhSL^{+}22, YK20b, ZZZ^{+}21a, ZLS^{+}20, ZZT^{+}22, dAdSM^{+}22]. AccessChain [SCE23]. accident [KLA22]. accommodation [LGL^{+}20a]. accompanied [YW21]. according [ORPPG20]. accountability [CCH21, ZBS23]. accounts [RVJMJ^{+}21]. accumulative [YLD^{+}23]. Accuracy [CGM^{+}23, ERL^{+}20, YXS23a, ZGL^{+}23]. accuracy-aware [ZGL^{+}23]. Accurate [WC22b, AOSA20b, CdST^{+}20, CBC^{+}20, CSY^{+}20, Jia21, LLF^{+}23, UYH21, ZLPZ21, ZLWH23]. Achieve [YLSL22a]. Achieving [GAdFGMA21, KRW^{+}20, LZZ^{+}23, YHC20, ZGL^{+}23, ZHJW20, YPL24]. ACMF [SYXL22]. acoustic [GZT^{+}21, ISD22, ZFZS23]. acquire [ADAHA^{+}21]. acquisition [JKS20c, bZSC^{+}23]. Across [WMU^{+}24, CPM^{+}23, GHW^{+}20, HCB^{+}20, HS24, KW20, MJB22, RAL^{+}24]. ACSIMCD [AOF21]. action [BEL20, BH21, DHC23, LSH^{+}20, LSB21, LZC21, MMU^{+}21, NN21, Wan21, Xu21, ZDC22]. actions [GIPS20]. activation [SZO^{+}20, ZRH^{+}23]. Active [BMD^{+}21, WTL^{+}20, XTL^{+}23, BQC22, GMB23, HZT^{+}22, HAH^{+}23, MPC^{+}24, PLHC24]. activities [FHGF20, IA20, MMPL20, RGDMMR^{+}23, SSS21, YWS21]. Activity [GK21, BNX22, ERL^{+}20, EKJ^{+}20, uHA20, HAqDE23, KBG20, LHF^{+}20, LZK21, PKLC23, PLHC24, PABBA20, PYL22, RAS^{+}20, RPdVR20, ZZP^{+}23, ELS20]. activity-based [PABBA20, ELS20]. Actor [SDA21, ZZLF21, KPA24]. actuated [LZZX20]. Actuator [MR23b]. Ad [FPMJ21, UJHN20, ZWX^{+}23, SKX^{+}20]. Ad-Hoc [ZWX^{+}23, FPMJ21, UJHN20]. Adaboost [TTD^{+}20]. adaptability [BYH^{+}20]. adaptable [SGSGGC^{+}23]. Adaptation [LKE22, AKE22, BEON24, FMN^{+}20, HTXW21, PPA^{+}24, PYL22, RJM^{+}21, WYDB24]. adaptation-based [WYDB24]. Adaptative [dAPHOMPJ20]. adapting [SEKS^{+}20]. Adaptive [uRBIBC20, BSM20, BBN^{+}20, BMZdP21, CdD20, DLH^{+}20, ENT^{+}22, EL23, FCGPSG^{+}21, JCK24, KYPJ20, LGYC20, LZL^{+}24c, MGX^{+}23, NGC24, Par20, RZA21, SME^{+}21, SAT20, SMS22, TZW^{+}22, WXC^{+}24, WDS^{+}23, ZFMB20, ZST^{+}20, AAGX^{+}22, ASYL22, AABKB22, AEZ22, BdL20, BCSS20, CWM^{+}20, CS23, DCGM20, DGY^{+}22, FBTJ23, GGK20, JYSH20, JRW^{+}20, JKS20a, KDX^{+}24, LFYH22, LLZ^{+}24a, LBY^{+}20, LZA^{+}20, LDX^{+}23, MZA23, MXS22, MA22, NNH^{+}20, yQhJL20, RKM23, SME^{+}19, SYG^{+}20, SUKN22, SDZ^{+}20, SZW^{+}23, SDV^{+}21, SZGB24, UPD^{+}20, WLZ^{+}20, WZHX23, WZS^{+}22, XCL^{+}20, YRV^{+}22, YLKK20, YGP^{+}24, YLTH22, YYB^{+}21, ZDC22, bZSC^{+}23, ZYW24, uRLW^{+}21, AAG23]. AdaptScale [SDZ^{+}20]. Adding [SNS^{+}20]. additional [WZXX21]. addressable [ISUC22]. addressed [PAS^{+}20]. addresses [PGMP23]. ADeLe [FSD^{+}20]. adherence [PRPPFRL20]. adjustment [LWLH20, MHA^{+}24]. administration [PMMG^{+}20]. admissible [ADAHA^{+}21]. Admission [FPMJ21, WLL24]. adolescents [BAGRB^{+}20]. adopted [SD24]. adoption [MIIN23]. ADS [WXC^{+}24]. ADS-CNN [WXC^{+}24]. Advanced [BOM^{+}22, CA21, EAA21, VKP22, BCB^{+}20, MMK^{+}20, NKG23, QRS^{+}21, SAAEK22, ZCWC20]. advancement [IB20]. Advancements [BGNM20, Dao23, MBJ^{+}20, RGPGP24, CDR24, SD22]. Advances [AMNZ20, BDF^{+}22, ICW21, ZLML20, RCJZ20, SACN^{+}21, WD24, ZXD^{+}20]. Advancing [BBSB21]. Adversarial [Ben23a, MJW23, NCLP21, RAA^{+}24, SCBP24, VMM20b, Ben23b, CSC23, CSS22, DBD^{+}23, DCD^{+}22, HRX^{+}21, HXWX23, JLK22, JWC22, KSH^{+}21, LZW^{+}22, LHLC23, PCK20, SHKW23, WWH^{+}21, WZC^{+}22, YYY^{+}23, ZCQ^{+}23, ZZ21a]. adverse [DFG^{+}21]. advisee [GWY^{+}20]. advisor [GWY^{+}20]. advisor-advisee [GWY^{+}20]. Aerial [HQLH20, LYW23, BCT^{+}21, XWK21, XW21, YLY^{+}23]. aerodynamics [BDGG^{+}20]. aerospace [DP20a, SMS^{+}24]. AES [HIMM20]. Aesthetic [LJ21, JL21]. AFCL [RPF21]. affect [DDM21]. affected [RAA^{+}20]. Affective [SG20, HZPS21, OOB^{+}21]. affinity [PCCX21]. Affordable [DLZ^{+}23]. after [CMF^{+}21]. against [CCL^{+}22, CSC23, DBD^{+}23, Elg20, EUEU24, FGG^{+}23, JLK22, LHXL22, LMW^{+}24, MV21, NMRK21, NCLP21, PCK20, QLJ21, RSQS21, RBMCLH22, VMM20b, WZHX23, YNVRPD23, ZZQ21, ZZG^{+}22]. Age [AKPT20, CLZ^{+}20, KWL^{+}23, POMK20]. aged [RSL24, TSKK23]. Agent [ASM^{+}22, CCW^{+}20a, NAC^{+}22, OÖA22, SGP^{+}20b, ADP^{+}22, CPJ^{+}21, DC21, DHD20, DQBS20, GRG20, LYYG20a, LYYG20b, MRM^{+}24, PR20, SCA22, SP21, WLP^{+}20, WYJ^{+}24, ZZLF21, ZLZ^{+}23b]. Agent-based [CCW^{+}20a, NAC^{+}22, SGP^{+}20b, CPJ^{+}21, DC21, SCA22]. agent-programming [GRG20]. AgentChain [HLZ^{+}22]. agents [AdSM^{+}22, BEL20]. Aggregate [XZZ^{+}20b, HDD20, SAT20]. aggregated [LQS^{+}20]. aggregation [AKF^{+}20, ABB^{+}21, CATD^{+}24, EKJ^{+}20, GJC^{+}20, MYL^{+}23, QCG^{+}24, SYYuR21, TGAP20, WHC^{+}22, WGS24, XGX20, XZZ^{+}20b, XJL^{+}24, XZC^{+}24]. Aggregator [MCT^{+}22, RPP^{+}20]. Aggression [SMU^{+}21, KSDR21]. aggressive [BMZdP21]. agile [SCX21, VZDS24]. agnostic [BLGCLA^{+}23, JLS^{+}23, PTZ^{+}20, TD21]. Agreement [MOU^{+}21, BGCL20, CDG^{+}20, YZW^{+}23]. agricultural [LBJ^{+}18, LBJ^{+}24, RWG21, ZP22, ZGZX21]. agriculture [SD22, TLN23]. ahead [YYXZ23]. AHP [ABMMC22, ABMMC18, WC22a]. AI [Ano24s, BEM^{+}24, CSD^{+}23, EBA^{+}22, HHW^{+}22, HKB^{+}24, JZK^{+}21, JLS^{+}23, KRA21, Kon21, KAJ^{+}24, LLW^{+}20, MLZ^{+}22, MJW^{+}22, QJS^{+}21, RKI^{+}23, SUKN22, TRB^{+}23a, VAKB23, WPX^{+}23, YWG^{+}20a, YXS23a, YZS^{+}21, ZZJC21, ZWZ^{+}23]. AI-accelerated [RKI^{+}23]. AI-assisted [HHW^{+}22, SUKN22]. AI-based [Kon21, YXS23a]. AI-empowered [Ano24s]. AI-enabled [JZK^{+}21, YWG^{+}20a, ZWZ^{+}23]. AI-guided [YZS^{+}21]. aid [ABM19, ABM21]. aided [BKV22, HLP21, LQYL21, LYH^{+}21, PZHD20, SLS^{+}20, SZM22, WX23]. aids [KAF^{+}23]. AIMDP [OCMJFB^{+}23]. AIOps [dATBMA23]. AIoT [WPPA22b, WPPA22a, GYAW22, HLL^{+}24]. air [CPT^{+}20, EJP22, Gur21a, Gur21b, LLT22, ZZG^{+}24, ZZ24]. air-gapped [Gur21a, Gur21b]. aircraft [WLN^{+}21]. AIRE [TWL23]. airplane [BDGG^{+}20]. Airports [MTT^{+}23]. AJ [YL20b]. AKA [BGCL20]. Akka [SDA21]. ALBUS [BHL^{+}21]. alcohol [BAGRB^{+}20]. alert [ABL22]. alerts [VGM24]. AlexNet [HZLH21]. algebra [HS21, MMAH22, TBB^{+}23]. algebraic [TBA23]. Algorithm [AWMM^{+}23, LLZ^{+}22, MSA^{+}24, PLBOC20, ABMESM18, ABMESM22, ASA23, AM20, Ben23b, AAT^{+}24, AJPM20, ANH^{+}21, CVdRA^{+}20, CFL^{+}20, CMM21, CCC^{+}23, DT21, DWL^{+}23, DZXS21, Elg20, EGD24, FWP21, FW22, GZXH24, GB20, GPR23, Gul22, HMSA^{+}23, HLP21, HSS20, HAqDE23, HRGL21, HX21, HY21, HL24, IAM^{+}22, IT20, JQZ^{+}22, JLP^{+}21, Jia21, JR22, JZM^{+}22, KSSR20, KS24, KHL20, KA22, KV22, LZH^{+}20, LCW^{+}20, LTX^{+}24, LDWZ20, LZA^{+}20, LGW^{+}21, LMZ^{+}22, LV24, LHW20, LZHS24, LEXH20, LCC^{+}24b, LCCP21, MDL^{+}23, MSG^{+}20, MISB22, MYM^{+}21, MLZ^{+}23a, MO24b, MDG^{+}22, PJBB20, Pan20, PWH^{+}22, PCACTMÁ24, PCG^{+}20, PFGDM22, QPL22a, yQhJL20, QZZ^{+}24, RK20, RDR^{+}24, RLML20, SCR20, STS^{+}20, SCXZ23, SSV24, SYXW21, SPL22, SDKM20, SVN20b, TDM^{+}22, WGG^{+}20, WGW^{+}21, WFL^{+}21, WZHX23, WSL^{+}23, WCS24, WGS24, Wei21, Wu22, XWM20, XLX^{+}21, YJLC20, YBX^{+}23, YPZ^{+}24, YXYH20, ZWH^{+}20, ZWL20, ZL21]. algorithm [ZLZ21, ZGK^{+}22, ZNX23, ZYW24, Ben23a, ZWL22]. algorithm-based [ASA23, QZZ^{+}24]. algorithmic [AB20]. Algorithms [BP20, BBSB21, CK20, FTM20, ACT24, CLC21a, CKL20, CS24b, DAMS23, DSW^{+}20, GCM21, GPGG23, GW20, HNV^{+}20, JT22, Kha24, LLC^{+}22, LOR22, LWJ^{+}23, LAT^{+}20, MEL^{+}23, NGCB20, RCJZ20, RTD24, ST20a, SMS^{+}24, VG21, VP20, WHW20, WLC^{+}20b, WD24, ZLZ^{+}20a, ZZJC21, ZHL24]. alignment [HMSA^{+}23, JZM^{+}22]. all-in-one [RCLEB20]. alliance [LHW^{+}23]. Allocation [BEM^{+}20, WSD^{+}22, AQN^{+}20, AKA20, BKV22, BKG^{+}20, CMX^{+}20, CWM^{+}20, CLY^{+}20, CWL20, GEN20, GMGV^{+}22, Hu20, HGY^{+}22, HYC^{+}23, HS24, JYP24, KMCJ20, KSMT24, LHC^{+}20, Li20, LYBS21, MGB24, MZX^{+}24, MM23, NGCB20, NPNC23, dSOFC^{+}23, PLS^{+}23, QHCH24, QL22, RNA^{+}22, SKTP24, SS21, SGDG23, WC22b, WSL^{+}23, gWLWZ21, XGS^{+}20, XGY^{+}23, XCSF20, YJF^{+}20, YZS^{+}21, ZWZ^{+}24, ZYX^{+}20, ZLXH20, ZLS^{+}22a, ZWZB24]. allocator [YYZ^{+}24]. allreduce [DCD^{+}24]. ALOHA [PZHD20]. ALOHA-based [PZHD20]. along [RGPGP24]. alpha [WLYL20]. ALPS [ZGL^{+}23]. alternating [CFL^{+}20, GBP23, ZLW^{+}24]. alternatives [CHKJ20]. Alzheimer [ESSS^{+}21]. am [CC21, AAG23]. AM-IF [AAG23]. ambience [AM21]. Ambient [GK21, SMY20, LHF^{+}20]. ambulatory [XLS^{+}21]. ameliorated [Kha24]. America [GDCGVG20]. among [CMF^{+}21, SZS^{+}21]. amplified [GdOAO20]. analyses [LHC21]. analysing [CFC^{+}20]. Analysis [CKL20, DKG^{+}22, JLP^{+}21, LQ20, MMR23a, PRPPFRL20, PFP^{+}22, TQC20, VPSC^{+}23, WMU^{+}24, WCWC20, ZWL21, ZZT^{+}22, AGdS^{+}21, ASL22, ASPG^{+}21, AM22, AIM23, AR20, BJ22, BSH^{+}24, BCCS20, BJP^{+}20, BNA^{+}21, BMBE20, BLMT20, BAGRB^{+}20, BHL^{+}20, Bo19, Bo20b, BDFR22, CLC21a, CDG^{+}20, CL21, CKZ^{+}22, DLdAR23, DNNG21, DVV^{+}20, DSW^{+}20, DK24, FIABC^{+}20, GWZ20, GFZ21, GDCGVG20, GDCPVG22, GMP20b, GHG^{+}21, GBP23, HTLM21, HZX^{+}19, HZX^{+}20, HOV20, JMA^{+}21, KAA^{+}21, KMK^{+}23, KBTM21, KZG^{+}22, KSE24, LZL^{+}20, LLW^{+}22a, LPS^{+}24, LZS^{+}22, zLsZjX20, LZJ^{+}20, LC20, MMP^{+}23, MZA23, MHH^{+}20, MMAH22, MBC22, MMC22, MNSL22, MS20, NRMI20, OMPSPL20, OCBO20, PWV^{+}21, PMMSE21, PSS^{+}23, PABBA20, PMT22, PS20, PP20, RGRV^{+}20, RLZW21, RHM20, SD24, SVFdA20, SBD^{+}24, SBF^{+}21, SUKN22, SuRMA^{+}23, SPL24, SDV^{+}21, SCC20, TWY^{+}23, TA23, TAM21, TA21]. analysis [TIA21, UCR21, UCO20, UAS^{+}20, VMM^{+}20a, VI21, VPA20, VCM^{+}21, WMU^{+}23, WGW^{+}20, WSL21, WCWC19, XCB^{+}20, XW21, YWS21, YK20a, YPX^{+}20, YJF^{+}20, ZMZ^{+}19, ZPLQ20, ZMZ^{+}20, ZZZ^{+}23, ZMJ^{+}22, ZHL24]. Analytical [ACG^{+}20b, DAMS23, dATBMA23, GPWL20, KPL22, YGE21]. Analytics [BP20, BOM^{+}22, DP20c, DP21a, DP21b, NGC24, Sha20, UUK^{+}21, CLC21b, CFÁA^{+}20, DMPS23, DP19, DZB23, EBA^{+}22, HBH21, HAVK22, IDM^{+}20, JKS20c, LCL^{+}20, MCBGSL24, MAC^{+}21, MHA^{+}24, MAA22, NFK^{+}20, QCY^{+}21, RBLD21, RGP^{+}22, SAT20, SKH20, SVN20b, TLW^{+}24, TCBF24, WKW^{+}22, YLL22, uRLW^{+}21]. analytics-based [YLL22]. analyze [CCP^{+}20, LOR22, Yan21]. Analyzer [ELS20, VMM20b]. Analyzing [GMF^{+}20, RPdVR20, WYG^{+}20, WW20, YDL^{+}20]. anatomy [YCG^{+}20]. anchor [SZO^{+}20, ZLWH23]. AND/OR [AB20]. Android [ATT^{+}20, DFF^{+}23, IuRJ^{+}21, KZG^{+}22, LWJ^{+}21, TGJ^{+}20, TLJ^{+}22]. Android-based [DFF^{+}23]. Aneka [TSB20]. Angiographic [WZC^{+}22]. angiography [ZZZ^{+}23]. angle [WSL21, Wu22]. animal [GAP24]. ankle [Bo19, Bo20b]. ANN [BRM^{+}20]. ANN-assisted [BRM^{+}20]. ANNA [LZL^{+}23]. annealing [BPUW24, OPOG23]. annealing-driven [BPUW24]. annotation [XXY^{+}23]. annotations [GSDGP21]. ANNProof [LWY^{+}24]. anomalies [MJSW21, OIG24, RGRV^{+}20, SXC^{+}23]. anomalous [WLC23]. Anomaly [GS20, MKC^{+}21, AEN^{+}23, ADP^{+}22, CCC^{+}21, FLF^{+}21, GKB^{+}20, GLP^{+}24, GZZG24, HN23, LS23a, LWW^{+}22, LCH^{+}23, LLD^{+}21, LGJ^{+}23, MRS^{+}22, MBC^{+}23, RCdF^{+}21, SDVC22, SHKW23, SDV^{+}21, UUH^{+}22, XCL^{+}20, ZCQ^{+}23]. Anonymous [ABC^{+}24, GCH^{+}22, NAT20, GZG20]. answering [XZ20]. Ant [IMuI^{+}21, PAC^{+}22, Ski20, SSB^{+}20, bHFF^{+}21, YXYH20]. Anti [WSD^{+}22, AKCP21]. anti-debugging [AKCP21]. Anti-Malicious [WSD^{+}22]. anti-virtualization [AKCP21]. antimicrobial [MASRAM^{+}22]. any [PFGDM22]. AOFAS [Bo20b, Bo19]. AoI [HGWC23]. AoI-aware [HGWC23]. Apache [LPSV22, RCP24, WGW^{+}20]. APapo [LJ24]. API [AAS^{+}20, CPH^{+}22, PMMG^{+}20]. API-based [PMMG^{+}20]. APIs [XWG^{+}21]. APK [MRMM20]. Apollon [LZB20]. app [YLL22]. appears [WM21]. Application [BBM^{+}22, BSM20, BDGG^{+}20, JLW^{+}23, LGL^{+}23, MIMS20, MBZ^{+}21, MJB22, PCACTMÁ24, PRF20, PVA^{+}20, SQ22, TSKK23, ASPG^{+}21, ADdMM20, ABGDT23, BLH^{+}24, BGR20, BBB22, BRM^{+}20, CG21, CDP20a, ÇYZZ21, CS24b, DHC23, Deh20, ELS20, GMB23, HAB^{+}20, KAK^{+}23, LLW^{+}20, LZ20a, LZH^{+}20, LFZJ21, LLW^{+}23b, MKB23, MDT^{+}20, MAB^{+}20, MBD^{+}20, MMH^{+}22, RLL^{+}22, SMBB^{+}24, SMC^{+}20, XZTC22, YC22, YBX^{+}23, Zha20, ZL21, ZST^{+}20]. Application-aware [LGL^{+}23]. application-defined [MAB^{+}20]. Application-driven [MIMS20]. Application-Level [PRF20]. application-specific [Deh20, HAB^{+}20]. Applications [CHS^{+}24, LXL^{+}21, LH21, PFP^{+}22, WPPA22a, YMY21, ZTP20, ZLML20, AAA20, AHMW23, AK20, ADRP23, AAM^{+}24, AFL23, ARA^{+}22, ARA^{+}23, BKHD20, CLC21b, CS24a, CdO20, DCZ20, DGT24, DML23, FTS^{+}24, GKB^{+}20, GRN20, GHEB^{+}18, GHEB^{+}23, HBK20, HLW^{+}23b, HS24, IDM^{+}20, JYSH20, JAS^{+}20, JLS^{+}23, KPL22, Kol22, KA24, KZG^{+}22, KSLC21, LHC21, LXH^{+}21, LZL^{+}23, LPS^{+}24, LLWJ24, LBDP23, LZC21, LWHW22, Liu24, LGM^{+}20, MHH^{+}20, MBF^{+}20, MFE^{+}20, MEC^{+}20, NLS23, OGO^{+}20, PBM^{+}22, PKB22, PSAL20, PKLC23, PPSC23, PLS^{+}23, PBC^{+}22, PPA^{+}24, RNA^{+}22, RHK^{+}23, RDR^{+}24, RCP24, dRRCGdC20, RGDMMR^{+}23, SAD24, SHZMA21, SACN^{+}21, SMS^{+}24, SZM^{+}21, SXC^{+}24, SDA21, SHY^{+}21, SLA^{+}23, TWI20, TLS^{+}21, TSB20, VAKB23, VMV20, VEH^{+}23, WLD^{+}20b, WWW^{+}24, XCGZ24, YTW^{+}20, YAZ^{+}20, YZS^{+}21, ZWH^{+}20, ZXY^{+}21, ZLL^{+}24b, ZYL^{+}22, ZLG^{+}24, uRLW^{+}21]. applied [ACG^{+}20b, MMAH22]. Applying [CLLCK20, ŠHDT21]. Approach [ASAM20, CMM^{+}23, FZ20, SK21a, VSV^{+}23, ARbL^{+}20, AMM^{+}20, AAGX^{+}22, AGYS20, AEN^{+}23, AAP21, AAT^{+}24, ACC20, ABL22, APC^{+}20, ATC^{+}24, ADAR22, ADAHA^{+}21, BK20, BDG23, BS20, BMM^{+}24, CPPY24, CLV24, CLZ^{+}20, CPT^{+}20, CLY^{+}20, CZZ^{+}23a, CZH^{+}24, CCL23, CS24a, CdO20, CDRS20, CHS^{+}24, dFCC23, DDMP^{+}23, DSW^{+}20, DCD^{+}22, Dut22, FPH^{+}21, GS20, GEN20, GDCGCPVG21, GMM22, GPRM21, GRN20, GHEB^{+}18, GHEB^{+}23, GGK20, GFM^{+}20, GDP20, GLW^{+}20, GMA^{+}22, GK21, HBEK20, HAK^{+}21, HOMD21, HLW^{+}23b, HHW^{+}22, HRY^{+}21, HHLZ20, ICBB20, JYSH20, JTGH21, JZK^{+}21, JVH^{+}20, Kha24, KSA^{+}20, KTIB22, Kho21b, KNV20, KLA22, LZHL23, LQW^{+}20, LWS^{+}23a, LLZ^{+}24a, LDM^{+}21, LLZ20, LDCZ20, LLT22, LSGA20, MZX^{+}24, MTT^{+}23, MK22, MBZ^{+}21, MJB22, MTA^{+}22, MFE^{+}20, MMBD20, MBGC20, MEC^{+}20, MDC^{+}24, MJZC21, MAQ^{+}20, NMRK21, NTY^{+}21, NKB^{+}20, NNUV20, NVS^{+}22, NCLP21, ORLV20]. approach [PB23, POBK21, QJS^{+}21, RMC20, RBSK23, RSQS21, RAS^{+}20, RPP^{+}20, RBW20, SMC23, SKB20, SGDK^{+}21, SZVVB^{+}23, SCÁB20, SEL^{+}22, SYHX23, SXW^{+}22, SXC^{+}23, SPL24, TOM^{+}20, TAM21, UPK^{+}23, Wan20, WLL24, WDSK21, XFJ^{+}20, XLL20a, XLZ^{+}22, YLM23, YYY^{+}23, YNK^{+}20, ZGW^{+}23b, uRKI^{+}21]. Approaches [DPN^{+}22, TSX^{+}24, KIJ^{+}24, LV24, OLP23, PDJS22]. Approaching [LZCGMVV20]. Approximate [BBSB21, DGK20, FSD^{+}20, FFAW20, HMA^{+}21, LWY^{+}24]. approximation [WYG^{+}20, YGP^{+}24]. apps [BSH^{+}24, BHL^{+}20, MSY20]. April [Ano21u, Ano20m, Ano22u, Ano23x, Ano24q]. Arabic [ALS21a, AJJ^{+}21, OCBO20]. architectural [ZZ21b]. Architecturally [Cha20]. Architecture [BBB22, CH24, JCW^{+}23, LBDP23, MSV^{+}20, SRP20, ACF^{+}21, AAA20, AK20, AM21, ACD^{+}20, BCC^{+}22, BÖ20a, BBM^{+}22, DJP^{+}24, BDGG^{+}20, CF24, CdRRdC^{+}24, CFC^{+}20, Deh20, EEN^{+}24, dRFRB24, GHB^{+}24, HSGY20, HLK^{+}23, KHH21, LBJ^{+}18, LBJ^{+}24, LZC^{+}23a, LZS^{+}24, LJW^{+}20, Liu21, zLsZjX20, MDDZ21, MKK^{+}24, MRM^{+}24, MMH^{+}22, OLLP24, PGSM^{+}24, PTZ^{+}20, QNM24, QHE^{+}20, QLHLB23, RSLL21, RCR21, RAS^{+}22, RHJ20, SIG24, SVN^{+}20a, SKH20, TK24b, VG21, VMM^{+}20a, VHP^{+}22, WXD^{+}23, WT24, XGS^{+}20, YCS^{+}20, ZFMB20, ZCQ^{+}23, Zha21, ZBS23, STH^{+}20]. Architectures [BBSB21, KKT^{+}23, CIS^{+}20, GMFC23, GOA23, KCJ23, LGM^{+}21, THA^{+}24, VZDS24, uRLW^{+}21]. archival [SGP^{+}20a]. ARD [FBTJ23]. ARD-Stream [FBTJ23]. Area [WHF^{+}20, HAB^{+}20, MAA22, bZSC^{+}23]. areas [RAA^{+}20, WCP23]. ARFED [IPPK23]. Argumentation [FGP23, GK21]. ARIES [BDM^{+}20]. ARM [LVLBB^{+}24, ZY20, MGGG^{+}20]. array [KSC20, TWY^{+}23, YSL^{+}24]. arrival [CL20a]. art [HHH22, IDM^{+}20, LZZX20, MZL^{+}22, MEL^{+}23, SGP^{+}20b, WXD^{+}23, ZWW^{+}23]. artery [ZZZ^{+}23]. arthritis [YTQ19, YTQ20a, YTQ20b]. arthroscopic [Bo19, Bo20b]. article [LZL^{+}20]. artifact [ODET21, XLS^{+}21]. artifact-centric [ODET21]. Artificial [BCT^{+}21, Elg20, EKK23, HIU^{+}22, IuRJ^{+}21, LCH^{+}23, LEWC24, LWHW22, LCLA21, OCMJFB^{+}23, Sha20, SRP20, UUH^{+}22, WPPA22b, WPPA22a, XWG^{+}21, ZZD22, AGdS^{+}21, BSH^{+}24, DSW^{+}20, GSSB24, MSLJ20, PWH^{+}22, RAA^{+}20, ZLZ21, WYHM21, WPPA22b, WPPA22a]. Artificial-Intelligence-Powered [WPPA22b]. Artist [ZWW^{+}23]. arts [VP20]. ASiMOV [CG21]. ask [CC21]. aspect [GDCGVG20, YPX^{+}20]. aspect-based [GDCGVG20]. aspects [TIA21]. ASR [MCV23]. assemble [LAS20]. assembly [DCGM20]. Assessing [GFPB23, MMFAB23, VCM^{+}21, GBC^{+}24, RJM^{+}21]. assessment [AB19, ABGMC19, AB21, ABGMC21, hAS24, BAGRB^{+}20, BOD^{+}24, CHJ^{+}20, GKA^{+}21, HZLH21, JZZD21, KNRI21, LOH^{+}23, LWZ^{+}20, LJ21, LLZ^{+}24b, MM21a, MBD^{+}20, PVA^{+}20, QC21, Zhu20, ZLWH23, GVŠ22]. assessment-based [LLZ^{+}24b]. asset [PMMG^{+}20]. assets [HMY^{+}23, WHJ20]. assignment [ACG^{+}20a, GCT^{+}20, GZF^{+}20b, Kha24, MA24, YHW^{+}20]. assimilating [ASA23]. assistants [ICGGAR22, VPSC^{+}23]. Assisted [GK21, AZA23, AQN^{+}20, AKF^{+}20, BKHD20, BRM^{+}20, CDG^{+}20, CHC^{+}20, CHS^{+}23, GZ22, GLF^{+}22, HN22, HXL^{+}23, HHW^{+}22, JYP24, JSP23, KMK^{+}23, LBGL20, LLT22, LMW^{+}24, MISS22, MR23b, NBJ21, NAT20, PLL^{+}24, PPSC23, QZZ^{+}24, RK20, RBSK23, SNM^{+}20, SUKN22, UUH^{+}22, WLAC20, WSL^{+}23, WLL24, WYJ^{+}24, XLCB20, ZA20, ZZQ21, ZGK^{+}22, ZGL^{+}23, ZLS23, ZSL^{+}23a, ZHLL24]. assisting [LSGA20]. associated [CHG^{+}20, YGS^{+}22]. association [LMNC22, SZW^{+}23]. associative [GCN^{+}24, TM20]. assumption [LYGF21]. Assurance [BOD^{+}24, AAB23, JRW^{+}20, MGC23]. astronomical [WYX^{+}23b]. Astrophysical [CIB^{+}20]. Asymmetric [HRX^{+}21, ZZXH20]. asynchronicity [AND24]. Asynchronous [NLS23, OÖA22, AHMW23, CMM21, LJ24, LZL^{+}24c, LL20, MA22, XJL^{+}24]. ATDNNS [MXS22]. ATHENA [dRFRB24]. atomistic [MBB^{+}20]. atrial [NED^{+}20, Wan20]. atrium [YCG^{+}20]. Attack [AWMM^{+}23, IPPK23, AZA23, ACG^{+}20a, BSF^{+}20, CYWS24, CZZ^{+}23b, DHC23, LHXL22, LWL23a, LWW24, TTD^{+}20, THB23, TSM24, VP20, WGG^{+}20, WML^{+}21, WLL22, ZZG^{+}22, ZWX^{+}23, ZZZX22, SN21]. attack-defense [WGG^{+}20]. Attack-Resistant [IPPK23]. attacking [JLW^{+}23]. attacks [AGYS20, ASASA^{+}20, Ben23b, Ben23a, BeKTK^{+}20, CKL20, CCL^{+}22, CSC23, DG21, DBD^{+}23, Elg20, EUEU24, FWY^{+}22, FD21, HMH^{+}22, HLL^{+}20, HXWX23, JLK22, Kho21b, KTC23, LWS^{+}23a, LZF^{+}24, MHH^{+}20, MLWA20, NMRK21, NCLP21, PCK20, PDT21, QHW^{+}20, RSQS21, RBMCLH22, SCBP24, STS^{+}20, SAAEK22, SRM^{+}23, TCW^{+}22, WZHX23, XLG^{+}23, YNVRPD23, ZZQ21, ZCWC20]. Attention [ACN^{+}21, BNA^{+}21, GSG^{+}23, HSGX22, SYXL22, UAS^{+}20, XLMC22, YLGG21, ASYL22, ASL22, FWX23, GWZ20, GFZ21, HZL^{+}21, JYSH23, KWL^{+}23, LTXL22, MSKG21, MMU^{+}21, PB23, WSL^{+}23, WLC23, WS23, WCL^{+}24, YCG^{+}20, ZXX^{+}20, ZZZ^{+}22]. Attention-based [BNA^{+}21, UAS^{+}20, XLMC22, WLC23]. attention-empowered [WS23]. attestation [RAB23]. attractive [ZT22b]. Attribute [ASSG22, BKHD20, SZM22, UPK^{+}23, CKV22, LHW^{+}23, MTHA24a, MTHA24b, NAT20, WXZ23, YLGG21, ZZQ21, ZZZ^{+}21a]. Attribute-Based [ASSG22, SZM22, NAT20, ZZZ^{+}21a]. attribute-order-preserving-free [CKV22]. attribution [ATT^{+}20, MTM21]. atypical [MCV23]. Auction [QL22, AKA20, AYY^{+}20, BKV22, HSR^{+}22, MXW22, SZdLZ22, SS21, TLL^{+}24, XGY^{+}23, YLX^{+}23, ZYX^{+}20, ZLS^{+}22a]. auction-based [BKV22, HSR^{+}22, XGY^{+}23]. audiovisual [HZPS21]. Audit [ZBF22]. auditable [GZF^{+}23, Yue20]. auditing [FBL^{+}20, GBK20, SHKW23, XCZ^{+}22, YYW^{+}21, ZWW^{+}20a]. auditor [CFM^{+}22]. augmentation [AADM21, PKLC22, WWH^{+}21, ZL22]. augmentations [YLGG21]. Augmented [MHH^{+}20, SJQ20, WCL^{+}24, ZLL^{+}24a]. augmenting [GKB^{+}20]. August [Ano20w, Ano21m, Ano22s, Ano23u, Ano24r]. Aurora [GMFC23]. Australia [AHN21]. authenticated [CDG^{+}20, LLS24, XWW^{+}20, YLH^{+}23, Yue20]. Authentication [FQH^{+}24, HOV20, ASDLS23, ANA24, BBTC20, BGCL20, CBN^{+}20, DAT21, DLL20, GZG20, HMT^{+}20, HZ20, JZZ^{+}23, KK20, LJL^{+}21, MAS23, RRAB24, TA21, WWY21, WSC^{+}23, YHC20, YZW^{+}23]. author [GDCPVG22]. authorisation [BdL20]. authorities [YYN^{+}20]. Authority [RMA21, CSB23, XRHS21, XZZ^{+}20b]. authorization [CBN^{+}20, MHL20]. authorship [MTM21]. Autism [XLCB20]. auto [CPH^{+}22, CS23, RSLL21, RSQS21, SSWW23, SPL24, TK24a, ZZ24]. auto-encoder [ZZ24]. auto-interaction [CPH^{+}22]. auto-resilient [RSQS21]. auto-scaling [CS23, RSLL21, SSWW23, SPL24, TK24a]. Autoencoder [HMH^{+}22, PLHC24]. Autoencoder-based [HMH^{+}22, PLHC24]. autoencoders [DK24]. AutoMan [CWYG23]. Automata [QHC24, AAG22]. Automated [BCM20, CCdS23, NVS^{+}22, hAS24, ASB^{+}23, ANH^{+}21, CDRS20, DRD20, EEN^{+}24, HZLH21, IA20, KBTT20, LCH^{+}22, LPL^{+}20, MISS22, MSKG21, PSvL^{+}20]. Automatic [CLZ^{+}20, FHGF20, HZL^{+}21, LGC^{+}21, MMP^{+}23, SGS24, TBG^{+}20, dlVGSB^{+}20, CBC^{+}20, CPH^{+}22, DVV^{+}20, DQBS20, FFAFD20, LCFM20, MRD^{+}20, RGP^{+}22, SVN20b, ZZZ21b]. Automating [LAT^{+}20, BFM23]. automation [CPM^{+}23, RCdF^{+}21]. AutoML [FGP23]. automotive [CHL23, MMH^{+}22]. Autonomic [DIB20, ZBTV^{+}20]. Autonomous [Kho21a, KA24, SXC^{+}24, DMSCA20, DP19, DP20c, DP21a, DP21b, GAP24, GMA^{+}22, IHA^{+}20, MKBT24, SZO^{+}20, YNVRPD23]. Autoscaler [PK24]. autoscaling [GMM22, GPR^{+}24]. AutoTrust [ADAR22]. autotuning [PSH^{+}20]. auxiliary [CPYY23]. availability [BLGCLA^{+}23, GMGV^{+}22, LWW^{+}22, MDZ^{+}21, SEL^{+}22]. availability-aware [SEL^{+}22]. available [WYS20]. average [KYY^{+}20]. average-utility [KYY^{+}20]. Averaged [BSF^{+}20]. averaging [IPPK23]. Avoidance [HAA^{+}20, CWM21, CCW^{+}20b, HS21, PZHD20, TLKX21]. avoiding [FLG^{+}20]. AVX [PK22]. AVX-2 [PK22]. Aware [ASAM20, AMR^{+}20, FNRP20, GPC21, HAA^{+}20, KRW^{+}20, MIMS20, ABMO24, ACG^{+}20a, AABKB22, AAP21, AÖ24, AKE22, BRK24, BLMT20, BKG^{+}20, CMX^{+}20, CCL^{+}20, CMM21, CYZ^{+}22, CLL^{+}23, CSAT24, CDV^{+}24, CMGS22, DCGM20, DCC22, DCD^{+}24, DATAA20, Deh20, DMPS23, GCT^{+}20, GAT^{+}20, GLP^{+}24, GB20, GOA23, HBEK20, HCWD21, HSS20, HdOP^{+}21, HRM20, HYRZ20, HQLH20, HGWC23, HKB^{+}24, HWR^{+}22, JTGH21, JCX^{+}21, KF22, KF23, KSH^{+}21, KNV20, KKL^{+}24, Kol22, LHC^{+}20, LWX22, LQW^{+}20, LYYG20a, LYY^{+}20a, LBGL20, LYYG20b, LZK21, LLZL21, LHH^{+}21, LDWZ20, LDDL21, LGL^{+}23, LLG^{+}20, LZA^{+}20, LWZ^{+}20, LZCH22, LEXH20, MAS23, MhCEANSM20, MMZI22, MISB22, MWK^{+}21, MMBD20, MBD21, MAA22, MKBT24, MJW^{+}24, NT22, OLP23, OLLP24, OMSL20, PJBB20, PKB22, PP24, yQhJL20, QHCH24, QCW^{+}24, RNA^{+}22, RMBMT21, RBW20, RKP^{+}21, RFP^{+}24, SEL^{+}22, SHB22, SCP^{+}21, SS22, SZL^{+}21, SHY^{+}21, SCW^{+}22, TWI20]. aware [TLW^{+}24, TSB20, WWH^{+}21, WCL^{+}24, WYG^{+}20, WML^{+}23, XCSF20, XWK21, XJL^{+}24, YWG^{+}20a, YZL^{+}23, YLTH22, YLX^{+}23, YYZ^{+}24, ZBTV^{+}20, ZGL^{+}23, ZHH^{+}23, ZrHhH^{+}23, ZZF^{+}24, ZGN^{+}20, ZLL^{+}23, ZLC^{+}21, ZH20, ZYW24, ZLG^{+}24]. awareness [CZCH24, CMF^{+}21, CSY^{+}20, CDP20a, DBBP24, EELB21, RJM^{+}21, SZVVB^{+}23, ZSZX24]. awareness-based [DBBP24]. AWS [MGZ^{+}20]. AxRAM [FFAW20].

B5G [ASH^{+}23, KKL^{+}24]. BA [SZM22]. BA-RMKABSE [SZM22]. back [GBdRACG20, Gra20, ZDC22]. back-style [ZDC22]. Backbone [MO24b]. backdoor [WZHX23]. Backoff [Par20]. backup [ZZF^{+}24]. backward [NBJ21]. bacterial [DT21, YHC^{+}22]. badminton [FS21]. Bag [NAC^{+}22]. Bag-of-Neural [NAC^{+}22]. Balance [LLP^{+}20b, LQNW20, SVN20b, YWH^{+}23]. balanced [GOA23, bHFF^{+}21, HXWX23, MAS23, MO24b, SHST20, XZK^{+}20, ZDLD24]. balancing [AZA23, ATC^{+}24, BHL^{+}21, BMM^{+}24, DGY^{+}22, DLW^{+}23, GPR23, KF22, KF23, Kha24, KNV20, LLP^{+}20b, LCL22, MYM^{+}21, MOU^{+}21, PR20, SZGB24, YYY^{+}23, ZZL^{+}22, ZCLL22, ZYW24, ZHJW20]. BalCon [GPR23]. bandwidth [ADITS20, CMX^{+}20]. bank [XHW20]. bank-level [XHW20]. banking [LGCY22, WMD^{+}20]. BARA [BKV22]. bark [AHN21]. barrier [CDP20c]. barrierless [LL20]. base [SGDG23]. Based [ASSG22, BKHD20, KRW^{+}20, KWL^{+}23, LCLW24, MAM^{+}24, SME^{+}21, SYXL22, AVK^{+}23, ABMESM18, ABMMC18, ABM19, ABM21, ABMESM22, ABMMC22, AAG22, AYA^{+}23, AZA23, AHSH22, ACF^{+}21, AMM^{+}20, ARIB22, ASA23, ACBT23, ALGMP^{+}21, hAS24, ASYL22, AACJ23, ArMA^{+}21, AM21, AKJJ20, AEN^{+}23, Ben23b, ABL23, ADRP23, ACI^{+}23, AADM21, ACC20, ABL22, AMZZ23, ACM^{+}21, ALS^{+}21b, AJPM20, ADAR22, ANH^{+}21, BZG23, BEM^{+}24, BYR^{+}20, BJ22, BAİP24, BÖ20a, BKV22, BQC23, BCT^{+}21, BNA^{+}21, BBTC20, Ben23a, BSH^{+}21, BeKTK^{+}20, BGNM20, BGCL20, BFM23, BS20, BYH^{+}20, BRM^{+}20, CHS22, CGFC20, CWB^{+}20, CZCH24, CMMST20, CdRRdC^{+}24, CREE^{+}24, CFC^{+}20, CCW^{+}20a, CLZ^{+}20, CCW^{+}20b, CZGS20, CWM^{+}20, CXWY21, CCL^{+}21, CL21, CCHD21, CFM^{+}22, CPH^{+}22, CTFW22, CXHC23, CQS^{+}23, CZZ^{+}23a, CDX^{+}23, CYWS24, CKZ^{+}22, CHL23, CCL23, CZZ^{+}23b, CWL20, CSD^{+}23, CSC23, CPJ^{+}21, CKFT20]. based [CHS^{+}24, ÇYZZ21, DP20a, DDMP^{+}23, DFF^{+}23, DG21, DHC23, DT21, DAM^{+}21, DCC22, DVV^{+}20, DBD^{+}23, DAT21, DMSCA20, DSPSNAHJ20, DC21, DGY^{+}22, DHA^{+}20, DLH^{+}20, DSW^{+}20, DFZ^{+}20, DCD^{+}22, DWL^{+}23, DML23, DQBS20, DLW^{+}23, DBBP24, ERL^{+}20, uHA20, ESSS^{+}21, EELB21, EKK23, EGDT20, EGD24, FLTQ20, FWZ^{+}20, FWP21, FWY^{+}22, FBTJ23, FW22, FFM^{+}20, FIABC^{+}20, GEN20, GSMF20, GWZ20, GWY^{+}20, GCN^{+}24, GCM21, GDCGVG20, GDCGCPVG21, GDCPVG22, GGCIV20, GSDGP21, GPR^{+}24, GPRM21, GRN20, GAT^{+}20, GNA^{+}21, GFM^{+}20, GAdFGMA21, GCH^{+}22, GLP^{+}24, GZXH24, GMGV^{+}22, GMAL23, GKA^{+}21, GMF^{+}20, GK21, GOA23, GZG20, GlRpG20, GTG^{+}21, GPC21, HLW^{+}23a, HWH^{+}23a, HRX^{+}21, HMH^{+}22, HIU^{+}22, HSGY20, HSR^{+}22, HLW^{+}23b, HBF24, HS21, HHW^{+}22, HJI24, HLL^{+}20, HJW^{+}20, HX21, Hu21, HY21, HZT^{+}22, HZ20, HMLS20, bHFF^{+}21, HGY^{+}22, HWH^{+}23b, HYC^{+}23, HKB^{+}24, HSvB20, IA23, IA24, IPPK23]. based [IHA^{+}20, IT20, JA20, JSV21, JCP^{+}20, JPMR21, JMA^{+}21, JYP24, JZZD21, JQZ^{+}22, JVH^{+}20, JGL^{+}20, JLP^{+}21, Jia21, JL21, JLT^{+}21, JR22, JLW^{+}23, JYSH23, JCW^{+}23, JPW20, JZL^{+}24, JKS20b, JMHB24, JZM^{+}22, KHHV21, KBTT20, KOM^{+}20, KOM^{+}22, Kad20, KSSR20, KSA^{+}20, KAK20, KSH^{+}21, KHH21, KIJ^{+}24, KYPJ20, KYY^{+}20, KHES21, KLW^{+}21, Kon21, KZG^{+}22, KMS20, KPA24, KSS^{+}21, KSLC21, KJ24, KTC23, LM20, LHC21, LMNC22, LCB^{+}20, LSN^{+}20, LBJ^{+}18, LBJ^{+}24, LLW^{+}20, LWW^{+}20, LHY^{+}20a, LDLS20, LHL20, LHF^{+}20, LSH^{+}20, LYFZ20, LP21a, LWLW21, LXL^{+}21, LLCH21, LLW^{+}22a, LWNH22, LWW^{+}22, LHXL22, LCH^{+}23, LWL23b, LQG^{+}23, LWS^{+}23a, LHW^{+}23, LZP23, LTX^{+}24, LCC^{+}24a, LRCL24, LYP^{+}24, LPS^{+}24, LZL^{+}24a, LDGS20, LLZ^{+}21, LFYH22, LGCY22, LLZ^{+}24a, LXZ^{+}20, LZS^{+}24, LBDP23, LDM^{+}21, LZA^{+}20, LZ20b, LLZ20, LDCZ20, LLY^{+}20, LJW^{+}20, LJ21, LYH^{+}21, LZW21, LGW^{+}21, LLD^{+}21, LYGF21, LWJ^{+}21, LLZ^{+}22, LMZ^{+}22, LLW^{+}22b, LWF^{+}23, LLF^{+}23, Liu23, LCO^{+}23, LGJ^{+}23, LMW^{+}24]. based [LLZ^{+}24b, LLLS24, Liu24, LZL^{+}24b, LDD^{+}22, LMCSE20, dTGC20, LV24, LZZ^{+}20, LZJ^{+}20, LH20, LGL^{+}20b, LZHS24, LC20, LGT^{+}20, LCLA21, LCLW21, MHL20, MSLJ20, MXL^{+}20, MLC^{+}20, MSZ^{+}20, MYT^{+}21, MXW22, MGW23, MLN24, MSLP24, MMFAB23, MMAH22, MTHA24a, MTHA24b, MSG^{+}20, MK20, MK21, MGGG^{+}20, MLP^{+}21, MGC23, MRS^{+}22, MBZ^{+}21, MISS22, MRMM20, MABK24, MSKG21, MGS21, MLWA20, MXS22, MJW23, MÖ22, MWL^{+}20, MYM^{+}21, MNSL22, MWS24, MM23, MRD^{+}20, MSA^{+}24, MNA^{+}23, MJZC21, MKK^{+}20, MMU^{+}21, NGCB20, NADY20, NTY^{+}21, NAC^{+}22, NRMI20, NAT20, NLS23, NPNC23, NFK^{+}20, NJB20, NN21, NIB^{+}21, OLP23, OLLP24, OMSL20, dSOFC^{+}23, OOB^{+}21, OPLB24, OMPSPL20, PBM^{+}22, PKR21, PKB22, PKLC22, PKLC23, PJLL23, PLHC24, PUMN^{+}24, PB23, PZHD20, PBL^{+}23, PCACTMÁ24, PPG^{+}20, PABBA20, PNL^{+}21, Pła24, PMMG^{+}20, PYL22, PS20, PR20, PBSS24, PTZ^{+}20, QNM24, QMCX20]. based [QHNL21, QPL22a, QGH^{+}22, yQhJL20, QWR^{+}20, QSZ^{+}24, QCWY23, QHC24, QZZ^{+}24, QCY^{+}21, RK20, RAN^{+}20, RBSK23, RCLEB20, RAA^{+}24, RRAB24, RAS^{+}20, RAA^{+}20, RJA^{+}22, RPP^{+}20, RLQ^{+}21, RLL^{+}22, RHWY23, RKG20, RGDMMR^{+}23, RPdVR20, RAB23, RWJ^{+}20, RBA^{+}22, SKB20, SME^{+}19, ST20a, SAM^{+}24, SW22, SYG^{+}20, SGDK^{+}21, SZVVB^{+}23, STK20, SCBP24, SCÁB20, SEL^{+}22, SCA22, SVFdA20, SGP^{+}20b, SP24, SP23, SHST20, SUKN22, ST20b, SJQ20, SP22, SSWW23, SCZ^{+}20, SWW^{+}20, SQ22, SZW^{+}23, SuRMA^{+}23, SN21, SMC^{+}20, SSA^{+}23, SK23, SKTP24, SS21, Ski20, SYXW21, SHKW23, SXC^{+}23, SXHD24, SPL24, SFC23, STK23, SDA21, SZM22, SACW23, SZO^{+}20, SCK^{+}22, SLA^{+}23, TGJ^{+}20, TSR^{+}20, TJG^{+}20, TWY^{+}23, TRB^{+}23a, TTD^{+}20, TLJ^{+}22, TCW^{+}22, TLL^{+}24, TKS^{+}23, TSX^{+}24, TWI20, TDMC23, TBH23, TLN23, TSM24, TLKX21, TA23, TQC20, TYR22, TLS^{+}21, TGAP20, TBG^{+}20, TCBF24, TPN^{+}21, UCR21]. based [UPD^{+}20, URN^{+}20, UYH21, UAACH21, UAS^{+}20, VG21, VS20, VPA20, VGL23, VSPM21, VZDS24, WMU^{+}23, WC22a, WMD^{+}20, WHA^{+}20, WGG^{+}20, WZL^{+}20, Wan20, WCHA20, WDG20, WLLF20, WHF^{+}20, WZB^{+}20, WGW^{+}20, WC20, WHZ^{+}20, WLL21, WGW^{+}21, WSXL21, WFL^{+}21, WWL21, WG21, WC22b, WYWS22, WCXW22, WCD^{+}22, WFLL22, WLL22, WDS^{+}23, WLC23, WSC^{+}23, WZW^{+}23, WHF^{+}23, WLZ^{+}23, WYDB24, WGGB24, WWC^{+}24, WYZ^{+}24, WJC^{+}24, WWZ^{+}20, Wei21, WFA20, WLLC20, WLR21, Wu22, WZH^{+}22, WXZ23, WWW^{+}24, XCH^{+}20, XLMC22, XW23, XGY^{+}23, XRHS21, XLS^{+}21, XLG^{+}23, XZD^{+}21, XZTC22, XCL^{+}20, XLZ^{+}22, XCZ^{+}22, XWZM24, XY20, YMAAH22, YC22, YLSL22b, YJH^{+}20, YHW^{+}20, YZC^{+}20, YLZL21, YLGZ21, YVSG22, YLL22, YXS23a, YLM23, YZZ^{+}23, YBX^{+}23, YPL24, YGP^{+}24, YWH^{+}23, YLY^{+}23, YCYO23, YLF^{+}23, YLZ^{+}24, YXYH20, YJB^{+}21, YGS^{+}22, YK20a, YJF^{+}20, YL20b, YDL^{+}20, Yu21, YYL22, YLX^{+}23, YYY^{+}23, YCS^{+}20, YYKK20, YNK^{+}20, YMY21, YNVRPD23, YD21, ZDC22, ZWC^{+}22]. based [ZJL^{+}22, ZWH^{+}20, ZWW^{+}20a, ZWL20, ZJW^{+}20, ZXX^{+}20, ZHD^{+}20, ZZLF21, Zha21, ZZ21a, ZCF21, ZLZ21, ZPQH21, ZZB^{+}22, ZL22, ZP22, ZZZ^{+}22, ZZG^{+}22, ZWZ^{+}23, ZLZ23a, ZNZ^{+}23, ZL23, ZWX^{+}23, ZFZS23, bZSC^{+}23, ZLF^{+}23b, ZG23, ZZG^{+}24, ZLL^{+}24b, ZWZB24, ZZ24, ZXW^{+}20, ZLP^{+}22, ZLST23, ZWL22, ZLL^{+}23, ZDLD24, ZCL24, ZMJ^{+}22, ZZZ^{+}21a, ZWW^{+}20b, ZZZ21b, ZWY^{+}21, ZY21, ZZZX22, ZBS23, ZTB23, ZSL^{+}23b, ZLS^{+}20, Zhu20, ZWWC21, ZYL^{+}22, ZCWC20, ZHJW20, dAdSM^{+}22, CECS20, ELS20, ASDLS23, BQI^{+}20, ETH20, NSJ^{+}24, TPD^{+}24, TK24b]. baseline [BCB^{+}20, LZJ^{+}20]. Basis [BKM^{+}22]. Basketball [WSL21, Liu21]. bat [GB20]. batch [BMBC20, FYHZ24, GJC^{+}20, HLK^{+}23, JJZ^{+}23, yQhJL20, SLS^{+}20, YLSL22a]. batch-stream [JJZ^{+}23]. batched [ZKL^{+}23]. BatchUp [YLSL22a]. battery [YZL^{+}23]. battery-powered [YZL^{+}23]. Bay [MBD^{+}20]. Bayes [GW22]. Bayesian [AEZ22, DG21, HMLS20, KAK^{+}23, RHM20, SZdLZ22, WLZ^{+}20, WGL^{+}24, XLZ^{+}22]. BC [Liu24]. BCFL [ADAHA^{+}21]. BDOLAP [TMT22]. BDOLAP-Bench [TMT22]. BDPM [FYHZ24]. be [PAS^{+}20]. beam [SGBC^{+}20, SWL^{+}20]. beamforming [WLZ^{+}23]. BEAST [HKS23]. beavers [DC21]. Bee [Elg20, PWH^{+}22]. beef [DVV^{+}20]. Bees [Ben23a, Ben23b]. behavior [AJJ^{+}21, BZG23, BSOK^{+}20, BDK^{+}20, CCML20, CZZ^{+}23a, DBC24, JWC22, KOM^{+}22, KK22, LWW^{+}20, LDLS20, MDP24, MS20, ODET21, PSMF21, SZW^{+}23, WYWS22, XLMC22, YRV^{+}22, HKS23]. behavior-based [LDLS20]. Behavioral [ZLPZ21, KHHV21, MLWA20, MTCS22]. behaviors [MTCS22, WW20, YZZ^{+}23, ZGW^{+}23b, ZZT^{+}22]. behaviour [KZF21, MMC22]. behaviours [MBC22]. Belief [VP20, ZZPK21]. Bench [MEL^{+}23, TMT22]. benchmark [HBK20, LVLBB^{+}24, MJW23, PSH^{+}20, TMT22]. Benchmarking [AB24, BPGL21, ORLV20]. Benchmarks [LGM^{+}21]. benefits [Dao23]. Bertrand [KAF^{+}20]. BES [TGAP20]. best [KMK^{+}23, Tau24]. Better [LRML21, CZ20, NLS23]. between [Gur21b, HWQ^{+}20, Par22, PP22, YTQ19, YTQ20a, YTQ20b]. Beyond [BEL20, TMT22, WWZW23]. BFT [FWP21, ZKL^{+}23]. bi [AEN^{+}23, LTX^{+}24, SACW23, MDL^{+}23]. bi-anomaly-based [AEN^{+}23]. bi-level [SACW23]. bi-objective [LTX^{+}24]. Bi-RRT [MDL^{+}23]. bias [GSSB24]. Bibliometric [SN23, GHG^{+}21]. bid [QL22]. bid-rigging [QL22]. bidding [AYY^{+}20]. Bidirectional [BNA^{+}21, CHW^{+}20, WYX^{+}23a]. Bifurcation [ZPLQ20]. Big [DP20c, DP21a, DP21b, EGD24, KSS^{+}20, LLW^{+}22a, MMH^{+}22, Sha20, Sun20, TMT22, UUH^{+}22, WYGP21, ZLS23, AT20, AAG^{+}20, BOL^{+}20, BDG23, CLLCK20, CDF^{+}22, DPN^{+}22, DP19, FPL24, GZF^{+}23, HSvB20, JTGH21, JKS20c, KPL22, LHC21, LYP^{+}24, LDDL21, LEWC24, LWZ^{+}20, LXY21, LQ20, MDT^{+}20, MDDZ21, NFK^{+}20, PCACTMÁ24, RHWY23, RBW20, SGP^{+}20a, SHW24, SW20, SPL24, SLX^{+}24, TSR^{+}20, WFLC22, WXD^{+}23, WLD^{+}20b, WT24, XW23, YJH^{+}20, YGP^{+}24, ZJL^{+}22, ZA20, ZCF21, ZYY^{+}23, AAB23, BP20, CHS22, EET20, IMM^{+}20, IDM^{+}20, MAB^{+}20, PGSM^{+}24, PMCP20, PP22, SGBC^{+}20, SKH20]. big-data [WXD^{+}23]. BigDEC [EGD24]. Bigraph [zLsZjX20, LZZ^{+}20]. BigTrustScheduling [RBW20]. bilateral [LZS^{+}24]. bile [HZX^{+}19, HZX^{+}20]. BiLSTM [GFZ21, WWS23b]. binarized [WCW^{+}21]. Binary [KSDR21, FW22, QLJ21, QZZ^{+}24, TGJ^{+}20, WZH^{+}22]. bio [JYSH20, RSQS21, BGCL20]. Bio-AKA [BGCL20]. bio-inspired [JYSH20, RSQS21]. bioabsorbable [HZL^{+}21]. biochemical [HZX^{+}19, HZX^{+}20]. BioinfoPortal [OGO^{+}20]. Bioinformatics [HBK20, SHH23, OGO^{+}20]. bioinspired [dAPHOMPJ20]. biological [ABB^{+}21, DC21, LXY21, XWM20]. biomarkers [DDMP^{+}23]. biomedical [CKZ^{+}22, LEWC24]. biometric [KK20, KAK^{+}23, WYDB24]. biometrics [KK20]. biosensors [ZHGX20]. bipartite [GEN20]. bit [Pan20]. Bitcoin [MHA^{+}24, WYZ^{+}20, WHJ20, WQHX20, DSPSNAHJ20, PSHW20]. bite [AHN21]. bitwise [DSW^{+}20]. black [CYWS24, LHXL22]. black-box [CYWS24, LHXL22]. BlackBox [CP22]. Blackmailing [SN21]. BLE [KAH^{+}23]. blind [MXW^{+}23]. blobs [MAB^{+}20]. Block [DSW^{+}20, PCCX21, ASA^{+}20, LZ20b, MAS23, MHA^{+}24, RMA^{+}20, XGX20, YZC^{+}20, ZG24]. block-based [YZC^{+}20]. Block-oriented [DSW^{+}20]. Blockchain [AVK^{+}23, AHWB20, BS20, BAR21, CFM^{+}22, CMM^{+}23, CH24, DS23, DMSCA20, FQH^{+}24, FZT^{+}23, GCH^{+}22, GMAL23, HTGW^{+}23, JGL^{+}20, KBTT20, LM20, LGCY22, LZS^{+}24, LTB^{+}22, LMW^{+}24, LLZ^{+}24b, MHL20, MXW22, MYL^{+}23, MÖ24a, MRR^{+}20, SRP20, SZM22, TLX^{+}23, WMD^{+}20, WLN^{+}21, WWZ^{+}20, XZH^{+}23, YYN^{+}20, ZZG^{+}24, ZWY^{+}21, AZA23, AAGX^{+}22, ACBT23, ANA24, BZG23, BYR^{+}20, BTF^{+}21, BAK22, BKV22, BVCH22, BFM23, CXWY21, CXHC23, CZH^{+}24, CQA^{+}24, CATD^{+}24, CCH21, DPN^{+}22, FBL^{+}20, FWP21, BAM^{+}24, FZC^{+}20, FLH^{+}24, GVCC24, GAdFGMA21, GKA^{+}21, Gra20, GLW^{+}20, GHG^{+}21, HSGY20, HYWY22, HZ20, JZZ^{+}23, KOM^{+}20, KOM^{+}22, KAK20, KSLC21, LCB^{+}20, LBJ^{+}18, LBJ^{+}24, LHL20, LJC^{+}20, LLCH21, LDLS22, LLS24, Liu24, LWY^{+}24, MISS22, MFE^{+}20, MHA^{+}24, NNPP23, OdVP20, PBL^{+}23, PSHW20, QNM24, QHNL21, QL22, RWG21, RLQ^{+}21, RHWY23, RRDSAML23, SCE23, SP24, SZdLZ22, SRA^{+}22, SYHX23, TRB^{+}23b, TDS^{+}22a, TDS^{+}22b]. blockchain [TLS^{+}21, WHA^{+}20, WLC^{+}20a, WSC^{+}23, WWC^{+}24, WXZ23, XRZ^{+}24, XCZ^{+}23, YC22, Yue20, ZWH^{+}20, ZZB^{+}22, ZLS^{+}22a, ZBF22, ZWZ^{+}23, ZKL^{+}23, ZWX^{+}23, ZLF^{+}23b, ZHL^{+}23, ZLL^{+}24b, dAdSM^{+}22, MBF^{+}20, NSJ^{+}24, NKG23, SMO^{+}24, SHB22]. blockchain-adaptive [AAGX^{+}22]. Blockchain-aided [SZM22, BKV22].

Blockchain-assisted [LMW^{+}24].

Blockchain-based

[AVK^{+}23, CFM^{+}22, DMSCA20, GMAL23, JGL^{+}20, KBTT20, LM20, LGCY22, LZS^{+}24, MHL20, WMD^{+}20, ZZG^{+}24, ZWY^{+}21, ACBT23, BZG23, BYR^{+}20, CXWY21, CXHC23, FWP21, GAdFGMA21, KOM^{+}20, KOM^{+}22, LLCH21, QHNL21, RHWY23, TLS^{+}21, WHA^{+}20, WSC^{+}23, YC22, ZWZ^{+}23, ZLF^{+}23b, dAdSM^{+}22, NSJ^{+}24]. blockchain-driven [TDS^{+}22a]. Blockchain-enabled [FZT^{+}23, LTB^{+}22, MXW22, SRP20, CZH^{+}24, JZZ^{+}23, MFE^{+}20, ZBF22]. Blockchain-escorted [MYL^{+}23]. Blockchain-secured [DS23]. Blockchain-supported [BAR21]. blockchains [DPLV23, KMS23, MMR23a, MÖ22, NAK^{+}22, SKA^{+}20, WZW^{+}20, XCZ^{+}22]. blocking [ZGN^{+}20]. BlockIoTIntelligence [SRP20]. blocks [FZC^{+}20, PRF22]. Blog [JZZD21]. blogging [BPCM21]. board [BJP^{+}20, Ano20a, Ano20b, Ano20c, Ano20d, Ano20e, Ano20f, Ano20g, Ano20h, Ano20i, Ano20j, Ano20k, Ano20l, Ano21a, Ano21b, Ano21c, Ano21d, Ano21e, Ano21f, Ano21g, Ano21h, Ano21i, Ano21j, Ano21k, Ano21l, Ano22a, Ano22b, Ano22c, Ano22d, Ano22e, Ano22f, Ano22g, Ano22h, Ano22i, Ano22j, Ano22k, Ano22l, Ano23a, Ano23b, Ano23c, Ano23d, Ano23e, Ano23f, Ano23g, Ano23h, Ano23i, Ano23j, Ano23k, Ano23l, Ano24a, Ano24b, Ano24c, Ano24d, Ano24e, Ano24f, Ano24g, Ano24h, Ano24i]. Body [WHF^{+}20, HWH^{+}23a, HAB^{+}20, SG20, bZSC^{+}23]. Boltzmann [LWJ^{+}21, JCP^{+}20]. bone [CLZ^{+}20, YTQ19, YTQ20a, YTQ20b]. bonus [YWDC23]. boost [SJVRS22]. Boosting [GCPM22, LYKK22, LMCSE20]. BOOTABLE [HBK20]. border [ST20b]. Bot [STS^{+}20]. both [ZAH^{+}20]. botnet [AJPM20, GPRM21, GMB23, RAA^{+}24]. botnets [ZLST23]. bottleneck [AÇP22, DZXS21]. bottom [LLZ^{+}21]. Bound [CGMT20, DJP^{+}24, BPUW24, DBSL23, YHC20]. boundary [GLP^{+}24]. bounded [RZIX20]. bowel [WMU^{+}23]. box [CYWS24, GSI22, LHXL22, LLC^{+}23, MAC^{+}21]. BP [ZL21]. BPSI [FLH^{+}24]. brain [HJW^{+}20, KSA^{+}20, MJW23, NAC^{+}22, QJZ^{+}20, SVFdA20, WLZ^{+}20]. Branch [CGMT20, BPUW24, DBSL23]. Branch-and-Bound [CGMT20, DBSL23]. Brazilian [OGO^{+}20]. breach [AHN21]. breaking [WXX^{+}24]. breast [AIM23]. breeding [YLZ^{+}24]. Bridge [HCL^{+}22, LXC^{+}24]. Bridging [GMI22]. Bringing [BQC22, SMT^{+}24]. Broadcast [FZ20, GLM21]. Broker [GSARS20, LCB^{+}23]. Broker-less [GSARS20]. Brokerage [AMR^{+}20]. Brokerage-Aware [AMR^{+}20]. brokers [ASAM20, HBH21]. browsing [WW20]. BS [JJZ^{+}23]. BS-Join [JJZ^{+}23]. BSEIFFS [DS23]. BSELA [CH24]. BSEM [ZLPZ21]. BSFCAN [DATAA20]. BSKM [GMAL23]. BSKM-FC [GMAL23]. BSMD [RHWY23]. BTG [HCL^{+}22]. buddy [YYZ^{+}24]. buddy-like [YYZ^{+}24]. Buffers [DGL^{+}20]. bugs [LOR22]. build [MWK^{+}21]. Building [LLFQ21, LCB^{+}23, LWY^{+}24, WWY^{+}24, ABMM18, ABMM22, BLGCLA^{+}23, BBB^{+}20, DRD20, IHA^{+}20, MJB22, SGSGGC^{+}23]. buildings [AOSA20b, CVdRA^{+}20, SKS22, PGCB23]. bulk [HWQ^{+}20]. Bulletproofing [SLH^{+}24]. Burst [DGL^{+}20]. Bus [CL20a, VCK^{+}20]. Bus-arrival [CL20a]. Business [CMM^{+}23, BKG^{+}20, KMS23, ODET21, PSHW20, QHC24, STK23, WLLC20, WBR20, YDL^{+}20, YFL^{+}24]. bwSlicer [ADITS20]. Byte [HT22, ATT^{+}20]. byte-code [ATT^{+}20]. Byzantine [DEJ20, Gra20, LWS^{+}23a, RBMCLH22].

C [Hu20, LGM^{+}21, SVD^{+}20, WWY^{+}24]. C-RANs [Hu20]. CaaSSET [MZZ20].

CABC [FQH^{+}24]. Cache [ARbL^{+}20, HdOP^{+}21, MIMS20, CWL20, LLZL21, PPGS20, RKM23, SMS22, WFLL22, WZS^{+}23, ZGN^{+}20, IA24, IA23]. Cache-Aware [MIMS20, HdOP^{+}21]. Cache-MAB [IA23]. Cache-MCDM} [IA24]. caches [GA22, GCN^{+}24]. Caching [BEB^{+}20, DHA^{+}20, GlRpG20, HGK20, IA23, IA24, TTTH20, TLW^{+}24, URN^{+}20, WLP^{+}20, WYX^{+}23a, WLS^{+}24, XCH^{+}20, XLG^{+}23, YLKK20, Yu21, ZAH^{+}20]. Cacol [TTTH20]. calculation [LLZ^{+}21, LYGF21, WSL21]. call [XCS^{+}22, ZLPZ21]. caller [AAR^{+}20]. calligraphy [LGC^{+}21]. calls [AAS^{+}20, SZM^{+}21]. CAM [ISUC22, ZLS^{+}20]. camera [SWL^{+}20]. campus [WZT^{+}20]. can [BEL20, RK20]. Canary [DCD^{+}24]. cancelable [KK20]. Cancer [SJD^{+}20, ABM19, ABM21, AIM23, GMH20, HIU^{+}22, JLC^{+}20, LYH^{+}21, LSGA20, QLHLB23, SWL^{+}20, YWG^{+}19, YWG^{+}20b, wZcZN^{+}19, wZcZN^{+}20]. candidate [bHFF^{+}21]. CANDIL [MCBGSL24]. canonical [GGK20]. capability [LGL^{+}20a, ZLW^{+}24]. capacities [TPD^{+}24]. capacity [HYC^{+}23, LCL22, XZD^{+}21]. capital [NFK^{+}20]. capping [KCP23]. CAPre [TQC20]. Capsule [MKK^{+}20, WMU^{+}24, JTGH21, SGL^{+}20a, WMU^{+}23]. capture [PWV^{+}21]. Capturing [GMF^{+}20]. car [GZG20, SSDC22]. Carbon [KXZW23, LXY21, YCS^{+}20]. carcass [DVV^{+}20]. Card [ZJL^{+}22, CDR24, GFM^{+}20, LPL^{+}20, YHC20]. cardiac [PZLL21]. cardinality [LYY^{+}20b, ŠHDT21]. cardiopulmonary [CHJ^{+}20]. cardiovascular [WFA20, ZYX^{+}23]. care [CHKJ20, TSKK23, UYH21, BSH^{+}21]. Carlo [MMAH22, SWL^{+}20, TBH23]. Carry [KCR20]. Cascaded [HIdAR^{+}20]. Case [VSV^{+}23, BZG23, CBC^{+}20, CHC^{+}20, FMN^{+}20, GDCGVG20, GdOAO20, JAAAZB20, LKJN^{+}20, OLLP24, OCMJFB^{+}23, PMMG^{+}20, RYL20, TRB^{+}23a, WG21, ZZD22]. case-based [WG21]. case-study [FMN^{+}20]. Cases [LFM^{+}22, AAM^{+}24, JRW^{+}20, MV21]. CAT [YLGG21]. categorization [KHB23, PGSM^{+}24, ZGZX21]. category [XWK21]. catering [RNA^{+}22]. Causal [ADMG20, GLM21, GSSB24]. cause [SDV^{+}21]. CBC [HIMM20]. CCoDaMiC [DSC20]. Cecoin [QHW^{+}20]. CEEMDAN [WHC^{+}24]. CEEMDAN-RF-LSTM [WHC^{+}24]. cell [JLC^{+}20, LZZ^{+}20, LZCGMVV20, SGDG23, SZO^{+}20, YWG^{+}19, YWG^{+}20b]. cell-edge [SGDG23]. cells [JLC^{+}20, YWG^{+}19, YWG^{+}20b, YGE21]. cellular [Par20, WLN^{+}21, YGE21]. cellular-connected [WLN^{+}21]. Censorship [dVIP24]. Censorship-resistant [dVIP24]. Center [FAA^{+}23, uRBIBC20, dMBPdSC20, CLL^{+}23, MLX23, MC20, MMBD20, RWJ^{+}20, ZZL^{+}22, ZHJW20]. centered [AdSM^{+}22, FGP23]. Centers [WMU^{+}24, ADITS20, ATZP21, AÖ24, APC^{+}20, AMT^{+}21, BJ22, CHS^{+}23, DLH^{+}20, HWQ^{+}20, HWR^{+}22, LDWZ20, LCY^{+}23b, LLT22, MBD21, MOW^{+}20, RFP22, SK20b, WHW20, WF21, XGX20, XZYH22, YZJ^{+}20, ZLW^{+}22]. central [BOM^{+}22]. centralized [LAHN22]. centres [CHKJ20, FCOJFM21, KTIB22, RMC20]. Centric [AMR^{+}20, BSM20, DHA^{+}20, NMR21, ABA24, CDY^{+}20, DML20, Dut22, GCCMK^{+}20, HBH21, HCK20a, KTIB22, Kho21b, MDDZ21, ODET21, QWR^{+}20, RMBMT21, SGDG23, UPK^{+}23, ZHLM20, uRLW^{+}21]. centroid [SWW^{+}20]. century [FIABC^{+}20]. certificate [CA21, DLL20, LCB^{+}20, LMW^{+}24]. certificate-based [LMW^{+}24]. Certificateless [FQH^{+}24, GWW^{+}22, LYP^{+}24]. Cervical [GMH20, LSGA20]. CFD [BDGG^{+}20, PMMSE21, RKI^{+}23, ZGW^{+}23a]. CH [WLAC20]. Chain [BAR21, PCG^{+}20, ABMM18, AB19, ABGMC19, AB21, ABGMC21, ABMM22, BFM23, Dho20, DC21, HLZ^{+}22, KOM^{+}22, LBJ^{+}18, LBJ^{+}24, LHW^{+}23, MFE^{+}20, RLZW21, RKG20, SEL^{+}22, SCE23, XLH^{+}24, XWD20, XQW^{+}24, YLF^{+}23, SLH^{+}24]. chaining [PCC21, WLZ^{+}23, WYJ^{+}24]. chains [AHSH22, AÇP22, DFF^{+}23, LZ20b, MSBAU24, RCR21, YXL^{+}20]. challenge [WFA20]. Challenges [DRC20, ERG^{+}22, BAM^{+}24, GZPZ20, LKJN^{+}20, PAS^{+}20, WMU^{+}23, WPPA22b, WPPA22a, ZXD^{+}20, AAA20, ABC^{+}20, AAM^{+}24, CFK^{+}20, CDR24, DFF21, FD21, HH22, HBSG21, LWHW22, LCO^{+}23, MGB24, MMC22, MKK^{+}20, NHY20, SD24, SGP^{+}20b, SD22, WXD^{+}23]. change [BBD^{+}21, SVD^{+}20]. changeload [BdL20]. changes [BGR20, LYG^{+}24]. changing [GST21, YYXZ23]. channel [GFZ21, Gur21b, HWH^{+}23a, KAA^{+}21, LF21, MSKG21, SAD24, WYX^{+}23b, XLH^{+}24, YZZ^{+}23, YLH^{+}23, ZZXH20]. channels [SMKC20, YL20b]. chaos [SZM^{+}21]. Chaotic [PKR21, KSSR20, WHF^{+}23]. Characterisation [CHG^{+}20]. Characteristics [WCWC19, WCWC20, CCW^{+}20c, LZ21b, LGT^{+}20, ZWL21, ZWL22, ZCWC20]. characterization [AJJ^{+}21, BAGRB^{+}20, DGL^{+}20, LCY^{+}23b, RVJMJ^{+}21, WCP23, ZZT^{+}22]. Characterizing [CMMST20, WWS^{+}23a, WYG^{+}20]. charging [AYY^{+}20, LDGS20]. CHChain [TDS^{+}22b]. Cheaper [YPL24]. Chebyshev [QJS^{+}21]. checking [GL20]. checkpoint [JHK20, MMK^{+}20]. checkpointing [MRD^{+}20]. ChEESE [FAA^{+}23]. cheque [KBTT20]. Chest [LHTSM^{+}23]. Chief [Fae21]. children [XLCB20]. Chinese [GFZ21, LGC^{+}21, LZW21, MYT^{+}21, TDLT20, XFJ^{+}20, XYH^{+}24]. chip [Deh20, TBB^{+}23]. choice [KHB20]. choices [FHGF20]. cholangiocarcinoma [ZMZ^{+}19, ZMZ^{+}20]. cholangiopancreatography [HZX^{+}19, HZX^{+}20]. Choreography [RPF21]. chorusing [GZT^{+}21]. chronic [MSLJ20, TA23, WCWC19, WCWC20]. chunk [ZPS^{+}24]. Chunks [BEB^{+}20]. CIMAR [LLC^{+}22]. cipher [SP24]. ciphers [RMA^{+}20, ZG24]. ciphertext [CSB23]. circle [Jia21]. circulating [JLC^{+}20, YWG^{+}19, YWG^{+}20b]. circulation [LCZB21]. Cities [ALR^{+}20, AAA20, BLH^{+}24, BOL^{+}20, CGFC20, CdO20, FCGPSG^{+}21, FFM^{+}20, IB20, IHA^{+}20, JAAAZB20, JKS20c, KMK^{+}23, KGO^{+}20, LCL^{+}20, LCB^{+}23, OCSCB22, PJBB20, PZHD20, SNM^{+}20, SACN^{+}21, SLS^{+}20, WHZ^{+}20, YWH^{+}21, BCM20]. citizen [DLGW^{+}20, LZB20, ZKD21]. City [LSL^{+}20, LCLW24, ABOS22, BÖ20a, CPS^{+}23, CLV24, CCW^{+}20a, CDF^{+}22, DML23, GNA^{+}21, GSG^{+}23, Kon21, LDLS20, Li20, LLW^{+}22a, LCZB21, LGW22, LCLW21, PSAL20, PBC^{+}22, QG20, RYL20, SKB20, SVN^{+}20a, STS^{+}20, SKH20, SKX^{+}20, TDL^{+}21, XWLC20, YHW^{+}20, Zha20, ZWH^{+}20, ZXW^{+}20]. cityscape [ZZ21b]. civil [CPS^{+}23]. civilians [CMF^{+}21]. CLARA [GMGV^{+}22]. Class [RFd20, ALGMP^{+}21, GLP^{+}24, JTGH21, LFM^{+}22, MKC^{+}21, RZIX20, ZA22, LFM^{+}22]. classes [ERL^{+}20]. classical [CDBD24, SSA^{+}23]. Classification [BM20, XYL^{+}20, ALGMP^{+}21, AGV23, ANH^{+}21, Ben23a, BCM20, CCC^{+}21, Che20, CYWS24, DVV^{+}20, DLR23, DCD^{+}22, DK24, DLW^{+}23, ERL^{+}20, GDCGVG20, GGK20, GMH20, HT22, HIU^{+}22, HLW^{+}23b, HQLH20, JTGH21, KLA22, LSN^{+}20, LFZJ21, LFM^{+}22, LLW^{+}22b, LPT22, LAT^{+}20, LP23, MRMM20, MS24, NKB^{+}20, ORPPG20, PZLL21, QCWY23, SCÁB20, SPWL23, SXC^{+}24, SK21b, TM20, TA23, VMCM^{+}20, VPA20, Wan20, WDL^{+}21, WWH^{+}21, WLC^{+}20b, XTL^{+}23, XLL20a, XWK21, YLG^{+}24, ZXX^{+}20, ZL21, ZZZX22, ZDZ21]. classification-aware [WWH^{+}21]. classified [WCY^{+}21]. Classifier [AYA^{+}23, ALGMP^{+}21, CCHA22, FPH^{+}21, KAH^{+}23, KP22, ZY21]. classifiers [BBB^{+}20, LRML21, RHM20]. classify [HAK^{+}21]. Classifying [SP22, VGM24]. classroom [SG20]. cleaned [MBC22]. clearance [KBTT20]. clickstream [HNV^{+}20]. Client [QCWY23, CCBFI^{+}23, PBS23]. Client-edge-cloud [QCWY23]. clients [SB24]. Clinical [WWP19, WWP20, MASRAM^{+}22]. clique [HPY20]. cliques [ARIB22]. clMF [CFL^{+}20]. CLoG [JPMR21]. Clone [LLZ^{+}22]. cloning [OTMN23]. closed [WFL^{+}20]. closure [MSK^{+}21, WGG^{+}20]. clothing [YWG^{+}20a]. Cloud [ADMG20, ABT20, ACY20, ACDY21, AMBD^{+}20, FPMJ21, FFB20, GZPZ20, GMM22, GBP23, HMO^{+}20, HJW^{+}20, JA20, KCR20, KVCY20, LKE22, MOW^{+}20, MOU^{+}21, MBB^{+}20, PFP^{+}22, PFS^{+}23, PVA^{+}20, SACW23, VGL23, WZB^{+}20, WLX^{+}24, YLZL21, ZTB23, ADITS20, ARB20, ASHO20, AKJJ20, ADdMM20, ATZP21, AdAHK20, ACG^{+}20b, ADAHA^{+}21, BJ22, BRK24, BLGCLA^{+}23, BPC^{+}24, BKHD20, BBF^{+}24, BKG^{+}20, BBB^{+}20, BMZdP21, CMX^{+}20, CK24, CdRRdC^{+}24, CFC^{+}20, CDG^{+}20, CWM^{+}20, CCL^{+}20, CPH^{+}22, CTFW22, CQS^{+}23, CHS^{+}23, CLL^{+}23, CP22, CS23, CS24a, CCBFI^{+}23, CKV22, DPPGCCA23, DLH^{+}20, DFZ^{+}20, DHD20, EAA21, FCOJFM21, FPL24, GEN20, GBK20, GPR^{+}24, GSKS20, GXS22, GZ22, HBEK20, HN22, HSS20, HTXW21, HXL^{+}23, HdOP^{+}21, HRM20, HJI24, HX21, HLL^{+}24, HHLZ20, HCK20a, HWR^{+}22, IT20, JHB22, KF22, KF23, KDX^{+}24, Kha24, KTIB22, Kho21b, KAF^{+}20, KYPJ20]. cloud [KMCJ20, KCKK24, KSMT24, LWX22, LBGL20, Li20, LLFQ21, LCL22, LTX^{+}24, LRCL24, LDGS20, LDWZ20, LFYH22, LLG^{+}20, LDW^{+}21, LPL22, LZCH22, LDX^{+}23, LCO^{+}23, LMW^{+}24, Liu24, LIP^{+}24, LZHS24, MXW22, MK22, MSS^{+}24, MISB22, MGS21, MWL^{+}20, MYM^{+}21, MS24, MAM^{+}24, NBJ21, NAT20, NPNC23, Ngu24, NIB^{+}21, OWK^{+}23, PKR21, PUMN^{+}24, PB23, PMMSE21, PK22, PDJS22, PPA^{+}24, QPL22a, QSZ^{+}24, QCWY23, QLHLB23, RBSK23, RJA^{+}22, RPP^{+}20, RLQ^{+}21, dRRCGdC20, RAL^{+}24, RBW20, RBA^{+}22, SD24, SEKS^{+}20, SLS^{+}20, SQ22, SZdLZ22, SCX21, SMC^{+}20, SS21, SXC^{+}23, SPL24, SHF23, STK23, SOKW^{+}20, SZO^{+}20, SLX^{+}24, TLN23, TDM^{+}22, TC23, UAACH21, VG21, VMM^{+}20a, WGG^{+}20, WZL^{+}20, WC20, WYS20, WZXX21, WGW^{+}21, WCY^{+}21, WKW^{+}22, WWZ^{+}20, WLD^{+}20b, WBR20, XZJ^{+}20, XYH^{+}24, XCGZ24, XZK^{+}20, YJLC20, YWH^{+}23, YSL^{+}22, YK20b, YYW^{+}21, YZJ^{+}20, YXS^{+}23b, ZWW^{+}20a, ZA20, ZZQ21, ZLW^{+}22, ZCLL22, ZLZ23a, ZHLL24]. cloud [ZWZB24, ZDLD24, ZBS23, ZLG^{+}24, ZA22, AALEF20, DATAA20, GFPB23, HCK20b, MGZ^{+}20, NGCB20, RFP^{+}24, VEH^{+}23, ZWZ^{+}24, CP22]. cloud-agnostic [BLGCLA^{+}23]. cloud-aided [SLS^{+}20]. cloud-assisted [CDG^{+}20, GZ22, HN22, LBGL20, NBJ21, NAT20, ZHLL24]. Cloud-based [JA20, YLZL21, CWM^{+}20, Liu24, MGS21, SMC^{+}20, STK23, TLN23, WGW^{+}21, DAT21]. Cloud-BlackBox [CP22]. cloud-computing [FCOJFM21]. Cloud-Edge [FFB20, WLX^{+}24, CFC^{+}20]. Cloud-Edges [ACY20, ACDY21, VMM^{+}20a]. cloud-enabled [CCBFI^{+}23, CKV22]. cloud-fog [EAA21, GEN20]. cloud-integrated [MS24]. Cloud-native [GBP23, VG21]. cloud-of-clouds [LPL22]. cloud-oriented [HRM20]. cloud-RANs [SZO^{+}20]. cloud-to-edge [BBF^{+}24, LIP^{+}24, RAL^{+}24, TC23, RFP^{+}24]. Cloud-to-Thing [VEH^{+}23]. Cloudification [PVA^{+}20]. Cloudlet [SHST20]. cloudlets [GCT^{+}20]. Clouds [SDZ^{+}20, CWB^{+}20, GBM20, GB20, KPL22, LYG^{+}24, LYP^{+}24, LPS^{+}24, LPL22, NT22, PSC^{+}21, RRHA21, SCZ^{+}20, SPL22, TSB20, WLY23, WWW^{+}24, XRHS21, YWG^{+}20a, YYXZ23, ZYX^{+}20]. club [TDLT20]. Cluster [CK24, CHS22, dFCC23, LCH^{+}21, MZX^{+}24, MGGG^{+}20, YSL^{+}24, YPL24]. cluster-based [CHS22]. Clustering [GAT^{+}20, LLZ^{+}22, LGS^{+}23, ASA23, ASYL22, AMBD^{+}20, ANH^{+}21, CIS^{+}20, CLV24, CHL23, CIJM20, DWL^{+}23, FBTJ23, GEN20, GMGV^{+}22, GPWL20, GLWP20, GYAW22, HCWD21, HMH^{+}22, HDN^{+}20, Hu21, HSvB20, IMM^{+}20, IAM^{+}22, JYSH20, JZM^{+}22, KLW^{+}21, LCY^{+}23b, LLZ^{+}24a, LZA^{+}20, LCH^{+}24, MSG^{+}20, PS20, yQhJL20, SGDG23, VGM24, WLR21, XW21, YHW^{+}20, YLKK20, YLZL21, YWS21, ZST^{+}20]. clustering-based [GMGV^{+}22]. Clusters [AMBGS21, BJP^{+}20, CREE^{+}24, DNNG21, DEJ20, FLF^{+}21, HZdLZ20, LWJ^{+}23, QSZ^{+}24, RK20, WCS24, XZTC22, LFYH22]. CMAC [LZS^{+}21]. CNN [AYA^{+}23, BNA^{+}21, GFZ21, HZS^{+}23, JZZD21, JL21, LYH^{+}21, Liu21, MPS21, MMU^{+}21, QLJ21, SK21b, TWY^{+}23, WXC^{+}24, YLTH22, ZXX^{+}20, ZH20]. CNN-based [JZZD21]. CNN-RNN [BNA^{+}21]. CNN2D [RGPGP24]. Co [CLZ21, HSGX22, LTX^{+}24, YLGG21, AdSM^{+}22, BDGG^{+}20, CCW^{+}20c, LZL^{+}23, MMPL20, NTY^{+}21, THB23, TDM^{+}22, VPBE22, WFL^{+}20, DWZ20, SN21]. Co-Attention [HSGX22, YLGG21]. co-design [LZL^{+}23, VPBE22]. Co-evolutionary [LTX^{+}24]. co-execution [BDGG^{+}20]. co-facilitation [AdSM^{+}22]. co-location [THB23]. Co-occurrence [CLZ21, MMPL20, WFL^{+}20]. co-offloading [TDM^{+}22]. Co-Operative [SN21]. co-scheduling [CCW^{+}20c]. CO-STAR [DWZ20]. co-training [NTY^{+}21]. Coalition [LWL23b, HBEK20]. Coarse [ZZB^{+}24]. Coarse-to-Fine [ZZB^{+}24]. coastal [YBC^{+}20]. Code [TQC20, ATT^{+}20, CIB^{+}20, DC21, LHLC23, MVLJ21, MRMM20, MTM21, PRF22, SNS^{+}20, WXZX23, WLYL20, XWR24, YLS21, HB21]. Code-Analysis [TQC20]. CODE-V [HB21]. coded [BWX20, PK22]. codes [CIB^{+}20]. coding [Dho20, HIMM20, MSKG21]. Coefficient [ArMA^{+}21]. coexistence [CDY^{+}20]. coflows [WS23]. cognition [GPWL20, ZLXH20]. Cognitive [Elg20, SKA^{+}20, WLP^{+}20, ZA20, ZLML20, ASA^{+}20, CHG^{+}20, EKK23, GZL^{+}22, KRA21, SP22, TPN^{+}21, UCR21, VAKB23, ZHGX20, ZTC20]. Cognitive-inspired [ZLML20]. coherence [HZL^{+}21]. Coherent [DSC20]. Cohort [SuRMA^{+}23]. Cohort-based [SuRMA^{+}23]. ColabNAS [GLF24]. Cold [LQG^{+}23, FSP^{+}24, Ngu24, RLZW21, ZZD22]. Cold-start [LQG^{+}23, Ngu24]. collaborating [RVJMJ^{+}21]. Collaboration [WMU^{+}24, CXS^{+}22, CPS^{+}23, HZS^{+}23, JYP24, SACW23, WLP^{+}20, YXS^{+}23b, ZWZ^{+}24, ZYL^{+}22]. collaboration-based [SACW23]. Collaborative [GVCC24, GCPM22, KRA21, LHH^{+}21, SYXL22, XWG^{+}21, ZZJC21, BAMR20, BKV^{+}20, CQS^{+}23, CWL20, DWZ20, GOA23, HLL^{+}24, KCKK24, LRCL24, LGYC20, LDX^{+}23, LWZ^{+}23a, MYL^{+}23, MDG^{+}22, MJW^{+}24, QCWY23, RGDMMR^{+}23, SCXZ23, SGLB22, TKS^{+}23, TDM^{+}22, WYHM21, XZZ^{+}20a, ZLF^{+}23a, HSR^{+}22, KAF^{+}23]. collaboratory [YBC^{+}20]. CoLLaRS [HLL^{+}24]. collect [LLG^{+}20]. collection [MLZ^{+}23a, SLX^{+}24, TLN23, WZL^{+}20, WLAC20]. Collective [EL21, ZZZ21b, ACT24, FLF^{+}21]. collectives [AHMW23]. college [WZT^{+}20]. Collision [HAA^{+}20, CWM21]. collocated [DSC24]. collusion [XRZ^{+}24]. collusion-resistant [XRZ^{+}24]. Colony [Elg20, IMuI^{+}21, PAC^{+}22, SSB^{+}20, bHFF^{+}21, PWH^{+}22, YXYH20]. Color [KSSR20, LLW^{+}20, SLH^{+}20]. Colored [YJF^{+}20, YFL^{+}24]. colorization [HWH^{+}23b]. combat [SSS21]. Combating [ZHZS23, KSS^{+}21]. combination [CDX^{+}23, LZP23, LYC^{+}22, RSLL21, RTD24, TIA21]. combinations [AKJJ20]. Combinatorial [AQN^{+}20, MXW22, SS21]. combine [ZZJC21]. combined [AMBD^{+}20, MRD^{+}20, WLZ^{+}20, YXLB20]. Combining [FQH^{+}24, RCLEB20, RAS^{+}22, XW21, YhSL^{+}22]. command [FZT^{+}23]. commerce [RLZW21, YDL^{+}20, Zha21, ZWY^{+}21]. committee [CFM^{+}22]. commodity [RWJ^{+}20, Zhu21]. common [GMI22, HZX^{+}19, HZX^{+}20]. communicate [SM20]. Communication [AMM^{+}22, CA21, KJ24, MJW^{+}22, AAM^{+}24, CMF^{+}21, CKW21, GPC21, KBTM21, LCB^{+}20, LDLS20, LHH^{+}21, LYY^{+}20b, MDC^{+}24, MJW^{+}24, NMR21, Par22, PSAL20, RHK^{+}23, RHJ20, UJHN20, WCD^{+}22, WYX^{+}23a, XLCB20, XLG^{+}23, YYL22, ZNZ^{+}23, ZZG^{+}24, ZGW^{+}23b]. communication-aware [MJW^{+}24]. Communication-efficient [AMM^{+}22, CA21, MJW^{+}22, ZGW^{+}23b]. communications [CDY^{+}20, RAA^{+}20]. communities [ARIB22, AOF21, GZF^{+}23, GMI22, IMuI^{+}21, PLBOC20, POBK21, RVJMJ^{+}21, SSMdS21]. Community [AMR^{+}20, FIABC^{+}20, LXL^{+}21, MBC22, MMPL20, VI21, ABAJ20, BPLFRL20, BR20, HLT^{+}21, LWW^{+}20, MLC^{+}20, WSXL21, WGL^{+}24, WLLC20, YHC^{+}22]. Community-Centric [AMR^{+}20]. commuting [CHKJ20]. CoMP [WLZ^{+}23]. Compact [SZW^{+}23, QNHB22]. comparative [AIM23, CHS22, DK24, LHLC23, MBB24]. Comparing [ASPG^{+}21, MSY20]. Comparison [CGM^{+}23, PRF20, SHH23, LKS^{+}21, MJSW21, See20, VP20]. compatibility [UAACH21]. compatible [HLZ^{+}22]. compensating [LKS^{+}21]. competitions [KAF^{+}20]. compiler [CPJ^{+}21]. complementarities [GMGV^{+}22]. complementary [BZG23]. complete [CDRS20]. completion [NZY^{+}23, WZZD23, WF21]. Complex [OOZ^{+}23, Zhu21, CCML20, CDF^{+}22, IMuI^{+}21, MMAH22, PGHS20, YZZ^{+}23, ZCWC20]. complex-valued [YZZ^{+}23]. complexity [BK20, LCH^{+}22]. compliance [ZBS23]. compliance-based [ZBS23]. Complicated [LZK21, Xu21, LZC21, Liu21, YWS21]. component [MPC^{+}24, SYG^{+}20, SuRMA^{+}23, ZMJ^{+}22]. component-based [SYG^{+}20]. components [AABKB22, GBC^{+}24]. composite [LPS^{+}24, LZC^{+}23b, NVS^{+}22, XYL^{+}20]. composition [AL20, CS24b, EEN^{+}24, KS24, LZCH22, LZZ^{+}20, SMC23, SBMN21, WGLH20, WLY^{+}20, ZGY20]. compositions [BYH^{+}20]. compound [TPN^{+}21]. comprehensive [AUJW22, ABGDT23, HAH^{+}23, KSH^{+}21, KAA^{+}21, QCG^{+}24, SSA^{+}23, SZZY22, ZG24]. compressed [Wu22, bZSC^{+}23]. compressible [WWP19, WWP20]. compression [CPT^{+}20, Dho20, DC21, HIMM20, MMPV22, PBSS24, RLML20, WFL^{+}21, WWY^{+}24, YK20b, ZWCS23, ZHL24]. compressive [LZL^{+}24b]. Comput [AB19, AB21, ABGMC21, ABM21, ABMESM22, ABMM22, ABMMC22, ARA^{+}23, Bo20b, DP20c, DP21a, DP21b, FGB21a, GHEB^{+}23, HZX^{+}20, JLC^{+}20, KF23, LYYG20b, SME^{+}21, WWP20, WCWC20, YWG^{+}20b, YTQ20a, YTQ20b, ZMZ^{+}20, wZcZN^{+}20]. Computation [BMS20, CZH^{+}24, DSC20, ZWH21a, AKA20, BJW22, GZ22, HXL^{+}23, HGWC23, HB21, LWNH22, LGL^{+}23, LDCZ20, LP21b, LWZ^{+}23b, MZA23, MISB22, MWL^{+}20, QCW^{+}24, QCY^{+}21, RDR^{+}24, RSL24, RNRA23, VDSB22, VPBE22, YW21]. Computational [DIB20, IDM^{+}20, RRHA21, ACT24, BBB^{+}24, CDP20b, GCCMK^{+}20, JAS^{+}20, KIM^{+}24, KSA^{+}20, LP21a, MMFAB23, WDSK21, XWR24]. computationally [PSAL20]. Computations [ACY20, ACDY21, CKW21]. Compute [BPC^{+}24, LWZ^{+}23b]. computed [MTHA24a, MTHA24b, SGBC^{+}20]. Computer [Fae21, HLP21, LBJ^{+}24, LCLW24, MTHA24a, Tau23, Tau24, BJP^{+}20, BGNM20, CDP20a, MJW23, SWL^{+}20, YMY21, SUKN22]. Computer-aided [HLP21]. computer-based [BGNM20]. computers [DA22, Gur21a]. Computing [ARbL^{+}20, AAP21, BBSB21, BBB22, BS20, GZPZ20, HGY^{+}22, HB21, LXC^{+}24, LCLW24, MR23b, NBB20, PAM21, QKG20, SN21, SDA21, TDC^{+}20, VPBE22, WZB^{+}20, WLX^{+}24, XLLL20, ZLML20, AALEF20, ASBT20, ANA24, ABT20, ACG^{+}20b, ATC^{+}24, BÖE24, BKV22, BPCM21, BM20, BMM^{+}24, CK24, CRdRR^{+}22, CdRRdC^{+}24, CMGI^{+}23, CLY^{+}20, CCL^{+}20, CQS^{+}23, CWL20, CF20, DLR23, DAT21, DPPGCCA23, DCZ20, DFZ^{+}20, DHD20, ERK^{+}24, uHA20, EAA21, FWP21, FCOJFM21, FTS^{+}24, GVCUGF20, GBK20, GMP^{+}20a, GBH^{+}23, GMAL23, GXS22, GZG20, GZ22, HHH22, HZPS21, HCWD21, HXL^{+}23, HCB^{+}20, HCG^{+}23, HX21, HZS^{+}23, rHZmH^{+}24, HKB^{+}24, IT20, JHB22, JYP24, KHHT21, Kha24, Kho21b, KMCJ20, KCP23, KGO^{+}20, LHC^{+}20, LMO^{+}22, LMNC22, LBGL20, LYFZ20, LLFQ21, LWNH22, LZW^{+}22, LDLS22, LLW^{+}23a, LZL^{+}24a, LGL^{+}23, LYBS21, LZ22, LDX^{+}23, LMZL24, zLsZjX20, LGL^{+}20b, LFHS23, LCLW21, MGB24, MZA23]. computing [MPC^{+}24, MKB23, dAPHOMPJ20, MGS21, MWL^{+}20, MWS24, MEC^{+}20, MBB^{+}20, Ngu24, NUB24, NLSY20, OGO^{+}20, PKB22, PJJ^{+}22, PK24, PNL^{+}21, PDJS22, QPL22a, QLJ21, QCW^{+}24, QL22, QZZ^{+}24, RNA^{+}22, RRD21, RAA^{+}21, RBW20, SJVRS22, SYYuR21, SIG24, SP23, SJQ20, SQ22, SW20, SCX21, SCP^{+}21, SSA^{+}23, SS21, SXW^{+}22, SFC23, SOKW^{+}20, SGL^{+}20b, SCW^{+}22, SGLB22, TDM^{+}22, TK24a, TBG^{+}20, URN^{+}20, VMM^{+}20a, VI21, VGL23, WLP^{+}20, WDG20, WLD^{+}20a, WWY21, WZXX21, WWL21, WCXW22, WYX^{+}23b, WX23, WGGB24, WYJ^{+}24, gWLWZ21, WPX^{+}23, XKK20, XCH^{+}20, XGY^{+}23, XLG^{+}23, XCZ^{+}23, YW21, YLKK20, YJLC20, YhSL^{+}22, YLL22, YZL^{+}23, YGD^{+}21, YCYO23, YK20a, Zha20, ZA20, ZLZ^{+}20a, ZZLF21, ZZQ21, ZNX23, ZZB^{+}24, ZHLL24, ZWZB24, ZBS23, ZTB23, ZLG^{+}24, HSR^{+}22, ZYW24]. computing-based [GZG20, SJQ20]. concentration [ZG23]. Concept [CFK^{+}20, BBB22, MKK^{+}24, RFd20, XCL^{+}20]. Conception [DFF21]. Concepts [RHK^{+}23, LWLW21, MGS21]. Conceptualization [MV21]. concerns [STK23]. Concurrent [AMBGS21, HZdLZ20, LZHL23, CCP^{+}20, CdO20, PMMSE21, WX24, ZGN^{+}20]. Conditional [JCP^{+}20, CCHA22, THB23]. conductance [LHW20]. conducting [DVEE^{+}20]. cone [SGBC^{+}20]. cone-beam [SGBC^{+}20]. CoNet [CLZ21]. conferencing [BSH^{+}24]. Configurable [HMA^{+}21]. configuration [LDD^{+}22, Man20, RMC20]. confirmation [MMR23a, XSW^{+}21]. conflict [MhCEANSM20]. conflict-aware [MhCEANSM20]. congenital [QPL^{+}22b]. Congestion [CCW^{+}20b, DCD^{+}24, Deh20, JLP^{+}21, SYW^{+}23]. Congestion-aware [DCD^{+}24, Deh20]. Connected [MLN24, AOKÖ24, CFK^{+}20, MMH^{+}22, PZHD20, WLN^{+}21]. Connecting [MLZ^{+}22]. Connection [XJL^{+}24]. Connection-density-aware [XJL^{+}24]. connectivity [GZG20, SSV24, ZYX^{+}23]. Conquer [SYYuR21, AAT^{+}24]. conscious [PAM21]. consensus [FWP21, FZC^{+}20, Gra20, WLC^{+}20a, ZWH^{+}20, ZKL^{+}23]. consent [PGCB23]. considering [HZS^{+}23, LCL22, YJLC20, YLM23]. Consistency [ADMG20, ZHZS23]. Consistently [AOSA20b]. consolidation [BRK24, DLH^{+}20, DHD20, GPR23, LYY^{+}20a, LCC^{+}24b, MOW^{+}20, VGL23, YWH^{+}23]. consortium [CATD^{+}24, PBL^{+}23, SJD^{+}20, WZW^{+}20, XCZ^{+}22, Yue20]. constant [BMS20]. constitute [TDLT20]. Constrained [SCR20, ASA^{+}20, BJW22, CS24a, GBM20, HWQ^{+}20, HLK^{+}23, HWR^{+}22, JHB22, LLP^{+}20a, LZL^{+}24a, MMMZ20, MKB23, RKP^{+}21, RZA21, VDSB22, WHC^{+}22, YZL^{+}24, ZG24]. constraint [GMT23]. constraints [BGMK22, HZdLZ20, KAK20, LLY^{+}20, qLhZ20, MDC^{+}24, SBD^{+}24, ŠHDT21, SW20, WLD^{+}20b, WWW^{+}24, YSL^{+}22]. constructed [PCCX21]. Constructing [GYAW22, DNNG21]. construction [Dut22, HMH^{+}22, LZL^{+}20, LDLS22, TSKK23]. consultation [KCY^{+}21]. consume [DML20]. consumption [LZL^{+}24a, MVLJ21, MGGG^{+}20, MOU^{+}21, PCACTMÁ24, SHKW23, TDM^{+}22]. Container [AMBGS21, CREE^{+}24, GMP^{+}20a, HZdLZ20, HOV20, RRD21, SSWW23, VG21, ZCLL22]. container-based [CREE^{+}24, SSWW23, VG21]. containerization [SXZZ23]. containerized [CSH^{+}23, QSZ^{+}24, SZM^{+}21, VAKB23]. Containers [JAS^{+}20, LAHN22, RKP^{+}21, SDGCB^{+}20, WWS^{+}23a]. containing [TLM21]. Content [BSM20, SCZ^{+}20, ABA24, CDY^{+}20, FGG^{+}23, GSMF20, GlRpG20, ISUC22, LGW^{+}21, MMZI22, MS20, PAP^{+}20, QWR^{+}20, SMS22, VCM^{+}21, XFJ^{+}20, YJH^{+}20, ZC22, ZAH^{+}20, ZLS^{+}20]. content-addressable [ISUC22]. Content-based [SCZ^{+}20, LGW^{+}21, YJH^{+}20]. Content-Centric [BSM20, ABA24, QWR^{+}20]. Context [CPH^{+}22, KRW^{+}20, LLG^{+}20, MGC23, MZZ20, OLP23, PTZ^{+}20, WZX^{+}21, CYZ^{+}22, CDV^{+}24, DATAA20, GAP24, LXY21, LYBS21, MTCS22, OLLP24, DCC22]. Context-Aware [KRW^{+}20, LLG^{+}20, OLP23, CYZ^{+}22, DATAA20, OLLP24, DCC22]. context-enriched [MTCS22]. contexts [uHA20]. Contextual [CDP20a, NRMI20, PP20]. contextualized [FWX23]. continent [GZB^{+}22]. Continual [LPT22, HZPS21, YZL^{+}24, ZLH^{+}24]. Continuous [NGC24, VZDS24, YZR23, ASDLS23, BBD^{+}21, DWZ20, KAH^{+}23, NTA^{+}22]. continuously [DLdAR23, GST21]. continuum [BBF^{+}24, CPM^{+}23, FTS^{+}24, LIP^{+}24, PPSC23, RAL^{+}24, RFP^{+}24, TC23]. Contour [WTL^{+}20]. contract [FBL^{+}20, KAJ^{+}24, LTB^{+}22, QHC24]. contracting [AAGX^{+}22]. contracts [DPLV23, DBC24, TLMP20, YC22, YhSL^{+}22, ZXD^{+}20]. contractual [AAGX^{+}22]. contrast [ALGMP^{+}21, LYG^{+}24]. contrastive [LWL23a, LCY^{+}23a, MSC^{+}23]. contributed [GMGV^{+}22]. contribution [ZLC^{+}21]. contribution-aware [ZLC^{+}21]. Control [ASSG22, AMR^{+}20, ADP^{+}22, BRM^{+}20, FPMJ21, KCR20, KRW^{+}20, NJB20, RMA21, SGS24, TSX^{+}24, AAG^{+}20, BBTC20, CGFC20, CWM21, CG21, CF21, DWL^{+}23, FLTQ20, HS21, HLL^{+}20, HGWC23, ICBB20, JLP^{+}21, JR22, LMNC22, LY23, LHW^{+}23, LGCY22, LZZX20, LLY^{+}20, LXY21, MZX^{+}24, MVLJ21, MWK^{+}21, MBD21, NAT20, NPNC23, QWR^{+}20, QCY^{+}21, RYL20, RCdF^{+}21, RWJ^{+}20, SCE23, SP24, SYW^{+}23, SKX^{+}20, TLMP20, VPSC^{+}23, WCXW22, WLL24, WXZ23, YhSL^{+}22, ZWC^{+}22, ZPLQ20, ZZJC21, ZZLF21, ZZZ^{+}21a, ZLS^{+}20]. control-theoretic [MZX^{+}24]. controllability [TDS^{+}22a]. controlled [KMK^{+}23, KCKK24]. controller [AZA23, HRY^{+}21, SDZ^{+}20, XHW20]. controllers [BSM20]. controlling [THB23]. convection [LWJ^{+}23]. convergence [EBA^{+}22]. Conversation [MYT^{+}21]. conversational [AdSM^{+}22]. convolution [CHW^{+}20, CZZ^{+}23b, DBSL23, SK21b]. Convolutional [HIdAR^{+}20, KSDR21, LSMT^{+}21, CMGS22, DKG^{+}22, GS20, GWZ20, GLF24, GMH20, HZX^{+}24, LYKK22, LZP23, MSKG21, NK20, NZY^{+}23, NED^{+}20, RSFB23, TLM21, UAS^{+}20, Wan20, XTL^{+}23, XY20, YYL22, ZN21, ZRH^{+}23]. cooled [LLT22]. cooling [MMBD20, PAM21]. cooperation [RBSK23, ZLZ^{+}23b]. Cooperative [BZG23, GZPZ20, GMA^{+}22, HAA^{+}20, ZZLF21, HLP21, HYRZ20, LLZL21, LZS^{+}24, PR20, SMO^{+}24, ZNZ^{+}23]. coordinate [PDA^{+}20]. Coordination [DSC20, DDM21, PCVN21]. CoPiFL [XRZ^{+}24]. copious [ERL^{+}20]. Cora [CTFW22]. CorClustST [HSvB20]. CorClustST-Correlation-based [HSvB20]. core [CIB^{+}20, JPJO22, KYPJ20, LQW^{+}20, LCFM20, QHE^{+}20, QNHB22, WLL24, XLX^{+}21]. cores [MNFQ24]. coronary [ZZZ^{+}23]. coronavirus [KSS^{+}21]. corpora [LZL^{+}20]. correct [BKG^{+}20, TSKK23]. correction [EGDT20, EGD24]. Correlation [HZX^{+}19, HZX^{+}20, HSvB20, DSW^{+}20, MLZ^{+}23b, YLD^{+}23]. correlations [CTFW22]. correlations-based [CTFW22]. correlative [SXC^{+}23]. correlative-GNN-based [SXC^{+}23]. Corrigendum [AB19, AB21, ARA^{+}23, FGB21a, KF23, MTHA24a, SME^{+}21]. Cosine [MSA^{+}24]. Cost [BGR20, LDGS20, SSWW23, WLY23, YZW22, YLG^{+}24, BWX20, CPT^{+}20, ESSS^{+}21, FCGPSG^{+}21, FPL24, GRN20, GCCMK^{+}20, HLH^{+}20, KIM^{+}24, KNV20, KSE24, LPL22, qLhZ20, LYY^{+}20b, NT22, PYL22, RMBMT21, RRHA21, SAD24, TDL^{+}21, WYD20, WHZ^{+}20, WLZ^{+}23, WPX^{+}23, YPL24, ZCLL22]. cost-aware [KNV20, NT22, RMBMT21]. Cost-driven [LDGS20]. Cost-effective [WLY23, YZW22, YLG^{+}24, ESSS^{+}21, LPL22, PYL22, WPX^{+}23, YPL24]. Cost-efficient [SSWW23, WLZ^{+}23]. costs [Par22]. COTS [CXHS20]. CoTwin [GVCC24]. count [BK20]. countermeasures [SAAEK22]. counters [ABOS22, LLC^{+}22]. counting [GSG^{+}23, WC23]. counts [ZJW^{+}20]. county [NLO^{+}20]. county-level [NLO^{+}20]. coupling [CWM21]. Cournot [KAF^{+}20]. course [SYXL22]. cover [WCP23]. coverage [ARHT20, KXZW23, SSV24, WSXL21]. Covert [Gur21b, POBK21, SMKC20, ZZXH20]. COVID [KSS^{+}21, BCT^{+}21, MMC^{+}23, VCM^{+}21]. COVID-19 [KSS^{+}21, BCT^{+}21, MMC^{+}23, VCM^{+}21]. CP [XRHS21, ZZQ21, ZPK^{+}23]. CP-ABE [ZPK^{+}23]. CPABE [CSB23]. CPS [CCW^{+}20a, CCL^{+}20, LSH^{+}20, QJS^{+}21]. CPU [BDGG^{+}20, Gur21b, MGGG^{+}20, MTA^{+}22, RNV^{+}21, SLH^{+}20, XWR24].

CPU-generated [Gur21b]. CPU-GPUs [SLH^{+}20]. CPU/GPU [BDGG^{+}20]. CPU2017 [ZZT^{+}22]. CQA [HPP20]. CR [WZTL20]. cracking [ZPS^{+}24]. CRDTs [BPSP23]. creation [DRD20, FZC^{+}20, RNA21]. credibility [KZB^{+}23]. credit [CDR24, GFM^{+}20, LPL^{+}20, WYGP21]. CredsCache [CSH^{+}23]. Crime [Sha20, DNNG21]. Criminal [LXL^{+}21, WLC^{+}20b]. Criteria [GSKS20, PSC^{+}21, IA24, SSDC22, YZX^{+}23]. Critic [KPA24, ZZLF21]. Critical [PFP^{+}22, PFS^{+}23, AB20, BGNM20, DP20a, GIPS20, LCO^{+}23, QSZ^{+}24, RHM20, ZLZ^{+}20a, ZLL^{+}24b, ZTP20]. criticality [CCDR22]. CRNs [KAA^{+}21]. Cross [FQH^{+}24, LHY^{+}20a, BWX20, CKZ^{+}22, DFG^{+}21, GZB^{+}22, GLP^{+}24, HZPS21, HLZ^{+}22, LQW^{+}20, ONK^{+}20, PBS23, SD20, WWH^{+}21, MSTN21]. cross-chain [HLZ^{+}22]. cross-continent [GZB^{+}22]. cross-core [LQW^{+}20]. cross-datacenter [BWX20]. Cross-dataset [LHY^{+}20a]. Cross-Domain [FQH^{+}24, SD20]. Cross-MapReduce [MSTN21]. cross-modal [WWH^{+}21]. cross-platform [ONK^{+}20]. cross-relating [DFG^{+}21]. cross-system [GLP^{+}24]. crossbar [JPW20]. crossing [YLL22, ZWW^{+}20b]. crossing-workflow [ZWW^{+}20b]. Crowd [GSG^{+}23, qLhZ20, WSC^{+}23, YZX^{+}23, NIB^{+}21, RGRV^{+}20, WDG20, ZKGB20]. Crowd-enabled [YZX^{+}23]. crowd-sensing [NIB^{+}21]. Crowdsensing [WSD^{+}22, ACBT23, KOM^{+}20, LMO^{+}22, LP23, LZL^{+}24b]. crowdsourced [BEM^{+}20, HYWY22, YZW22]. Crowdsourcing [JSP23, CCHD21, FLH^{+}24, KOM^{+}22, MHL20, RLL^{+}22, SLX^{+}24, TDS^{+}22b, XRZ^{+}24, XCSF20]. Crypto [ArMA^{+}21]. Crypto-ransomware [ArMA^{+}21]. cryptocurrencies [WHC^{+}24]. cryptocurrency [CCH21, SM20].

Cryptographic [AHWB20, RMA21, DAMS23, HRX^{+}21, ZPK^{+}23]. cryptography [RMI22, SMKA23, WHJ20]. CShield} [YLS21]. CSSaaS [HCK20b]. CT [SVFdA20, Wu22, WZC^{+}22, ZHP^{+}21]. Cube [CHS22]. CUDA [GDEBC20, PSH^{+}20, ZGW^{+}23a]. CUDA-JMI [GDEBC20]. cues [GSMF20]. Cultural [WGLH20]. cultures [HZPS21]. cumulus [LWJ^{+}23]. curative [Bo19, Bo20b]. Curious [JAS^{+}20]. Currency [AHWB20, BZG23]. Current [HBSG21, SACN^{+}21, MMC22, TLX^{+}23]. Curriculum [DLW^{+}23]. curve [WHJ20]. custom [MVLJ21]. customer [ASL22, HCK20a, WBR20]. customer-centric [HCK20a]. cut [LCC^{+}24b]. cut-and-solve [LCC^{+}24b]. CV [CKZ^{+}22]. CVFuzz [LCH^{+}22]. Cyber [AABKB22, DAAW20, GCPM22, IDM^{+}20, KVCY20, LLG^{+}20, LGKA21, MR23b, NHY20, RCJZ20, Sha20, SUKN22, VKP22, YLS21, ZYL^{+}20, ZGY20, ASASA^{+}20, CFC^{+}20, CDG^{+}20, CLQS20, DG21, HRM20, HMLS20, HBSG21, IA20, JSV21, Kho21a, KYPJ20, LCLA21, MLWA20, MV21, RPP^{+}20, SZVVB^{+}23, SVN^{+}20a, SWW^{+}20, SRM^{+}23, TCMV20, VZDS24, WGLH20, XZJ^{+}20, XZK^{+}20, XWW^{+}20, YXYH20, YD21, ZXL^{+}20]. Cyber-Physical [DAAW20, IDM^{+}20, KVCY20, MR23b, LLG^{+}20, ASASA^{+}20, CDG^{+}20, CLQS20, HRM20, KYPJ20, MLWA20, TCMV20, XZJ^{+}20, XZK^{+}20, XWW^{+}20, YXYH20]. Cyber-Physical-Social [RCJZ20, SWW^{+}20, HMLS20, WGLH20]. cyber-situational [SZVVB^{+}23]. Cyber-Threat [GCPM22]. cyberattack [BAR21, SSS21]. cyberattacks [uRKI^{+}21]. cyberbullying [LVNCC21]. cyberinfrastructure [QRS^{+}21]. Cybersecurity [GMBdF^{+}23, GMP20b, GAdFGMA21, KAJ^{+}24, RNA21, YLX^{+}23]. CyberShip [SME^{+}21, SME^{+}19].

CyberShip-IoT [SME^{+}21, SME^{+}19].

cyberspace [GTG^{+}21, HRGL21, YBX^{+}23]. cycle [ADdMM20, CBC^{+}20, PZLL21, YCS^{+}20]. cycle-accurate [CBC^{+}20]. CycleGAN [HWH^{+}23b]. cyclone [MXS22].

D [MTHA24a, GLF^{+}22, HIdAR^{+}20,

HYRZ20, JZZD21, JL21, JPJO22, KSSR20, LGW^{+}21, NMRK21, NQH^{+}20, PB23, QJZ^{+}20, TJG^{+}20, VHP^{+}22]. D-UNet [QJZ^{+}20]. D2D [AQN^{+}20, LHH^{+}21]. D3QN [CHS^{+}23]. DACCA [HAA^{+}20]. DAD [MKC^{+}21]. DADIM [LWLH20]. DAFL [WYDB24]. DAG [AZA23, KPA24, MAS23, TLW^{+}24]. DAG-aware [TLW^{+}24]. DAG-based [KPA24]. DagOnStar [SGDK^{+}21]. DAI [MJW^{+}24]. DAI-NET [MJW^{+}24]. damage [ÇYZZ21, MM21a, YYB^{+}21]. damping [LYGF21]. dangerous [ZY20]. DApps [KSE24]. dark [PGSM^{+}24]. Data [ABC^{+}24, AAP21, AKPT20, BFG^{+}22, BeKTK^{+}20, CHS22, CTFW22, DSC20, DP20b, DP20c, DP21a, DP21b, EGD24, GAA^{+}21, HDN^{+}20, KSE24, LMO^{+}22, LCL^{+}20, MSTN21, NJB20, OCMJFB^{+}23, RVJMJ^{+}21, Sha20, SSC^{+}20, SHH23, TSX^{+}24, TMT22, UUH^{+}22, UPK^{+}23, UUK^{+}21, WMU^{+}24, WZL^{+}20, YMY21, IA24, AMM^{+}22, AAG22, AAG23, ADITS20, AOSA20b, ARB20, AHN21, ATZP21, AESI^{+}21, AÖ24, AAM^{+}24, APC^{+}20, ACM^{+}21, AMT^{+}21, AAG^{+}20, BJ22, BSH^{+}24, uRBIBC20, BCC^{+}22, BBM^{+}22, BWX20, BQC22, BLGCLA^{+}23, BMBE20, BLMT20, BOM^{+}22, BOL^{+}20, dMBPdSC20, BDG23, BBP22, BMZdP21, CHG^{+}20, CXS^{+}22, CLLCK20, CGFC20, CWB^{+}20, CVdRA^{+}20, CL20a, CFC^{+}20, Che20, CPT^{+}20, CXWY21, CMM21, CHS^{+}23, CLL^{+}23, CDX^{+}23, CDX^{+}24, CKZ^{+}22, CDV^{+}24, CHJK22, CFÁA^{+}20, CDF^{+}22, CKFT20, CIJM20, CKV22, DLdAR23, DCC22, DSC24, DK20, DBC24, DAT21, DPN^{+}22, DSPSNAHJ20, DML20, DP19]. data [DLH^{+}20, DWZ20, DCD^{+}22, DWL^{+}23, DSRG22, DLC^{+}22, DKD22, EBA^{+}22, ESSS^{+}21, FAAS20, FPH^{+}21, FBL^{+}20, FWP21, FLG^{+}20, FCOJFM21, FSBŞ^{+}20, FFAW20, FGG^{+}21, FPL24, GZF^{+}23, GZF^{+}20a, GBK20, GGK20, GHW^{+}20, GMT23, GSSB24, GPGG23, GJC^{+}20, GLWP20, GYAW22, GZZG24, Gur21a, HN22, HIMM20, HYWY22, HLW^{+}23b, HCB^{+}20, HWQ^{+}20, HYL^{+}20, HLH^{+}20, Hu21, HY21, HHLZ20, HAH^{+}23, HWR^{+}22, IA23, JA20, JTGH21, JZK^{+}21, JPMR21, JAAAZB20, JJZ^{+}23, JQZ^{+}22, JKS20c, KPL22, KSS^{+}20, KTIB22, KIJ^{+}24, KSC20, Kol22, KA24, KHB23, KJ24, LHC21, LSN^{+}20, LLP^{+}20a, LLZL21, LZ21a, LLW^{+}22a, LFM^{+}22, LWW^{+}22, LCL22, LZZ^{+}23, LYP^{+}24, LDWZ20, LCY^{+}23b, LEWC24, LLG^{+}20, LZA^{+}20, LWZ^{+}20, LXY21, LLT22, LYY^{+}22, LGW22, LDD^{+}22, LZZ^{+}20, LCH^{+}24, LQ20, LWH^{+}22, MHL20, MZL^{+}22, MLX23, MSLP24, MBC22, MSG^{+}20, MBM^{+}20, MC20, MDT^{+}20, MLZ^{+}22, MCBGSL24, MDDZ21, MBZ^{+}21, MABK24, MLZ^{+}23a, MWS24, MMBD20]. data [MBD21, MLZ^{+}23b, MOW^{+}20, MBB24, MMH^{+}22, MSBAU24, MSC^{+}23, NBJ21, NLO^{+}20, NFK^{+}20, OIG24, ONK^{+}20, OTMN23, OWK^{+}23, Pan20, PWV^{+}21, PSMF21, PKLC22, PCCX21, PBC^{+}22, PCACTMÁ24, Pła24, PDJS22, PS20, QHCH24, QRS^{+}21, RMC20, RCP24, RRD21, RZIX20, RGRV^{+}20, RWG21, RHWY23, RKM23, RBW20, RFP22, RLML20, RWJ^{+}20, RSFB23, RHM20, ST20a, SGSGGC^{+}23, SDGCB^{+}20, SGP^{+}20a, SCE23, SYYuR21, SYYuR22, SBD^{+}24, SIG24, SVN^{+}20a, SP24, SEKS^{+}20, SBF^{+}21, SHW24, SDZ^{+}20, SK20b, SW20, SNS^{+}20, SM20, SRA^{+}22, SK23, SYHX23, SKTP24, SXHD24, SPL24, SDV^{+}21, SGL^{+}20b, SVN20b, Sun20, SZGB24, SLX^{+}24, SLA^{+}23, TSR^{+}20, THA^{+}24, TCW^{+}22, Tao23, TM20, TAM21, TDS^{+}22a, TLW^{+}24, TSB20, UCR21, UJHN20, VPSC^{+}23, VPBE22, WMD^{+}20, WLAC20, WWY21, WWH^{+}21, WLW^{+}21, WFLC22, WHC^{+}22, WYX^{+}23b, WXD^{+}23, WWC^{+}24, WWZ^{+}20, WHW20, WWY^{+}24, WLD^{+}20b, WYGP21]. data [WT24, WWS23b, WDSK21, WWS20, WF21, WC23, WWW^{+}24, XKK20, XGX20, XCH^{+}20, XZYH22, XW23, XZH^{+}23, XCB^{+}20, XZZ^{+}20b, XLL^{+}20b, XWW^{+}20, XCZ^{+}22, XXL^{+}24, YJH^{+}20, YHW^{+}20, YLM23, YGP^{+}24, YJB^{+}21, YK20a, YWH^{+}21, YLX^{+}23, YZJ^{+}20, YYY^{+}23, YLG^{+}24, YVW^{+}20, ZBTV^{+}20, ZJL^{+}22, ZWH^{+}20, ZA20, ZCF21, ZT22a, ZL22, ZZL^{+}22, ZLW^{+}22, ZLS23, ZXX23, ZSL^{+}23a, ZLF^{+}23b, ZHL^{+}23, ZYY^{+}23, ZG23, ZHLL24, ZDLD24, ZCL24, ZXZ^{+}23, ZCS20, ZHJW20, dlVGSB^{+}20, uRLW^{+}21, AAB23, BP20, EET20, IMM^{+}20, IDM^{+}20, MAB^{+}20, PGSM^{+}24, PMCP20, PP22, SGBC^{+}20, SKH20]. data-assisted [LLT22, ZLS23]. data-aware [Kol22, TSB20]. data-based [WWZ^{+}20]. data-centric [MDDZ21]. Data-Driven [TSX^{+}24, LMO^{+}22, RVJMJ^{+}21, BSH^{+}24, FPH^{+}21, LSN^{+}20, PSMF21, SEKS^{+}20, YJB^{+}21]. Data-flow [GAA^{+}21]. data-intensive [HLH^{+}20, HHLZ20, RRD21]. data-locality [BLMT20]. data-parallel [LCY^{+}23b]. data-processing [THA^{+}24]. data-sets [OWK^{+}23]. data-smart [SVN^{+}20a]. database [BAMR20, CYH20, SG20, ŠHDT21, WYS20, ZSZ^{+}24, ZTQ^{+}20]. databases [CKV22, KYY^{+}20, MSZ^{+}20, SCL20, WZXX21, YNK^{+}20]. datacenter [BWX20, CMX^{+}20]. Datacenters [SZZY22, AABB24, BMBC20, TPD^{+}20]. Dataflow [WXC^{+}24, JMHB24, YTW^{+}20]. dataflows [RCLEB20]. dataframes [PSS^{+}23]. dataset [AHH20, BPC^{+}24, JMA^{+}21, LHY^{+}20a, LFM^{+}22, MSA^{+}24, TA23, BPC^{+}24]. datasets [ABB^{+}21, DVEE^{+}20, HSvB20, LHTSM^{+}23, SXF22, TLN23, VCG^{+}23, WSJ^{+}21, ZIOT^{+}20]. DATAVIEW [LXL^{+}23]. dating [LGT^{+}20]. DAVINCI [LKE22]. day [CHKJ20]. DBN [YHW^{+}20]. DC [KLW^{+}21]. dCCPI [LQW^{+}20]. dCCPI-predictor [LQW^{+}20]. DCEM [LZZ^{+}20]. DCN [DGY^{+}22]. DDFPN [WZX^{+}21]. DDoS [uRKI^{+}21, BeKTK^{+}20, DG21, FD21, KCB20, LZS^{+}22, MSLP24, NCLP21, UPD^{+}20, VP20, YNVRPD23, ZZZX22]. Deadline [HWR^{+}22, NGCB20, GBM20, HSR^{+}22, HWQ^{+}20, SW20, WWW^{+}24]. Deadline-based [NGCB20]. Deadline-constrained [HWR^{+}22, GBM20, HWQ^{+}20]. deal [AADM21]. debug [SNS^{+}20]. debugging [AKCP21]. decay [WDL^{+}21]. DecChain [BS20]. December [Ano20u, Ano21x, Ano22r, Ano23w]. decentralised [ACT24, FGG^{+}23, MAQ^{+}20, TOM^{+}20, TLS^{+}21]. Decentralized [DCGM20, FBL^{+}20, HCG^{+}23, BSB^{+}22, BS20, FGP20, HSGY20, JGL^{+}20, LLLS24, NAT20, PSC^{+}21, QHW^{+}20, SMKA23, XZC^{+}24, ZLL^{+}24a, ZWY^{+}21]. Decision [GSKS20, YMY21, AAG22, BÖ20a, FMM^{+}20, GSSB24, GK21, HRGL21, IA24, JCK24, KP22, Kol22, KGO^{+}20, LMO^{+}22, LLW^{+}20, LV24, PSvL^{+}20, SCXZ23, Sun20, YZX^{+}23, YZS^{+}21]. Decision-Making [GSKS20, FMM^{+}20, GSSB24, Sun20, YZX^{+}23]. decisions [ABMMC18, ABMMC22, CCHA22]. decline [KRA21]. Decoder [QHE^{+}20]. decoding [PK22, YJB^{+}21]. Decompiled [MRMM20]. decompose [AÇP22]. decomposition [KHB20, LZJ^{+}20, MXS22, XYL^{+}20]. decreased [BK20, WCWC19, WCWC20]. decryption [ZZQ21]. deduplication [ZSL^{+}23a]. DEEDS [HCB^{+}20]. Deep [ASM^{+}22, BYW^{+}21, BNA^{+}21, GZL^{+}22, GDS^{+}20, GTG^{+}21, HTLM21, IuRJ^{+}21, JAS^{+}20, JHB22, KPA24, LYH^{+}21, LZW21, LTXL22, LXL^{+}23, LHTSM^{+}23, MKB23, NRMI20, NN21, NCLP21, PAM21, RAA^{+}24, RYL20, RSB20, SP23, SSS21, VP20, WMU^{+}24, WFLC22, WLC23, WGGB24, WFA20, WTL^{+}20, XLCB20, ZWC^{+}22, ZLZ23a, ZZZ21b, AGdS^{+}21, ASYL22, ASL22, AHL^{+}23, ATT^{+}20, ABL22, Ben23a, BMM^{+}24, CLZ^{+}20, CQS^{+}23, CHS^{+}23, CZH^{+}24, CDC^{+}24, CHL23, CYH20, CMA^{+}22, DAM^{+}21, DML23, DK24, GS20, GRN20, HLP21, HIU^{+}22, HAqDE23, HJI24, rHZmH^{+}24, JTGH21, JZL^{+}20, JLW^{+}23, JCX^{+}21, JSP23, KNRI21, KMK^{+}23, KS24, KLW^{+}21, KMS20, KCY^{+}21, LYYG20a, LYYG20b, LZK21, LZ21a, LLW^{+}22a, LFM^{+}22, LZW^{+}22, LCY^{+}23b, LBY^{+}20, LGL^{+}20a, LH21, Liu21, LZS^{+}22, LWF^{+}23, LMZL24, LGL^{+}20b, LP23, MSLJ20, MYL^{+}23, MK20, MK21, MK22, MS24, MKBT24, Ngu24, PDFV21, RRAB24, SMU^{+}21, SW22, SMS^{+}24, TWM^{+}23, TBG^{+}20, TK24b]. deep [UPD^{+}20, Wan20, WCHA20, WLLF20, WLLY20, WLL21, WWH^{+}21, WYWS22, WSL^{+}23, WWZW23, WS23, WLL24, WWS23b, XLMC22, XW23, XLL^{+}20b, XY20, YJH^{+}20, YCG^{+}20, Yan21, YCYO23, Zha21, ZYF^{+}22, ZZB^{+}22, ZGK^{+}22, ZWM^{+}23, ZLZ^{+}23b, ZWZB24, ZZ24, ZLZ^{+}20b, ZH20, ZZPK21, KWL^{+}23, NRBC23, SHB22, ZWX^{+}23]. deep-attention [ASL22]. deep-learning-based [DAM^{+}21]. DeepAMD [IuRJ^{+}21]. Deeply [LZS^{+}21, LSB21, Wan21, LSL^{+}20, LF21, WSL21]. Deeply-learned [Wan21, LSL^{+}20, WSL21]. DEEPSEL [ARA^{+}23, ARA^{+}22]. DeepVulSeeker [WXZX23]. default [ZYX^{+}23]. defect [LFZJ21, LAT^{+}20]. Defending [CCL^{+}22, LWS^{+}23a, PCK20, QLJ21, TTZ^{+}21, WZHX23, WMCH22]. Defense [Elg20, DG21, HBSG21, LWW24, NCLP21, RBMCLH22, SUKN22, TSM24, WGG^{+}20, YNVRPD23, YD21, ZZG^{+}22, ZFZS23]. defenses [SCBP24]. defensive [SD24]. deferred [KP22]. Defined [FD21, HYL^{+}20, HRY^{+}21, MGM^{+}20, GZF^{+}20b, HAB^{+}20, HZZ^{+}20, JAAAZB20, LWNH22, LZS^{+}22, MAB^{+}20, MNA^{+}23, PCC21, SW22, SMS22, YZJ^{+}20, ZTP20, ZHX^{+}20, ZWZ^{+}21]. Deflated [MZLT21]. Degradation [PFS^{+}23]. degree [BR20]. degrees [ZJW^{+}20]. DeGTeC [LCY^{+}23b]. Delay [WZTL20, GCT^{+}20, LZA^{+}20, OPOG23, XLLL20]. delay-aware [LZA^{+}20]. Delay-tolerant [WZTL20]. Delegated [PAP^{+}20]. delegation [TLMP20]. deleting [YNK^{+}20]. deletion [WZXX21, XCB^{+}20]. delineation [WLLF20]. delirium [ZYX^{+}23]. Delivery [BSM20, BBP22, CPH^{+}22, HMY^{+}23, QRS^{+}21, SBMN21]. Demand [CCBFI^{+}23, ASAM20, BYW^{+}21, CMX^{+}20, JKS20b, JKS20c, LLZ^{+}24a, OPLB24, RPP^{+}20, SCA22, SPL22, ZA22]. demand-aware [CMX^{+}20]. demand-responsive [SCA22]. demote [ZZPK21]. demotion [KHL20]. Dempster [RTD24, UYH21]. deniable [CC21]. Denial [BAR21, GdOAO20, uRKI^{+}21, PDT21]. denoising [XLS^{+}21]. Dense [CKZ^{+}22, LZCGMVV20, NTA^{+}22, SGDG23]. Density [LGS^{+}23, ANH^{+}21, CLV24, FBTJ23, QNRA23, XJL^{+}24, YLZL21]. density-based [ANH^{+}21, FBTJ23]. dependence [BSF^{+}20]. Dependent [GBH^{+}23, LXC^{+}24, GYAW22, LZ22, SXW^{+}22]. deployed [ABOS22]. Deploying [LWHW24]. Deployment [WLX^{+}24, LYKK22, LDGS20, dTGC20, PPA^{+}24, PVA^{+}20, SSV24, TCMV20, VAKB23, WLD^{+}20a, WPX^{+}23, ZWC^{+}22]. Deployments [FFB20, NSJ^{+}24]. DEPO [KCP23]. depth [LTXL22, OMPSPL20, PSS^{+}23, ZG23]. derivation [XCB^{+}20]. derived [BLH^{+}24]. DESC [CHL23]. DESC-IDS [CHL23]. DeScan [dVIP24]. descent [PDA^{+}20]. description [FSD^{+}20, LWLW21, ZGK^{+}22]. Design [CDG^{+}20, HX21, LZZX20, MSR20, RBLD21, SK21a, THA^{+}24, Zha20, AK20, AABKB22, ADRP23, BGNBH^{+}20, BLGCLA^{+}23, Deh20, DGT24, DBBP24, GZF^{+}20a, GCN^{+}24, GMMR24, HLP21, LLW^{+}20, LZL^{+}23, LJW^{+}20, LJ21, MSZ^{+}20, MGC23, MÖ22, NLS23, NVS^{+}22, PLL^{+}24, PLMZ23, SCP^{+}21, TCMV20, TSKK23, VPBE22, ZYL^{+}20, ZJL^{+}22]. design-driven [BLGCLA^{+}23]. Designated [WHJ20]. Designated-verifier [WHJ20]. designed [YJLC20]. Designing [AMBD^{+}20, GCM21, KMS23, RLML20, YPEK23]. designs [ASPG^{+}21, JPW20]. destructive [DP20a]. Desynchronization [SSB^{+}20]. detail [XCS^{+}22]. Detect [WW20, CLV24, MS20, OHÁV20, RZIX20, RGRV^{+}20, SGL^{+}20a]. Detecting [AJPM20, GDCGCPVG21, HLL^{+}20, Kho21b, LCH^{+}22, MLWA20, MJSW21, MJZC21, PSMF21, PFS^{+}23, POBK21, YXL^{+}21, ZLST23, ARIB22, AOF21, AGYS20, GPWL20, GIPS20, ISD22, JSV21, KSA^{+}20, MKK^{+}24, NAC^{+}22, PLBOC20, XWM20]. Detection [AAH^{+}23, CSC23, FWY^{+}22, GCPM22, HTAY21, IuRJ^{+}21, MSLJ20, MKC^{+}21, SOKW^{+}20, UUK^{+}21, VMM20b, VSPM21, XY20, AZA23, AHH20, ArMA^{+}21, AAS^{+}20, AEN^{+}23, AJJ^{+}21, ATT^{+}20, AADM21, ADP^{+}22, ADAR22, BSF^{+}20, BAMR20, BSOK^{+}20, BCB^{+}20, BNX22, BR20, CCC^{+}21, CKL20, CDR24, CXHS20, CCC^{+}23, CSAT24, CHL23, CZZ^{+}23b, DFF^{+}23, DAM^{+}21, DAA^{+}21, DBD^{+}23, DRD20, DCD^{+}22, ESSS^{+}21, FLF^{+}21, FMN^{+}20, dRFRB24, FHGF20, FIABC^{+}20, GS20, GKB^{+}20, GPRM21, GFM^{+}20, GDGK20, GLP^{+}24, GST21, GMB23, GTG^{+}21, GZZG24, HN23, HPY20, HLT^{+}21, HYWY22, HCL^{+}22, JPMR21, Jia21, JZL^{+}24, JWC22, KMR^{+}22, KIJ^{+}24, KP22, KTC23, LS23a, LWW^{+}20, LQS^{+}20, LCH^{+}23, LWCC23, LZP23, LLY^{+}20, LLD^{+}21, Liu21, LWJ^{+}21, LTB^{+}22, LZS^{+}22, LLF^{+}23, LGJ^{+}23, LVNCC21, LPL^{+}20, MCT^{+}22, MLC^{+}20, MSLP24, MK20, MK21, MSV^{+}20, MMC22, MRS^{+}22, MJB22, MBC^{+}23, MRD^{+}20, MNA^{+}23, NADY20, NK20, NCLP21]. detection [NED^{+}20, OIG24, PJL^{+}24, PCK20, PFGDM22, PZLL21, RAA^{+}24, RAA^{+}20, RCdF^{+}21, RLCB22, RVJMJ^{+}21, SMU^{+}21, SDVC22, SPWW21, SN21, SHKW23, SRM^{+}23, SDV^{+}21, TGJ^{+}20, TTD^{+}20, TLJ^{+}22, TCW^{+}22, TLN23, TSM24, UPD^{+}20, UUH^{+}22, VFOV20, VP20, WLL22, WLC23, WZX^{+}21, XLS^{+}21, XCL^{+}20, YHC^{+}22, YLL22, YLF^{+}23, YLZ^{+}24, YPX^{+}20, YYL22, YFL^{+}24, YYB^{+}21, ZCQ^{+}23, ZLLD21, ZLPZ21, ZL22, ZLS22b, ZZG^{+}22, ZLP^{+}22, ZMJ^{+}22, ZZZ21b, ZY21, Zhu21, ZHP^{+}21, ZCWC20, ZCS20, uRKI^{+}21, AWMM^{+}23]. detection/tracking [ZLS22b]. detector [HZL22, JLK22, JLW^{+}23, MS20]. deterability [LWZ^{+}23a]. determination [AP20, CLZ^{+}20]. Determining [ARB20]. developer [BPLFRL20]. Developing [CdST^{+}20, MhCEANSM20, PBM^{+}22, PBC^{+}22]. Development [ELS20, GBC^{+}24, MZZ20, CFÁA^{+}20, CCP^{+}22, JM20, PPM24, RGDMMR^{+}23]. deviation [WGL^{+}24]. Device [NRBC23, uRLW^{+}21, AK20, CHG^{+}20, ENT^{+}22, FLG^{+}20, FZT^{+}23, HWH^{+}23b, LHL20, LZC^{+}23a, PBS23, SCBP24, SSV24, YW21, YHC20]. device-based [HWH^{+}23b]. Device-centric [uRLW^{+}21]. Deviceless [MTD^{+}24]. devices [ACG^{+}20a, BJW22, CFK^{+}20, CMGI^{+}23, CSB23, DFF^{+}23, FFAFD20, HLH^{+}20, HKS23, JCX^{+}21, KAH^{+}23, LQS^{+}20, LLFQ21, LTXL22, MA22, MDC^{+}24, OMSL20, RZA21, RAB23, SWW^{+}20, VDSB22, WHC^{+}22, WZW^{+}23, WWY^{+}24, XRHS21, XCB^{+}20, YZL^{+}24]. devil [XCS^{+}22]. DevOps [CHS^{+}24]. DEVp2p [HZT^{+}22]. DGA [ZLST23]. DGA-based [ZLST23]. DGSD [SHT^{+}21].

DGT [ZLC^{+}21]. DGX [GDS^{+}20]. DGX-1 [GDS^{+}20]. DGX-1/Pascal [GDS^{+}20]. Diabetes [ADP^{+}22, WZH^{+}22, OOB^{+}21]. diabetic [DDMP^{+}23, PRPPFRL20]. diagnose [DDMP^{+}23]. diagnosing [SXC^{+}24]. Diagnosis [TBG^{+}20, DKG^{+}22, DRD20, GZL^{+}22, KRA21, LCZB21, LYH^{+}21, LSMT^{+}21, LSGA20, MSLJ20, MMP^{+}23, OOB^{+}21, QPL^{+}22b, WFA20, XCGZ24, XY20, ZHD^{+}20, ZSL^{+}23b]. diagnostic [MSKG21]. Diagram [GGCIV20]. dialogues [LSS^{+}22]. DID [SXHD24]. DID-HVC-based [SXHD24]. DIDDOS [uRKI^{+}21]. diet [FHGF20]. diet-related [FHGF20]. difference [YXL^{+}21]. differences [KHL20]. different [CKW21, MRD^{+}20, NLS23, TG20]. Differentiable [JCW^{+}23]. Differential [BKM^{+}22, QNM24, WWZW23, CHC^{+}20, EL21, EL23, LCC^{+}24a, SCR20, WFL^{+}20, WZS^{+}22, ZHH^{+}23, ZXX23, ZLC^{+}21]. Differentially [CSS22, TGAP20, ZXL^{+}20]. differentiate [LBY^{+}20]. differently [RSL24]. differently-aged [RSL24]. Digital [CDR24, JYP24, KAJ^{+}24, NSJ^{+}24, RSR^{+}24, SMT^{+}24, SK20a, SMBB^{+}24, WLL24, WYJ^{+}24, AVK^{+}23, ADAHA^{+}21, BSH^{+}24, DGT24, DBBP24, ERG^{+}22, GVCC24, GMBdF^{+}23, GHB^{+}24, HMY^{+}23, JLS^{+}23, KPGD24, KSLC21, LLCH21, LLW^{+}22a, LEWC24, LCB^{+}23, MAC^{+}21, MSM^{+}22, MTCS22, OPLB24, OHÁV20, PMMG^{+}20, PPM24, QZZ^{+}24, RCR21, SAM^{+}24, SMS^{+}24, SQ22, SKTP24, VZDS24, WG21, YPEK23, YLX^{+}23, ZSL^{+}23b, BBF^{+}24, RD23]. dilated [GWZ20, GFZ21, MMU^{+}21]. dimension [GYAW22]. dimensional [CCW^{+}20b, DWL^{+}23, EUEU24, PCCX21, SYHX23, SXF22, ZWZ^{+}24, ZLLD21, ZCS20]. Dimensionality [OIG24, WDHY20]. Dimensioning [ABA24]. dimensions [BCM20]. DIRAC [BHSH22]. direct [MMPV22]. directed [RDR^{+}24]. Direction [HAA^{+}20, LP23, RSBM20].

Direction-Aware [HAA^{+}20]. directions [AAA20, CLLCK20, DPN^{+}22, GBP23, KSS^{+}20, TRB^{+}23b]. Director [BHSH22]. directory [MMKS22, WYZ^{+}24]. Dirichlet [ASA23]. disaggregated [TLW^{+}24]. Disaster [SKS22, CPS^{+}23, MM21a, RAA^{+}20]. DISCERNER [FCOJFM21]. discharge [MSKG21]. disclosure [NNUV20]. Discovering [DFG^{+}21, Gas22, IMuI^{+}21, CKFT20, HHLZ20, Xu21]. discovery [ASHO20, BSB^{+}22, BHH22, CCC^{+}24, LDDL21, LTB^{+}22, MLZ^{+}23b, NMRK21, SNMWC21, TD21, TDL^{+}21, WLLC20, YC22, ZWW^{+}20b]. Discrete [SGS24, YHC^{+}22, MYM^{+}21]. discretization [GBM20]. Discriminant [Che20]. discriminative [LSB21, XWK21]. discussion [BBB22]. disease [AYHA20, AMZZ23, ESSS^{+}21, GW22, KNRI21, KSS^{+}21, LSN^{+}20, MSLJ20, QPL^{+}22b, TLN23, TA23, WCWC19, WCWC20, YXLB20, ZZZ21b]. Diseases [TBG^{+}20, GDCGVG20, YGS^{+}22]. disentangled [ZHZS23]. disjoint [IMuI^{+}21]. disjunctive [TSR^{+}20]. disk [HLH^{+}20, LGJ^{+}23, YK20a]. disorder [SSMdS21]. disorders [BAGRB^{+}20, UCR21, XLCB20]. dispatching [KHHT21]. displacement [YLM23]. display [FHGF20, MFMSG20, XSW^{+}21]. disruption [AHSH22]. Dissecting [BCCS20]. Distance [GGCIV20, ZZZ^{+}23, ACN^{+}21, CZGS20, LWLH20, TGJ^{+}20, WGLH20]. distance-based [CZGS20]. Distance-Join [GGCIV20]. distillation [THVL24]. distortion [WM21]. distress [VCM^{+}21]. Distributed [BP20, CK20, CAC^{+}22, FTM20, GdOAO20, LWF^{+}23, MKC^{+}21, SHT^{+}21, SWW^{+}20, WZJ^{+}22, ZLL^{+}24a, uRKI^{+}21, AÖ24, AdAHK20, BEM^{+}24, BMS20, BSM20, BQC23, BVFGSF20, CCW^{+}20a, CBN^{+}20, CF20, dATBMA23, DKD22, EMHE21, EGDT20, FGB21a, FGB21b, GMI22, Gul22, HWR^{+}22, IAM^{+}22, JHK20, JCX^{+}21, KHES21, KNV20, LY23, Li20, LCL22, LCH^{+}23, LS23b, LYS^{+}20, LJW^{+}20, Liu23, LZL^{+}24b, LPSV22, MYL^{+}23, MMAH22, MSTN21, MYM^{+}21, MS24, MDG^{+}22, MCF20, MJW^{+}22, OPLB24, POMK20, PBS23, PPA^{+}24, QLHLB23, RSLL21, RHJ20, SMC23, SIG24, SK20a, SHW24, SQ22, SDA21, SSB^{+}20, SCC20, SCW^{+}22, TCMV20, VHP^{+}22, WGW^{+}20, WYHM21, WLS^{+}24, WYZ^{+}24, WLLC20, WWW^{+}24, XZYH22, YXYH20, ZWM^{+}23, ZLC^{+}21, SYW^{+}23]. distributed-memory [EGDT20]. distributing [RFP22]. distribution [BN21, CL21, GCM21, GAA^{+}21, HAH^{+}23, NLO^{+}20, RPP^{+}20, SHST20, ZHZS23, Zhu21]. distributional [ZC22]. distributions [NLS23]. disturbances [RYL20]. Diverse [WMU^{+}24, Cha20, KXZW23, YSL^{+}22]. diversification [YZR23]. Diversified [HPY20]. Diversity [AKJJ20, CATD^{+}24, GZXH24]. Divide [AAT^{+}24, SYYuR21]. Divide-and-Conquer [SYYuR21]. DL [RFd20]. DL-Focl [RFd20]. DL-Foil [RFd20]. DLP [FAAS20]. DMEIRL [rHZmH^{+}24]. DNA [EGDT20, YW21]. DNN [ERK^{+}24, HLK^{+}23, KCKK24, LCH^{+}21, LDX^{+}23, LJ24, ZZB^{+}24]. DNNs [JCW^{+}23]. DNNShifter [ERK^{+}24]. DNS [GdOAO20, HSGY20, YL20a]. Do [CC21, GSI22, TDLT20, DCC22, BSH^{+}21]. Do-Care [BSH^{+}21]. Docker [SZM^{+}21]. docking [SXZZ23]. doctor [DKG^{+}22]. document [QG20]. documents [Gas22]. does [MDP24]. Domain [FQH^{+}24, WYDB24, AR20, CHS^{+}24, DOR^{+}21, Kad20, KCB20, LP21b, LZC^{+}23b, PSvL^{+}20, PYL22, SD20, SNS^{+}20, TDS^{+}22a, WCHA20, WYJ^{+}24, XLS^{+}21]. domain-independent [SD20]. domains [NQB^{+}23, VI21]. dominance [SMC23]. dominated [KSSR20, WM21]. DONS [BAK22]. double [LBJ^{+}18, LBJ^{+}24, MXW22, RWG21, SS21, TTTH20, WHF^{+}23]. double-blockchain [RWG21]. Down [BMD^{+}21, AHMW23]. DP [ZHH^{+}23, ZXX23]. DP-TrajGAN [ZHH^{+}23]. DPLRS [WZJ^{+}22]. DQN [LMNC22]. DR [DLC^{+}22, FWP21]. drainage [ZMZ^{+}19, ZMZ^{+}20]. Dredas [FBL^{+}20]. drift [MKK^{+}24, XCL^{+}20]. drifts [ST20a]. drilling [LCZB21, SYXW21]. drive [SMC^{+}20]. Driven [AWMM^{+}23, CHS^{+}24, MZZ20, TSX^{+}24, AOKÖ24, ATC^{+}24, BSH^{+}24, BLGCLA^{+}23, BPUW24, CF24, dATBMA23, FPH^{+}21, FTS^{+}24, GDCGVG20, GAA^{+}21, KA21, LMO^{+}22, LSN^{+}20, LDGS20, MIMS20, PLBOC20, PSMF21, RAS^{+}20, RVJMJ^{+}21, SMC23, SYG^{+}20, SEKS^{+}20, TDS^{+}22a, TDS^{+}22b, WYGP21, WWS20, XHW20, YLF^{+}23, YJB^{+}21, HDN^{+}20, MGX^{+}23]. Driver [BQI^{+}20, SSDC22, XLMC22]. Driverless [DRC20]. driving [BQI^{+}20, SCXZ23, WFLC22, ZGW^{+}23b]. drone [KSS^{+}21, MZA23, NMRK21, QCY^{+}21, SBMN21]. drone-based [KSS^{+}21]. DroneCOCoNet [QCY^{+}21]. drones [ISD22, NKG23]. drought [BBM^{+}22]. drug [CCC^{+}24, DFG^{+}21, LP21a]. DS [TMT22]. DSI [HMA^{+}21]. DSL [MMFAB23]. DSM [RPP^{+}20]. Dual [WSL^{+}23, RPP^{+}20, TLL^{+}24, ZZZ^{+}23, WWY^{+}24]. Dual-attention [WSL^{+}23]. DUAL-C [WWY^{+}24]. dual-reputation [TLL^{+}24]. dual-view [ZZZ^{+}23]. duct [HZX^{+}19, HZX^{+}20]. duration [MSK^{+}21]. during [MM21a, MSK^{+}21, RHM20, TPF^{+}20, VCM^{+}21]. DVFS [HSS20]. DVFS-enabled [HSS20]. DWT [CKZ^{+}22]. DWT-CV [CKZ^{+}22]. DX [JLP^{+}21]. Dynamic [BAK22, DHD20, FTM20, FCOJFM21, FZ20, GWY^{+}20, HSGX22, HZX^{+}24, KCP23, LLS24, MMP^{+}23, MGX^{+}23, MA22, OLLP24, PNL^{+}21, RBMCLH22, SME^{+}21, SYXW21, SGL^{+}20b, TWL23, UUK^{+}21, XCW20, ZHGX20, ZWM^{+}23, AM20, AOF21, AKE22, BEM^{+}24, BLH^{+}24, BRK24, BSB^{+}22, BSH^{+}21, BGMK22, CECS20, CDX^{+}23, CZZ^{+}23b, CDF^{+}22, DCD^{+}24, DFZ^{+}20, EBA^{+}22, FWP21, FYHZ24, GW22, GZF^{+}20b, HTAY21, HS21, HX21, bHFF^{+}21, IT20, JA20, JWC22, KF22, KF23, Kha24, KYY^{+}20, LHC^{+}20, LWW^{+}20, LDLS22, LQYL21, LZC^{+}23b, LWLH20, LZCGMVV20, LEXH20, MLN24, MECRFD20, MFMSG20, MHA^{+}24, MSBAU24, NGCB20, NNH^{+}20, PBM^{+}22, PLBOC20, PLS^{+}23, PSH^{+}20, Pła24, QGH^{+}22, QNRA23, RSFB23, SME^{+}19, ST20a, SGSGGC^{+}23, SAT20, TYR22, TSB20, WLY^{+}20, WKW^{+}22, XWD20, XJL^{+}24, YYW^{+}21, Yu21, YFL^{+}24, YNK^{+}20, ZWC^{+}22, WMNV20, LKE22]. dynamics [BBB^{+}24, GW22, KNRI21, LGZ^{+}24, MEL^{+}23, MMFAB23, RGRV^{+}20, TBO20, TA21, XZD^{+}21, XWR24]. DYVERSE [WMNV20].

E-business [QHC24, YFL^{+}24].

E-commerce

[Zha21, ZWY^{+}21, RLZW21, YDL^{+}20]. e-health [GPC21]. E-healthcare [WHF^{+}23, ZHGX20]. e-learning [SG20]. e-voting [KAK20]. EagleMine [FLF^{+}21]. EAIS [YLTH22]. EANDC [ASYL22]. Early [LVNCC21, ArMA^{+}21, ARA^{+}22, ARA^{+}23, ESSS^{+}21, KNRI21, KRA21, MASRAM^{+}22, PGSM^{+}24, RGRV^{+}20, WYGP21, ZLS23, ZHP^{+}21]. early-stage [ZHP^{+}21]. Earth [FAA^{+}23, MBZ^{+}21]. earthquake [CMF^{+}21]. Easy [CPJ^{+}21]. EB [ZKL^{+}23]. EB-BFT [ZKL^{+}23]. ECDSA [WYZ^{+}20]. ECG [HIMM20, OMPSPL20, WLLF20, XLS^{+}21, bZSC^{+}23]. echocardiograms [PZLL21]. Economic [AB19, ABGMC19, AB21, ABGMC21, RNA21, SCGVP20, ABT20]. economical [SS21]. Economics [ABT20]. ecosystem [ADAHA^{+}21, Dho20, LGCY22, RNA21]. Ecosystems [MÖ24a, PLMZ23, RKG20, XXY^{+}23]. ECSNeT [AdAHK20]. Edge [AAH^{+}23, AAP21, BBB22, BS20, ETH20, FPMJ21, FFB20, HSR^{+}22, HGY^{+}22, HWH^{+}23b, LXC^{+}24, LCLW24, MRMB24, MKK^{+}24, MR23b, NGC24, PAM21, QKG20, SB24, SN21, TPD^{+}24, WWL21, WLX^{+}24, WML^{+}23, XLLL20, XZC^{+}24, ZWZ^{+}24, AZA23, ASL22, AM21, AdAHK20, ATC^{+}24, BEB^{+}20, BÖE24, BKV22, BPC^{+}24, BBF^{+}24, BEON24, CMGI^{+}23, CFC^{+}20, CLY^{+}20, CWL20, CF20, DCC22, DS23, DLR23, DPPGCCA23, DCZ20, DML23, DLZ^{+}23, ERK^{+}24, FWP21, GCT^{+}20, GCCMK^{+}20, GBH^{+}23, GZG20, GlRpG20, GYAW22, GZ22, HCWD21, HAVK22, HTXW21, HXL^{+}23, HCG^{+}23, HLL^{+}24, HLK^{+}23, HZS^{+}23, HKB^{+}24, JPMR21, JHB22, JYP24, JXYC24, JCX^{+}21, Kha24, KCKK24, KA21, KA24, KCJ23, KSMT24, KGO^{+}20, LHC^{+}20, LMO^{+}22, LMNC22, LBGL20, LLZL21, LWNH22, LDLS22, LZL^{+}23, LLW^{+}23a, LRCL24, LGL^{+}23, LYBS21, LZ22, LDX^{+}23, LLLS24, LMZL24, LIP^{+}24, LGL^{+}20b, LFHS23, LCLW21, MA24, MISB22, MWL^{+}20, MGX^{+}23, MV21]. edge [MA22, MKC^{+}21, MCV23, MJW^{+}22, MJW^{+}24, NPNC23, NLSY20, PLL^{+}24, PSC^{+}21, PUMN^{+}24, PPSC23, PK24, PPA^{+}24, QLJ21, QHCH24, QSZ^{+}24, QCWY23, QCW^{+}24, QL22, QZZ^{+}24, QCY^{+}21, RDR^{+}24, RRD21, RAA^{+}21, RSL24, RAL^{+}24, RGP^{+}22, SHB22, SCP^{+}21, SSV24, SXW^{+}22, SFC23, SACW23, SGDG23, SAF23, SLX^{+}24, TDM^{+}22, TK24a, THVL24, TC23, TCBF24, TPN^{+}21, URN^{+}20, VHP^{+}22, WLP^{+}20, WLD^{+}20a, WC20, WKW^{+}22, WCXW22, WX23, WLZ^{+}23, WZS^{+}23, WGGB24, WLS^{+}24, WLY23, WWY^{+}24, WPX^{+}23, XKK20, XCH^{+}20, XGY^{+}23, XXY^{+}23, XCZ^{+}23, YWG^{+}20a, YLKK20, YhSL^{+}22, YLL22, YZL^{+}23, YZL^{+}24, YYW^{+}24, YCYO23, YXS^{+}23b, ZBTV^{+}20, ZLZ^{+}20a, ZZLF21, ZPQH21, ZLZ^{+}23b, ZZB^{+}24, ZHLL24, ZWZB24, ZWH^{+}21b, ZLL^{+}23, ZDLD24, ZCL24, ZZZ^{+}21a, ZH20, ZHLM20, ZLG^{+}24, ZLH^{+}24, MPC^{+}24, RFP^{+}24, WMNV20, ZYW24]. edge-aware [ZBTV^{+}20]. edge-based [AM21, DCC22, DML23, ZPQH21, ZCL24]. edge-centric [ZHLM20]. edge-cloud [KCKK24, NPNC23, QSZ^{+}24, SLX^{+}24]. edge-computing [QZZ^{+}24]. Edge-enabled [WML^{+}23, LLLS24, THVL24, ZZZ^{+}21a]. edge-intelligent [DS23]. Edge/Fog [QKG20]. edge/hub/cloud [KSMT24]. EdgeABC [XGS^{+}20]. EdgeOptimizer} [QSZ^{+}24]. Edges [ACY20, ACDY21, VMM^{+}20a]. EdgeSimPy [SFC23]. EdgeVPN [SAF23]. EDIT [DBC24]. Editor [Fae21]. Editor-in-Chief [Fae21]. Editorial [AMB^{+}21, ACDY21, BBSB21, BDF^{+}22, CCH21, DAAW20, ICW21, LEWC24, NMR21, PJJ^{+}22, ZWH21a, ZTP20, Ano20a, Ano20b, Ano20c, Ano20d, Ano20e, Ano20f, Ano20g, Ano20h, Ano20i, Ano20j, Ano20k, Ano20l, Ano21a, Ano21b, Ano21c, Ano21d, Ano21e, Ano21f, Ano21g, Ano21h, Ano21i, Ano21j, Ano21k, Ano21l, Ano22a, Ano22b, Ano22c, Ano22d, Ano22e, Ano22f, Ano22g, Ano22h, Ano22i, Ano22j, Ano22k, Ano22l, Ano23a, Ano23b, Ano23c, Ano23d, Ano23e, Ano23f, Ano23g, Ano23h, Ano23i, Ano23j, Ano23k, Ano23l, Ano24a, Ano24b, Ano24c, Ano24d, Ano24e, Ano24f, Ano24g, Ano24h, Ano24i]. editors [Tau24]. EDOM [ZTQ^{+}20]. Education [BDF^{+}22, MBC22, MIIN23]. educational [Gas22, GZB^{+}22]. EEG [CHG^{+}20, LF21, MJW23]. EEG-based [MJW23]. Effeclouds [LPL22]. Effect [PFP^{+}22, Bo19, Bo20b, CL21, EKJ^{+}20, JLC^{+}20, KAF^{+}23, MC20, WWP19, WWP20, YWG^{+}19, YWG^{+}20b]. Effective [CKV22, KIM^{+}24, ABB^{+}21, BGR20, ESSS^{+}21, GMB23, HAK^{+}21, HJX^{+}23, IMM^{+}20, LRCL24, LSS^{+}22, LPL22, MDG^{+}22, NHTH20, PYL22, RAA^{+}20, STS^{+}20, WLY23, WPX^{+}23, XXY^{+}23, YPL24, YZW22, YLG^{+}24, ZYX^{+}23, ZLST23]. effectively [LQW^{+}20]. effectiveness [SB24]. Effects [MPS21, CF21, LOR22]. efficiency [AOKÖ24, BAİP24, DHD20, FWZ^{+}20, GBC^{+}24, KAA^{+}21, LWL23b, MDZ^{+}21, RMA^{+}20, RZA21, SK20b, XCW20, ZHL^{+}23, ZTQ^{+}20]. Efficient [ASSG22, ABB^{+}21, CCHD21, CDY^{+}20, CFÁA^{+}20, HYC^{+}21, HYL^{+}20, HNV^{+}20, JT22, KJYC23, LLZL21, LS23b, LWS^{+}23b, LRCL24, LLF^{+}23, LLLS24, NQH^{+}20, QNHB22, RCP24, RAA^{+}21, RTD24, RNV^{+}21, SYYuR22, SLS^{+}20, STH^{+}20, SYHX23, TKS^{+}23, WWY21, WCW^{+}21, WZXX21, XGX20, XWM20, YSL^{+}24, YZL^{+}24, YYW^{+}21, YNK^{+}20, ABMM18, ABMM22, AMM^{+}22, AMM^{+}20, AHL^{+}23, AÖ24, ABGDT23, AKF^{+}20, ACA^{+}23, AMBD^{+}20, ANH^{+}21, BYR^{+}20, BGCL20, CXS^{+}22, CWYG23, CA21, CSY^{+}20, CDX^{+}24, CHL23, CYG22, CPJ^{+}21, CKV22, DLR23, DC21, DVEE^{+}20, DHA^{+}20, DFZ^{+}20, DLZ^{+}23, DLL20, ERK^{+}24, FBL^{+}20, GCN^{+}24, GBK20, GLW^{+}20, GMB23, Gul22, GZZG24, HMSA^{+}23, HAVK22, HDN^{+}20, HXL^{+}23, HJI24, HRY^{+}21, IuRJ^{+}21, JJZ^{+}23, JXYC24, KHHT21, KSH^{+}21, KIJ^{+}24, KA22, KCP23, KJ24, LYY^{+}20a, LZZ^{+}23, LYP^{+}24, LLG^{+}20, LGW^{+}21, LHY^{+}20b, LFHS23, LWY^{+}24, MSZ^{+}20, MXW^{+}23, MK20, MLZ^{+}23a, MOW^{+}20, MS24, MDG^{+}22, MJW^{+}22, NADY20, NQB^{+}23]. efficient [OLLP24, PSAL20, PKLC23, PSH^{+}20, PMCP20, PPGS20, RDR^{+}24, RJA^{+}22, RKM23, SKB20, SAM^{+}24, SGDK^{+}21, SGSGGC^{+}23, SB24, SSWW23, SW20, SKH20, SK23, SCW^{+}22, TLN23, TBB^{+}23, UAACH21, WLL21, WSL^{+}23, WYX^{+}23b, WX23, WLZ^{+}23, WCS24, WGS24, WHW20, WWY^{+}24, XW23, XLZ^{+}22, YZL^{+}23, YWDC23, YYW^{+}24, YWH^{+}23, YTW^{+}20, YLC23, YXS^{+}23b, ZGY20, ZSZX24, ZGW^{+}23b, ZZZ^{+}21a, ZWWC21, ZXZ^{+}23, MRR^{+}20]. efficiently [BHL^{+}21, GIPS20]. EGNN [GZZG24]. EH [ZL23]. eHealth [GVCUGF20, WSJ^{+}21]. eHealthcare [WHA^{+}20]. EHRs [YhSL^{+}22]. EKnad [BNX22]. Elastic [BCSS20, LBGL20, SSWW23, SGL^{+}20b, SGLB22, ZKL^{+}23, PK24]. elasticity [dRRCGdC20, WDS^{+}23]. elderly [EKK23]. ELECT} [YXS^{+}23b]. election [WLAC20]. electric [AYY^{+}20, PJBB20, UKY^{+}20]. electricity [PCACTMÁ24]. electrocardiogram [NED^{+}20]. Electroencephalography [NAC^{+}22]. electromagnetic [FWY^{+}22]. electromyogram [XYL^{+}20]. Electronic [PSAL20, FWZ^{+}20, SQ22]. element [ZRH^{+}23]. element-wise [ZRH^{+}23]. elements [WZL^{+}20, ZZ21b]. elephant [dMBPdSC20, VP20]. elevated [WCWC19, WCWC20]. ELIB [MRR^{+}20]. elicitation [BGNM20, DP20a, MMPL20]. elimination [IPPK23, ZZF^{+}24]. Elite [LTX^{+}24]. elliptic [WHJ20]. Elman [Wan20]. email [FAAS20]. Embedded [AGV23, AACJ23, BCSS20, LTXL22, LDD^{+}22, NQH^{+}20, RZA21, SKH20, SMC^{+}20, ZXX^{+}20]. Embedding [CHS^{+}23, MOW^{+}20, NRMI20, ODET21, WZZD23, ASA23, CPH^{+}22, GWY^{+}20, HMLS20, LY21, LLWJ24, MTM21, PJLL23, POBK21, TYR22, WCW^{+}21, XTL^{+}23, ZGK^{+}22, ZLW^{+}24]. embeddings [AR20, GDCGCPVG21, GSDGP21, Gas22, HT22]. emergencies [GW22]. emergency [SKS22]. emergent [BDK^{+}20]. Emerging [AAM^{+}24, WMCH22, YMS20a, IB20, LWHW22, VKP22, VI21]. emissions [ISD22]. Emotion [LF21, dAMVULM20, UCR21, XLCB20, XLL20a, YWG^{+}20a]. emotion-aware [YWG^{+}20a]. emotional [HZPS21, SSMdS21]. emotions [AGdS^{+}21, BEKF21]. EMPAIA [JLS^{+}23]. Empathic [AdSM^{+}22]. Empirical [DP20b, TA23, ATC^{+}24, GMP20b, LZJ^{+}20, PLL^{+}24, ZMJ^{+}22]. Employing [NLS23, DOR^{+}21]. Empowered [CSD^{+}23, AVK^{+}23, Ano24s, MMZI22, WS23]. Empowering [LFM^{+}22, WLP^{+}20]. Emulation [Elg20]. enable [GMMR24, LAHN22]. Enabled [LXL^{+}23, AUJW22, ADP^{+}22, BKV22, BCT^{+}21, BKV^{+}20, CZH^{+}24, CCBFI^{+}23, CMA^{+}22, CKV22, FMM^{+}20, FZT^{+}23, GVCC24, GKB^{+}20, GLW^{+}20, GK21, GPC21, HN23, HSS20, HGK20, JZK^{+}21, JZZ^{+}23, KDX^{+}24, KAJ^{+}24, LTB^{+}22, LLLS24, LGKA21, MXW22, MFE^{+}20, MNA^{+}23, Par20, RRAB24, SCE23, SYYuR22, SRP20, SKTP24, TRB^{+}23a, Tao23, THVL24, WLN^{+}21, WML^{+}23, YWG^{+}20a, YZX^{+}23, ZXY^{+}21, ZBF22, ZWZ^{+}23, ZZZ^{+}21a]. enabler [NMR21]. Enabling [CPS^{+}23, CDV^{+}24, CYG22, CCC^{+}24, CHS^{+}24, EBA^{+}22, GZB^{+}22, HN22, MRMB24, PBK^{+}22, WMU^{+}24, WZW^{+}20, YLS21, ZLXH20, ZLF^{+}23a, ZXZ^{+}23, ZLW^{+}24, ARB20, ASH^{+}23, CCP^{+}22, EELB21]. encapsulated [SDGCB^{+}20]. encoded [LCCP21]. Encoder [QHE^{+}20, ŠTI24, ZZ24]. encoder-only [ŠTI24]. encoding [PK22]. Encompassing [BOD^{+}24]. encrypted [CKV22, FRGBHPPS23, LZZ^{+}23, NBJ21, SCZ^{+}20]. Encryption [SZM22, ATK^{+}22, CC21, DDT^{+}23, HN22, HIMM20, HWH^{+}23a, HLH^{+}20, KSSR20, LYP^{+}24, LMW^{+}24, MAS23, MTHA24a, MTHA24b, MTA^{+}22, NNH^{+}20, Pan20, PBSS24, TSR^{+}20, UAACH21, WHF^{+}20, WCXW22, WHF^{+}23, WC23, XYH^{+}24, XCB^{+}20, XRZ^{+}24, YC22, YLH^{+}23]. End [BÖE24, MGM^{+}20, MSBAU24, PWV^{+}21, TC23, ATT^{+}20, BBB22, GBdRACG20, JM20, KHH21, LDX^{+}23, LGJ^{+}23, LLT20, PSvL^{+}20, RAB23, RSFB23, ZYF^{+}22, ZWZ^{+}24]. end-of-life [PSvL^{+}20]. End-to-end [BÖE24, MGM^{+}20, PWV^{+}21, TC23, ATT^{+}20, BBB22, KHH21, LGJ^{+}23, RSFB23, ZYF^{+}22]. end-user [JM20, LLT20]. Endoscopy [WMU^{+}24, WMU^{+}23, MKK^{+}20]. endpoints [ZZXH20]. EneA [ABMO24]. EneA-FL [ABMO24]. Energy [ABMO24, ACA^{+}23, CKW21, GB20, GZZG24, HCWD21, HXL^{+}23, HRGL21, HJI24, KVCY20, LYYG20a, LYY^{+}20a, LYYG20b, LHH^{+}21, LLZ^{+}22, LHY^{+}20b, LFHS23, MDZ^{+}21, MISB22, PAM21, RJA^{+}22, SVD^{+}20, SW20, SK23, SZZY22, SHKW23, TSX^{+}24, TWI20, WYD20, WLL21, XZJ^{+}20, YLTH22, YXS^{+}23b, ZWH21a, ABAJ20, APC^{+}20, BJ22, BRK24, DLR23, DHA^{+}20, DFZ^{+}20, DHD20, GZF^{+}23, GA22, GCN^{+}24, GLW^{+}20, GBC^{+}24, Gul22, HBEK20, HSS20, HRM20, HLH^{+}20, HGWC23, HWR^{+}22, JHB22, JXYC24, JKS20b, KHHT21, KSH^{+}21, KTIB22, KCP23, LZL^{+}24a, LDWZ20, LLG^{+}20, LGL^{+}20a, LGKA21, LEXH20, MC20, MGGG^{+}20, MDT^{+}20, MR23b, MOW^{+}20, MOU^{+}21, MDG^{+}22, NQB^{+}23, PJBB20, PKLC22, PWH^{+}22, yQhJL20, RZA21, SAM^{+}24, SHZMA21, SHB22, SKH20, SCW^{+}22, TPD^{+}20, TDM^{+}22, TBB^{+}23, WC22a, WFLL22, WSL^{+}23, WX23, WHW20, XLG^{+}23, XZK^{+}20, YJLC20, YZL^{+}23, YWH^{+}23, YSL^{+}22, ZAH^{+}20, ZGY20, ZLXH20, ZL23, ZTQ^{+}20, Zhu20]. ENergy [RNA^{+}22]. Energy-aware [ABMO24, GB20, HCWD21, LYYG20a, LYYG20b, LHH^{+}21, YLTH22, HRM20, HWR^{+}22, PJBB20, yQhJL20, SHB22]. energy-based [JKS20b]. Energy-conscious [PAM21]. Energy-efficient [GZZG24, HXL^{+}23, LYY^{+}20a, LHY^{+}20b, LFHS23, RJA^{+}22, SW20, SK23, WLL21, YXS^{+}23b, DFZ^{+}20, GCN^{+}24, JXYC24, MDG^{+}22, SAM^{+}24, TBB^{+}23, WSL^{+}23, WX23, WHW20, YWH^{+}23]. energy-harvesting [WFLL22]. energy-neutral [ZL23]. Energy-Optimal [TSX^{+}24]. energy-performance [XZK^{+}20]. energy-saving [BJ22]. Energy-SLA-aware [MISB22]. Enforcement [SME^{+}21, SME^{+}19, SNM^{+}20]. Enforcing [MWK^{+}21]. engine [BGNBH^{+}20, MBD^{+}20, TLX^{+}23, TBB^{+}23]. EngineCL [NBB20]. Engineering [CHS^{+}24, RMA21, BDK^{+}20, FPH^{+}21, LH21, LPL^{+}20, NN21, SW22, SQ22, SZM^{+}21, Wan21, WXD^{+}23, ZGZX21, CECS20]. English [LC20]. enhance [ARbL^{+}20, CMGI^{+}23, KHB20, SMKC20, SZS^{+}21, ZGZX21]. Enhanced [GMP20b, KIJ^{+}24, RKG20, VMM20b, ADAR22, CATD^{+}24, GWY^{+}20, GBdRACG20, LY21, SNM^{+}20, SZdLZ22, WZX^{+}21, ZrHhH^{+}23, ZLL^{+}24b, ZMJ^{+}22]. enhancement [LSH^{+}20, SCX21, ZYX^{+}23, ZWZ^{+}24, ZLPZ21, ZWH^{+}21b]. enhancements [DFF21, UCO20]. Enhancing [BQC23, MBB24, NTI24, dRRCGdC20, SVN^{+}20a, TDMC23, VMCM^{+}20, YFL^{+}24, EKK23, GFPB23, GLWP20, XLCB20, XHW20, ZL22]. enigma [PJL^{+}24]. enriched [MTCS22]. enrichment [XKK20]. Ensemble [AAH^{+}23, JWC22, DLW^{+}23, HAK^{+}21, KCY^{+}21, LWS^{+}23a, LLZ20, LGJ^{+}23, LSGA20, MGX^{+}23, MKC^{+}21, TWL23, TBG^{+}20, YFQ^{+}22, YLZ^{+}24, ZHD^{+}20, ZLLD21, ZKGB20, ZA22]. Ensemble-Learning [AAH^{+}23]. Ensembles [LSMT^{+}21, PBK^{+}22]. Entanglement [PCI^{+}24]. enterprise [AK20]. entities [AAGX^{+}22, ALS21a, HHLZ20, LZW21]. Entity [JTGH21, WZZD23, GWZ20, YZW22]. Entity-aware [JTGH21]. Entropy [LWNH22, ERL^{+}20, HLL^{+}20, LZS^{+}22, OMPSPL20, XYL^{+}20, XLG^{+}23]. Entropy-based [LWNH22, HLL^{+}20]. EnTruVe [RNA^{+}22]. enumeration [KHB20]. Environment [ADMG20, FSBŞ^{+}20, XZH^{+}23, AWMM^{+}23, Dho20, DCZ20, ENT^{+}22, FFM^{+}20, GS20, GEN20, GRG20, GMAL23, HSS20, HJI24, HCK20a, HCK20b, JAC^{+}23, KF22, KF23, KMK^{+}23, KAF^{+}20, KJ24, LMO^{+}22, LDLS20, LSH^{+}20, LLZL21, zLsZjX20, MA24, MK20, MYM^{+}21, MAM^{+}24, NGCB20, PKR21, RDR^{+}24, SKB20, SN21, SSB^{+}20, TK24b, WDG20, WYS20, XWD20, XSW^{+}21, YWH^{+}23, ZLM^{+}23, ZY20, ZZ21b, ZWY^{+}21, FLTQ20]. Environment-and-Blockchain-supported [XZH^{+}23]. environmental [JAAAZB20, LZB20, ZT22a]. environments [AM20, AdAHK20, AMBGS21, BPSP23, BBP22, CPS^{+}23, DRD20, DFZ^{+}23, EAA21, EET20, FYHZ24, FSP^{+}24, GSCP22, GST21, HTXW21, HGK20, JHB22, KHHT21, Kho21b, LYFZ20, LS23b, LWF^{+}23, MMMZ20, MPC^{+}24, MKB23, MSS^{+}24, MFMSG20, MKK^{+}24, MWS24, MOU^{+}21, NCLP21, OWK^{+}23, PKB22, PUMN^{+}24, PAS^{+}20, PPA^{+}24, RBW20, SG20, SYW^{+}23, SCP^{+}21, SLH^{+}24, SXC^{+}23, SPL24, SZGB24, TCMV20, TPF^{+}20, TBG^{+}20, WMNV20, WYX^{+}23b, WGGB24, WPX^{+}23, XYH^{+}24, ZDLD24, ZG24, KAF^{+}23, ASDLS23]. Envisioned [ZWC^{+}22]. EOS [MXW^{+}23]. epidemic [MSR20]. epileptic [AEZ22]. episodes [LLP^{+}20a]. epistatic [PFGDM22]. epistemic [ACG^{+}20b]. EPPSQ [LZZ^{+}23]. equation [ARB20]. equilibrium [LXZ^{+}20]. equipment [LHW^{+}23]. era [DFG^{+}21, RKP^{+}21]. erasure [BWX20, PK22, PAP^{+}20]. erasure-coded [BWX20, PK22]. erosion [YTQ19, YTQ20a, YTQ20b]. Erratum [DP20c, DP21a]. error [EGDT20, EGD24, WYG^{+}20, WWY^{+}24]. errors [hAS24, LRML21, MRD^{+}20]. escorted [MYL^{+}23]. EsPADA [VMM20b]. essential [MMP^{+}23]. Establishing [RAA^{+}20]. estimate [ZJW^{+}20]. Estimating [PFS^{+}23, VFOV20]. Estimation [LZJ^{+}20, MSK^{+}21, WHC^{+}24, uRBIBC20, CDP20c, DLZ^{+}23, HL24, JWZ^{+}22, KTIB22, KWL^{+}23, LRCL24, LTXL22, LZC^{+}23b, RSBM20, RAS^{+}22, RLML20, SHZMA21, SHF23, WCD^{+}22, ZST^{+}20]. estimations [SCÁB20]. estimators [BSF^{+}20]. ETERS [KSH^{+}21]. Ethereum [BCCS20, HZT^{+}22, KSE24, LDM^{+}21, LTB^{+}22]. Ethereum-based [LDM^{+}21]. EU [FAA^{+}23]. European [BDM^{+}20, CDP20b]. Evaluate [YTQ19, YTQ20b, RZH21, UADD21, YTQ20a]. Evaluating [AL20, CF21, GA22, LGL^{+}20a, MMC^{+}23, MWS24, RJM^{+}21, WDG20, ARIB22, LGM^{+}21, RNA21]. Evaluation [BDM^{+}20, BAR21, GPR^{+}24, GMFC23, Par22, XKK20, AB19, ABGMC19, AB21, ABGMC21, AKJJ20, BdL20, dMBPdSC20, DKG^{+}22, DP20b, GSARS20, GBC^{+}24, LBDP23, LGKA21, NFK^{+}20, PLL^{+}24, RAN^{+}20, See20, VMV20, WYGP21, YWS21, ZLQ23, ZDZ21]. Evaluations [LZL^{+}21]. evasion [DBD^{+}23, PCK20, RAA^{+}24]. Event [CH24, MLP^{+}21, ACC20, CF24, FMN^{+}20, LZ20a, SCÁB20, SGL^{+}20a, TGAP20, Zhu21, ZYL^{+}22]. event-driven [CF24]. Event-Layered [CH24]. events [DFG^{+}21, FAŞ^{+}20, MJZC21, RHM20]. evidence [ADAHA^{+}21, LLCH21, RTD24, WG21, TDLT20]. Evolution [BRM^{+}20, LXL^{+}21, RBSK23, BPLFRL20, CHC^{+}20, HTXW21, HX21, LCC^{+}24a, LYG^{+}24, LQNW20, zLsZjX20, LZZ^{+}20, SCR20]. evolution-assisted [CHC^{+}20]. Evolutionary [LCC^{+}24a, MMMZ20, EJP22, GBH^{+}23, HMO^{+}20, HOMD21, IT20, JCK24, LTX^{+}24, LHW20, MLP^{+}21, MJZC21, NVS^{+}22, yQhJL20, SMC23, ST20a, WCY^{+}20, gWLWZ21, YPZ^{+}24]. eVolvable [LKE22]. Evolving [AMNZ20, MSG^{+}20, MTD^{+}24, CHL23]. Exact [WHW20]. exam [KZF21]. example [GRG20]. examples [MMH^{+}22]. Exascale [CIB^{+}20, FAA^{+}23, MGM^{+}20, ABC^{+}20, LGM^{+}20]. Excellence [FAA^{+}23]. Exchange [CSD^{+}23, BMZdP21, CXHC23, HLZ^{+}22, MSBAU24, PCI^{+}24, QG20, TDS^{+}22a, WHJ20]. exclusive [LLY^{+}20]. ExDe [THA^{+}24]. executable [NADY20]. executing [Hu20, SDA21]. Execution [CMM^{+}23, FLTQ20, XZH^{+}23, AHMW23, AM21, AFL23, BDGG^{+}20, CCP^{+}20, CYH20, DA22, DFZ^{+}23, GVCUGF20, HSS20, MGZ^{+}20, MJSW21, SOT24, SNS^{+}20, VG21, XSW^{+}21, ZY20, ZGN^{+}20]. Executors [SHH23]. Exfiltrating [Gur21a]. existing [FD21]. Exogenous [DDM21]. exoskeleton [ZWL21]. Expanding [LIP^{+}24]. expansion [ARHT20, JQZ^{+}22]. experience [HBSG21, SGDK^{+}21, YPEK23]. Experiences [CIB^{+}20]. Experimental [ZLQ23, JAAAZB20, ZWL21]. experimentation [LZHL23, LOR22, PPA^{+}24]. Experimenting [TLMP20]. Experiments [MGZ^{+}20, KIM^{+}24, MGC23]. expert [SDV^{+}21]. expertise [ORPPG20]. explainability [ERG^{+}22, LV24]. Explainable [CSD^{+}23, DKG^{+}22, HIU^{+}22, KAH^{+}23, KMR^{+}22, ASYL22, Ano24s, BEM^{+}24, BSH^{+}24, DDMP^{+}23, GSSB24, KRA21, KAJ^{+}24, MCT^{+}22, QPL^{+}22b, SXC^{+}23, TRB^{+}23a, TK24b, WZC^{+}22]. explanation [CCHA22]. Explicit [TIA21, KW20]. Exploit [BNX22, BDG23]. Exploiting [BBF^{+}24, JZL^{+}20, KKT^{+}23, SJVRS22, URN^{+}20, GZF^{+}23, GMGV^{+}22]. Exploration [AKPT20, BDG23, DLGW^{+}20, IMM^{+}20, RFd20, THA^{+}24, YWDC23, ZIOT^{+}20].

Exploratory [YRV^{+}22, CF24, YWDC23]. Explore [XW23, VCG^{+}23]. Exploring [GCCMK^{+}20, GZF^{+}20b, LWW^{+}20, ZTC20, ASL22]. expression [LWLW21, RFd20, XLCB20]. expressions [XCSF20]. Extended [SYXL22, DAMS23, LZH^{+}20]. Extending [Gul22, MMK^{+}20]. extensible [XLZ^{+}22]. external [CCL^{+}21, SXF22]. Extracting [HTLM21]. Extraction [CGM^{+}23, CLZ^{+}20, CCC^{+}23, HAK^{+}21, HY21, TIA21, WCP23, WJC^{+}24, Wei21, XFJ^{+}20, ZDC22]. extractive [CZ20]. extractor [HLW^{+}23a]. EXTraS [DLGW^{+}20]. Extreme [LYC^{+}22, DK20, GMH20, KAK^{+}23, LYKK22, Par20].

FaaS [BPGL21, FSP^{+}24, HKB^{+}24, MTD^{+}24]. FaaS-based [HKB^{+}24]. FaaSification [PRF22]. FaaVPP [ABAJ20]. Fabric [LGKA21, LCL^{+}20, MCBGSL24]. Fabric-enabled [LGKA21]. face [SNM^{+}20, SZS^{+}21, ZZ21a, ZLZ^{+}20b, MBD^{+}20]. faces [SG20]. Facial [LWLW21, CSS22]. Facilitating [CCL^{+}21]. facilitation [AdSM^{+}22]. facility [SCK^{+}22]. factor [BGCL20, GOA23, SD20, YPX^{+}20]. Factorization [SYXL22, CFL^{+}20]. factors [ARB20]. factory [CG21]. Faculty [LZL^{+}21]. failure [AIM23, BMD^{+}21, LQML22, LGJ^{+}23, LZC^{+}23b, SYG^{+}20]. failures [HRY^{+}21]. Fair [GSSB24, MAM^{+}24, NBJ21, ZGN^{+}20, CDY^{+}20, DSPSNAHJ20, SS21]. Fairness [GZ22, MÖ24a, GZB^{+}22, MÖ22, WF21]. Fairness-oriented [GZ22]. fake [HAK^{+}21, YXL^{+}21]. Fall [GK21, CXHS20, LZP23]. False [HAH^{+}23]. Farming [CMA^{+}22, GAP24]. farms [MBD^{+}20]. Fast [IMM^{+}20, SHZMA21, SVN20b, BR20, Jia21, KDX^{+}24, LYKK22, LJL^{+}21, LCH^{+}23, dTGC20, Wu22, YLSL22a]. fast-iterative [Wu22]. FastAiAlloc [dSOFC^{+}23]. Faster [YPL24, AHMW23, JLT^{+}21]. Faster-RCNN [JLT^{+}21]. Fat [BMZdP21]. fatigue [ZGW^{+}23b]. fattening [GAP24]. faucet [MÖ22]. Fault [BVCH22, LGM^{+}20, PRF20, WYWS22, DRD20, DEJ20, GCN^{+}24, HYL^{+}20, KYPJ20, MA22, NKB^{+}20, PRF22, SXC^{+}24, TWY^{+}23, ZHD^{+}20, ZLW^{+}24]. fault-tolerance [GCN^{+}24]. Fault-tolerant [BVCH22, MA22, ZLW^{+}24]. FC [GMAL23]. FDMS [BQI^{+}20]. FDMSs [BQI^{+}20]. feasibility [NKG23, SCGVP20]. Feature [ArMA^{+}21, DIB20, LXL^{+}21, LYC^{+}22, MSA^{+}24, XLL^{+}20b, ZZZ21b, ABL23, ARA^{+}22, ARA^{+}23, BK20, BH21, CLZ^{+}20, CPH^{+}22, CYZ^{+}22, CCC^{+}23, CZZ^{+}23b, FPH^{+}21, GDEBC20, HAK^{+}21, HMH^{+}22, KMK^{+}23, LZS^{+}21, LSB21, LWJ^{+}21, LPL^{+}20, MHH^{+}20, MASRAM^{+}22, NNUV20, NN21, ST20a, SGL^{+}20a, TA23, Wan21, WFL^{+}21, WLL22, WWF^{+}23, WJC^{+}24, Wei21, WZH^{+}22, YLZ^{+}24, ZDC22, ZLLD21, ZGZX21, ZZZX22, Zhu20]. feature-based [ABL23]. featured [LAS20]. features [AMM^{+}20, DAM^{+}21, GDCGCPVG21, HZT^{+}22, JL21, JLW^{+}23, LSL^{+}20, LF21, MSY20, MMK^{+}20, MMU^{+}21, SZL^{+}21, TLM21, TGJ^{+}20, TTD^{+}20, TLJ^{+}22, WML^{+}21, XWK21, YLG^{+}24, ZKGB20, ZLST23, ZY21]. features-based [JLW^{+}23]. February [Ano21w, Ano20r, Ano22w, Ano23n, Ano24j]. Fed [DLC^{+}22]. Fed-DR-Filter [DLC^{+}22]. FedBnR [WXX^{+}24]. FedDCS [PBS23]. FedDOVe [SP23]. Federated [ABC^{+}24, AWMM^{+}23, AAH^{+}23, ADP^{+}22, BBB^{+}20, CMM21, DML23, FRGBHPPS23, HAVK22, IPPK23, MMZI22, MRMB24, PK24, SP23, SGS24, SuRMA^{+}23, SRA^{+}22, SLH^{+}24, WMU^{+}24, ZCL24, ZLH^{+}24, AMM^{+}22, ABMO24, ASL22, AB24, AIM23, CXS^{+}22, CYWS24, DFF^{+}23, DLZ^{+}23, DLC^{+}22, EUEU24, EL23, FXZZ24, GFPB23, HN23, HOMD21, HL24, JZZ^{+}23, JZL^{+}24, KJYC23, KAF^{+}20, KJ24, LWL23b, LWS^{+}23a, LLW^{+}23b, LLW^{+}23a, LZTM24, LLZ^{+}24a, LLLS24, LIP^{+}24, LZL^{+}24c, LCH^{+}24, MCBGSL24, MA22, MDC^{+}24, MPP^{+}21, MSC^{+}23, NTI24, NT22, PBS23, Par22, PUMN^{+}24, QHNL21, QCG^{+}24, QHCH24, QCWY23, QLHLB23, RHK^{+}23, RBMCLH22, RRHA21, SCX^{+}24, SB24, SZdLZ22, TLL^{+}24, TSM24, TWM^{+}23, WXZZ22, WZHX23, WYDB24, WXX^{+}24, WGS24, WML^{+}23, WZS^{+}22, XW23, XRZ^{+}24, XPT^{+}22, XJL^{+}24, XZC^{+}24, XXL^{+}24, YLM23, YZL^{+}23, YZL^{+}24, ZZG^{+}22, ZLZ23a, ZNX23, ZGW^{+}23b, ZWCS23, ASH^{+}23, MZL^{+}22, MBB24, YYB^{+}21]. FederatedTrust [SCX^{+}24]. federation [AAG^{+}20, HMO^{+}20, HS24]. federations [FXZZ24, YMAAH22]. FedHEONN [FRGBHPPS23]. FedProc [MSC^{+}23]. FedRD [YYB^{+}21]. FedSA [CMM21]. FedSup [ZGW^{+}23b]. FedViT [ZLH^{+}24]. feed [GSMF20]. Feedback [MDP24, MMH^{+}22, ZFMB20, DWL^{+}23, GW20, LWW^{+}22, MXW22, SS21]. feedback- [LWW^{+}22]. feedback-based [MXW22, SS21]. FellowMe [ARbL^{+}20]. femtocells [YGE21]. Fetal [PZLL21, QPL^{+}22b, ZLP^{+}22]. fever [LBY^{+}20]. few [LLW^{+}22b]. few-shot [LLW^{+}22b]. FGCS [MTHA24a, AMB^{+}21, BDF^{+}22, ICW21, ZWH21a, ZTP20]. FGCS-D-23-00424 [MTHA24a]. FGMD [JLK22]. FI [TWY^{+}23, CZGS20, CDY^{+}20, YZZ^{+}23]. FIB [Dut22]. fibrillation [NED^{+}20, Wan20]. fidelity [CPT^{+}20, KIM^{+}24]. FIDO2 [XSW^{+}21]. field [PLL^{+}24]. Fields [WPPA22a, Gur21b]. fifth [AKA20]. File [PMMG^{+}20, HT22, KHES21, MMKS22, MWK^{+}21, PBL^{+}23, WYZ^{+}24, YZC^{+}20, YYZ^{+}24]. File- [PMMG^{+}20]. files [JAC^{+}23, NADY20]. filter [RSB20, DLC^{+}22, ZJL^{+}22]. filtering [CDP20a, GOA23, TKS^{+}23, XZZ^{+}20a, ZYF^{+}22]. FilterLSTM [PKLC23]. finance [WYGP21]. financial [KBTT20, LWH^{+}22, MMM^{+}20, Yan21]. Find [RAN^{+}20]. Finding [DAA^{+}21, WGL^{+}24, LHW20, SP21]. Findings [CMMST20]. Fine [FLTQ20, ZZB^{+}24, CFL^{+}20, HZZ^{+}20, JZL^{+}20, JVH^{+}20, MHL20, MPS21, SK21b, WCXW22, XCGZ24]. Fine-grained [FLTQ20, CFL^{+}20, HZZ^{+}20, JZL^{+}20, JVH^{+}20, MHL20, WCXW22, XCGZ24]. fine-tune [SK21b]. fine-tuning [MPS21]. Fingerprint [WLC^{+}20b, ZGZX21, BGCL20, FWY^{+}22, JQZ^{+}22]. fingerprint-based [JQZ^{+}22]. fingerprinting [CZGS20, HN23, HZZ^{+}20, NHTH20, SCBP24]. finite [XWR24, ZGW^{+}23a]. FinTech [YDK20]. fire [DS23, MNA^{+}23]. Firehawk [MSA^{+}24]. firewall [ZPK^{+}23]. Firewalls [UADD21]. first [LLT22]. first-principle [LLT22]. fish [ZLZ21]. Fisher [Che20]. fit [XLL^{+}20b]. fitting [ZLZ21]. fixed [BK20]. FL [ABMO24, JZL^{+}24]. FL-IIDS [JZL^{+}24]. FLAGS [SDV^{+}21]. flame [Kha24]. FLAS [RSLL21]. flash [KHES21]. flash-based [KHES21]. flask [ZY20]. Flat [WLD^{+}20a, PPGS20]. flattened [ZSZX24]. fleet [UKY^{+}20]. Flex [DGY^{+}22]. flexibility [KRW^{+}20]. Flexible [CMM^{+}23, GZF^{+}20a, ACC20, HYL^{+}20, PPA^{+}24, TBB^{+}23, WKW^{+}22, WLL24, XCSF20]. FlexVF [ENT^{+}22]. Flink [JJZ^{+}23]. flipped [SP22]. floating [BBB^{+}24, NRBC23]. floating-point [NRBC23]. flooding [BeKTK^{+}20]. flooding-based [BeKTK^{+}20]. FLoRa [HLW^{+}23a]. flow [ABMESM18, ABMESM22, AB20, ABOS22, BK20, dMBPdSC20, BN21, Deh20, DBSL23, GAA^{+}21, HZX^{+}24, JR22, LDW^{+}21, MDG^{+}22, QHNL21, QWR^{+}20, YFQ^{+}22, ZZLF21, ZZZX22]. flow-shop [MDG^{+}22]. flower [GHEB^{+}18, GHEB^{+}23]. FlowGraph [RNRA23]. Flowlet [YLSL22b, DGY^{+}22, ZHJW20]. flowlet-based [ZHJW20]. Flowlet-level [YLSL22b, DGY^{+}22]. flows [DDM21, GDP20]. Floyd [LCCP21]. FLPM [XXL^{+}24]. fluctuating [SXC^{+}23]. Fluid [XWR24, BBB^{+}24, MMFAB23]. Fluidity [PPA^{+}24]. Flux [ABC^{+}20]. fly [MSMJ22, WWY^{+}24]. Focl [RFd20]. focus [RLML20]. focused [MEL^{+}23]. foetal [LZJ^{+}20]. Fog [ARbL^{+}20, AMBD^{+}20, BHH22, CMA^{+}22, CHS^{+}24, DK20, ETH20, FGB21a, FGB21b, GZPZ20, HGK20, HB21, KRW^{+}20, PKB22, QKG20, SDA21, VPBE22, WZL^{+}20, WZB^{+}20, ZGY20, ASBT20, ANA24, BM20, BMM^{+}24, CdRRdC^{+}24, DPPGCCA23, EAA21, FMM^{+}20, GS20, GEN20, GVCUGF20, GRN20, GSCP22, GMAL23, HHH22, HOMD21, HBH21, HS24, KF23, KHHT21, KS24, KJ24, LZL^{+}24a, MZA23, MMZI22, MWS24, NIB^{+}21, PNL^{+}21, PDJS22, RNA^{+}22, RJA^{+}22, RMBMT21, RBA^{+}22, SYYuR21, SYYuR22, SIG24, SP23, SSV24, TBG^{+}20, VHP^{+}22, VS20, WWY21, WGGB24, gWLWZ21, Zha20, AALEF20, ABAJ20, ABL23, NGCB20, KF22]. Fog-Based [KRW^{+}20, GRN20, KJ24, NIB^{+}21, RJA^{+}22]. fog-centric [HBH21]. Fog-enabled [CMA^{+}22, FMM^{+}20, SYYuR22]. Fog-to-Cloud [AMBD^{+}20]. fog-to-fog [PNL^{+}21]. fogging [DATAA20]. fogs [RAN^{+}20]. Foil [RFd20]. Follow [HJW^{+}20]. following [GLF24, ZYX^{+}23]. food [FHGF20, MBD^{+}20]. foot [DDMP^{+}23]. footprint [KXZW23]. footprints [MSM^{+}22]. foraging [DT21, YHC^{+}22]. forcing [WW20, WLZ^{+}23]. forecast [VGL23, ZZD22]. Forecasting [LYG^{+}24, LH21, SK20b, ACF^{+}21, DBSL23, EL23, EJP22, HSGX22, KTIB22, LMCSE20, MMM^{+}20, MXS22, PB23, PCACTMÁ24, WZW^{+}23]. forecasts [AOSA20b]. forensic [KMS20, SBF^{+}21, KSLC21]. Forensics [UUK^{+}21, ASASA^{+}20, ADAHA^{+}21, DAT21, KSLC21, LKJN^{+}20, LLCH21, OHÁV20, RAA^{+}21, UAACH21, WG21]. forest [DS23, MNA^{+}23, WDL^{+}21, ZT22b]. ForestEyes [DFF21]. forests [AIM23]. form [BMS20, QJZ^{+}20]. Formal [NMRK21, RSQS21, RHJ20, SK21a, AM22]. formation [FFM^{+}20, HBEK20, HMO^{+}20, HOMD21]. formats [JVH^{+}20]. Forward [KCR20, LYY^{+}22, NBJ21]. forwarding [AAG22, BBN^{+}20, CCW^{+}20b, DCC22, Dut22, FGG^{+}23, NPNC23, ZHX^{+}20, AAG23]. foundation [SN23]. founded [TDMC23]. four [LHY^{+}20a]. four-stage [LHY^{+}20a]. FP [RK20, TZW^{+}22]. FP-GNN [TZW^{+}22]. FP-growth [RK20]. FPGA [HMSA^{+}23, PBM^{+}22, RNV^{+}21, SXF22, SHF23, TZW^{+}22, ZJL^{+}22]. FPGAs [ISUC22, JCP^{+}20, SHF23, WXC^{+}24]. fractal [LYC^{+}22]. fractals [NQH^{+}20, QNHB22]. fractional [MSR20]. fractures [WWP19, WWP20]. fragment [HT22, ZWW^{+}20b]. fragmentation [WLLC20, ZZF^{+}24]. fragmentation-aware [ZZF^{+}24]. Framework [CHC^{+}20, MZZ20, RMA21, SME^{+}21, ABMM18, ABMMC18, AB19, ABGMC19, ABM19, AB21, ABGMC21, ABM21, ABMM22, ABMMC22, AHSH22, AGdS^{+}21, ACBT23, hAS24, AOF21, ADITS20, AHH20, ASHO20, AESI^{+}21, ANA24, AABB24, AYB^{+}22, ACT24, BLH^{+}24, BGNBH^{+}20, BDM^{+}20, BFM23, CCC^{+}21, CCW^{+}20a, CCHA22, CS23, CCP^{+}22, CDP20b, DAMS23, DSC20, DKG^{+}22, dATBMA23, DLH^{+}20, DLZ^{+}23, DFZ^{+}23, EUEU24, FWP21, FSBŞ^{+}20, FGP23, GS20, GBdRACG20, GNA^{+}21, GSCP22, GMP20b, HJX^{+}23, HDN^{+}20, HTXW21, HLL^{+}24, HZS^{+}23, rHZmH^{+}24, HCK20b, IA20, IHA^{+}20, JRW^{+}20, JCP^{+}20, JAS^{+}20, JAAAZB20, JYP24, JZZD21, JCX^{+}21, JKS20c, KBTT20, KOM^{+}20, KMK^{+}23, KCKK24, KMS20, KSMT24, KSLC21, KAJ^{+}24, LAHN22, LSN^{+}20, LYFZ20, LCY^{+}23b, LGCY22, LLZ20, LYH^{+}21, LPL22, LDX^{+}23, LLLS24, Liu24, LGKA21, LCB^{+}23, MK21, MMZI22, MS24, NAC^{+}22, NUB24, dSOFC^{+}23, OPLB24]. framework [ONK^{+}20, PBS23, PUMN^{+}24, PMMSE21, PMCP20, PP22, PZLL21, QCWY23, RAN^{+}20, RSR^{+}24, RSQS21, RGDMMR^{+}23, SME^{+}19, ST20a, SKS22, SNM^{+}20, SHZMA21, SMKA23, SCE23, SHST20, SXZZ23, SZW^{+}23, SSS21, SRA^{+}22, SKTP24, SDA21, SCW^{+}22, SGLB22, TBO20, TWY^{+}23, TG20, UKY^{+}20, WC20, WLAC20, WC22b, WKW^{+}22, WXZX23, WYX^{+}23b, WCP23, WWF^{+}23, XLS^{+}21, XCGZ24, Yan21, YLZ^{+}24, YYB^{+}21, YNVRPD23, ZrHhH^{+}23, ZZB^{+}24, ZYL^{+}22, ZWCS23, ZHL24, XQW^{+}24]. frameworks [LGM^{+}21, MLZ^{+}22, ORLV20, SGBC^{+}20, TLW^{+}24]. fraud [CDR24, GFM^{+}20, HCL^{+}22, LTB^{+}22, LPL^{+}20, ZZB^{+}22]. free [CECS20, CKV22, ICBB20, LLT22, PJL^{+}24, XCW20, YLH^{+}23]. free-cooled [LLT22]. Freight [TSX^{+}24]. frequency [ASB^{+}23, LLP^{+}20a, Liu23, MXS22, RZA21, TA21, YXLB20, Zhu21]. frequency-switchable [Liu23]. frequent [AM22, LMZ^{+}22]. freshness [DKD22]. friendly [MTHA24a, MTHA24b]. frog [GZT^{+}21]. frontal [MNSL22]. fronthaul [YK20b]. FSRM [LHY^{+}20a]. FSRM-STS [LHY^{+}20a]. FTLink [HYL^{+}20]. fulfilling [SEKS^{+}20]. Fully [HIdAR^{+}20, CPYY23, CDRS20, LYY^{+}22]. Function [BKM^{+}22, HTAY21, MDZ^{+}21, MABK24, NUB24, RPF21, BPC^{+}24, CS24b, DSDV20, GMA^{+}22, KKL^{+}24, PCC21, RZIX20, SEL^{+}22, WGW^{+}20, WLZ^{+}23, WYJ^{+}24, XWD20, YXL^{+}20, YCYO23, MA24]. Function-as-a-Service [MABK24, NUB24, BPC^{+}24]. Functional [ZHP^{+}21, DDT^{+}23, ZZZ^{+}23]. Functional-realistic [ZHP^{+}21]. Functions [MGZ^{+}20, HRX^{+}21, LLC^{+}23, RFP^{+}24, TCBF24, WBR20, ZLL^{+}23]. fusing [GPWL20, JL21, SZL^{+}21, SDV^{+}21, ZDZ21]. Fusion [TRB^{+}23b, BH21, CZZ^{+}23b, CS24b, ESSS^{+}21, GHEB^{+}18, GHEB^{+}23, JZK^{+}21, JAAAZB20, JYSH23, KMK^{+}23, KLA22, LQG^{+}23, LZA^{+}20, LJ21, LXY21, LLW^{+}22b, MMP^{+}23, MR23a, Pan20, PCCX21, SK23, WLZ^{+}20, WFL^{+}21, WJC^{+}24, XLL20a, XWG^{+}21, YWG^{+}20a, YGS^{+}22, ZLLD21, ZG23, Zhu20]. fusion-based [YGS^{+}22]. Future [AB19, AB21, ABGMC21, ABM21, ABMESM22, ABMM22, ABMMC22, ALR^{+}20, ARA^{+}23, Bo20b, BDFR22, DP20c, DP21a, DP21b, FGB21a, GHEB^{+}23, HZX^{+}20, JLC^{+}20, KF23, LBJ^{+}24, LYYG20b, LWW24, LCLW24, MTHA24a, MSM^{+}22, NHY20, SME^{+}21, WWP20, WCWC20, YW21, YGD^{+}21, YWG^{+}20b, YTQ20a, YTQ20b, ZMZ^{+}20, wZcZN^{+}20, AAA20, CLLCK20, DRC20, DPN^{+}22, BAM^{+}24, JAAAZB20, KSS^{+}20, MGB24, MSR20, PJJ^{+}22, PPM24, RHK^{+}23, SACN^{+}21, SMT^{+}24, SAAEK22, TLX^{+}23, TRB^{+}23b, VEH^{+}23, YGE21, YMS20b, uRLW^{+}21, Fae21, Tau23, Tau24]. Future-generation [LWW24, MSM^{+}22, PJJ^{+}22]. fuzzing [DFZ^{+}23, LCFM20]. Fuzzy [AM20, VP20, BÖ20a, CCTZ22, CIS^{+}20, GEN20, HLW^{+}23a, HCK20b, JKS20a, JR22, LWLW21, RAN^{+}20, SN21, UYH21, gWLWZ21, XYH^{+}24, ZNZ^{+}23, BQI^{+}20]. fuzzy-based [RAN^{+}20, BQI^{+}20]. fuzzy-decision [BÖ20a]. Fuzzy-Genetic [AM20]. Fuzzy-Taylor-elephant [VP20].

G [WYG^{+}20, PKR21]. G-SEAP

[WYG^{+}20]. GA [ASA23]. GABC [Elg20]. gain [WSXL21]. gait [KAK^{+}23, KMK^{+}23]. Game [GFM^{+}20, MXL^{+}20, CLY^{+}20, DG21, EKK23, HMO^{+}20, HOMD21, KNV20, LZS^{+}24, MMMZ20, RBSK23, SZdLZ22, TBO20, TPD^{+}20, WGG^{+}20, WCY^{+}20, YJLC20, ZNZ^{+}23]. game-based [RBSK23]. game-enhanced [SZdLZ22]. game-theoretic [KNV20]. games [AUJW22]. GAN [AADM21, LGJ^{+}23, WZC^{+}22, YLY^{+}23]. GAN-based [LGJ^{+}23, YLY^{+}23]. Gap [MAM^{+}24, GMP20b, TLKX21]. Gap-Priority [MAM^{+}24]. gapped [Gur21a, Gur21b]. GARLSched [LZW^{+}22]. Gas [PRD^{+}22]. Gated [uRKI^{+}21, LZP23]. gateway [ARHT20, CBN^{+}20, DLGW^{+}20, OGO^{+}20, SJD^{+}20, YVW^{+}20, MO24b]. Gateways [DPG20, CHJK22]. gathering [DSRG22, PGMP23]. GDPR [ZBS23]. gearbox [SDGCB^{+}20]. gem5 [CBC^{+}20]. gem5-x86 [CBC^{+}20]. gen [MR23b, RSR^{+}24]. GenArchBench [LVLBB^{+}24]. gene [JLC^{+}20, JZM^{+}22, YWG^{+}19, YWG^{+}20b]. Gener [AB19, AB21, ABGMC21, ABM21, ABMESM22, ABMM22, ABMMC22, ARA^{+}23, Bo20b, DP20c, DP21a, DP21b, FGB21a, GHEB^{+}23, HZX^{+}20, JLC^{+}20, KF23, LYYG20b, SME^{+}21, WWP20, WCWC20, YWG^{+}20b, YTQ20a, YTQ20b, ZMZ^{+}20, wZcZN^{+}20]. general [ACT24, CCHA22, LZC^{+}23b, ST20a]. generalization [KLA22, MBB24]. generated [BOL^{+}20, Gur21b]. Generating [DOR^{+}21, OTMN23, XCS^{+}22]. Generation [DNNG21, Fae21, LBJ^{+}24, LCLW24, MTHA24a, Tau23, Tau24, UADD21, AHH20, AKA20, AKF^{+}20, AAG^{+}20, BDFR22, CCdS23, DCC22, GSSB24, HLW^{+}23a, rHZmH^{+}24, LCH^{+}22, LWW24, LGC^{+}21, LGYC20, LGL^{+}20a, MYT^{+}21, MSR20, MSM^{+}22, PJJ^{+}22, PCCX21, PPM24, RCdF^{+}21, YGD^{+}21, ZZ21a, ZHH^{+}23, dlVGSB^{+}20]. generations [WD24]. Generative [LZW^{+}22, RAA^{+}24, CCHA22, CSS22, DCD^{+}22, HRX^{+}21, SHKW23, WFL^{+}21, WZC^{+}22, YYY^{+}23, ZZ21a]. GENerator [MMR^{+}23b]. generators [CCdS23]. Genetic [AM20, Elg20, NK20, PLBOC20, ASA23, AEZ22, DP20a, DDMP^{+}23, DSW^{+}20, DZXS21, HMO^{+}20, KSSR20, KA22, LEXH20, MISB22, Pan20, RDR^{+}24, XLX^{+}21]. genome [GLF^{+}22, WXD^{+}23]. genomic [YC22]. Genomics [SHH23, AIM23, HMSA^{+}23, LVLBB^{+}24]. genotyping [WLW^{+}21]. Geo [WCL^{+}24, KXZW23, MSTN21, WWW^{+}24, XZYH22, CDX^{+}23]. Geo-aware [WCL^{+}24]. geo-distributed [MSTN21, WWW^{+}24, XZYH22]. geo-diverse [KXZW23]. geocoded [NLO^{+}20]. geographic [TD21, YYW^{+}21]. geographically [HWR^{+}22, LCL22, LS23b]. GeoPM [rHZmH^{+}24]. GeoPM-DMEIRL [rHZmH^{+}24]. geospatial [LPSV22, WFLC22]. gesture [RSFB23, SZL^{+}21]. gestures [SG20]. Ghost [GSG^{+}23]. Glaucoma [DKG^{+}22]. gliomas [MXL^{+}20]. Global [MLC^{+}20, CZCH24, DLC^{+}22, GAA^{+}21, KV22, LLZ20, OWK^{+}23, SMKA23, SZW^{+}23, SZL^{+}21, WDSK21, YLGG21]. Globus [BPC^{+}24, CPM^{+}23]. GLOR [NNH^{+}20]. GLR [LH20]. GMM [ZST^{+}20]. GMPLS [FNRP20]. GNN [SXC^{+}23, TZW^{+}22, ZWW^{+}23]. go [DCC22]. goal [BEL20, dATBMA23, LHA20]. goal-driven [dATBMA23]. goal-rationality [BEL20]. good [GZF^{+}23, LSS^{+}22, OMPSPL20]. GoodFATR [CGM^{+}23]. Google [MGZ^{+}20]. gossip [OCSCB22]. gouty [YTQ19, YTQ20a, YTQ20b]. governance [ARB20, XWLC20, ZLS23]. government [ZLS23]. GOZDE [KV22]. GP [HCG^{+}23]. GP-NFSP [HCG^{+}23]. GPARS [WCS24]. GPGPUs [WYG^{+}20]. Gpipe [ZLQ23]. GPS [PZHD20]. GPS-aided [PZHD20]. GPU [ASB^{+}23, BDGG^{+}20, CMGI^{+}23, CDC^{+}24, HYC^{+}21, bHFF^{+}21, JMHB24, JPJO22, JCK24, KCP23, LAHN22, LCH^{+}21, LWJ^{+}23, LGS^{+}23, LXL^{+}23, NQH^{+}20, NQB^{+}23, QNRA23, RTD24, RKP^{+}21, RLML20, Ski20, SCK^{+}22, WCS24, XWR24, ZGW^{+}23a, ZGN^{+}20, ZHLM20]. GPU-accelerated [JCK24]. GPU-algorithms [RTD24]. GPU-based [Ski20]. GPU-Enabled [LXL^{+}23]. GPU-parallel [RLML20]. GPUs [CCP^{+}20, KW20, PDA^{+}20, QNHB22, SLH^{+}20, SCA22, TBA23, YLTH22]. gradient [LYKK22, LMCSE20, ZLC^{+}21]. grading [MXL^{+}20]. Gradual [ArMA^{+}21, SCL20]. grained [CFL^{+}20, FLTQ20, HZZ^{+}20, JZL^{+}20, JVH^{+}20, MHL20, SGL^{+}20a, WCXW22, XCGZ24]. grams [ZXX^{+}20]. granularity [GPWL20, TLJ^{+}22]. Graph [FWX23, GTG^{+}21, HCL^{+}22, LZL^{+}21, MMR^{+}23b, PJJ^{+}22, PJLL23, TLM21, TZW^{+}22, WCS24, XCH^{+}20, XWZM24, YLGG21, YZL^{+}20, AB20, BLH^{+}24, CYH20, DNNG21, DBSL23, GEN20, GYAW22, GZZG24, HTLM21, HDD20, Hu21, HZX^{+}24, IAM^{+}22, JWYÍ21, JWZ^{+}22, KLW^{+}21, LSN^{+}20, LP21a, LXH^{+}21, LWL23a, LWS^{+}23b, LYG^{+}24, LCY^{+}23b, LDCZ20, LMZL24, LH20, LCCP21, MSLP24, MJZC21, NZY^{+}23, RLZW21, SHT^{+}21, ŠHDT21, SCL20, SPWW21, SVN20b, SZO^{+}20, WZZD23, WXZX23, WCL^{+}24, WLS^{+}24, WLR21, WC23, XTL^{+}23, XZD^{+}21, Xu21, YLSL22b, YLF^{+}23, YYKK20, ZN21, ZJW^{+}20, ZLP^{+}22, ZHLM20, ZYL^{+}22, DMPS23, PJL^{+}24, YLGG21, ZWW^{+}23]. graph-augmented [WCL^{+}24]. Graph-based [XCH^{+}20, LH20, XZD^{+}21]. Graph-CAT [YLGG21]. graph-clustering-based [Hu21]. graph-encoded [LCCP21]. graph-evolution [LYG^{+}24]. graph-kernel-based [MJZC21]. Graph-powered [PJJ^{+}22]. graph-temporal [LCY^{+}23b]. graphical [JM20, LOH^{+}23]. graphics [MFMSG20, XWM20]. Graphs [AFMG^{+}22, AMNZ20, AD21, BFG^{+}22, FTM20, BEKF21, CDF^{+}22, CYG22, DOR^{+}21, LPT22, LHW20, LL20, MBGC20, MAQ^{+}20, SCC20, TLM21, TD21, WCW^{+}21, WSXL21, WWF^{+}23, ZLW^{+}24]. Grassroots [BEON24]. gray [XCW20]. Great [HIdAR^{+}20]. Greedy [JZM^{+}22]. green [FNRP20, HBEK20, RMC20, SCX21, TPD^{+}20, WYD20, ZGY20, ZWZ^{+}21, GBC^{+}24]. greenhouse [RSR^{+}24]. grid [BHSH22, CFD^{+}20, DGT24, GPWL20, GJC^{+}20, LXZ^{+}20, LGL^{+}20a, TAM21, YGP^{+}24]. gridding [WYX^{+}23b]. GridMesa [YGP^{+}24]. grids [APC^{+}20, LZZ^{+}23]. grippers [AGV23]. ground [CDX^{+}23, GMA^{+}22, XJL^{+}24, ZZG^{+}24]. Group [AND24, DKD22, GCH^{+}22, ANA24, FFM^{+}20, KJYC23, LWCC23, LGYC20, MMH^{+}22, ZWW^{+}20a, ZLW^{+}24, GPC21, LFM^{+}22]. Group-based [GPC21]. Group-of-Single-Class-predictors [LFM^{+}22]. Grouping [CZ20]. groups [ZIOT^{+}20]. growing [MSZ^{+}20]. growth [RK20]. GRU [uRKI^{+}21, ABL22, SKTP24]. GRU-based [ABL22, SKTP24]. gShare [LAHN22]. GSM [LOH^{+}23]. GSMM [GBC^{+}24]. GTFS [VCK^{+}20]. Guadiana [RGPGP24]. guarantee [MDZ^{+}21, PP24]. Guaranteeing [KHH21]. guarantees [CWYG23]. guessing [ZZQ21]. Guest [BDF^{+}22, PJJ^{+}22]. guide [AABKB22]. Guided [SGS24, BH21, FLF^{+}21, VGM24, YPZ^{+}24, YZS^{+}21]. guiding [MXW^{+}23].

H2M [KKT^{+}23]. HA-D3QN [CHS^{+}23]. HABits [SVD^{+}20]. Hadoop [PS20, RK20]. HAFLoop [ZFMB20]. hailing [LLZ^{+}24a]. HAMLET [FGP23]. Hamming [TGJ^{+}20]. hand [HZLH21, SG20, SZL^{+}21, XYL^{+}20]. handle [BGR20, TM20]. Handling [AND24, See20, SCL20, ZT22a, CDF^{+}22, ST20a]. handover [GPC21, KKL^{+}24]. Handwriting [KNRI21, MMP^{+}23]. hard [YWDC23]. hard-transiting [YWDC23]. hardness [AGV23]. hardware [DJP^{+}24, BCSS20, CF20, FSD^{+}20, HBK20, LLC^{+}22, LZL^{+}23, MSR20, SCBP24, SHF23, TLC^{+}20, WD24, XHW20]. hardware-based [SCBP24]. hardware-bound [DJP^{+}24]. hardware-driven [XHW20]. harmonic [SLH^{+}20]. harmonizing [TLW^{+}24]. harms [OOZ^{+}23]. harness [MEL^{+}23]. Harris [CHC^{+}20]. harvesting [MR23b, WC22a, WFLL22, ZLXH20]. hash [MSZ^{+}20, NADY20, RAB23, YLF^{+}23]. hash-based [MSZ^{+}20, RAB23]. hashing [CSC23, GDGK20, SZO^{+}20]. Hasse [WLR21, ZrHhH^{+}23]. Hatch [RFP22]. hawks [CHC^{+}20]. hazard [AABB24]. HazardNet [AABB24]. hazardous [KHB23, YYB^{+}21]. HBM [RNV^{+}21]. HCI [SZS^{+}21]. HCL [AYA^{+}23]. HCL-Classifier [AYA^{+}23]. head [ZLP^{+}22]. headsets [FHGF20]. Health [DP20c, DP21a, DP21b, PSAL20, PRPPFRL20, SZM22, WWC^{+}24, ZDZ21, ASYL22, BDFR22, DP19, GW22, GPC21, MAC^{+}21, NLO^{+}20, OCMJFB^{+}23, SGSGGC^{+}23, UCR21, UYH21, ZXZ^{+}23]. health-care [UYH21]. Healthcare [TBG^{+}20, ZXY^{+}21, AdSM^{+}22, AMB^{+}21, AESI^{+}21, AIM23, BEM^{+}24, BSH^{+}21, CdRRdC^{+}24, GLWP20, HHW^{+}22, JAAAZB20, LLFQ21, LLW^{+}23b, LEWC24, LLZ^{+}24b, LQ20, MKK^{+}20, QMCX20, SRA^{+}22, SK23, SXHD24, SACW23, TAM21, WC20, WHF^{+}23, WZH^{+}22, XWW^{+}20, YLH^{+}23, YJB^{+}21, ZHGX20, ZLF^{+}23a, ZSL^{+}23a, ZZZ^{+}21a, ZSL^{+}23b, UPK^{+}23, ZXZ^{+}23]. healthcare-oriented [YJB^{+}21]. HealthFog [TBG^{+}20]. HealthXAI [KRA21]. Heart [AYHA20, HIdAR^{+}20, TBG^{+}20, AIM23, LZJ^{+}20, QPL^{+}22b, SYG^{+}20, YXLB20]. heating [BKV^{+}20]. HEGrid [WYX^{+}23b]. height [MXS22]. Hellinger [PCCX21]. hemodialysis [CXHS20]. hepatitis [WCWC19, WCWC20]. herd [VP20]. Hermes [BVCH22]. hetero [LMO^{+}22]. hetero-crowdsensing [LMO^{+}22]. Heterogeneity [LBGL20, MDC^{+}24, YZL^{+}23]. Heterogeneity-aware [LBGL20, YZL^{+}23]. Heterogeneous [BDGG^{+}20, KKT^{+}23, LWJ^{+}23, LXC^{+}24, LP21b, NLSY20, NBB20, ORPPG20, RNRA23, AMM^{+}22, AACJ23, ACG^{+}20a, AQN^{+}20, AMBGS21, ATC^{+}24, BLT^{+}24, BBM^{+}22, CZCH24, CIS^{+}20, CQS^{+}23, CF20, DAMS23, FFAFD20, GAA^{+}21, GHW^{+}20, GDEBC20, GTG^{+}21, HJX^{+}23, HZdLZ20, HZS^{+}23, JWYÍ21, JYSH23, JXYC24, JCX^{+}21, LQG^{+}23, LS23b, LZC^{+}23a, LLWJ24, LZ22, LWF^{+}23, Liu24, LHY^{+}20b, LEXH20, MSLJ20, MMMZ20, MBD21, MBB24, PBM^{+}22, PJLL23, PP24, PBC^{+}22, RNV^{+}21, SJVRS22, SYXL22, TLC^{+}20, TDMC23, TPD^{+}24, WYD20, WYX^{+}23b, WWF^{+}23, WCS24, XWR24, YSL^{+}24, YHC20, YZL^{+}24, YGE21, YLF^{+}23, ZCL24, ZTB23, ZLG^{+}24]. heterogeneous-graph-driven [YLF^{+}23]. HeterPS [LWF^{+}23]. Heuristic [KHHV21, CHS^{+}23, DFZ^{+}23, KIJ^{+}24, WGL^{+}24]. Heuristic-based [KHHV21]. heuristics [SCK^{+}22]. HFuzz [LCFM20]. HGC [GAT^{+}20]. HGRBOL [KAK^{+}23]. HGRNet [KMK^{+}23]. Hidden [LGT^{+}20, PDFV21, AÇP22, CKL20, FLG^{+}20, MPS21, PGHS20, WXZ23, ZZQ21]. HIDE [UPK^{+}23]. Hiding [SZM22, WSJ^{+}21, PSAL20, ST20b]. Hier [QCWY23]. Hier-SFL [QCWY23]. Hierarchical [MTM21, AMM^{+}22, CXS^{+}22, FXZZ24, GWZ20, GPWL20, LMNC22, Li20, LDCZ20, QHCH24, QCWY23, SYG^{+}20, SZS^{+}21, WLD^{+}20a, WLR21, XY20, ZZB^{+}24]. hierarchy [LZHS24]. High [CXS^{+}22, CIS^{+}20, DWL^{+}23, GHW^{+}20, bHFF^{+}21, LZC^{+}23a, LLZ^{+}22, RNV^{+}21, TDC^{+}20, XWLC20, YTQ19, YTQ20a, YTQ20b, ZZ21a, ASPG^{+}21, BLGCLA^{+}23, CGMT20, CPT^{+}20, CPH^{+}22, CFÁA^{+}20, Den20, FWZ^{+}20, IuRJ^{+}21, KYY^{+}20, LFZJ21, LZW^{+}22, Liu21, MMFAB23, MTA^{+}22, MEC^{+}20, OGO^{+}20, PSS^{+}23, PK22, RCP24, SAD24, SXF22, SKX^{+}20, UCO20, WYX^{+}23b, YW21, YXLB20, YLTH22, YK20a, YYL22, YYKK20, YNK^{+}20, ZDC22, ZLPZ21, ZLW^{+}24, ZHJW20]. high-availability [BLGCLA^{+}23]. High-dimensional [DWL^{+}23]. high-efficiency [FWZ^{+}20]. High-efficient [CXS^{+}22, IuRJ^{+}21]. high-frequency [YXLB20]. High-performance [GHW^{+}20, MMFAB23, OGO^{+}20, PSS^{+}23, PK22, SAD24, YLTH22, YYKK20]. high-productivity [CGMT20]. High-quality [XWLC20, ZZ21a]. high-reliability [RCP24]. High-resolution [YTQ19, YTQ20a, YTQ20b]. High-speed [LZC^{+}23a]. High-throughput [CIS^{+}20, YW21]. Higher [LXL^{+}21, MIIN23, YXS23a]. Higher-Order [LXL^{+}21]. Highly [XLX^{+}21, ABGDT23, BSB^{+}22, BPSP23, DC21, JCP^{+}20, PSH^{+}20, WYS20, WX24, ZFMB20]. highly-efficient [PSH^{+}20]. historical [DLdAR23, SYHX23]. history [ZWW^{+}23]. hitching [LYW23]. HIVE [ONK^{+}20]. HMMs [LPL^{+}20]. Hoc [ZWX^{+}23, FPMJ21, SKX^{+}20, UJHN20]. HOG [ZY21]. Holistic [MBD21, Ngu24, MBB^{+}20]. home [DCC22, FMM^{+}20, GZG20, MAC^{+}21, OOZ^{+}23, QNM24, SKTP24, YXLB20, ZTC20, RGDMMR^{+}23]. homes [JKS20b, RAS^{+}20, RLQ^{+}21]. homogeneous [CVdRA^{+}20, ZTB23]. homomorphic [FWZ^{+}20, HN22, MTT^{+}23, MTA^{+}22, XRZ^{+}24, YC22]. homomorphically [FRGBHPPS23]. honest [FZC^{+}20]. honeynet [RZH21]. honeypot [AZA23]. Honing [CIB^{+}20]. hooks [AKCP21]. hop [HB21, MR23b]. Hopfield [CL21]. horizon [HSGX22]. horizontal [WLY23]. hospital [TDLT20, YXLB20]. hospitals [Kon21]. Host [MRS^{+}22, ZLPZ21]. Host-based [MRS^{+}22]. hosts [YZJ^{+}20]. HostWatcher [YZJ^{+}20]. Hot [MSY20, DAA^{+}21]. hotline [ZLS23]. hotness [ZSZX24]. hotspots [CLV24]. housing [BKV^{+}20]. HPC [LFYH22, ACA^{+}23, CKW21, dFCC23, EBA^{+}22, LVLBB^{+}24, MDP24, MGGG^{+}20, MAB^{+}20, MBC^{+}23, NKB^{+}20, NSR^{+}23, PMMSE21, PBK^{+}22, PP22, SSB^{+}20, SLA^{+}23, VSV^{+}23, WD24]. HPC-cloud [PMMSE21]. HPCG [GMFC23]. HPCLS [Liu24]. HPCLS-BC [Liu24]. HPCP [LLZ^{+}22]. HPCP-QCWOA [LLZ^{+}22]. HSAS [YSL^{+}24]. HSCFC [DWL^{+}23]. HSCFHA [MSA^{+}24]. HSE [FWZ^{+}20]. HSE-Voting [FWZ^{+}20]. HTTP [DVEE^{+}20]. hub [KSMT24, SJD^{+}20, TLC^{+}20]. Huffman [HIMM20]. Human [BH21, FGP23, HAqDE23, KAK^{+}23, LHA20, LMZ^{+}22, MMU^{+}21, NFK^{+}20, YZZ^{+}23, YMY21, AR20, AAR^{+}20, DHC23, uHA20, GCCMK^{+}20, KMK^{+}23, LSS^{+}22, LZC21, MTCS22, PKLC23, PLHC24, PYL22, RAS^{+}22, SZL^{+}21, Wan21, WXS^{+}22, Xu21]. Human-centered [FGP23]. human-centric [GCCMK^{+}20]. human-in-the-loop [WXS^{+}22]. human-robot [AR20]. Human-to-human [YZZ^{+}23]. hunt [RBSK23]. HVAC [DRD20]. HVC [SXHD24]. Hybrid [ADMG20, AMT^{+}21, DBSL23, DZXS21, HGY^{+}22, KCR20, MYT^{+}21, RCLEB20, RSL24, ZGW^{+}23a, ABMESM18, ABMESM22, AYA^{+}23, AALEF20, AM20, CS24a, CDBD24, DCD^{+}22, GGK20, GWW^{+}22, ICBB20, IA23, IA24, KAA^{+}21, KIJ^{+}24, KP22, LWX22, LLW^{+}20, LP21a, LQML22, LCC^{+}24a, LDGS20, LLS24, LZA^{+}20, LDCZ20, LYGF21, LHW20, MDG^{+}22, NNH^{+}20, OPLB24, PSMF21, PP20, PZLL21, RGRV^{+}20, RAS^{+}22, RSFB23, SVN^{+}20a, TLN23, TA23, TDS^{+}22b, TK24a, TSB20, TK24b, WSC^{+}23, WWC^{+}24, XRHS21, YBX^{+}23, YLZ^{+}24, ZLZ21, ZL23, ZKGB20, ZLP^{+}22, MSA^{+}24]. hybrid-storage [LLS24]. Hyper [PFP^{+}22, FCOJFM21]. hyper-scale [FCOJFM21]. Hyper-Threading [PFP^{+}22]. HyperFlow [MGZ^{+}20]. hypergraph [JSV21, MBZ^{+}21, SOT24, GAT^{+}20]. hypergraph-based [MBZ^{+}21]. Hyperledger [LGKA21]. hyperspectral [DLR23]. hypothesis [TBH23]. hyrax [Kha24]. hysteretic [LYGF21].

I-Scheduler [EMHE21]. I/O [BBN^{+}20, EELB21, KHES21, LFHS23]. I/Oclouds [MTD^{+}24]. IaaS [PVA^{+}20, SPL22, TBO20]. IBoNN [NAC^{+}22]. Icing [CFD^{+}20]. ICN [BBP22, HGK20, RHK^{+}23, URN^{+}20, XZZ^{+}20b, ZLS^{+}20]. ICN-based [URN^{+}20]. ICN-IoT [BBP22, RHK^{+}23]. ICT [PPM24]. ICTBioMed [SJD^{+}20]. ICU [WDL^{+}21]. IDDANet [MGX^{+}23]. ideation [MJB22]. Identification [BCCS20, AHSH22, AYHA20, ARA^{+}22, ARA^{+}23, BH21, GWY^{+}20, GDCPVG22, IuRJ^{+}21, KNRI21, KSDR21, LZ20a, PGSM^{+}24, SCBP24, STS^{+}20, SSS21, WML^{+}21, WXZX23, WLC^{+}20b, YLD^{+}23, ZZ21b, ZYF^{+}22, Zhu21, uRKI^{+}21]. identify [SB24, SBF^{+}21]. Identifying [AOSA20a, DDMP^{+}23, HZT^{+}22, SXC^{+}23, ZZB^{+}22, MM21a, MSY20, MTCS22, QMCX20, YXL^{+}21]. identity [BDM^{+}20, LGCY22, LGT^{+}20, UAACH21, WSC^{+}23, ZWWC21]. identity-based [UAACH21, ZWWC21]. ideology [GDCPVG22]. IDH [CXHS20]. IDLists [HNV^{+}20]. IDS [CHL23, MRS^{+}22]. IF [AAG23]. iFlask [ZY20]. Ignis [PMCP20]. II [WPPA22b]. IID [CXS^{+}22, CMM21, LCH^{+}24, MZL^{+}22, MSC^{+}23, WXX^{+}24].

IIDS [JZL^{+}24]. III [GW20]. IIoT [DMPS23, FWY^{+}22, FQH^{+}24, GLW^{+}20, HN22, LZF^{+}24, PMMG^{+}20, WLC^{+}20a, ZGK^{+}22, ZBF22, ZWZ^{+}23]. IIoT-enabled [GLW^{+}20]. IIVIFS [GSKS20]. IIVIFS-WASPAS [GSKS20]. illegal [MCF20]. illness [KCY^{+}21]. Image [GXS22, JL21, VFOV20, YJH^{+}20, Zhu20, ASPG^{+}21, Ben23a, BBB^{+}20, CLZ^{+}20, CDX^{+}24, CSC23, DVV^{+}20, DLR23, DKG^{+}22, GHEB^{+}18, GHEB^{+}23, HQLH20, JPJO22, KSSR20, LLW^{+}20, LYFZ20, LZ21b, LHXL22, LWZ^{+}20, LYH^{+}21, MK21, MPC^{+}24, QZZH21, QC21, SGBC^{+}20, SCZ^{+}20, WCHA20, Wei21, WFA20, XTL^{+}23, XWK21, YBX^{+}23, YJQ^{+}23, ZZ21a, ZCF21, ZH20, ZHP^{+}21]. image-aided [LYH^{+}21]. Image/video [JL21]. Images [PSAL20, YTQ19, YTQ20b, FS21, HIU^{+}22, HZLH21, HZL^{+}21, LFM^{+}22, LYC^{+}22, PBSS24, QLHLB23, SLH^{+}20, SVFdA20, XW21, YXL^{+}21, YTQ20a, YJQ^{+}23, ZDC22, ZWH^{+}21b]. imaging [AADM21, BCT^{+}21, dAdSM^{+}22]. imaging-based [AADM21]. imbalance [AADM21]. Imbalanced [DCD^{+}22, AMM^{+}22, LHTSM^{+}23, WWS23b]. Imitation [KDX^{+}24]. immersion [PAM21]. immigrant [PLBOC20]. immunity [LCH^{+}23]. impact [ABMM18, ABMM22, BCCS20, CF24, DLC^{+}22, GSDGP21, MDP24, BCC^{+}22]. impacts [LKS^{+}21]. imperfect [ZHGX20]. impersonation [Kho21b]. implants [MSR20]. Implementation [BQI^{+}20, DSDV20, FAA^{+}23, SK21a, SDKM20, HWH^{+}23b, LOR22, LZC^{+}23a, SCP^{+}21, Yue20, Zha21, ZLF^{+}23b]. Implementation-independent [DSDV20]. implementations [CHS22, DAMS23, HKB^{+}24, LCFM20, RMA^{+}20, SLH^{+}20]. Implementing [Ski20, ARB20]. implicit [FFAW20, LP23]. Importance [QHCH24, NZY^{+}23]. Importance-aware [QHCH24]. improve [FSP^{+}24, PP22, RZA21, SHY^{+}21]. Improved [AM20, dAMVULM20, QZZ^{+}24, XCL^{+}20, uRBIBC20, BK20, CVdRA^{+}20, DCZ20, HZLH21, JLT^{+}21, LLP^{+}20b, LMZ^{+}22, MR23a, PKR21, QZZH21, TRB^{+}23a, TTD^{+}20, YLZ^{+}24, YXYH20, ZG23, ZWL22]. improvement [GVCC24, JLP^{+}21, LZH^{+}20]. improvements [CKW21]. Improving [BHSH22, DGK20, FPH^{+}21, GGCIV20, GW20, JCW^{+}23, KHES21, KW20, LDLS20, LWW^{+}22, LZL^{+}21, LGS^{+}23, PPG^{+}20, SOT24, SSM22, ZRH^{+}23, ZTQ^{+}20, BBD^{+}21, BDT21, CCW^{+}20a, FLG^{+}20, MOU^{+}21, NKG23, SDZ^{+}20]. imputation [LHC21, LWW^{+}22, ZZ24]. In-depth [PSS^{+}23, OMPSPL20]. in-memory [JA20, JCK24, ZSZ^{+}24]. in-network [CF20, DCD^{+}24, URN^{+}20, XGX20]. in-orbit [CDX^{+}24]. in-situ [LS23b]. in-the-wild [uHA20]. in-transit [ZBTV^{+}20]. in-vehicle [BDG23]. Incentive [LYW23, WCY^{+}20, WYHM21, HSR^{+}22, LLZ^{+}24b, WLC^{+}20a]. incentivized [DG21]. Inception [LYH^{+}21]. Inception-v3 [LYH^{+}21]. incident [SD24]. Incomplete [Che20, Tao23, WCW^{+}21]. incompletely [SWW^{+}20]. incorporated [MECRFD20]. Incorporating [KZB^{+}23, Zhu21]. incorporation [dAMVULM20]. Incremental [JZM^{+}22, SP21, ANH^{+}21, IAM^{+}22, JZL^{+}24, WWY21]. independent [BK20, DSDV20, HCG^{+}23, SD20, dlVGSB^{+}20]. Indeterminacy [HMT^{+}20, BN21]. Index [CC21, KHL20, XYL^{+}20, ZSZX24]. Indexing [OWK^{+}23, JA20, MSZ^{+}20, SYHX23, ZSZ^{+}24, dVIP24]. Indian [SG20]. Indicator [CGM^{+}23]. indicators [HZX^{+}19, HZX^{+}20, LZL^{+}20, SK20b]. indices [ACG^{+}20b, GZT^{+}21, NLO^{+}20]. Indistinguishability [HZ20]. individual [WCL^{+}24]. individualized [MOW^{+}20].

Indoor [NHTH20, ASH^{+}23, Ano24s, KAH^{+}23, LQS^{+}20, LWS^{+}23b, MDL^{+}23, MR23a, SDKM20, TRB^{+}23a, TK24b, WJC^{+}24, YVSG22]. indoor/outdoor [LQS^{+}20]. inducing [DZB23]. induction [JCK24, Liu23, RFd20]. induction-based [Liu23]. inductive [WZZD23]. Industrial [AVK^{+}23, DMPS23, HL24, MBF^{+}20, BKV22, CHJK22, CCDR22, FBL^{+}20, FYHZ24, GMP^{+}20a, HLL^{+}20, HGWC23, LZF^{+}24, LCLA21, MJW^{+}24, QJS^{+}21, QCW^{+}24, RCdF^{+}21, SMC^{+}20, ZWH^{+}20, FZT^{+}23, HN23, LCH^{+}23, LWZ^{+}23a, LMW^{+}24, WSL^{+}23, XZC^{+}24]. industrial-based [LCLA21]. industries [KHB23]. industry [BAM^{+}24, GHB^{+}24, HS24, AEN^{+}23, ASH^{+}23, MDDZ21]. infant [QJZ^{+}20]. infection [GPRM21]. infectious [GW22, GDCGVG20, LBY^{+}20]. infer [ZLM^{+}23]. Inference [CMGS22, BEM^{+}24, JKS20a, Kol22, LZ20a, LLZ20, LDX^{+}23, MTA^{+}22, MAQ^{+}20, RSFB23, Tao23, UYH21, YYW^{+}24, YLTH22, ZZB^{+}24]. Inference-aware [CMGS22]. InferFair [PP24]. inferring [uHA20, JVH^{+}20]. Influence [WSXL21, LWLH20, MLC^{+}20, XZ20]. influencer [WM21]. influences [LZL^{+}20]. influential [QMCX20, XWM20]. influenza [KCY^{+}21]. influenza-like [KCY^{+}21]. infodemiological [GDCGVG20]. Informatics [PRPPFRL20]. Information [ArMA^{+}21, BJW22, CSD^{+}23, DHA^{+}20, GZPZ20, LZ20b, LDW^{+}21, NMR21, BLH^{+}24, CFC^{+}20, DDM21, Dut22, GW20, HJX^{+}23, HGY^{+}22, JYSH23, LHL20, LP21a, LQG^{+}23, LLWJ24, LLW^{+}22b, LZS^{+}22, PSAL20, RNA21, SMKC20, SQ22, SKX^{+}20, TD21, VMCM^{+}20, WDHY20, WC22b, WZW^{+}23, WJC^{+}24, WSJ^{+}21, XFJ^{+}20, XWG^{+}21, YHC20, YCG^{+}20, YLGZ21, Yan21, YZZ^{+}23]. Information-Centric [DHA^{+}20, NMR21]. informational [AP20]. informative [CABB20]. Informed [PGCB23]. infos [XCW20]. infrared [HWH^{+}23b, wZcZN^{+}19, wZcZN^{+}20]. Infrastructure [CA21, LXL^{+}23, BBB^{+}20, DCC22, GVCUGF20, HTXW21, KS24, MBD21, SAAEK22]. Infrastructure-level [LXL^{+}23]. Infrastructures [ZTP20, BLT^{+}24, BdL20, CBC^{+}20, FGB21a, FGB21b, JLS^{+}23, LFYH22, PJBB20, VI21, VEH^{+}23, WDSK21, LKE22]. infrequent [CCML20]. Inhibition [JLC^{+}20, YWG^{+}19, YWG^{+}20b]. INHIBITOR [WGW^{+}21]. injection [HAH^{+}23, TWY^{+}23]. injury [Bo19, Bo20b]. inland [KBG20]. inline [ZZF^{+}24]. innovation [PPM24, YPEK23]. innovative [RBLD21]. Input [KWL^{+}23, LCH^{+}22, MGX^{+}23]. Input-Driven [MGX^{+}23]. inputs [BLMT20, CPYY23, PCCX21]. inquiry [LZW21]. insert [WGS24]. insertion [WZXX21]. insider [MLWA20, RSQS21]. insiders [SD24]. insights [DGT24]. inspection [CP22, DBC24]. Inspiration [DA22]. inspired [JYSH20, RSBM20, RSQS21, VP20, ZLML20, SUKN22]. instability [Bo19, Bo20b]. Instance [AD21, MCT^{+}22, PJLL23, PYL22, TPF^{+}20, WLLY20]. Instance-based [PJLL23]. instances [BMM^{+}24, SPL22, ZA22]. instantaneous [WHF^{+}23]. institutions [KBTT20]. instruction [hAS24]. instruction-level [hAS24]. insurance [ZZB^{+}22]. Integer [XZTC22]. Integrated [LLZ^{+}22, ANA24, ÇYZZ21, GSKS20, HCB^{+}20, KAJ^{+}24, MISB22, MRR^{+}20, MS24, RNV^{+}21, TBG^{+}20, WLW^{+}21, WKW^{+}22, YK20a, ZZG^{+}24]. Integrating [BBM^{+}20, WLY^{+}20, YPX^{+}20, BGNBH^{+}20, BCSS20, JLS^{+}23, OGO^{+}20, PBC^{+}22, PP20, WC20]. Integration [CA21, SKH20, UYH21, XCZ^{+}23, ACI^{+}23, APC^{+}20, BBD^{+}21, BDT21, DMPS23, FGP20, LKS^{+}21, LYH^{+}21, MBZ^{+}21, MGS21, PCVN21, RHK^{+}23, SGP^{+}20a, SMBB^{+}24].

Integrity [ZBF22, CFM^{+}22, FWP21, FMN^{+}20, GBK20, HOV20, SIG24, WWZ^{+}20, XWW^{+}20, YLZL21]. integrity-protecting [YLZL21]. Intel [GDS^{+}20]. Intelligence [AT20, DIB20, DAAW20, GCPM22, HIU^{+}22, IDM^{+}20, LWHW22, OCMJFB^{+}23, Sha20, SRP20, UUH^{+}22, WYHM21, WPPA22b, WPPA22a, XWG^{+}21, ZWH21a, ZBF22, ZZD22, AGdS^{+}21, ASL22, AM21, AMZZ23, ADP^{+}22, BSH^{+}24, BCT^{+}21, DSW^{+}20, EKK23, GSSB24, HY21, KCKK24, LEWC24, LCLA21, MGX^{+}23, RAA^{+}20, TPN^{+}21]. intelligence-based [EKK23]. intelligence-enabled [BCT^{+}21]. Intelligence-of-Things [WPPA22b, WPPA22a]. Intelligent [AMB^{+}21, AAS^{+}20, BAMR20, BDF^{+}22, ETH20, LQYL21, LZA^{+}20, LCLW21, LCLW24, MLX23, MWL^{+}20, NAC^{+}22, NRMI20, QKG20, SRP20, SZL^{+}21, UUK^{+}21, WLX^{+}24, ZXW^{+}20, ASHO20, AESI^{+}21, BQI^{+}20, CXHS20, CMA^{+}22, DHC23, DS23, DAT21, DLHD22, EBA^{+}22, HLP21, HTAY21, ICGGAR22, LHD^{+}20, LLZL21, LY21, LLW^{+}23a, LCZB21, LYBS21, MDT^{+}20, MASRAM^{+}22, QC21, RAA^{+}21, RLL^{+}22, SACN^{+}21, SCXZ23, SMC^{+}20, SACW23, SLX^{+}24, TRB^{+}23a, UPK^{+}23, WC22b, WX23, WZS^{+}23, Wei21, YLL22, YPX^{+}20, YWH^{+}21, YYB^{+}21, YXS^{+}23b, ZWC^{+}22, ZLXH20, ZLS23, ZYL^{+}22]. inTensive [ASDLS23, HLH^{+}20, HHLZ20, RRD21, SEKS^{+}20, TSB20]. Intent [FTS^{+}24, MGC23]. Intent-based [MGC23]. Intent-driven [FTS^{+}24]. Inter [XWK21, GLP^{+}24, PZHD20, LCB^{+}23]. inter-class [GLP^{+}24]. inter-microcell [PZHD20]. Inter/intra [XWK21]. Inter/intra-category [XWK21]. interacting [HHLZ20]. interaction [ABGDT23, AR20, CPH^{+}22, DBBP24, GLF^{+}22, JRW^{+}20, KHHV21, XLCB20, YZZ^{+}23, YGE21, YMY21]. interaction-based [YMY21]. interactions [BPLFRL20, CYZ^{+}22, PFGDM22]. Interactive [ZCS20, CCHD21, DLdAR23, DDT^{+}23, DML20, DLL20, WHC^{+}22, XPT^{+}22]. intercomparison [YCS^{+}20]. intercontinental [BBB^{+}20]. Interest [AAG23, BeKTK^{+}20, QWR^{+}20, ABA24, Dut22, GMI22, HMLS20]. Interface [ZJL^{+}22, FFAW20]. interfaces [JM20, MJW23]. interference [BMBC20, FWY^{+}22, LQW^{+}20, LFHS23, PZHD20, ZPLQ20]. interferon [WCWC19, WCWC20]. International [DPG20]. Internet [DP20c, DP21a, DP21b, FZT^{+}23, HN23, HPY20, LCH^{+}23, LWZ^{+}23a, LMW^{+}24, Sun20, WSL^{+}23, XZC^{+}24, BKV22, FYHZ24, uRLW^{+}21, AVK^{+}23, ARbL^{+}20, ABMM18, ABMM22, AND24, AYA^{+}23, AALEF20, ARHT20, AUJW22, AWMM^{+}23, ADRP23, ANA24, Ano24s, ADAR22, BCT^{+}21, BOL^{+}20, BCSS20, CPPY24, CRdRR^{+}22, CPJ20, CDP20b, Dao23, DP19, DLHD22, ETH20, FHGF20, GS20, GWY^{+}20, GCCMK^{+}20, GWW^{+}22, GLW^{+}20, Gul22, HZPS21, HRX^{+}21, HH22, HRGL21, JZK^{+}21, JKS20b, JM20, KMR^{+}22, KMS20, KSLC21, LAS20, LZH^{+}20, LLFQ21, LLW^{+}22a, LWL23b, LZ20b, LDW^{+}21, Liu24, MO24b, NBJ21, NAC^{+}22, NKG23, NLSY20, NMR21, PAS^{+}20, RJA^{+}22, RWG21, SN23, SAD24, SGP^{+}20b, STS^{+}20, SBF^{+}21, SCXZ23, SMS22, SM20, SD22, SDKM20, SK21b, TLX^{+}23, TLL^{+}24, TDL^{+}21, TG20, TWM^{+}23, TSKK23, TDC^{+}20, TK24b, UAACH21, VMV20, VSPM21, WCHA20, WXZZ22, WPPA22b, WYGP21, XGS^{+}20, XXY^{+}23, XY20, XCZ^{+}23]. Internet [YNN^{+}20, YZW^{+}23, YAZ^{+}20, Yu21, ZC22, ZCQ^{+}23, Zha20, ZWH^{+}20, ZA20, ZWH21a, ZZZ^{+}22, ZLS^{+}22a, ZNX23, ZZG^{+}24, ZXW^{+}20, ZCL24, Zhu20, Zhu21, ZLWH23]. Internet-of-Drones [NKG23]. Internet-of-Forensic [KSLC21].

Internet-of-Medical-Things [WCHA20]. Internet-of-Things [ARHT20, SM20, TG20, XXY^{+}23]. Internet-of-Things-enabled [AUJW22]. Internet-scale [ZC22]. interoperability [BVCH22, DSRG22, PP22, PMMG^{+}20]. interoperable [CDV^{+}24, CCBFI^{+}23, LLY^{+}20, PSvL^{+}20]. interoperating [WHA^{+}20]. interplay [RBA^{+}22]. interpolation [VFOV20]. interpretability [MXL^{+}20]. Interpretable [MASRAM^{+}22]. Interpretation [CDRS20]. Interpulse [OMPSPL20]. intersection [LYY^{+}20b, SCXZ23, FLH^{+}24]. interval [LDGS20]. interval-based [LDGS20]. Intervals [OMPSPL20, JHK20]. interventions [FHGF20]. intra-category [XWK21]. Intraoperative [ZYX^{+}23]. Intravascular [HZL^{+}21]. introducing [MHA^{+}24]. Introduction [SHW24]. Intrusion [GCPM22, HTAY21, RCdF^{+}21, VSPM21, ZMJ^{+}22, AHH20, AEN^{+}23, AADM21, ABL22, ADAR22, CHL23, DBD^{+}23, DCD^{+}22, ICBB20, JZL^{+}24, KP22, KTC23, NK20, PCK20, RLCB22, TSM24, WGW^{+}21, YLZ^{+}24, YYL22, ZLLD21, ZLPZ21, ZL22, ZZG^{+}22, ZCS20, LWHW24]. intrusions [JSV21]. intrusive [TTTH20]. invariance [HZL22]. invasive [DVV^{+}20, MSK^{+}21, RGPGP24]. inverse [rHZmH^{+}24, MDG^{+}22]. Investigating [KAF^{+}23, KAK20, THB23]. Investigation [LZL^{+}20, LXL^{+}21, Sha20, UAACH21, WLC^{+}20b]. investment [CL21, Sun20]. invocations [BVFGSF20]. IoF [KSLC21]. IoMT [KSA^{+}20, KMR^{+}22]. IoMT-based [KSA^{+}20]. IoNT [AT20]. IoP [AM21]. IoT [ABMM18, ABMM22, AYA^{+}23, ARHT20, ASDLS23, ETH20, KF23, MTHA24a, SN23, SME^{+}21, UPK^{+}23, XZH^{+}23, LJL^{+}21, MBF^{+}20, SN21, AMM^{+}22, AK20, AMB^{+}21, AHH20, ABAD22, ASA^{+}20, ACI^{+}23, AKF^{+}20, AL20, BEM^{+}24, BSF^{+}20, BTF^{+}21, BBTC20, BKHD20, BBP22, BJW22, BCSS20, CVdRA^{+}20, CCC^{+}21, CA21, CSB23, CYZ^{+}22, CZZ^{+}23a, CDV^{+}24, CSD^{+}23, CATD^{+}24, CHJK22, CL20b, CDP20c, DFF^{+}23, DCC22, DSC24, DATAA20, DAT21, DMPS23, DCZ20, DHA^{+}20, DK24, DLL20, EAA21, FBL^{+}20, FFM^{+}20, GFPB23, GKB^{+}20, GRN20, GNA^{+}21, GMB23, GK21, GLWP20, GPC21, HBH21, HMH^{+}22, HBF24, HMT^{+}20, HHW^{+}22, HGK20, HGWC23, HKS23, JYSH20, JLK22, JCX^{+}21, KF22, KS24, KBTM21, KSLC21, KJ24, LHC21, LMNC22, LY23, LDLS20, LHL20, LWW^{+}22, LRCL24, LBDP23, LCFM20, LYY^{+}22, LIP^{+}24, LMCSE20, MHH^{+}20, MTHA24b, MK20, MSMJ22, MRS^{+}22, MTD^{+}24, MFE^{+}20]. IoT [MRR^{+}20, MS24, MJW^{+}22, NADY20, NAT20, NSJ^{+}24, NNUV20, NGC24, OMSL20, OIG24, PKB22, Par20, PKLC22, PKLC23, PGCB23, PCG^{+}20, Pła24, PDJS22, PLMZ23, QCW^{+}24, RHK^{+}23, RDR^{+}24, RMI22, RRAB24, RZIX20, RLL^{+}22, dRRCGdC20, RKG20, RAB23, RLCB22, SME^{+}19, SCBP24, SVFdA20, SYYuR22, SMT^{+}24, SHB22, SEKS^{+}20, SDVC22, STS^{+}20, SMBB^{+}24, STH^{+}20, SRP20, SMC^{+}20, SRA^{+}22, SK23, SKTP24, TRB^{+}23a, TRB^{+}23b, TLMP20, TDS^{+}22a, TGAP20, TBG^{+}20, TCBF24, TPN^{+}21, UYH21, VAKB23, VS20, WC20, WWL21, WX23, WWS23b, XKK20, XCB^{+}20, XCL^{+}20, YMAAH22, YHW^{+}20, YLF^{+}23, YXS^{+}23b, ZLM^{+}23, ZHD^{+}20, ZZJC21, ZXY^{+}21, ZZQ21, ZPQH21, ZL22, ZYL^{+}22, ZXZ^{+}23, ZLG^{+}24, GZZG24]. IoT-based [DHA^{+}20, GNA^{+}21, HBF24, LHC21, TGAP20, UYH21, VS20]. IoT-BSFCAN [DATAA20]. IoT-Cloud [DATAA20]. IoT-edge [WX23]. IoT-edge-cloud [ZLG^{+}24]. IoT-enabled [GKB^{+}20, GPC21, RRAB24, SKTP24, ZXY^{+}21]. IoT-friendly [MTHA24a, MTHA24b]. IoT-systems [BCSS20]. IoT/Cloud [DAT21]. IoT/Cloud-based [DAT21]. IOTA [MRMB24, SM20]. IoTMD [ADP^{+}22]. IoTMD-enabled [ADP^{+}22]. IoTs [TG20, JPMR21]. IP [UUK^{+}21]. IPFS [PAP^{+}20]. IPS [LWHW24]. IPsec [UCO20]. IPU [RKI^{+}23]. iRECOVer [MSMJ22]. iris [AMM^{+}20]. irregular [LYC^{+}22]. irrigation [CMA^{+}22]. ISAG [GPC21]. isogeometric [GBP23]. Isolate [ZY20]. Isolating [AB20]. isolation [PP24, SSM22, WWS^{+}23a, ZZL^{+}22]. isomorphic [CYG22]. ISP [KCB20]. Issue [AFMG^{+}22, AHWB20, BDF^{+}22, ETH20, GZPZ20, TDC^{+}20, VKP22, WPPA22b, WPPA22a, YMS20a, YDK20, AMB^{+}21, Ano24s, ACY20, ACDY21, AMNZ20, AKPT20, BBSB21, BDT21, BDFR22, DPG20, ICW21, LLWJ24, LWHW22, MM21b, MBJ^{+}20, PPM24, VPBE22, WD24, YMS20b, ZA20, ZWH21a, ZTP20, dSGST21]. Issues [HH22, KSS^{+}20, KHH21, CCH21, CCDR22, HHH22, LH21, MRS^{+}22, RHK^{+}23, YNN^{+}20]. Italian [FIABC^{+}20]. Italy [MBD^{+}20]. item [JT22, LQG^{+}23, GVŠ22]. itemset [LMZ^{+}22]. itemsets [Den20]. iteration [BQC23]. Iterative [EMHE21, CWM^{+}20, SGBC^{+}20, Wu22].

J [RSFB23]. J/inference [RSFB23]. JACC [SHF23]. JACC-FPGA [SHF23]. Jaccard [SHF23]. January [Ano21v, Ano20n, Ano22p, Ano23o, Ano24m]. Janus [GRG20]. Java [BPLFRL20, ORPPG20]. JMI [GDEBC20]. Job [BHSH22, KPL22, LYS^{+}20, KMCJ20, LCY^{+}23b, LEXH20, MAM^{+}24, TLW^{+}24, WF21]. Jobs [VSV^{+}23, BPUW24, FPL24]. JOI [HBH21]. Join [GGCIV20, JJZ^{+}23, SHB22, SZGB24, YZW22, JJZ^{+}23]. Joint [HBH21, MMBD20, WXZZ22, YZL^{+}23, YXL^{+}20, ZWZ^{+}24, AOKÖ24, ATZP21, AYY^{+}20, PBSS24, TDL^{+}21, TPD^{+}20, WML^{+}21]. JointCloud [SYW^{+}23]. Jointly [NPNC23]. July [Ano20s, Ano21n, Ano22n, Ano23p, Ano24k]. jump [TLKX21, ZDC22]. junctions [BN21]. June [Ano21o, Ano20o, Ano22m, Ano23s, Ano24n]. Jupyter [CAC^{+}22]. JVM [MVLJ21].

K-means-based [ZP22]. K/MAPE [JRW^{+}20]. Kafka [MLZ^{+}22, RCP24]. Kafka-ML [MLZ^{+}22]. Keccak [VDSB22]. Keep [LLC^{+}23]. KELM [CFD^{+}20]. Kernel [WWS^{+}23a, CCP^{+}20, MJZC21, SuRMA^{+}23, ZGN^{+}20, PSH^{+}20]. kernels [LCH^{+}22, PSH^{+}20]. key [AND24, BGCL20, CDG^{+}20, GCM21, GMAL23, HLW^{+}23a, LYP^{+}24, SK20b, WCXW22, XCB^{+}20, YZW^{+}23, YLH^{+}23]. keystroke [TA21]. keyword [HDD20, JVH^{+}20, JGL^{+}20, LLS24, LYY^{+}22, SZM22, VI21, XYH^{+}24, YLH^{+}23]. keyword-based [JVH^{+}20]. kidney [MSLJ20]. Killer [YLL22]. Kinect [ÇYZZ21]. Kinect-based [ÇYZZ21]. kinematics [YJB^{+}21]. Kinesiology [RSBM20]. Kinesiology-inspired [RSBM20]. kinetics [YJB^{+}21]. Kit [BNX22]. klepto [BHL^{+}20]. Kmeans [NTY^{+}21]. knapsack [FW22, TPD^{+}24]. Knapsack-based [TPD^{+}24]. knative [TK24a]. Knights [GDS^{+}20]. KNN [DCD^{+}22, SXF22]. know [AAR^{+}20]. Knowledge [AFMG^{+}22, AD21, BFG^{+}22, DP20a, DT21, DMPS23, LZL^{+}21, MMM^{+}20, MBZ^{+}21, SD20, YPZ^{+}24, BLH^{+}24, BGNM20, BEKF21, CKFT20, DOR^{+}21, GSDGP21, HTLM21, HDD20, LSN^{+}20, MLP^{+}21, NFK^{+}20, NZY^{+}23, PSMF21, dAMVULM20, RBLD21, RAS^{+}20, ST20b, SDV^{+}21, TD21, THVL24, WZZD23, WLLF20, WCW^{+}21, WML^{+}23, XLS^{+}21, XCZ^{+}22, YZL^{+}24, YZW22, ZGK^{+}22, ZLP^{+}22, CPJ20]. knowledge- [PSMF21]. knowledge-aware [WML^{+}23]. knowledge-based [GSDGP21, NFK^{+}20, WLLF20]. knowledge-driven [RAS^{+}20]. Knowledge-guided [YPZ^{+}24]. Known [HMT^{+}20]. Kohonen [JSV21]. KRAS [JLC^{+}20, YWG^{+}20b, YWG^{+}19]. Kriper [RRDSAML23]. KubeAdaptor [SXZZ23]. kubeFlower [PUMN^{+}24]. Kubernetes} [QSZ^{+}24, DEJ20, PUMN^{+}24, PK24, SXZZ23]. Kubernetes}-based [QSZ^{+}24, PUMN^{+}24].

LA-MDPF [AAG22]. label [HXWX23, ORPPG20, WLZ^{+}20, XTL^{+}23, YGS^{+}22]. Labeled [WLS^{+}24]. labels [DLC^{+}22, ZHZS23]. laboratories [LZHL23]. labs [RGDMMR^{+}23]. LAFED [JZZ^{+}23]. Lagrange [SS22]. Lambda [MGZ^{+}20]. land [WCP23, XWLC20]. Landing [GDS^{+}20]. landmark [JSP23]. landmark-assisted [JSP23]. landslide [YLM23]. Language [DOR^{+}21, RPF21, CZ20, CGMT20, CPJ^{+}21, DBBP24, FSD^{+}20, GRG20, OCBO20, PMCP20, QG20, YGR21]. languages [ABL23, PP22, RMA^{+}20]. Large [AMNZ20, BHL^{+}20, LXH^{+}21, MM21b, TLC^{+}20, CLLCK20, CABB20, CCL^{+}21, DRD20, DZB23, GCM21, GW20, IAM^{+}22, JPJO22, KHL20, LOR22, LZW^{+}22, LCY^{+}23a, LHW20, MMAH22, MAQ^{+}20, ONK^{+}20, QMCX20, RZIX20, SGDK^{+}21, SDGCB^{+}20, SEL^{+}22, SW20, TDL^{+}21, TGAP20, WF21, XWM20, YSL^{+}24, YNK^{+}20, ZC22, ZJW^{+}20, dSGST21, FAŞ^{+}20]. Large-Scale [MM21b, BHL^{+}20, LXH^{+}21, TLC^{+}20, CCL^{+}21, GW20, IAM^{+}22, LZW^{+}22, LCY^{+}23a, ONK^{+}20, SGDK^{+}21, SEL^{+}22, SW20, TDL^{+}21, TGAP20, WF21, dSGST21, FAŞ^{+}20]. larger [HXWX23]. laryngeal [MSK^{+}21]. lasso [AOSA20b]. latencies [YPL24]. Latency [HLK^{+}23, PFP^{+}22, PFS^{+}23, AALEF20, BGMK22, CWYG23, FSP^{+}24, JLP^{+}21, KAF^{+}23, LLZL21, SKX^{+}20, SHY^{+}21, ASDLS23]. latency-aware [LLZL21]. latency-based [JLP^{+}21, ASDLS23]. Latency-constrained [HLK^{+}23]. Latency-Critical [PFP^{+}22, PFS^{+}23]. latent [BK20, GOA23, LH20, SD20]. Latin [GDCGVG20]. Lattice [WCXW22, LYP^{+}24, ZWWC21]. Lattice-based [WCXW22, LYP^{+}24]. law [EKJ^{+}20, LPT22, SNM^{+}20]. law-enforcement [SNM^{+}20]. lawful [LLCH21]. Layer [AHL^{+}23, BBN^{+}20, DLHD22, FWP21, FRGBHPPS23, HLW^{+}23a, JZL^{+}20, KS24, MPS21, MAB^{+}20, POBK21, SYG^{+}20, SAF23, XKK20, YL20b, ZWM^{+}23, IMuI^{+}21]. layer-2 [SAF23]. Layer-wise [AHL^{+}23, JZL^{+}20, ZWM^{+}23]. Layered [CH24]. layers [CVdRA^{+}20, TG20]. layout [SCK^{+}22]. lazy [ZSZX24]. LBP [LRML21]. LBSN [RGRV^{+}20]. LDC [ZWH^{+}20]. LDoS [TTD^{+}20, TCW^{+}22]. leaf [TLN23]. leakage [ATK^{+}22, CPYY23, DSW^{+}20, QLJ21, RPŠŠ23]. leaning [YBX^{+}23]. LearnChain [SMO^{+}24]. learned [CLLCK20, LSL^{+}20, LF21, WSL21, Wan21]. Learning [ABC^{+}24, AWMM^{+}23, ASM^{+}22, AFMG^{+}22, AAH^{+}23, BEM^{+}20, BP20, BOL^{+}20, CMM21, dFCC23, DOR^{+}21, GPR^{+}24, GPRM21, GDS^{+}20, HAVK22, JAS^{+}20, KOM^{+}22, LWNH22, LYC^{+}22, LXL^{+}23, LCO^{+}23, LLT20, LCLW24, MBC22, MIIN23, MRMB24, NCLP21, PLS^{+}23, PAM21, QCY^{+}21, SGS24, SuRMA^{+}23, SPRA21, SRA^{+}22, SLH^{+}24, SZS^{+}21, UUK^{+}21, VSV^{+}23, WMU^{+}24, WLLY20, WGGB24, WZJ^{+}22, ZrHhH^{+}23, ACN^{+}21, AMM^{+}22, AAG22, AHSH22, AMM^{+}20, ARIB22, ABMO24, AYHA20, ASL22, AHL^{+}23, AB24, ABL22, AEZ22, AÇP22, BEM^{+}24, BAİP24, BFG^{+}22, Ben23a, BBN^{+}20, BMM^{+}24, CXS^{+}22, CZCH24, CCP^{+}20, CdD20, CKL20, CLZ^{+}20, CQS^{+}23, CHS^{+}23, CZH^{+}24, CYWS24, CWL20, CYH20, CKFT20, DCGM20, DAM^{+}21, DPLV23, DBD^{+}23, DDT^{+}23, DFZ^{+}20, DML23, DLZ^{+}23, DK24, DLW^{+}23, DLC^{+}22, EUEU24, EL23, FXZZ24, FW22, FRGBHPPS23, GS20, GZT^{+}21, GMM22, GRN20, GMH20, GDP20, GCCMK^{+}20, GMB23, GSI22]. learning [HAK^{+}21, HN23, HJX^{+}23, HIU^{+}22, HAqDE23, HCG^{+}23, HJI24, HCL^{+}22, HLL^{+}24, HLK^{+}23, HQLH20, rHZmH^{+}24, HL24, IA23, JTGH21, JMA^{+}21, JHB22, JZZ^{+}23, JLW^{+}23, JYSH23, JCX^{+}21, JZL^{+}24, JSP23, KJYC23, KDX^{+}24, KAK^{+}23, KMK^{+}23, KS24, KWL^{+}23, KMS20, KJ24, LWHW24, LSN^{+}20, LOR22, LYYG20a, LDLS20, LSH^{+}20, LYYG20b, LZK21, LZS^{+}21, LZ21a, LLW^{+}22a, LFM^{+}22, LZW^{+}22, LWL23b, LQG^{+}23, LWS^{+}23a, LWL23a, LLW^{+}23b, LCY^{+}23a, LWS^{+}23b, LLW^{+}23a, LTX^{+}24, LYG^{+}24, LLZ^{+}21, LZTM24, LLZ^{+}24a, LBY^{+}20, LGL^{+}20a, LYS^{+}20, LLZ20, LJW^{+}20, LSB21, LYH^{+}21, LH21, LWJ^{+}21, LZS^{+}22, LWF^{+}23, LLLS24, LMZL24, LHTSM^{+}23, LIP^{+}24, LPT22, LAT^{+}20, LSGA20, LGL^{+}20b, LZL^{+}24c, LP23, LCH^{+}24, LCLW21, MSLJ20, MXL^{+}20, MYL^{+}23, MLN24, MSLP24, MK20, MK21, MKB23, MMZI22, MFE^{+}20, MBD21, MM23, MOW^{+}20, MA22, MDC^{+}24, MPP^{+}21, MNA^{+}23, MKC^{+}21, MSC^{+}23, MKBT24, NTI24, NKB^{+}20, Ngu24, NVS^{+}22, OOB^{+}21, OHÁV20, PJJ^{+}22, PBS23, Par22, PKLC22]. learning [PLHC24, PUMN^{+}24, PCCX21, PBK^{+}22, QHNL21, QCG^{+}24, QHCH24, QPL^{+}22b, QCWY23, RHK^{+}23, RSR^{+}24, RAA^{+}24, RRAB24, RYL20, RLL^{+}22, RSL24, RBMCLH22, RD23, RSB20, SG20, SKB20, ŠTI24, SW22, SMO^{+}24, SYG^{+}20, SCX^{+}24, SEL^{+}22, SAT20, SMS^{+}24, SP23, STS^{+}20, SB24, SBF^{+}21, SP22, SPWL23, SSS21, SXW^{+}22, SPL24, SDV^{+}21, SK21b, SZO^{+}20, TLC^{+}20, TLL^{+}24, TWL23, TSM24, TWM^{+}23, TBG^{+}20, TK24b, UPD^{+}20, VGM24, WMU^{+}23, Wan20, WCHA20, WLY^{+}20, WLLF20, WLL21, WYHM21, WML^{+}21, WYWS22, WXZZ22, WZHX23, WSL^{+}23, WWZW23, WS23, WLZ^{+}23, WYDB24, WLL24, WYJ^{+}24, WXX^{+}24, WYZ^{+}24, WGS24, WFA20, WZS^{+}22, WXS^{+}22, XW23, XTL^{+}23, XRZ^{+}24, XPT^{+}22, XLZ^{+}22, XJL^{+}24, XZC^{+}24, XXL^{+}24, XWZM24, XY20, YRV^{+}22, YJH^{+}20, YCG^{+}20, Yan21, YLM23, YWDC23, YZL^{+}24, YPL24, YCYO23, YYB^{+}21, YNVRPD23, ZHD^{+}20, Zha21, ZYF^{+}22, ZZB^{+}22, ZL22, ZGK^{+}22, ZZG^{+}22, ZWM^{+}23, ZLZ23a, ZZZ^{+}23, ZLZ^{+}23b, ZL23]. learning [ZLL^{+}24b, ZWZB24, ZKGB20, ZLP^{+}22, ZGW^{+}23b, ZDLD24, ZLC^{+}21, ZHZS23, ZWCS23, ZCWC20, ZCS20, ZLH^{+}24, ASH^{+}23, Liu23, MZL^{+}22, MBB24, NRBC23, PK24, SHB22, ZWX^{+}23, ZCL24]. Learning-based [GPRM21, QCY^{+}21, WGGB24, AHSH22, BEM^{+}24, DBD^{+}23, IA23, JZL^{+}24, LTX^{+}24, LLLS24, MSLJ20, MK20, MK21, OOB^{+}21, RRAB24, SEL^{+}22, TSM24, WLZ^{+}23, WYZ^{+}24, WFA20, XY20, ZZG^{+}22]. learning-ready [PBK^{+}22]. Learningo [KPA24]. least [CFL^{+}20]. LEChain [LLCH21]. ledger [Liu24]. ledgers [MCF20, SK20a]. left [YCG^{+}20]. legacy [OWK^{+}23]. legal [LHD^{+}20]. legitimate [FAAS20]. LENTO [ASDLS23]. lesion [WLZ^{+}20]. Less [LRML21, VCG^{+}23, GSARS20, RZA21, ZWZ^{+}21]. Less-is-Better [LRML21]. Lessons [CLLCK20]. Level [ASSG22, MOU^{+}21, PRF20, ACF^{+}21, hAS24, BSOK^{+}20, BCB^{+}20, CFD^{+}20, CZZ^{+}23b, DAM^{+}21, DGY^{+}22, KSC20, LXL^{+}23, MVLJ21, NLO^{+}20, NN21, ORPPG20, SJVRS22, SGL^{+}20a, SACW23, SGLB22, TWY^{+}23, XHW20, YLSL22b, ZYL^{+}20, ZrHhH^{+}23, ZDZ21]. leveling [YYZ^{+}24]. levels [MRD^{+}20]. Leveraging [CCHA22, DBBP24, QRS^{+}21, YWDC23, CS24b, LZ21b, LSL^{+}20, Liu21, LCCP21, ZHGX20]. lexical [dAMVULM20]. LGAN [ZXX23]. LGAN-DP [ZXX23]. LGIEM [MLC^{+}20]. libraries [YPEK23]. library [HZT^{+}22]. Life [CK20, ADdMM20, CK24, JAS^{+}20, PSvL^{+}20]. life-cycle [ADdMM20]. lifecycle [HCB^{+}20, TCBF24]. Lifelong [LGJ^{+}23, HLL^{+}24]. lifestyle [PRPPFRL20]. Lifetime [NTA^{+}22, Gul22]. LiGen [CCC^{+}24]. light [DAMS23, PLL^{+}24, THVL24]. light-weight [THVL24]. lighting [GHB^{+}24]. Lightning [RKG20]. Lightweight [CQA^{+}24, DAT21, FGB21a, FGB21b, GJC^{+}20, MRR^{+}20, RMI22, SYYuR21, WHC^{+}22, ZYY^{+}23, ZG24, ANA24, BTF^{+}21, FFAW20, GLF24, GWW^{+}22, GSG^{+}23, HZL22, JZK^{+}21, JZZ^{+}23, LQS^{+}20, LGL^{+}20b, MAS23, MCV23, RMA^{+}20, RAB23, RLCB22, SYHX23, TLN23, WXC^{+}24, XWW^{+}20, ZWH^{+}20]. like [CREE^{+}24, KCY^{+}21, YYZ^{+}24]. likelihood [FAAS20]. limb [LSH^{+}20, YJB^{+}21, ZWL21]. limit [BQC22]. limitation [LCL22]. Limitations [YNN^{+}20]. limited [XLG^{+}23, XZD^{+}21, ZLZ^{+}20a]. linear [DSW^{+}20, MMAH22, MNSL22, TBB^{+}23]. lines [CFD^{+}20]. linguistic [GSDGP21]. linguistics [GDCGCPVG21]. Link [TD21, CDX^{+}23, HYL^{+}20, HRY^{+}21, LZ20a, LQML22, MAQ^{+}20, SMC23, TYR22, XCW20, XZ20]. link-dominance [SMC23]. link-weighted [XCW20]. Linked [DP20b, DML20]. Linking [TD21]. links [ACA^{+}23, UCO20]. Linux [CMMST20, PBM^{+}22, XCS^{+}22]. Linux-based [CMMST20, PBM^{+}22]. literacy [YPEK23]. literature [BHH22, CRdRR^{+}22, KPGD24, MBB24, MKK^{+}20, OLP23, SDVC22]. Live [ASASA^{+}20, MK22, BEM^{+}20, FCGPSG^{+}21, SS22, TPF^{+}20]. liver [wZcZN^{+}19, wZcZN^{+}20]. livestock [GAP24]. Living [GK21, LJL^{+}21, RGDMMR^{+}23]. LMMA [LLC^{+}22]. lncRNA [ABGDT23]. Load [ATC^{+}24, BHL^{+}21, GOA23, MOU^{+}21, ZZL^{+}22, ZYW24, AZA23, BMM^{+}24, CHG^{+}20, DGY^{+}22, JZK^{+}21, KF22, KF23, Kha24, KNV20, LLP^{+}20b, LCL22, MAS23, MYM^{+}21, MO24b, PR20, dRRCGdC20, SHST20, SVN20b, SZGB24, WGGB24, YWH^{+}23, ZCLL22, ZHJW20]. load-adaptive [DGY^{+}22]. Load-aware [ZYW24]. load-balanced [MO24b]. Local [CYWS24, HAVK22, ABMESM18, ABMESM22, BR20, EL21, GZF^{+}23, KSSR20, MLC^{+}20, MGS21, PLBOC20, Par20, SZL^{+}21, WDHY20, XW21, YLGG21, YPX^{+}20]. localisation [SDKM20]. locality [BQC22, BLMT20, GDGK20, GMT23, GOA23, KW20, WWW^{+}24]. locality-aware [GOA23]. localization [ASM^{+}22, GKA^{+}21, JQZ^{+}22, JSP23, KAH^{+}23, dTGC20, PGHS20, TK24b, WYWS22, WJC^{+}24, YXS23a, ZLWH23]. Locally [BLT^{+}24]. Location [AYY^{+}20, EL21, FLH^{+}24, HMLS20, LHL20, LRCL24, OMSL20, RAS^{+}22, TDL^{+}21, THB23, THVL24, WHZ^{+}20, WG21, YVSG22, YYW^{+}21, ZGL^{+}23, ZKGB20]. location-based [HMLS20, OMSL20, YVSG22]. log [GLP^{+}24, LLD^{+}21]. log-based [GLP^{+}24]. LogETA [GLP^{+}24]. logging [ADAHA^{+}21]. logic [CCTZ22, UYH21, ZNZ^{+}23, ZYL^{+}22]. logical [YFL^{+}24]. logistic [WHF^{+}20, WZH^{+}22, ZSL^{+}23b]. logistics [NSJ^{+}24, ZWC^{+}22]. LogNADS [LLD^{+}21]. logs [CDP20a]. LogSC [WCD^{+}22]. Long [ACF^{+}21, BMZdP21, XCL^{+}20, CHW^{+}20, GMFC23, HMSA^{+}23, MZLT21, SHZMA21, WLL22, XWZM24]. long-running [SHZMA21]. long-term [XWZM24]. long-vector [GMFC23]. Look [YYXZ23]. Look-ahead [YYXZ23]. looking [CC21]. lookup [ZC22]. loop [WXS^{+}22]. Loops [ZFMB20]. LoPrO [AYY^{+}20]. LoRa [ABAD22, HN23, OCSCB22]. LoRa-enabled [HN23]. LoRaMoto [CMF^{+}21]. LoRaWAN [RPŠŠ23]. loss [DLW^{+}23, FAAS20, FLG^{+}20, GDP20, LCZB21, RZIX20]. lossless [HIMM20]. lossy [KCKK24]. Lotaru [BLT^{+}24]. loud [MRM^{+}24]. Louder [AHN21]. Low [EUEU24, SAD24, BR20, CPT^{+}20, DLR23, FCGPSG^{+}21, GMGV^{+}22, LDWZ20, LHW20, LYY^{+}20b, MMPV22, MR23b, Par20, PPGS20, RAB23, SKX^{+}20, TDL^{+}21, YPL24]. low-availability [GMGV^{+}22]. Low-cost [SAD24, FCGPSG^{+}21, TDL^{+}21]. low-end [RAB23]. low-energy [DLR23, LDWZ20]. low-rank [MMPV22]. lower [YJB^{+}21, ZWL21]. lower-limb [YJB^{+}21]. LPP [FLH^{+}24]. LPP-BPSI [FLH^{+}24]. LPWAN [HLW^{+}23a]. Lr [SHY^{+}21]. Lr-Stream [SHY^{+}21]. LSMD [BR20]. LSTM [AYA^{+}23, AACJ23, JQZ^{+}22, MMU^{+}21, WHC^{+}24, ZZ24]. LSTMs [SK20b]. LTE [AKF^{+}20, CDY^{+}20, RRAB24]. LTE-U [CDY^{+}20]. LTE-WiFi [AKF^{+}20]. LTP [ZY21]. lung [JLC^{+}20, YWG^{+}19, YWG^{+}20b]. Lyapunov [QPL22a].

M

[BJ22, MGW23, MGW23, MIIN23, OÖA22]. M-Learning [MIIN23]. m-RENDEZVOUS [OÖA22]. M/M/ [MGW23]. M2FaaS [PRF22]. M2M [GPC21, YHC20]. M2M-device [YHC20]. MAB [IA23]. MAC [ABAD22]. Machine [AFMG^{+}22, BP20, DOR^{+}21, GPRM21, JCP^{+}20, KOM^{+}22, LYC^{+}22, LLT20, LSGA20, LCLW24, PKLC22, PAC^{+}22, RNA^{+}22, SPRA21, UUK^{+}21, VSV^{+}23, WMU^{+}23, AMM^{+}20, AYHA20, ATZP21, AEZ22, BFG^{+}22, BPUW24, CCP^{+}20, CCL^{+}20, CLL^{+}23, CKFT20, DLH^{+}20, DHD20, GZT^{+}21, GDP20, GCCMK^{+}20, GSI22, HAK^{+}21, HCL^{+}22, JMA^{+}21, KAK^{+}23, LWHW24, LSN^{+}20, LOR22, LZLY20, LFM^{+}22, LLZ^{+}21, LYS^{+}20, LJW^{+}20, LAT^{+}20, LFHS23, LCC^{+}24b, LCLW21, MBD21, MOW^{+}20, NKB^{+}20, NVS^{+}22, OHÁV20, OPOG23, PJJ^{+}22, PBK^{+}22, QG20, QHC24, RSR^{+}24, SKB20, STS^{+}20, SBF^{+}21, SPWL23, SDV^{+}21, VPA20, VGL23, WHW20, WXS^{+}22, XZK^{+}20, YRV^{+}22, YBX^{+}23, YWH^{+}23, YGR21, ZHD^{+}20, ZLW^{+}22, ZLC^{+}21, ZCS20]. machine-learning [SKB20]. machine-readable [QHC24]. machines [GMH20, MGW23, OPOG23, RZIX20, SS22, YYKK20, LWJ^{+}21]. macro [YGE21]. macro-cells [YGE21]. Macroscope [YVW^{+}20]. MAG [PB23]. MAG-D [PB23]. MagBox [LLC^{+}23]. magic [LLC^{+}23]. magnetic [Gur21b, HZX^{+}19, HZX^{+}20, Liu23]. MAGNETO [Gur21b]. main [YYKK20]. maintenance [BDG23, KCJ23, YWH^{+}21]. Major [PFP^{+}22, AOSA20a]. makespan [LCH^{+}21, ZTB23]. Making [AHMW23, CSH^{+}23, FGG^{+}21, GSKS20, XLL^{+}20b, FMM^{+}20, GSSB24, GK21, IA24, Kol22, LMO^{+}22, Sun20, YZX^{+}23]. maladaptive [YRV^{+}22]. Malaysia [RYL20]. Malicious [BdL20, WSD^{+}22, ACG^{+}20a, FPH^{+}21, JLW^{+}23, MMC22, MRMM20, PSMF21, SB24, SSS21]. Malware [SK21b, UUK^{+}21, AYA^{+}23, AAS^{+}20, ATT^{+}20, ARA^{+}22, ARA^{+}23, CMMST20, DFF^{+}23, DAM^{+}21, IuRJ^{+}21, JWC22, KZG^{+}22, LWJ^{+}21, LLF^{+}23, MWK^{+}21, MS24, NADY20, TGJ^{+}20, TLJ^{+}22, VMM20b]. malware-resilient [MWK^{+}21]. manage [RCR21]. managed [RPP^{+}20]. Management [ADP^{+}22, DP20c, DP21a, DP21b, GZPZ20, KVCY20, AB19, ABGMC19, AB21, ABGMC21, AND24, AK20, ADdMM20, ABAJ20, ASBT20, AÖ24, BLT^{+}24, BDM^{+}20, BJW22, CdST^{+}20, CA21, CDC^{+}24, DP19, FYHZ24, GSCP22, GMAL23, HAB^{+}20, HMY^{+}23, HSGY20, HBF24, Hu21, HKS23, IA20, JJZ^{+}23, JL21, JKS20b, JKS20c, KSS^{+}20, KAF^{+}20, KYPJ20, KCB20, KA21, KA24, KSE24, LM20, LAHN22, LYYG20a, LYYG20b, LLZL21, LLCH21, LDWZ20, LGCY22, LH21, MLX23, MZX^{+}24, MDT^{+}20, MSS^{+}24, MISS22, MRM^{+}24, MMBD20, MBD21, MLZ^{+}23b, MOU^{+}21, Ngu24, OLP23, OLLP24, OOB^{+}21, PGCB23, PS20, RBLD21, RSR^{+}24, RPP^{+}20, RKM23, SNMWC21, SZVVB^{+}23, STK20, SCÁB20, SKH20, SFC23, TDMC23, TPD^{+}20, VPBE22, WMD^{+}20, WYD20, WXZZ22, WFLL22, WWS20, XZJ^{+}20, YMAAH22, YJLC20, YZL^{+}24, YGP^{+}24, YK20a, ZCLL22, ZFZS23, ZA22, uRLW^{+}21, UPK^{+}23]. manager [FCOJFM21]. Managing [CBN^{+}20, GFM^{+}20]. mandatory [AHN21]. MANET [SN21]. manipulating [LZS^{+}21]. manipulation [TTZ^{+}21]. manipulator [HS21]. manipulators [SHKW23]. Manufacturing [CHS^{+}24, LLLS24, OLP23, OLLP24]. Many [CCL23, CIB^{+}20, GW20, KYPJ20, XLX^{+}21]. many-core [CIB^{+}20, KYPJ20, XLX^{+}21]. Many-objective [CCL23, GW20]. many-task [CCL23]. Map [RK20, BDK^{+}20, CZGS20, IHA^{+}20, JSV21, KSSR20, LWLH20, NQB^{+}23, QJS^{+}21, LLP^{+}20b]. Map-Balance-Reduce [LLP^{+}20b]. Map-optimize-reduce [RK20]. MAPE [JRW^{+}20, JRW^{+}20]. mapping [BSOK^{+}20, CL20b, JPW20, JPJO22, MMMZ20, NQH^{+}20, RGPGP24, SW20, VCK^{+}20, WHF^{+}20, YXL^{+}20]. mapping-based [WHF^{+}20]. MapReduce [LLP^{+}20b, MSTN21, ST20b, TM20]. March [Ano21r, Ano20t, Ano22x, Ano23t, Ano24l]. margin [DLW^{+}23]. marine [HAqDE23, QZZ^{+}24]. maritime [CSAT24, BAM^{+}24]. market [BMBE20, HGY^{+}22, MMM^{+}20, TBO20]. market-based [HGY^{+}22]. markets [MXW22]. Markov [AAG22, CKL20, DFF^{+}23, PCG^{+}20]. MARN [SJQ20]. MARS [KSC20]. MAS [MRM^{+}24]. MAS-loud [MRM^{+}24]. mask [Pan20, LYH^{+}21]. Massive [Hu21, LY23, LXY21, LZC21, RLZW21, ZCF21, ABB^{+}21, BBM^{+}22, LLP^{+}20a, LWCC23, SHB22, Xu21]. Massive-scale [Hu21, LXY21, LZC21, RLZW21, LWCC23, Xu21]. master [DEJ20]. Mastering [BBD^{+}21]. MAT [WZW^{+}23]. MAT-transformer-based [WZW^{+}23]. Matching [ASAM20, AD21, GVŠ22, XQW^{+}24, CZGS20, GHW^{+}20, KZG^{+}22, LRCL24, LZS^{+}24, LGT^{+}20, YLC23, YJQ^{+}23]. matching-based [LRCL24]. materials [WXD^{+}23]. maternal [LLFQ21]. mathematical [BEON24, TJG^{+}20]. Matrix [SYXL22, CFL^{+}20, LLY^{+}20, RNV^{+}21, TJG^{+}20, TBB^{+}23]. matters [MNSL22]. matting [WLYL20]. maturity [HZLH21]. Max [MÖ22, MÖ24a, Ski20]. Maximal [ARIB22, HPY20, SSV24]. maximization [WSXL21]. maximize [NT22]. Maximizing [XGY^{+}23, SHST20, XWD20]. MaxMin [ST20b]. MaxMin-scalable [ST20b]. May [Ano21s, Ano20v, Ano22o, Ano23r, Ano24o]. MCDM} [IA24]. MCFT [SK21b]. MCOTM [QCW^{+}24]. MCS [CZH^{+}24, YLX^{+}23]. MD [MEL^{+}23]. MD-Bench [MEL^{+}23]. MDPF [AAG22]. me [CC21, HJW^{+}20]. meaningful [AOF21]. means [SLA^{+}23, THVL24, ZP22]. measure [ZXY^{+}21]. Measurement [DSRG22, GBC^{+}24]. measures [MOU^{+}21]. Measuring [YLGZ21]. MEC [LQYL21, MISS22, WYX^{+}23a]. MEC-assisted [MISS22]. Mechanism [WSD^{+}22, AVK^{+}23, ASA^{+}20, BBTC20, CCW^{+}20b, DG21, Dho20, FNRP20, FZC^{+}20, GFZ21, GMGV^{+}22, HIMM20, HSR^{+}22, HRGL21, JZZ^{+}23, KTC23, LTXL22, LGW22, LLZ^{+}24b, LYW23, MHL20, MA24, QWR^{+}20, RMA21, RLQ^{+}21, RHWY23, RLCB22, SYW^{+}23, TLL^{+}24, TLKX21, WZB^{+}20, WYHM21, WXZX23, WWZ^{+}20, WWY^{+}24, XGX20, YZC^{+}20, YLX^{+}23, ZTC20, ZYX^{+}20, ZLLD21, ZWZ^{+}23, ZXX23, ZWX^{+}23, ZFZS23, ZLC^{+}21, ZLS^{+}20]. mechanisms [AL20, KCKK24, LKS^{+}21, NMRK21, THA^{+}24, ZLS^{+}22a]. Media [YVW^{+}20, ALS21a, CLC21b, DFG^{+}21, KK22, LVNCC21, MJZC21, dAMVULM20, RAA^{+}20, UCR21, VMCM^{+}20, Yu21]. mediation [QG20]. Medical [ANA24, KMR^{+}22, NAC^{+}22, XY20, YNN^{+}20, BBB^{+}20, CXWY21, LYFZ20, LFM^{+}22, LHXL22, LHW^{+}23, LGYC20, LZW21, MLWA20, Pan20, PBSS24, QJS^{+}21, QZZH21, SZS^{+}21, TAM21, WCHA20, WWS23b, XLL^{+}20b, YJF^{+}20, YZS^{+}21, ZZB^{+}22, ZXW^{+}20, ZWL22, ZHZS23, ZDZ21, dAdSM^{+}22]. medicine [AACJ23, LBY^{+}20]. medium [SKX^{+}20]. medoid [GCM21]. medoid-based [GCM21]. Mellitus [WZH^{+}22, OOB^{+}21]. meltdown [AGYS20]. MeltdownDetector [AGYS20]. Memory [ACF^{+}21, KKT^{+}23, LDWZ20, XCL^{+}20, CHW^{+}20, CDC^{+}24, EGDT20, GMT23, ISUC22, JA20, JCK24, LLC^{+}22, LAHN22, LRML21, LGM^{+}21, LL20, MSZ^{+}20, MZLT21, MMPV22, SXF22, TWI20, WLL22, WX24, XHW20, YZC^{+}20, YYW^{+}24, YK20a, YYZ^{+}24, YYKK20, ZSZ^{+}24, ZZT^{+}22]. Memory-aware [LDWZ20, TWI20]. memory-disk [YK20a]. memory-efficient [YYW^{+}24]. memristor [JPW20]. mental [ASYL22, BDFR22, SSMdS21, UCR21]. mentor [ÇYZZ21]. mer [EGD24]. Mergesort [AAT^{+}24]. merging [AHL^{+}23]. meritocratic [FFM^{+}20]. Merlin [PBK^{+}22]. Mesh [MO24b, GZF^{+}20a, LGW^{+}21, NNH^{+}20]. mesoscale [ZZD22]. message [HBH21, JVH^{+}20, WZTL20]. messages [BCM20]. messaging [FGG^{+}23]. Meta [LCO^{+}23]. Meta-Learning [LCO^{+}23]. metadata [CATD^{+}24, POMK20]. metagenomics [VMM^{+}20a]. Metaheuristic [TPD^{+}24, ATZP21, GGK20, KV22, RMC20, SSV24]. Metaheuristics [DA22, SJVRS22]. metamodel [MDDZ21]. Metamorphic [DML20]. Metapath [PJLL23]. Metaverse [KTC23, CXHC23, DHC23, HTGW^{+}23, SP24, ZSL^{+}23a, ZPK^{+}23]. metaverse-oriented [ZPK^{+}23]. meteorological [SGL^{+}20a]. Metering [CA21, SAAEK22]. MeterSSD [HZL22]. Method [FQH^{+}24, ARIB22, ATZP21, ASB^{+}23, BHL^{+}21, CQS^{+}23, CSAT24, CSC23, CYH20, CDP20b, DHC23, EGD24, FRGBHPPS23, GZT^{+}21, HWH^{+}23a, HCWD21, HTXW21, HLL^{+}20, JJZ^{+}23, JQZ^{+}22, JWC22, JPJO22, KHHT21, LYY^{+}20a, LCW^{+}20, LFZJ21, LZC^{+}23a, LJ24, LMCSE20, MECRFD20, MM21a, MGX^{+}23, NHTH20, PSC^{+}21, QMCX20, QL22, QZZ^{+}24, RZH21, RJM^{+}21, SB24, SPWL23, TLJ^{+}22, TCW^{+}22, WLLF20, WG21, WLL22, WZW^{+}23, WGL^{+}24, WWC^{+}24, WLLC20, WLR21, XZK^{+}20, XCL^{+}20, YFQ^{+}22, YPL24, YLY^{+}23, YK20a, ZDC22, ZHD^{+}20, ZXY^{+}21, ZLPZ21, ZNZ^{+}23, ZZZX22]. Methodologies [BBSB21]. Methodology [CGM^{+}23, FMN^{+}20, ASA23, ACI^{+}23, DVEE^{+}20, LZB20, SDV^{+}21, TCMV20, ZYL^{+}20, HCK20a]. methods [AB24, AÖ24, CLC21a, CdD20, HBSG21, KW20, KGO^{+}20, KSS^{+}21, LHLC23, LYGF21, MR23a, YDL^{+}20, ZT22a, ZTB23]. metric [PCCX21, See20]. Metrics [MOU^{+}21, BSOK^{+}20, GPGG23, Hu21, LCL^{+}20]. METSM [JXYC24]. MEV [PJL^{+}24]. MF [TTD^{+}20]. MF-Adaboost [TTD^{+}20]. MFE [WCP23]. MFE-ResNet [WCP23]. Micro [FLF^{+}21, PWH^{+}22, BPCM21, HS24, PSC^{+}21]. micro-blogging [BPCM21]. micro-clouds [PSC^{+}21]. Micro-clusters [FLF^{+}21]. micro-service [HS24]. microarchitecture [TWY^{+}23]. microarchitecture-level [TWY^{+}23]. microbes [YGS^{+}22]. microblogging [VCM^{+}21]. microcell [PZHD20]. microcontrollers [NRBC23]. microgrid [KBTM21, PWH^{+}22, RMC20]. micropayment [RKG20]. Microservice [BMM^{+}24, HTXW21, KS24, SXC^{+}24, WPX^{+}23]. Microservices [TCBF24, BBF^{+}24, CWYG23, DCC22, LBDP23, MJSW21, PKB22]. microservices-based [PKB22]. microsimulation [PABBA20]. middleware [BVCH22, CCW^{+}20a, CBN^{+}20, GSARS20, MFMSG20]. Migration [DSC20, GZXH24, HJW^{+}20, NT22, PPSC23, QCW^{+}24, SJQ20, SS22, ZCLL22, ZYW24, MK22]. migrations [LLC^{+}22]. MIGTNet [PJLL23]. military [CPS^{+}23]. mimicked [AR20]. MIMO [LY23]. min [MÖ22, MÖ24a, Ski20]. Mind [HJW^{+}20]. miners [FZC^{+}20]. Mini [FXZZ24]. minimization [WYD20]. minimize [AALEF20]. Minimizing [Hu20, ZTB23]. minimum [MNFQ24]. Mining [Den20, LLP^{+}20a, LSS^{+}22, LWH^{+}22, PRPPFRL20, YMY21, AJJ^{+}21, CATD^{+}24, GZF^{+}20a, HNV^{+}20, IB20, JT22, KHHV21, KYY^{+}20, LMZ^{+}22, MSG^{+}20, RK20, RLZW21, SCL20, TOM^{+}20, TCW^{+}22, WFL^{+}20, WYWS22, WCP23, XFJ^{+}20, YNK^{+}20, ZCF21]. MiniPFL [FXZZ24]. minor [AOSA20a]. MIOV [LYFZ20]. Mirror [GdOAO20]. misogyny [GDCGCPVG21]. MISOSE [YL20b]. missing [LHC21, SCL20, TLM21, ZT22a, ZZ24]. Mission [MAB^{+}20]. Mitigating [ACG^{+}20a, GSSB24, WXX^{+}24, QHW^{+}20]. Mitigation [BeKTK^{+}20, FZ20, DG21, GMP20b, Kho21a, KCB20, MSLP24, RWJ^{+}20, TTTH20]. MitM [QHW^{+}20]. mix [SZZY22]. mixed [CCDR22, FHGF20, JJZ^{+}23, TBB^{+}23, YPL24, XZTC22]. mixed-criticality [CCDR22]. Mixed-Integer [XZTC22]. mixed-precision [TBB^{+}23]. mixture [ASA23, MSG^{+}20, NTI24]. ML [KTIB22, LLF^{+}23, MLZ^{+}22, SCBP24]. ML- [SCBP24]. ML-based [LLF^{+}23]. ML-centric [KTIB22]. ML/AI [MLZ^{+}22]. MLS [ASSG22]. MLS-ABAC [ASSG22]. MMAP [BJ22]. MMDS [ZHLL24]. MMSparse [TJG^{+}20]. mmWave [PMT22]. MobiGyges [FLG^{+}20]. Mobile [AAP21, HGY^{+}22, LZL^{+}24b, qLhZ20, PLL^{+}24, WSC^{+}23, XLLL20, ZKGB20, ACBT23, AAS^{+}20, AB20, ARA^{+}22, ARA^{+}23, CHG^{+}20, CLY^{+}20, CWL20, DATAA20, DCZ20, FLG^{+}20, GBH^{+}23, HCG^{+}23, HLH^{+}20, LMNC22, LLF^{+}23, LMZL24, zLsZjX20, LHY^{+}20b, LGL^{+}20b, LFHS23, LGT^{+}20, MGB24, MSY20, MWL^{+}20, PPA^{+}24, QL22, RSR^{+}24, RJA^{+}22, RSL24, SLX^{+}24, Tao23, TDMC23, UJHN20, WLP^{+}20, WZL^{+}20, XRHS21, XXY^{+}23, ZLZ^{+}20a, ZLL^{+}24a, IA24, SJQ20]. mobile-device [CHG^{+}20]. mobile-edge [MWL^{+}20, RSL24]. mobile-fogging [DATAA20]. Mobility [AOKÖ24, KKL^{+}24, NNPP23, QCW^{+}24, DPPGCCA23, EL23, LM20, MMZI22, RPdVR20, WXZZ22, WCL^{+}24]. Mobility-as-a-Service [NNPP23]. Mobility-aware [KKL^{+}24, QCW^{+}24, MMZI22]. Mobility-driven [AOKÖ24]. MobiPCR [LLF^{+}23]. mobiTopp [BBM^{+}20]. modal [HMLS20, KHB23, KSDR21, LZ21b, MMP^{+}23, WWH^{+}21, WJC^{+}24]. Mode [AAH^{+}23, LYGF21, LZJ^{+}20, XYL^{+}20, ZYX^{+}23]. Model [AKJJ20, ADMG20, BNA^{+}21, CHS^{+}24, GL20, LXL^{+}21, LCB^{+}23, MMKS22, MMK^{+}20, MIMS20, MZZ20, QCG^{+}24, SYXL22, WCD^{+}22, WTL^{+}20, ZWCS23, ACF^{+}21, ASA23, ASYL22, AWMM^{+}23, AMZZ23, AMT^{+}21, AAG^{+}20, BÖ20a, BEON24, BH21, CL20a, CKL20, CSY^{+}20, CWM^{+}20, CLQS20, CKZ^{+}22, CF20, DT21, DMPS23, ESSS^{+}21, GFZ21, GZL^{+}22, GW22, GBC^{+}24, HZLH21, HQLH20, HZS^{+}23, ICBB20, JKS20a, JHK20, JLT^{+}21, JYSH23, JXYC24, JMHB24, KOM^{+}22, KLW^{+}21, Kon21, LOH^{+}23, LLW^{+}20, LHY^{+}20a, LLP^{+}20b, LP21a, LHXL22, LWL23a, LZP23, LZF^{+}24, LLZ^{+}21, LBY^{+}20, LWLH20, LZZ^{+}20, LH20, LL20, MYT^{+}21, MYL^{+}23, MGW23, MSR20, MSKG21, MS20, MRR^{+}20, NTA^{+}22, dSOFC^{+}23, PJL^{+}24, PP24, PABBA20, PRD^{+}22, PMMG^{+}20, QNM24, QC21, RCLEB20, RNA21, RGP^{+}22, RKI^{+}23, SMU^{+}21, ŠTI24, SMKA23, SDGCB^{+}20, SB24, SZdLZ22, SS21, SYXW21, SDA21, Sun20, TWL23, TLN23, TA23, TDS^{+}22a, VPA20, WLZ^{+}20, WGG^{+}20]. model [WCY^{+}20, WSL21, WYWS22, WLL22, WGS24, WWS23b, WZH^{+}22, XLL20a, XWK21, YJH^{+}20, YHW^{+}20, YGP^{+}24, ZYF^{+}22, ZLW^{+}22, ZHH^{+}23, ZFZS23, ZZ24, ZLP^{+}22, ZXL^{+}20, Zhu20, Zhu21, KWL^{+}23]. Model-based [AKJJ20, WCD^{+}22]. Model-Driven [CHS^{+}24, MZZ20]. model-free [ICBB20]. Modeling [CYZ^{+}22, FTM20, GIPS20, MhCEANSM20, QNRA23, TSX^{+}24, WBR20, YJF^{+}20, ZHX^{+}20, ZCWC20, AGdS^{+}21, ABL23, ACG^{+}20b, BYW^{+}21, BFM23, DBC24, DC21, KLA22, LZ21b, LDW^{+}21, LLT22, MASRAM^{+}22, PLMZ23, SFC23, Tao23, TSKK23, WLC23, YRV^{+}22]. Modelling [RMC20, YGE21, ARB20, KHB23]. models [ACI^{+}23, BMS20, BOM^{+}22, BMBC20, CCML20, CCHA22, EELB21, GSSB24, GW20, HMO^{+}20, KHHV21, KHB20, LHLC23, LH21, LJ24, LSMT^{+}21, dAPHOMPJ20, MRM^{+}24, MOW^{+}20, ODET21, PP22, QHE^{+}20, RHJ20, SXC^{+}24, WD24, YCS^{+}20, ZLM^{+}23, dlVGSB^{+}20]. Modern [OCMJFB^{+}23, SK20b]. modernization [CIB^{+}20, WLYL20]. modification [XXL^{+}24]. modified [KS24, MSLJ20, Wan20]. Modular [SPRA21, DLdAR23, ONK^{+}20]. module [DJP^{+}24]. molecular [LGZ^{+}24, MEL^{+}23]. monetizing [BMBE20]. Monitoring [BQI^{+}20, DP20c, DP21a, DP21b, ACN^{+}21, AdSM^{+}22, AESI^{+}21, ACD^{+}20, AMZZ23, BBM^{+}22, BMD^{+}21, BSH^{+}21, CPT^{+}20, DP19, FGB21a, FGB21b, dRFRB24, GFPB23, HIMM20, JAAAZB20, KBG20, LZB20, MFE^{+}20, MBD^{+}20, RAS^{+}22, SAM^{+}24, SMC^{+}20, SSC^{+}20, VS20, VEH^{+}23, WC20, YXLB20, ZHD^{+}20, ZZJC21, ZT22a, ZLS23, ZHL24]. monocular [LTXL22]. monolith [PRF22]. monotonic [MM23]. Monte [MMAH22, SWL^{+}20, TBH23]. MOOC [SYXL22]. MOOCs [RVJMJ^{+}21]. morphable [LDD^{+}22]. morphing [GZF^{+}20a]. morphology [TJG^{+}20]. Mortadelo [dlVGSB^{+}20]. MOSES [FGG^{+}21]. MOSS [HCK20a]. moth [FW22, Kha24]. Motion [FS21, HYRZ20, XLS^{+}21, YLD^{+}23, ZWL22]. motional [LZ21b]. motivated [ABT20]. motivation [ZWX^{+}23]. mould [LCW^{+}20]. move [LSS^{+}22]. movement [CL20b, ZLZ^{+}20a]. movements [XYL^{+}20]. movie [BEKF21]. Moving [JAC^{+}23, SWW^{+}20, AM20, KHL20]. MPdist [LHC21]. MPdist-based [LHC21]. MPI [AHMW23, LGM^{+}20, MTA^{+}22, ZGW^{+}23a]. MPSoCs [TWI20]. MQDS [YSL^{+}22]. MQTT [MWS24]. MR [ST20b]. MR-I [ST20b]. MRI [GBdRACG20, HIU^{+}22, WLZ^{+}20]. MS [PK22]. MSGC [GPWL20]. MSIDN [BeKTK^{+}20]. MSPPIR [GXS22]. MSSN [ACC20]. MSSN-Onto [ACC20]. MT [LGZ^{+}24]. MT-3000 [LGZ^{+}24]. MTD [AKJJ20]. MudraChain [KBTT20]. Multi [AAG23, ASSG22, ASM^{+}22, AR20, BBB22, CHC^{+}20, CMM^{+}23, DK20, DK24, EJP22, GWZ20, GHEB^{+}18, GHEB^{+}23, GSKS20, GPWL20, HMLS20, HB21, JMHB24, KCR20, KWL^{+}23, KSDR21, LYYG20a, LZLY20, LHF^{+}20, LHD^{+}20, LYYG20b, LXC^{+}24, LLW^{+}22b, MMM^{+}20, MSKG21, MR23b, OÖA22, PSC^{+}21, PLBOC20, SMC23, ŠTI24, SuRMA^{+}23, SZM22, TLKX21, TWM^{+}23, WFL^{+}21, WLL22, WWW^{+}24, XGY^{+}23, XW21, YGS^{+}22, YJQ^{+}23, ZLLD21, ZLZ^{+}23b, ZSL^{+}23a, ZLZ^{+}20b, ZTB23, AZA23, ACF^{+}21, AAGX^{+}22, AK20, ASA^{+}20, BYW^{+}21, BRK24, BAMR20, BH21, CLV24, CKL20, CSB23, CZZ^{+}23b, CKW21, DBD^{+}23, DDM21, DEJ20, DZXS21, DHD20, DQBS20, DLW^{+}23, EGD24, FWP21, FZC^{+}20, FMN^{+}20, GFZ21, GKB^{+}20, GBH^{+}23, GMA^{+}22, HBEK20, HLP21, HYC^{+}21, HSGX22, HXL^{+}23, HRGL21, HZdLZ20, HYRZ20, HX21, Hu21, HZX^{+}24, IA24, IT20, JTGH21, KF22, KF23, KK20, KSC20, KHB23, KCJ23, LLW^{+}20, LXH^{+}21, LCH^{+}21, LZ21b, LWNH22]. multi [LFM^{+}22, LWL23b, LZC^{+}23a, LQYL21, LZL^{+}21, LF21, LZCH22, LWZ^{+}23a, LZHS24, LPL^{+}20, LZCGMVV20, MMP^{+}23, MLX23, MECRFD20, MVLJ21, MKB23, MRM^{+}24, MNSL22, MDC^{+}24, NBJ21, NTY^{+}21, NLS23, dSOFC^{+}23, OWK^{+}23, PCC21, PKR21, PLL^{+}24, PWH^{+}22, PMCP20, POBK21, PR20, PZLL21, QGH^{+}22, yQhJL20, QHE^{+}20, RAS^{+}22, ST20a, SCR20, SYG^{+}20, SNMWC21, SJVRS22, SP21, SCZ^{+}20, SYHX23, SSDC22, SXW^{+}22, SGLB22, TTD^{+}20, TLJ^{+}22, TDS^{+}22a, VG21, WLP^{+}20, WMNV20, WYS20, WYX^{+}23b, WYJ^{+}24, WJC^{+}24, gWLWZ21, XLH^{+}24, XRHS21, XLS^{+}21, XTL^{+}23, XYH^{+}24, XLL20a, XCZ^{+}22, YLH^{+}23, YPZ^{+}24, YZX^{+}23, YYW^{+}21, YYL22, ZWZ^{+}24, ZZLF21, ZLW^{+}22, ZLF^{+}23a, ZGW^{+}23a, ZYY^{+}23, ZWZB24, ZZ24, ZLP^{+}22, ZLL^{+}23, ZZZ21b, Zhu20, ZCWC20, ZA22, GXS22, IMuI^{+}21, LFYH22]. multi-accelerator [KCJ23]. Multi-Access [MR23b, BBB22, HXL^{+}23, LWNH22, SXW^{+}22]. multi-adversarial [DBD^{+}23]. Multi-Agent [ASM^{+}22, OÖA22, LYYG20a, LYYG20b, ZLZ^{+}23b, DHD20, DQBS20, MRM^{+}24, PR20, SP21, WLP^{+}20, WYJ^{+}24, ZZLF21]. multi-algorithm [EGD24]. multi-authority [CSB23, XRHS21]. multi-blocks [FZC^{+}20]. Multi-channel [MSKG21, GFZ21, LF21, WYX^{+}23b]. multi-class [JTGH21, LFM^{+}22, ZA22]. multi-cloud [KF22, KF23, OWK^{+}23, WYS20, ZWZB24]. multi-contracting [AAGX^{+}22]. multi-controller [AZA23]. multi-core [QHE^{+}20]. Multi-Criteria [GSKS20, PSC^{+}21, IA24, SSDC22, YZX^{+}23]. multi-density [CLV24]. multi-device [AK20, LZC^{+}23a]. Multi-dimensional [ZLLD21, SYHX23, ZWZ^{+}24]. multi-DNN [LCH^{+}21]. Multi-domain [AR20, TDS^{+}22a, WYJ^{+}24]. Multi-entity [GWZ20]. Multi-feature [WFL^{+}21, WLL22, Zhu20]. multi-features [TTD^{+}20]. multi-frontal [MNSL22]. multi-function [GMA^{+}22]. Multi-GPU [JMHB24, HYC^{+}21, ZGW^{+}23a]. multi-granularity [TLJ^{+}22]. Multi-Heterogeneous [LXC^{+}24]. multi-hierarchy [LZHS24]. Multi-hop [HB21, MR23b]. multi-horizon [HSGX22]. multi-HPC-clusters [LFYH22]. multi-image [LLW^{+}20]. Multi-information [LLW^{+}22b]. Multi-Input [KWL^{+}23]. Multi-keyword [SZM22, XYH^{+}24, YLH^{+}23]. multi-label [XTL^{+}23]. multi-language [PMCP20]. multi-layer [FWP21, POBK21, SYG^{+}20, IMuI^{+}21]. Multi-Level [ASSG22, CZZ^{+}23b, KSC20, SJVRS22, SGLB22]. multi-loss [DLW^{+}23]. multi-LSTM [ZZ24]. multi-master [DEJ20]. multi-metrics [Hu21]. Multi-modal [HMLS20, KSDR21, KHB23, LZ21b, MMP^{+}23, WJC^{+}24]. multi-model [ACF^{+}21, BH21, XLL20a]. multi-node [CKW21]. Multi-Objective [PLBOC20, DK20, EJP22, LZLY20, SMC23, WWW^{+}24, GBH^{+}23, HBEK20, HLP21, HRGL21, HX21, IT20, LZCGMVV20, MECRFD20, MDC^{+}24, PKR21, PWH^{+}22, yQhJL20, ST20a, SXW^{+}22, gWLWZ21]. multi-operator [SCR20]. multi-output [ZLW^{+}22]. multi-owner [NBJ21]. Multi-party [CMM^{+}23, MMM^{+}20, LWZ^{+}23a, XLH^{+}24, XCZ^{+}22]. Multi-Path [AAG23, SuRMA^{+}23, QGH^{+}22]. multi-perspective [LPL^{+}20]. multi-phase [NLS23]. Multi-population [CHC^{+}20, VG21]. multi-provision [LZCH22]. multi-replica [YYW^{+}21]. Multi-resident [LHF^{+}20]. multi-resource [BRK24, HZdLZ20, PZLL21]. multi-resources [dSOFC^{+}23]. Multi-robot [TLKX21]. Multi-round [XGY^{+}23]. multi-satellite [YPZ^{+}24]. Multi-scale [GPWL20, YJQ^{+}23, DDM21, DZXS21, HZX^{+}24, WJC^{+}24, YYL22]. Multi-search-routes-based [ZTB23]. multi-sensor [FMN^{+}20, RAS^{+}22]. Multi-server [ZSL^{+}23a, KK20, ZLF^{+}23a]. Multi-similarity [YGS^{+}22]. multi-site [PCC21]. multi-source [SCZ^{+}20, GXS22]. Multi-spectral [GHEB^{+}18, GHEB^{+}23]. multi-stage [ASA^{+}20, CKL20, GKB^{+}20, XLS^{+}21, ZLL^{+}23]. multi-stages [ZCWC20]. multi-step [ACF^{+}21, BYW^{+}21]. multi-strategy [PWH^{+}22]. Multi-Stream [KWL^{+}23]. Multi-task [DK24, LHD^{+}20, ŠTI24, LWL23b, MMP^{+}23, ZLP^{+}22]. multi-tasks [AK20]. multi-tenant [MLX23, MVLJ21, MKB23, SNMWC21, WMNV20]. multi-terminal [ZYY^{+}23]. Multi-type [TWM^{+}23]. multi-UAV [HYRZ20]. multi-user [LQYL21]. Multi-view [XW21, ZLZ^{+}20b, BAMR20, LXH^{+}21, LZL^{+}21, NTY^{+}21, PLL^{+}24, ZZZ21b]. multicast [UADD21]. Multiclass [PLHC24, TA23]. MultiCNN [PKLC23]. MultiCNN-FilterLSTM [PKLC23]. multicore [GOA23, LDD^{+}22, ZTQ^{+}20]. Multidimensional [AKA20, BBM^{+}22, FW22, ZYX^{+}20]. multidisciplinary [SRM^{+}23]. multidrug [MASRAM^{+}22]. multifaceted [CP22]. multigrid [TBA23]. multihop [FCGPSG^{+}21]. multilabel [LHTSM^{+}23]. Multimedia [BOL^{+}20, HOV20, SAM^{+}24, WZW^{+}20, ZHLL24, ACC20]. Multimodal [WYX^{+}23a, ESSS^{+}21, LJ21, MTCS22, ZZP^{+}23]. multinomial [ASA23]. Multiobjective [JXYC24, LCC^{+}24a]. multipath [YLSL22b]. Multiple [CCC^{+}21, LXL^{+}21, PYL22, QZZH21, RLQ^{+}21, WMU^{+}24, YhSL^{+}22, CLY^{+}20, CdO20, DDMP^{+}23, DZXS21, GBdRACG20, HKB^{+}24, KOM^{+}20, KGO^{+}20, LYYG20a, LYYG20b, LTX^{+}24, LCC^{+}24a, MRM^{+}24, dSOFC^{+}23, SDZ^{+}20, SS22, SPL22, WLZ^{+}20, WLLY20, WFL^{+}20, XZJ^{+}20, YZX^{+}23, ZDZ21]. Multiple-Feature-based [LXL^{+}21]. Multiple-instance [PYL22]. Multiplex [CZCH24]. multiplicative [MZLT21]. multiplier [HMA^{+}21, SS22]. multiprocessing [AFL23]. multiprocessor [JXYC24]. multiscale [JLT^{+}21]. multisensor [KLA22]. multisite [HdOP^{+}21]. MultiSLA [FNRP20]. MultiSLA-Aware [FNRP20]. multitasking [LCC^{+}24a]. Multithreaded [PFP^{+}22, PPGS20]. multitude [THVL24]. multitype [LXZ^{+}20]. multitype-users [LXZ^{+}20]. Multivariate [ZCQ^{+}23, CIJM20, GZZG24, HSGX22, PB23, PCACTMÁ24]. multiview [YCG^{+}20]. muscles [LZZX20]. mussel [MBD^{+}20]. mutation [DFZ^{+}23, JLC^{+}20, YWG^{+}19, YWG^{+}20b]. Mutual [ArMA^{+}21, ZN21, GLWP20, JYSH23, LLY^{+}20]. Mutually [BK20]. my [DCC22, RAN^{+}20, XZZ^{+}20b]. myocardial [WZC^{+}22].

N [ZXX^{+}20]. N-grams [ZXX^{+}20]. Named [ABC^{+}24, BeKTK^{+}20, IA24, NJB20, AAG22, AAG23, AAM^{+}24, IA23, RKM23]. Named-Data [ABC^{+}24]. Nano [Gul22, Gul22, SAD24]. Nano-Sensor [Gul22]. Nano-Things [Gul22, SAD24]. nanoprobe [wZcZN^{+}19, wZcZN^{+}20]. Napoli [MBD^{+}20]. NAS [LGM^{+}21]. national [OGO^{+}20, OCMJFB^{+}23]. native [GBP23, PMMSE21, VG21]. Natural [DOR^{+}21, QG20, YGR21]. Nature [SUKN22]. Nature-Inspired [SUKN22]. navigation [ASH^{+}23, Ano24s, IHA^{+}20, KAF^{+}23, SSDC22, TRB^{+}23a, YVSG22]. NB [LCFM20]. NCIP [SJD^{+}20]. NCSLab [LZHL23]. NDN [WC20, WWL21]. NDN-based [WC20, WWL21]. NDT [SMBB^{+}24]. Near [KCJ23, SW22, CF20, dRFRB24, ICBB20, ZT22a, wZcZN^{+}19, wZcZN^{+}20]. near-edge [CF20]. Near-optimal [KCJ23, SW22, ICBB20]. nearby [Gur21b]. nearest [LWY^{+}24]. NEC [GMFC23]. need [CPH^{+}22, VCG^{+}23]. needs [TLS^{+}21]. negative [HZX^{+}19, HZX^{+}20, LKS^{+}21]. Negotiation [TPD^{+}20, FZC^{+}20, STK20]. Neighbor [BAK22, LWY^{+}24]. neighborhood [DZXS21, SYG^{+}20]. Neo4j [ŠHDT21]. net [YJF^{+}20, YDL^{+}20, MJW^{+}24, ZG23]. net-based [YDL^{+}20]. Nets [YFL^{+}24]. Network [ASDLS23, AAH^{+}23, BKM^{+}22, Elg20, HTAY21, IuRJ^{+}21, KBG20, KAJ^{+}24, KSDR21, LKE22, LXL^{+}21, LY21, LZL^{+}23, LLD^{+}21, LSMT^{+}21, MDZ^{+}21, MR23b, MGM^{+}20, MAA22, NRBC23, NAC^{+}22, NMR21, QHC24, RKG20, SUKN22, SMBB^{+}24, TYR22, VP20, YZL^{+}20, ASYL22, AACJ23, ACG^{+}20a, ACA^{+}23, ABL22, ACM^{+}21, BÖE24, BWX20, BCM20, BAGRB^{+}20, BNX22, CLC21a, CZCH24, CVdRA^{+}20, CL20a, CKL20, CHW^{+}20, CL21, CCC^{+}23, CDC^{+}24, CZZ^{+}23b, CMGS22, CKW21, CF20, DVV^{+}20, DCD^{+}24, DBD^{+}23, Deh20, DDT^{+}23, DKG^{+}22, DBSL23, DLW^{+}23, Dut22, ENT^{+}22, FWX23, FIABC^{+}20, GWZ20, GWY^{+}20, GTG^{+}21, GZZG24, GPC21, HJX^{+}23, HAVK22, HY21, HZX^{+}24, HZL^{+}21, HYC^{+}23, IT20, JTGH21, JWYÍ21, JLK22, JCW^{+}23, JPW20, JWZ^{+}22, KNRI21, KAF^{+}23, KP22, KCB20, KMS20, LZ20a, LP21a, LZS^{+}21, LXH^{+}21, LZP23, LLWJ24, LCFM20, LJW^{+}20, LZW21, LC20, MSLJ20, MMP^{+}23, MK22, MCBGSL24, MGC23]. network [MXS22, MHA^{+}24, MS24, MNA^{+}23, MMU^{+}21, NK20, NZY^{+}23, OGO^{+}20, PCC21, PB23, PCK20, PLS^{+}23, POBK21, PP20, PBSS24, QMCX20, RWG21, RSL24, RKP^{+}21, RRDSAML23, SCGVP20, SMKC20, SAM^{+}24, SHB22, SVN^{+}20a, SHST20, SPWW21, SGL^{+}20a, SSV24, SHKW23, SSM22, TRB^{+}23a, TDLT20, TLX^{+}23, TDMC23, UADD21, URN^{+}20, UCO20, UUH^{+}22, UAS^{+}20, VCK^{+}20, Wan20, WFL^{+}21, WLL24, WCL^{+}24, WM21, WWS23b, WZX^{+}21, XGX20, XLMC22, XWD20, XTL^{+}23, XZ20, XLZ^{+}22, XY20, YLSL22b, YHW^{+}20, YZZ^{+}23, YYW^{+}24, YXL^{+}20, YYL22, YJQ^{+}23, YYY^{+}23, ZAH^{+}20, ZYX^{+}23, ZWC^{+}22, ZXX^{+}20, ZL21, ZZLF21, ZLLD21, ZGK^{+}22, ZZG^{+}22, bZSC^{+}23, ZG23, ZZG^{+}24, ZLL^{+}24a, ZMJ^{+}22, ZZPK21, ZCS20, GSG^{+}23, HTAY21, MGX^{+}23, PJLL23, RAA^{+}24, SMBB^{+}24, ZJL^{+}22, ZWX^{+}23, Zhu21]. Network-aware [MAA22, RKP^{+}21]. network-based [GTG^{+}21, LZW21, MXS22, QMCX20]. network-on-chip [Deh20]. Networked [MGM^{+}20, BCC^{+}22, GCT^{+}20, JAC^{+}23, KSS^{+}21, WMCH22]. Networking [ABC^{+}24, ALR^{+}20, BeKTK^{+}20, DHA^{+}20, HYL^{+}20, NJB20, AAG22, AAG23, AESI^{+}21, AAM^{+}24, dMBPdSC20, CDY^{+}20, MNA^{+}23, QWR^{+}20, QCY^{+}21, WLN^{+}21, YZJ^{+}20, ZWZ^{+}21]. Networks [AMR^{+}20, BSM20, BMZdP21, Gul22, HIdAR^{+}20, HDN^{+}20, HSGX22, HTLM21, HRY^{+}21, MBJ^{+}20, OMSL20, SKA^{+}20, SJQ20, TZW^{+}22, WHF^{+}20, WLX^{+}24, YLGG21, YMS20a, ABA24, ACF^{+}21, AAA20, ARIB22, AOF21, AOKÖ24, ABAD22, AQN^{+}20, AKA20, ASA^{+}20, AKF^{+}20, BZG23, BEB^{+}20, BSF^{+}20, BAK22, Ben23a, BR20, CCTZ22, CABB20, CECS20, CLZ21, CSD^{+}23, CDF^{+}22, CMA^{+}22, CSS22, DGK20, DCD^{+}22, FCGPSG^{+}21, FZC^{+}20, FRGBHPPS23, GLF24, GCM21, GMI22, GAT^{+}20, GMH20, GDGK20, GLM21, GZG20, GZF^{+}20b, GlRpG20, HRX^{+}21, HAB^{+}20, HLT^{+}21, HZZ^{+}20, HMLS20, IMuI^{+}21, IA23, JZL^{+}20, JYSH23, KMR^{+}22, KKL^{+}24, KMS20, LYKK22, LY23, LS23a, LWW^{+}20, LY21, LDLS22, LQG^{+}23, LWW24, LQYL21, LGL^{+}23, LQNW20, LZA^{+}20, LGW^{+}21, LTXL22, Liu23, LMCSE20, LVNCC21, LHY^{+}20b, LYW23, LZCGMVV20, LGT^{+}20, MCT^{+}22, MZLT21, MYL^{+}23, MMAH22, MhCEANSM20, MMZI22, MISB22, MJB22, MLWA20, MLZ^{+}23a, MKC^{+}21, NNH^{+}20]. networks [NPNC23, NTA^{+}22, NED^{+}20, PGHS20, Pła24, POBK21, QGH^{+}22, QLJ21, QZZH21, RSR^{+}24, RBSK23, RMI22, RKM23, RPdVR20, RLCB22, RSFB23, RHM20, SW22, SYG^{+}20, SNMWC21, SEL^{+}22, SMY20, SMBB^{+}24, SKTP24, SGDG23, SAF23, SK21b, SKX^{+}20, TLM21, TLKX21, TPD^{+}24, TYR22, URN^{+}20, UJHN20, VSPM21, WC22a, WYD20, WCY^{+}20, WLY^{+}20, WLAC20, WWF^{+}23, WLZ^{+}23, WYJ^{+}24, WGL^{+}24, WZC^{+}22, XWM20, XCW20, XLL^{+}20b, YHC^{+}22, YGE21, YMS20b, YK20b, ZN21, ZWL20, ZZ21a, ZZL^{+}22, ZLS^{+}22a, ZRH^{+}23, ZLZ^{+}23b, ZFZS23, ZLZ^{+}20b, ZCWC20, ZDZ21, ZHJW20, ACC20, FD21, IA24, MO24b, PJL^{+}24, ZWW^{+}23]. networks-based [DCD^{+}22]. Neural [HTLM21, IuRJ^{+}21, IT20, KSDR21, LZL^{+}23, LSMT^{+}21, NRBC23, NAC^{+}22, PJL^{+}24, TZW^{+}22, YGR21, YZL^{+}20, ZWW^{+}23, ACF^{+}21, CZCH24, CHW^{+}20, CLZ21, CL21, CDC^{+}24, CMGS22, CMA^{+}22, DGK20, DDT^{+}23, DKG^{+}22, DBSL23, FRGBHPPS23, GWZ20, GLF24, GMH20, GTG^{+}21, GZZG24, HRX^{+}21, JWYÍ21, JZL^{+}20, JPW20, JWZ^{+}22, KNRI21, LYKK22, LS23a, LP21a, LXH^{+}21, LHLC23, LWW24, LZW21, LTXL22, LMCSE20, LC20, MSLJ20, MZLT21, MMP^{+}23, MK22, MSKG21, MXS22, MS24, NK20, NED^{+}20, QLJ21, QHE^{+}20, RSFB23, SMU^{+}21, SPWW21, SK21b, UUH^{+}22, UAS^{+}20, Wan20, WLY^{+}20, WWF^{+}23, WWS23b, XLMC22, XLL^{+}20b, XY20, YLSL22b, YZZ^{+}23, YYW^{+}24, YJB^{+}21, YYL22, ZWC^{+}22, ZWL20, ZL21, ZRH^{+}23, ZLZ^{+}20b]. neuro [JKS20a]. neuro-fuzzy [JKS20a]. Neurofibromatosis [WTL^{+}20]. neutral [MR23b, ZL23]. neutrosophic [ABMMC18, ABM19, ABM21, ABMMC22]. news [ALS21a, HAK^{+}21, JWYÍ21]. Next [AAG^{+}20, MR23b, UADD21, AKF^{+}20, RSR^{+}24, RCdF^{+}21, THVL24, WD24]. Next-gen [MR23b, RSR^{+}24]. Next-generation [AAG^{+}20, RCdF^{+}21]. next-generations [WD24]. Nextflow [SHH23]. NFSP [HCG^{+}23]. NFTs [HMY^{+}23]. NFV-based [YNVRPD23]. nGIA [JZM^{+}22]. NIC [ZJL^{+}22]. NIC-QF [ZJL^{+}22]. NIDS [VGM24]. NiFi [MWS24]. NIMAR [LLC^{+}22]. NIST [OMPSPL20]. NLOS [PMT22, YXLB20]. NNs [LRML21]. no [AHN21]. NoC [TWI20]. NoC-based [TWI20]. node [BHH22, CKW21, HYWY22, HZT^{+}22, Liu24, LPT22, MLC^{+}20, Pła24, RMBMT21, WSXL21, ZJW^{+}20, ZWX^{+}23, STH^{+}20]. Node-RED [STH^{+}20]. Node.js [PRF22]. nodes [BR20, GMGV^{+}22, HZS^{+}23, LS23a, LWW^{+}20, NLS23, SN21, XWM20]. nodule [ZHP^{+}21]. Noise2Weight [ISD22]. noisy [DLC^{+}22, KIM^{+}24, ZHZS23]. NOMA [Kad20, PMT22, YL20b]. Non [DVV^{+}20, DDT^{+}23, DLL20, WXX^{+}24, XPT^{+}22, BJ22, CXS^{+}22, CMM21, DP20a, GGK20, HPP20, ICBB20, JLC^{+}20, KSSR20, LYYG20a, LYYG20b, LCH^{+}24, MSZ^{+}20, MZL^{+}22, MSC^{+}23, Par20, TTTH20, WHC^{+}22, YWG^{+}19, YWG^{+}20b, ZWL20]. non-canonical [GGK20]. non-cellular [Par20]. non-CQA [HPP20]. non-dominated [KSSR20]. Non-IID [WXX^{+}24, CXS^{+}22, CMM21, LCH^{+}24, MZL^{+}22, MSC^{+}23]. Non-interactive [DDT^{+}23, DLL20, XPT^{+}22, WHC^{+}22]. non-intrusive [TTTH20]. Non-invasive [DVV^{+}20]. non-orthogonal [LYYG20a, LYYG20b]. non-preemptive [BJ22]. non-small [JLC^{+}20, YWG^{+}19, YWG^{+}20b]. non-stationary [ICBB20, ZWL20]. non-volatile [MSZ^{+}20]. nondominated [CCL23]. nonlinear [RZIX20]. nonvolatile [WFLL22]. norm [CCL^{+}22]. normalising [BMS20]. normalization [HLK^{+}23]. Northwest [ZZD22]. NoSQL [CF21, SCL20, YGP^{+}24, dlVGSB^{+}20]. NoSQL-based [YGP^{+}24]. Note [Fae21, RCJZ20]. notice [AB21, ABGMC21, ABM21, ABMESM22, ABMM22, ABMMC22, Bo20b, DP21a, DP21b, GHEB^{+}23, HZX^{+}20, JLC^{+}20, LBJ^{+}24, LYYG20b, LCLW24, WWP20, WCWC20, YWG^{+}20b, YTQ20a, YTQ20b, ZMZ^{+}20, wZcZN^{+}20]. notification [AHN21]. Novel [CGM^{+}23, LLC^{+}22, PS20, WZH^{+}22, YLF^{+}23, ZLML20, ABM19, ABM21, AMM^{+}20, ALGMP^{+}21, ABAD22, AEN^{+}23, ATK^{+}22, ARA^{+}22, ARA^{+}23, BAMR20, DJP^{+}24, BAR21, CZZ^{+}23a, CYH20, GZT^{+}21, GPRM21, GMGV^{+}22, HCK20a, HCK20b, IA20, JJZ^{+}23, JZL^{+}24, KTC23, KV22, LSN^{+}20, LYFZ20, LDLS22, LHW^{+}23, LXZ^{+}20, Liu24, MM21a, MKK^{+}24, MRM^{+}24, NMRK21, NTY^{+}21, QG20, SYW^{+}23, WZB^{+}20, WXZX23, XFJ^{+}20, XYH^{+}24, Yan21, YZW^{+}23, YBX^{+}23, ZXY^{+}21, ZXX23, wZcZN^{+}19, wZcZN^{+}20, ZH20, JZM^{+}22]. Novelty [GST21]. November [Ano20q, Ano21t, Ano22q, Ano23v]. NS [SMBB^{+}24]. NSGA [GW20]. NSGA-III [GW20]. NTA [QHC24]. NTP [GdOAO20]. Nuclei [ZWH^{+}21b]. Nudge [WMD^{+}20]. NUMA [LLC^{+}22]. number [AAR^{+}20, BPUW24]. NVIDIA [GDS^{+}20]. NVM [BQC22, LDD^{+}22].

O [BBN^{+}20, EELB21, KHES21, LFHS23]. obfuscation [CSS22, NMRK21, TLJ^{+}22]. Object [AAA20, TQC20, ZLS22b, AGV23, BQC22, CTFW22, CSAT24, JLT^{+}21, WLLY20, WZX^{+}21]. Objective [PLBOC20, CCL23, DK20, EJP22, GBH^{+}23, GW20, HBEK20, HLP21, HRGL21, HX21, IT20, LZLY20, LTX^{+}24, LZCGMVV20, MECRFD20, MDC^{+}24, PKR21, PWH^{+}22, yQhJL20, RPP^{+}20, SMC23, ST20a, SXW^{+}22, gWLWZ21, WWW^{+}24]. Objects [Liu21, CPPY24, FMM^{+}20, GMI22, KHL20, ZZZ^{+}22]. oblivious [WGS24, YVSG22]. OBPP [GNA^{+}21]. Observability [SZM^{+}21]. Observation [MBZ^{+}21]. observatories [QRS^{+}21]. Observatory [FAŞ^{+}20, ZKD21]. Obstacle [HS21, MXW^{+}23, TLKX21]. obstacles [AM20]. Obtaining [GLF24]. Occam [GLF24]. occupancy [Den20, dTGC20, RPŠŠ23]. occurrence [CLZ21, MMPL20, WFL^{+}20]. occurring [ZZD22]. Ocean [ZZD22]. oceanic [ZZD22]. Oclouds [MTD^{+}24]. October [Ano20x, Ano21q, Ano22v, Ano23q]. off [ACA^{+}23, LZCGMVV20, Par22, RKG20, XLH^{+}24, ZDC22]. off-chain [RKG20, XLH^{+}24]. office [ZZP^{+}23]. Offloading [KGO^{+}20, LHH^{+}21, SP23, SXW^{+}22, AOKÖ24, AAP21, AKA20, CdRRdC^{+}24, CZZ^{+}23a, CZH^{+}24, GBH^{+}23, GZ22, HHH22, HXL^{+}23, HCG^{+}23, HX21, HGWC23, HB21, KJ24, LWNH22, LZL^{+}24a, LGL^{+}23, LZ22, LMZL24, LHY^{+}20b, LGL^{+}20b, MZA23, MMZI22, MISB22, MWL^{+}20, PNL^{+}21, QCW^{+}24, QZZ^{+}24, QCY^{+}21, RDR^{+}24, RSL24, RFP^{+}24, SHB22, TDM^{+}22, TWM^{+}23, XGS^{+}20, YCYO23, ZWZ^{+}24, uRLW^{+}21]. offs [AP20, XZK^{+}20]. offshore [SYXW21, YZL^{+}20]. oil [SYXW21]. OIPSO [SYXW21]. OLAP [TMT22]. oligopoly [TDLT20]. Omicron [MMC^{+}23]. Omnibus [GDGK20]. On-chain [KOM^{+}22, XQW^{+}24, XLH^{+}24]. on-chip [TBB^{+}23]. on-demand [ASAM20, SPL22, ZA22]. On-Device [NRBC23]. on-site [PKLC22]. on-the-fly [MSMJ22, WWY^{+}24]. on/off [ACA^{+}23]. once [LWZ^{+}23b]. One [Ben23b, KYY^{+}20, ALGMP^{+}21, FRGBHPPS23, KW20, MKC^{+}21, RCLEB20, RZIX20, WCD^{+}22]. one-class [ALGMP^{+}21, MKC^{+}21, RZIX20]. one-layer [FRGBHPPS23]. One-pixel [Ben23b]. one-sided [WCD^{+}22]. one-step [KW20]. onion [PGMP23]. Online [AYY^{+}20, AMR^{+}20, BEM^{+}20, BDF^{+}22, FPL24, GPR^{+}24, WCY^{+}21, WX23, WS23, AOSA20a, AMT^{+}21, CIJM20, HSR^{+}22, Kon21, KCP23, LKE22, LZHL23, LS23a, LYFZ20, LXH^{+}21, LY21, LZW21, MJB22, MM23, NKB^{+}20, OOZ^{+}23, PWV^{+}21, QPL22a, SSMdS21, SPL22, VPA20, WLC23, YYN^{+}20, YJB^{+}21, YLX^{+}23, ZYX^{+}20, ZZQ21, ZLZ^{+}23b]. only [LWZ^{+}23b, ŠTI24]. Onto [ACC20]. ontological [EEN^{+}24]. Ontologies [KPGD24]. Ontology [ACM^{+}21, GDCGVG20, SZVVB^{+}23, ACC20, BSH^{+}21, GNA^{+}21, MLP^{+}21, PSvL^{+}20, PS20, Tao23, TDMC23]. Ontology-based [SZVVB^{+}23, ACC20, GNA^{+}21, PS20]. Ontology-driven [GDCGVG20]. opcode [TLJ^{+}22, DAM^{+}21]. OpCode-level [DAM^{+}21]. opcodes [ZXX^{+}20]. Open [KHH21, BPC^{+}24, HHH22, HH22, LGCY22, MRS^{+}22, NLSY20, WMD^{+}20, FSBŞ^{+}20, MLX23]. OpenABL [CPJ^{+}21]. OpenCHK [MMK^{+}20]. OpenCL [JCP^{+}20, LCH^{+}22, PSH^{+}20]. OpenFlow [AYB^{+}22, YLSL22b]. OpenFlow-based [YLSL22b]. OpenFOAM [BBB^{+}24].

OpenMP [SPWL23, WLYL20]. OpenStreetMap [TD21]. operated [WLX^{+}24, ZLS22b]. Operating [RZA21, WCWC19, WCWC20, RAS^{+}22]. operation [CVdRA^{+}20, FZT^{+}23, Par20, TBB^{+}23, YWH^{+}21, ZL23]. operational [MBD^{+}20]. operationalization [dATBMA23]. operations [CPS^{+}23, MV21, RCR21, WZXX21, ZTQ^{+}20, SUKN22]. Operative [SN21]. operator [BEON24, HS21, SCR20, SZGB24]. Operators [NGC24, HBH21, SCGVP20]. opinion [LZL^{+}20, TTZ^{+}21, WM21, ZLS23]. opinions [VMCM^{+}20]. Opportunistic [uHA20, HYC^{+}23, GLM21, WCY^{+}20]. Opportunities [GZPZ20, WPPA22b, WPPA22a, CFK^{+}20, CDR24, DRC20, DPN^{+}22, LWHW22, MKK^{+}20, WMU^{+}23]. OPRA [BEM^{+}20]. opt [SCK^{+}22]. optical [HZL^{+}21, LYS^{+}20]. Optimal [AAP21, HCK20a, KMCJ20, LCL22, TSX^{+}24, UKY^{+}20, ASB^{+}23, BKG^{+}20, CDC^{+}24, DGL^{+}20, GCT^{+}20, GAA^{+}21, HZS^{+}23, ICBB20, KCJ23, SW22, SCX21, SACW23, TBH23, TA23, TIA21, VPA20, YZX^{+}23]. optimality [AMBGS21, BLMT20]. optimally [SPL22]. optimisation [RMC20]. optimise [DHD20]. Optimization [AWMM^{+}23, ASM^{+}22, AP20, Ben23a, KSDR21, LTX^{+}24, LLZ^{+}22, LGL^{+}20b, PSC^{+}21, PAC^{+}22, See20, SS22, SSB^{+}20, YHW^{+}20, YK20b, ABMESM18, ABMESM22, Ben23b, APC^{+}20, AJPM20, BÖE24, BÖ20a, BGNBH^{+}20, BM20, BRM^{+}20, CWB^{+}20, CHC^{+}20, CZZ^{+}23a, CLL^{+}23, CCL23, CS24a, CdO20, DT21, DGK20, DBSL23, EAA21, FPL24, GHEB^{+}18, GHEB^{+}23, GMFC23, GBH^{+}23, GLP^{+}24, GZXH24, GB20, GW20, HLP21, JZK^{+}21, JHK20, Kha24, KAK^{+}23, KSMT24, KV22, LZLY20, Li20, LCW^{+}20, LFM^{+}22, LZW^{+}22, LWL23b, LCC^{+}24a, LGZ^{+}24, LDDL21, LDW^{+}21, LJ24, LLLS24, LZHS24, LZCGMVV20, MECRFD20, MBM^{+}20, MSA^{+}24, NRBC23, NTA^{+}22, dSOFC^{+}23, PKR21, PWH^{+}22, PAM21, QPL22a, QKG20, RLZW21, SCX21, SLA^{+}23, VG21, VP20, WX23, WWW^{+}24, XCH^{+}20, XLLL20, XW23, XZTC22, YLSL22a, YHC^{+}22, YXL^{+}20, YCYO23, YPZ^{+}24, YLZ^{+}24, YXYH20, ZWZ^{+}24, ZHGX20, ZWL20, ZLZ^{+}20a, ZZ21a, ZNX23, ZHX^{+}20, ZWZ^{+}21, ZWL22, ZLL^{+}23, IMuI^{+}21]. optimization-based [LZHS24, QPL22a, WWW^{+}24, YLZ^{+}24, ZZ21a]. optimizations [LFYH22]. optimize [RK20]. Optimized [BAK22, CSD^{+}23, LY23, RRD21, TK24a, CCC^{+}23, CHJK22, Deh20, HTXW21, JHB22, LZS^{+}22, MK21, NPNC23, PVA^{+}20, WYZ^{+}24, WX24]. optimizer [AEZ22, FFAFD20]. Optimizing [ERL^{+}20, LCH^{+}21, LDD^{+}22, PGHS20, RKI^{+}23, SSV24, SCK^{+}22, WGW^{+}20, WF21, YZC^{+}20, QHE^{+}20, VMCM^{+}20, WWS^{+}23a, WGGB24, XWZM24, ZCLL22]. optimum [HY21]. option [CDP20b, CDP20c]. Oracle [PRD^{+}22]. orbit [CDX^{+}24]. orchestrate [TC23]. orchestration [ABMO24, AK20, ALS^{+}21b, FTS^{+}24, GMP^{+}20a, MABK24, SEKS^{+}20, STH^{+}20, VS20]. Order [LXL^{+}21, CKV22, MSR20, PFGDM22]. ordered [WLR21]. ordering [HTAY21, XLZ^{+}22]. ordering-based [XLZ^{+}22]. Ordinal [KWL^{+}23]. ore [LZH^{+}20]. Orfeon [dATBMA23]. organising [FGB21a, FGB21b, PCVN21]. organization [PPGS20]. organizational [AAGX^{+}22]. Organizations [RMA21, NFK^{+}20, SD24]. organizing [SAF23]. orientation [OMSL20]. orientation-aware [OMSL20]. oriented [AT20, DSW^{+}20, GZ22, HRM20, LGC^{+}21, Pan20, PRD^{+}22, VAKB23, WFLL22, YGD^{+}21, YJB^{+}21, YCS^{+}20, ZPK^{+}23]. orienteering [PPG^{+}20]. orienteering-based [PPG^{+}20]. orthogonal [LYYG20a, LYYG20b, LAT^{+}20]. OSM [VCK^{+}20]. Osmotic [SJQ20]. osteochondral [Bo19, Bo20b]. osteoporotic [WWP19, WWP20]. out-of-core [JPJO22]. out-of-the-box [GSI22]. Outage [PMT22]. outcome [MBC22]. outdoor [LQS^{+}20]. outlet [LLT22]. outlier [GDGK20, IPPK23, YPX^{+}20]. outliers [RZIX20]. output [ZLW^{+}22]. Outsourced [BKHD20, XCSF20, DKD22, FBL^{+}20, LYY^{+}22, LWY^{+}24, MTHA24a, MTHA24b]. outsourcing [TKS^{+}23, ZZZ^{+}21a]. outsourcing-supported [ZZZ^{+}21a]. Outstanding [Tau23, Tau24]. OVE [FSBŞ^{+}20]. Overcoming [ABC^{+}20]. Overflow [BPLFRL20]. overhead [TTTH20]. overlapping [DNNG21, HLT^{+}21]. overlay [Kad20]. OverlayFS [CSH^{+}23]. oversubscription [MC20]. overview [AT20, ABT20, SN23, ZXD^{+}20]. owner [NBJ21]. ownership [HMY^{+}23].

P [HAA^{+}20, HL24]. P-DACCA [HAA^{+}20]. P-PFedSGD [HL24]. P4 [PPSC23]. P4-assisted [PPSC23]. PaaS [SSM22]. PAC [ZHL24]. PAChain [Yue20]. Pacific [ZZD22]. package [MTM21, ZLL^{+}23]. package-aware [ZLL^{+}23]. Packet [DHA^{+}20, Dut22, GDP20, XLG^{+}23, ZHX^{+}20]. packets [ACG^{+}20a, QWR^{+}20]. PACO [PAC^{+}22]. padding [QHE^{+}20]. PAGCL [LWL23a]. Page [GPGG23]. Pages [Ano20m, Ano20w, Ano20u, Ano20r, Ano20n, Ano20o, Ano20s, Ano20t, Ano20v, Ano20q, Ano20x, Ano20p, Ano21u, Ano21m, Ano21x, Ano21w, Ano21v, Ano21o, Ano21n, Ano21r, Ano21s, Ano21t, Ano21q, Ano21p, Ano22u, Ano22s, Ano22r, Ano22w, Ano22p, Ano22m, Ano22n, Ano22x, Ano22o, Ano22q, Ano22v, Ano22t, Ano23x, Ano23u, Ano23w, Ano23n, Ano23o, Ano23s, Ano23p, Ano23r, Ano23t, Ano23v, Ano23q, Ano23m, Ano24q, Ano24r, Ano24j, Ano24m, Ano24n, Ano24k, Ano24o, Ano24l, Ano24p]. painting [Wei21]. panchromatic [GHEB^{+}18, GHEB^{+}23]. pancreatic [XY20]. PANDA [WHC^{+}22]. pandemic [KSS^{+}21, VCM^{+}21]. Panoramic [ZZ21b]. Pap [ZWH^{+}21b]. paper [Tau24]. paradigm [AM21, DML20, GMP20b, KMK^{+}23, KSMT24, ZLG^{+}24]. paradigms [CLLCK20]. paradox [ERG^{+}22]. Parallel [BP20, CK20, DZB23, LDDL21, LGM^{+}21, LV24, PAC^{+}22, ABGDT23, BRM^{+}20, CFL^{+}20, GOA23, HTAY21, HYC^{+}21, HHLZ20, LLP^{+}20b, LPS^{+}24, LCY^{+}23b, LJ24, LL20, LEXH20, dAPHOMPJ20, MGS21, OPOG23, PSS^{+}23, RLML20, SJVRS22, SHW24, SNS^{+}20, Ski20, SVN20b, TDS^{+}22b, XWM20, XLX^{+}21]. Parallel-inducing [DZB23]. paralleled [ZGW^{+}23a]. parallelism [BPGL21, JZL^{+}20, QNRA23, SPWW21, SLA^{+}23, XHW20, ZLQ23]. parallelity [TDS^{+}22a]. parallelized [TSR^{+}20]. Parallelizing [LLT20, QHE^{+}20]. Parameter [ZLZ21, BK20, LJW^{+}20, ZST^{+}20]. parameterizable [JCP^{+}20]. parameterized [KHL20]. Parametric [WTL^{+}20]. Pareto [YZX^{+}23]. Pareto-optimal [YZX^{+}23]. parity [PK22]. parking [RPŠŠ23, SKB20]. Parkinson [AMZZ23, KNRI21]. Part [WPPA22b, WPPA22a]. partially [WLR21, ZZQ21]. partially-ordered [WLR21]. Participatory [YCS^{+}20, KSS^{+}20]. Particle [KSDR21, ZWL22, AJPM20, CZZ^{+}23a, GZXH24, KMS20, LZHS24, ZLZ^{+}20a, ZZ21a, ZLZ21]. partition [LDCZ20, NSR^{+}23]. partition-based [LDCZ20]. Partitioned [PCVN21, HMA^{+}21, MLZ^{+}23a]. partitioning [AHL^{+}23, AB20, GGCIV20, KMCJ20, KCKK24, RCP24, TJG^{+}20, WLS^{+}24]. partner [YPEK23]. party [CMM^{+}23, LWZ^{+}23a, MMM^{+}20, XLH^{+}24, XCZ^{+}22]. Parzen [SVFdA20]. Pascal [GDS^{+}20]. passenger [BYW^{+}21, LLZ^{+}24a]. passive [dTGC20]. passively [YRV^{+}22]. Password [ZPS^{+}24, FYHZ24, HZ20]. password-based [HZ20]. Patch [MSMJ22, ZXX^{+}20]. patch-based [ZXX^{+}20]. Path [AAG23, CZGS20, LQML22, PCG^{+}20, AB20, KHB20, LGW^{+}21, LYW23, LCCP21, MDL^{+}23, QGH^{+}22, SP21, SuRMA^{+}23, TLKX21, XWZM24]. pathological [LCH^{+}22]. Pathologies [WMU^{+}24]. pathology [ERG^{+}22, JLS^{+}23]. paths [GZF^{+}20b, PMT22, RPdVR20]. patient [AdSM^{+}22, AYHA20, MSKG21, TDLT20, WCWC20]. patient-centered [AdSM^{+}22]. patients [ABM19, ABM21, CXHS20, HZX^{+}19, HZX^{+}20, KIJ^{+}24, PRPPFRL20, WCWC19]. Pattern [MABK24, PCACTMÁ24, YK20a, ALGMP^{+}21, ATK^{+}22, uHA20, dRFRB24, KYY^{+}20, KHES21, LWCC23, LYG^{+}24, PCVN21, SGDK^{+}21, WGW^{+}20, YYKK20, YNK^{+}20, ZL21]. Pattern-based [MABK24, ALGMP^{+}21, SGDK^{+}21, YYKK20]. patterns [CABB20, DDMP^{+}23, GPWL20, HNV^{+}20, LWH^{+}22, MTCS22, PSS^{+}23, PLMZ23, QRS^{+}21, SGSGGC^{+}23, SCL20, SOKW^{+}20, TSKK23, VFOV20, WFL^{+}20]. Paving [CDBD24]. Payload [VMM20b, ISD22]. payment [XLH^{+}24]. PBC4occ [ALGMP^{+}21]. PBFS [MAM^{+}24]. PBRL [ZLL^{+}24b]. PBRL-TChain [ZLL^{+}24b]. PCCP [BEB^{+}20]. PDAS [JCW^{+}23]. PDF [FPH^{+}21, JLW^{+}23]. PDR [LWS^{+}23b, MR23a]. peak [JR22]. Peaks [LGS^{+}23, YLZL21]. pedagogy [SP22]. Pedestrian [HH22, BAMR20, LHY^{+}20a, RSBM20, ZY21]. pediatric [LSMT^{+}21]. peer [Liu24, LGKA21, PBL^{+}23, ŠTI24, ZC22]. peer-node [Liu24]. peer-to-peer [LGKA21, PBL^{+}23, ŠTI24, ZC22]. pegylated [WCWC19, WCWC20]. pegylated-interferon [WCWC19, WCWC20]. Pelican [AWMM^{+}23]. pellet [LZH^{+}20, LYC^{+}22]. pending [ABA24]. Penetralium [YYW^{+}24]. people [HZPS21, FHGF20]. Perception [LSH^{+}20, CZZ^{+}23a, DGK20, LDW^{+}21, XLCB20, ZZ21b]. Perception-enhancement [LSH^{+}20]. perceptual [CSC23]. percutaneous [WWP19, WWP20]. Performability [MZA23]. Performance [BJ22, BJP^{+}20, BLMT20, dMBPdSC20, CDC^{+}24, DGL^{+}20, LPS^{+}24, LLZ^{+}22, LGKA21, MGGG^{+}20, MNSL22, NBB20, RNV^{+}21, SHH23, TDC^{+}20, UCO20, VMV20, WFLL22, YCYO23, AM22, ACG^{+}20b, ATC^{+}24, BQC23, BK20, CF24, CCP^{+}20, CCW^{+}20c, CFÁA^{+}20, CKW21, CCC^{+}24, DLH^{+}20, DLC^{+}22, GSARS20, GA22, GZL^{+}22, GHW^{+}20, bHFF^{+}21, KAF^{+}23, KAK20, KHES21, KBTM21, KSE24, KCP23, LQW^{+}20, LZW^{+}22, LBDP23, MEL^{+}23, MMFAB23, MDP24, MMPV22, MTA^{+}22, MEC^{+}20, OGO^{+}20, PWV^{+}21, Par22, PP24, PSS^{+}23, PK22, dRRCGdC20, RLML20, SAD24, See20, SDZ^{+}20, SK20b, SPRA21, SXC^{+}23, TCBF24, UADD21, WCD^{+}22, XCGZ24, XZK^{+}20, YLTH22, YK20a, YYKK20, ZLQ23, ZLL^{+}24b, ZHX^{+}20, ZHL24]. performance-driven [ATC^{+}24]. performance-enhanced [ZLL^{+}24b]. performance-focused [MEL^{+}23]. Performance-oriented [WFLL22]. performance-security [AM22]. performance-to-power [DLH^{+}20]. perfusion [WZC^{+}22]. period [JR22]. periodicity [SOKW^{+}20]. periods [CL20a]. permission [AAS^{+}20, CCHD21, KZG^{+}22]. permission-based [CCHD21]. permissioned [HSGY20, RRDSAML23, ZLL^{+}24b]. permutation [ABMESM18, ABMESM22, HLL^{+}20, MDG^{+}22, XYL^{+}20, XLZ^{+}22]. Persistent [TQC20, ZCWC20, BCB^{+}20, WX24, XHW20, YZC^{+}20, YYZ^{+}24]. person [YLD^{+}23]. personal [BMBE20, SP24, VPSC^{+}23]. Personality [GVŠ22, GSMF20, MSM^{+}22, YRV^{+}22]. Personalized [TSM24, BAGRB^{+}20, CHJ^{+}20, EKK23, FXZZ24, HJW^{+}20, KKL^{+}24, MKK^{+}20, OOB^{+}21, PP20, YZL^{+}23]. personnel [SZW^{+}23]. perspective [GSKS20, GTG^{+}21, LPL^{+}20, RMA^{+}20, Sun20, VPSC^{+}23]. perspectives [BAM^{+}24, RHK^{+}23, ZWW^{+}23]. perturbation [CYWS24, HXWX23]. perturbation-based [CYWS24]. Pervasive [DAAW20, YMS20a, Kol22, KA24, LGKA21, SCP^{+}21, YMS20b]. Peter [Fae21]. Petri [YFL^{+}24, YJF^{+}20, YDL^{+}20]. petroleum [LCZB21]. PEWOBS [XLZ^{+}22]. PFedSGD [HL24]. PFPMine [HHLZ20]. Pharmacovigilance [DFG^{+}21]. phase [AOF21, FAA^{+}23, JPW20, LZH^{+}20, NLS23, PAM21, ST20b]. phenomena [ZZD22]. phenomenon [TDLT20]. phenotyping [MTCS22]. PHEVs [JKS20b]. phishing [LWS^{+}23a]. phonics [ZLST23]. phonics-based [ZLST23]. photodynamic [ZMZ^{+}19, ZMZ^{+}20]. photographs [ZLS22b]. photography [YLY^{+}23]. photon [PCI^{+}24]. photovoltaic [PKLC22]. phylogenetics [SJVRS22]. Physical [AABKB22, DAAW20, DLHD22, IDM^{+}20, Kho21a, KVCY20, LGKA21, MR23b, RCJZ20, SWW^{+}20, SRM^{+}23, VKP22, YLS21, YL20b, ZGY20, ASASA^{+}20, CDG^{+}20, CLQS20, EKK23, HLW^{+}23a, HMY^{+}23, HRM20, HMLS20, JSV21, KYPJ20, LLG^{+}20, LCLA21, MLWA20, RPP^{+}20, SVN^{+}20a, TCMV20, TDL^{+}21, VZDS24, WGLH20, XZJ^{+}20, XZK^{+}20, XWW^{+}20, YXYH20, ZYL^{+}20, ZLW^{+}22, ZXL^{+}20]. Physiological [WHF^{+}23, dRFRB24, KHB23]. physiotherapy [ÇYZZ21]. Pi [SNM^{+}20]. PICEA [PKR21]. PICEA-g-based [PKR21]. picture [ZCF21]. PIHA [CSC23].

Pilot [BHSH22]. Pilot-Job [BHSH22]. pipeline [CCC^{+}24, PDJS22, SDGCB^{+}20]. pipelines [dATBMA23, MABK24, MWS24, NVS^{+}22]. pipelining [TWI20]. pixel [Ben23b]. Pkg2Vec [MTM21]. PKI [QHW^{+}20]. Placement [PAC^{+}22, ATZP21, CCL^{+}20, HBH21, HRY^{+}21, HKB^{+}24, LZLY20, LCL22, LDD^{+}22, MPC^{+}24, Man20, MO24b, MAA22, PKB22, PGHS20, PSC^{+}21, SEL^{+}22, SMS22, THB23, TBH23, WYZ^{+}24, WHW20, WLY23, XWD20, XZC^{+}24, ZDLD24, TPD^{+}24]. Plane [MECRFD20, HYL^{+}20, MSLP24, ZLP^{+}22]. planes [NPNC23]. planning [HYRZ20, LS23b, LYW23, LCCP21, MDL^{+}23, MLN24, PPG^{+}20, TLKX21]. plant [ABAJ20, RGPGP24]. Platform [CHS^{+}24, OCMJFB^{+}23, BLGCLA^{+}23, DJP^{+}24, BCSS20, CGFC20, CREE^{+}24, CDF^{+}22, EET20, GZF^{+}23, GRG20, HCB^{+}20, JLS^{+}23, LLFQ21, MAC^{+}21, MLZ^{+}23b, ONK^{+}20, PBC^{+}22, RBLD21, RAA^{+}20, RKI^{+}23, SVFdA20, SYXW21, WXD^{+}23, XY20, YLKK20, ZKD21, Zha20, dlVGSB^{+}20]. platform-independent [dlVGSB^{+}20]. platforms [BPGL21, DSC20, DPPGCCA23, FGG^{+}21, GSI22, HBH21, HPP20, NLSY20, PBM^{+}22, PSMF21, PDT21, TLC^{+}20, VAKB23, YAZ^{+}20, ZXD^{+}20]. players [FS21]. Plethora [ALR^{+}20]. Plummeting [SMS^{+}24]. PM [ZG23]. PMKT [MMM^{+}20]. PMMSA [KZG^{+}22]. pneumatic [LZZX20]. pneumonia [LSMT^{+}21]. Pod [PK24]. POI [LLZ^{+}24a, LH20]. POI-based [LLZ^{+}24a]. point [BBB^{+}24, HMLS20, NRBC23, WGW^{+}20]. point-of-interest [HMLS20]. points [CCL23]. poisoned [LWL23a]. Poisoning [TSM24, EUEU24, RBMCLH22, ZZG^{+}22]. polar [SLH^{+}20]. polarized [AR20]. policies [ACD^{+}20, CTFW22, CF21, PLS^{+}23, RFP^{+}24, SFC23, WDS^{+}23].

Policy [ASM^{+}22, CGFC20, SME^{+}21, SZM22, BKHD20, DG21, MM23, MSBAU24, PPA^{+}24, RSQS21, SME^{+}19, WYZ^{+}24, WXZ23, ZZQ21]. Policy-based [CGFC20]. political [FAŞ^{+}20, GDCPVG22]. politician [GDCPVG22]. pollination [GHEB^{+}18, GHEB^{+}23]. polling [UPD^{+}20]. pollution [EJP22, LXY21, ZZ24]. POMF [XQW^{+}24]. pommDNN [CDC^{+}24]. Ponzi [BCCS20]. pool [GFM^{+}20]. Pools [PRF20]. Poor [LTX^{+}24]. Population [WZJ^{+}22, CHC^{+}20, KA22, VG21]. pornography [MSV^{+}20]. portability [CCC^{+}24]. portable [ASB^{+}23, CFL^{+}20, NADY20]. portal [MBD^{+}20]. PoRX [WLC^{+}20a]. pose [LRCL24, RAS^{+}22]. positional [JVH^{+}20]. Positioning [NHTH20, CZGS20, MR23a, SYXW21]. possible [MAB^{+}20]. Post [LYKK22, LWZ^{+}23a, OHÁV20, SK20a]. post-processing [OHÁV20]. post-quantum [SK20a]. Post-training [LYKK22]. posted [GDCPVG22]. postoperative [ZYX^{+}23]. posts [AYHA20]. posture [HHW^{+}22, ZWL21]. postures [SG20]. potency [RZH21]. Potential [BAİP24, LXL^{+}21, DLR23, JZL^{+}20, YXL^{+}21]. Potential-based [BAİP24]. potentially [YGS^{+}22]. Poverty [CL21, WZT^{+}20]. poverty-stricken [WZT^{+}20]. Poverty/investment [CL21]. Power [CLL^{+}23, ZLXH20, ZWZ^{+}21, ABAJ20, BMD^{+}21, DLH^{+}20, DSW^{+}20, EKJ^{+}20, GAT^{+}20, Kad20, KCP23, LY23, LPT22, MC20, OPLB24, Par20, PPGS20, SK20b, SMC^{+}20, WYJ^{+}24, XZTC22, YZL^{+}20, ZWC^{+}22, ZHGX20]. power-domain [Kad20]. power-efficiency [SK20b]. power-law [LPT22]. POWER9 [BDGG^{+}20]. Powered [WPPA22b, LP21a, PJJ^{+}22, SZZY22, YZL^{+}23, ZGY20, ZLXH20, ZL23, ZZD22]. powerful [ABM19, ABM21]. Powertrain [SRM^{+}23]. PP [LOH^{+}23]. PPCensor [MSV^{+}20]. PPG [HWH^{+}23a]. PPSFL [ZCL24]. PPUP [YLY^{+}23]. PPUP-GAN [YLY^{+}23]. PR [CFD^{+}20]. PR-KELM [CFD^{+}20]. Practical [CPYY23, dTGC20, CP22, LFZJ21, HLZ^{+}22]. practice [BFG^{+}22]. practices [GZF^{+}23]. pre [GlRpG20, MTHA24a, MTHA24b, WXZX23, WZS^{+}23, YNK^{+}20, ZZZ^{+}22, ZGZX21]. pre-cache [WZS^{+}23]. pre-caching [GlRpG20]. pre-computed [MTHA24a, MTHA24b]. pre-large [YNK^{+}20]. pre-processing [ZGZX21]. pre-training [WXZX23, ZZZ^{+}22]. Prebaking [FSP^{+}24]. precedence [JHB22]. precedence-constrained [JHB22]. precise [CCW^{+}20c, LRCL24]. precision [BBB^{+}24, GAP24, LDLS20, NRBC23, TBB^{+}23, TBA23, YYL22]. predator [Dho20, QZZ^{+}24]. predators [HAqDE23]. predict [NLO^{+}20, XLG^{+}23, Yan21]. Predictable [BVFGSF20, SWW^{+}20]. Predicting [FAAS20, GSMF20, KK22, LLZ^{+}24a, SSMdS21, BLT^{+}24, GZL^{+}22, LQW^{+}20, YLM23, YGS^{+}22]. Prediction [Pła24, VSV^{+}23, ABOS22, ABGDT23, ABL22, AABB24, BYW^{+}21, BAR21, BMBC20, CL20a, CKL20, CSY^{+}20, CFD^{+}20, CWM^{+}20, CLQS20, CPH^{+}22, CYZ^{+}22, CCW^{+}20c, CYH20, DBC24, DWZ20, GW22, GDP20, GLF^{+}22, HZX^{+}24, KOM^{+}22, KCY^{+}21, KZF21, LGJ^{+}23, MBC22, MLP^{+}21, MASRAM^{+}22, MSM^{+}22, MWL^{+}20, MOW^{+}20, MBD^{+}20, POMK20, PKLC22, PCG^{+}20, PRD^{+}22, PDFV21, PTZ^{+}20, QHNL21, dRRCGdC20, RRHA21, SYG^{+}20, SPRA21, TYR22, THVL24, WCY^{+}21, WLL22, WCL^{+}24, WWS20, WZH^{+}22, YFQ^{+}22, YJB^{+}21, YZL^{+}20, ZWL20, ZL21, ZPQH21, ZLW^{+}22, ZG23, ZKGB20]. Prediction-based [Pła24]. predictions [AMT^{+}21]. Predictive [HBSG21, TAM21, WDL^{+}21, HAVK22, KCJ23, MHA^{+}24, dSOFC^{+}23, RSL24, WCS24]. predictor [LQW^{+}20]. predictors [dMBPdSC20, LFM^{+}22]. predicts [ZYX^{+}23]. preemptive [BJ22, YXL^{+}20]. Preface [Ano24s, LLWJ24, LWHW22, PPM24, VKP22, WD24]. preference [KK22]. preferences [MECRFD20]. Prefetching [TQC20, ZSZ^{+}24]. Pregel [LDCZ20]. preliminary [MGC23]. preoperative [MXL^{+}20]. preprocessing [DSC24]. prerequisite [Gas22]. presence [PMT22]. presentation [CZZ^{+}23b]. preservation [MTT^{+}23, QJS^{+}21, SCZ^{+}20, SRA^{+}22]. preserve [ADAHA^{+}21]. preserved [XQW^{+}24]. Preserving [GXS22, TPF^{+}20, WQHX20, ACBT23, AB24, AYY^{+}20, BYR^{+}20, BKHD20, BDM^{+}20, CXWY21, CYG22, CKV22, DFF^{+}23, DDT^{+}23, FLH^{+}24, GNA^{+}21, GJC^{+}20, HLT^{+}21, JT22, KJYC23, KK20, LOH^{+}23, LZZ^{+}23, LLW^{+}23a, LLY^{+}20, LYY^{+}22, LCH^{+}24, MBGC20, NNUV20, PUMN^{+}24, QMCX20, QHNL21, SYYuR21, SYYuR22, SLX^{+}24, TOM^{+}20, WDHY20, WHC^{+}22, WYDB24, WZS^{+}22, XZYH22, XRZ^{+}24, XZZ^{+}20b, XWW^{+}20, XPT^{+}22, XCZ^{+}22, YVSG22, YYW^{+}24, YYB^{+}21, ZWW^{+}20a, ZPQH21, ZLF^{+}23a, ZZP^{+}23, ZSL^{+}23b, ZWCS23, ZLWH23, ZCL24, MMM^{+}20]. pressure [AMT^{+}21, LLZ^{+}21]. prevent [NNUV20, QL22]. preventing [FLG^{+}20]. Prevention [LWHW24, Sha20, MK21, UADD21]. prey [Dho20]. pRIblast [ABGDT23]. price [CDP20b, PRD^{+}22, ZL21]. prices [LZL^{+}20]. Pricing [CDP20c, CLY^{+}20, LQYL21, LXZ^{+}20, SKB20, WZW^{+}20]. Primary [Elg20]. principal [SuRMA^{+}23]. principle [LLT22]. Principled [EEN^{+}24]. prior [GYAW22, YZW22]. prior-dependent [GYAW22]. priorities [OPOG23]. prioritization [CDP20a, CdO20]. prioritized [GMP20b, KAA^{+}21].

Prioritizing [BSH^{+}24]. Priority

[MAM^{+}24, BJ22, CdRRdC^{+}24, HXL^{+}23, LRCL24, RWJ^{+}20]. priority-based [CdRRdC^{+}24, RWJ^{+}20]. Privacy [AHN21, AYY^{+}20, BKHD20, BKM^{+}22, CSD^{+}23, DFF^{+}23, GXS22, KK20, LOH^{+}23, LWX22, LLY^{+}20, LZCH22, LCH^{+}24, NHY20, PGCB23, QHNL21, RPŠŠ23, XZYH22, XZZ^{+}20b, XWW^{+}20, YYW^{+}24, YDK20, YYB^{+}21, ZPQH21, ZZP^{+}23, ACBT23, AB24, ADRP23, ACD^{+}20, AP20, ADAR22, BYR^{+}20, BDM^{+}20, CXWY21, CDV^{+}24, CCH21, CYG22, DDT^{+}23, EL21, EL23, FLH^{+}24, GNA^{+}21, GJC^{+}20, GLWP20, JT22, KJYC23, LZZ^{+}23, LLW^{+}23a, LWHW22, LYY^{+}22, MTT^{+}23, MRR^{+}20, MPP^{+}21, NBJ21, NNUV20, PUMN^{+}24, QNM24, QMCX20, QLJ21, RHK^{+}23, RLL^{+}22, SMKC20, STK20, SYYuR21, SYYuR22, SCZ^{+}20, SRA^{+}22, SXHD24, STK23, SLX^{+}24, TOM^{+}20, TG20, UPK^{+}23, WMD^{+}20, WQHX20, WFL^{+}20, WHZ^{+}20, WHC^{+}22, WWZW23, WYDB24, WLD^{+}20b, WZS^{+}22, XRZ^{+}24, XCSF20, XPT^{+}22, XCZ^{+}22, YL20a, YLZL21, YLS21, YVSG22, YLY^{+}23, ZWW^{+}20a, ZZQ21, ZGL^{+}23, ZHH^{+}23, ZLF^{+}23a, ZrHhH^{+}23, ZNX23, ZSL^{+}23b, ZWCS23, ZLWH23, MMM^{+}20, XQW^{+}24, ZCL24]. privacy- [YLZL21]. privacy-aware [CDV^{+}24, XCSF20, ZHH^{+}23]. privacy-enhanced [ADAR22, ZrHhH^{+}23]. privacy-preservation [SRA^{+}22]. Privacy-preserved [XQW^{+}24]. Privacy-Preserving [GXS22, AYY^{+}20, BKHD20, DFF^{+}23, LOH^{+}23, LLY^{+}20, LCH^{+}24, QHNL21, XZYH22, XZZ^{+}20b, XWW^{+}20, YYW^{+}24, YYB^{+}21, ZPQH21, ZZP^{+}23, ACBT23, AB24, BYR^{+}20, BDM^{+}20, CYG22, DDT^{+}23, FLH^{+}24, GNA^{+}21, JT22, KJYC23, LZZ^{+}23, LLW^{+}23a, NNUV20, PUMN^{+}24, QMCX20, SYYuR21, SYYuR22, SLX^{+}24, WHC^{+}22, WYDB24, XRZ^{+}24, XPT^{+}22, XCZ^{+}22, YVSG22, ZLF^{+}23a, ZSL^{+}23b, ZLWH23, ZCL24, MMM^{+}20]. privacy-protected [RLL^{+}22]. privacy-protecting [YLY^{+}23]. privacy-protective [ZZQ21]. privacy-utility [AP20]. Private [XZZ^{+}20a, Yue20, CHS^{+}23, CSS22, FGG^{+}23, JGL^{+}20, LGT^{+}20, LYY^{+}20b, TGAP20, ZXX23, ZLF^{+}23b, ZXL^{+}20, FLH^{+}24]. Pro [MPC^{+}24]. Pro-active [MPC^{+}24]. Proactive [BEB^{+}20, BEM^{+}20, KA21, MSS^{+}24, ZAH^{+}20, Kol22, KA24, MMZI22, RSLL21]. Probabilistic [HAA^{+}20, LZ20a, WLZ^{+}20]. probability [MDL^{+}23]. problem [ABMESM18, ABMMC18, ABMESM22, ABMMC22, BPUW24, DZXS21, LZCGMVV20, LEXH20, LCC^{+}24b, yQhJL20, SCK^{+}22, WXX^{+}24]. problems [BM20, FW22, HH22, MECRFD20, SCR20, YWDC23]. Process [CMM^{+}23, DMPS23, AAG22, AAB23, CCML20, CPM^{+}23, ODET21, WLLC20, ZTC20]. Process-aware [DMPS23]. processes [AHMW23, ACT24, BKG^{+}20, KMS23, LDGS20, LDDL21, PSHW20, SSWW23, YDL^{+}20, YFL^{+}24]. Processing [BEB^{+}20, DOR^{+}21, LL20, XWM20, AdAHK20, ACC20, BVFGSF20, BHL^{+}21, BGMK22, BMBC20, CHG^{+}20, CWB^{+}20, CKV22, DKG^{+}22, EMHE21, GGCIV20, JHK20, KIM^{+}24, KYPJ20, LLZL21, LZ21a, MABK24, OHÁV20, OWK^{+}23, PSS^{+}23, QLHLB23, RBA^{+}22, SDGCB^{+}20, SHW24, SQ22, THA^{+}24, VMV20, WT24, XKK20, XW23, XLL^{+}20b, YLSL22a, YYKK20, ZBTV^{+}20, ZGZX21, ZHLM20]. processors [LVLBB^{+}24, PPGS20, WFLL22, XLX^{+}21]. produce [BK20]. product [BMS20, LLY^{+}20, VPA20]. product-form [BMS20]. production [CdST^{+}20]. productivity [CGMT20]. products [BSH^{+}24, GBC^{+}24, Zha21, ZP22, ZGZX21]. Prof. [Fae21]. profile [MJB22, RNV^{+}21].

Profiling [SLA^{+}23, Kho21b, KLW^{+}21, MLWA20, RAS^{+}20, SNS^{+}20]. profit [NT22]. programmable [LWHW24, MSLP24, QSZ^{+}24]. programmers [ORPPG20]. Programming [HTXW21, DP20a, DDMP^{+}23, EELB21, GRG20, GBdRACG20, JPW20, JMHB24, LLP^{+}20b, LGM^{+}21, MBC22, MGS21, PP22, RCLEB20, RMA^{+}20, XZTC22]. programming-based [DDMP^{+}23]. programs [SPWW21]. progression [ESSS^{+}21]. Progressive [YLD^{+}23, CXS^{+}22, JCW^{+}23]. Project [SCR20, DSRG22, LH21, YPEK23, DFF21]. PROMENADE [CDF^{+}22]. promising [WFA20]. promote [GZF^{+}23]. promotion [KHL20]. promotion/demotion [KHL20]. PROMPT [PLS^{+}23]. Proof [FZC^{+}20, BBB22, HMY^{+}23, LGW22, MMR23a, WHJ20]. proof-of-concept [BBB22]. Proof-of-negotiation [FZC^{+}20]. proof-of-work [MMR23a]. proofing [CIB^{+}20]. proofs [XCZ^{+}22]. propagation [YGS^{+}22]. properties [DDM21, GL20, SHT^{+}21]. property [XXL^{+}24]. proportionality [MC20]. proposal [CPPY24, dSOFC^{+}23]. prospect [TLX^{+}23]. Prostate [HIU^{+}22, LYH^{+}21]. protect [SCE23]. protected [RLL^{+}22]. Protecting [SK21a, YZJ^{+}20, CG21, ICW21, JLW^{+}23, SP24, YLZL21, YLY^{+}23]. Protection [LRML21, AHN21, LQML22, SXHD24, WHZ^{+}20, WWZ^{+}20, WLD^{+}20b, XXL^{+}24]. protective [ZZQ21]. PROTECTOR [MK21]. protein [GLF^{+}22]. Protocol [LLZ^{+}22, AOKÖ24, ABAD22, BMZdP21, CECS20, CXHC23, DSPSNAHJ20, DLL20, GBK20, GCH^{+}22, LGYC20, Liu23, LYY^{+}20b, MNA^{+}23, NNH^{+}20, NIB^{+}21, RRAB24, RWJ^{+}20, SWW^{+}20, SKX^{+}20, YYN^{+}20, ZC22, ZWW^{+}20a, ZKL^{+}23, dAdSM^{+}22]. protocols [GdOAO20, KBTM21, LCFM20]. proton [SWL^{+}20]. Prototype [PLL^{+}24]. Prototypical [MSC^{+}23]. prototyping [MEL^{+}23]. PROUD [BKHD20]. PROV [MBGC20]. proven [TSKK23]. Provenance [GMF^{+}20, BCB^{+}20, MBGC20, SSC^{+}20]. provide [BMZdP21, CMF^{+}21, GSI22]. Provider [PVA^{+}20, GSKS20, GZB^{+}22]. providers [MRM^{+}24, QMCX20, ZAH^{+}20]. Providing [PPA^{+}24, CdRRdC^{+}24, NSR^{+}23]. Proving [FTM20]. provision [LZCH22, OPLB24, SKH20]. provisioning [BHSH22, BGMK22, CWYG23, CS23, HWQ^{+}20, KHH21, KKL^{+}24, LBGL20, NGCB20, QPL22a, SPL24, TSB20]. Proximal [ASM^{+}22, MCT^{+}22]. proxy [LYFZ20, ZWWC21]. pruning [BEM^{+}24, CMGS22, ERK^{+}24, JCW^{+}23]. PS [BPSP23, LWHW24]. PS-CRDTs [BPSP23]. PS-IPS [LWHW24]. PSDCE [WHF^{+}23]. pseudo [HNV^{+}20]. pseudo-IDLists [HNV^{+}20]. PSNet [LJW^{+}20]. PSO [MYM^{+}21, TLKX21, ZP22, ZT22b]. PSO-based [MYM^{+}21]. PSO-weighted [ZT22b]. Psychographic [GDCPVG22]. psychological [VCM^{+}21]. PTCP [RWJ^{+}20]. pub [HBH21, ACBT23]. Pub-SubMCS [ACBT23]. pub/sub [HBH21]. Public [PFS^{+}23, YPEK23, ARB20, BBM^{+}20, GW22, IB20, LZL^{+}20, LYP^{+}24, SVN^{+}20a, WCXW22, YLH^{+}23, ZLS^{+}22a, ZLS23]. publication [ZXX23]. Publicly [YYN^{+}20]. publish [ACBT23, LPSV22]. publish/subscribe [LPSV22]. Publisher [Fae21]. publishing [TRB^{+}23b, ZXL^{+}20]. PubMed [DFG^{+}21]. PUC [DHA^{+}20]. PUDT [SMS^{+}24]. PUFs [RAB23]. pulmonary [ZHP^{+}21]. pure [XHW20]. pvp [WWP20, WWP19]. Pyramid [GSG^{+}23, HQLH20, LWH^{+}22]. Python [AFL23, CFÁA^{+}20, SFC23, SLA^{+}23]. Python-based [SFC23, SLA^{+}23].

Q [CWL20, DFZ^{+}20, GMM22, GPR^{+}24, Liu23, SP23, YPL24]. Q-learning [CWL20, DFZ^{+}20, GPR^{+}24, GMM22, YPL24]. Q-Learning-based [Liu23, SP23]. QA [HPP20]. QAOC [PS20]. QASEC [UJHN20]. QCWOA [LLZ^{+}22]. QF [ZJL^{+}22]. QFaaS [NUB24]. QFD [ABMMC22, ABMMC18]. QHSE [BYR^{+}20]. QoE [ARbL^{+}20, LHC^{+}20, ZWZ^{+}24]. QoE-aware [LHC^{+}20]. QoS [BRK24, BMZdP21, CSY^{+}20, CWM^{+}20, CPH^{+}22, CYZ^{+}22, DCGM20, HWQ^{+}20, HKB^{+}24, KHH21, LHC^{+}20, LDDL21, PKB22, PP24, PFS^{+}23, RFP^{+}24, SSM22, YSL^{+}22, ZLG^{+}24]. QoS-aware [DCGM20, HKB^{+}24, LDDL21, PKB22, PP24, RFP^{+}24, ZLG^{+}24]. Quality [BOD^{+}24, AAP21, BH21, CPT^{+}20, CPH^{+}22, CDV^{+}24, GSDGP21, GPGG23, HQLH20, JZZD21, JL21, LLW^{+}20, LYY^{+}20a, LZK21, LWZ^{+}20, LZC^{+}23b, LC20, MBD^{+}20, NKG23, QC21, SMC^{+}20, SZL^{+}21, WDG20, WLLY20, WWS20, XWLC20, XWK21, YJH^{+}20, YWS21, ZZ21a, ZPQH21, Zhu20, ZH20, ZLWH23, MOU^{+}21]. quality-aware [AAP21, HQLH20, LYY^{+}20a, LZK21, SZL^{+}21, XWK21, ZH20]. quality-guided [BH21]. Quantifying [BYH^{+}20, AB19, ABGMC19, AB21, ABGMC21, THB23]. Quantised [MÖ24a]. quantization [LYKK22]. Quantum [BPUW24, GZXH24, LLZ^{+}22, NUB24, OPOG23, CZZ^{+}23a, CDBD24, DA22, GCM21, KIM^{+}24, KA22, LWZ^{+}23a, PCI^{+}24, SK20a]. Quantune [LYKK22]. quaternion [SLH^{+}20]. queries [CQA^{+}24, KA21, LZZ^{+}23, MNFQ24, YZW22, YZX^{+}23, ZLF^{+}23b]. Query [GGCIV20, LHXL22, ZJL^{+}22, CSC23, CYH20, CYG22, CKV22, HDD20, MBM^{+}20, PS20, SOT24, TSR^{+}20, ZHL^{+}23]. Query-based [LHXL22, CSC23]. querying [MBZ^{+}21]. question [XZ20]. question-answering [XZ20]. questions [WDG20]. queue [BJ22, MLX23]. queueing [MGW23]. Queuing [VSV^{+}23]. quick [LGT^{+}20].

R [LYH^{+}21]. R-CNN [LYH^{+}21]. Radial [BKM^{+}22]. Radio [Elg20, HN23, SKA^{+}20, ASA^{+}20, MMZI22, WYX^{+}23b, YK20b, Zhu21]. radiological [LFM^{+}22]. radius [FBTJ23]. rainbow [LZC^{+}23a]. rainy [ZPLQ20]. RAM [GA22]. random [AD21, CFM^{+}22, FZC^{+}20, Jia21, LS23a, NHTH20, WDL^{+}21, ZT22b]. random-honest [FZC^{+}20]. randomization [NTI24]. Randomized [RZIX20]. randomness [WYZ^{+}20]. range [CSY^{+}20, MEL^{+}23, MNFQ24, SWL^{+}20, XCZ^{+}22]. rank [ABB^{+}21, MMPV22, XYH^{+}24]. Ranked [SZM22]. RANs [Hu20, SZO^{+}20]. Ransomware [ZXX^{+}20, ArMA^{+}21, DMSCA20]. RaP [GMF^{+}20]. Raphtory [SCC20]. Raspberry [SNM^{+}20]. Rate [WZJ^{+}22, KCY^{+}21, LFZJ21, LZJ^{+}20, YK20b, YNVRPD23]. rating [KZB^{+}23, ZIOT^{+}20]. ratio [DLH^{+}20]. rationality [BEL20]. Ratios [Che20]. raw [RLML20, WWY^{+}24]. ray [CLZ^{+}20, DLGW^{+}20, HZLH21, LHTSM^{+}23, MNFQ24]. rays [MSK^{+}21]. razor [GLF24]. RCNN [JLT^{+}21]. re [YLD^{+}23]. re-identification [YLD^{+}23]. reactions [GSMF20]. reactive [LBDP23, ODET21, RSLL21, SCP^{+}21]. read [EGD24, HMSA^{+}23]. readable [QHC24]. reading [LHD^{+}20]. readmission [WDL^{+}21]. ready [PBK^{+}22]. Real [AHH20, AEZ22, EET20, KP22, KBTM21, SNMWC21, XZTC22, AdSM^{+}22, AGV23, AYB^{+}22, CHL23, DA22, DSRG22, dRFRB24, FHGF20, GA22, HWH^{+}23b, JPJO22, LXZ^{+}20, MSV^{+}20, MXS22, dSOFC^{+}23, PBM^{+}22, RCP24, RAS^{+}22, SZVVB^{+}23, SGL^{+}20b, TCBF24, VMV20, XZJ^{+}20, YHW^{+}20, YLL22, ZT22a, ZLS23]. Real-time [AEZ22, EET20, KP22, SNMWC21, XZTC22, AdSM^{+}22, AGV23, AYB^{+}22, CHL23, DSRG22, dRFRB24, FHGF20, HWH^{+}23b, LXZ^{+}20, MSV^{+}20, MXS22, dSOFC^{+}23, PBM^{+}22, RCP24, RAS^{+}22, SZVVB^{+}23, SGL^{+}20b, TCBF24, VMV20, XZJ^{+}20, YHW^{+}20, YLL22, ZT22a, ZLS23]. real-world [GA22]. realistic [ZHP^{+}21]. realistically [MEC^{+}20]. Reality [SJQ20, FHGF20, YLKK20, YPEK23, ZZ21b, ZLL^{+}24a]. realize [PAS^{+}20]. realized [WHC^{+}24]. realizing [LCZB21]. realtime [ZLL^{+}24a]. reasoning [ACD^{+}20, BFG^{+}22, BSH^{+}21, BKV^{+}20, MRM^{+}24, WG21]. reasoning-enabled [BKV^{+}20]. rebalancing [LZLY20]. REBATE [CECS20]. ReBEC [GCN^{+}24]. reboot [FLG^{+}20]. recall [LFZJ21]. Receiver [WCWC20, WCWC19]. recharging [UKY^{+}20]. Recognition [GK21, Tau23, Tau24, AMM^{+}20, AEZ22, DHC23, uHA20, Fae21, FS21, FLF^{+}21, GZT^{+}21, HAqDE23, HHW^{+}22, JLT^{+}21, KAK^{+}23, LHF^{+}20, LZK21, LWLW21, LSL^{+}20, LZW21, LF21, LZC21, LYC^{+}22, MMU^{+}21, NN21, PKLC23, PLHC24, dAMVULM20, PYL22, RAS^{+}20, RSFB23, SNM^{+}20, SZL^{+}21, WYDB24, XLMC22, ZDC22, Zha21, ZZP^{+}23, ZLZ^{+}20b]. Recognize [HXWX23]. recommendation [AACJ23, CLZ21, FWX23, HMLS20, JWYÍ21, JYSH23, LY21, LQG^{+}23, LMZ^{+}22, LH20, OPLB24, OMPSPL20, TWL23, WML^{+}23, XWG^{+}21, ZT22b, ZZZ^{+}22, ZFZS23, ZWW^{+}23]. recommendations [BEKF21, YNN^{+}20, SVD^{+}20]. recommender [ACM^{+}21, GZB^{+}22, JKS20a, KZB^{+}23, OMSL20, PP20, SD20, TKS^{+}23, XZZ^{+}20a]. Reconfigurable [ISUC22, LJW^{+}20, DLHD22, GSCP22, TRB^{+}23a, YSL^{+}22, ZKD21]. reconfiguration [LZC^{+}23b, PBM^{+}22]. reconfigurations [AM22]. reconstructing [WXX^{+}24]. reconstruction [BN21, PBSS24, SGBC^{+}20, VHP^{+}22, Wu22, bZSC^{+}23]. reconstructions [WZC^{+}22]. Record [PSAL20]. recording [CP22]. recovery [MRD^{+}20]. recruitment [qLhZ20]. Recurrent [uRKI^{+}21, KMR^{+}22, LZP23, LC20, UAS^{+}20, WLY^{+}20, ZWL20, KLW^{+}21]. Recurrent-DC [KLW^{+}21]. recursive [KAH^{+}23]. RED [STH^{+}20]. Redesigning [TPN^{+}21]. redirection [SGL^{+}20b]. RedMule [TBB^{+}23]. reduce [ATK^{+}22, BEL20, DLC^{+}22, RK20, SB24, LLP^{+}20b]. Reduced [NRBC23, CSB23]. Reducing [BWX20, SXF22]. reduction [GYAW22, MSTN21, MDC^{+}24, OIG24, Pła24, SSC^{+}20, WDHY20]. Redundancy [ArMA^{+}21, ZZF^{+}24]. refactorings [TPF^{+}20]. reference [AAG^{+}20, CCL23, GBC^{+}24, TDMC23, WT24]. reference-points-based [CCL23]. referral [TDLT20]. refinement [XW21]. reflecting [WX23]. reflection [GdOAO20]. regarding [GDCGVG20]. Region [KF22, KF23, LZL^{+}24b, ZZZ21b]. Region-based [LZL^{+}24b, ZZZ21b]. Regional [JR22]. regions [PCVN21, ZGL^{+}23]. registration [LWS^{+}23b]. regression [AOSA20b, AMZZ23, WZH^{+}22, ZSL^{+}23b]. regression-based [ZSL^{+}23b]. regularization [ZHZS23]. regularized [LYC^{+}22]. regulation [AHN21, LZCH22, RSQS21]. REHAB [SVD^{+}20]. REHAB-C [SVD^{+}20]. rehabilitation [AMB^{+}21, KIJ^{+}24, LZS^{+}21, LZZX20]. Reinforcement [ASM^{+}22, BEM^{+}20, BEM^{+}24, dFCC23, DPLV23, KPA24, LWNH22, LCO^{+}23, MM23, PAM21, SHB22, SGS24, WLZ^{+}23, WGGB24, ZrHhH^{+}23, ZL23, ZWX^{+}23, AHSH22, AÇP22, BAİP24, BBN^{+}20, BMM^{+}24, CQS^{+}23, CHS^{+}23, CZH^{+}24, GRN20, HCG^{+}23, HJI24, rHZmH^{+}24, IA23, JHB22, JLW^{+}23, KS24, LYYG20a, LYYG20b, LZK21, LZW^{+}22, LGL^{+}20a, LWF^{+}23, LMZL24, LGL^{+}20b, MLN24, MKB23, MFE^{+}20, MNA^{+}23, MKBT24, RAA^{+}24, RYL20, RLL^{+}22, SMO^{+}24, SEL^{+}22, SXW^{+}22, SZO^{+}20, TWM^{+}23, WLY^{+}20, WLL21, WSL^{+}23, WS23, WWF^{+}23, WLL24, WYJ^{+}24, XW23, XWZM24, YWDC23, YCYO23, YNVRPD23, ZGK^{+}22, ZLZ23a, ZLZ^{+}23b, ZLL^{+}24b, ZWZB24, ZDLD24]. reject [HXWX23, LLZ20]. related [FHGF20, SK20b]. relating [DFG^{+}21]. relations [Gas22]. Relationship [YTQ19, YTQ20b, GWY^{+}20, YTQ20a, ZZZ^{+}22]. relationships [ZLM^{+}23]. Relative [MMR^{+}23b]. Relativistic [SMT^{+}24]. Relaxed [LWCC23, CdD20]. relaxing [AHMW23]. relay [AQN^{+}20, CDX^{+}23, WC22a, YL20b]. release [EL21]. releasing [SPL22]. Reliability [BPCM21, CZZ^{+}23b, CL20b, HSS20, RCP24, TWY^{+}23, YLX^{+}23, ZHL^{+}23]. Reliable [ASL22, BBP22, TLL^{+}24, BDM^{+}20, FBL^{+}20, RAA^{+}21, SKX^{+}20, ZDLD24]. relief [CPS^{+}23]. Remedy [RMA21]. remote [BVFGSF20, CMGI^{+}23, CDX^{+}23, CDX^{+}24, FBL^{+}20, GHEB^{+}18, GHEB^{+}23, HIMM20, KK20, LCZB21, Liu21, RGPGP24, RAB23, YXS^{+}23b]. remotely [ZLS22b]. rendering [FSBŞ^{+}20, ZLL^{+}24a]. Rendezvous [OÖA22, OÖA22]. Renewable [SZZY22, KHHT21, LGL^{+}20a, ZAH^{+}20]. repair [BWX20]. repairing [XGX20]. replacement [GCN^{+}24, GPGG23]. replacement-based [GCN^{+}24]. Replay [MDP24]. replica [YYW^{+}21, ZDLD24]. replicated [BLMT20]. replication [MRD^{+}20, SYYuR22]. report [DNNG21]. repositioning [LP21a]. repositories [BOM^{+}22]. Representation [ZZZ21b, CZCH24, DLC^{+}22, HJX^{+}23, JCK24, KSC20, KLW^{+}21, LQG^{+}23, LCY^{+}23a, LLD^{+}21, LPT22, LH20, SHT^{+}21, SOT24, SZS^{+}21, WLLY20, WXX^{+}24]. representations [QHC24]. represented [MMKS22]. Representing [BEKF21]. reproducibility [GSI22, JAS^{+}20]. reproducible [VG21]. REpulsive [CECS20]. REpulsive-BAsed [CECS20]. Reputation [UUK^{+}21, DG21, GL20, LWS^{+}23a, LLZ^{+}24b, MZLT21, TLL^{+}24, WLC^{+}20a, ZWX^{+}23, ZWY^{+}21]. reputation-based [LWS^{+}23a]. request [BBN^{+}20, KHHT21, PZHD20, ZLZ^{+}23b]. request-transmission [PZHD20]. requesters [KOM^{+}20]. requests [AALEF20, AAS^{+}20]. requirements [BSH^{+}24, DBBP24, SDZ^{+}20, VEH^{+}23]. Rescheduling [RAL^{+}24]. rescue [YZS^{+}21, NKG23]. Research [ALR^{+}20, CPM^{+}23, HRGL21, HY21, LBJ^{+}18, LBJ^{+}24, LZH^{+}20, LLZ^{+}21, LWJ^{+}21, MGB24, SCX21, Sun20, Wei21, WLR21, ZWL22, ZY21, AAA20, BDK^{+}20, CLC21a, CCP^{+}22, DRC20, FD21, HCB^{+}20, HBSG21, LCZB21, OCBO20, SGP^{+}20b, SJD^{+}20, TLX^{+}23, YBC^{+}20, YVW^{+}20, ZL21, ZWL21, wZcZN^{+}19, wZcZN^{+}20]. reservation [KAA^{+}21]. Reservoir [ZJW^{+}20]. Reservoir-based [ZJW^{+}20]. reshaping [KHES21]. resident [LHF^{+}20]. residual [QPL^{+}22b, QZZH21]. resilience [BdL20, PDT21, YBC^{+}20]. Resilient [SBMN21, CPYY23, MWK^{+}21, RSQS21]. resistance [MASRAM^{+}22]. Resistant [IPPK23, OdVP20, XRZ^{+}24, dVIP24]. ResNet [WCP23]. resolution [ASPG^{+}21, MFMSG20, PBSS24, QZZH21, YTQ19, YTQ20a, YTQ20b, YZW22, ZHP^{+}21]. resonance [HZX^{+}19, HZX^{+}20]. Resource [BEM^{+}20, BGMK22, CWYG23, HS24, KAF^{+}20, Li20, LYBS21, PKLC23, AOKÖ24, AQN^{+}20, ASA^{+}20, BLT^{+}24, uRBIBC20, BRK24, BKV22, BKG^{+}20, BSOK^{+}20, BJW22, CCW^{+}20a, CWM^{+}20, CS23, EET20, FCOJFM21, GSCP22, GMGV^{+}22, GPR23, GBC^{+}24, HZdLZ20, Hu20, HYC^{+}23, JYP24, KF22, KF23, KYPJ20, LYYG20a, LYYG20b, LYFZ20, LCH^{+}21, LYG^{+}24, LDWZ20, LMZ^{+}22, LZ22, LCO^{+}23, MZX^{+}24, MMMZ20, MVLJ21, MKB23, MSS^{+}24, MRM^{+}24, MM23, NGCB20, NPNC23, PLS^{+}23, PZLL21, QHCH24, QL22, RZA21, RRHA21, SJQ20, SCX21, SS21, SFC23, SGDG23, SHY^{+}21, TK24a, TSB20, TCBF24, WLP^{+}20, WZW^{+}20, WCY^{+}21, WC22b, WXZZ22, WWS^{+}23a, WSL^{+}23, WCS24, WLY23, gWLWZ21, XGS^{+}20, XGY^{+}23, YZL^{+}24, YPZ^{+}24, YJF^{+}20, YZS^{+}21, ZWZ^{+}24, ZYX^{+}20, ZLXH20, ZLZ^{+}20a, ZLW^{+}22, ZGK^{+}22, ZLS^{+}22a, ZWZB24, ZG24, SCR20]. resource-allocation [GMGV^{+}22]. resource-constrained [ASA^{+}20, BJW22, MKB23, RZA21, YZL^{+}24, ZG24]. Resource-efficient [CWYG23, PKLC23]. resource-limited [ZLZ^{+}20a]. Resources [KCR20, PFP^{+}22, AKE22, BHSH22, CLY^{+}20, HGY^{+}22, dSOFC^{+}23, TC23, ZTB23]. respiratory [CHJ^{+}20]. Response [CCBFI^{+}23, TDM^{+}22, ICBB20, IA20, JKS20b, JKS20c, NAC^{+}22, SKS22, SCX21, WGGB24, WCWC19, WCWC20]. Responsibility [NJB20]. Responsibility-based [NJB20]. responsibly [GMMR24]. responsive [SW22, SCA22]. RESTful [SKH20]. restrained [LYFZ20]. Restricted [JCP^{+}20, BCC^{+}22, LWJ^{+}21]. restructuring [ZSZX24]. result [YZR23]. results [BLMT20, FAA^{+}23]. resupply [PCI^{+}24]. Resurrecting [AKCP21]. retention [ERL^{+}20, WFLL22]. retirement [Fae21]. Retraction [AB21, ABGMC21, ABM21, ABMESM22, ABMM22, ABMMC22, Bo20b, DP21a, DP21b, GHEB^{+}23, HZX^{+}20, JLC^{+}20, LBJ^{+}24, LYYG20b, LCLW24, WWP20, WCWC20, YWG^{+}20b, YTQ20a, YTQ20b, ZMZ^{+}20, wZcZN^{+}20]. retransmission [WW20]. Retrieval [HDN^{+}20, See20, CDX^{+}24, LHY^{+}20a, MYT^{+}21, SCZ^{+}20, TLC^{+}20, Tao23, YJH^{+}20, ZCF21, GXS22]. reusable [RHJ20, ZKD21]. reuse [DSDV20, VAKB23]. reuse-oriented [VAKB23]. Revenue [XWD20, XWZM24]. Revenue-maximizing [XWD20]. reverse [TLL^{+}24, ZPK^{+}23]. reversible [PSAL20]. Review [HTGW^{+}23, LLW^{+}23b, BHH22, CRdRR^{+}22, Cha20, HJI24, ICGGAR22, KPGD24, MRS^{+}22, MBB24, MKK^{+}20, NAK^{+}22, OLP23, OOZ^{+}23, OCBO20, PGHS20, SN23, SDVC22, SSA^{+}23, SZZY22, STK23, YPX^{+}20, ZT22a]. Reviewers [Tau23, Tau24]. reviews [AOSA20a, LDM^{+}21, VPA20]. Revised [SZO^{+}20]. Revisiting [BQC22]. revocation [CA21, LCB^{+}20]. reward [BAİP24]. Rewarding [LDM^{+}21]. RF [WHC^{+}24, CXHS20]. RF-IDH [CXHS20]. RFID [CXHS20]. RICDS [YLL22]. Rich [LTX^{+}24, TDLT20, VCG^{+}23]. rich-club [TDLT20]. ride [LLZ^{+}24a]. ride-hailing [LLZ^{+}24a]. ridesharing [LHA20]. ridge [AOSA20b]. rigging [QL22]. rigorous [hAS24]. ring [NTA^{+}22]. ring-sector [NTA^{+}22]. RIOMS [YWH^{+}21]. Ripley [WGW^{+}20]. RISC [GMFC23]. RISC-V [GMFC23]. Rise [CGM^{+}23]. rising [DAA^{+}21]. risk [AB19, ABGMC19, AB21, ABGMC21, AHSH22, BAGRB^{+}20, LLC^{+}23, PDFV21, RJM^{+}21, SYG^{+}20, SZVVB^{+}23, SCÁB20, THB23, WCWC19, WCWC20]. risks [AB19, ABGMC19, AB21, ABGMC21, Kho21a]. River [RGPGP24]. RKD [ZGK^{+}22]. RKD-VNE [ZGK^{+}22]. RL [BEM^{+}20, GPR^{+}24]. RL-based [GPR^{+}24]. RL-OPRA [BEM^{+}20]. RLDS [QPL^{+}22b]. RMA [CSB23]. RMA-CPABE [CSB23]. RMKABSE [SZM22]. RNA [ABGDT23]. RNN [BNA^{+}21]. Road [AOKÖ24, BDG23, CIB^{+}20, FCGPSG^{+}21, WC22b, YL20a, ZPLQ20, YYB^{+}21]. roadmap [FAA^{+}23]. roads [YWH^{+}21]. robo [AAR^{+}20]. Robot [LCCP21, AR20, HJW^{+}20, IHA^{+}20, LZK21, LSB21, MDL^{+}23, TLKX21, XLCB20, YWG^{+}20a, ZTC20, ZWL21, HJW^{+}20]. robot-assisted [XLCB20]. robotic [AGV23, LSH^{+}20, LGC^{+}21, SHKW23]. robotics [GMA^{+}22, YWG^{+}20a, ZWL22, ZLG^{+}24]. robots [LZS^{+}21, LZF^{+}24, LZZX20, SZS^{+}21]. Robust [LZTM24, NED^{+}20, WMU^{+}24, WJC^{+}24, YFQ^{+}22, ZZG^{+}22, ASAM20, ABB^{+}21, AMZZ23, BR20, FGG^{+}23, HRY^{+}21, JLK22, LSN^{+}20, SXC^{+}23, VMM20b, WLLY20, WZHX23, WWF^{+}23, WPX^{+}23, XCGZ24, ZPLQ20, ZLPZ21, ZYF^{+}22]. Robustness [LCO^{+}23, AB24, GL20, MJW23, NTI24, XCW20, ZRH^{+}23]. ROC [WCWC19, WCWC20]. rock [Kha24]. ROI [PBSS24]. Role [RMA21, BOM^{+}22, GZF^{+}20b, LHW^{+}23, LLY^{+}20, RGPGP24]. role-based [LLY^{+}20]. roles [LLY^{+}20]. Roofline [MIMS20]. RoofSplit [HZS^{+}23]. root [HSGY20, SDV^{+}21, HSGY20]. roots [SBF^{+}21]. ROS [OLLP24]. ROS-based [OLLP24]. rotate [HZL22]. round [XGY^{+}23]. rounds [WYZ^{+}24]. route [KHB20, LWS^{+}23b, MLN24]. routes [ZTB23]. Routing [OCSCB22, SuRMA^{+}23, AM20, FNRP20, Gul22, KSH^{+}21, LGW^{+}21, Liu23, MAS23, MNA^{+}23, NNH^{+}20, QGH^{+}22, SWW^{+}20, SZW^{+}23, WZTL20, XWZM24, YLSL22b]. ROVs [ZLS22b]. row [SCK^{+}22]. RPL [VSPM21]. RPL-based [VSPM21]. RRT [MDL^{+}23]. RSSI [KAH^{+}23, MR23a]. RTGEN [MMR^{+}23b]. RUAD [MBC^{+}23]. Rule [CKFT20, HTAY21]. Rule-based [CKFT20]. rules [GFM^{+}20, TIA21, YLC23]. Run [GRG20, XCGZ24]. Run-time [GRG20, XCGZ24]. running [FPL24, LLC^{+}23, SHZMA21]. Runtime [PSHW20, AGYS20, FSP^{+}24, GMT23, SCW^{+}22, ZBTV^{+}20]. runtime-aware [SCW^{+}22]. runtimes [BLT^{+}24].

S [LYH^{+}21]. S-Mask [LYH^{+}21]. SaaS [MVLJ21]. SAC [JRW^{+}20]. Saca [TWY^{+}23]. Saca-FI [TWY^{+}23]. safe [BCC^{+}22, BQI^{+}20]. safely [LLC^{+}23]. Safety [FMN^{+}20, CMF^{+}21, HH22, IB20, SSDC22, TLKX21]. Saiph [MMFAB23]. salesmen [DZXS21]. SAMbA [GMF^{+}20]. SAMbA-RaP [GMF^{+}20]. sampled [RWG21]. samples [JQZ^{+}22, ZL22]. Sampling [CABB20, Jia21, UPD^{+}20, WCY^{+}21, WLYL20, ZJW^{+}20]. sanitization [DVEE^{+}20]. SAR [RLML20]. Sarafu [BZG23]. SARL [GRG20]. SARSA [GPR^{+}24]. satellite [ASPG^{+}21, CDX^{+}23, QGH^{+}22, XJL^{+}24, YPZ^{+}24]. satellite-ground [CDX^{+}23, XJL^{+}24]. satisfaction [LHA20]. saturation [GdOAO20]. Saudi [ARB20]. Saving [GRN20, BJ22, YSL^{+}22]. SBAC [SP24]. scalability [AL20, KRW^{+}20, TM20]. Scalable [BP20, GFZ21, HDD20, LFYH22, NAK^{+}22, AHL^{+}23, ABAD22, ANA24, BBTC20, CdST^{+}20, CLV24, CZH^{+}24, CSH^{+}23, CCBFI^{+}23, DPPGCCA23, EGD24, FSBŞ^{+}20, HTAY21, KS24, LZHL23, LJL^{+}21, MFMSG20, NQB^{+}23, OdVP20, PMCP20, PK22, ST20b, WCY^{+}21, WS23, XLX^{+}21]. Scale [HZL22, MM21b, BHL^{+}20, CLLCK20, CGMT20, CECS20, CCL^{+}21, DRD20, DDM21, DZXS21, FCOJFM21, GW20, GPWL20, Hu21, HZX^{+}24, IAM^{+}22, LXH^{+}21, LZW^{+}22, LWCC23, LCY^{+}23a, LXY21, LZC21, ONK^{+}20, PDA^{+}20, RZIX20, RLZW21, SGDK^{+}21, SEL^{+}22, SW20, TLC^{+}20, TDL^{+}21, TGAP20, WG21, WJC^{+}24, WF21, XCW20, Xu21, YSL^{+}24, YYL22, YJQ^{+}23, ZC22, dSGST21, FAŞ^{+}20]. scale-free [CECS20, XCW20]. Scaling [AIM23, GDS^{+}20, WMNV20, CS23, EET20, LZTM24, RSLL21, RZA21, SSWW23, SDZ^{+}20, SPL24, TK24a, YPL24, ZRH^{+}23]. scan [KYY^{+}20]. scar [YCG^{+}20]. scavenging [SBF^{+}21]. Scenario [GW22, CDP20c, NNUV20, Par22, ZGK^{+}22]. scenarios [AMBD^{+}20, CCC^{+}21, JPJO22]. scene [YJQ^{+}23]. SCERM [IA20]. Schedule [WZJ^{+}22]. scheduler [PK22, QSZ^{+}24, THA^{+}24, gWLWZ21, XHW20, EMHE21]. schedulers [dFCC23]. schedules [EKJ^{+}20]. Scheduling [AALEF20, AMBGS21, LXC^{+}24, MAM^{+}24, NGC24, WXC^{+}24, WLD^{+}20b, ABMESM18, ABMESM22, ABC^{+}20, ATZP21, BGR20, BLMT20, BBN^{+}20, BPUW24, BGMK22, CCL^{+}20, CQS^{+}23, CLL^{+}23, CCW^{+}20c, DT21, DK20, DFZ^{+}20, EAA21, EMHE21, GRN20, GBM20, GMT23, GB20, HCWD21, HSS20, HdOP^{+}21, HJI24, HZdLZ20, Hu20, HKB^{+}24, HWR^{+}22, IT20, JHB22, JXYC24, KF22, KF23, KPL22, KDX^{+}24, Kha24, KPA24, LWX22, LSH^{+}20, Li20, LZW^{+}22, LTX^{+}24, LDGS20, LYS^{+}20, LH21, LYBS21, LWF^{+}23, LCO^{+}23, LFHS23, LZHS24, LEXH20, MGW23, MA24, MhCEANSM20, MKB23, MR23b, MDG^{+}22, OPOG23, PKR21, PP24, QPL22a, yQhJL20, QWR^{+}20, RCJZ20, RRD21, RBW20, RKP^{+}21, SJQ20, SCX21, SACW23, SHY^{+}21, TLW^{+}24, TSB20, VI21, WGG^{+}20, WZTL20, WLL21, WGW^{+}21, WS23, WGGB24, WCS24, WWW^{+}24, XZYH22, XZK^{+}20, YSL^{+}24, YJLC20, YYXZ23, YXL^{+}20, YLTH22, YPZ^{+}24, YXYH20, YSL^{+}22, ZLZ^{+}20a, ZLZ23a, ZLZ^{+}23b, ZGN^{+}20, ZLL^{+}23, SCR20]. Schema [CSD^{+}23, TD21]. schema-agnostic [TD21]. Scheme [HAA^{+}20, WLX^{+}24, ASSG22, AYY^{+}20, ATK^{+}22, BYR^{+}20, BGCL20, CDG^{+}20, CSB23, CDX^{+}24, DKD22, FBL^{+}20, FWZ^{+}20, FYHZ24, FLH^{+}24, GCT^{+}20, GKB^{+}20, GAT^{+}20, GWW^{+}22, GKA^{+}21, GZG20, GPC21, HYL^{+}20, HGK20, IA23, IA24, JPW20, JKS20b, Kad20, KSH^{+}21, Kol22, LCB^{+}20, LJL^{+}21, LLCH21, LWJ^{+}23, LLW^{+}23a, LLD^{+}21, MSZ^{+}20, MK20, MISS22, MLZ^{+}23b, PLBOC20, PKLC22, PZHD20, RKM23, SIG24, SK20a, SLS^{+}20, SN21, SK23, SXHD24, SLX^{+}24, TSR^{+}20, TKS^{+}23, TDL^{+}21, UAACH21, UJHN20, WHF^{+}20, WLC^{+}20a, WHZ^{+}20, WFLL22, WSC^{+}23, WZS^{+}23, WLS^{+}24, WZS^{+}22, WXZ23, XLH^{+}24, XRHS21, XYH^{+}24, XCB^{+}20, XRZ^{+}24, XXL^{+}24, YLKK20, YVSG22, YhSL^{+}22, YZW^{+}23, ZZQ21, ZBF22, ZZF^{+}24, ZHLL24, ZSZ^{+}24, ZPK^{+}23, ZDLD24, ZZZ^{+}21a, ZSL^{+}23b]. Schemes [PRF20, BCCS20, LWH^{+}22]. scholar [XFJ^{+}20]. scholarly [DOR^{+}21, PP20]. Scholars [GWY^{+}20]. Science [BOM^{+}22, DPG20, MM21b, MGM^{+}20, BLGCLA^{+}23, BFG^{+}22, CBN^{+}20, DLGW^{+}20, GDP20, LZB20, SGDK^{+}21, SGSGGC^{+}23, SJD^{+}20, YVW^{+}20, dSGST21]. Sciences [CK20, CK24, JAS^{+}20, SLA^{+}23]. Scientific [GMF^{+}20, CLLCK20, CCP^{+}22, CCdS23, CDBD24, DGL^{+}20, DK20, GSDGP21, GMM22, GPR^{+}24, HdOP^{+}21, MMKS22, MGZ^{+}20, MEC^{+}20, OGO^{+}20, PWV^{+}21, QRS^{+}21, RRHA21, SPRA21, SSC^{+}20, STK23, TRB^{+}23b, WGG^{+}20, WGW^{+}21, YTW^{+}20]. sclerosis [WLZ^{+}20]. scoping [OOZ^{+}23]. score [DG21]. scores [Bo19, Bo20b]. scoring [NADY20]. screening [BCT^{+}21]. SCTD [MLZ^{+}23b]. SD [WLX^{+}24]. SD-SRF [WLX^{+}24]. SDC [hAS24]. SDD [XSW^{+}21]. SDG4 [MIIN23]. SDN [SME^{+}21, AZA23, AKF^{+}20, AYB^{+}22, BÖ20a, BSM20, GCT^{+}20, KHH21, MSLP24, MGC23, NPNC23, NCLP21, PNL^{+}21, QGH^{+}22, RBA^{+}22, SME^{+}19, SHB22, SHST20, UPD^{+}20, XWZM24, YLSL22a, YLSL22b, YNVRPD23, YD21]. SDN-assisted [AKF^{+}20]. SDN-Based [SME^{+}21, NPNC23, PNL^{+}21, QGH^{+}22, RBA^{+}22, SME^{+}19, YD21]. SDN-Blockchain [SHB22]. SDN/NFV [YNVRPD23]. SDN/NFV-based [YNVRPD23]. SDNs [LQML22]. SDS [SK20a]. sea [ACF^{+}21]. Seamless [GMP^{+}20a, PPSC23]. SEAP [WYG^{+}20]. Search [OÖA22, ABMESM18, ABMESM22, BEON24, DZXS21, FW22, HYC^{+}21, HY21, JGL^{+}20, KSSR20, LCC^{+}24a, LLS24, LYY^{+}22, LWY^{+}24, NBJ21, PLBOC20, RDR^{+}24, SXF22, TTZ^{+}21, TLX^{+}23, WCW^{+}21, XLZ^{+}22, YZL^{+}23, YLH^{+}23, YZR23, ZLF^{+}23a, ZHLL24, ZTB23, dVIP24, JCW^{+}23, NKG23]. Searchable [SZM22, WZS^{+}23, ATK^{+}22, LMW^{+}24, TSR^{+}20, WCXW22, XYH^{+}24]. Searchain [JGL^{+}20]. seasonal [YLG^{+}24]. SecBPMN2BC [KMS23]. seccomp [XCS^{+}22]. SecFedNIDS [ZZG^{+}22]. second [FAA^{+}23]. secondary [ZHGX20]. Secrecy [GPC21]. secret [KK20, SLS^{+}20]. SeCrowd [CCHD21]. sector [ARB20, NTA^{+}22]. Secure [FFB20, JPMR21, Man20, NIB^{+}21, NSR^{+}23, Pan20, RMA21, TDS^{+}22b, VS20, WC20, WYS20, YLH^{+}23, ABMM18, ABMM22, AZA23, AAM^{+}24, BBB^{+}20, CCHD21, CCBFI^{+}23, EUEU24, FWZ^{+}20, FYHZ24, GWW^{+}22, GKA^{+}21, GLW^{+}20, HN22, KAK20, KBTM21, KMS23, LYY^{+}22, LWZ^{+}23a, MSBAU24, NNH^{+}20, PSAL20, RJA^{+}22, RHWY23, SVN^{+}20a, SK23, TKS^{+}23, TBH23, UAACH21, VMM^{+}20a, WZW^{+}20, WLAC20, WZS^{+}23, XLH^{+}24, XCB^{+}20, ZSL^{+}23a, ZLF^{+}23b, ZZG^{+}24, ZHLL24, ZXZ^{+}23, NNH^{+}20]. Secure-GLOR [NNH^{+}20]. secured [DS23, GMAL23, UJHN20, WXD^{+}23]. SecurePrivChain [SMKA23]. Securing [SKA^{+}20, YNN^{+}20, LCLA21, SMKA23, TAM21]. Security [ASSG22, ADdMM20, Elg20, KZG^{+}22, LHH^{+}21, MBJ^{+}20, QJS^{+}21, SME^{+}21, STK23, TG20, YDK20, ZWZB24, AT20, AM22, BS20, CFC^{+}20, CZZ^{+}23b, CDP20a, DRC20, DLHD22, FZT^{+}23, GKB^{+}20, GIPS20, HRGL21, rHZmH^{+}24, JRW^{+}20, KRW^{+}20, LOH^{+}23, LWX22, LJC^{+}20, LLW^{+}23b, LYP^{+}24, LZF^{+}24, LZ20b, LWHW22, MCT^{+}22, MTT^{+}23, MS20, MLZ^{+}23b, MRR^{+}20, MPP^{+}21, MCF20, MBB^{+}20, ORLV20, QKG20, RHK^{+}23, RWG21, RCR21, SME^{+}19, SCÁB20, SVN^{+}20a, SAAEK22, SXHD24, VPSC^{+}23, WWC^{+}24, YLM23, YBX^{+}23, YL20b, YDL^{+}20, YYL22, YFL^{+}24, ZY20, ZLZ^{+}20a, ZWZ^{+}23, ZHL^{+}23, NHY20, SUKN22]. security-aware [LWX22]. security-critical [GIPS20, ZLZ^{+}20a]. security-privacy [RHK^{+}23]. seed [DFZ^{+}23]. SEENS [ZWH^{+}21b]. Segmentation [WTL^{+}20, BAİP24, HIdAR^{+}20, HZL^{+}21, JLT^{+}21, LHXL22, LC20, MXW^{+}23, QJZ^{+}20, QC21, WLZ^{+}20, WCHA20, ZP22, ZWH^{+}21b]. segmentations [YCG^{+}20]. seismic [KBG20, LYGF21]. seizure [AEZ22]. selected [ZT22a]. selecting [ASB^{+}23]. Selection [ArMA^{+}21, BAK22, DIB20, HCK20a, HCK20b, LHY^{+}20a, STS^{+}20, XZD^{+}21, ABMMC18, ABMMC22, ARA^{+}22, ARA^{+}23, BMM^{+}24, BHH22, CKZ^{+}22, FZC^{+}20, FCOJFM21, GSKS20, GDEBC20, HYWY22, JT22, LQML22, LZCH22, MHH^{+}20, MASRAM^{+}22, MSA^{+}24, PBS23, Pła24, PBSS24, QHCH24, RMBMT21, SXC^{+}24, SVN20b, TA23, WC22a, WLW^{+}21, WYZ^{+}24, XWLC20, XWK21, XWZM24, YLZ^{+}24, ZZZX22]. selection-based [ZZZX22]. selective [ZWH^{+}21b]. selectively [hAS24]. Self [BBD^{+}21, CWM^{+}20, JCX^{+}21, KCB20, LAS20, LCY^{+}23a, RFP22, SEKS^{+}20, SK21a, SAF23, bZSC^{+}23, AGdS^{+}21, AABKB22, AP20, AKE22, BdL20, BDT21, BCSS20, CG21, DCGM20, FGP20, FW22, FMN^{+}20, FGB21a, FGB21b, FWX23, GWZ20, GAP24, HTXW21, ICW21, MR23b, PCVN21, QNRA23, SGSGGC^{+}23, WFLC22, WCL^{+}24, ZFMB20, ZXX^{+}20, ZWW^{+}23]. self-adaptable [SGSGGC^{+}23]. self-adaptation [AKE22, FMN^{+}20, HTXW21]. Self-adapting [SEKS^{+}20]. Self-adaptive [CWM^{+}20, bZSC^{+}23, AABKB22, BdL20, BCSS20, DCGM20, ZFMB20]. Self-assemble-featured [LAS20]. self-attention [FWX23, GWZ20, WCL^{+}24, ZXX^{+}20]. Self-aware [JCX^{+}21]. self-determination [AP20]. Self-distributing [RFP22]. self-driving [WFLC22]. Self-improving [BBD^{+}21, BDT21]. self-integrating [BCSS20]. self-integration [FGP20]. self-learning [FW22]. self-organising [FGB21a, FGB21b, PCVN21]. Self-organizing [SAF23]. Self-Protecting [SK21a, CG21, ICW21]. Self-supervised [KCB20, LCY^{+}23a, ZWW^{+}23]. self-supervision [GAP24]. self-sustainable [MR23b]. selfish [ZWX^{+}23]. Semantic [CDRS20, JLT^{+}21, LWZ^{+}20, PLMZ23, STK20, SGP^{+}20a, Tao23, ZLM^{+}23, BAGRB^{+}20, DMPS23, KHB23, LZ20a, LWLW21, LC20, PSvL^{+}20, QG20, QHC24, RGDMMR^{+}23, WYX^{+}23a, XKK20, XZ20, XXY^{+}23, YLGZ21, ZC22, ZP22, OLP23]. Semantic-aware [LWZ^{+}20]. Semantic-based [STK20, QHC24]. semantically [CCBFI^{+}23]. SemanticPeer [ZC22]. Semantics [ADRP23, AD21, LLD^{+}21, POMK20, ZLPZ21]. Semantics-based [ADRP23]. Semi [CBC^{+}20, WWH^{+}21, DMSCA20, JMA^{+}21, LLZ20, XZC^{+}24, ZCWC20]. Semi-automatic [CBC^{+}20]. semi-autonomous [DMSCA20]. semi-decentralized [XZC^{+}24]. Semi-supervised [WWH^{+}21, LLZ20, ZCWC20]. semitrailer [ZPLQ20]. senile [WWP19, WWP20]. SenseChain [KOM^{+}20]. sensed [YRV^{+}22]. Sensing [qLhZ20, WSC^{+}23, Ano24s, CDX^{+}23, CDX^{+}24, CdO20, DP19, DP20c, DP21a, DP21b, uHA20, GHEB^{+}18, GHEB^{+}23, KSS^{+}20, MKBT24, NIB^{+}21, PPG^{+}20, RGPGP24, TRB^{+}23a, Wu22, YZZ^{+}23, ZLM^{+}23, ZHGX20, bZSC^{+}23]. sensitive [GDGK20, JWYÍ21, WSJ^{+}21]. sensitivity [ZrHhH^{+}23]. sensitivity-aware [ZrHhH^{+}23]. Sensor [ACC20, Gul22, KBG20, LZP23, MR23a, MR23b, WZB^{+}20, Zhu21, AAA20, AOSA20b, CCTZ22, CHJK22, DVV^{+}20, FMN^{+}20, GAT^{+}20, GDGK20, LS23a, LZA^{+}20, Liu23, MLZ^{+}23a, MNA^{+}23, NTA^{+}22, PKLC23, PLHC24, Pła24, PYL22, RAS^{+}22, RSFB23, SAM^{+}24, SDV^{+}21, TLKX21, WC22a, WLAC20, ZLZ23a, ZFZS23, ZMJ^{+}22]. Sensor-based [LZP23, PKLC23, PLHC24, PYL22]. Sensor-Cloud [WZB^{+}20, ZLZ23a]. sensors [AESI^{+}21, HH22, LHF^{+}20, PGHS20, RPŠŠ23, ZZP^{+}23]. sensory [SEKS^{+}20]. sentence [SGL^{+}20a]. sentence-level [SGL^{+}20a]. sentences [CZ20]. senti [BDFR22]. senti-mental [BDFR22]. Sentiment [RHM20, ALS21a, AR20, BNA^{+}21, BDFR22, GWZ20, GFZ21, GDCGVG20, NRMI20, OCBO20, TIA21, UAS^{+}20]. sentimental [VPA20]. separate [GMI22]. separating [PCI^{+}24]. Separation [MECRFD20]. September [Ano20p, Ano21p, Ano22t, Ano23m, Ano24p]. Seq2Seq [RKI^{+}23]. Sequence [BMBC20, BYW^{+}21, JZM^{+}22, PCACTMÁ24, WLC23, ZLPZ21]. sequence-based [PCACTMÁ24]. Sequence-to-sequence [BMBC20]. sequences [GIPS20, LLP^{+}20a]. Sequential [HLW^{+}23a, FWX23, PDFV21, SPWW21, TWL23, WJC^{+}24, WML^{+}23]. serial [LLP^{+}20a]. Series [IB20, CYWS24, CIJM20, GZZG24, HN22, HSGX22, LWH^{+}22, MASRAM^{+}22, Ngu24, PCACTMÁ24, SDVC22, SOKW^{+}20, Yan21, ZCQ^{+}23, ZWL20, ZZ24]. serious [AUJW22, EKK23]. Server [TPD^{+}24, KK20, LJW^{+}20, LLT22, LMW^{+}24, MC20, XZZ^{+}20a, XZTC22, ZLF^{+}23a, ZSL^{+}23a]. Serverless [CRdRR^{+}22, CREE^{+}24, CS24b, MGZ^{+}20, MTD^{+}24, NUB24, PDJS22, WLX^{+}24, ABMO24, ALS^{+}21b, AFL23, ATC^{+}24, BBF^{+}24, CdRRdC^{+}24, EET20, FTS^{+}24, rHZmH^{+}24, HS24, KS24, LIP^{+}24, MKB23, MABK24, MWS24, Ngu24, PPSC23, PPA^{+}24, RRD21, RAL^{+}24, RPF21, RFP^{+}24, THA^{+}24, TK24a, TCBF24, WT24, YCYO23, ZWZB24, ZLL^{+}23]. Serverless-like [CREE^{+}24]. Serverless-operated [WLX^{+}24]. servers [QWR^{+}20, ZTQ^{+}20]. Service [BOD^{+}24, GdOAO20, HKS23, HCK20a, HCK20b, MABK24, MZZ20, NNPP23, NUB24, SuRMA^{+}23, WLX^{+}24, uRKI^{+}21, ASA23, ASHO20, ABAJ20, ASAM20, AL20, AAR^{+}20, BCC^{+}22, BSB^{+}22, BPC^{+}24, BSOK^{+}20, BYH^{+}20, CSY^{+}20, DCGM20, DWZ20, FWX23, GFPB23, GSKS20, GPGG23, HJW^{+}20, HS24, KHHV21, KXZW23, LOH^{+}23, Li20, LWNH22, LPS^{+}24, LDGS20, LDDL21, LQYL21, LLW^{+}22b, LZCH22, LLT20, LZZ^{+}20, MDZ^{+}21, NKG23, ORLV20, OPLB24, OCMJFB^{+}23, PCC21, PSC^{+}21, PSvL^{+}20, PDT21, QMCX20, QPL22a, RCR21, RRHA21, SMC23, SEL^{+}22, SJQ20, SYXL22, SKH20, WGLH20, WLY^{+}20, WLZ^{+}23, WYJ^{+}24, YGD^{+}21, YXL^{+}20, YCS^{+}20, ZAH^{+}20, ZGY20, Zha20, ZTC20, ZGL^{+}23, MA24, MOU^{+}21]. service-based [LPS^{+}24]. service-oriented [YGD^{+}21, YCS^{+}20]. Services [ETH20, TDC^{+}20, ABT20, ACI^{+}23, BLGCLA^{+}23, BPCM21, BOD^{+}24, BBB^{+}20, CPM^{+}23, CWM^{+}20, CYZ^{+}22, CSH^{+}23, DSC24, ENT^{+}22, GVCUGF20, GZG20, GPC21, LKE22, LMNC22, LHD^{+}20, LLFQ21, LDDL21, LDW^{+}21, LZC^{+}23b, LLZ^{+}24b, dTGC20, MBM^{+}20, MDT^{+}20, MSS^{+}24, MGC23, RBSK23, RSLL21, SNM^{+}20, SGSGGC^{+}23, SEKS^{+}20, SBMN21, SZdLZ22, SPL24, TDLT20, TC23, VS20, WHF^{+}23, XCL^{+}20, XXY^{+}23, YZX^{+}23, YXS^{+}23b, ZPQH21]. serving [PP24]. Set [FLH^{+}24, bHFF^{+}21, LYY^{+}20b, PSH^{+}20, WLW^{+}21]. sets [ABMMC18, ABM19, ABM21, ABMMC22, ONK^{+}20, OWK^{+}23]. setting [See20]. settings [STK20]. SFL [QCWY23]. sFlow [UPD^{+}20]. SG [MAM^{+}24]. SG-PBFS [MAM^{+}24]. SGD [GOA23]. SHA [VDSB22]. shadow [LYFZ20]. Shafer [RTD24, UYH21]. shallow [JL21, LWJ^{+}23]. SHAMC [WYS20]. shaping [BAİP24]. Shapley [DLZ^{+}23]. share [TBO20]. Shared [KKT^{+}23, TSB20, DCC22, DSC24, LZS^{+}24, LGM^{+}21, LL20, MVLJ21, QRS^{+}21, WLYL20, dAdSM^{+}22]. shared-memory [LGM^{+}21, LL20]. shared-use [QRS^{+}21]. Sharing [KCR20, BCC^{+}22, CGFC20, CFC^{+}20, CXWY21, CDV^{+}24, HN22, JMHB24, Kad20, KK20, LAHN22, PBL^{+}23, RNA21, SMKC20, SIG24, SLS^{+}20, WXD^{+}23, WWC^{+}24, WLY23, XZH^{+}23, YVSG22, ZSL^{+}23a, ZXW^{+}20, ZXZ^{+}23]. Sharpening [DCZ20]. shell [PMMG^{+}20]. Ships [SME^{+}21, SME^{+}19]. shooting [WSL21]. shop [ABMESM18, ABMESM22, LEXH20, MDG^{+}22]. Short [ACF^{+}21, XCL^{+}20, CHW^{+}20, CLQS20, DWZ20, EGD24, HMSA^{+}23, MZLT21, MEL^{+}23, NED^{+}20, WLL22, YFQ^{+}22, YLGZ21]. short-range [MEL^{+}23]. short-read [EGD24]. Short-Term [ACF^{+}21, XCL^{+}20, CHW^{+}20, CLQS20, DWZ20, MZLT21, NED^{+}20, WLL22, YFQ^{+}22]. Shortest [MAM^{+}24, LGW^{+}21]. shot [LLW^{+}22b]. shoulder [ÇYZZ21]. Shuffle [AKJJ20]. SI4IoT [ACI^{+}23]. side [LP21a, LQG^{+}23, OPLB24, RPP^{+}20]. sided [BMBE20, WCD^{+}22]. sign [CdRRdC^{+}24]. signal [BÖ20a, dRFRB24, HWH^{+}23a, NAC^{+}22, OMPSPL20, PGHS20, RYL20, SQ22, WLLF20, WHF^{+}23, XLS^{+}21, YJB^{+}21]. signal-based [WHF^{+}23]. signaling [ZNZ^{+}23]. signals [AMZZ23, HZPS21, dTGC20, Wan20, XYL^{+}20]. Signature [FQH^{+}24, GCH^{+}22, CXHC23, KJYC23, LWZ^{+}23a, SK20a]. Signatures [SK20a, CPYY23, CCHD21, RAB23]. signcryption [FWZ^{+}20, GWW^{+}22, ZWWC21, BKHD20]. signed [HLT^{+}21, LQNW20, MAQ^{+}20]. silo [OCMJFB^{+}23]. SIMD [BHL^{+}21, PFGDM22]. similar [QNRA23]. Similarity [HLT^{+}21, NADY20, TGJ^{+}20, KZG^{+}22, LP21a, PP20, SHF23, YLGZ21, YLF^{+}23, YGS^{+}22, YZR23, ZXY^{+}21, ZPS^{+}24]. Similarity-based [TGJ^{+}20]. SIMPA [GVŠ22]. Simple [LZH^{+}20, KMR^{+}22]. simplex [NQB^{+}23]. Simulated [MÖ24a]. simulating [MEC^{+}20]. Simulation [SMBB^{+}24, BRM^{+}20, CBC^{+}20, CHKJ20, DA22, DC21, KSC20, LGZ^{+}24, PMMSE21, SUKN22, SZZY22, SFC23, SWL^{+}20, TBH23, UKY^{+}20]. simulation-based [BRM^{+}20]. simulations [BDGG^{+}20, CPJ^{+}21, MYM^{+}21, RKI^{+}23, SCA22, ZGW^{+}23a]. Simulator [CH24, AdAHK20, DPPGCCA23]. simulators [CdST^{+}20, dAPHOMPJ20]. Simultaneous [YCG^{+}20]. since [SN23]. Sine [MSA^{+}24]. Singapore [PABBA20]. Single [LFM^{+}22, RMA21, BJP^{+}20, BPUW24, CABB20, NTY^{+}21, SCK^{+}22, YYKK20]. single-row [SCK^{+}22]. single-view [NTY^{+}21]. singular [LZJ^{+}20]. Site [BHSH22, PCC21, PKLC22]. sites [KXZW23, PGSM^{+}24]. situ [LS23b, WDS^{+}23]. situated [BEL20]. situation [MWK^{+}21]. situation-aware [MWK^{+}21]. situational [CDP20a, SZVVB^{+}23]. Size [KA22, CSB23, LDD^{+}22, MHA^{+}24, QWR^{+}20]. size-based [QWR^{+}20]. Size-efficient [KA22]. sizing [MPS21]. skeletal [HZLH21]. skeleton [DHC23]. skeleton-based [DHC23]. skew [HLW^{+}23b]. skew-based [HLW^{+}23b]. skewed [SZGB24, WXX^{+}24]. skewness [TM20]. sky [DLGW^{+}20]. SkyFlow [RNRA23]. skyline [CKV22, LDDL21, RNRA23]. SkySwapping [PCI^{+}24]. SLA [LZL^{+}24a, MISB22, SS22]. SLA-based [LZL^{+}24a]. Sleep [CHW^{+}20]. slice [PCC21, WLL24]. slices [TC23]. slicing [ADITS20, BÖE24, BBB22, CVdRA^{+}20, HYC^{+}23, MGC23, SCGVP20, YFL^{+}24]. sliding [uRBIBC20]. Slime [LCW^{+}20]. slippery [ZPLQ20]. Slitheen [WW20]. SLO [AKE22]. SLO-aware [AKE22]. Sloot [Fae21]. slotted [ABAD22, PZHD20]. Slow [ZLZ^{+}20a, CL21, YNVRPD23]. Slow-movement [ZLZ^{+}20a]. slow-rate [YNVRPD23]. Slowing [AHMW23]. SLRTA [TDM^{+}22]. small [HXWX23, JLC^{+}20, JAC^{+}23, WMU^{+}23, YWG^{+}19, YWG^{+}20b]. Smart [ALR^{+}20, ADAR22, BKV^{+}20, CFK^{+}20, CCW^{+}20c, CMA^{+}22, CHS^{+}24, DAAW20, DRD20, DP19, DP20c, DP21a, DP21b, LKJN^{+}20, LCLW24, SMBB^{+}24, SZM22, TBG^{+}20, VKP22, WWS20, YJB^{+}21, ZZZ^{+}22, ABMM18, ABMM22, AAA20, AAT^{+}24, APC^{+}20, AMBGS21, BLH^{+}24, BAK22, BDG23, CPS^{+}23, CGFC20, CVdRA^{+}20, CG21, CLV24, CCW^{+}20a, CFD^{+}20, CdO20, DCC22, DPLV23, DBC24, DATAA20, DGT24, DML23, uHA20, FBL^{+}20, FCGPSG^{+}21, FMM^{+}20, FFM^{+}20, GVCUGF20, GNA^{+}21, GJC^{+}20, GSG^{+}23, HSS20, HBF24, IB20, IHA^{+}20, JAAAZB20, JKS20b, JKS20c, KMK^{+}23, Kon21, KGO^{+}20, KAJ^{+}24, LMO^{+}22, LCL^{+}20, LDLS20, LQS^{+}20, Li20, LLFQ21, LJL^{+}21, LLW^{+}22a, LZZ^{+}23, LCZB21, LXZ^{+}20, LBY^{+}20, LGYC20, LGL^{+}20a, LTB^{+}22, LGW22, LLLS24, LGKA21, LCLW21, MDT^{+}20, MAC^{+}21, MKK^{+}24, MKK^{+}20, OLP23, OMSL20, OOZ^{+}23, OCSCB22, PJBB20, PSAL20, PAS^{+}20, PZHD20, PBC^{+}22, QNM24, QG20, RSR^{+}24, RSBM20, RAS^{+}20]. smart [RPP^{+}20, RLQ^{+}21, RPŠŠ23, SKB20, SKS22, SNM^{+}20, SVN^{+}20a, STS^{+}20, SLS^{+}20, SKH20, SKTP24, SD22, SKX^{+}20, TCMV20, TLMP20, TDL^{+}21, TAM21, VPSC^{+}23, WHZ^{+}20, WZT^{+}20, XWLC20, YC22, YWG^{+}20a, YHW^{+}20, YhSL^{+}22, YWH^{+}21, Yu21, Zha20, ZTC20, ZWH^{+}20, ZLF^{+}23a, ZZP^{+}23, ZXD^{+}20, ZZZ^{+}21a, BCM20, LMO^{+}22, MTT^{+}23, PGCB23, RGDMMR^{+}23, SK20a]. Smart-3DM [LMO^{+}22]. Smart-troubleshooting [CFK^{+}20]. SmartAuction [ZLF^{+}23b]. smarter [BOL^{+}20, FGG^{+}21]. Smartgrid [OPLB24]. Smartgrid-based [OPLB24]. smartphone [HOV20]. smartphones [Gur21b]. SMDP [dFCC23]. smear [ZWH^{+}21b]. smell [Ben23b, Ben23a]. smishing [MS20, MS20]. smooth [ZYF^{+}22]. smoothening [ZH20]. smoothing [HXWX23, MDL^{+}23]. SMR [CCL^{+}21]. SMR-based [CCL^{+}21]. SMS [MS20, RSB20]. SMusket [EGDT20]. SneakLeak [BHL^{+}20]. SNMP [GdOAO20]. SOA [KHH21]. Social [AJJ^{+}21, AMR^{+}20, BAGRB^{+}20, GK21, HPP20, OMSL20, RCJZ20, SACN^{+}21, SWW^{+}20, YLKK20, AGdS^{+}21, ARIB22, AYHA20, AOF21, ALS21a, AESI^{+}21, ACM^{+}21, BCM20, BPLFRL20, BR20, CLC21a, CLC21b, DFG^{+}21, FIABC^{+}20, GSMF20, GMI22, HMLS20, JYSH23, KK22, LY21, LVNCC21, LGT^{+}20, MJB22, MJZC21, PSMF21, PPG^{+}20, dAMVULM20, PP20, PPM24, RAA^{+}20, RPdVR20, SGL^{+}20a, SLA^{+}23, UCR21, VMCM^{+}20, WGLH20, WCY^{+}20, WSXL21, WM21, XGY^{+}23, XLL20a, YVW^{+}20, ZYL^{+}20, ZZZ^{+}22, ZDZ21, HPY20, YVW^{+}20]. Social-viewport [YLKK20]. socialization [CPPY24]. socially [BEL20]. socio [FGP20]. socio-technical [FGP20]. sociome [PRPPFRL20]. Socioscope [AAR^{+}20]. Soft [MRD^{+}20, HSR^{+}22, LZZX20, SSA^{+}23, WYG^{+}20, WWY^{+}24]. soft-deadline [HSR^{+}22]. soft-error [WYG^{+}20]. Software [BOD^{+}24, HAB^{+}20, HYL^{+}20, HRY^{+}21, LZS^{+}22, MGM^{+}20, ZTP20, ASASA^{+}20, BDK^{+}20, CWM^{+}20, FWX23, GBC^{+}24, GZF^{+}20b, HZZ^{+}20, HLH^{+}20, JAAAZB20, LWNH22, LZL^{+}23, zLsZjX20, MMC22, Man20, MNA^{+}23, RMA^{+}20, SW22, SMS22, TWI20, WYWS22, WD24, YZJ^{+}20, ZHX^{+}20, ZWZ^{+}21, FD21]. Software-Defined [HYL^{+}20, HRY^{+}21, MGM^{+}20, HAB^{+}20, LZS^{+}22, ZTP20, GZF^{+}20b, HZZ^{+}20, LWNH22, MNA^{+}23, SMS22, YZJ^{+}20, ZHX^{+}20]. software-hardware [LZL^{+}23]. Softwarized [VSPM21]. solar [ZLXH20]. solar-powered [ZLXH20]. Solid [FAA^{+}23, HBF24]. solution [AYHA20, DSC24, FMM^{+}20, IHA^{+}20, LCO^{+}23, LGM^{+}20, MKK^{+}24, MCV23, OPOG23, RWG21, SCX^{+}24, TM20, TSKK23, ZAH^{+}20, ZYY^{+}23]. Solutions [WPPA22b, WPPA22a, ABT20, FD21, PAS^{+}20, PPM24, SD24, SDVC22]. solve [AÇP22, LCC^{+}24b, MECRFD20]. solver [GBP23, MMAH22, XWR24]. solvers [MMPV22, MNSL22]. Solving [LEXH20, BN21, MZL^{+}22, SCR20, SCK^{+}22]. Sonnet [MZX^{+}24]. Sophisticated [BeKTK^{+}20]. Sort [WGS24, DZB23]. Sort-then-insert [WGS24]. sorting [CCL^{+}21, CCL23, KSSR20]. sound [YXS23a]. source [LHLC23, LGW22, OMPSPL20, PGHS20, SCZ^{+}20, WHZ^{+}20, WWS20, YXS23a, GXS22]. sources [DML20, GHW^{+}20, SGP^{+}20a]. SP [LYFZ20]. SP-MIOV [LYFZ20]. space [BAİP24, BK20, BEON24, CPM^{+}23, CCL^{+}22, CL20b, HS21, LLC^{+}22, LSL^{+}20, PCCX21, PDFV21, RFd20, THA^{+}24, WGW^{+}20, WCW^{+}21, WML^{+}21, WGS24, ZZG^{+}24]. space-time [WGW^{+}20]. spaces [ZC22]. SPAE [LGJ^{+}23]. Spain [MMC^{+}23]. Spam [RSB20, MK20, MK21, YPX^{+}20, ZZPK21]. spammer [GTG^{+}21]. Spanish [GDCGCPVG21, GDCPVG22, OCMJFB^{+}23, dAMVULM20]. spanning [PPA^{+}24]. spark [TA23, AYHA20, EGDT20, GMF^{+}20, LPSV22, WDG20, WGW^{+}20, ZHL24]. Spark-based [EGDT20, GMF^{+}20]. SPARQL [YGR21]. Sparse [LS23a, DHC23, JQZ^{+}22, KA22, LZ21a, MMPV22, MNSL22, PK22, TJG^{+}20, WLZ^{+}20, Wu22]. sparsification [WGS24, ZWM^{+}23]. SPARTAN [SGP^{+}20a]. Spatial [AÖ24, NN21, ASAM20, BYW^{+}21, FLH^{+}24, GPWL20, HZX^{+}24, HQLH20, JYSH20, JA20, JWZ^{+}22, KWL^{+}23, LSB21, Wan21, XWLC20, YLD^{+}23, YGP^{+}24]. Spatial/temporal [NN21]. Spatial/temporal-level [NN21]. SpatialHadoop [GGCIV20]. spatio [DWZ20, EJP22, HSvB20, KHL20, RHWY23, SGP^{+}20a, WZZD23, YWH^{+}21, YJQ^{+}23, ZWW^{+}23, ZG23, ZZ24]. spatio-temporal [DWZ20, EJP22, HSvB20, KHL20, RHWY23, SGP^{+}20a, WZZD23, YWH^{+}21, YJQ^{+}23, ZWW^{+}23, ZG23, ZZ24]. spatiotemporal [CPT^{+}20, DLdAR23, GHW^{+}20, JJZ^{+}23, MLZ^{+}23b, NLO^{+}20, SHW24, WGW^{+}20, YLC23, ZYY^{+}23]. SPEC [ZZT^{+}22]. Special [AFMG^{+}22, ACY20, ACDY21, AMNZ20, AKPT20, AHWB20, BBSB21, BDT21, BDF^{+}22, BDFR22, DPG20, ETH20, GZPZ20, MM21b, MBJ^{+}20, TDC^{+}20, VKP22, VPBE22, WPPA22b, WPPA22a, YMS20a, YMS20b, YDK20, ZA20, ZTP20, dSGST21, AMB^{+}21, Ano24s, ICW21, LLWJ24, LWHW22, PPM24, WD24, ZWH21a]. specialization [bHFF^{+}21, RGP^{+}22]. specialized [DGK20, HJI24]. specific [Deh20, GLF24, HAB^{+}20, LP21b]. Specification [DQBS20, zLsZjX20, RPF21, RHJ20]. SpecMiner [KHHV21]. spectral [GHEB^{+}18, GHEB^{+}23, XW21]. Spectrum [KAA^{+}21, XLCB20, ASA^{+}20, EGD24, Kad20, LZJ^{+}20, ZHGX20]. speculative [SOT24]. speech [AMZZ23, MCV23]. speed [LZC^{+}23a, UCO20]. speeding [GBM20]. speeding-up [GBM20]. speeds [VDSB22]. sphere [CIJM20]. spinal [WWP19, WWP20]. spine [WWP19, WWP20]. Split [ZCL24, QCWY23]. splitting [HZS^{+}23, PZHD20]. SPM [LDD^{+}22]. SpMV [BHL^{+}21]. sport [YWS21]. Sports [LZ21b, FS21, Liu21]. spot [ZA22]. Spotlight [AYB^{+}22]. spots [ZT22b]. spread [MMC^{+}23]. spreading [XZD^{+}21]. SPsync [ZYY^{+}23]. squares [CFL^{+}20]. squatting [ZWL21]. Squeeze [QNHB22]. SRF [WLX^{+}24]. SSAR [ZWW^{+}23]. SSAR-GNN [ZWW^{+}23]. SSDP [GdOAO20]. SSDs [KHES21]. SSH [GPRM21]. stability [ZWL21]. Stable [HOMD21]. stack [YW21, BPLFRL20]. stacked [KLA22]. Stackelberg [CLY^{+}20]. stacking [ZLLD21]. stacks [MAB^{+}20]. stage [ASA^{+}20, BÖ20a, CKL20, GKB^{+}20, LHY^{+}20a, PSC^{+}21, SCR20, TBO20, XLS^{+}21, ZCLL22, ZNX23, ZLL^{+}23, ZHP^{+}21]. stages [KW20, MJZC21, ZCWC20]. staging [CHW^{+}20, WDS^{+}23]. staging-based [WDS^{+}23]. staleness [CMM21]. staleness-aware [CMM21]. standard [WGL^{+}24, ZLP^{+}22]. STAR [DWZ20]. stars [DAA^{+}21]. start [FSP^{+}24, LQG^{+}23, Ngu24]. starting [BR20]. State [HHH22, LZZX20, SGP^{+}20b, WXD^{+}23, BAİP24, IDM^{+}20, KTIB22, LQW^{+}20, MZL^{+}22, MEL^{+}23, SP22, TPF^{+}20, VP20, WFLL22, WZW^{+}23, YZZ^{+}23, ZWZ^{+}21]. state-aware [LQW^{+}20]. State-of-the-art [HHH22, SGP^{+}20b, WXD^{+}23, MZL^{+}22, MEL^{+}23]. state-of-the-arts [VP20]. stateful [SZGB24]. Statement [GVŠ22]. Statement-to-Item [GVŠ22]. states [AÇP22, RSL24]. Static [ATT^{+}20, bHFF^{+}21, KYY^{+}20, LLY^{+}20, MYM^{+}21, TGJ^{+}20]. station [SGDG23]. stationary [GHEB^{+}18, GHEB^{+}23, ICBB20, ZWL20]. Statistical [ZMZ^{+}19, ZMZ^{+}20, JZZD21, KTC23, MKC^{+}21, NHTH20, SHT^{+}21]. statistics [EL21, LZZ^{+}23]. status [BMD^{+}21, TLX^{+}23]. Stealthy [WLL22, HLL^{+}20, ZZXH20]. steering [SSC^{+}20]. steganography [Pan20, YBX^{+}23]. stencil [CKW21, LWZ^{+}23b]. stent [ZMZ^{+}19, ZMZ^{+}20]. stents [HZL^{+}21]. step [ACF^{+}21, BYW^{+}21, KW20]. steps [KW20]. STF [ZG23]. STF-Net [ZG23]. STGNN [JWZ^{+}22]. stimulated [Dho20]. Stochastic [CCL^{+}20, HGY^{+}22, KIJ^{+}24, LCW^{+}20, WLY23]. stock [VGL23, ZL21]. stone [HZX^{+}19, HZX^{+}20]. stop [VCK^{+}20]. Stopping [AAP21]. storage [BWX20, CCL^{+}21, CTFW22, FLG^{+}20, JGL^{+}20, JKS20c, LHL20, LYFZ20, LLS24, LZS^{+}24, LZ20b, LPL22, Liu24, LZL^{+}24b, MAB^{+}20, PBL^{+}23, PK22, PPGS20, QLHLB23, RLQ^{+}21, RHWY23, RRDSAML23, SCP^{+}21, SKTP24, TLC^{+}20, YSL^{+}22, YLG^{+}24, ZZXH20]. store [CHJK22, FPL24]. Stores [TQC20, BQC22, dlVGSB^{+}20]. Storm [CWB^{+}20, FZ20]. Storm-based [CWB^{+}20]. STPCyber [NHY20]. STPD [CCL^{+}22]. strategies [LSS^{+}22, MJW^{+}22, dSOFC^{+}23, QL22, RMBMT21, SAD24, THB23]. Strategy [LDLS22, ABMESM18, ABMESM22, AAG22, AMBD^{+}20, BJ22, CMX^{+}20, CDX^{+}23, CKZ^{+}22, DLR23, GBH^{+}23, GZXH24, GMA^{+}22, GLWP20, GlRpG20, HXL^{+}23, bHFF^{+}21, KIJ^{+}24, LY23, LCL22, LCC^{+}24a, LCZB21, LQYL21, LGL^{+}20b, MGW23, PWH^{+}22, QGH^{+}22, RZA21, SZGB24, TBH23, WC22a, WML^{+}21, XZD^{+}21, YWH^{+}23, YSL^{+}22, YCS^{+}20, ZLZ23a, ZNZ^{+}23, ZLL^{+}23]. Stream [KWL^{+}23, VGM24, AdAHK20, BGR20, BVFGSF20, BGMK22, CWB^{+}20, CHL23, EMHE21, FBTJ23, GGK20, GPGG23, JHK20, JJZ^{+}23, KMK^{+}23, LFYH22, MSG^{+}20, MLZ^{+}22, RBA^{+}22, SGL^{+}20b, SCW^{+}22, SGLB22, SZGB24, UUH^{+}22, VMV20, WKW^{+}22, XKK20, uRLW^{+}21, FBTJ23, SHY^{+}21]. stream-based [LFYH22]. Streaming [SCC20, BGNBH^{+}20, DWL^{+}23, FCGPSG^{+}21, MSV^{+}20, MAA22, RCP24, RNRA23, SGP^{+}20a, SHY^{+}21, TWI20, TA23, XWW^{+}20, YLKK20, Yu21, ZBTV^{+}20]. streams [CIJM20, LLP^{+}20a, ST20a, SDV^{+}21, SGL^{+}20b, WFL^{+}20, ZJW^{+}20]. Street [LP23]. stricken [WZT^{+}20]. Strict [ADMG20, LLF^{+}23]. strictly [KA22]. string [DZB23]. stroke [LGC^{+}21]. strong [SCX21]. structural [ARB20, BGR20, JLW^{+}23, LQNW20, NZY^{+}23, VMCM^{+}20, YDL^{+}20, ZZZ^{+}23]. Structure [MHH^{+}20, MNSL22, HOV20, ORPPG20, SVN20b, TM20, WSXL21, WXZX23, WGL^{+}24, ZSZX24, ZWL22]. Structured [FGP23, WC23, KA22, KGO^{+}20, LFM^{+}22]. structures [LYGF21, XWLC20, XWW^{+}20]. structuring [AGdS^{+}21]. STS [LHY^{+}20a]. STT [GA22]. STT-RAM [GA22]. Student [KZF21, GZL^{+}22, PDFV21, SG20]. students [SP22, WZT^{+}20]. studies [CHC^{+}20]. study [AIM23, BZG23, BPLFRL20, CF24, CHS22, CHKJ20, FMN^{+}20, GDCGVG20, GDCPVG22, GdOAO20, KPL22, Kon21, LKJN^{+}20, LHLC23, MMC22, MV21, NKG23, OLLP24, OCSCB22, PMMG^{+}20, RYL20, Yan21, YGE21]. style [LGC^{+}21, Wei21, ZDC22]. style-oriented [LGC^{+}21]. stylometric [AMM^{+}20]. Sub [WZZD23, HBH21, KLA22, NTY^{+}21]. sub-classification [KLA22]. Sub-Entity [WZZD23]. sub-Kmeans [NTY^{+}21]. subdivision [GZF^{+}20a]. subgraph [CYG22]. subjective [WDG20]. SubMCS [ACBT23]. submission [MDP24]. subscribe [ACBT23, LPSV22]. subscription [ZLS^{+}20]. Substitution [SP24]. subtalar [Bo19, Bo20b]. success [KZF21]. successive [LH20]. suggested [UADD21]. suicidal [MJB22]. suitable [DNNG21]. suite [HBK20, LVLBB^{+}24]. suites [RAS^{+}22]. summaries [MSKG21]. summarization [CZ20]. Sunway [RYL20, XLX^{+}21]. super [PBSS24, QZZH21, WM21, ZHP^{+}21]. super-influencer [WM21]. super-resolution [PBSS24, QZZH21, ZHP^{+}21]. Superposition [YZL^{+}20, LYGF21]. supervisable [XLH^{+}24]. supervised [CdD20, HZL^{+}21, JMA^{+}21, KCB20, LCY^{+}23a, LLZ20, RLCB22, VGM24, WWH^{+}21, ZCWC20, ZWW^{+}23]. supervision [GAP24, ZGW^{+}23b]. supplier [ABMMC18, ABMMC22]. supply [ABMM18, AB19, ABGMC19, AB21, ABGMC21, ABMM22, AHSH22, BFM23, LBJ^{+}18, LBJ^{+}24, LHW^{+}23, SCE23, YLF^{+}23]. Support [HCB^{+}20, LXL^{+}23, MM21b, YMY21, AM21, BCSS20, DPPGCCA23, JAS^{+}20, Kol22, KA24, LLW^{+}20, PBM^{+}22, PSvL^{+}20, RZIX20, SOT24, VPA20, YPEK23, dSGST21]. supported [BAR21, WXZZ22, XZH^{+}23, ZZZ^{+}21a]. Supporting [FHGF20, WZT^{+}20, BKHD20, KRA21, LHC21, SGDK^{+}21, XCL^{+}20, ZFMB20, ZSL^{+}23a, ZBS23]. suppress [ZWX^{+}23]. surface [DLHD22, WX23, XYL^{+}20]. surface-aided [WX23]. surfaces [TRB^{+}23a]. surgery [WWP19, WWP20, ZYX^{+}23]. surplus [ZAH^{+}20]. surrogate [AMT^{+}21]. surveillance [CSAT24, DS23, FCGPSG^{+}21, RSBM20, UUH^{+}22, ZZXH20]. Survey [ASH^{+}23, DGT24, AUJW22, ABL23, CCDR22, DPN^{+}22, BAM^{+}24, FD21, GAdFGMA21, HAH^{+}23, ISUC22, KHH21, LJC^{+}20, MZL^{+}22, MPP^{+}21, QCG^{+}24, RMI22, SD24, SAAEK22, SD22, VI21, WLC^{+}20b, WXS^{+}22, XCZ^{+}23, YD21, ZYL^{+}20, ZHL^{+}23, ZG24]. survival [AIM23]. Susceptible [TTZ^{+}21]. sustainability [MV21]. sustainable [GZF^{+}20a, HBF24, Kon21, MR23b, MHA^{+}24, PPM24, TLN23, YBC^{+}20]. SVC [Yu21]. SVC-based [Yu21]. SVD [TKS^{+}23]. SVD-based [TKS^{+}23]. SVM [JKS20b, ZY21]. swallowing [MSK^{+}21]. swapping [PCI^{+}24]. SwapQt [JA20]. Swarm [KSDR21, ZWL22, AJPM20, CZZ^{+}23a, GHEB^{+}18, GHEB^{+}23, GZXH24, HY21, LZHS24, ZLZ^{+}20a, ZZ21a, ZLZ21]. Swarms [CIJM20, CP22]. switch [GZF^{+}20b, LWHW24, LZCGMVV20]. switch-off [LZCGMVV20]. switchable [Liu23]. switching [OPOG23, RMBMT21]. SX [GMFC23]. SX-Aurora [GMFC23]. Sybil [OdVP20]. Sybil-resistant [OdVP20]. SYCL [RNRA23]. Symbiotic [ZLG^{+}24]. Symbolic [STH^{+}20]. synchronisation [BMS20]. synchronization [CdD20, YZC^{+}20, ZYY^{+}23]. Synchronizing [MSLP24]. synchronous [ACN^{+}21, HY21]. syndrome [Bo19, Bo20b]. syndromes [LBY^{+}20]. synergy [DKG^{+}22]. syntactic [YLGZ21]. Synthesis [wZcZN^{+}19, wZcZN^{+}20, ZZT^{+}22]. SyRoC [ZLG^{+}24]. Syscall [ZLPZ21]. Syst [AB19, AB21, ABGMC21, ABM21, ABMESM22, ABMM22, ABMMC22, ARA^{+}23, Bo20b, DP20c, DP21a, DP21b, FGB21a, GHEB^{+}23, HZX^{+}20, JLC^{+}20, KF23, LYYG20b, SME^{+}21, WWP20, WCWC20, YWG^{+}20b, YTQ20a, YTQ20b, ZMZ^{+}20, wZcZN^{+}20]. System [APC^{+}20, BQI^{+}20, CL20b, DP20c, DP21a, DP21b, GZPZ20, KCR20, LLZ^{+}22, MR23b, NKG23, NHTH20, PFP^{+}22, RZA21, Ski20, SZM22, TBG^{+}20, VSPM21, WZB^{+}20, XLLL20, YMY21, AACJ23, AEN^{+}23, ASBT20, ATT^{+}20, ACM^{+}21, AJPM20, BCT^{+}21, BBD^{+}21, BSH^{+}21, BCB^{+}20, CWM21, CMF^{+}21, CXHS20, CPT^{+}20, CXWY21, CHL23, CMA^{+}22, CHJ^{+}20, DATAA20, DDM21, DP19, DWL^{+}23, DHD20, ERK^{+}24, EEN^{+}24, ELS20, FWP21, FGG^{+}23, GAP24, GLP^{+}24, HYC^{+}21, HLZ^{+}22, HBF24, HHW^{+}22, HRM20, bHFF^{+}21, JKS20a, JZZ^{+}23, JXYC24, JZL^{+}24, KAK20, KBG20, KKL^{+}24, KZG^{+}22, KSS^{+}21, LBJ^{+}18, LBJ^{+}24, LKS^{+}21, LLW^{+}20, LHL20, LQS^{+}20, Li20, LHW^{+}23, LPSV22, dTGC20, LWY^{+}24, MGB24, MDP24, MMKS22, MRS^{+}22, MR23a, MXS22, MKC^{+}21, NSR^{+}23, OMSL20, OOB^{+}21, PLL^{+}24, PBL^{+}23, PP20, PR20, QPL^{+}22b, RPP^{+}20, SKB20, SCP^{+}21, SZM^{+}21, SMC^{+}20, SK23, SRM^{+}23, SWL^{+}20, TKS^{+}23, TSM24, TTTH20, TK24a, TLS^{+}21]. system [UYH21, WHA^{+}20, WLD^{+}20a, WYS20, WLN^{+}21, WGW^{+}21, WGGB24, XCS^{+}22, XZ20, XWW^{+}20, YXLB20, YLL22, YXS23a, YGP^{+}24, YXYH20, YK20a, YZS^{+}21, YWH^{+}21, YYL22, YYKK20, ZYL^{+}20, ZWC^{+}22, ZY20, ZHD^{+}20, ZZJC21, ZCF21, ZLPZ21, ZT22a, ZZG^{+}22, ZLS23, ZGW^{+}23a, ZZZ21b, ZWY^{+}21, ZA22, ADP^{+}22, HTAY21, LWHW24, XZH^{+}23]. system-level [BCB^{+}20, ZYL^{+}20]. Systematic [BHH22, SAAEK22, BDK^{+}20, CRdRR^{+}22, ICGGAR22, KPGD24, NAK^{+}22, SDVC22, STK23]. Systems [AABKB22, APC^{+}20, BAMR20, BDF^{+}22, DAAW20, IDM^{+}20, Kho21a, KVCY20, LBJ^{+}24, LGKA21, LCLW24, MTHA24a, SWW^{+}20, SRM^{+}23, VKP22, WSD^{+}22, YMS20a, ZGY20, ABMM18, ABMM22, AMM^{+}22, AMB^{+}21, AHH20, ASASA^{+}20, ACI^{+}23, ACG^{+}20b, AL20, BEM^{+}24, BMS20, BSOK^{+}20, BGNM20, BFM23, BBB22, BAR21, BCSS20, BDFR22, CdST^{+}20, CDG^{+}20, Cha20, CLQS20, CCL^{+}21, CZH^{+}24, CSAT24, CIB^{+}20, CCDR22, CKW21, CF21, DP20a, DBD^{+}23, DDM21, DSRG22, DQBS20, DKD22, EEN^{+}24, EMHE21, EGDT20, FWY^{+}22, FGP20, GAA^{+}21, GL20, GAdFGMA21, GMP^{+}20a, GZB^{+}22, GMT23, GDEBC20, GK21, GLWP20, Gur21b, HTXW21, HLL^{+}20, HLK^{+}23, HGY^{+}22, HYC^{+}23, HAH^{+}23, ICBB20, ICW21, JSV21, JRW^{+}20, JAAAZB20, JHK20, KZB^{+}23, KYPJ20, KHES21, KNV20, LHC^{+}20, LLC^{+}22, LMNC22, LCB^{+}20, LBGL20, LJC^{+}20, LZW^{+}22, LZL^{+}23, LLG^{+}20, LDD^{+}22, LGM^{+}20, LL20, LCLA21, MABK24, MWK^{+}21]. systems [MEC^{+}20, MBC^{+}23, NKB^{+}20, NK20, OPLB24, PJJ^{+}22, PSMF21, PCK20, PP24, PNL^{+}21, PCG^{+}20, PK22, QMCX20, RRAB24, RFP22, RAS^{+}22, RCdF^{+}21, RHJ20, SD20, SDGCB^{+}20, SVN^{+}20a, SMY20, SW20, SK21a, SSDC22, SACW23, SCW^{+}22, SGLB22, SLA^{+}23, TCMV20, TWL23, TMT22, TAM21, TGAP20, VMV20, VZDS24, VEH^{+}23, WGLH20, WLD^{+}20a, WFLC22, WZS^{+}23, WYZ^{+}24, WMCH22, WWS23b, WZH^{+}22, WD24, XZJ^{+}20, XLG^{+}23, XZK^{+}20, YNN^{+}20, YZC^{+}20, YLS21, YLH^{+}23, YGD^{+}21, YMS20b, YSL^{+}22, YYZ^{+}24, ZFMB20, ZYL^{+}20, ZJL^{+}22, ZWZ^{+}23, ZLZ23a, ZSL^{+}23a, ZHL^{+}23, ZZF^{+}24, ZSZ^{+}24, ZLS^{+}20, ZXL^{+}20, ZXZ^{+}23, Fae21, Tau23, Tau24]. systolic [TWY^{+}23, YSL^{+}24].

T [ZLS^{+}20]. T-CAM [ZLS^{+}20]. Table [CDRS20, ABA24, LZC^{+}23a]. tactical [MV21]. Tactile [YAZ^{+}20]. TAD [DBD^{+}23]. TagSNP [WLW^{+}21]. TagSNP-set [WLW^{+}21]. tail [CWYG23, YPL24]. take [ZDC22]. take-off [ZDC22]. tallying [YYN^{+}20]. Taming [GMT23, HLH^{+}20]. tampering [VFOV20]. Tangle [MRMB24, SM20]. tardy [BPUW24]. Target [ASM^{+}22, JLT^{+}21, LZH^{+}20, Pła24]. targeting [wZcZN^{+}19, wZcZN^{+}20]. Task [HHH22, LHH^{+}21, LMZL24, PRF20, WSD^{+}22, YXYH20, AOKÖ24, ATZP21, AAP21, AM22, AKA20, BLT^{+}24, BQC23, CZZ^{+}23a, CCL23, CYH20, DFZ^{+}20, DK24, EELB21, GCT^{+}20, GEN20, GLF24, GMT23, HXL^{+}23, HCG^{+}23, HJI24, Hu20, HX21, JXYC24, JMHB24, KF22, KF23, KDX^{+}24, KSMT24, LHD^{+}20, LSH^{+}20, LZW^{+}22, LWL23b, LZL^{+}24a, LZ22, LGL^{+}20b, LFHS23, MMP^{+}23, MMFAB23, MMMZ20, MWL^{+}20, MNSL22, NLS23, OPOG23, QCW^{+}24, RCLEB20, RBW20, ŠTI24, SHB22, SCX21, SLA^{+}23, TWM^{+}23, TSB20, WXX^{+}24, XGS^{+}20, XW23, XCSF20, YSL^{+}24, YJLC20, ZWZ^{+}24, ZTC20, ZLP^{+}22, ZYW24]. task-based [BQC23, EELB21, JMHB24, MMFAB23, MNSL22, NLS23, RCLEB20]. task-specific [GLF24]. Tasks [LXC^{+}24, AK20, AOSA20a, AM21, AÇP22, BLMT20, CHG^{+}20, DGK20, EAA21, GAA^{+}21, GBH^{+}23, HSR^{+}22, JHB22, KGO^{+}20, MTA^{+}22, QSZ^{+}24, SXW^{+}22, ZLZ^{+}20a]. taxonomies [SD24, VI21]. Taxonomy [AAA20, MGB24, MMC22, MKK^{+}20, TRB^{+}23b, KPL22, MBB^{+}20, NMRK21]. Taylor [VP20]. TBAN [ZL23]. TCAM [YLSL22a]. TChain [ZLL^{+}24b]. TCLBS [PR20]. TCP [TPN^{+}21]. TCSLP [WHZ^{+}20]. TD [HSGY20]. TD-Root [HSGY20]. Teaching [RD23, ZN21]. TEBDS [XZH^{+}23]. Technical [LCZB21, FGP20]. Technique [OÖA22, ArMA^{+}21, CCTZ22, CATD^{+}24, EAA21, LZ21b, LZ21a, MAM^{+}24, NNH^{+}20, OHÁV20, PSAL20, PK22, SYG^{+}20, ST20b, THVL24, WCHA20, Xu21]. techniques [AMM^{+}20, AKJJ20, CCP^{+}20, CMGI^{+}23, DOR^{+}21, MMC22, QCG^{+}24, RDR^{+}24, SSA^{+}23, VKP22, VGL23]. Technologies [TDC^{+}20, YAZ^{+}20, BGNBH^{+}20, BAGRB^{+}20, Dao23, SMY20, WFLC22]. Technology [AHWB20, AAM^{+}24, DLdAR23, FS21, BAM^{+}24, GKA^{+}21, LBJ^{+}18, LBJ^{+}24, LZH^{+}20, SQ22, SRA^{+}22, ZXW^{+}20, ZY21]. TEEFuzzer [DFZ^{+}23]. teens [GMMR24]. telecommunications [HCL^{+}22]. temperature [AOSA20b, AMT^{+}21, LLT22]. Temporal [JWYÍ21, MMR^{+}23b, NZY^{+}23, ZZT^{+}22, BYW^{+}21, DBC24, DWZ20, EJP22, HZX^{+}24, HSvB20, JWZ^{+}22, KHL20, LWW^{+}20, LZP23, LCY^{+}23b, LSB21, RHWY23, SGP^{+}20a, SCC20, WZZD23, Wan21, YLD^{+}23, YWH^{+}21, YJQ^{+}23, ZWW^{+}23, ZG23, ZZ24]. temporal-level [NN21]. Temporal-structural [NZY^{+}23]. Tenant [PCC21, CMX^{+}20, MLX23, MVLJ21, MKB23, SNMWC21, WMNV20]. Tenant-defined [PCC21]. Tennis [NN21]. Tennis-based [NN21]. tensor [QNHB22]. Tera [PDA^{+}20]. Tera-scale [PDA^{+}20]. Term [ACF^{+}21, XCL^{+}20, CHW^{+}20, CLQS20, DWZ20, MZLT21, NED^{+}20, WLL22, XWZM24, YFQ^{+}22]. terminal [HLL^{+}24, ZYY^{+}23]. ternarized [RSFB23]. terrestrial [YCS^{+}20]. territories [DC21]. terroristic [AJJ^{+}21]. test [OTMN23, YAZ^{+}20]. testbed [BQI^{+}20, FGP20, SUKN22, UADD21]. testbeds [GFPB23]. testing [DP20a, LCFM20, TBH23, ZZQ21]. text [CZ20, GVŠ22, JZZD21, LC20, XFJ^{+}20, YLGZ21, YJQ^{+}23, YLG^{+}24]. texts [LZW21]. Textual [JMA^{+}21, GFZ21, LF21]. texture [YXL^{+}21]. their [CHG^{+}20, HMY^{+}23, ORPPG20, ZYL^{+}20]. theoretic [KNV20, MXL^{+}20, MZX^{+}24]. theoretical [HMO^{+}20, HOMD21, RZH21, TBO20]. Theory [AAP21, CS24b, GFM^{+}20, LZC21, UYH21, Zha21, CL20b, DG21, LYC^{+}22, LZZ^{+}20, RTD24, WMD^{+}20, YJLC20]. therapeutic [PRPPFRL20]. therapy [HRM20, ZMZ^{+}19, ZMZ^{+}20]. thermal [AÖ24, AABB24, BCT^{+}21, CLL^{+}23, MMBD20, MBD21]. thermal-aware [CLL^{+}23, MMBD20, MBD21]. thing [ZZZ^{+}22, VEH^{+}23]. Things [ABMM22, AYA^{+}23, ANA24, DP20c, DP21a, DP21b, SN23, Sun20, UUH^{+}22, WPPA22b, WPPA22a, XWG^{+}21, BKV22, Dao23, GS20, LZH^{+}20, SBF^{+}21, SKH20, AVK^{+}23, ABMM18, AND24, AALEF20, ARHT20, AUJW22, AWMM^{+}23, ADRP23, Ano24s, ADAR22, BCT^{+}21, BOL^{+}20, BCSS20, CPPY24, CRdRR^{+}22, CDP20b, DP19, DLHD22, ETH20, FZT^{+}23, GCCMK^{+}20, GWW^{+}22, Gul22, HN23, HPY20, HRX^{+}21, HH22, JZK^{+}21, JM20, KMR^{+}22, KMS20, LAS20, LLFQ21, LLW^{+}22a, LCH^{+}23, LZ20b, LDW^{+}21, LWZ^{+}23a, LMW^{+}24, Liu24, NBJ21, NAC^{+}22, NLSY20, NMR21, PAS^{+}20, RJA^{+}22, RWG21, SAD24, SGP^{+}20b, STS^{+}20, SMS22, SM20, SD22, SDKM20, SK21b, TLX^{+}23, TLL^{+}24, TDL^{+}21, TG20, TWM^{+}23, TSKK23, TK24b, UAACH21, VMV20, VSPM21, WCHA20, WYHM21, WSL^{+}23, WPPA22b, WYGP21, XGS^{+}20, XXY^{+}23, XZC^{+}24, XY20, XCZ^{+}23, YNN^{+}20, Yu21, ZCQ^{+}23, Zha20, ZWH^{+}20, ZA20, ZZZ^{+}22, ZNX23, ZZG^{+}24, ZXW^{+}20, ZCL24, Zhu20]. Things [Zhu21, ZLWH23]. Things-assisted [UUH^{+}22]. Things-based [BCT^{+}21]. Things-generated [BOL^{+}20]. Thread [MVLJ21, LLC^{+}22, NQH^{+}20, NQB^{+}23]. Thread-level [MVLJ21]. Threading [PFP^{+}22]. Threat [GCPM22, BCB^{+}20, CFC^{+}20, HZZ^{+}20, IA20, KMR^{+}22, SCÁB20, UADD21, ZBF22]. threats [HAH^{+}23, ZCWC20, VMM20b]. Three [ABMMC18, ABMMC22, TBA23, ZCS20]. three-dimensional [ZCS20]. Three-precision [TBA23]. Three-way [ABMMC18, ABMMC22]. Throughput [CWB^{+}20, CIS^{+}20, RKI^{+}23, SHST20, SHY^{+}21, YW21]. ThunderX2 [MGGG^{+}20]. thwart [NMRK21, RSQS21]. Thyme [SCP^{+}21]. tick [RZA21]. tick-less [RZA21]. tier [GEN20, LPL22, LHY^{+}20b]. TIIA [ZBF22]. Tikiri [BTF^{+}21]. tiling [KW20]. Time [ASA^{+}20, GBM20, GLP^{+}24, RKP^{+}21, TA21, VSV^{+}23, ZLS^{+}20, AdSM^{+}22, AHH20, ABAD22, AGV23, AEZ22, AYB^{+}22, BZG23, BKG^{+}20, CL20a, CPM^{+}23, CYWS24, CHL23, CYH20, CIJM20, DSRG22, EET20, dRFRB24, FHGF20, GRG20, GRN20, GZZG24, HN22, HSGX22, Hu20, HWH^{+}23b, JHB22, JWZ^{+}22, KHL20, KP22, KBTM21, KW20, KK22, LLP^{+}20a, LXZ^{+}20, LQNW20, LCO^{+}23, qLhZ20, LWH^{+}22, MMR23a, MSV^{+}20, MASRAM^{+}22, MXS22, Ngu24, dSOFC^{+}23, PBM^{+}22, PCACTMÁ24, QSZ^{+}24, RCP24, RAS^{+}22, SNMWC21, SZVVB^{+}23, SDVC22, SCP^{+}21, SOKW^{+}20, SGL^{+}20b, TDM^{+}22, TCBF24, VMV20, WGW^{+}20, WML^{+}21, WGGB24, WF21, XZJ^{+}20, XCGZ24, XZTC22, YHW^{+}20, Yan21, YLL22, ZCQ^{+}23, ZWL20, ZYX^{+}20, ZT22a, ZLS23, ZLL^{+}24b, ZZ24, ZTP20]. Time-aware [GLP^{+}24, BKG^{+}20, SCP^{+}21]. Time-based [ZLS^{+}20]. Time-constrained [RKP^{+}21, LLP^{+}20a]. time-critical [LCO^{+}23, QSZ^{+}24, ZLL^{+}24b, ZTP20]. Time-discretization [GBM20]. time-parameterized [KHL20]. time-series [HN22, MASRAM^{+}22]. time-slotted [ABAD22]. Time-spectrum [ASA^{+}20]. time-varying [LQNW20, ZYX^{+}20]. Timed [ADMG20, QHC24, YJF^{+}20]. timeout [DGY^{+}22, RWJ^{+}20, See20]. timing [DDM21]. tiny [MSR20]. Tinycubes [DLdAR23]. TinyML [TBB^{+}23]. tire [LFZJ21]. TireNet [LFZJ21]. TLS [HLW^{+}23b]. TMHD [LXC^{+}24]. tokenization [GZF^{+}23]. Tokenizer [MBF^{+}20]. tokens [LDM^{+}21]. Tolerance [PRF20, DEJ20, GCN^{+}24, HYL^{+}20, KYPJ20, LGM^{+}20]. tolerant [BVCH22, JPW20, MA22, PRF22, WZTL20, WGW^{+}21, ZLW^{+}24]. tolerating [hAS24]. tomography [SGBC^{+}20]. tone [SSMdS21]. tool [AB19, ABGMC19, AB21, ABGMC21, DBC24, EGD24, HBK20, HL24, KCP23, YW21]. toolkit [PVA^{+}20, PSH^{+}20]. Tools [CGM^{+}23, CK20, DP20b, NLSY20, SNS^{+}20, TRB^{+}23b]. toolset [PSvL^{+}20]. Top [MBM^{+}20, WFL^{+}20, CQA^{+}24, HPY20, WGS24]. Top- [MBM^{+}20, WFL^{+}20, CQA^{+}24, HPY20, WGS24]. tophi [YTQ19, YTQ20a, YTQ20b]. Topic [ZWW^{+}20b, RCP24]. Topic-based [ZWW^{+}20b]. topics [DAA^{+}21, WMCH22]. topological [IHA^{+}20, LSL^{+}20, XWLC20]. topologies [ACA^{+}23, CKW21]. topology [CDX^{+}23, LJW^{+}20, SNMWC21]. Tor [PGMP23]. total [BPUW24]. Touch [KAJ^{+}24]. tourism [ZT22b]. tourist [PPG^{+}20]. TPC [TMT22]. TPC-DS [TMT22]. trace [MJSW21, WHZ^{+}20]. Traceability [GCH^{+}22, LHW^{+}23]. traces [KHHV21]. tracing [BFM23, KCP23, MNFQ24]. Tracking [ALS21a, CWM21, AAA20, CP22, LZH^{+}20, WLLY20, WFL^{+}21, ZLS22b, dAdSM^{+}22]. tracklet [YLD^{+}23]. TrackPuzzle [LWS^{+}23b]. tractor [ZPLQ20]. tractor-semitrailer [ZPLQ20]. trade [AP20, Par22, XZK^{+}20]. trade-off [Par22]. trade-offs [AP20, XZK^{+}20]. tradeoffs [AM22, KXZW23]. Trading [MMPV22, DSPSNAHJ20, GLW^{+}20, LZ22, LGKA21, LWH^{+}22, VGL23]. Traditional [BEL20, SK21b]. Traffic [BN21, CECS20, ABOS22, BÖ20a, CLQS20, CCC^{+}23, DBSL23, DK24, FCGPSG^{+}21, GS20, HAB^{+}20, HLW^{+}23b, HZT^{+}22, HZX^{+}24, JSV21, KAA^{+}21, KCB20, LMCSE20, LHY^{+}20b, LYW23, MISS22, OIG24, QHNL21, QCWY23, RYL20, SW22, STS^{+}20, XLG^{+}23, YFQ^{+}22, YHW^{+}20, ZZLF21, ZZL^{+}22]. traffic-based [XLG^{+}23]. train [CWM21]. trained [ZY21]. training [CDC^{+}24, CHJ^{+}20, DVEE^{+}20, HXWX23, LYKK22, LCH^{+}21, LHLC23, MJW^{+}24, NTY^{+}21, RKI^{+}23, TBB^{+}23, WXZX23, ZZZ^{+}22]. Trains [TSX^{+}24]. traitor [JMA^{+}21]. traitor-based [JMA^{+}21]. traits [GDCPVG22, YRV^{+}22]. trajectories [LWCC23, LCY^{+}23a, LWS^{+}23b, PDFV21]. trajectory [GHW^{+}20, rHZmH^{+}24, PTZ^{+}20, TOM^{+}20, Tao23, WLC23, YZR23, ZHH^{+}23, ZrHhH^{+}23, ZXX23]. TrajGAN [ZHH^{+}23]. transaction [WQHX20, XSW^{+}21, YNK^{+}20]. transactions [BYR^{+}20, DSPSNAHJ20]. Transfer [DBD^{+}23, CKZ^{+}22, LFM^{+}22, MSTN21, SD20, SSS21, SK21b, MMM^{+}20]. transfer-based [CKZ^{+}22]. transferred [ZZ24]. transfers [HWQ^{+}20]. transform [GHEB^{+}18, GHEB^{+}23, SLH^{+}20, ZZZ^{+}23]. Transformation [PJLL23, CCL^{+}22, KAH^{+}23, PMMG^{+}20, WZH^{+}22]. Transformer [NRMI20, WML^{+}23, ZHD^{+}20, ŠTI24, WZW^{+}23, ZCQ^{+}23, ZLH^{+}24]. Transforming [NNPP23, NTY^{+}21]. transient [DLGW^{+}20]. transit [ZBTV^{+}20]. transiting [YWDC23]. transition [LWW^{+}20, YWDC23, ZWZ^{+}21]. transitions [AÇP22]. translating [YGR21]. Translation [LHY^{+}20a]. Transmission [BRM^{+}20, CCW^{+}20b, CFD^{+}20, CPT^{+}20, CDX^{+}23, JPMR21, LDLS22, Pan20, PZHD20, PMT22, ZLC^{+}21]. transmissions [SKX^{+}20]. transmitting [Hu20]. transparency [NNPP23]. Transparent [AFL23, PRF22, SMO^{+}24, GSSB24]. Transplantation [LGZ^{+}24]. Transport [AAH^{+}23, NJB20, BBM^{+}20, BMZdP21, RWJ^{+}20, SCA22]. Transportation [BAMR20, LGYC20, MhCEANSM20, NSJ^{+}24, THVL24]. Travel [JWZ^{+}22, JR22]. Traveler [ELS20]. traveling [DZXS21]. treatment [ASYL22, Bo19, Bo20b, SWL^{+}20, WCWC19, WCWC20]. tree [BÖ20a, LV24, MMKS22, ORPPG20, RK20, WX24]. tree-based [LV24]. trees [DCD^{+}24, JCK24]. tremor [MMP^{+}23]. trend [YYXZ23]. trends [CLC21b, CMMST20, CCDR22, DWZ20, IDM^{+}20, KSS^{+}20, MMC22, NHY20, VKP22, VI21, YMS20a]. triage [Kon21]. trials [YAZ^{+}20]. triangle [WC23, ZJW^{+}20]. Trigger [ALS^{+}21b]. Trigger-based [ALS^{+}21b]. Triggerflow [ALS^{+}21b]. trip [PPG^{+}20]. tropical [MXS22]. troubleshooting [AYB^{+}22, CFK^{+}20]. true [LGT^{+}20]. truly [Cha20]. Trust [FFB20, MAS23, NHY20, PR20, RBW20, UPK^{+}23, YMAAH22, AVK^{+}23, ADdMM20, ASBT20, ADAR22, BJW22, DQBS20, FFM^{+}20, GL20, GKA^{+}21, HBEK20, KSH^{+}21, NNPP23, RAN^{+}20, SLX^{+}24, TG20, TLS^{+}21, WZB^{+}20, ZWZ^{+}23, ZFZS23, RNA^{+}22, HKS23]. Trust-aware [RBW20, HBEK20, RNA^{+}22]. Trust-based [YMAAH22, ADAR22, DQBS20, FFM^{+}20, KSH^{+}21]. Trustable [BSB^{+}22]. TrustChain [OdVP20]. Trusted [ACY20, ACDY21, DJP^{+}24, DFZ^{+}23, HYWY22, LGW22, LLZ^{+}24b, XZZ^{+}20b, XSW^{+}21, YYN^{+}20, FLTQ20, XZH^{+}23]. TrustedMaaS [NNPP23]. trustless [MBF^{+}20, TLS^{+}21]. trustworthiness [AAB23]. trustworthy [CGFC20, HSGY20, LZF^{+}24, RAN^{+}20, SCX^{+}24]. TrustZone [ZY20]. truth [MLZ^{+}23b, YLX^{+}23]. truth-aware [YLX^{+}23]. Truthful [LZ22, ZLS^{+}22a]. TS [KMK^{+}23]. TTE [JWZ^{+}22]. Tuberculosis [BRM^{+}20]. tumor [JLC^{+}20, KSA^{+}20, XY20, YWG^{+}19, YWG^{+}20b]. tune [SK21b]. Tuning [PSH^{+}20, MPS21]. Tunnel [DZB23]. Tuoris [MFMSG20]. TurboStencil [LWZ^{+}23b]. tutorial [ISUC22, LSS^{+}22]. tutoring [LSS^{+}22]. Tweet [JKS20a, POMK20]. tweets [AJJ^{+}21, GDCGCPVG21, GDCPVG22, MM21a]. Twin [LXC^{+}24, SMT^{+}24, AVK^{+}23, CDR24, GHB^{+}24, JYP24, QZZ^{+}24, RSR^{+}24, SAM^{+}24, SKTP24, WLL24, WYJ^{+}24, NSJ^{+}24, SMBB^{+}24]. twin-assisted [JYP24, QZZ^{+}24, WLL24, WYJ^{+}24]. twin-based [SAM^{+}24]. Twin-Bridge [LXC^{+}24]. Twins [KAJ^{+}24, DGT24, DBBP24, GVCC24, HMY^{+}23, KPGD24, LLW^{+}22a, LCB^{+}23, OPLB24, PMMG^{+}20, SMS^{+}24, VZDS24, BBF^{+}24].

Twins-enabled [KAJ^{+}24]. Twitter [FAŞ^{+}20, DFG^{+}21, GSMF20, NRMI20, NLO^{+}20, PRPPFRL20, RHM20, SMU^{+}21]. Two [TBO20, ASBT20, BMBE20, BGCL20, BQI^{+}20, CCW^{+}20b, GEN20, JPW20, KMK^{+}23, LPL22, LHY^{+}20b, PSC^{+}21, PCCX21, PAM21, SCR20, ST20b, UUH^{+}22, WYZ^{+}24, ZCLL22, ZNX23]. two-dimensional [CCW^{+}20b, PCCX21]. two-phase [JPW20, PAM21, ST20b]. two-rounds [WYZ^{+}24]. two-sided [BMBE20]. Two-stage [TBO20, PSC^{+}21, SCR20, ZCLL22, ZNX23]. two-stream [UUH^{+}22]. two-tier [GEN20, LPL22, LHY^{+}20b]. two-way [ASBT20]. type [ERL^{+}20, GSMF20, LHF^{+}20, LFZJ21, LSL^{+}20, LMCSE20, TWM^{+}23, UYH21, WZH^{+}22]. type-2 [UYH21]. type-of-traffic [LMCSE20]. types [HWQ^{+}20]. typhoon [ZZD22]. typicalities [MSG^{+}20].

U [CDY^{+}20]. UAV [LYW23, HYRZ20, MISS22, WLAC20, XLG^{+}23, ZWC^{+}22]. UAV-assisted [WLAC20]. UAV-based [MISS22]. UAVs [TDL^{+}21, ZLXH20]. Ubiquitous [LLW^{+}23a, YMS20a]. ULAMA [ASAM20]. ULFM [LGM^{+}20]. ultimate [LHA20]. ultra [CGMT20, LZCGMVV20, MR23b, SGDG23]. ultra-dense [LZCGMVV20, SGDG23]. ultra-low [MR23b]. ultra-scale [CGMT20]. ultrasound [HIU^{+}22, LYH^{+}21, YTQ19, YTQ20a, YTQ20b, ZLP^{+}22]. unauthorised [FGG^{+}23]. Unbalanced [LZ21a, LYY^{+}20b, ZL22]. uncertain [LZ20a, LGL^{+}20a]. uncertainties [SMS^{+}24]. Uncertainty [MKBT24, PABBA20, AABKB22, ACG^{+}20b, KA21]. Uncertainty-aware [MKBT24, AABKB22]. uncertainty-driven [KA21]. underground [Liu23]. Understanding [BPLFRL20, CCML20, SD24, LSB21, LHTSM^{+}23, SZW^{+}23, Wan21, WFLC22, XWLC20, Xu21]. underwater [ZLS22b, ZFZS23]. UNet [QJZ^{+}20]. unextractablility [HZ20]. Unhooking [AKCP21]. Unified [MGS21, JKS20c, PP22, WLC23]. Unify [MAB^{+}20]. Union [HXWX23]. unipath [CZCH24]. UniPreCIS} [DSC24]. unit [CZ20, LZP23, XWM20]. Units [uRKI^{+}21, BKV^{+}20, KMR^{+}22]. University [LZL^{+}21, KLW^{+}21]. unknown [HLW^{+}23b, WML^{+}21]. unknowns [HMT^{+}20]. unlabeled [LWS^{+}23b]. unlinkability [PDT21]. Unlocking [DLR23]. unmanned [WLN^{+}21, LYW23]. unobtrusive [MGS21]. Unpredictable [ASDLS23]. Unraveling [PJL^{+}24]. unrelated [OPOG23]. unresectable [ZMZ^{+}19, ZMZ^{+}20]. unstructured [MSKG21, ZGW^{+}23a]. Unsupervised [MBC^{+}23, BCB^{+}20, LWL23a, LWJ^{+}21, YLD^{+}23]. untrusted [LMW^{+}24, XZZ^{+}20a, YL20b]. Unused [KCR20]. Up-weighting-based [ArMA^{+}21]. Update [BKHD20, DHA^{+}20, DLL20, LGJ^{+}23, YLSL22a]. updated [DLdAR23, WWY21]. updates [MSBAU24]. uplink [LYYG20a, LYYG20b, YK20b]. urban [BÖ20a, LCL^{+}20, LZB20, MhCEANSM20, RPdVR20, WWS20, YWH^{+}21, ZLS23]. URL [MS20]. Usability [NBB20]. usage [CGFC20, LYG^{+}24, SCX21, TK24a, ZLW^{+}22]. Use [VSV^{+}23, AAM^{+}24, BAGRB^{+}20, CdD20, DGL^{+}20, GMI22, KHHT21, POMK20, QRS^{+}21, SOT24, TRB^{+}23a, TDL^{+}21, OCMJFB^{+}23]. used [LYGF21]. User [Elg20, SGDG23, BSH^{+}24, BGCL20, CZZ^{+}23a, DML20, GSMF20, JM20, KZB^{+}23, KAF^{+}23, Kho21b, LHC^{+}20, LLC^{+}22, LMNC22, LQYL21, LLT20, MDP24, POMK20, PRD^{+}22, QRS^{+}21, RMBMT21, RPdVR20, SD20, SSMdS21, SSC^{+}20, TTZ^{+}21, TA21, VPSC^{+}23, ZIOT^{+}20, ZWW^{+}20a, dAdSM^{+}22].

User-centric [SGDG23, DML20, RMBMT21]. user-oriented [PRD^{+}22]. users [KGO^{+}20, LXZ^{+}20, LHY^{+}20b, SGDG23]. uses [MCF20]. USIM [YHC20]. Using [AÇP22, CCP^{+}20, CMGI^{+}23, DLC^{+}22, HMY^{+}23, KHB20, KCY^{+}21, MBD^{+}20, NLO^{+}20, SHH23, SHY^{+}21, XLG^{+}23, ZZPK21, AM20, ASL22, AOSA20b, ASM^{+}22, AM21, AAS^{+}20, AKA20, AESI^{+}21, ABOS22, ANA24, Ano24s, AEZ22, ACM^{+}21, AJPM20, BSH^{+}24, BAİP24, BBM^{+}22, BCT^{+}21, BAGRB^{+}20, BSOK^{+}20, BEON24, BBN^{+}20, BHL^{+}21, BMM^{+}24, BBB^{+}24, CGMT20, CL20a, CCW^{+}20a, CCL23, CSC23, CMA^{+}22, DLGW^{+}20, DDMP^{+}23, DAM^{+}21, DIB20, DCD^{+}24, DATAA20, DDT^{+}23, DC21, DP19, DP20c, DP21a, DP21b, DHD20, uHA20, Elg20, EEN^{+}24, FLH^{+}24, FMM^{+}20, GS20, GWZ20, GZF^{+}20a, GW22, GBdRACG20, GRN20, GHEB^{+}18, GHEB^{+}23, GMH20, GDGK20, GMP^{+}20a, GdOAO20, GKA^{+}21, GB20, GZZG24, HN23, HIMM20, HMO^{+}20, HLP21, HH22, HIU^{+}22, HTAY21, HZLH21, HYWY22, HAqDE23, HLL^{+}20, HLK^{+}23, HZL^{+}21, HNV^{+}20, IMuI^{+}21, IuRJ^{+}21, JKS20a, JAAAZB20, JZZD21, JLT^{+}21, JKS20b, KAH^{+}23, KNRI21, KSSR20, KK20]. using [KRW^{+}20, Kha24, KAF^{+}23, KAK^{+}23, KMK^{+}23, Kho21b, KAF^{+}20, KYPJ20, KP22, KPA24, KCP23, KLA22, KSDR21, KZF21, LMO^{+}22, LLC^{+}22, LYKK22, LLW^{+}20, LQS^{+}20, LZ21a, LZP23, LGL^{+}20a, LLZ20, Liu24, LCB^{+}23, LAT^{+}20, LPL^{+}20, LP23, MSLJ20, MZLT21, MMP^{+}23, MMPV22, MISS22, MBD21, MOU^{+}21, MHA^{+}24, MSA^{+}24, MA22, MKC^{+}21, MMU^{+}21, NAC^{+}22, NHTH20, NTA^{+}22, NED^{+}20, OMSL20, Pan20, PLBOC20, PJL^{+}24, PMMG^{+}20, PZLL21, QJS^{+}21, QHC24, RSR^{+}24, RDR^{+}24, RAA^{+}20, RAA^{+}21, RAB23, RNRA23, RSFB23, SG20, SKA^{+}20, ŠTI24, SMKA23, SGSGGC^{+}23, SDGCB^{+}20, SMS^{+}24, SBF^{+}21, ST20b, SZdLZ22, SK20b, STH^{+}20, SPRA21, SRA^{+}22, SS22, SSV24, SHF23, SSC^{+}20, SK21b, TGJ^{+}20, TA23, TIA21, TSB20, UAS^{+}20, UUK^{+}21, WHJ20, WZW^{+}20, WDL^{+}21, WLW^{+}21, WHC^{+}24, XYL^{+}20, XRZ^{+}24, XWW^{+}20, XLL20a, XY20, YC22, YHW^{+}20, YZC^{+}20, YZZ^{+}23, YWH^{+}21, YVW^{+}20, YNVRPD23, ZAH^{+}20]. using [ZY20, ZXX^{+}20, ZHD^{+}20, ZZ21a, ZRH^{+}23, ZZP^{+}23, ZKGB20, ZLZ^{+}20b, ZMJ^{+}22, ZPS^{+}24, ZWWC21, ZA22, ZLWH23, uRKI^{+}21]. utilisation [TCBF24]. utility [AP20, JT22, KYY^{+}20, LWL23b, WBR20, YNK^{+}20]. Utilization [ASAM20, PFP^{+}22, AOKÖ24, ASA^{+}20, uRBIBC20, CCW^{+}20a, FLG^{+}20, JHK20, KAA^{+}21, LCH^{+}21, ZHJW20]. Utilization-Aware [ASAM20]. Utilizing [YFL^{+}24, PSHW20]. UWPEE [XLG^{+}23].

V [GMFC23, HB21]. V2X [SLH^{+}24]. v3 [LYH^{+}21]. vacation [MGW23]. Vadalog [BFG^{+}22]. valid [OTMN23]. validation [BDK^{+}20, CBC^{+}20, CKZ^{+}22]. Validity [SBD^{+}24, MBGC20, SZZY22]. validity-preserving [MBGC20]. value [DLZ^{+}23, MSBAU24, RNA21, THB23, ZZ24]. valued [YZZ^{+}23]. values [SCL20, ZZQ21]. VANET [BÖ20a]. VANETs [BQI^{+}20, FZ20, HAA^{+}20, WZTL20]. variable [DZXS21, SPWL23, WG21]. variance [MSA^{+}24]. variant [DAM^{+}21, MMC^{+}23, QJZ^{+}20, YC22]. variants [TLJ^{+}22, VDSB22]. variation [CCL^{+}20, JPW20]. variation-aware [CCL^{+}20]. variational [XYL^{+}20]. variogram [ZLZ21]. Various [WPPA22a, HWQ^{+}20]. varying [LQNW20, ZYX^{+}20]. vascular [HZL^{+}21]. vector [GS20, GMFC23, RZIX20, VPA20, STH^{+}20]. vegetable [LZL^{+}20]. Vehicle [KLA22, LYW23, SGS24, AOKÖ24, BDG23, DRC20, LKJN^{+}20, LQYL21, LP23, MLN24, PJBB20, QC21, UKY^{+}20, ZNZ^{+}23, RNA^{+}22]. vehicle-aided [LQYL21]. vehicles [AYY^{+}20, MMH^{+}22, PZHD20, TDL^{+}21, WC22b, ZLS22b, ARbL^{+}20, LWL23b, SCXZ23, WXZZ22, YZW^{+}23, ZLS^{+}22a]. Vehicular [AAH^{+}23, ABOS22, HDN^{+}20, HB21, SP23, SCXZ23, SKX^{+}20, CLQS20, GZG20, GlRpG20, HHH22, JYP24, KMCJ20, LCB^{+}20, MISB22, NIB^{+}21, RBSK23, ZWX^{+}23]. velocity [HYRZ20, KHL20]. velocity-aware [HYRZ20]. vendor [JLS^{+}23]. vendor-agnostic [JLS^{+}23]. venue [PP20]. Verifiable [BKHD20, CQA^{+}24, LYY^{+}22, LMW^{+}24, LWY^{+}24, SIG24, SLS^{+}20, TKS^{+}23, WZXX21, XYH^{+}24, XPT^{+}22, YYN^{+}20, ZHLL24]. verification [BDK^{+}20, CFM^{+}22, DQBS20, GJC^{+}20, LRCL24, PSHW20, RJM^{+}21, RHJ20, SLS^{+}20, SWL^{+}20, WLN^{+}21, XWW^{+}20]. verifier [WHJ20]. Versatility [SZZY22]. versus [Che20]. vertebroplasty [WWP19, WWP20]. vertex [ŠHDT21]. VERtical [WMNV20, PK24, EL23, LZL^{+}23, MGC23, TM20, WLY23]. Vessel [HIdAR^{+}20, WZC^{+}22]. Vessel-GAN [WZC^{+}22]. Vesta [MAC^{+}21]. VFL [SLH^{+}24]. VFL-Chain [SLH^{+}24]. via [ABC^{+}24, ACG^{+}20a, BEM^{+}24, BÖ20a, CXHS20, CCHD21, CPH^{+}22, CDY^{+}20, CSS22, DLHD22, DLZ^{+}23, FGP23, FHGF20, GCM21, GSSB24, GCPM22, GSG^{+}23, Gur21a, Gur21b, HN22, HZPS21, JPW20, JWZ^{+}22, LCH^{+}22, LWW^{+}22, LYG^{+}24, LZL^{+}21, LGJ^{+}23, LZC^{+}23b, LMZL24, PSMF21, PCVN21, QLJ21, SCGVP20, SHT^{+}21, SYXL22, VS20, WDHY20, WCY^{+}21, WXZX23, XJL^{+}24, YJH^{+}20, YCG^{+}20, YLGG21, ZGL^{+}23, ZZZ21b]. viable [HCK20b]. ViBrAtIoNs [Gur21a]. victim [JT22]. Video [BEB^{+}20, UUH^{+}22, YWS21, YLL22, BSH^{+}24, FS21, FCGPSG^{+}21, HWH^{+}23b, JL21, MSV^{+}20, PLL^{+}24, QCY^{+}21, RGP^{+}22, TLC^{+}20, WMU^{+}23, WWY^{+}24, YLD^{+}23]. videos [BEM^{+}20, HOV20, OHÁV20]. view [BAMR20, LXH^{+}21, LZL^{+}21, MBB^{+}20, NTY^{+}21, PLL^{+}24, SYXL22, VEH^{+}23, XW21, ZZZ^{+}23, ZLZ^{+}20b, ZZZ21b]. viewport [YLKK20]. VIMAC [SKX^{+}20]. violation [LLY^{+}20]. Violations [MOU^{+}21]. Virtual [KAF^{+}23, PAC^{+}22, RNA^{+}22, ZGK^{+}22, LKE22, ATZP21, ABAJ20, CWM21, CCTZ22, CVdRA^{+}20, CCL^{+}20, CHS^{+}23, CLL^{+}23, Dho20, DLH^{+}20, DHD20, LZLY20, LFHS23, LCC^{+}24b, MGW23, SCGVP20, SDGCB^{+}20, SS22, SAF23, VGL23, WHW20, XZK^{+}20, YLKK20, YWH^{+}23, YPEK23, ZZ21b]. virtual-reality [ZZ21b]. virtualise [DJP^{+}24]. Virtualization [HTAY21, MDZ^{+}21, AKCP21, CMGI^{+}23, KF22, KF23, YXL^{+}20]. virtualized [BSOK^{+}20, ENT^{+}22, XWD20]. Virtualizing [CCDR22]. Vision [FLF^{+}21, MKK^{+}20, LZH^{+}20, RSFB23, SAAEK22, WSL21, Wei21, ZLH^{+}24]. Vision-based [MKK^{+}20]. Vision-guided [FLF^{+}21]. VISO [CPPY24]. Visual [ZIOT^{+}20, DLdAR23, DBBP24, GPWL20, LRCL24, ZZ21b]. Visualization [DAM^{+}21, FSBŞ^{+}20, DP20b, GHG^{+}21, LYFZ20, ONK^{+}20, PMMSE21, WDS^{+}23, ZCS20]. visualizations [FSBŞ^{+}20]. visualizing [MFMSG20]. vital [CdRRdC^{+}24]. Vivace [SYW^{+}23]. Vivace-Distributed [SYW^{+}23]. VM [BRK24, CCW^{+}20c, CP22, GPR23, LYY^{+}20a, LDGS20, MK22, MOW^{+}20, NT22, THB23, WLL21]. VMP [PAC^{+}22]. VNE [ZGK^{+}22]. VoIP [SMKC20]. volatile [BPSP23, MSZ^{+}20]. volatility [BEL20, WHC^{+}24]. VoLTE [ZZXH20]. volume [FLG^{+}20, KIM^{+}24, MTHA24a, XWR24, ZGW^{+}23a]. volumes [SDGCB^{+}20]. Volumetric [HIdAR^{+}20]. voluntarily [GMGV^{+}22]. volunteer [BPCM21, MGB24, MA22]. volunteered [TD21]. Voronoi [GGCIV20]. Voronoi-Diagram [GGCIV20]. voting [AJPM20, FWZ^{+}20, KAK20, LWW^{+}22, LZTM24, YYN^{+}20, FWZ^{+}20]. voting-based [LWW^{+}22]. vSwitch [MLX23]. Vulnerabilities [SAAEK22, LCH^{+}22]. vulnerability [hAS24, ACG^{+}20a, GMP20b, MCT^{+}22, VCM^{+}21, WXZX23, YLF^{+}23, YFL^{+}24]. vulnerability-aware [ACG^{+}20a].

wake [ZZD22]. walk [RSBM20]. walks [AD21]. WAN [LYS^{+}20, ZHX^{+}20]. warehouse [GPGG23, ZWC^{+}22]. warning [WC22b, WYGP21, YYB^{+}21, ZLS23]. warp [bHFF^{+}21, ZGN^{+}20]. wars [FIABC^{+}20]. WASPAS [GSKS20]. waste [HBF24, LGYC20]. water [WWS20]. waterbirds [SAM^{+}24]. watermarking [SSA^{+}23]. wave [MXS22]. waveform [WCY^{+}21]. wavefront [HMSA^{+}23]. wavelet [GHEB^{+}18, GHEB^{+}23, KAH^{+}23, XLG^{+}23]. way [ABMMC18, ABMMC22, ASBT20, CDBD24, XZZ^{+}20b]. WBANs [HWH^{+}23a, WHA^{+}20]. WBATimeNet [MK22]. WBSNs [ZHGX20]. weak [FWY^{+}22, WYZ^{+}20]. weakly [HZL^{+}21]. Wear [YYZ^{+}24, HL24]. Wear-leveling-aware [YYZ^{+}24]. Wearable [XLCB20, AESI^{+}21, CHJ^{+}20, Dao23, DP19, DP20c, DP21a, DP21b, GA22, HHW^{+}22, KZG^{+}22, LLFQ21, ZHGX20, ZWL21, ZWL22]. wearable-based [HHW^{+}22]. Weather [CSAT24, CKFT20, PCCX21]. Weather-aware [CSAT24]. Web [AKPT20, AGdS^{+}21, CSY^{+}20, FSBŞ^{+}20, MK20, ORLV20, PGSM^{+}24, QWR^{+}20, SKH20, XFJ^{+}20, ZLL^{+}24a, ZZPK21, LLW^{+}22b, OLP23, SMC23]. Web-based [OLP23]. Web3 [SXHD24, dVIP24]. webpage [See20]. website [LJ21]. WeChat [LZL^{+}20]. weight [CKZ^{+}22, ISD22, JPW20, Kha24, THVL24, WDL^{+}21]. weighted [BPUW24, NZY^{+}23, TWL23, XCW20, ZT22b]. weighting [ArMA^{+}21]. Welch [Che20]. welfare [LXZ^{+}20, XGY^{+}23]. well [LLZ^{+}21, TDMC23]. well-founded [TDMC23]. Wellbeing [RD23]. wellness [ZLM^{+}23]. WFA [HMSA^{+}23]. WFA-FPGA [HMSA^{+}23]. WfChef [CCdS23]. WfCommons [CCP^{+}22]. WGSDMM [ASA23]. Whale [LLZ^{+}22, ABMESM18, ABMESM22, AEZ22, MHH^{+}20]. Where [DCC22]. whistleblower [PDT21]. White [PFS^{+}23]. whitelist [XCS^{+}22]. who [AAR^{+}20]. Whole [HIdAR^{+}20]. Whole-Heart [HIdAR^{+}20]. Wi [CZGS20, CDY^{+}20, YZZ^{+}23]. Wi-Fi [CZGS20, CDY^{+}20, YZZ^{+}23]. Wide [SGL^{+}20a, CSY^{+}20, MAA22, WYWS22]. wide-area [MAA22]. Wide-grained [SGL^{+}20a]. wide-range [CSY^{+}20]. width [YYXZ23]. WIFI [Ano24s, AKF^{+}20, NHTH20, TRB^{+}23a, TK24b, TPN^{+}21]. WiFi-based [TK24b, TPN^{+}21]. wild [uHA20]. WiMAX [CSD^{+}23]. WiMAX-based [CSD^{+}23]. win [ZAH^{+}20]. win-win [ZAH^{+}20]. wind [YZL^{+}20]. window [WCY^{+}21]. windowed [GDGK20]. windows [uRBIBC20]. Wireless [Gul22, KBG20, MR23b, MO24b, WMU^{+}24, WHF^{+}20, YXLB20, Deh20, FCGPSG^{+}21, GCT^{+}20, GAT^{+}20, HAB^{+}20, HYC^{+}23, LZA^{+}20, Liu23, dTGC20, LHY^{+}20b, MLZ^{+}23a, MNA^{+}23, NNH^{+}20, NTA^{+}22, SAM^{+}24, TLKX21, TWM^{+}23, WC22a, WZW^{+}20, WLAC20, WZS^{+}23, XLG^{+}23, YGE21, bZSC^{+}23, ZMJ^{+}22, MKK^{+}20, Zhu21]. wise [AHL^{+}23, JZL^{+}20, ZWM^{+}23, ZRH^{+}23]. within [AOSA20b, AOSA20a, CF21, DOR^{+}21, KCB20, MOW^{+}20, RNA21]. without [MSK^{+}21, QHE^{+}20, XZZ^{+}20b, XSW^{+}21, YYN^{+}20, YZW22]. WolfGraph [ZHLM20]. word [ASA23, AR20, GDCGCPVG21, Gas22, ZXY^{+}21]. work [JMHB24, MGB24, MMR23a]. Worker [qLhZ20, DT21, MAA22]. workers [KOM^{+}20]. workflow [ASPG^{+}21, AB20, BLT^{+}24, BGR20, BYH^{+}20, CLLCK20, CdST^{+}20, CCP^{+}22, CCdS23, GB20, HSS20, HWR^{+}22, IT20, Kha24, LWX22, LS23b, LZHS24, MA24, MBD^{+}20, NSR^{+}23, OLP23, OLLP24, PKR21, RPF21, RAS^{+}22, SGDK^{+}21, SXZZ23, SNS^{+}20, VI21, WGG^{+}20, WGW^{+}21, WWW^{+}24, XZYH22, YYXZ23, ZWW^{+}20b, ZA22]. workflow-based [BYH^{+}20]. Workflows [GMF^{+}20, RCLEB20, ABC^{+}20, ALS^{+}21b, BQC23, BSB^{+}22, BGMK22, CQS^{+}23, CAC^{+}22, CDBD24, DGL^{+}20, DK20, EBA^{+}22, GMM22, GPR^{+}24, GBM20, HdOP^{+}21, HHLZ20, KPA24, LTX^{+}24, LZL^{+}24a, LFYH22, MMKS22, MGZ^{+}20, MM21b, PWV^{+}21, PMMSE21, RCJZ20, RRHA21, SBD^{+}24, SEKS^{+}20, SW20, STH^{+}20, SPRA21, SSC^{+}20, STK23, WLD^{+}20b, XZJ^{+}20, dSGST21]. Workload [CHJK22, KTIB22, KHB23, AÖ24, BMBC20, MMBD20, MBD21, PB23]. Workload-optimized [CHJK22]. Workloads [PFS^{+}23, EET20, GA22, GDS^{+}20, JCP^{+}20, MGGG^{+}20, PPGS20, QNRA23, RAL^{+}24, SAT20, SOKW^{+}20, WCY^{+}21, ZZT^{+}22]. workspaces [DML20]. world [GA22, GMBdF^{+}23, JPJO22]. wormhole [KTC23]. WRENCH [CdST^{+}20]. write [WX24]. write-optimized [WX24]. WSAN [GSARS20]. WSN [MAS23]. WSNs [GKA^{+}21, KSH^{+}21, WZL^{+}20, WHZ^{+}20]. WUR [Par20]. WWW [See20].

X [CLZ^{+}20, DLGW^{+}20, HZLH21, LHTSM^{+}23, MSK^{+}21]. X-ray [CLZ^{+}20, DLGW^{+}20, HZLH21, LHTSM^{+}23]. X-rays [MSK^{+}21]. x86 [CBC^{+}20]. xAI [ERG^{+}22]. Xel} [BLGCLA^{+}23]. XGBoost [KAH^{+}23, WZH^{+}22]. XSRU [KMR^{+}22].

YARN [CREE^{+}24]. year [BPLFRL20]. YOLO [PBSS24]. YOLO-based [PBSS24]. young [GMMR24].

Zero [KAJ^{+}24, TTTH20, WLZ^{+}23, XCZ^{+}22]. zero-forcing [WLZ^{+}23]. zeroknowledge [XCZ^{+}22]. zerorized [HLK^{+}23]. Zipf [EKJ^{+}20]. zkrpChain [XCZ^{+}22]. zone [GlRpG20]. zone-based [GlRpG20].

References

Adam:2020:OTS

[AAA20] Mohammed Sani Adam, Mohammad Hossein Anisi, and Ihsan Ali. Object tracking sensor networks in smart cities: Taxonomy, architecture, applications, research challenges and future directions. Future Generation Computer Systems, 107(??):909–923, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                                       www.sciencedirect.com/
                                                       science/article/pii/S0167739X17310385.

Anisetti:2023:APB

[AAB23] Marco Anisetti, Claudio A. Ardagna, and Filippo Berto. An assurance process for Big Data trustworthiness. Future Generation Computer Systems, 146(??):34–46, September 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                                       www.sciencedirect.com/
                                                       science/article/pii/S0167739X23001371.

Ardebili:2024:HTH

[AABB24] Mohsen Seyedkazemi Ardebili, Andrea Acquaviva, Luca Benini, and Andrea Bartolini. HazardNet: a thermal hazard prediction framework for datacenters. Future Generation Computer Systems, 155(??): 340–353, June 2024. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24000347.

Al-Ali:2022:GDU

[AABKB22] Rima Al-Ali, Lubomír Bulej, Jan Kofroň, and Tomáš Bureš. A guide to design uncertainty-aware self-adaptive components in cyber–physical systems. Future Generation Computer Systems, 128(??): 466–489, March 2022. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21004234.

Ahmed:2023:HNE

[AACJ23] Imran Ahmed, Misbah Ahmad, Abdellah Chehri, and Gwanggil Jeon. A heterogeneous network embedded medicine recommendation system based on LSTM. Future Generation Computer Systems, 149(??):1– 11, December 2023. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23002522.

Andresini:2021:GAD

[AADM21] Giuseppina Andresini, Annalisa Appice, Luca De Rose, and Donato Malerba. GAN augmentation to deal with imbalance in imaging-based intrusion detection. Future Generation Computer Systems, 123(??):108–127, October 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21001382.

Awaysheh:2020:NGB

[AAG^{+}20] Feras M. Awaysheh, Mamoun Alazab, Maanak Gupta, Tomás F. Pena, and José C. Cabaleiro. Next-generation big data federation access control: a reference model. Future Generation Computer Systems, 108(??): 726–741, July 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19315948.

Abdi:2022:MFS

[AAG22] Fatemeh Abdi, Mahmood Ahmadi, and Montajab Ghanem. LA-MDPF: a forwarding strategy based on learning automata and Markov decision process in named data networking. Future Generation Computer Systems, 134(??):22– 39, September 2022. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22001194.

Abdi:2023:IAM

[AAG23] Fatemeh Abdi, Mahmood Ahmadi, and Montajab Ghanem. AM-IF: Adaptive Multi-Path Interest Forwarding in named data networking. Future Generation Computer Systems, 148(??):564–583, November 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23002479.

Adjei-Arthur:2022:BAC

[AAGX^{+}22] Bonsu Adjei-Arthur, Jianbin Gao, Qi Xia, Eliezer da Silva Tavares, Hu Xia, Sandro Amofa, and Yu Wang. A blockchain-adaptive contractual approach for multicontracting organizational entities. Future Generation Computer Systems, 132(??):93–107, July 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22000462.

Alam:2023:FEL

[AAH^{+}23] Md. Mustakin Alam, Tanjim Ahmed, Meraz Hossain, Mehedi Hasan Emo, Md. Kausar Islam Bidhan, Md. Tanzim Reza, Md. Golam Rabiul Alam, Mohammad Mehedi Hassan, Francesco Pupo, and Giancarlo Fortino. Federated ensemble-learning for transport mode detection in vehicular edge network. Future Generation Computer Systems, 149(??):89– 104, December 2023. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23002674.

Aburukba:2020:SIT

[AALEF20] Raafat O. Aburukba, Mazin AliKarrar, Taha Landolsi, and Khaled El-Fakih. Scheduling Internet of Things requests to minimize latency in hybrid Fog– Cloud computing. Future Generation Computer Systems, 111(??):539–551, October 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18303327.

Anjum:2024:TND

[AAM^{+}24] Afia Anjum, Paul Agbaje, Arkajyoti Mitra, Emmanuel Oseghale, Ebelechukwu Nwafor, and Habeeb Olufowobi. Towards named data networking technology: Emerging applications, use cases, and challenges for secure data communication. Future Generation Computer Systems, 151(??):12–31, February 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23003631.

Alghamdi:2021:DQA

[AAP21] Ibrahim Alghamdi, Christos Anagnostopoulos, and Dimitrios P. Pezaros. Data quality-aware task offloading in mobile edge computing: an optimal stopping theory approach. Future Generation Computer Systems, 117(??):462–479, April 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2033079X.

Azad:2020:SKW

[AAR^{+}20] Muahammad Ajmal Azad, Mamoun Alazab, Farhan Riaz, Junaid Arshad, and Tariq Abullah. Socioscope: I know who you are, a robo, human caller or service number. Future Generation Computer Systems, 105(??):297–307, April 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19315110.

Alazab:2020:IMM

[AAS^{+}20] Moutaz Alazab, Mamoun Alazab, Andrii Shalaginov, Abdelwadood Mesleh, and Albara Awajan. Intelligent mobile malware detection using permission requests and API calls. Future Generation Computer Systems, 107(??):509–521, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19321223.

Amrahov:2024:NAM

[AAT^{+}24] Sahin Emrah Amrahov, Yilmaz Ar, Bulent Tugrul, Bekir Emirhan Akay, and Nermin Kartli. A new approach to Mergesort algorithm: Divide smart and conquer. Future Generation Computer Systems, 157(??):330–343, August 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24001225.

Abdel-Basset:2019:CFR

[AB19] Mohamed Abdel-Basset.

Corrigendum to “A framework for risk assessment, management and evaluation: Economic tool for quantifying risks in supply chain” [Future Gener. Comput. Syst. 90 (2019) 489–502]. Future Generation Computer Systems, 93(??):1076–1077, April 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18331959.

See [ABGMC19] and retraction notice [AB21].

Ali:2020:ICF

[AB20] Ihtisham Ali and Susmit Bagchi. Isolating critical flow path and algorithmic partitioning of the AND/ OR mobile workflow graph. Future Generation Computer Systems, 103(??):28– 43, February 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19319636.

Abdel-Basset:2021:RNC

[AB21] Mohamed Abdel-Basset.

Retraction notice to “Corrigendum to “A framework for risk assessment, management and evaluation: Economic tool for quantifying risks in supply chain”’ [Future Gener. Comput. Syst. (2019) 1076–1077]. Future Generation Computer Systems, 120(??):127, July 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21000972.

See [AB19].

Alebouyeh:2024:BRP

[AB24] Zeinab Alebouyeh and Amir Jalaly Bidgoly. Benchmarking robustness and privacy-preserving methods in federated learning. Future Generation Computer Systems, 155(??):18– 38, June 2024. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24000086.

Abu:2024:DPI

[ABA24] Amuda James Abu, Brahim Bensaou, and Ahmed M. Abdelmoniem. Dimensioning the pending interest table in contentcentric networks. Future Generation Computer Systems, 152(??):179–192, March 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23003837.

Alahmadi:2022:NTS

[ABAD22] Hanan Alahmadi, Fatma Bouabdallah, and Ahmed Al-Dubai. A novel timeslotted LoRa MAC protocol for scalable IoT networks. Future Generation Computer Systems, 134(??):287–302, September 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22001261.

Aldegheishem:2020:FFV

[ABAJ20] Abdulaziz Aldegheishem, Rasool Bukhsh, Nabil Alrajeh, and Nadeem Javaid. FaaVPP: Fog as a virtual power plant service for community energy management. Future Generation Computer Systems, 105(??):675–683, April 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19322988.

Andrieu:2021:ERE

[ABB^{+}21] Pierre Andrieu, Bryan Brancotte, Laurent Bulteau, Sarah Cohen-Boulakia, Alain Denise, Adeline Pierrot, and Stéphane Vialette. Efficient, robust and effective rank aggregation for massive biological datasets. Future Generation Computer Systems, 124(??):406–421, November 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21002089.

Ahn:2020:FOS

[ABC^{+}20] Dong H. Ahn, Ned Bass, Albert Chu, Jim Garlick, Mark Grondona, Stephen Herbein, Helgi I. Ingólfsson, Joseph Koning, Tapasya Patki, Thomas R. W. Scogland, Becky Springmeyer, and Michela Taufer. Flux: Overcoming scheduling challenges for exascale workflows. Future Generation Computer Systems, 110(??):202–213, September 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19317169.

Agiollo:2024:AFL

[ABC^{+}24] Andrea Agiollo, Enkeleda Bardhi, Mauro Conti, Nicolò Dal Fabbro, and Riccardo Lazzeretti. Anonymous federated learning via named-data networking. Future Generation Computer Systems, 152(??):

288–303, March 2024. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23004144.

Amatria-Barral:2023:PHE

[ABGDT23] Iñaki Amatria-Barral, Jorge González-Domínguez, and Juan Touriño. pRIblast: a highly efficient parallel application for comprehensive lncRNA–RNA interaction prediction. Future Generation Computer Systems, 138(??):270–279, January 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22002758.

Abdel-Basset:2019:FRA

[ABGMC19] Mohamed Abdel-Basset, M. Gunasekaran, Mai Mohamed, and Naveen Chilamkurti. A framework for risk assessment, management and evaluation: Economic tool for quantifying risks in supply chain. Future Generation Computer Systems, 90(??):489– 502, January 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18312172.

See corrigendum [AB19] and retraction notice [AB21].

Abdel-Basset:2021:RNF

[ABGMC21] Mohamed Abdel-Basset, M. Gunasekaran, Mai Mohamed, and Naveen Chilamkurti. Retraction notice to “A framework for risk assessment, management and evaluation: Economic tool for quantifying risks in supply chain” [Future Gener. Comput. Syst. (2018) 489– 502]. Future Generation Computer Systems, 120 (??):126, July 2021. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21000960.

See [ABGMC19].

Ansari:2022:GBD

[ABL22] Mohammad Samar Ansari, Václav Bartoš, and Brian Lee. GRU-based deep learning approach for network intrusion alert prediction. Future Generation Computer Systems, 128(??):235–247, March 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21003861.

Alidra:2023:FBS

[ABL23] Abdelghani Alidra, Hugo Bruneliere, and Thomas Ledoux. A feature-based survey of Fog modeling languages. Future Generation Computer Systems, 138(??):104–119, January 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22002710.

Abdel-Basset:2019:NPF

[ABM19] Mohamed Abdel-Basset and Mai Mohamed. A novel and powerful framework based on neutrosophic sets to aid patients with cancer. Future Generation Computer Systems, 98(??):144– 153, September 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18325664.

See retraction notice [AB21].

Abdel-Basset:2021:RNN

[ABM21] Mohamed Abdel-Basset and Mai Mohamed. Retraction notice to “A novel and powerful framework based on neutrosophic sets to aid patients with cancer” [Future Gener. Comput. Syst. (2019) 144–153]. Future Generation Computer Systems, 120(??):125, July 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/S0167739X21000959.

See [ABM19].

Abdel-Basset:2018:HWO

[ABMESM18] Mohamed Abdel-Basset, Gunasekaran Manogaran, Doaa El-Shahat, and Seyedali Mirjalili. A hybrid whale optimization algorithm based on local search strategy for the permutation flow shop scheduling problem. Future Generation Computer Systems, 85(??):129–145, August 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18302735.

See retraction notice [ABMMC22].

Abdel-Basset:2022:RNH

[ABMESM22] Mohamed Abdel-Basset, Gunasekaran Manogaran, Doaa El-Shahat, and Seyedali Mirjalili. Retraction notice to “A hybrid whale optimization algorithm based on local search strategy for the permutation flow shop scheduling problem” [Future Gener. Comput. Syst. 85 (2018) 129–145]. Future Generation Computer Systems, 128(??):567, March 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21004453.

See [ABMESM18].

Abdel-Basset:2018:ITI

[ABMM18] Mohamed Abdel-Basset, Gunasekaran Manogaran, and Mai Mohamed. Internet of Things (IoT) and its impact on supply chain: A framework for building smart, secure and efficient systems. Future Generation Computer Systems, 86(??):614–628, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X1830400X.

See retraction notice [ABMMC22].

Abdel-Basset:2022:RNI

[ABMM22] Mohamed Abdel-Basset, Gunasekaran Manogaran, and Mai Mohamed. Retraction notice to “Internet of Things (IoT) and its impact on supply chain: a framework for building smart, secure and efficient systems” [Future Gener. Comput. Syst. 86 (2018) 614–628]. Future Generation Computer Systems, 128(??):568, March 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21004465.

See [ABMM18].

Abdel-Basset:2018:TWD

[ABMMC18] Mohamed Abdel-Basset, Gunasekaran Manogaran, Mai Mohamed, and Naveen Chilamkurti. Three-way decisions based on neutrosophic sets and AHP–QFD framework for supplier selection problem. Future Generation Computer Systems, 89(??):19–30, December 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18310859.

See retraction notice [ABMMC22].

Abdel-Basset:2022:RNT

[ABMMC22] Mohamed Abdel-Basset, Gunasekaran Manogaran, Mai Mohamed, and Naveen Chilamkurti. Retraction notice to “Three-way decisions based on neutrosophic sets and AHP–QFD framework for supplier selection problem” [Future Gener. Comput. Syst. 89 (2018) 19–30]. Future Generation Computer Systems, 128 (??):569, March 2022. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21004477.

See [ABMMC18].

Agiollo:2024:EFE

[ABMO24] Andrea Agiollo, Paolo Bellavista, Matteo Mendula, and Andrea Omicini. EneA-FL: Energy-aware orchestration for serverless federated learning. Future Generation Computer Systems, 154(??):219–234, May 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24000074.

Almeida:2022:VTF

[ABOS22] Ana Almeida, Susana Brás, Ilídio Oliveira, and Susana Sargento. Vehicular traffic flow prediction using deployed traffic counters in a city. Future Generation Computer Systems, 128(??):429–442, March 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21004180.

Altmann:2020:ECS

[ABT20] Jörn Altmann, José Ángel Bañares, and Konstantinos Tserpes. Economics of computing services: an overview of economic motivated solutions to cloud computing. Future Generation Computer Systems, 111(??):931–933, October 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2030563X.

Andujar:2023:EEH

[ACA^{+}23] Francisco J. Andújar, Salvador Coll, Marina Alonso, Juan-Miguel Martínez, Pedro López, José L. Sánchez, and Francisco J. Alfaro. Energy efficient HPC network topologies with on/ off links. Future Generation Computer Systems, 139(??):126–138, February 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22002977.

Agrawal:2023:PSP

[ACBT23] Ankit Agrawal, Sarthak Choudhary, Ashutosh Bhatia, and Kamlesh Tiwari. Pub-SubMCS: a privacy-preserving publish– subscribe and blockchainbased mobile crowdsensing framework. Future Generation Computer Systems, 146(??):234–249, September 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23001516.

Angsuchotmetee:2020:MOB

[ACC20] Chinnapong Angsuchotmetee, Richard Chbeir, and Yudith Cardinale. MSSN-Onto: an ontology-based approach for flexible event processing in Multimedia Sensor Networks. Future Generation Computer Systems, 108(??):1140–1158, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18301419.

Amato:2020:ARA

[ACD^{+}20] Flora Amato, Luigi Coppolino, Salvatore D’Antonio, Nicola Mazzocca, Francesco Moscato, and Luigi Sgaglione. An abstract reasoning architecture for privacy policies monitoring. Future Generation Computer Systems, 106(??):393–400, May 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19324549.

Ardagna:2021:ESI

[ACDY21] Claudio A. Ardagna, Mauro Conti, Ernesto Damiani, and Chia-Mu Yu. Editorial: Special issue on trusted cloud-edges computations. Future Generation Computer Systems, 114(??):661–664, January 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20326819.

Accarino:2021:MMA

[ACF^{+}21] Gabriele Accarino, Marco Chiarelli, Sandro Fiore, Ivan Federico, Salvatore Causio, Giovanni Coppini, and Giovanni Aloisio. A multi-model architecture based on long shortterm memory neural networks for multi-step sea level forecasting. Future Generation Computer Systems, 124(??):1–9, November 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21001588.

Ai:2020:MMP

[ACG^{+}20a] Jianjian Ai, Hongchang Chen, Zehua Guo, Guozhen Cheng, and Thar Baker. Mitigating malicious packets attack via vulnerabilityaware heterogeneous network devices assignment. Future Generation Computer Systems, 111(??): 841–852, October 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0167739X18330991.

Antonelli:2020:AMP

[ACG^{+}20b] Fabio Antonelli, Vittorio Cortellessa, Marco Gribaudo, Riccardo Pinciroli, Kishor S. Trivedi, and Catia Trubiani. Analytical modeling of performance indices under epistemic uncertainty applied to cloud computing systems. Future Generation Computer Systems, 102(??):746–761, January 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19308428.

Alulema:2023:SMB

[ACI^{+}23] Darwin Alulema, Javier Criado, Luis Iribarne, Antonio Jesús Fernández-García, and Rosa Ayala. SI4IoT: a methodology based on models and services for the integration of IoT systems. Future Generation Computer Systems, 143(??):132–151, June 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23000316.

Arafeh:2021:OBR

[ACM^{+}21] Mohamad Arafeh, Paolo Ceravolo, Azzam Mourad, Ernesto Damiani, and Emanuele Bellini. Ontology based recommender system using social network data. Future Generation Computer Systems, 115(??):769–779, February 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20305963.

Abate:2021:AMS

[ACN^{+}21] Andrea F. Abate, Lucia Cascone, Michele Nappi, Fabio Narducci, and Ignazio Passero. Attention monitoring for synchronous distance learning. Future Generation Computer Systems, 125(??):774–784, December 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21002880.

Aydin:2022:UCB

[AÇP22] Hüseyin Aydın, Erkin Çilden, and Faruk Polat. Using chains of bottleneck transitions to decompose and solve reinforcement learning tasks with hidden states. Future Generation Computer Systems, 133(??):153–168, August 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22000917.

Audrito:2024:GFD

[ACT24] Giorgio Audrito, Roberto Casadei, and Gianluca Torta. A general framework and decentralised algorithms for collective computational processes. Future Generation Computer Systems, 158(??):11–27, September 2024. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24001493.

Ardagna:2020:SIT

[ACY20] Claudio A. Ardagna, Mauro Conti, and Chia-Mu Yu. Special issue on trusted cloud-edges computations. Future Generation Computer Systems, 112(??): 1180–1183, November 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167739X20324183.

Assi:2021:IMK

[AD21] Ali Assi and Wajdi Dhifli. Instance matching in knowledge graphs through random walks and semantics. Future Generation Computer Systems, 123(??):73–84, October 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21001369.

Awuson-David:2021:BLA

[ADAHA^{+}21] Kenny Awuson-David, Tawfik Al-Hadhrami, Mamoun Alazab, Nazaraf Shah, and Andrii Shalaginov. BCFL logging: an approach to acquire and preserve admissible digital forensics evidence in cloud ecosystem. Future Generation Computer Systems, 122(??):1– 13, September 2021. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21000807.

Amarasinghe:2020:ESD

[AdAHK20] Gayashan Amarasinghe, Marcos D. de Assunção, Aaron Harwood, and Shanika Karunasekera. ECSNeT++: a simulator for distributed stream processing on edge and cloud environments. Future Generation Computer Systems, 111(??): 401–418, October 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167739X19300494.

Awan:2022:APE

[ADAR22] Kamran Ahmad Awan, Ikram Ud Din, Ahmad Almogren, and Joel J. P. C. Rodrigues. AutoTrust: a privacy-enhanced trustbased intrusion detection approach for Internet of Smart Things. Future Generation Computer Systems, 137(??):288–301, December 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22002606.

Albanese:2020:STC

[ADdMM20]  Massimiliano  Albanese,

Alessandra De Benedictis, Douglas D. J. de Macedo, and Fabrizio Messina. Security and trust in cloud application life-cycle management. Future Generation Computer Systems, 111(??):934–936, October 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20302107.

Al-Dulaimy:2020:BBS

[ADITS20] Auday Al-Dulaimy, Wassim Itani, Javid Taheri, and Maha Shamseddine. bwSlicer: a bandwidth slicing framework for cloud data centers. Future Generation Computer Systems, 112(??):767–784, November 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19328055.

Aldin:2020:STC

[ADMG20] Hesam Nejati Sharif Aldin, Hossein Deldari, Mohammad Hossein Moattar, and Mostafa Razavi Ghods. Strict timed causal consistency as a hybrid consistency model in the cloud environment. Future Generation Computer Systems, 105(??):259–274, April 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19321053.

Astillo:2022:FIA

[ADP^{+}22] Philip Virgil Astillo, Daniel Gerbi Duguma, Hoonyong Park, Jiyoon Kim, Bonam Kim, and Ilsun You. Federated intelligence of anomaly detection agent in IoTMDenabled Diabetes Management Control System. Future Generation Computer Systems, 128(??):395–405, March 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21004192.

Alkhariji:2023:SBP

[ADRP23] Lamya Alkhariji, Suparna De, Omer Rana, and Charith Perera. Semanticsbased privacy by design for Internet of Things applications. Future Generation Computer Systems, 138(??):280–295, January 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22002746.

Adikari:2022:ECA

[AdSM^{+}22] Achini Adikari, Daswin de Silva, Harsha Moraliyage, Damminda Alahakoon, Jiahui Wong, Mathew Gancarz, Suja Chackochan, Bomi Park, Rachel Heo, and Yvonne Leung. Empathic conversational agents for real-time monitoring and co-facilitation of patient-centered healthcare. Future Generation Computer Systems, 126(??):318–329, January 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2100323X.

Alem:2023:NBA

[AEN^{+}23] Salwa Alem, David Espes, Laurent Nana, Eric Martin, and Florent De Lamotte. A novel bi-anomaly-based intrusion detection system approach for Industry 4.0. Future Generation Computer Systems, 145(??): 267–283, August 2023. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23001012.

Ali:2021:IHM

[AESI^{+}21] Farman Ali, Shaker El-Sappagh, S. M. Riazul Islam, Amjad Ali, Muhammad Attique, Muhammad Imran, and Kyung-Sup Kwak. An intelligent healthcare monitoring framework using wearable sensors and social networking data. Future Generation Computer Systems, 114(??):23–43, January 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1931605X.

Anter:2022:RTE

[AEZ22] Ahmed M. Anter, Mohamed Abd Elaziz, and Zhiguo Zhang. Real-time epileptic seizure recognition using Bayesian genetic whale optimizer and adaptive machine learning. Future Generation Computer Systems, 127(??):426–434, February 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21003782.

Arjona:2023:TSE

[AFL23] Aitor Arjona, Gerard Finol, and Pedro García López.

Transparent serverless execution of Python multiprocessing applications. Future Generation Computer Systems, 140(??):436–449, March 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22003612.

Alam:2022:SIM

[AFMG^{+}22] Mehwish Alam, Anna Fensel, Jorge Martinez-Gil, Bernhard Moser, Diego Reforgiato Recupero, and Harald Sack. Special issue on machine learning and knowledge graphs. Future Generation Computer Systems, 129(??):50– 53, April 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21004660.

Adikari:2021:SSA

[AGdS^{+}21] Achini Adikari, Gihan Gamage, Daswin de Silva, Nishan Mills, Sze-Meng Jojo Wong, and Damminda Alahakoon. A self structuring artificial intelligence framework for deep emotions modeling and analysis on the social web. Future Generation Computer Systems, 116(??):302–315, March 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20330053.

Amin:2023:ERT

[AGV23] Youssef Amin, Christian Gianoglio, and Maurizio Valle. Embedded realtime objects’ hardness classification for robotic grippers. Future Generation Computer Systems, 148(??):211–224, November 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23002182.

Akyildiz:2020:MRA

[AGYS20] Taha Atahan Akyildiz, Can Berk Guzgeren, Cemal Yilmaz, and Erkay Savas. MeltdownDetector: a runtime approach for detecting meltdown attacks. Future Generation Computer Systems, 112(??):136–147, November 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19318370.

Al-Hadhrami:2020:RTD

[AHH20] Yahya Al-Hadhrami and Farookh Khadeer Hussain. Real time dataset generation framework for intrusion detection systems in IoT. Future Generation Computer Systems, 108(??):414–423, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19322678.

Akintoye:2023:LWP

[AHL^{+}23] S. B. Akintoye, L. Han, H. Lloyd, X. Zhang, D. Dancey, H. Chen, and D. Zhang. Layer-wise partitioning and merging for efficient and scalable deep learning. Future Generation Computer Systems, 149(??):432–444, December 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2300300X.

Afzal:2023:MAF

[AHMW23] Ayesha Afzal, Georg Hager, Stefano Markidis, and Gerhard Wellein. Making applications faster by asynchronous execution: Slowing down processes or relaxing MPI collectives. Future Generation Computer Systems, 148(??):472–487, November 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2300239X.

Alazab:2021:LBN

[AHN21] Mamoun Alazab, Seung-Hun Hong, and Jenny Ng. Louder bark with no bite: Privacy protection through the regulation of mandatory data breach notification in Australia. Future Generation Computer Systems, 116(??):22– 29, March 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20329940.

Aboutorab:2022:RLB

[AHSH22] Hamed Aboutorab, Omar K. Hussain, Morteza Saberi, and Farookh Khadeer Hussain. A reinforcement learning-based framework for disruption risk identification in supply chains. Future Generation Computer Systems, 126(??):110–122, January 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21003034.

Au:2020:SIC

[AHWB20] Man Ho Au, Jinguang Han, Qianhong Wu, and Colin Boyd. Special issue on cryptographic currency and blockchain technology. Future Generation Computer Systems, 107(??):758–759, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19314840.

Archetti:2023:SSA

[AIM23] Alberto Archetti, Francesca Ieva, and Matteo Matteucci. Scaling survival analysis in healthcare with federated survival forests: a comparative study on heart failure and breast cancer genomics. Future Generation Computer Systems, 149(??):343–358, December 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23002935.

Alhalabi:2021:SMT

[AJJ^{+}21] Wadee Alhalabi, Jari Jussila, Kamal Jambi, Anna Visvizi, Hafsa Qureshi, Miltiadis Lytras, Areej Malibari, and Raniah Samir Adham. Social mining for terroristic behavior detection through Arabic tweets characterization. Future Generation Computer Systems, 116(??):132–144, March 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20330041.

Asadi:2020:DBU

[AJPM20] Mehdi Asadi, Mohammad Ali Jabraeil Jamali, Saeed Parsa, and Vahid Majidnezhad. Detecting botnet by using particle swarm optimization algorithm based on voting system. Future Generation Computer Systems, 107(??):95–111, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19323350.

Ahmad:2020:MDM

[AK20] Shabir Ahmad and Do-Hyeun Kim. A multidevice multi-tasks management and orchestration architecture for the design of enterprise IoT applications. Future Generation Computer Systems, 106(??): 482–500, May 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19318035.

Ali:2020:MAT

[AKA20] Eslam B. Ali, Sherif Kishk, and Ehab H. Abdelhay. Multidimensional auction for task allocation using computation offloading in fifth generation networks. Future Generation Computer Systems, 108(??):

717–725, July 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1832795X.

Apostolopoulos:2021:RAV

[AKCP21] Theodoros Apostolopoulos, Vasilios Katos, Kim-Kwang Raymond Choo, and Constantinos Patsakis. Resurrecting antivirtualization and antidebugging: Unhooking your hooks. Future Generation Computer Systems, 116(??):393–405, March 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20330284.

Awad:2022:SAD

[AKE22] Mirna Awad, Nadjia Kara, and Claes Edstrom. SLOaware dynamic self-adaptation of resources. Future Generation Computer Systems, 133(??):266–280, August 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22000930.

Anbalagan:2020:SAE

[AKF^{+}20] Sudha Anbalagan, Dhananjay Kumar, Mercy Faustina J, Gunasekaran Raja, Waleed Ejaz, and Ali Kashif Bashir. SDN-assisted efficient LTE-WiFi aggregation in next generation IoT networks. Future Generation Computer Systems, 107(??): 898–908, June 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17310907.

Alavizadeh:2020:MBE

[AKJJ20] Hooman Alavizadeh, Dong Seong Kim, and Julian Jang-Jaccard. Model-based evaluation of combinations of shuffle and diversity MTD techniques on the cloud. Future Generation Computer Systems, 111(??):507–522, October 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19315183.

Atzori:2020:SID

[AKPT20] Maurizio Atzori, Georgia Koutrika, Barbara Pes, and Letizia Tanca. Special issue on “Data Exploration in the Web 3.0 Age”. Future Generation Computer Systems, 112(??):1177–1179, November 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20324171.

Arellanes:2020:EIS

[AL20] Damian Arellanes and Kung-Kiu Lau. Evaluating IoT service composition mechanisms for the scalability of IoT systems. Future Generation Computer Systems, 108(??):827–848, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19320291.

Aguilar:2021:PNC

[ALGMP^{+}21] Diana Laura Aguilar, Octavio Loyola-González, Miguel Angel Medina-Pérez, Leonardo Cañete-Sifuentes, and Kim-Kwang Raymond Choo. PBC4occ: a novel contrast pattern-based classifier for one-class classification. Future Generation Computer Systems, 125(??):71–90, December 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21002417.

Ahmed:2020:FNR

[ALR^{+}20] Syed Hassan Ahmed, Jaime Lloret, Danda B. Rawat, Mohsen Guizani, and Wael Guibene. Future networking research plethora for smart cities. Future Generation Computer Systems, 107(??):863–864, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19319739.

Al-Laith:2021:TST

[ALS21a] Ali Al-Laith and Muhammad Shahbaz. Tracking sentiment towards news entities from Arabic news on social media. Future Generation Computer Systems, 118(??):467–484, May 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X2100025X.

Arjona:2021:TTB

[ALS^{+}21b] Aitor Arjona, Pedro García López, Josep Sampé, Aleksander Slominski, and Lionel Villard. Triggerflow: Trigger-based orchestration of serverless workflows. Future Generation Computer Systems, 124(??):215–229, November 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21001989.

Aghda:2020:IRD

[AM20] Seyed Ali Fatemi Aghda and Mahdi Mirfakhraei.

Improved routing in dynamic environments with moving obstacles using a hybrid fuzzy-genetic algorithm. Future Generation Computer Systems, 112(??):250–257, November 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19326676.

Alanezi:2021:EBA

[AM21] Khaled Alanezi and Shivakant Mishra. An edgebased architecture to support the execution of ambience intelligence tasks using the IoP paradigm. Future Generation Computer Systems, 114(??):349–357, January 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20304337.

Alhozaimy:2022:FAP

[AM22] Sarah Alhozaimy and Daniel A.

Menascé. A formal analysis of performance-security tradeoffs under frequent task reconfigurations. Future Generation Computer Systems, 127(??):252–262, February 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21003514.

Ai:2021:EFS

[AMB^{+}21] Qingsong Ai, Wei Meng, Faycal Bensaali, Xiaojun Zhai, Lu Liu, and Nasser Alaraje. Editorial for FGCS special issue: Intelligent IoT systems for healthcare and rehabilitation. Future Generation Computer Systems, 125(??):770–773, December 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21002910.

Asensio:2020:DEC

[AMBD^{+}20] A. Asensio, X. Masip-Bruin, R. J. Durán, I. de Miguel, G. Ren, S. Daijavad, and A. Jukan. Designing an efficient clustering strategy for combined fog-to-cloud scenarios. Future Generation Computer Systems, 109(??):392–406, August 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19329541.

Asensio:2021:OCC

[AMBGS21] A. Asensio, X. Masip-Bruin, J. Garcia, and S. Sánchez. On the optimality of concurrent container clusters scheduling over heterogeneous smart environments. Future Generation Computer Systems, 118(??):157–169, May 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21000054.

Adamovic:2020:ENA

[AMM^{+}20] Saša Adamović, Vladislav Miškovic, Nemanja Maček, Milan Milosavljević, Marko Šarac, Muzafer Saračević, and Milan Gnjatović. An efficient novel approach for iris recognition based on stylometric features and machine learning techniques. Future Generation Computer Systems, 107 (??):144–157, June 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167739X19314463.

Abdellatif:2022:CEH

[AMM^{+}22] Alaa Awad Abdellatif, Naram Mhaisen, Amr Mohamed, Aiman Erbad, Mohsen Guizani, Zaher Dawy, and Wassim Nasreddine. Communicationefficient hierarchical federated learning for IoT heterogeneous systems with imbalanced data. Future Generation Computer Systems, 128(??):406–419, March 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X2100412X.

Aridhi:2020:SIA

[AMNZ20] Sabeur Aridhi, José Macedo, Engelbert Mephu Nguifo, and Karine Zeitouni. Special issue on “Advances on Large Evolving Graphs”. Future Generation Computer Systems, 110(??):310, September 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20312760.

Asim:2020:CCB

[AMR^{+}20] Yousra Asim, Ahmad Kamran Malik, Basit Raza, Wajeeha Naeem, and Saima Rathore. Communitycentric brokerage-aware access control for online social networks. Future Generation Computer Systems, 109(??):469–478, August 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17321817.

Asgari:2021:HSM

[AMT^{+}21] Sahar Asgari, Hosein Moazamigoodarzi, Peiying Jennifer Tsai, Souvik Pal, Rong Zheng, Ghada Badawy, and Ishwar K. Puri. Hybrid surrogate model for online temperature and pressure predictions in data centers. Future Generation Computer Systems, 114(??):531–547, January 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19329036.

Anter:2023:RIR

[AMZZ23] Ahmed M. Anter, Ali W.

Mohamed, Min Zhang, and Zhiguo Zhang. A robust intelligence regression model for monitoring Parkinson’s disease based on speech signals. Future Generation Computer Systems, 147(??):316–327, October 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23001942.

Alsaeed:2024:SLG

[ANA24] Norah Alsaeed, Farrukh Nadeem, and Faisal Albalwy. A scalable and lightweight group authentication framework for Internet of medical things using integrated blockchain and fog computing. Future Generation Computer Systems, 151(??):162–181, February 2024. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23003643.

Abdmeziem:2024:GKM

[AND24] Mohammed Riyadh Abdmeziem, Amina Ahmed Nacer, and Nawfel Moundji Deroues. Group key management in the Internet of Things: Handling asynchronicity. Future Generation Computer Systems, 152(??):273–287, March 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23003965.

Azhir:2021:EAI

[ANH^{+}21] Elham Azhir, Nima Jafari Navimipour, Mehdi Hosseinzadeh, Arash Sharifi, and Aso Darwesh. An efficient automated incremental density-based algorithm for clustering and classification. Future Generation Computer Systems, 114(??):665–678, January 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19327050.

Anonymous:2020:EBa

[Ano20a]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 102 (??):ii, January 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19329966.

Anonymous:2020:EBb

[Ano20b]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 103 (??):ii, February 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19329267.

Anonymous:2020:EBc

[Ano20c]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 104(??):ii, March 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19330171.

Anonymous:2020:EBd

[Ano20d]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 105(??):ii, April 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20301758.

Anonymous:2020:EBe

[Ano20e]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 106(??):ii, May 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20305756.

Anonymous:2020:EBf

[Ano20f]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 107(??):ii, June 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20307986.

Anonymous:2020:EBg

[Ano20g]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 108(??):ii, July 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20311638.

Anonymous:2020:EBh

[Ano20h]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 109 (??):ii, August 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20313145.

Anonymous:2020:EBi

[Ano20i]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 110(??):ii, September 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20314709.

Anonymous:2020:EBj

[Ano20j]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 111 (??):ii, October 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20316939.

Anonymous:2020:EBk

[Ano20k]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 112(??):ii, November 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20324225.

Anonymous:2020:EBl

[Ano20l]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 113(??):ii, December 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20326571.

Anonymous:2020:PAa

[Ano20m] Anonymous. Pages 1–1042 (April 2020). Future Generation Computer Systems, 105(??):1–1042, April 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2020:PJa

[Ano20n] Anonymous. Pages 1–1062 (January 2020). Future Generation Computer Systems, 102(??):1–1062, January 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2020:PJb

[Ano20o] Anonymous. Pages 1–1154 (June 2020). Future Generation Computer Systems, 107(??):1–1154, June 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2020:PS

[Ano20p] Anonymous. Pages 1–1166 (September 2020). Future Generation Computer Systems, 110(??):1–1166, September 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2020:PN

[Ano20q] Anonymous. Pages 1–1184 (November 2020). Future Generation Computer Systems, 112(??):1–1184, November 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2020:PF

[Ano20r] Anonymous. Pages 1–122 (February 2020). Future Generation Computer Systems, 103(??):1–122, February 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2020:PJc

[Ano20s] Anonymous. Pages 1–1360 (July 2020). Future Generation Computer Systems, 108(??):1–1360, July 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2020:PMa

[Ano20t] Anonymous. Pages 1–224 (March 2020). Future Generation Computer Systems, 104(??):1–224, March 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2020:PD

[Ano20u] Anonymous. Pages 1–634 (December 2020). Future Generation Computer Systems, 113(??):1–634, December 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2020:PMb

[Ano20v] Anonymous. Pages 1–690 (May 2020). Future Generation Computer Systems, 106(??):1–690, May 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2020:PAb

[Ano20w] Anonymous. Pages 1–712 (August 2020). Future Generation Computer Systems, 109(??):1–712, August 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2020:PO

[Ano20x] Anonymous. Pages 1–940 (October 2020). Future Generation Computer Systems, 111(??):1–940, October 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2021:EBa

[Ano21a]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 114 (??):ii, January 2021. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20329721.

Anonymous:2021:EBb

[Ano21b]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 115 (??):ii, February 2021. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20330090.

Anonymous:2021:EBc

[Ano21c]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 116(??):ii, March 2021. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2033051X.

Anonymous:2021:EBd

[Ano21d]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 117(??):ii, April 2021. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2100011X.

Anonymous:2021:EBe

[Ano21e]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 118(??):ii, May 2021. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21000534.

Anonymous:2021:EBf

[Ano21f]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 119(??):ii, June 2021. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21000753.

Anonymous:2021:EBg

[Ano21g]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 120(??):ii, July 2021. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21001060.

Anonymous:2021:EBh

[Ano21h]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 121 (??):ii, August 2021. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21001254.

Anonymous:2021:EBi

[Ano21i]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 122(??):ii, September 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21001527.

Anonymous:2021:EBj

[Ano21j]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 123 (??):ii, October 2021. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21001928.

Anonymous:2021:EBk

[Ano21k]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 124(??):ii, November 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2100265X.

Anonymous:2021:EBl

[Ano21l]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 125(??):ii, December 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21003083.

Anonymous:2021:PAb

[Ano21m] Anonymous. Pages 1–122 (August 2021). Future Generation Computer Systems, 121(??):1–122, August 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2021:PJc

[Ano21n] Anonymous. Pages 1–128 (July 2021). Future Generation Computer Systems, 120(??):1–128, July 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2021:PJb

[Ano21o] Anonymous. Pages 1–200 (June 2021). Future Generation Computer Systems, 119(??):1–200, June 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2021:PS

[Ano21p] Anonymous. Pages 1–234 (September 2021). Future Generation Computer Systems, 122(??):1–234, September 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2021:PO

[Ano21q] Anonymous. Pages 1–262 (October 2021). Future Generation Computer Systems, 123(??):1–262, October 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2021:PMa

[Ano21r] Anonymous. Pages 1–454 (March 2021). Future Generation Computer Systems, 116(??):1–454, March 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2021:PMb

[Ano21s] Anonymous. Pages 1–496 (May 2021). Future Generation Computer Systems, 118(??):1–496, May 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2021:PN

[Ano21t] Anonymous. Pages 1–496 (November 2021). Future Generation Computer Systems, 124(??):1–496, November 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2021:PAa

[Ano21u] Anonymous. Pages 1–520 (April 2021). Future Generation Computer Systems, 117(??):1–520, April 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2021:PJa

[Ano21v] Anonymous. Pages 1–692 (January 2021). Future Generation Computer Systems, 114(??):1–692, January 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2021:PF

[Ano21w] Anonymous. Pages 1–880 (February 2021). Future Generation Computer Systems, 115(??):1–880, February 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2021:PD

[Ano21x] Anonymous. Pages 1–954 (December 2021). Future Generation Computer Systems, 125(??):1–954, December 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2022:EBa

[Ano22a]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 126 (??):ii, January 2022. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21003460.

Anonymous:2022:EBb

[Ano22b]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 127 (??):ii, February 2022. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2100409X.

Anonymous:2022:EBc

[Ano22c]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 128(??):ii, March 2022. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21004623.

Anonymous:2022:EBd

[Ano22d]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 129(??):ii, April 2022. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22000048.

Anonymous:2022:EBe

[Ano22e]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 130(??):ii, May 2022. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22000395.

Anonymous:2022:EBf

[Ano22f]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 131(??):ii, June 2022. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22000681.

Anonymous:2022:EBg

[Ano22g]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 132(??):ii, July 2022. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X2200111X.

Anonymous:2022:EBh

[Ano22h]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 133 (??):ii, August 2022. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22001455.

Anonymous:2022:EBi

[Ano22i]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 134(??):ii, September 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22001832.

Anonymous:2022:EBj

[Ano22j]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 135 (??):ii, October 2022. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22002138.

Anonymous:2022:EBk

[Ano22k]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 136(??):ii, November 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22002497.

Anonymous:2022:EBl

[Ano22l]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 137(??):ii, December 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22002825.

Anonymous:2022:PJb

[Ano22m] Anonymous. Pages 1–308 (June 2022). Future Generation Computer Systems, 131(??):1–308, June 2022.

CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2022:PJc

[Ano22n] Anonymous. Pages 1–310 (July 2022). Future Generation Computer Systems, 132(??):1–310, July 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2022:PMb

[Ano22o] Anonymous. Pages 1–320 (May 2022). Future Generation Computer Systems, 130(??):1–320, May 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2022:PJa

[Ano22p] Anonymous. Pages 1–340 (January 2022). Future Generation Computer Systems, 126(??):1–340, January 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2022:PN

[Ano22q] Anonymous. Pages 1–378 (November 2022). Future Generation Computer Systems, 136(??):1–378, November 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2022:PD

[Ano22r] Anonymous. Pages 1–394 (December 2022). Future Generation Computer Systems, 137(??):1–394, December 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2022:PAb

[Ano22s] Anonymous. Pages 1–410 (August 2022). Future Generation Computer Systems, 133(??):1–410, August 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2022:PS

[Ano22t] Anonymous. Pages 1–430 (September 2022). Future Generation Computer Systems, 134(??):1–430, September 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2022:PAa

[Ano22u] Anonymous. Pages 1–464 (April 2022). Future Generation Computer Systems, 129(??):1–464, April 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2022:PO

[Ano22v] Anonymous. Pages 1–466 (October 2022). Future Generation Computer Systems, 135(??):1–466, October 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2022:PF

[Ano22w] Anonymous. Pages 1–502 (February 2022). Future Generation Computer Systems, 127(??):1–502, February 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2022:PMa

[Ano22x] Anonymous. Pages 1–570 (March 2022). Future Generation Computer Systems, 128(??):1–570, March 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2023:EBa

[Ano23a]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 138 (??):ii, January 2023. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22003132.

Anonymous:2023:EBb

[Ano23b]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 139 (??):ii, February 2023. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22003442.

Anonymous:2023:EBc

[Ano23c]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 140(??):ii, March 2023. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22003909.

Anonymous:2023:EBd

[Ano23d]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 141(??):ii, April 2023. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23000079.

Anonymous:2023:EBe

[Ano23e]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 142(??):ii, May 2023. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23000390.

Anonymous:2023:EBf

[Ano23f]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 143(??):ii, June 2023. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23000766.

Anonymous:2023:EBg

[Ano23g]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 144(??):ii, July 2023. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23001139.

Anonymous:2023:EBh

[Ano23h]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 145 (??):ii, August 2023. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23001590.

Anonymous:2023:EBi

[Ano23i]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 146(??):ii, September 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23001887.

Anonymous:2023:EBj

[Ano23j]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 147 (??):ii, October 2023. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23002273.

Anonymous:2023:EBk

[Ano23k]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 148(??):ii, November 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23002819.

Anonymous:2023:EBl

[Ano23l]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 149(??):ii, December 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23003424.

Anonymous:2023:PS

[Ano23m] Anonymous. Pages 1–272 (September 2023). Future Generation Computer Systems, 146(??):1–272, September 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2023:PF

[Ano23n] Anonymous. Pages 1–280 (February 2023). Future Generation Computer Systems, 139(??):1–280, February 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2023:PJa

[Ano23o] Anonymous. Pages 1–338 (January 2023). Future Generation Computer Systems, 138(??):1–338, January 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2023:PJc

[Ano23p] Anonymous. Pages 1–344 (July 2023). Future Generation Computer Systems, 144(??):1–344, July 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2023:PO

[Ano23q] Anonymous. Pages 1–370 (October 2023). Future Generation Computer Systems, 147(??):1–370, October 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2023:PM

[Ano23r] Anonymous. Pages 1–410 (May 2023). Future Generation Computer Systems, 142(??):1–410, May 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2023:PJb

[Ano23s] Anonymous. Pages 1–422 (June 2023). Future Generation Computer Systems, 143(??):1–422, June 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2023:PMa

[Ano23t] Anonymous. Pages 1–466 (March 2023). Future Generation Computer Systems, 140(??):1–466, March 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2023:PAb

[Ano23u] Anonymous. Pages 1–590 (August 2023). Future Generation Computer Systems, 145(??):1–590, August 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2023:PN

[Ano23v] Anonymous. Pages 1–636 (November 2023). Future Generation Computer Systems, 148(??):1–636, November 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2023:PD

[Ano23w] Anonymous. Pages 1–732 (December 2023). Future Generation Computer Systems, 149(??):1–732, December 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2023:PAa

[Ano23x] Anonymous. Pages 1–776 (April 2023). Future Generation Computer Systems, 141(??):1–776, April 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2024:EBa

[Ano24a]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 150 (??):ii, January 2024. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23003710.

Anonymous:2024:EBb

[Ano24b]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 151 (??):ii, February 2024. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23004107.

Anonymous:2024:EBc

[Ano24c]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 152(??):ii, March 2024. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23004430.

Anonymous:2024:EBd

[Ano24d]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 153(??):ii, April 2024. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24000281.

Anonymous:2024:EBe

[Ano24e]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 154(??):ii, May 2024. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2400044X.

Anonymous:2024:EBf

[Ano24f]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 155(??):ii, June 2024. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X24001766.

Anonymous:2024:EBg

[Ano24g]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 156(??):ii, July 2024. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X24001821.

Anonymous:2024:EBh

[Ano24h]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 157 (??):ii, August 2024. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X2400205X.

Anonymous:2024:EBi

[Ano24i]  Anonymous.  Editorial

Board. Future Generation Computer Systems, 158(??):ii, September 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X2400267X.

Anonymous:2024:PF

[Ano24j] Anonymous. Pages 1–272 (February 2024). Future Generation Computer Systems, 151(??):1–272, February 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2024:PJc

[Ano24k] Anonymous. Pages 1–338 (July 2024). Future Generation Computer Systems, 156(??):1–338, July 2024. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2024:PMa

[Ano24l] Anonymous. Pages 1–372 (March 2024). Future Generation Computer Systems, 152(??):1–372, March 2024. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2024:PJa

[Ano24m] Anonymous. Pages 1–450 (January 2024). Future Generation Computer Systems, 150(??):1–450, January 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2024:PJb

[Ano24n] Anonymous. Pages 1–486 (June 2024). Future Generation Computer Systems, 155(??):1–486, June 2024. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2024:PM

[Ano24o] Anonymous. Pages 1–490 (May 2024). Future Generation Computer Systems, 154(??):1–490, May 2024. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2024:PS

[Ano24p] Anonymous. Pages 1–570 (September 2024). Future Generation Computer Systems, 158(??):1–570, September 2024. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2024:PAa

[Ano24q] Anonymous. Pages 1–586 (April 2024). Future Generation Computer Systems, 153(??):1–586, April 2024. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2024:PAb

[Ano24r] Anonymous. Pages 1–638 (August 2024). Future Generation Computer Systems, 157(??):1–638, August 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2024:PSI

[Ano24s] Anonymous. Preface of special issue on explainable AI-empowered Internet of Things for indoor navigation using WIFI sensing. Future Generation Computer Systems, 153(??): 249–250, April 2024. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23004399.

Ali:2024:STA

[AÖ24] Ahsan Ali and Öznur Özkasap. Spatial and thermal aware methods for efficient workload management in distributed data centers. Future Generation Computer Systems, 153 (??):360–374, April 2024. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167739X23004685.

Akachar:2021:APF

[AOF21] Elyazid Akachar, Brahim Ouhbi, and Bouchra Frikh. ACSIMCD: a 2-phase framework for detecting meaningful communities in dynamic social networks. Future Generation Computer Systems, 125(??):399–420, December 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2100251X.

Akyildiz:2024:REM

[AOKÖ24] Oğuzhan Akyıldız, Feyza Yıldırım Okay, İbrahim Kök, and Suat Özdemir. Road to efficiency: Mobilitydriven joint task offloading and resource utilization protocol for connected vehicle networks. Future Generation Computer Systems, 156(??):157–167, July 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X24000372.

Al-Obeidat:2020:IMT

[AOSA20a] Feras Al-Obeidat, Bruce Spencer, and May Al Taei. Identifying major tasks and minor tasks within online reviews. Future Generation Computer Systems, 110(??):413–421, September 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17319015.

Al-Obeidat:2020:CAF

[AOSA20b] Feras Al-Obeidat, Bruce Spencer, and Omar Alfandi. Consistently accurate forecasts of temperature within buildings from sensor data using ridge and lasso regression. Future Generation Computer Systems, 110(??):382–392, September 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17327541.

Asikis:2020:OPU

[AP20] Thomas Asikis and Evangelos Pournaras. Optimization of privacy-utility tradeoffs under informational self-determination. Future Generation Computer Systems, 109(??):488–499, August 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17322252.

Antal:2020:SSA

[APC^{+}20] Marcel Antal, Claudia Pop, Tudor Cioara, Ionut Anghel, Ioan Salomie, and Florin Pop. A system of systems approach for data centers optimization and integration into smart energy grids. Future Generation Computer Systems, 105(??):948–963, April 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17310129.

Ali:2020:CRA

[AQN^{+}20] Mudassar Ali, Saad Qaisar, Muhammad Naeem, Shahid Mumtaz, and Joel J. P. C. Rodrigues. Combinatorial resource allocation in D2D assisted heterogeneous relay networks. Future Generation Computer Systems, 107(??):956–964, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17307112.

Atzeni:2020:MDS

[AR20] Mattia Atzeni and Diego Reforgiato Recupero. Multidomain sentiment analysis with mimicked and polarized word embeddings for human-robot interaction. Future Generation Computer Systems, 110(??):984–999, September 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19309719.

Azad:2022:DNF

[ARA^{+}22] Muhammad Ajmal Azad, Farhan Riaz, Anum Aftab, Syed Khurram Jah Rizvi, Junaid Arshad, and Hany F. Atlam. DEEPSEL: a novel feature selection for early identification of malware in mobile applications. Future Generation Computer Systems, 129(??):54– 63, April 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21004258.

See corrigendum [ARA^{+}23].

Azad:2023:CDN

[ARA^{+}23] Muhammad Ajmal Azad, Farhan Riaz, Anum Aftab, Syed Khurram Jah Rizvi, Junaid Arshad, and Hany F. Atlam. Corrigendum to “DEEPSEL: a novel feature selection for early identification of malware in mobile applications” [Future Gener. Comput. Syst. 129 (2022) 54–63]. Future Generation Computer Systems, 143(??):422, June 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23000675.

See [ARA^{+}22].

Al-Ruithe:2020:DEF

[ARB20] Majid Al-Ruithe and Elhadj Benkhelifa. Determining the enabling factors for implementing cloud data governance in the Saudi public sector by structural equation modelling. Future Generation Computer Systems, 107(??):1061–1076, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17321489.

Abar:2020:FCF

[ARbL^{+}20] Tasnim Abar, Abderrezak Rachedi, Asma ben Letaifa, Philippe Fabian, and Sadok el Asmi. FellowMe cache: Fog computing approach to enhance (QoE) in Internet of Vehicles. Future Generation Computer Systems, 113(??):170–182, December 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19328870.

Adhinugraha:2020:ITI

[ARHT20] Kiki Adhinugraha, Wenny Rahayu, Takahiro Hara, and David Taniar. On Internet-of-Things (IoT) gateway coverage expansion. Future Generation Computer Systems, 107 (??):578–587, June 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0167739X19323064.

Adraoui:2022:MCB

[ARIB22] Meriem Adraoui, Asmaâ Retbi, Mohammed Khalidi Idrissi, and Samir Bennani. Maximal cliques based method for detecting and evaluating learning communities in social networks. Future Generation Computer Systems, 126(??):1–14, January 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2100296X.

Al-rimy:2021:RCG

[ArMA^{+}21] Bander Ali Saleh Alrimy, Mohd Aizaini Maarof, Mamoun Alazab, Syed Zainudeen Mohd Shaid, Fuad A. Ghaleb, Abdulmohsen Almalawi, Abdullah Marish Ali, and Tawfik Al-Hadhrami. Redundancy coefficient gradual up-weighting-based mutual information feature selection technique for cryptoransomware early detection. Future Generation Computer Systems, 115(??):641–658, February 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20329794.

Aloqaily:2020:MSR

[ASA^{+}20] Moayad Aloqaily, Haythem Bany Salameh, Ismaeel Al Ridhawi, Khalaf Batieha, and Jalel Ben Othman. A multistage resource-constrained spectrum access mechanism for cognitive radio IoT networks: Time-spectrum block utilization. Future Generation Computer Systems, 110(??):254–266, September 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19324860.

Agarwal:2023:WGA

[ASA23] Neha Agarwal, Geeta Sikka, and Lalit Kumar Awasthi. WGSDMM+GA: a genetic algorithm-based service clustering methodology assimilating Dirichlet multinomial mixture model with word embedding. Future Generation Computer Systems, 145(??):254–266, August 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X2300105X.

Ali:2020:UUA

[ASAM20] Reem Y. Ali, Shashi Shekhar, Shounak Athavale, and Eric Marsman. ULAMA: a utilization-aware matching approach for robust on-demand spatial service brokers. Future Generation Computer Systems, 108(??):1030–1048, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17312177.

Al-Sharif:2020:LFS

[ASASA^{+}20] Ziad A. Al-Sharif, Mohammed I. Al-Saleh, Luay M. Alawneh, Yaser I. Jararweh, and Brij Gupta. Live forensics of software attacks on cyber-physical systems. Future Generation Computer Systems, 108(??): 1217–1229, July 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17330534.

Ali:2023:APM

[ASB^{+}23] Ghazanfar Ali, Mert Side, Sridutt Bhalachandra, Nicholas J. Wright, and Yong Chen. An automated and portable method for selecting an optimal GPU frequency. Future Generation Computer Systems, 149(??):71– 88, December 2023. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23002613.

Alemneh:2020:TWT

[ASBT20] Esubalew Alemneh, Sidi-Mohammed Senouci, Philippe Brunet, and Tesfa Tegegne. A two-way trust management system for fog computing. Future Generation Computer Systems, 106(??):206–220, May 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167739X19316437.

Al-Sadi:2023:LUL

[ASDLS23] Mohammed Al-Sadi, Roberto Di Pietro, Flavio Lombardi, and Matteo Signorini. LENTO: Unpredictable Latency-based continuous authEntication for Network inTensive IoT envirOnments. Future Generation Computer Systems, 139(??):151–166, February 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22003077.

Alsamhi:2023:SFL

[ASH^{+}23] Saeed Hamood Alsamhi, Alexey V. Shvetsov, Ammar Hawbani, Svetlana V. Shvetsova, Santosh Kumar, and Liang Zhao. Survey on Federated Learning enabling indoor navigation for Industry 4.0 in B5G. Future Generation Computer Systems, 148(??):250–265, November 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23002169.

Al-Sayed:2020:ICS

[ASHO20] Mustafa M. Al-Sayed, Hesham A. Hassan, and Fatma A. Omara. An intelligent cloud service discovery framework. Future Generation Computer Systems, 106(??):438–466, May 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19322630.

Ahmed:2022:RCA

[ASL22] Usman Ahmed, Gautam Srivastava, and Jerry Chun-Wei Lin. Reliable customer analysis using federated learning and exploring deep-attention edge intelligence. Future Generation Computer Systems, 127(??):70–79, February 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21003368.

Alagha:2022:TLU

[ASM^{+}22] Ahmed Alagha, Shakti Singh, Rabeb Mizouni, Jamal Bentahar, and Hadi Otrok. Target localization using multi-agent deep reinforcement learning with proximal policy optimization. Future Generation Computer Systems, 136(??):342–357, November 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22002266.

Al-Saadi:2021:CWA

[ASPG^{+}21] Aymen Al-Saadi, Ioannis Paraskevakos, Bento Collares Gonçalves, Heather J. Lynch, Shantenu Jha, and Matteo Turilli. Comparing workflow application designs for high resolution satellite image analysis. Future Generation Computer Systems, 124(??):315–329, November 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21001448.

Aghili:2022:PMA

[ASSG22] Seyed Farhad Aghili, Mahdi Sedaghat, Dave Singelée, and Maanak Gupta. MLS-ABAC: Efficient multi-level security attribute-based access control scheme. Future Generation Computer Systems, 131(??):75–90, June 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22000115.

Ahmed:2022:EEA

[ASYL22] Usman Ahmed, Gautam Srivastava, Unil Yun, and Jerry Chun-Wei Lin. EANDC: an explainable attention network based deep adaptive clustering model for mental health treatment. Future Generation Computer Systems, 130(??): 106–113, May 2022. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21004891.

Al-Turjman:2020:ISB

[AT20]  Fadi Al-Turjman.  In-

telligence and security in big 5G-oriented IoNT: an overview. Future Generation Computer Systems, 102(??):357–368, January 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19301074.

Aslanpour:2024:LBH

[ATC^{+}24] Mohammad Sadegh Aslanpour, Adel N. Toosi, Muhammad Aamir Cheema, Mohan Baruwal Chhetri, and Mohsen Amini Salehi. Load balancing for heterogeneous serverless edge computing: a performancedriven and empirical approach. Future Generation Computer Systems, 154(??):266–280, May 2024. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167739X24000207.

Awais:2022:NSE

[ATK^{+}22] Muhammad Awais, Shahzaib Tahir, Fawad Khan, Hasan Tahir, Ruhma Tahir, Rabia Latif, and Mir Yasir Umair. A novel searchable encryption scheme to reduce the access pattern leakage. Future Generation Computer Systems, 133(??):338–350, August 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22001066.

Amin:2020:SMD

[ATT^{+}20] Muhammad Amin, Tamleek Ali Tanveer, Mohammad Tehseen, Murad Khan, Fakhri Alam Khan, and Sajid Anwar. Static malware detection and attribution in Android bytecode through an end-to-end deep system. Future Generation Computer Systems, 102(??):112–126, January 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19308490.

Alboaneen:2021:MMJ

[ATZP21] Dabiah Alboaneen, Hugo Tianfield, Yan Zhang, and Bernardi Pranggono. A metaheuristic method for joint task scheduling and virtual machine placement in cloud data centers. Future Generation Computer Systems, 115(??):201–212, February 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20310967.

Ahmad:2022:ITE

[AUJW22] Shabir Ahmad, Sabina Umirzakova, Faisal Jamil, and Taeg Keun Whangbo. Internet-of-Things-enabled serious games: a comprehensive survey. Future Generation Computer Systems, 136(??):67–83, November 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22001972.

A:2023:BBT

[AVK^{+}23] Sasikumar A., Subramaniyaswamy Vairavasundaram, Ketan Kotecha, Indragandhi V., Logesh Ravi, Ganeshsree Selvachandran, and Ajith Abraham. Blockchainbased trust mechanism for digital twin empowered industrial Internet of Things. Future Generation Computer Systems, 141(??):16– 27, April 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22003636.

Al-Wesabi:2023:POA

[AWMM^{+}23] Fahd N. Al-Wesabi, Hanan Abdullah Mengash, Radwa Marzouk, Nuha Alruwais, Randa Allafi, Rana Alabdan, Meshal Alharbi, and Deepak Gupta. Pelican Optimization Algorithm with Federated Learning Driven Attack Detection model in Internet of Things environment. Future Generation Computer Systems, 148(??):118–127, November 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23002121.

Abdullah:2023:HCC

[AYA^{+}23] Muhammed Amin Abdullah, Yongbin Yu, Kwabena Adu, Yakubu Imrana, Xiangxiang Wang, and Jingye Cai. HCL-classifier: CNN and LSTM based hybrid malware classifier for Internet of Things (IoT). Future Generation Computer Systems, 142(??):41– 58, May 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22004381.

Aryan:2022:SSR

[AYB^{+}22] Ramtin Aryan, Anis Yazidi, Frode Brattensborg, Øivind Kure, and Paal Einar Engelstad. SDN Spotlight: a real-time OpenFlow troubleshooting framework. Future Generation Computer Systems, 133(??):364–377, August 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22000899.

Ahmed:2020:HDI

[AYHA20] Hager Ahmed, Eman M. G.

Younis, Abdeltawab Hendawi, and Abdelmgeid A. Ali. Heart disease identification from patients’ social posts, machine learning solution on Spark. Future Generation Computer Systems, 111(??):714–722, October 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19315523.

An:2020:LLP

[AYY^{+}20] Dou An, Qingyu Yang, Wei Yu, Donghe Li, and Wei Zhao. LoPrO: Location privacy-preserving online auction scheme for electric vehicles joint bidding and charging. Future Generation Computer Systems, 107(??):394–407, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19316188.

Abdulqadder:2023:DBS

[AZA23] Ihsan H. Abdulqadder, Deqing Zou, and Israa T. Aziz. The DAG blockchain: a secure edge assisted honeypot for attack detection and multi-controller based load balancing in SDN 5G. Future Generation Computer Systems, 141(??):339–354, April 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22003697.

Benitez-Andrades:2020:SNA

[BAGRB^{+}20] José Alberto Benítez-Andrades, Isaías García-Rodríguez, Carmen Benavides, Héctor Alaiz-Moretón, and Alejandro Rodríguez-González. Social network analysis for personalized characterization and risk assessment of alcohol use disorders in adolescents using semantic technologies. Future Generation Computer Systems, 106(??):154–170, May 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19316796.

Bal:2024:PBR

[BAİP24] Melis İlayda Bal, Hüseyin Aydın, Cem İyigün, and Faruk Polat. Potentialbased reward shaping using state–space segmentation for efficiency in reinforcement learning. Future Generation Computer Systems, 157(??):469–484, August 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24001262.

Baniata:2022:DDO

[BAK22] Hamza Baniata, Ahmad Anaqreh, and Attila Kertesz. DONS: Dynamic optimized neighbor selection for smart blockchain networks. Future Generation Computer Systems, 130(??):75–90, May 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2100491X.

Farah:2024:SBT

[BAM^{+}24] Mohamed Ben Farah, Yussuf Ahmed, Haithem Mahmoud, Syed Attique Shah, M. Omar Al-kadri, Sandy Taramonli, Xavier Bellekens, Raouf Abozariba, Moad Idrissi, and Adel Aneiba. A survey on blockchain technology in the maritime industry: Challenges and future perspectives. Future Generation Computer Systems, 157(??):618–637, August 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24001213.

BenKhalifa:2020:NMV

[BAMR20] Anouar Ben Khalifa, Ihsen Alouani, Mohamed Ali Mahjoub, and Atika Rivenq. A novel multi-view pedestrian detection database for collaborative intelligent transportation systems. Future Generation Computer Systems, 113(??):506–527, December 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20300340.

Bordel:2021:DCE

[BAR21] Borja Bordel, Ramón Alcarria, and Tomás Robles. Denial of chain: Evaluation and prediction of a novel cyberattack in blockchainsupported systems. Future Generation Computer Systems, 116(??):426–439, March 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20330375.

Blanquer:2020:FSC

[BBB^{+}20] Ignacio Blanquer, Francisco Brasileiro, Andrey Brito, Amanda Calatrava, André Carvalho, Christof Fetzer, Flavio Figueiredo, Ronny Petterson Guimarães, Leandro Marinho, Wagner Meira, Altigran Silva, Ángel Alberich-Bayarri, Eduardo Camacho-Ramos, Ana Jimenez-Pastor, Antonio Luiz L. Ribeiro, Bruno Ramos Nascimento, and Fábio Silva. Federated and secure cloud services for building medical image classifiers on an intercontinental infrastructure. Future Generation Computer Systems, 110(??):119–134, September 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19318047.

Bolettieri:2022:TEE

[BBB22] Simone Bolettieri, Dinh Thai Bui, and Raffaele Bruno. Towards end-to-end application slicing in multiaccess edge computing systems: Architecture discussion and proof-of-concept. Future Generation Computer Systems, 136(??):110–127, November 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22001984.

Brogi:2024:FPP

[BBB^{+}24] F. Brogi, S. Bnà, G. Boga, G. Amati, T. Esposti Ongaro, and M. Cerminara. On floating point precision in computational fluid dynamics using OpenFOAM. Future Generation Computer Systems, 152(??):1– 16, March 2024. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23003813.

Bellman:2021:SISa

[BBD^{+}21] Kirstie Bellman, Jean Botev, Ada Diaconescu, Lukas Esterle, Christian Gruhl, Christopher Landauer, Peter R. Lewis, Phyllis R. Nelson, Evangelos Pournaras, Anthony Stein, and Sven Tomforde. Self-improving system integration: Mastering continuous change. Future Generation Computer Systems, 117(??):29– 46, April 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20330430.

Bellavista:2024:EMS

[BBF^{+}24] Paolo Bellavista, Nicola Bicocchi, Mattia Fogli, Carlo Giannelli, Marco Mamei, and Marco Picone. Exploiting microservices and serverless for Digital Twins in the cloud-to-edge continuum. Future Generation Computer Systems, 157(??):275–287, August 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24001249.

Briem:2020:IPT

[BBM^{+}20] Lars Briem, H. Sebastian Buck, Nicolai Mallig, Peter Vortisch, Ben Strasser, Dorothea Wagner, and Tobias Zündorf. Integrating public transport into mobiTopp. Future Generation Computer Systems, 107(??):1089–1096, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17320071.

Balti:2022:MAU

[BBM^{+}22] Hanen Balti, Ali Ben Abbes, Nedra Mellouli, Imed Riadh Farah, Yanfang Sang, and Myriam Lamolle. Multidimensional architecture using a massive and heterogeneous data: Application to drought monitoring. Future Generation Computer Systems, 136(??):1–14, November 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22001753.

Bez:2020:ARS

[BBN^{+}20] Jean Luca Bez, Francieli Zanon Boito, Ramon Nou, Alberto Miranda, Toni Cortes, and Philippe O. A. Navaux. Adaptive request scheduling for the I/O forwarding layer using reinforcement learning. Future Generation Computer Systems, 112(??):1156–1169, November 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20300856.

Borgia:2022:RDD

[BBP22] Eleonora Borgia, Raffaele Bruno, and Andrea Passarella. Reliable data delivery in ICN-IoT environments. Future Generation Computer Systems, 134(??):271–286, September 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22001273.

Barbareschi:2021:ESI

[BBSB21] Mario Barbareschi, Alberto Bosio, Lukas Sekanina, and Claus Braun. Editorial: Special issue on advancing on approximate computing: Methodologies, architectures and algorithms. Future Generation Computer Systems, 124(??):54– 55, November 2021. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21001771.

Behrad:2020:NSA

[BBTC20] Shanay Behrad, Emmanuel Bertin, Stéphane Tuffin, and Noel Crespi. A new scalable authentication and access control mechanism for 5g-based IoT. Future Generation Computer Systems, 108(??):46– 61, July 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19310143.

Berrada:2020:BUA

[BCB^{+}20] Ghita Berrada, James Cheney, Sidahmed Benabderrahmane, William Maxwell, Himan Mookherjee, Alec Theriault, and Ryan Wright. A baseline for unsupervised advanced persistent threat detection in system-level provenance. Future Generation Computer Systems, 108(??): 401–413, July 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19320448.

Baldin:2022:INS

[BCC^{+}22] Ilya Baldin, Jeff Chase, Jonathan Crabtree, Thomas Nechyba, Laura Christopherson, Michael Stealey, Charley Kneifel, Victor Orlikowski, Rob Carter, Erik Scott, Akio Sone, and Don Sizemore. ImPACT: a networked service architecture for safe sharing of restricted data. Future Generation Computer Systems, 129(??):269–285, April 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21004702.

Bartoletti:2020:DPS

[BCCS20] Massimo Bartoletti, Salvatore Carta, Tiziana Cimoli, and Roberto Saia. Dissecting Ponzi schemes on Ethereum: Identification, analysis, and impact. Future Generation Computer Systems, 102(??):259–277, January 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18301407.

Bencke:2020:ACS

[BCM20] Luciana Bencke, Cristian Cechinel, and Roberto Munoz. Automated classification of social network messages into Smart Cities dimensions. Future Generation Computer Systems, 109(??):218–237, August 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19303000.

Burger:2020:EIT

[BCSS20] Alwyn Burger, Christopher Cichiwskyj, Stephan Schmeißer, and Gregor Schiele. The elastic Internet of Things — a platform for self-integrating and self-adaptive IoT-systems with support for embedded adaptive hardware. Future Generation Computer Systems, 113(??):607–619, December 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20304763.

Barnawi:2021:AIE

[BCT^{+}21] Ahmed Barnawi, Prateek Chhikara, Rajkumar Tekchandani, Neeraj Kumar, and Bander Alzahrani. Artificial intelligence-enabled Internet of Things-based system for COVID-19 screening using aerial thermal imaging. Future Generation Computer Systems, 124(??):119–132, November 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21001692.

Bonnin:2022:GEF

[BDF^{+}22] Geoffray Bonnin, Danilo Dessì, Gianni Fenu, Martin Hlosta, Mirko Marras, and Harald Sack. Guest editorial of the FGCS special issue on advances in intelligent systems for online education. Future Generation Computer Systems, 127(??):331–333, February 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X2100368X.

Buscaldi:2022:SIS

[BDFR22] Davide Buscaldi, Mauro Dragoni, Flavius Frasincar, and Diego Reforgiato Recupero. Special issue on senti-mental health: Future generation sentiment analysis systems. Future Generation Computer Systems, 129(??):170–173, April 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21004775.

Bianchini:2023:BDE

[BDG23] Devis Bianchini, Valeria De Antonellis, and Massimiliano Garda. A big data exploration approach to exploit in-vehicle data for smart road maintenance. Future Generation Computer Systems, 149(??):701–716, December 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23003060.

Borrell:2020:HCG

[BDGG^{+}20] R. Borrell, D. Dosimont, M. Garcia-Gasulla, G. Houzeaux, O. Lehmkuhl, V. Mehta, H. Owen, M. Vázquez, and G. Oyarzun. Heterogeneous CPU/GPU coexecution of CFD simulations on the POWER9 architecture: Application to airplane aerodynamics. Future Generation Computer Systems, 107(??):31– 48, June 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1930994X.

Brings:2020:SMV

[BDK^{+}20] Jennifer Brings, Marian Daun, Kevin Keller, Patricia Aluko Obe, and Thorsten Weyer. A systematic map on verification and validation of emergent behavior in software engineering research. Future Generation Computer Systems, 112(??):1010–1037, November 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19333606.

Bailey:2020:MCR

[BdL20] Christopher Bailey and Rogério de Lemos. Malicious changeload for the resilience evaluation of selfadaptive authorisation infrastructures. Future Generation Computer Systems, 113(??):113–131, December 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19319326.

Bernabe:2020:AER

[BDM^{+}20] Jorge Bernal Bernabe, Martin David, Rafael Torres Moreno, Javier Presa Cordero, Sébastien Bahloul, and Antonio Skarmeta. ARIES: Evaluation of a reliable and privacypreserving European identity management framework. Future Generation Computer Systems, 102(??):409–425, January 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1930843X.

Bellman:2021:SISb

[BDT21] Kirstie L. Bellman, Ada Diaconescu, and Sven Tomforde. Special issue on “selfimproving self integration”. Future Generation Computer Systems, 119(??): 136–139, June 2021. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21000625.

Baccour:2020:PPV

[BEB^{+}20] Emna Baccour, Aiman Erbad, Kashif Bilal, Amr Mohamed, and Mohsen Guizani. PCCP: Proactive video chunks caching and processing in edge networks. Future Generation Computer Systems, 105 (??):44–60, April 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18331674.

Breitfuss:2021:REK

[BEKF21] Arno Breitfuss, Karen Errou, Anelia Kurteva, and Anna Fensel. Representing emotions with knowledge graphs for movie recommendations. Future Generation Computer Systems, 125(??):715–725, December 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21001953.

Benmoussa:2020:MMS

[BeKTK^{+}20] Ahmed Benmoussa, Abdou el Karim Tahari, Chaker Abdelaziz Kerrache, Nasreddine Lagraa, Abderrahmane Lakas, Rasheed Hussain, and Farhan Ahmad. MSIDN: Mitigation of sophisticated interest floodingbased DDoS attacks in named data networking. Future Generation Computer Systems, 107(??): 293–306, June 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19328729.

Barnes:2020:BGR

[BEL20] Chloe M. Barnes, Anikó Ekárt, and Peter R. Lewis. Beyond goal-rationality: Traditional action can reduce volatility in socially situated agents. Future Generation Computer Systems, 113(??):579–596, December 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20304714.

Baccour:2020:ROR

[BEM^{+}20] Emna Baccour, Aiman Erbad, Amr Mohamed, Fatima Haouari, Mohsen Guizani, and Mounir Hamdi. RL-OPRA: Reinforcement learning for online and proactive resource allocation of crowdsourced live videos. Future Generation Computer Systems, 112(??):982–995, November 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20306269.

Baccour:2024:RLB

[BEM^{+}24] Emna Baccour, Aiman Erbad, Amr Mohamed, Mounir Hamdi, and Mohsen Guizani. Reinforcement learning-based dynamic pruning for distributed inference via explainable AI in healthcare IoT systems. Future Generation Computer Systems, 155(??):1– 17, June 2024. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X24000232.

BenAli:2023:AAD

[Ben23a] Yamina Mohamed Ben Ali. Adversarial attacks on deep learning networks in image classification based on Smell Bees Optimization Algorithm. Future Generation Computer Systems, 140(??):185–195, March 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22003375.

Ali:2023:OPX

[Ben23b] Yamina Mohamed Ben Ali. One-pixel and X-pixel adversarial attacks based on smell bees optimization algorithm. Future Generation Computer Systems, 149(??):562–576, December 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23002856.

Benmeziane:2024:GOS

[BEON24] Hadjer Benmeziane, Kaoutar El Maghraoui, Hamza Ouarnoughi, and Smail Niar. Grassroots operator search for model edge adaptation using mathematical search space. Future Generation Computer Systems, 157(??):29–40, August 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24001018.

Bellomarini:2022:DSV

[BFG^{+}22] Luigi Bellomarini, Ruslan R. Fayzrakhmanov, Georg Gottlob, Andrey Kravchenko, Eleonora Laurenza, Yavor Nenov, Stéphane Reissfelder, Emanuel Sallinger, Evgeny Sherkhonov, Sahar Vahdati, and Lianlong Wu. Data science with Vadalog: Knowledge graphs with machine learning and reasoning in practice. Future Generation Computer Systems, 129(??):407–422, April 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21004179.

Bistarelli:2023:CFA

[BFM23] Stefano Bistarelli, Francesco Faloci, and Paolo Mori. *-chain: a framework for automating the modeling of blockchain based supply chain tracing systems. Future Generation Computer Systems, 149(??):679–700, December 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23002625.

Bian:2020:BAE

[BGCL20] Weixin Bian, Prosanta Gope, Yongqiang Cheng, and Qingde Li. Bio-AKA: an efficient fingerprint based two factor user authentication and key agreement scheme. Future Generation Computer Systems, 109(??):45–55, August 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19332467.

Brown:2022:RSP

[BGMK22] Alexander Brown, Saurabh Garg, James Montgomery, and Ujjwal KC. Resource scheduling and provisioning for processing of dynamic stream workflows under latency constraints. Future Generation Computer Systems, 131(??):166–182, June 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22000280.

Barba-Gonzalez:2020:DFI

[BGNBH^{+}20] Cristóbal Barba-González, Antonio J. Nebro, Antonio Benítez-Hidalgo, José García-Nieto, and José F. Aldana-Montes. On the design of a framework integrating an optimization engine with streaming technologies. Future Generation Computer Systems, 107(??):538–550, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19315699.

Bernardi:2020:AKE

[BGNM20] Simona Bernardi, Ugo Gentile, Roberto Nardone, and Stefano Marrone. Advancements in knowledge elicitation for computerbased critical systems. Future Generation Computer Systems, 110(??):311–313, September 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20309183.

Barika:2020:CES

[BGR20] Mutaz Barika, Saurabh Garg, and Rajiv Ranjan. Cost effective stream workflow scheduling to handle application structural changes. Future Generation Computer Systems, 112(??):348–361, November 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20302788.

Bi:2021:HAI

[BH21] Zhuo Bi and Wenju Huang. Human action identification by a quality-guided fusion of multi-model feature. Future Generation Computer Systems, 116(??):13– 21, March 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20329885.

Huang:2021:HPA

[bHFF^{+}21] Zhi bin Huang, Guang-Tao Fu, Tian-Hao Fa, Dan-Yang Dong, Peng Bai, and Chen Xiao. High performance ant colony system based on GPU warp specialization with a static– dynamic balanced candidate set strategy. Future Generation Computer Systems, 125(??):136–150, December 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21002363.

Bukhari:2022:FND

[BHH22] Afnan Bukhari, Farookh Khadeer Hussain, and Omar K. Hussain. Fog node discovery and selection: a systematic literature review. Future Generation Computer Systems, 135(??):114–128, October 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22001625.

Bhandari:2020:SLS

[BHL^{+}20] Shweta Bhandari, Frederic Herbreteau, Vijay Laxmi, Akka Zemmari, Manoj Singh Gaur, and Partha S. Roop. Sneak-Leak+: Large-scale klepto apps analysis. Future Generation Computer Systems, 109(??):593–603, August 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17321672.

Bian:2021:AME

[BHL^{+}21] Haodong Bian, Jianqiang Huang, Lingbin Liu, Dongqiang Huang, and Xiaoying Wang. ALBUS: a method for efficiently processing SpMV using SIMD and load balancing. Future Generation Computer Systems, 116(??):371–392, March 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X2033020X.

Boyer:2022:DSD

[BHSH22] Alexandre F. Boyer, Christophe Haen, Federico Stagni, and David R. C. Hill. DIRAC Site Director: Improving pilot-job provisioning on grid resources. Future Generation Computer Systems, 133(??):23–38, August 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22000772.

Bai:2022:PAE

[BJ22] Xiaojun Bai and Shunfu Jin. Performance analysis of an energy-saving strategy in cloud data centers based on a MMAP[K]/M[K]/N_{1}+ N_{2} non-preemptive priority queue. Future Generation Computer Systems, 136(??):205–220, November 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22002035.

Basford:2020:PAS

[BJP^{+}20] Philip J. Basford, Steven J. Johnston, Colin S. Perkins, Tony Garnock-Jones, Fung Po Tso, Dimitrios Pezaros, Robert D. Mullins, Eiko Yoneki, Jeremy Singer, and Simon J. Cox. Performance analysis of single board computer clusters. Future Generation Computer Systems, 102(??):278–291, January 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1833142X.

Bradbury:2022:IMT

[BJW22] Matthew Bradbury, Arshad Jhumka, and Tim Watson. Information management for trust computation on resource-constrained IoT devices. Future Generation Computer Systems, 135(??):348–363, October 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22001698.

Bawa:2020:MIF

[BK20] Vivek Singh Bawa and Vinay Kumar. Mutually independent feature flow: an approach to produce fixed complexity latent space for improved performance and decreased parameter count. Future Generation Computer Systems, 110(??):1067–1078, September 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19306818.

BenHalima:2020:TCO

[BKG^{+}20] Rania Ben Halima, Slim Kallel, Walid Gaaloul, Zakaria Maamar, and Mohamed Jmaiel. Toward a correct and optimal timeaware cloud resource allocation to business processes. Future Generation Computer Systems, 112(??):751–766, November 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19333679.

Belguith:2020:PVP

[BKHD20] Sana Belguith, Nesrine Kaaniche, Mohammad Hammoudeh, and Tooska Dargahi. PROUD: Verifiable privacy-preserving outsourced attribute based SignCryption supporting access policy update for cloud assisted IoT applications. Future Generation Computer Systems, 111(??):899–918, October 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1930130X.

Bugshan:2022:RBF

[BKM^{+}22] Neda Bugshan, Ibrahim Khalil, Nour Moustafa, Mahathir Almashor, and Alsharif Abuadbba. Radial basis function network with differential privacy. Future Generation Computer Systems, 127(??):473–486, February 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21003599.

Bourelos:2020:SHC

[BKV^{+}20] P. Bourelos, G. Kousiouris, O. Voutyras, A. Marinakis, T. Varvarigou, and V. Moulos. Smart heating in collaborative and reasoningenabled housing units. Future Generation Computer Systems, 109(??):644–656, August 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18306009.

Baranwal:2022:BBA

[BKV22] Gaurav Baranwal, Dinesh Kumar, and Deo Prakash Vidyarthi. BARA: a blockchainaided auction-based resource allocation in edge computing enabled industrial internet of things. Future Generation Computer Systems, 135(??):333–347, October 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22001728.

Barron-Lugo:2023:XCA

[BLGCLA^{+}23] J. Armando Barron-Lugo, J. L. Gonzalez-Compean, Ivan Lopez-Arevalo, Jesus Carretero, and Jose L. Martinez-Rodriguez. Xel: a cloud-agnostic data platform for the design-driven building of high-availability data science services. Future Generation Computer Systems, 145(??):87–103, August 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2300095X.

Bai:2024:DIF

[BLH^{+}24] Jiaru Bai, Kok Foong Lee, Markus Hofmeister, Sebastian Mosbach, Jethro Akroyd, and Markus Kraft. A derived information framework for a dynamic knowledge graph and its application to smart cities. Future Generation Computer Systems, 152(??): 112–126, March 2024. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23003825.

Beaumont:2020:PAO

[BLMT20] Olivier Beaumont, Thomas Lambert, Loris Marchal, and Bastien Thomas. Performance analysis and optimality results for datalocality aware tasks scheduling with replicated inputs. Future Generation Computer Systems, 111(??):582–598, October 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18325469.

Bader:2024:LLP

[BLT^{+}24] Jonathan Bader, Fabian Lehmann, Lauritz Thamsen, Ulf Leser, and Odej Kao. Lotaru: Locally predicting workflow task runtimes for resource management on heterogeneous infrastructures. Future Generation Computer Systems, 150(??):171–185, January 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23003229.

Bellendorf:2020:COP

[BM20] Julian Bellendorf and Zoltán Ádám Mann. Classification of optimization problems in fog computing. Future Generation Computer Systems, 107(??):158–176, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19323568.

Buchaca:2020:SSM

[BMBC20] David Buchaca, Joan Marcual, Josep LLuis Berral, and David Carrera. Sequenceto-sequence models for workload interference prediction on batch processing datacenters. Future Generation Computer Systems, 110(??):155–166, September 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19310921.

Bataineh:2020:TMP

[BMBE20] Ahmed Saleh Bataineh, Rabeb Mizouni, Jamal Bentahar, and May El Barachi. Toward monetizing personal data: a twosided market analysis. Future Generation Computer Systems, 111(??):435–459, October 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19301414.

Bayat:2021:FAP

[BMD^{+}21] Niloofar Bayat, Kunal Mahajan, Sam Denton, Vishal Misra, and Dan Rubenstein. Down for failure: Active power status monitoring. Future Generation Computer Systems, 125(??):629–640, December 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21002508.

Boudieb:2024:MIS

[BMM^{+}24] Wassim Boudieb, Abdelhamid Malki, Mimoun Malki, Ahmed Badawy, and Mahmoud Barhamgi. Microservice instances selection and load balancing in fog computing using deep reinforcement learning approach. Future Generation Computer Systems, 156 (??):77–94, July 2024. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24000815.

Balsamo:2020:CNC

[BMS20] Simonetta Balsamo, Andrea Marin, and Ivan Stojic. Computation of the normalising constant for product-form models of distributed systems with synchronisation. Future Generation Computer Systems, 111(??):475–490, October 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19310313.

Briones:2021:AAT

[BMZdP21] Alan Briones, Adrià Mallorquí, Agustín Zaballos, and Ramon Martin de Pozuelo. Adaptive and aggressive transport protocol to provide QoS in cloud data exchange over long fat networks. Future Generation Computer Systems, 115(??):34–44, February 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1832497X.

Bilotta:2021:TFR

[BN21] Stefano Bilotta and Paolo Nesi. Traffic flow reconstruction by solving indeterminacy on traffic distribution at junctions. Future Generation Computer Systems, 114(??):649–660, January 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20308359.

Basiri:2021:AAB

[BNA^{+}21] Mohammad Ehsan Basiri, Shahla Nemati, Moloud Abdar, Erik Cambria, and U. Rajendra Acharya. ABCDM: an attentionbased bidirectional CNN-RNN deep model for sentiment analysis. Future Generation Computer Systems, 115(??):279–294, February 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20309195.

Bountakas:2022:EEK

[BNX22] Panagiotis Bountakas, Christoforos Ntantogian, and Christos Xenakis. EKnad: Exploit kits’ network activity detection. Future Generation Computer Systems, 134(??):219–235, September 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22001248.

Bo:2019:ASC

[Bo19] Lei Bo. AOFAS scores for curative effect analysis on arthroscopic treatment of subtalar ankle instability with osteochondral injury syndrome. Future Generation Computer Systems, 91(??):506–510, February 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/

                                        science/article/pii/S0167739X18318119.

See retraction notice [Bo20b].

Balta:2020:SFD

[BÖ20a] Musa Balta and Ibrahim Özçelik. A 3-stage fuzzydecision tree model for traffic signal optimization in urban city via a SDN based VANET architecture. Future Generation Computer Systems, 104(??):142–158, March 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19319028.

Bo:2020:RNA

[Bo20b] Lei Bo. Retraction notice to “AOFAS scores for curative effect analysis on arthroscopic treatment of subtalar ankle instability with osteochondral injury syndrome” [future gener. comput. syst. 91 (2019) 506– 510]. Future Generation Computer Systems, 107 (??):1144, June 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20306373.

See [Bo19].

Bernardo:2024:SQA

[BOD^{+}24] Samuel Bernardo, Pablo Orviz, Mario David, Jorge Gomes, David Arce, Diana Naranjo, Ignacio Blanquer, Isabel Campos, Germán Moltó, and Joao Pina. Software quality assurance as a service: Encompassing the quality assessment of software and services. Future Generation Computer Systems, 156(??):254–268, July 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24000955.

Baktir:2024:EEN

[BÖE24] Ahmet Cihat Baktır, Atay Özgövde, and Cem Ersoy. End-to-end network slicing for edge computing optimization. Future Generation Computer Systems, 157(??):516–528, August 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24000724.

Bellavista:2020:TSC

[BOL^{+}20] Paolo Bellavista, Kaoru Ota, Zhihan Lv, Irfan Mehmood, and Seungmin Rho. Towards smarter cities: Learning from Internet of Multimedia Thingsgenerated big data. Future Generation Computer Systems, 108(??):879–881, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19314864.

Bellatreche:2022:CRD

[BOM^{+}22] Ladjel Bellatreche, Carlos Ordonez, Dominique Méry, Matteo Golfarelli, and El Hassan Abdelwahed. The central role of data repositories and data models in data science and advanced analytics. Future Generation Computer Systems, 129(??):13– 17, April 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21004714.

Bal:2020:PDM

[BP20] Henri Bal and Arindam Pal. Parallel and distributed machine learning algorithms for scalable Big Data analytics. Future Generation Computer Systems, 108(??):1159–1161, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19318084.

Bauer:2024:GCD

[BPC^{+}24] André Bauer, Haochen Pan, Ryan Chard, Yadu Babuji, Josh Bryan, Devesh Tiwari, Ian Foster, and Kyle Chard. The Globus Compute Dataset: an open function-as-a-service dataset from the edge to the cloud. Future Generation Computer Systems, 153(??):558–574, April 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23004703.

Bayliss:2021:RVC

[BPCM21] Christopher Bayliss, Javier Panadero, Laura Calvet, and Joan Manuel Marquès. Reliability in volunteer computing micro-blogging services. Future Generation Computer Systems, 115(??):857–871, February 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2031102X.

Barcelona-Pons:2021:BPF

[BPGL21] Daniel Barcelona-Pons and Pedro García-López. Benchmarking parallelism in FaaS platforms. Future Generation Computer Systems, 124(??):268–284, November 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21001990.

Blanco:2020:USE

[BPLFRL20] Guillermo Blanco, Roi Pérez-López, Florentino Fdez-Riverola, and Anália Maria Garcia Lourenço. Understanding the social evolution of the Java community in Stack Overflow: a 10-year study of developer interactions. Future Generation Computer Systems, 105(??):446–454, April 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19311884.

Barreto:2023:PCC

[BPSP23] António Barreto, Hervé Paulino, João A. Silva, and Nuno Preguiça. PS-CRDTs: CRDTs in highly volatile environments. Future Generation Computer Systems, 141(??):755–767, April 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22004186.

Bozejko:2024:QAD

[BPUW24] Wojciech Bożejko, Jarosław Pempera, Mariusz Uchroński, and Mieczysław Wodecki. Quantum annealing-driven branch and bound for the single machine total weighted number of tardy jobs scheduling problem.

Future Generation Computer Systems, 155(??): 245–255, June 2024. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2400061X.

Barcelo:2022:RAO

[BQC22] Alex Barcelo, Anna Queralt, and Toni Cortes. Revisiting active object stores: Bringing data locality to the limit with NVM. Future Generation Computer Systems, 129(??):425–439, April 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21004210.

Barcelo:2023:EIP

[BQC23] Alex Barcelo, Anna Queralt, and Toni Cortes. Enhancing iteration performance on distributed taskbased workflows. Future Generation Computer Systems, 149(??):359–375, December 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23002911.

Bylykbashi:2020:FBD

[BQI^{+}20] Kevin Bylykbashi, Ermioni Qafzezi, Makoto Ikeda, Keita Matsuo, and Leonard Barolli. Fuzzy-based Driver Monitoring System (FDMS): Implementation of two intelligent FDMSs and a testbed for safe driving in VANETs. Future Generation Computer Systems, 105(??):665–674, April 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19322927.

Bouyer:2020:LFR

[BR20] Asgarali Bouyer and Hamid Roghani. LSMD: a fast and robust local community detection starting from low degree nodes in social networks. Future Generation Computer Systems, 113(??):41–57, December 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2030724X.

Banerjee:2024:TEQ

[BRK24] Sounak Banerjee, Sarbani Roy, and Sunirmal Khatua. Towards energy and QoS aware dynamic VM consolidation in a multiresource cloud. Future Generation Computer Systems, 157(??):376–391, August 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24001274.

Briffoteaux:2020:ECP

[BRM^{+}20] G. Briffoteaux, R. Ragonnet, M. Mezmaz, N. Melab, and D. Tuyttens. Evolution control for parallel ANN-assisted simulationbased optimization application to tuberculosis transmission control. Future Generation Computer Systems, 113(??):454–467, December 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19308635.

Bonnah:2020:DDS

[BS20] Ernest Bonnah and Ju Shiguang.

DecChain: a decentralized security approach in edge computing based on blockchain. Future Generation Computer Systems, 113(??):363–379, December 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19323933.

Barclay:2022:TSD

[BSB^{+}22] Iain Barclay, Chris Simpkin, Graham Bent, Tom La Porta, Declan Millar, Alun Preece, Ian Taylor, and Dinesh Verma. Trustable service discovery for highly dynamic decentralized workflows. Future Generation Computer Systems, 134(??):236–246, September 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22001169.

Baig:2020:ADE

[BSF^{+}20] Zubair A. Baig, Surasak Sanguanpong, Syed Naeem Firdous, Van Nhan Vo, Tri Gia Nguyen, and Chakchai So-In. Averaged dependence estimators for DoS attack detection in IoT networks. Future Generation Computer Systems, 102(??):198–209, January 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19300834.

BenElhadj:2021:DCD

[BSH^{+}21] Hadda Ben Elhadj, Farag Sallabi, Amira Henaien, Lamia Chaari, Khaled Shuaib, and Maryam Al Thawadi. Do-Care: a dynamic ontology reasoning based healthcare monitoring system. Future Generation Computer Systems, 118(??):417–431, May 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21000017.

Bai:2024:PUR

[BSH^{+}24] Shizhen Bai, Songlin Shi, Chunjia Han, Mu Yang, Brij B. Gupta, and Varsha Arya. Prioritizing user requirements for digital products using explainable artificial intelligence: a datadriven analysis on video conferencing apps. Future Generation Computer Systems, 158(??):167–182, September 2024. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24001663.

Bannour:2020:ADS

[BSM20] Fetia Bannour, Sami Souihi, and Abdelhamid Mellouk. Adaptive distributed SDN controllers: Application to content-centric delivery networks. Future Generation Computer Systems, 113(??):78–93, December 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19328316.

Benmakrelouf:2020:ABD

[BSOK^{+}20] Souhila Benmakrelouf, Cédric St-Onge, Nadjia Kara, Hanine Tout, Claes Edstrom, and Yves Lemieux. Abnormal behavior detection using resource level to service level metrics mapping in virtualized systems. Future Generation Computer Systems, 102(??):680–700, January 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19306600.

Bandara:2021:TTL

[BTF^{+}21] Eranga Bandara, Deepak Tosh, Peter Foytik, Sachin Shetty, Nalin Ranasinghe, and Kasun De Zoysa. Tikiri

— towards a lightweight blockchain for IoT. Future Generation Computer Systems, 119(??):154–165, June 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21000583.

Belchior:2022:HFT

[BVCH22] Rafael Belchior, André Vasconcelos, Miguel Correia, and Thomas Hardjono. Hermes: Faulttolerant middleware for blockchain interoperability. Future Generation Computer Systems, 129(??): 236–251, April 2022. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21004337.

Basanta-Val:2020:PRI

[BVFGSF20] P. Basanta-Val, N. Fernández-García, and L. Sánchez-Fernández. Predictable remote invocations for distributed stream processing. Future Generation Computer Systems, 107(??): 716–729, June 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17318368.

Bao:2020:RNC

[BWX20] Han Bao, Yijie Wang, and Fangliang Xu. Reducing network cost of data repair in erasure-coded crossdatacenter storage. Future Generation Computer Systems, 102(??):494–506, January 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19311422.

Botangen:2020:QAW

[BYH^{+}20] Khavee Agustus Botangen, Jian Yu, Yanbo Han, Quan Z. Sheng, and Jun Han. Quantifying the adaptability of workflowbased service compositions. Future Generation Computer Systems, 102(??):95– 111, January 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19301517.

Bai:2020:QEP

[BYR^{+}20] Shuangjie Bai, Geng Yang, Chunming Rong, Guoxiu Liu, and Hua Dai. QHSE: an efficient privacy-preserving scheme for blockchainbased transactions. Future Generation Computer Systems, 112(??):930–944, November 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19334508.

Bai:2021:DST

[BYW^{+}21] Lei Bai, Lina Yao, Xianzhi Wang, Can Li, and Xiang Zhang. Deep spatial– temporal sequence modeling for multi-step passenger demand prediction. Future Generation Computer Systems, 121(??):25–34, August 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21000832.

Ba:2023:CBB

[BZG23] Cheick Tidiane Ba, Matteo Zignani, and Sabrina Gaito. Cooperative behavior in blockchain-based complementary currency networks through time: the Sarafu case study. Future Generation Computer Systems, 148(??):266–279, November 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23002066.

Zhang:2023:SAR

[bZSC^{+}23] Li bo Zhang, Shuang Sun, Junxin Chen, Yue Teng, and Zhihan Lv. Selfadaptive reconstruction for compressed sensing based ECG acquisition in wireless body area network. Future Generation Computer Systems, 142(??):228–236, May 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22004174.

Cebe:2021:CEC

[CA21] Mumin Cebe and Kemal Akkaya. Communicationefficient certificate revocation management for advanced metering infrastructure and IoT integration. Future Generation Computer Systems, 115(??):267–278, February 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20304891.

Chehreghani:2020:SIP

[CABB20] Mostafa Haghir Chehreghani, Talel Abdessalem, Albert Bifet, and Meriem Bouzbila. Sampling informative patterns from large single networks. Future Generation Computer Systems, 106(??):653–658, May 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19303437.

Colonnelli:2022:DWJ

[CAC^{+}22] Iacopo Colonnelli, Marco Aldinucci, Barbara Cantalupo, Luca Padovani, Sergio Rabellino, Concetto Spampinato, Roberto Morelli, Rosario Di Carlo, Nicolò Magini, and Carlo Cavazzoni. Distributed workflows with Jupyter. Future Generation Computer Systems, 128(??): 282–298, March 2022. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21003976.

Chithaluru:2024:ECB

[CATD^{+}24] Premkumar Chithaluru, Fadi Al-Turjman, Raman Dugyala, Thompson Stephan, Manoj Kumar, and Jagjit Singh Dhatterwal. An enhanced consortium blockchain diversity mining technique for IoT metadata aggregation. Future Generation Computer Systems, 152(??):239–253, March 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23003941.

Cebrian:2020:SAV

[CBC^{+}20] Juan M. Cebrian, Adrián Barredo, Helena Caminal, Miquel Moretó, Marc Casas, and Mateo Valero. Semi-automatic validation of cycle-accurate simulation infrastructures: the case for gem5-x86. Future Generation Computer Systems, 112(??):832–847, November 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19304972.

Christie:2020:MAA

[CBN^{+}20] Marcus A. Christie, Anuj Bhandar, Supun Nakandala, Suresh Marru, Eroma Abeysinghe, Sudhakar Pamidighantam, and Marlon E. Pierce. Managing authentication and authorization in distributed science gateway middleware. Future Generation Computer Systems, 111(??):780–785, October 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18314729.

Chi:2021:DAM

[CC21] Po-Wen Chi and Yu-Lun Chang. Do not ask me what I am looking for: Index deniable encryption. Future Generation Computer Systems, 122(??):28– 39, September 2021. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21001199.

Cimmino:2023:SSS

[CCBFI^{+}23] Andrea Cimmino, Juan Cano-Benito, Alba Fernández-Izquierdo, Christos Patsonakis, Apostolos C. Tsolakis, Raúl García-Castro, Dimosthenis Ioannidis, and Dimitrios Tzovaras. A scalable, secure, and semantically interoperable client for cloud-enabled demand response. Future Generation Computer Systems, 141(??):54–66, April 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22003648.

Cauteruccio:2021:FAD

[CCC^{+}21] Francesco Cauteruccio, Luca Cinelli, Enrico Corradini, Giorgio Terracina, Domenico Ursino, Luca Virgili, Claudio Savaglio, Antonio Liotta, and Giancarlo Fortino. A framework for anomaly detection and classification in multiple IoT scenarios. Future Generation Computer Systems, 114(??):322–335, January 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19335253.

Chen:2023:OFE

[CCC^{+}23] Jinfu Chen, Yuhao Chen, Saihua Cai, Shang Yin, Lingling Zhao, and Zikang Zhang. An optimized feature extraction algorithm for abnormal network traffic detection. Future Generation Computer Systems, 149(??):330–342, December 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23002959.

Crisci:2024:EPP

[CCC^{+}24] Luigi Crisci, Lorenzo Carpentieri, Biagio Cosenza, Gianmarco Accordi, Davide Gadioli, Emanuele Vitali, Gianluca Palermo, and Andrea Rosario Beccari. Enabling performance portability on the LiGen drug discovery pipeline. Future Generation Computer Systems, 158(??):44– 59, September 2024. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24001195.

Cinque:2022:VMC

[CCDR22] Marcello Cinque, Domenico Cotroneo, Luigi De Simone, and Stefano Rosiello. Virtualizing mixed-criticality systems: a survey on industrial trends and issues. Future Generation Computer Systems, 129(??):315–330, April 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21004787.

Coleman:2023:AGS

[CCdS23] Tainã Coleman, Henri Casanova, and Rafael Ferreira da Silva. Automated generation of scientific workflow generators with WfChef. Future Generation Computer Systems, 147(??):16–29, October 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23001711.

Chow:2021:EAP

[CCH21] Sherman S. M. Chow, Kim-Kwang Raymond Choo, and Jinguang Han. Editorial for accountability and privacy issues in blockchain and cryptocurrency. Future Generation Computer Systems, 114(??):647–648, January 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20326637.

Charachon:2022:LCG

[CCHA22] Martin Charachon, Paul-Henry Cournède, Céline Hudelot, and Roberto Ardon. Leveraging conditional generative models in a general explanation framework of classifier decisions. Future Generation Computer Systems, 132(??):223–238, July 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22000632.

Chen:2021:SES

[CCHD21] Jing Chen, Jiong Chen, Kun He, and Ruiying Du. SeCrowd: Efficient secure interactive crowdsourcing via permission-based signatures. Future Generation Computer Systems, 115(??):448–458, February 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20305264.

Chen:2020:SSV

[CCL^{+}20] Yunliang Chen, Xiaodao Chen, Wangyang Liu, Yuchen Zhou, Albert Y. Zomaya, Rajiv Ranjan, and Shiyan Hu. Stochastic scheduling for variationaware virtual machine placement in a cloud computing CPS. Future Generation Computer Systems, 105(??):779–788, April 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17320101.

Chen:2021:FES

[CCL^{+}21] Chih-Hsuan Chen, Shuo-Han Chen, Yu-Pei Liang, Tseng-Yi Chen, Tsan sheng Hsu, Hsin-Wen Wei, and Wei-Kuan Shih. Facilitating external sorting on SMR-based large-scale storage systems. Future Generation Computer Systems, 116(??):333–348, March 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20330168.

Chen:2022:SDA

[CCL^{+}22] Jinlin Chen, Jiannong Cao, Zhixuan Liang, Xiaohui Cui, Lequan Yu, and Wei Li. STPD: Defending against ℓ_{0}-norm attacks with space transformation. Future Generation Computer Systems, 126(??):225–236, January 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21003174.

Cheng:2023:MOM

[CCL23] Yan-Yang Cheng, Zheng-Yi Chai, and Ya-Lun Li. Many-objective manytask optimization using reference-points-based nondominated sorting approach. Future Generation Computer Systems, 145(??):496–510, August 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2300119X.

Chapela-Campa:2020:UCP

[CCML20]  David  Chapela-Campa,

Manuel Mucientes, and Manuel Lama. Understanding complex process models by abstracting infrequent behavior. Future Generation Computer Systems, 113(??):428–440, December 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19331206.

Carvalho:2020:UML

[CCP^{+}20] Pablo Carvalho, Esteban Clua, Aline Paes, Cristiana Bentes, Bruno Lopes, and Lúcia Maria de A. Drummond. Using machine learning techniques to analyze the performance of concurrent kernel execution on GPUs. Future Generation Computer Systems, 113(??):528–540, December 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19312658.

Coleman:2022:WFE

[CCP^{+}22] Tainã Coleman, Henri Casanova, Loïc Pottier, Manav Kaushik, Ewa Deelman, and Rafael Ferreira da Silva. WfCommons: a framework for enabling scientific workflow research and development. Future Generation Computer Systems, 128(??):16– 27, March 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21003897.

Caroprese:2022:FLT

[CCTZ22] Luciano Caroprese, Carmela Comito, Domenico Talia, and Ester Zumpano. A fuzzy logic technique for virtual sensor networks. Future Generation Computer Systems, 137(??):302–322, December 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22002564.

Chang:2020:ABM

[CCW^{+}20a] Kuo-Chi Chang, Kai-Chun Chu, Hsiao-Chuan Wang, Yuh-Chung Lin, and Jeng-Shyang Pan. Agent-based middleware framework using distributed CPS for improving resource utilization in smart city. Future Generation Computer Systems, 108(??):445–453, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19334648.

Chen:2020:CAT

[CCW^{+}20b] Wenlong Chen, Heyang Chen, Zhiliang Wang, Chengan Zhao, Mingwei Xu, Ke Xu, and Yingya Guo. Congestion avoidance transmission mechanism based on two-dimensional forwarding. Future Generation Computer Systems, 102(??):1–13, January 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18331509.

Cheng:2020:SVC

[CCW^{+}20c] Yuxia Cheng, Wenzhi Chen, Zonghui Wang, Zhongxian Tang, and Yang Xiang. Smart VM coscheduling with the precise prediction of performance characteristics. Future Generation Computer Systems, 105(??):1016–1027, April 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16306616.

Cranganore:2024:PWH

[CDBD24] Sandeep Suresh Cranganore, Vincenzo De Maio, Ivona Brandic, and Ewa Deelman. Paving the way to hybrid quantum–classical scientific workflows. Future Generation Computer Systems, 158(??):346–366, September 2024. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24001596.

Chen:2024:PPO

[CDC^{+}24] Weiduo Chen, Xiaoshe Dong, Xinhang Chen, Song Liu, Qin Xia, and Qiang Wang. pommDNN: Performance optimal GPU memory management for deep neural network training. Future Generation Computer Systems, 152(??): 160–169, March 2024. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23003989.

CavalcantiBueno:2020:ARS

[CdD20] Andre Luis Cavalcanti Bueno, Noemi de La Rocque Rodriguez, and Elisa Dominguez Sotelino. Adaptive relaxed synchronization through the use of supervised learning methods. Future Generation Computer Systems, 106 (??):260–269, May 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167739X19300469.

Colarusso:2022:PBD

[CDF^{+}22] Carmine Colarusso, Antonio De Iasio, Angelo Furno, Lorenzo Goglia, Mohammed Amine Merzoug, and Eugenio Zimeo.

PROMENADE: a big data platform for handling city complex networks with dynamic graphs. Future Generation Computer Systems, 137(??):129–145, December 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22002357.

Challa:2020:DAA

[CDG^{+}20] Sravani Challa, Ashok Kumar Das, Prosanta Gope, Neeraj Kumar, Fan Wu, and Athanasios V. Vasilakos. Design and analysis of authenticated key agreement scheme in cloudassisted cyber-physical systems. Future Generation Computer Systems, 108(??):1267–1286, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17326328.

Costa:2020:PAO

[CdO20] Daniel G. Costa and Felipe P. de Oliveira. A prioritization approach for optimization of multiple concurrent sensing applications in smart cities. Future Generation Computer Systems, 108(??):228–243, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18329455.

Cinque:2020:CFP

[CDP20a] Marcello Cinque, Raffaele Della Corte, and Antonio Pecchia. Contextual filtering and prioritization of computer application logs for security situational awareness. Future Generation Computer Systems, 111(??):668–680, October 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19306454.

Cuomo:2020:CME

[CDP20b] Salvatore Cuomo, Vittorio Di Somma, and Francesco Piccialli. A computational method for the European option price in an Internet of Things framework. Future Generation Computer Systems, 107(??):730–735, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17316680.

Cuomo:2020:PEB

[CDP20c] Salvatore Cuomo, Vittorio Di Somma, and Francesco Piccialli. Pricing estimation of a barrier option in an IoT scenario. Future Generation Computer Systems, 110(??):407–412, September 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17325074.

Chatterjee:2024:DTC

[CDR24] Pushpita Chatterjee, Debashis Das, and Danda B. Rawat. Digital twin for credit card fraud detection: opportunities, challenges, and fraud detection advancements. Future Generation Computer Systems, 158(??):410–426, September 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24001997.

Cassel:2024:TPP

[CdRRdC^{+}24] Gustavo André Setti Cassel, Rodrigo da Rosa Righi, Cristiano André da Costa, Marta Rosecler Bez, and Marcelo Pasin. Towards providing a priority-based vital sign offloading in healthcare with serverless computing and a fog– cloud architecture. Future Generation Computer Systems, 157(??):51–66, August 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24000979.

Cremaschi:2020:FAA

[CDRS20] Marco Cremaschi, Flavio De Paoli, Anisa Rula, and Blerina Spahiu. A fully automated approach to a complete semantic table interpretation. Future Generation Computer Systems, 112(??):478–500, November 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19302663.

Casanova:2020:DAS

[CdST^{+}20] Henri Casanova, Rafael Ferreira da Silva, Ryan Tanaka, Suraj Pandey, Gautam Jethwani, William Koch, Spencer Albrecht, James Oeth, and Frédéric Suter. Developing accurate and scalable simulators of production workflow management systems with WRENCH. Future Generation Computer Systems, 112(??):162–175, November 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19317431.

Chhetri:2024:EPA

[CDV^{+}24] Tek Raj Chhetri, Chinmaya Kumar Dehury, Blesson Varghese, Anna Fensel, Satish Narayana Srirama, and Rance J. DeLong. Enabling privacy-aware interoperable and quality IoT data sharing with context. Future Generation Computer Systems, 157(??): 164–179, August 2024. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24001109.

Chen:2023:RSD

[CDX^{+}23] Jing Chen, Xiaoqiang Di, Rui Xu, Hui Qi, Ligang Cong, Kehan Zhang, Ziyang Xing, Xiongwen He, Wenping Lei, and Shiwei Zhang. A remote sensing data transmission strategy based on the combination of satellite-ground link and GEO relay under dynamic topology. Future Generation Computer Systems, 145(??):337–353, August 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23000572.

Chen:2024:ESO

[CDX^{+}24] Jing Chen, Xiaoqiang Di, Rui Xu, Hao Luo, Hui Qi, Panpan Zhan, and Yuming Jiang. An efficient scheme for in-orbit remote sensing image data retrieval. Future Generation Computer Systems, 150(??):103–114, January 2024. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23003175.

Chi:2020:EFW

[CDY^{+}20] Kuo Chi, Xiaojiang Du, Guisheng Yin, Jie Wu, Mohsen Guizani, Qilong Han, and Yaling Yang. Efficient and fair Wi-Fi and LTE-U coexistence via communications over content centric networking. Future Generation Computer Systems, 112(??):297–306, November 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20301795.

Chemodanov:2020:RRB

[CECS20] Dmitrii Chemodanov, Flavio Esposito, Prasad Calyam, and Andrei Sukhov. RE-BATE: a REpulsive-BAsed Traffic Engineering protocol for dynamic scalefree networks. Future Generation Computer Systems, 108(??):624–635, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19321004.

Cooke:2020:MDN

[CF20] Ryan A. Cooke and Suhaib A. Fahmy. A model for distributed in-network and near-edge computing with heterogeneous hardware. Future Generation Computer Systems, 105(??): 395–409, April 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19312130.

Colombo:2021:EEA

[CF21] Pietro Colombo and Elena Ferrari. Evaluating the effects of access control policies within NoSQL systems. Future Generation Computer Systems, 114(??):491–505, January 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19330389.

Cabane:2024:IED

[CF24] Hebert Cabane and Kleinner Farias. On the impact of event-driven architecture on performance: an exploratory study. Future Generation Computer Systems, 153(??):52–69, April 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23003977.

Cid-Fuentes:2020:EDH

[CFÁA^{+}20] Javier Álvarez Cid-Fuentes, Pol Álvarez, Ramon Amela, Kuninori Ishii, Rafael K. Morizawa, and Rosa M. Badia. Efficient development of high performance data analytics in Python. Future Generation Computer Systems, 111(??):570–581, October 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18321393.

Chadwick:2020:CEB

[CFC^{+}20] David W. Chadwick, Wenjun Fan, Gianpiero Costantino, Rogerio de Lemos, Francesco Di Cerbo, Ian Herwono, Mirko Manea, Paolo Mori, Ali Sajjad, and Xiao-Si Wang. A cloud-edge based data security architecture for sharing and analysing cyber threat information. Future Generation Computer Systems, 102(??):710–722, January 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19300895.

Chen:2020:PKI

[CFD^{+}20] Yunliang Chen, Junqing Fan, Ze Deng, Bo Du, Xiaohui Huang, and Qirui Gui. PR-KELM: Icing level prediction for transmission lines in smart grid. Future Generation Computer Systems, 102(??):75–83, January 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17323993.

Caporuscio:2020:STC

[CFK^{+}20] Mauro Caporuscio, Francesco Flammini, Narges Khakpour, Prasannjeet Singh, and Johan Thornadtsson. Smarttroubleshooting connected devices: Concept, challenges and opportunities. Future Generation Computer Systems, 111(??):681–697, October 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19306491.

Chen:2020:CFG

[CFL^{+}20] Jing Chen, Jianbin Fang, Weifeng Liu, Tao Tang, and Canqun Yang. clMF: a finegrained and portable alternating least squares algorithm for parallel matrix factorization. Future Generation Computer Systems, 108(??):1192–1205, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17318198.

Chen:2022:BBR

[CFM^{+}22] Lanxiang Chen, Qingxiao Fu, Yi Mu, Lingfang Zeng, Fatemeh Rezaeibagha, and Min-Shiang Hwang. Blockchain-based random auditor committee for integrity verification. Future Generation Computer Systems, 131(??): 183–193, June 2022. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22000267.

Casalicchio:2021:ASP

[CG21] Emiliano Casalicchio and Gabriele Gualandi. ASi-MOV: a self-protecting control application for the smart factory. Future Generation Computer Systems, 115(??):213–235, February 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19320485.

Cao:2020:PBU

[CGFC20] Quyet H. Cao, Madhusudan Giyyarpuram, Reza Farahbakhsh, and Noel Crespi. Policy-based usage control for a trustworthy data sharing platform in smart cities. Future Generation Computer Systems, 107(??):998–1010, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1731107X.

Caballero:2023:RGN

[CGM^{+}23] Juan Caballero, Gibran Gomez, Srdjan Matic, Gustavo Sánchez, Silvia Sebastián, and Arturo Villacañas. The rise of GoodFATR: a novel accuracy comparison methodology for indicator extraction tools. Future Generation Computer Systems, 144 (??):74–89, July 2023. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23000535.

Carneiro:2020:TUS

[CGMT20] Tiago Carneiro, Jan Gmys, Nouredine Melab, and Daniel Tuyttens. Towards ultra-scale branchand-bound using a highproductivity language. Future Generation Computer Systems, 105(??):196–209, April 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1930946X.

Cui:2024:BBS

[CH24] Bo Cui and Yun Hu.

BSELA: a blockchain simulator with event-layered architecture. Future Generation Computer Systems, 151(??):182–195, February 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23003679.

Chamberlain:2020:ATD

[Cha20] Roger D. Chamberlain.

Architecturally truly diverse systems: a review. Future Generation Computer Systems, 110(??):33– 44, September 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19313184.

Chen:2020:MPD

[CHC^{+}20] Hao Chen, Ali Asghar Heidari, Huiling Chen, Mingjing Wang, Zhifang Pan, and Amir H. Gandomi. Multi-population differential evolution-assisted Harris hawks optimization: Framework and case studies. Future Generation Computer Systems, 111(??):175–198, October 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19313263.

Chen:2020:IDC

[Che20] Bo-Wei Chen. Incomplete data classification — Fisher Discriminant Ratios versus Welch Discriminant Ratios. Future Generation Computer Systems, 108(??): 894–908, July 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17322501.

Cabanero:2020:CMD

[CHG^{+}20] Luis Cabañero, Ramón Hervás, Iván González, Jesús Fontecha, Tania Mondéjar, and José Bravo. Characterisation of mobiledevice tasks by their associated cognitive load through EEG data processing. Future Generation Computer Systems, 113(??):380–390, December 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20305112.

Cui:2020:WSC

[CHJ^{+}20] Jiajia Cui, Zhipei Huang, Dina Jiaerken, Ye Fan, Shuxia Zhao, Lingyan Zhang, and Jiankang Wu. A wearable system for cardiopulmonary assessment and personalized respiratory training. Future Generation Computer Systems, 112(??):1131–1140, November 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20300881.

Choi:2022:WOS

[CHJK22] Kihan Choi, Hyuck Han, Hyungsoo Jung, and Sooyong Kang. Workloadoptimized sensor data store for industrial IoT gateways. Future Generation Computer Systems, 135(??):394–408, October 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22001777.

Cich:2020:SSC

[CHKJ20] Glenn Cich, Irith Ben-Arroyo Hartman, Luk Knapen, and Davy Janssens. A simulation study of commuting alternatives for day care centres. Future Generation Computer Systems, 110(??):323–337, September 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18301493.

Cheng:2023:DIT

[CHL23] Pengzhou Cheng, Mu Han, and Gongshen Liu. DESC-IDS: Towards an efficient real-time automotive intrusion detection system based on deep evolving stream clustering. Future Generation Computer Systems, 140(??):266–281, March 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22003351.

Caetano:2022:CSC

[CHS22] André Francisco Morielo Caetano, Celso Massaki Hirata, and Rodrigo Rocha Silva. A comparative study of cluster-based Big Data Cube implementations. Future Generation Computer Systems, 133(??):240–253, August 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22000991.

Chen:2023:DEV

[CHS^{+}23] Meng Chen, Jiaxin Hou, Yongpan Sheng, Yingbo Wu, Sen Wang, Jianyuan Lu, and Qilin Fan. HA-D3QN: Embedding virtual private cloud in cloud data centers with heuristic assisted deep reinforcement learning. Future Generation Computer Systems, 148(??):1–14, November 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2300208X.

Cuadra:2024:EDF

[CHS^{+}24] Julen Cuadra, Ekaitz Hurtado, Isabel Sarachaga, Elisabet Estévez, Oskar Casquero, and Aintzane Armentia. Enabling DevOps for fog applications in the smart manufacturing domain: a modeldriven based platform engineering approach. Future Generation Computer Systems, 157(??):360–375, August 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24001286.

Chen:2020:SSB

[CHW^{+}20] Xueyan Chen, Jie He, Xiaoqiang Wu, Wei Yan, and Wei Wei. Sleep staging by bidirectional long shortterm memory convolution neural network. Future Generation Computer Systems, 109(??):188–196, August 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20302211.

Cielo:2020:HPA

[CIB^{+}20] Salvatore Cielo, Luigi Iapichino, Fabio Baruffa, Matteo Bugli, and Christoph Federrath. Honing and proofing astrophysical codes on the road to exascale. Experiences from code modernization on many-core systems. Future Generation Computer Systems, 112(??):93–107, November 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19308350.

Crnkic:2020:SSO

[CIJM20] Aladin Crnkić, Igor Ivanović, Vladimir Jaćimović, and Nevena Mijajlović. Swarms on the 3-sphere for online clustering of multivariate time series and data streams. Future Generation Computer Systems, 112(??):11–17, November 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19310726.

Cebrian:2020:HTF

[CIS^{+}20] Juan M. Cebrian, Baldomero Imbernón, Jesus Soto, José M. García, and José M. Cecilia. Highthroughput fuzzy clustering on heterogeneous architectures. Future Generation Computer Systems, 106(??):401–411, May 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167739X19313883.

Carretero:2020:NPD

[CK20] Jesus Carretero and Dagmar Krefting. New parallel and distributed tools and algorithms for life sciences. Future Generation Computer Systems, 112(??):1174–1176, November 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2032416X.

Carretero:2024:CCC

[CK24] Jesus Carretero and Dagmar Krefting. Cluster and cloud computing for life sciences. Future Generation Computer Systems, 152(??):254–256, March 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23003904.

Coulibaly:2020:RBM

[CKFT20] Lassana Coulibaly, Bernard Kamsu-Foguem, and Fana Tangara. Rule-based machine learning for knowledge discovering in weather data. Future Generation Computer Systems, 108 (??):861–878, July 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0167739X19324355.

Chadza:2020:AHM

[CKL20] Timothy Chadza, Konstantinos G. Kyriakopoulos, and Sangarapillai Lambotharan. Analysis of hidden Markov model learning algorithms for the detection and prediction of multistage network attacks. Future Generation Computer Systems, 108(??):636–649, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1932847X.

Cuzzocrea:2022:EES

[CKV22] Alfredo Cuzzocrea, Panagiotis Karras, and Akrivi Vlachou. Effective and efficient skyline query processing over attributeorder-preserving-free encrypted data in cloudenabled databases. Future Generation Computer Systems, 126(??):237–251, January 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21003137.

Ciznicki:2021:EPI

[CKW21] Miłosz Ciżnicki, Krzysztof Kurowski, and Jan Weglarz. Energy and performance improvements in stencil computations on multinode HPC systems with different network and communication topologies. Future Generation Computer Systems, 115(??):45–58, February 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19326433.

Cheng:2022:DCD

[CKZ^{+}22] Jianhong Cheng, Hulin Kuang, Qichang Zhao, Yahui Wang, Lei Xu, Jin Liu, and Jianxin Wang. DWT-CV: Dense weight transfer-based cross validation strategy for model selection in biomedical data analysis. Future Generation Computer Systems, 135(??):20–29, October 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22001546.

Celan:2020:BAT

[CL20a] Marko Celan and Marjan Lep. Bus-arrival time prediction using bus network data model and time periods. Future Generation Computer Systems, 110(??):364–371, September 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17327085.

Cui:2020:SMS

[CL20b] Tiejun Cui and Shasha Li. System movement space and system mapping theory for reliability of IoT. Future Generation Computer Systems, 107(??):70– 81, June 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19315298.

Chen:2021:PIS

[CL21] Hao Chen and Qiongfu Lian. Poverty/investment slow distribution effect analysis based on Hopfield neural network. Future Generation Computer Systems, 122(??):63–68, September 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21001138.

Camacho:2021:NRM

[CLC21a] David Camacho, Ma. Victoria Luzón, and Erik Cambria. New research methods & algorithms in social network analysis. Future Generation Computer Systems, 114(??):290–293, January 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20324912.

Camacho:2021:NTA

[CLC21b] David Camacho, Ma. Victoria Luzón, and Erik Cambria. New trends and applications in social media analytics. Future Generation Computer Systems, 114(??):318–321, January 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20324924.

Chen:2023:PTA

[CLL^{+}23] Rui Chen, Bo Liu, WeiWei Lin, JianPeng Lin, Hui-Wen Cheng, and KeQin Li. Power and thermal-aware virtual machine scheduling optimization in cloud data center. Future Generation Computer Systems, 145(??):578–589, August 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23001346.

Caino-Lores:2020:ABD

[CLLCK20] S. Caíno-Lores, A. Lapin, J. Carretero, and P. Kropf. Applying big data paradigms to a large scale scientific workflow: Lessons learned and future directions. Future Generation Computer Systems, 110(??):440–452, September 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16308214.

Chen:2020:STT

[CLQS20] Chen Chen, Xiaomin Liu, Tie Qiu, and Arun Kumar Sangaiah. A short-term traffic prediction model in the vehicular cyberphysical systems. Future Generation Computer Systems, 105(??):894–903, April 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17311846.

Cesario:2024:SMD

[CLV24] Eugenio Cesario, Paolo Lindia, and Andrea Vinci. A scalable multi-density clustering approach to detect city hotspots in a smart city. Future Generation Computer Systems, 157(??):226–236, August 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24001122.

Chen:2020:SGA

[CLY^{+}20] Yifan Chen, Zhiyong Li, Bo Yang, Ke Nai, and Keqin Li. A Stackelberg game approach to multiple resources allocation and pricing in mobile edge computing. Future Generation Computer Systems, 108(??):273–287, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19311653.

Chen:2020:AFE

[CLZ^{+}20] Xu Chen, Jianjun Li, Yanchao Zhang, Yu Lu, and Shaoyu Liu. Automatic feature extraction in X-ray image based on deep learning approach for determination of bone age. Future Generation Computer Systems, 110(??):795–801, September 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19312026.

Chen:2021:CCO

[CLZ21] Ming Chen, Yunhao Li, and Xiuze Zhou. CoNet: Cooccurrence neural networks for recommendation. Future Generation Computer Systems, 124(??):308–314, November 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X2100203X.

Cordeiro:2022:TSF

[CMA^{+}22] Matheus Cordeiro, Catherine Markert, Sayonara S. Araújo, Nídia G. S. Campos, Rubens S. Gondim, Ticiana L. Coelho da Silva, and Atslands R. da Rocha. Towards smart farming: Fog-enabled intelligent irrigation system using deep neural networks. Future Generation Computer Systems, 129(??):115–124, April 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21004428.

Centelles:2021:LCS

[CMF^{+}21] Roger Pueyo Centelles, Roc Meseguer, Felix Freitag, Leandro Navarro, Sergio F. Ochoa, and Rodrigo M. Santos. LoRaMoto: a communication system to provide safety awareness among civilians after an earthquake. Future Generation Computer Systems, 115(??):150–170, February 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20306063.

Cecilia:2023:URG

[CMGI^{+}23] José M. Cecilia, Juan Morales-García, Baldomero Imbernón, Javier Prades, Juan-Carlos Cano, and Federico Silla. Using remote GPU virtualization techniques to enhance edge computing devices. Future Generation Computer Systems, 142(??):14–24, May 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22004423.

Choudhary:2022:IAC

[CMGS22] Tejalal Choudhary, Vipul Mishra, Anurag Goswami, and Jagannathan Sarangapani. Inference-aware convolutional neural network pruning. Future Generation Computer Systems, 135(??):44–56, October 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22001601.

Chen:2021:FSA

[CMM21] Ming Chen, Bingcheng Mao, and Tianyi Ma. FedSA: a staleness-aware asynchronous federated learning algorithm with non-IID data. Future Generation Computer Systems, 120(??):1–12, July 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21000649.

Corradini:2023:FAM

[CMM^{+}23] Flavio Corradini, Alessandro Marcelletti, Andrea Morichetta, Andrea Polini, Barbara Re, and Francesco Tiezzi. A flexible approach to multi-party business process execution on blockchain. Future Generation Computer Systems, 147(??):219–234, October 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23001814.

Carrillo-Mondejar:2020:CLB

[CMMST20] J. Carrillo-Mondéjar, J. L.

Martínez, and G. Suarez-Tangil. Characterizing Linux-based malware: Findings and recent trends. Future Generation Computer Systems, 110(??): 267–281, September 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167739X19325002.

Cao:2020:TTD

[CMX^{+}20] Jiuxin Cao, Zhuo Ma, Jue Xie, Xiangying Zhu, Fang Dong, and Bo Liu. Towards tenant demandaware bandwidth allocation strategy in cloud datacenter. Future Generation Computer Systems, 105 (??):904–915, April 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167739X17311834.

Choi:2022:CBT

[CP22] Sang-Hoon Choi and Ki-Woong Park. Cloud-BlackBox: Toward practical recording and tracking of VM swarms for multifaceted cloud inspection. Future Generation Computer Systems, 137(??):219–233, December 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22002321.

Chen:2022:CAI

[CPH^{+}22] Zhen Chen, Maosheng Pan, Pengfei He, Wenchao Qi, Linlin Liu, Limin Shen, and Dianlong You. Context and auto-interaction are all you need: Towards context embedding based QoS prediction via automatic feature interaction for high quality cloud API delivery. Future Generation Computer Systems, 128(??):265–281, March 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21004040.

Choi:2020:IK

[CPJ20] Chang Choi, Francesco Piccialli, and Jason J. Jung. Internet of Knowledge. Future Generation Computer Systems, 102(??):948–949, January 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19319752.

Cosenza:2021:EEA

[CPJ^{+}21] Biagio Cosenza, Nikita Popov, Ben Juurlink, Paul Richmond, Mozhgan Kabiri Chimeh, Carmine Spagnuolo, Gennaro Cordasco, and Vittorio Scarano. Easy and efficient agent-based simulations with the Open-ABL language and compiler. Future Generation Computer Systems, 116 (??):61–75, March 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167739X20329915.

Chard:2023:GAS

[CPM^{+}23] Ryan Chard, Jim Pruyne, Kurt McKee, Josh Bryan, Brigitte Raumann, Rachana Ananthakrishnan, Kyle Chard, and Ian T. Foster. Globus automation services: Research process automation across the space–time continuum. Future Generation Computer Systems, 142(??):393–409, May 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23000183.

Camargo:2024:VAS

[CPPY24] L. S. Camargo, J. F.

Pauletti, A. M. Pernas, and A. Yamin. VISO approach: a socialization proposal for the Internet of Things objects. Future Generation Computer Systems, 150(??):326–340, January 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23003345.

Campioni:2023:ECM

[CPS^{+}23] Lorenzo Campioni, Filippo Poltronieri, Cesare Stefanelli, Niranjan Suri, Mauro Tortonesi, and Konrad Wrona. Enabling civil–military collaboration for disaster relief operations in smart city environments. Future Generation Computer Systems, 139(??):181–195, February 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22003041.

Chen:2020:SDC

[CPT^{+}20] Hsing-Chung Chen, Karisma Trinanda Putra, Shian-Shyong Tseng, Chin-Ling Chen, and Jerry Chun-Wei Lin. A spatiotemporal data compression approach with low transmission cost and high data fidelity for an air quality monitoring system. Future Generation Computer Systems, 108(??):488–500, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19324197.

Cai:2023:PFL

[CPYY23] Cailing Cai, Shimin Pan, Tsz Hon Yuen, and Siu-Ming Yiu. Practical fully leakage resilient signatures with auxiliary inputs. Future Generation Computer Systems, 141(??):448–461, April 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2200396X.

Cheng:2024:LVB

[CQA^{+}24] Jingxian Cheng, Saiyu Qi, Bochao An, Yong Qi, Jianfeng Wang, and Yanan Qiao. Lightweight verifiable blockchain top-k queries. Future Generation Computer Systems, 156(??): 105–115, July 2024. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24000827.

Chen:2023:CSM

[CQS^{+}23] Genxin Chen, Jin Qi, Ying Sun, Xiaoxuan Hu, Zhenjiang Dong, and Yanfei Sun. A collaborative scheduling method for cloud computing heterogeneous workflows based on deep reinforcement learning. Future Generation Computer Systems, 141(??):284–297, April 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22004034.

Cassel:2022:SCI

[CRdRR^{+}22] Gustavo André Setti Cassel, Vinicius Facco Rodrigues, Rodrigo da Rosa Righi, Marta Rosecler Bez, Andressa Cruz Nepomuceno, and Cristiano André da Costa. Serverless computing for Internet of Things: a systematic literature review. Future Generation Computer Systems, 128(??):299–316, March 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21004167.

Castellanos-Rodriguez:2024:SLP

[CREE^{+}24] Óscar Castellanos-Rodríguez, Roberto R. Expósito, Jonatan Enes, Guillermo L. Taboada, and Juan Touriño. Serverlesslike platform for containerbased YARN clusters. Future Generation Computer Systems, 155(??):256–271, June 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2400058X.

Chouliaras:2023:AAS

[CS23] Spyridon Chouliaras and Stelios Sotiriadis. An adaptive auto-scaling framework for cloud resource provisioning. Future Generation Computer Systems, 148(??):173–183, November 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23002005.

Chouliaras:2024:TCO

[CS24a] Spyridon Chouliaras and Stelios Sotiriadis. Towards constrained optimization of cloud applications: a hybrid approach. Future Generation Computer Systems, 151(??):100–110, February 2024. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23003539.

Czentye:2024:SAC

[CS24b] János Czentye and Balázs Sonkoly. Serverless application composition leveraging function fusion: Theory and algorithms. Future Generation Computer Systems, 153(??):403–418, April 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23004648.

Chen:2024:WAO

[CSAT24] Mingkang Chen, Jingtao Sun, Kento Aida, and Atsuko Takefusa. Weatheraware object detection method for maritime surveillance systems. Future Generation Computer Systems, 151(??):111–123, February 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X2300362X.

Chaudhary:2023:RCM

[CSB23] Chandan Kumar Chaudhary, Richa Sarma, and Ferdous Ahmed Barbhuiya. RMA-CPABE: a multiauthority CPABE scheme with reduced ciphertext size for IoT devices. Future Generation Computer Systems, 138(??):226–242, January 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22002783.

Choi:2023:PDM

[CSC23] Seok-Hwan Choi, Jinmyeong Shin, and Yoon-Ho Choi. PIHA: Detection method using perceptual image hashing against query-based adversarial attacks. Future Generation Computer Systems, 145(??):563–577, August 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23001395.

Chithaluru:2023:OPI

[CSD^{+}23] Premkumar Chithaluru, Aman Singh, Jagjit Singh Dhatterwal, Ali Hassan Sodhro, Marwan Ali Albahar, Anca Jurcut, and Ahmed Alkhayyat. An optimized privacy information exchange schema for explainable AI empowered WiMAX-based IoT networks. Future Generation Computer Systems, 148(??):225–239, November 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23002170.

Choi:2023:CMO

[CSH^{+}23] Kihan Choi, Hyungseok Seo, Hyuck Han, Minsoo Ryu, and Sooyong Kang. CredsCache: Making OverlayFS scalable for containerized services. Future Generation Computer Systems, 147(??):44–58, October 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X2300167X.

Croft:2022:DPF

[CSS22] William L. Croft, Jörg-Rüdiger Sack, and Wei Shi. Differentially private facial obfuscation via generative adversarial networks. Future Generation Computer Systems, 129(??):358–379, April 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21004763.

Chen:2020:AEW

[CSY^{+}20] Zhen Chen, Yuanhao Sun, Dianlong You, Feng Li, and Limin Shen. An accurate and efficient web service QoS prediction model with wide-range awareness. Future Generation Computer Systems, 109(??):275–292, August 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19329061.

Chen:2022:CDC

[CTFW22] Yu Chen, Wei Tong, Dan Feng, and Zike Wang. Cora: Data correlationsbased storage policies for cloud object storage. Future Generation Computer Systems, 129(??):331–346, April 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21004672.

Casado-Vara:2020:INS

[CVdRA^{+}20] Roberto Casado-Vara, Angel Martin del Rey, Soffiene Affes, Javier Prieto, and Juan M. Corchado. IoT network slicing on virtual layers of homogeneous data for improved algorithm operation in smart buildings. Future Generation Computer Systems, 102(??):965–977, January 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19304819.

Cao:2020:TOS

[CWB^{+}20] Huiyan Cao, Chase Q. Wu, Liang Bao, Aiqin Hou, and Wei Shen. Throughput optimization for storm-based processing of stream data on clouds. Future Generation Computer Systems, 112(??):567–579, November 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19313093.

Chien:2020:QLB

[CWL20] Wei-Che Chien, Hung-Yen Weng, and Chin-Feng Lai. Q-learning based collaborative cache allocation in mobile edge computing. Future Generation Computer Systems, 102(??):603–610, January 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19305473.

Chen:2020:SAR

[CWM^{+}20] Xing Chen, Haijiang Wang, Yun Ma, Xianghan Zheng, and Longkun Guo. Selfadaptive resource allocation for cloud-based software services based on iterative QoS prediction model. Future Generation Computer Systems, 105(??): 287–296, April 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19302894.

Cao:2021:TCA

[CWM21] Yuan Cao, Jiakun Wen, and Lianchuan Ma. Tracking and collision avoidance of virtual coupling train control system. Future Generation Computer Systems, 120(??):76–90, July 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21000662.

Cai:2023:ARE

[CWYG23] Binlei Cai, Bin Wang, Meihong Yang, and Qin Guo. AutoMan: Resourceefficient provisioning with tail latency guarantees for microservices. Future Generation Computer Systems, 143(??):61–75, June 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23000213.

Chen:2023:BBS

[CXHC23] Jiahui Chen, Hang Xiao, Muchuang Hu, and Chien-Ming Chen. A blockchainbased signature exchange protocol for metaverse. Future Generation Computer Systems, 142(??):237–247, May 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22004356.

Chen:2020:RII

[CXHS20] Yi Chen, Fu Xiao, Haiping Huang, and Lijuan Sun. RF-IDH: an intelligent fall detection system for hemodialysis patients via COTS RFID. Future Generation Computer Systems, 113(??):13–24, December 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19327074.

Cai:2022:HEH

[CXS^{+}22] Yunyun Cai, Wei Xi, Yuhao Shen, Youcheng Peng, Shixuan Song, and Jizhong Zhao. High-efficient hierarchical federated learning on non-IID data with progressive collaboration. Future Generation Computer Systems, 137(??):111–128, December 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22002394.

Chen:2021:BBP

[CXWY21] Zeng Chen, Weidong Xu, Bingtao Wang, and Hua Yu. A blockchain-based preserving and sharing system for medical data privacy. Future Generation Computer Systems, 124(??):338–350, November 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21001734.

Cong:2022:EEP

[CYG22] Linhao Cong, Jia Yu, and Xinrui Ge. Enabling efficient privacypreserving subgraph isomorphic query over graphs. Future Generation Computer Systems, 132(??):1– 10, July 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22000358.

Chu:2020:NDL

[CYH20] Zheng Chu, Jiong Yu, and Askar Hamdulla. A novel deep learning method for query task execution time prediction in graph database. Future Generation Computer Systems, 112(??):534–548, November 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19321156.

Chen:2024:LPB

[CYWS24] Shengbo Chen, Jidong Yuan, Zhihai Wang, and Yongqi Sun. Local perturbationbased black-box federated learning attack for time series classification. Future Generation Computer Systems, 158(??):488–500, September 2024. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24001894.

Chen:2022:MFI

[CYZ^{+}22] Yuanyi Chen, Peng Yu, Zengwei Zheng, Jiaxing Shen, and Minyi Guo. Modeling feature interactions for context-aware QoS prediction of IoT services. Future Generation Computer Systems, 137(??):173–185, December 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X2200245X.

Cubukcu:2021:KBI

[ÇYZZ21] Burakhan Çubukçu, Uğur Yüzgeç, Ahu Zileli, and Raif Zileli. Kinect-based integrated physiotherapy mentor application for shoulder damage. Future Generation Computer Systems, 122(??):105–116, September 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21001187.

Cao:2020:GSB

[CZ20] Mengyun Cao and Hai Zhuge. Grouping sentences as better language unit for extractive text summarization. Future Generation Computer Systems, 109(??):331–359, August 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19318989.

Cao:2024:MHN

[CZCH24] Yuehang Cao, Xiang Zhao, Dong Chen, and Hongbin Huang. Multiplex heterogeneous network representation learning with unipath based global awareness neural network. Future Generation Computer Systems, 150(??):317–325, January 2024. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23003291.

Chen:2020:PDB

[CZGS20] Pan Chen, Xiaoping Zheng, Fuqiang Gu, and Jianga Shang. Path distance-based map matching for Wi-Fi fingerprinting positioning. Future Generation Computer Systems, 107(??):82– 94, June 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19309240.

Chen:2024:COB

[CZH^{+}24] Zheyi Chen, Junjie Zhang, Zhiqin Huang, Pengfei Wang, Zhengxin Yu, and Wang Miao. Computation offloading in blockchainenabled MCS systems: a scalable deep reinforcement learning approach. Future Generation Computer Systems, 153(??):301–311, April 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23004673.

Chen:2023:NOA

[CZZ^{+}23a] Lu Chen, De-Gan Zhang, Jie Zhang, Ting Zhang, Wen-Jing Wang, and Ya-Hui Cao. A novel offloading approach of IoT user perception task based on quantum behavior particle swarm optimization. Future Generation Computer Systems, 141(??):577–594, April 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22004149.

Cheng:2023:PAD

[CZZ^{+}23b] Xin Cheng, Jingmei Zhou, Xiangmo Zhao, Hongfei Wang, and Yuqi Li. A presentation attack detection network based on dynamic convolution and multi-level feature fusion with security and reliability. Future Generation Computer Systems, 146(??):114–121, September 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23001449.

Dahi:2022:MQC

[DA22] Zakaria Abdelmoiz Dahi and Enrique Alba. Metaheuristics on quantum computers: Inspiration, simulation and real execution. Future Generation Computer Systems, 130(??):164–180, May 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21004969.

Daud:2021:FRS

[DAA^{+}21] Ali Daud, Faizan Abbas, Tehmina Amjad, Abdulrahman A. Alshdadi, and Jalal S. Alowibdi. Finding rising stars through hot topics detection. Future Generation Computer Systems, 115(??):798–813, February 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20329903.

Delicato:2020:ESC

[DAAW20] Flavia C. Delicato, Adnan Al-Anbuky, and Kevin I-Kai Wang. Editorial: Smart cyber-physical systems: Toward pervasive intelligence systems. Future Generation Computer Systems, 107(??):1134–1139, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19316619.

deAguiar:2022:BBP

[dAdSM^{+}22] Erikson J. de Aguiar, Alyson J. dos Santos, Rodolfo I. Meneguette, Robson E. De Grande, and Jó Ueyama. A blockchainbased protocol for tracking user access to shared medical imaging. Future Generation Computer Systems, 134(??):348–360, September 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22001406.

Darem:2021:VDL

[DAM^{+}21] Abdulbasit Darem, Jemal Abawajy, Aaisha Makkar, Asma Alhashmi, and Sultan Alanazi. Visualization and deep-learning-based malware variant detection using OpCode-level features. Future Generation Computer Systems, 125(??):314–323, December 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21002272.

Damaj:2023:EAF

[DAMS23] Issam W. Damaj, Hadi Al-Mubasher, and Mahmoud Saadeh. An extended analytical framework for heterogeneous implementations of light cryptographic algorithms. Future Generation Computer Systems, 141(??):154–172, April 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22003673.

Plaza-del-Arco:2020:IER

[dAMVULM20] Flor Miriam Plaza del Arco, M. Teresa Martín-Valdivia, L. Alfonso Ureña-López, and Ruslan Mitkov. Improved emotion recognition in Spanish social media through incorporation of lexical knowledge. Future Generation Computer Systems, 110(??):1000–1008, September 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1931163X.

Dao:2023:IWT

[Dao23] Nhu-Ngoc Dao. Internet of wearable things: Advancements and benefits from 6G technologies. Future Generation Computer Systems, 138(??):172–184, January 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22002345.

Martinez-del-Amor:2020:APS

[dAPHOMPJ20] Miguel Á. Martínez del Amor, Ignacio Pérez-Hurtado, David Orellana-Martín, and Mario J. Pérez-Jiménez. Adaptative parallel simulators for bioinspired computing models. Future Generation Computer Systems, 107(??):469–484, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19308817.

Deebak:2021:LAI

[DAT21] B. D. Deebak and Fadi Al-Turjman. Lightweight authentication for IoT/Cloudbased forensics in intelligent data computing. Future Generation Computer Systems, 116(??):406–425, March 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X2033034X.

Deebak:2020:IBS

[DATAA20] B. D. Deebak, Fadi Al-Turjman, Moayad Aloqaily, and Omar Alfandi. IoT-BSFCAN: a smart contextaware system in IoT-Cloud using mobile-fogging. Future Generation Computer Systems, 109(??):368–381, August 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19320874.

Diaz-de-Arcaya:2023:OAF

[dATBMA23] Josu Díaz de Arcaya, Ana I.

Torre-Bastida, Raúl Miñón, and Aitor Almeida. Orfeon: an AIOps framework for the goal-driven operationalization of distributed analytical pipelines. Future Generation Computer Systems, 140(??):18–35, March 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22003223.

Duque:2024:LVL

[DBBP24] Rafael Duque, Crescencio Bravo, Santos Bringas, and Daniel Postigo. Leveraging a visual language for the awareness-based design of interaction requirements in digital twins. Future Generation Computer Systems, 153(??):41– 51, April 2024. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23004296.

DeSalve:2024:EDI

[DBC24] Andrea De Salve, Alessandro Brighente, and Mauro Conti. EDIT: a data inspection tool for smart contracts temporal behavior modeling and prediction. Future Generation Computer Systems, 154(??): 413–425, May 2024. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X24000013.

Debicha:2023:TTL

[DBD^{+}23] Islam Debicha, Richard Bauwens, Thibault Debatty, Jean-Michel Dricot, Tayeb Kenaza, and Wim Mees. TAD: Transfer learning-based multiadversarial detection of evasion attacks against network intrusion detection systems. Future Generation Computer Systems, 138(??):185–197, January 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22002722.

Djenouri:2023:HGC

[DBSL23] Youcef Djenouri, Asma Belhadi, Gautam Srivastava, and Jerry Chun-Wei Lin. Hybrid graph convolution neural network and branchand-bound optimization for traffic flow forecasting. Future Generation Computer Systems, 139(??):100–108, February 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22003028.

Dhou:2021:HEC

[DC21] Khaldoon Dhou and Christopher Cruzen. A highly efficient chain code for compression using an agentbased modeling simulation of territories in biological beavers. Future Generation Computer Systems, 118 (??):1–13, May 2021. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20330788.

Das:2022:WDA

[DCC22] Anirban Das, Sandip Chakraborty, and Suchetana Chakraborty. Where do all my smart home data go? Contextaware data generation and forwarding for edge-based microservices over shared IoT infrastructure. Future Generation Computer Systems, 134(??):204–218, September 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22001029.

Ding:2022:IDC

[DCD^{+}22] Hongwei Ding, Leiyang Chen, Liang Dong, Zhongwang Fu, and Xiaohui Cui. Imbalanced data classification: a KNN and generative adversarial networksbased hybrid approach for intrusion detection. Future Generation Computer Systems, 131(??):240–254, June 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22000346.

DeSensi:2024:CCA

[DCD^{+}24] Daniele De Sensi, Edgar Costa Molero, Salvatore Di Girolamo, Laurent Vanbever, and Torsten Hoefler. Canary: Congestionaware in-network allreduce using dynamic trees. Future Generation Computer Systems, 152(??):70– 82, March 2024. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23003850.

DAngelo:2020:DLS

[DCGM20] Mirko D’Angelo, Mauro Caporuscio, Vincenzo Grassi, and Raffaela Mirandola. Decentralized learning for self-adaptive QoS-aware service assembly. Future Generation Computer Systems, 108(??):210–227, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19312439.

Diaz:2020:SET

[DCZ20] Cristian Mateos Diaz, Kim-Kwang Raymond Choo, and Alejandro Zunino. Sharpening the edge: Towards improved edge computing environment for mobile and IoT applications. Future Generation Computer Systems, 107(??): 1130–1133, June 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19315353.

Diaconescu:2021:ECM

[DDM21] Ada Diaconescu, Louisa Jane Di Felice, and Patricia Mellodge. Exogenous coordination in multi-scale systems: How information flows and timing affect system properties. Future Generation Computer Systems, 114(??):403–426, January 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20304726.

DAngelo:2023:IPM

[DDMP^{+}23] Gianni D’Angelo, David Della-Morte, Donatella Pastore, Giulia Donadel, Alessandro De Stefano, and Francesco Palmieri. Identifying patterns in multiple biomarkers to diagnose diabetic foot using an explainable genetic programmingbased approach. Future Generation Computer Systems, 140(??):138–150, March 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X2200334X.

Deng:2023:NIP

[DDT^{+}23] Guoqiang Deng, Xuefeng Duan, Min Tang, Yuhao Zhang, and Ying Huang. Non-interactive and privacy-preserving neural network learning using functional encryption. Future Generation Computer Systems, 145(??):454–465, August 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23001218.

Dehghani:2020:DFO

[Deh20]  Abbas Dehghani.  A

design flow for an optimized congestion-aware application-specific wireless network-on-chip architecture. Future Generation Computer Systems, 106(??):234–249, May 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167739X19306594.

Diouf:2020:BFT

[DEJ20] Gor Mack Diouf, Halima Elbiaze, and Wael Jaafar. On Byzantine fault tolerance in multi-master Kubernetes clusters. Future Generation Computer Systems, 109(??):407–419, August 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19307939.

Deng:2020:MHO

[Den20]  Zhi-Hong Deng.  Min-

ing high occupancy itemsets. Future Generation Computer Systems, 102(??):222–229, January 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17302662.

Cunha:2023:SAR

[dFCC23] Renato Luiz de Freitas Cunha and Luiz Chaimowicz. An SMDP approach for reinforcement learning in HPC cluster schedulers. Future Generation Computer Systems, 139(??):239–252, February 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22003090.

Dallaqua:2021:FPC

[DFF21] Fernanda B. J. R. Dallaqua, Álvaro L. Fazenda, and Fabio A. Faria. ForestEyes Project: Conception, enhancements, and challenges. Future Generation Computer Systems, 124(??):422–435, November 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21001965.

DAngelo:2023:PPM

[DFF^{+}23] Gianni D’Angelo, Eslam Farsimadan, Massimo Ficco, Francesco Palmieri, and Antonio Robustelli. Privacy-preserving malware detection in Android-based IoT devices through federated Markov chains. Future Generation Computer Systems, 148(??):93–105, November 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23002054.

DeRosa:2021:PES

[DFG^{+}21] Michela De Rosa, Giuseppe Fenza, Alessandro Gallo, Mariacristina Gallo, and Vincenzo Loia. Pharmacovigilance in the era of social media: Discovering adverse drug events cross-relating Twitter and PubMed. Future Generation Computer Systems, 114(??):394–402, January 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19331188.

Ding:2020:QLB

[DFZ^{+}20] Ding Ding, Xiaocong Fan, Yihuan Zhao, Kaixuan Kang, Qian Yin, and Jing Zeng. Q-learning based dynamic task scheduling for energy-efficient cloud computing. Future Generation Computer Systems, 108(??):361–371, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19313858.

Duan:2023:TFF

[DFZ^{+}23] Guoyun Duan, Yuanzhi Fu, Boyang Zhang, Peiyao Deng, Jianhua Sun, Hao Chen, and Zhiwen Chen. TEEFuzzer: a fuzzing framework for trusted execution environments with heuristic seed mutation. Future Generation Computer Systems, 144(??): 192–204, July 2023. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23000857.

Dahiya:2021:RSP

[DG21] Amrita Dahiya and Brij B. Gupta. A reputation score policy and Bayesian game theory based incentivized mechanism for DDoS attacks mitigation and cyber defense. Future Generation Computer Systems, 117(??):193–204, April 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20330600.

DelaParra:2020:IAN

[DGK20] Cecilia De la Parra, Andre Guntoro, and Akash Kumar. Improving approximate neural networks for perception tasks through specialized optimization. Future Generation Computer Systems, 113(??):597–606, December 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20301576.

Daley:2020:PCS

[DGL^{+}20] C. S. Daley, D. Ghoshal, G. K. Lockwood, S. Dosanjh, L. Ramakrishnan, and N. J. Wright. Performance characterization of scientific workflows for the optimal use of burst buffers. Future Generation Computer Systems, 110(??):468–480, September 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16308287.

Djebali:2024:SID

[DGT24] Sonia Djebali, Guillaume Guerard, and Ihab Taleb. Survey and insights on digital twins design and smart grid’s applications. Future Generation Computer Systems, 153(??):234–248, April 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23004466.

Diao:2022:FFL

[DGY^{+}22] Xinglong Diao, Huaxi Gu, Xiaoshan Yu, Liang Qin, and Changyun Luo. Flex: a flowlet-level load balancing based on load-adaptive timeout in DCN. Future Generation Computer Systems, 130(??):219–230, May 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21005021.

Din:2020:PPU

[DHA^{+}20] Ikram Ud Din, Suhaidi Hassan, Ahmad Almogren, Farrukh Ayub, and Mohsen Guizani. PUC: Packet update caching for energy efficient IoTbased information-centric networking. Future Generation Computer Systems, 111(??):634–643, October 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19325087.

Dai:2023:SAM

[DHC23] Cheng Dai, Yinqin Huang, and Wei-Che Chien. A sparse attack method on skeleton-based human action recognition for intelligent metaverse application. Future Generation Computer Systems, 143(??):51– 60, June 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22004472.

Donnell:2020:DVM

[DHD20] Nicola Mc Donnell, Enda Howley, and Jim Duggan. Dynamic virtual machine consolidation using a multi-agent system to optimise energy efficiency in cloud computing. Future Generation Computer Systems, 108(??):288–301, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19314591.

Dhou:2020:NCC

[Dho20] Khaldoon Dhou. A new chain coding mechanism for compression stimulated by a virtual environment of a predator– prey ecosystem. Future Generation Computer Systems, 102(??):650–669, January 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1832630X.

Day:2020:AFS

[DIB20] Patrick Day, Stefano Iannucci, and Ioana Banicescu. Autonomic feature selection using computational intelligence. Future Generation Computer Systems, 111(??):68–81, October 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19320564.

Benedictis:2024:NAV

[DJP^{+}24] Marco De Benedictis, Ludovic Jacquin, Ignazio Pedone, Andrea Atzeni, and Antonio Lioy. A novel architecture to virtualise a hardware-bound trusted platform module. Future Generation Computer Systems, 150(??):21–36, January 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2300314X.

DeMaio:2020:MOS

[DK20] Vincenzo De Maio and Dragi Kimovski. Multiobjective scheduling of extreme data scientific workflows in fog. Future Generation Computer Systems, 106(??):171–184, May 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19309197.

Dong:2024:MTL

[DK24] Huiyao Dong and Igor Kotenko. Multi-task learning for IoT traffic classification: a comparative analysis of deep autoencoders. Future Generation Computer Systems, 158(??):242–254, September 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24001341.

Dwivedi:2022:GDF

[DKD22] Amit Kumar Dwivedi, Naveen Kumar, and Manik Lal Das. Group data freshness scheme for outsourced data in distributed systems. Future Generation Computer Systems, 133(??):141–152, August 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22000863.

Deperlioglu:2022:EFG

[DKG^{+}22] Omer Deperlioglu, Utku Kose, Deepak Gupta, Ashish Khanna, Fabio Giampaolo, and Giancarlo Fortino. Explainable framework for glaucoma diagnosis by image processing and convolutional neural network synergy: Analysis with doctor evaluation. Future Generation Computer Systems, 129(??):152–169, April 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21004556.

Duan:2022:FDF

[DLC^{+}22] Shaoming Duan, Chuanyi Liu, Zhengsheng Cao, Xiaopeng Jin, and Peiyi Han. Fed-DR-Filter: Using global data representation to reduce the impact of noisy labels on the performance of federated learning. Future Generation Computer Systems, 137(??):336–348, December 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22002412.

Damasceno:2023:TMT

[DLdAR23] Nilson L. Damasceno, Marcos Lage, and Antônio A. de A. Rocha. Tinycubes: a modular technology for interactive visual analysis of historical and continuously updated spatiotemporal data. Future Generation Computer Systems, 143(??):378–391, June 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23000146.

DAgostino:2020:CSE

[DLGW^{+}20] Daniele D’Agostino, Duncan Law-Green, Mike Watson, Giovanni Novara, Andrea Tiengo, Stefano Sandrelli, Andrea Belfiore, Ruben Salvaterra, and Andrea De Luca. A citizen science exploration of the X-ray transient sky using the EXTraS science gateway. Future Generation Computer Systems, 111(??):806–818, October 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19301372.

Ding:2020:AVM

[DLH^{+}20] Weichao Ding, Fei Luo, Liangxiu Han, Chunhua Gu, Haifeng Lu, and Joel Fuentes. Adaptive virtual machine consolidation framework based on performance-to-power ratio in cloud data centers. Future Generation Computer Systems, 111(??):254–270, October 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19307769.

Do:2022:PLS

[DLHD22] Dinh-Thuan Do, Anh-Tu Le, Nhat-Duy Xuan Ha, and Nhu-Ngoc Dao. Physical layer security for Internet of Things via reconfigurable intelligent surface. Future Generation Computer Systems, 126(??):330–339, January 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21003204.

Duan:2020:NIC

[DLL20] Li Duan, Yong Li, and Lijun Liao. Non-interactive certificate update protocol for efficient authentication in IoT. Future Generation Computer Systems, 113(??):132–144, December 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X2030412X.

DeLucia:2023:UPE

[DLR23] Gianluca De Lucia, Marco Lapegna, and Diego Romano. Unlocking the potential of edge computing for hyperspectral image classification: an efficient lowenergy strategy. Future Generation Computer Systems, 147(??):207–218, October 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23001802.

delaVega:2020:MAG

[dlVGSB^{+}20] Alfonso de la Vega, Diego García-Saiz, Carlos Blanco, Marta Zorrilla, and Pablo Sánchez. Mortadelo: Automatic generation of NoSQL stores from platform-independent data models. Future Generation Computer Systems, 105(??):455–474, April 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167739X19312063.

Du:2023:CCN

[DLW^{+}23] Shaoyi Du, Yuying Liu, Xijing Wang, Yuting Chi, Nanning Zheng, and Yucheng Guo. Curriculum classification network based on margin balancing multi-loss and ensemble learning. Future Generation Computer Systems, 145(??):150–163, August 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23000900.

Dong:2023:AFE

[DLZ^{+}23] Liguo Dong, Zhenmou Liu, Kejia Zhang, Abdulsalam Yassine, and M. Shamim Hossain. Affordable federated edge learning framework via efficient Shapley value estimation. Future Generation Computer Systems, 147(??):339–349, October 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23001826.

Bezerra:2020:PEE

[dMBPdSC20] Jeandro de M. Bezerra, Antonio Janael Pinheiro, Críston P. de Souza, and Divanilson R. Campelo. Performance evaluation of elephant flow predictors in data center networking. Future Generation Computer Systems, 102(??):952–964, January 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18330309.

Desolda:2020:MDS

[DML20] Giuseppe Desolda, Maristella Matera, and Rosa Lanzilotti. Metamorphic data sources: a usercentric paradigm to consume linked data in interactive workspaces. Future Generation Computer Systems, 102(??):992–1015, January 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19303607.

Djenouri:2023:FDL

[DML23] Youcef Djenouri, Tomasz P.

Michalak, and Jerry Chun-Wei Lin. Federated deep learning for smart city edge-based applications. Future Generation Computer Systems, 147(??):350–359, October 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23001747.

Diamantini:2023:PAI

[DMPS23] Claudia Diamantini, Alex Mircoli, Domenico Potena, and Emanuele Storti. Processaware IIoT Knowledge Graph: a semantic model for industrial IoT integration and analytics. Future Generation Computer Systems, 139(??):224–238, February 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2200320X.

Delgado-Mohatar:2020:BBS

[DMSCA20] Oscar Delgado-Mohatar, José María Sierra-Cámara, and Eloy Anguiano. Blockchainbased semi-autonomous ransomware. Future Generation Computer Systems, 112(??):589–603, November 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19317406.

Das:2021:GOC

[DNNG21] Ankur Das, Janmenjoy Nayak, Bighnaraj Naik, and Uttam Ghosh. Generation of overlapping clusters constructing suitable graph for crime report analysis. Future Generation Computer Systems, 118(??): 339–357, May 2021. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21000376.

Dessi:2021:GKG

[DOR^{+}21] Danilo Dessì, Francesco Osborne, Diego Reforgiato Recupero, Davide Buscaldi, and Enrico Motta. Generating knowledge graphs by employing natural language processing and machine learning techniques within the scholarly domain. Future Generation Computer Systems, 116(??):253–264, March 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2033003X.

Din:2019:SHM

[DP19] Sadia Din and Anand Paul.

Smart health monitoring and management system: Toward autonomous wearable sensing for Internet of Things using big data analytics. Future Generation Computer Systems, 91(??):611–619, February 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X17315078.

See erratum [DP20c] and retraction notice [DP21b].

DAngelo:2020:KEB

[DP20a] Gianni D’Angelo and Francesco Palmieri. Knowledge elicitation based on genetic programming for non destructive testing of critical aerospace systems. Future Generation Computer Systems, 102(??):633–642, January 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19306193.

Desimoni:2020:EEL

[DP20b] Federico Desimoni and Laura Po. Empirical evaluation of linked data visualization tools. Future Generation Computer Systems, 112(??):258–282, November 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19303474.

Din:2020:ESH

[DP20c] Sadia Din and Anand Paul. Erratum to “Smart Health Monitoring and Management System: Toward autonomous wearable sensing for Internet of Things using Big Data Analytics” [Future Gener. Comput. Syst. 91 (2019) 611–619]. Future Generation Computer Systems, 111(??):939, October 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19317005.

See [DP19] and retraction notice [DP21a].

Din:2021:RNE

[DP21a] Sadia Din and Anand Paul. Retraction notice to Erratum to “Smart Health Monitoring and Management System: Toward autonomous wearable sensing for Internet of Things using Big Data Analytics [Future Gener. Comput. Syst. 91 (2019) 611–619]”’ [Future Gener. Comput. Syst. 108 (2019) 1350–1359]. Future Generation Computer Systems, 124(??):496, November 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21002600.

See [DP20c].

Din:2021:RNS

[DP21b] Sadia Din and Anand Paul. Retraction notice to “Smart Health Monitoring and Management System: Toward autonomous wearable sensing for Internet of Things using Big Data Analytics” [Future Gener. Comput. Syst. 91 (2019) 611–619]. Future Generation Computer Systems, 124(??):495, November 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21002612.

See [DP19].

Dahan:2020:ISG

[DPG20] Maytal Dahan, Rebecca Pirzl, and Sandra Gesing. International Science Gateways 2017 special issue. Future Generation Computer Systems, 110(??):320–322, September 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20313686.

Davarakis:2023:RLS

[DPLV23] Theodoros-Thirimachos Davarakis, Georgios Palaiokrassas, Antonios Litke, and Theodora Varvarigou. Reinforcement learning with smart contracts on blockchains. Future Generation Computer Systems, 148(??):550–563, November 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23002406.

Deepa:2022:SBB

[DPN^{+}22] N. Deepa, Quoc-Viet Pham, Dinh C. Nguyen, Sweta Bhattacharya, B. Prabadevi, Thippa Reddy Gadekallu, Praveen Kumar Reddy Maddikunta, Fang Fang, and Pubudu N. Pathirana. A survey on blockchain for big data: Approaches, opportunities, and future directions. Future Generation Computer Systems, 131(??):209–226, June 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22000243.

Del-Pozo-Punal:2023:SSC

[DPPGCCA23] Elías Del-Pozo-Puñal, Félix García-Carballeira, and Diego Camarmas-Alonso. A scalable simulator for cloud, fog and edge computing platforms with mobility support. Future Generation Computer Systems, 144(??):117–130, July 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23000511.

Drawel:2020:SAV

[DQBS20] Nagat Drawel, Hongyang Qu, Jamal Bentahar, and Elhadi Shakshuki. Specification and automatic verification of trust-based multiagent systems. Future Generation Computer Systems, 107(??):1047–1060, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17321349.

DeLaTorre:2020:DVS

[DRC20] Gonzalo De La Torre, Paul Rad, and Kim-Kwang Raymond Choo. Driverless vehicle security: Challenges and future research opportunities. Future Generation Computer Systems, 108(??):1092–1111, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17315066.

Dey:2020:SBC

[DRD20] Maitreyee Dey, Soumya Prakash Rana, and Sandra Dudley. Smart building creation in large scale HVAC environments through automated fault detection and diagnosis. Future Generation Computer Systems, 108(??):950–966, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17324287.

Frohlich:2024:AAN

[dRFRB24] William da Rosa Fröhlich, Sandro José Rigo, and Marta Rosecler Bez. ATHENA I — an architecture for near real-time physiological signal monitoring and pattern detection. Future Generation Computer Systems, 150(??):395–411, January 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23003357.

Righi:2020:EPI

[dRRCGdC20] Rodrigo da Rosa Righi, Everton Correa, Márcio Miguel Gomes, and Cristiano André da Costa. Enhancing performance of IoT applications with load prediction and cloud elasticity. Future Generation Computer Systems, 109(??):689–701, August 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17329229.

Datta:2023:BBS

[DS23] Sreemana Datta and Ditipriya Sinha. BSEIFFS: Blockchain-secured edgeintelligent forest fire surveillance. Future Generation Computer Systems, 147(??):59–76, October 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23001474.

Dehury:2020:CFC

[DSC20] Chinmaya Kumar Dehury, Satish Narayana Srirama, and Tek Raj Chhetri. CCo-DaMiC: a framework for coherent coordination of data migration and computation platforms. Future Generation Computer Systems, 109(??):1–16, August 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19330924.

Das:2024:UDP

[DSC24] Anirban Das, Navlika Singh, and Suchetana Chakraborty. UniPreCIS: a data preprocessing solution for collocated services on shared IoT. Future Generation Computer Systems, 153(??):543–557, April 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23004375.

DeMeester:2020:IIF

[DSDV20] Ben De Meester, Tom Seymoens, Anastasia Dimou, and Ruben Verborgh. Implementationindependent function reuse. Future Generation Computer Systems, 110(??): 946–959, September 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0167739X19303723.

daSilva:2021:SIW

[dSGST21] Rafael Ferreira da Silva, Sandra Gesing, Rizos Sakellariou, and Ian Taylor. Special issue on workflows in support of large-scale science. Future Generation Computer Systems, 118 (??):73–74, May 2021. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21000078.

Oliveira:2023:FRT

[dSOFC^{+}23] Marcos de S. Oliveira, Francisco Erivaldo Fernandes, Lukas Cerveny, Flávia Akemi Miyazaki, Leonardo Valeriano Neri, Alan da Silva, Beatriz Leandro Bonafini, Victor Medeiros Outtes Alves, and Órion Darshan Winter de Lima. FastAiAlloc: a real-time multi-resources allocation framework proposal based on predictive model and multiple optimization strategies. Future Generation Computer Systems, 149(??):622–636, December 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23003126.

Delgado-Segura:2020:FPD

[DSPSNAHJ20] Sergi Delgado-Segura, Cristina Pérez-Solà, Guillermo Navarro-Arribas, and Jordi Herrera-Joancomartí. A fair protocol for data trading based on Bitcoin transactions. Future Generation Computer Systems, 107(??):832–840, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17318344.

Divan:2022:MPI

[DSRG22] Mario José Diván, María Laura Sánchez-Reynoso, and Silvio Miguel Gonnet. Measurement project interoperability for real-time data gathering systems. Future Generation Computer Systems, 129(??):298–314, April 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21004738.

Ding:2020:BOC

[DSW^{+}20] Yaoling Ding, Ying Shi, An Wang, Yongjuan Wang, and Guoshuang Zhang. Block-oriented correlation power analysis with bitwise linear leakage: an artificial intelligence approach based on genetic algorithms. Future Generation Computer Systems, 106 (??):34–42, May 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19312816.

Dan:2021:KWS

[DT21] Yufang Dan and Jianwen Tao. Knowledge worker scheduling optimization model based on bacterial foraging algorithm. Future Generation Computer Systems, 124(??):330–337, November 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21001783.

Lopez-de-Teruel:2020:PPL

[dTGC20] Pedro E. Lopez de Teruel, Felix J. Garcia, and Oscar Canovas. Practical passive localization system based on wireless signals for fast deployment of occupancy services. Future Generation Computer Systems, 107(??):692–704, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17320083.

Dutta:2022:AFC

[Dut22]  Nitul Dutta.  An ap-

proach for FIB construction and interest packet forwarding in information centric network. Future Generation Computer Systems, 130(??):269–278, May 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22000139.

Diaz-Verdejo:2020:MCE

[DVEE^{+}20] Jesús E. Díaz-Verdejo, Antonio Estepa, Rafael Estepa, German Madinabeitia, and Fco. Javier Muñoz-Calle. A methodology for conducting efficient sanitization of HTTP training datasets. Future Generation Computer Systems, 109(??):67–82, August 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19322629.

deVos:2024:DCR

[dVIP24] Martijn de Vos, Georgy Ishmaev, and Johan Pouwelse. DeScan : Censorshipresistant indexing and search for Web3. Future Generation Computer Systems, 152(??):257–272, March 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23004132.

DeLaIglesia:2020:NIA

[DVV^{+}20] Daniel H. De La Iglesia, Gabriel Villarrubia González, Marcelo Vallejo García, Alfonso José López Rivero, and Juan F. De Paz. Non-invasive automatic beef carcass classification based on sensor network and image analysis. Future Generation Computer Systems, 113(??):318–328, December 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19317492.

Ding:2023:HHD

[DWL^{+}23] Guohui Ding, Yankai Wang, Chenyang Li, Haohan Sun, Cailong Li, Lei Wang, Haijun Yin, and Tiantian Huang. HSCFC: Highdimensional streaming data clustering algorithm based on feedback control system. Future Generation Computer Systems, 146(??):156–165, September 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23001425.

Ding:2020:CSC

[DWZ20] Weilong Ding, Xuefei Wang, and Zhuofeng Zhao. CO-STAR: a collaborative prediction service for shortterm trends on continuous spatio-temporal data. Future Generation Computer Systems, 102(??):481–493, January 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18332369.

Du:2023:TPI

[DZB23] Zhihui Du, Sen Zhang, and David A. Bader. Tunnel: Parallel-inducing sort for large string analytics. Future Generation Computer Systems, 149(??):650–663, December 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23003096.

Dong:2021:HGA

[DZXS21] Xueshi Dong, Hong Zhang, Min Xu, and Fanfan Shen. Hybrid genetic algorithm with variable neighborhood search for multi-scale multiple bottleneck traveling salesmen problem. Future Generation Computer Systems, 114(??):229–242, January 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19330742.

Elaziz:2021:AOT

[EAA21] Mohamed Abd Elaziz, Laith Abualigah, and Ibrahim Attiya. Advanced optimization technique for scheduling IoT tasks in cloud-fog computing environments. Future Generation Computer Systems, 124(??):142–154, November 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2100176X.

Ejarque:2022:EDI

[EBA^{+}22] Jorge Ejarque, Rosa M. Badia, Loïc Albertin, Giovanni Aloisio, Enrico Baglione, Yolanda Becerra, Stefan Boschert, Julian R. Berlin, Alessandro D’Anca, Donatello Elia, François Exertier, Sandro Fiore, José Flich, Arnau Folch, Steven J. Gibbons, Nikolay Koldunov, Francesc Lordan, Stefano Lorito, Finn Løvholt, Jorge Macías, Fabrizio Marozzo, Alberto Michelini, Marisol Monterrubio-Velasco, Marta Pienkowska, Josep de la Puente, Anna Queralt, Enrique S. Quintana-Ortí, Juan E. Rodríguez, Fabrizio Romano, Riccardo Rossi, Jedrzej Rybicki, Miroslaw Kupczyk, Jacopo Selva, Domenico Talia, Roberto Tonini, Paolo Trunfio, and Manuela Volpe. Enabling dynamic and intelligent workflows for HPC, data analytics, and AI convergence. Future Generation Computer Systems, 134(??):414–429, September 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/S0167739X22001364.

Elshazly:2021:TEA

[EELB21] Hatem Elshazly, Jorge Ejarque, Francesc Lordan, and Rosa M. Badia. Towards enabling I/O awareness in task-based programming models. Future Generation Computer Systems, 121(??):74–89, August 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21000893.

Elhabbash:2024:PAS

[EEN^{+}24] Abdessalam Elhabbash, Yehia Elkhatib, Vatsala Nundloll, Vicent Sanz Marco, and Gordon S. Blair. Principled and automated system of systems composition using an ontological architecture. Future Generation Computer Systems, 157(??):499–515, August 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X24001055.

Enes:2020:RTR

[EET20] Jonatan Enes, Roberto R. Expósito, and Juan Touriño. Real-time resource scaling platform for Big Data workloads on serverless environments. Future Generation Computer Systems, 105 (??):361–379, April 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167739X19310015.

Exposito:2024:BMA

[EGD24] Roberto R. Expósito and Jorge González-Domínguez. BigDEC: a multi-algorithm big data tool based on the k-mer spectrum method for scalable short-read error correction. Future Generation Computer Systems, 154(??):314–329, May 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24000116.

Exposito:2020:SSB

[EGDT20] Roberto R. Expósito, Jorge González-Domínguez, and Juan Touriño. SMusket: Spark-based DNA error correction on distributedmemory systems. Future Generation Computer Systems, 111(??):698–713, October 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19313731.

Espinosa:2022:MOE

[EJP22] Raquel Espinosa, Fernando Jiménez, and José Palma. Multi-objective evolutionary spatio-temporal forecasting of air pollution. Future Generation Computer Systems, 136(??):15– 33, November 2022. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22001911.

Ectors:2020:ZPL

[EKJ^{+}20] Wim Ectors, Bruno Kochan, Davy Janssens, Tom Bellemans, and Geert Wets. Zipf’s power law in activity schedules and the effect of aggregation. Future Generation Computer Systems, 107(??):1014–1025, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17321891.

Eun:2023:AIB

[EKK23] Sung-Jong Eun, Eun Joung Kim, and JungYoon Kim. Artificial intelligence-based personalized serious game for enhancing the physical and cognitive abilities of the elderly. Future Generation Computer Systems, 141(??):713–722, April 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22004216.

Errounda:2021:CLS

[EL21] Fatima Zahra Errounda and Yan Liu. Collective location statistics release with local differential privacy. Future Generation Computer Systems, 124(??):174–186, November 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21001709.

Errounda:2023:ADP

[EL23] Fatima Zahra Errounda and Yan Liu. Adaptive differential privacy in vertical federated learning for mobility forecasting. Future Generation Computer Systems, 149(??):531–546, December 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2300290X.

Elghamrawy:2020:SCR

[Elg20] Sally M. Elghamrawy. Security in cognitive radio network: Defense against primary user emulation attacks using genetic artificial bee colony (GABC) algorithm. Future Generation Computer Systems, 109(??):479–487, August 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17321246.

Eom:2020:DAA

[ELS20] Jin Ki Eom, Kwang-Sub Lee, and Myeong-Eon Seong. Development and application of the Activity-BAsed Traveler Analyzer (ABATA) system. Future Generation Computer Systems, 106(??):135–153, May 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19308842.

Eskandari:2021:SIS

[EMHE21] Leila Eskandari, Jason Mair, Zhiyi Huang, and David Eyers. I-Scheduler: Iterative scheduling for distributed stream processing systems. Future Generation Computer Systems, 117(??):219–233, April 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20330351.

Ekane:2022:FAN

[ENT^{+}22] Brice Ekane, Tu Dinh Ngoc, Boris Teabe, Daniel Hagimont, and Noel De Palma.

FlexVF: Adaptive network device services in a virtualized environment. Future Generation Computer Systems, 127(??):14–22, February 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21003198.

Evans:2022:EPC

[ERG^{+}22] Theodore Evans, Carl Orge Retzlaff, Christian Geißler, Michaela Kargl, Markus Plass, Heimo Müller, Tim-Rasmus Kiehl, Norman Zerbe, and Andreas Holzinger. The explainability paradox: Challenges for xAI in digital pathology. Future Generation Computer Systems, 133(??):281–296, August 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22000838.

Eccles:2024:DED

[ERK^{+}24] Bailey J. Eccles, Philip Rodgers, Peter Kilpatrick, Ivor Spence, and Blesson Varghese. DNNShifter: an efficient DNN pruning system for edge computing. Future Generation Computer Systems, 152(??):43– 54, March 2024. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23003576.

Ectors:2020:OCA

[ERL^{+}20] Wim Ectors, Sofie Reumers, Won Do Lee, Bruno Kochan, Davy Janssens, Tom Bellemans, and Geert Wets. Optimizing copious activity type classes based on classification accuracy and entropy retention. Future Generation Computer Systems, 110(??):338–349, September 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18301675.

El-Sappagh:2021:ADP

[ESSS^{+}21] Shaker El-Sappagh, Hager Saleh, Radhya Sahal, Tamer Abuhmed, S. M. Riazul Islam, Farman Ali, and Eslam Amer. Alzheimer’s disease progression detection model based on an early fusion of cost-effective multimodal data. Future Generation Computer Systems, 115(??):680–699, February 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20329824.

Enokido:2020:SII

[ETH20] Tomoya Enokido, David Taniar, and Omar Khadeer Hussain. Special issue: Intelligent edge, fog and Internet of Things (IoT)- based services. Future Generation Computer Systems, 109(??):710–711, August 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20309687.

Erdol:2024:LDS

[EUEU24] Eda Sena Erdol, Beste Ustubioglu, Hakan Erdol, and Guzin Ulutas. Low dimensional secure federated learning framework against poisoning attacks. Future Generation Computer Systems, 158(??):183–199, September 2024. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24001444.

Folch:2023:ECE

[FAA^{+}23] Arnau Folch, Claudia Abril, Michael Afanasiev, Giorgio Amati, Michael Bader, Rosa M. Badia, Hafize B. Bayraktar, Sara Barsotti, Roberto Basili, Fabrizio Bernardi, Christian Boehm, Beatriz Brizuela, Federico Brogi, Eduardo Cabrera, Emanuele Casarotti, Manuel J. Castro, Matteo Cerminara, Antonella Cirella, Alexey Cheptsov, Javier Conejero, Antonio Costa, Marc de la Asunción, Josep de la Puente, Marco Djuric, Ravil Dorozhinskii, Gabriela Espinosa, Tomaso Esposti-Ongaro, Joan Farnós, Nathalie Favretto-Cristini, Andreas Fichtner, Alexandre Fournier, Alice-Agnes Gabriel, Jean-Matthieu Gallard, Steven J. Gibbons, Sylfest Glimsdal, José Manuel González-Vida, Jose Gracia, Rose Gregorio, Natalia Gutierrez, Benedikt Halldorsson, Okba Hamitou, Guillaume Houzeaux, Stephan Jaure, Mouloud Kessar, Lukas Krenz, Lion Krischer, Soline Laforet, Piero Lanucara, Bo Li, Maria Concetta Lorenzino, Stefano Lorito, Finn Løvholt, Giovanni Macedonio, Jorge Macías, Guillermo Marín, Beatriz Martínez Montesinos, Leonardo Mingari, Geneviève Moguilny, Vadim Montellier, Marisol Monterrubio-Velasco, Georges Emmanuel Moulard, Masaru Nagaso, Massimo Nazaria, Christoph Niethammer, Federica Pardini, Marta Pienkowska, Luca Pizzimenti, Natalia Poiata, Leonhard Rannabauer, Otilio Rojas, Juan Esteban Rodriguez, Fabrizio Romano, Oleksandr Rudyy, Vittorio Ruggiero, Philipp Samfass, Carlos Sánchez-Linares, Sabrina Sanchez, Laura Sandri, Antonio Scala, Nathanael Schaeffer, Joseph Schuchart, Jacopo Selva, Amadine Sergeant, Angela Stallone, Matteo Taroni, Solvi Thrastarson, Manuel Titos, Nadia Tonelllo, Roberto Tonini, Thomas Ulrich, Jean-Pierre Vilotte, Malte Vöge, Manuela Volpe, Sara Aniko Wirp, and Uwe Wössner. The EU Center of Excellence for Exascale in Solid Earth (ChEESE): Implementation, results, and roadmap for the second phase. Future Generation Computer Systems, 146(??):47–61, September 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23001401.

Faiz:2020:PLL

[FAAS20] Mohamed Falah Faiz, Junaid Arshad, Mamoun Alazab, and Andrii Shalaginov. Predicting likelihood of legitimate data loss in email DLP. Future Generation Computer Systems, 110(??):744–757, September 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/S0167739X19314943.

Faez:2021:FGC

[Fae21] Telli Faez. Future Generation Computer Systems: Publisher’s note: In recognition of Prof. Peter Sloot on his retirement as Editorin-Chief. Future Generation Computer Systems, 116 (??):iii, March 2021. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20330545.

Fernando:2020:TLS

[FAŞ^{+}20] Senaka Fernando, Julio Amador Díaz López, Ovidiu Şerban, Juan Gómez-Romero, Miguel Molina-Solana, and Yike Guo. Towards a Large-Scale Twitter Observatory for political events. Future Generation Computer Systems, 110(??):976–983, September 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19309720.

Fan:2020:DDR

[FBL^{+}20] Kuan Fan, Zijian Bao, Mingxi Liu, Athanasios V. Vasilakos, and Wenbo Shi. Dredas: Decentralized, reliable and efficient remote outsourced data auditing scheme with blockchain smart contract for industrial IoT. Future Generation Computer Systems, 110(??):665–674, September 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19315560.

Faroughi:2023:ASA

[FBTJ23] Azadeh Faroughi, Reza Boostani, Hadi Tajalizadeh, and Reza Javidan. ARD-Stream: an adaptive radius density-based stream clustering. Future Generation Computer Systems, 149(??):416–431, December 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23002777.

Felici-Castell:2021:ALV

[FCGPSG^{+}21] Santiago Felici-Castell, Miguel García-Pineda, Jaume Segura-Garcia, Rafael Fayos-Jordan, and Jesus Lopez-Ballester. Adaptive live video streaming on low-cost wireless multihop networks for road traffic surveillance in smart cities. Future Generation Computer Systems, 115(??):741–755, February 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20329873.

Fernandez-Cerero:2021:DDS

[FCOJFM21] Damián Fernández-Cerero, F. Javier Ortega, Agnieszka Jakóbik, and Alejandro Fernández-Montes. DIS-CERNER: Dynamic selection of resource manager in hyper-scale cloudcomputing data centres. Future Generation Computer Systems, 116(??): 190–199, March 2021. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20330156.

FayezEliyan:2021:DAS

[FD21] Lubna Fayez Eliyan and Roberto Di Pietro. DoS and DDoS attacks in Software Defined Networks: a survey of existing solutions and research challenges. Future Generation Computer Systems, 122(??):149–171, September 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21000911.

Fernandez-Fabeiro:2020:AOH

[FFAFD20] Jorge Fernández-Fabeiro, Diego Andrade, Basilio B. Fraguela, and Ramón Doallo. An automatic optimizer for heterogeneous devices. Future Generation Computer Systems, 106(??):572–584, May 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167739X18324294.

Filho:2020:ALI

[FFAW20] João Fabrício Filho, Isaías B.

Felzmann, Rodolfo Azevedo, and Lucas F. Wanner. AxRAM: a lightweight implicit interface for approximate data access. Future Generation Computer Systems, 113(??):556–570, December 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20301060.

Forti:2020:SCE

[FFB20] Stefano Forti, Gian-Luigi Ferrari, and Antonio Brogi. Secure cloud-edge deployments, with trust. Future Generation Computer Systems, 102(??):775–788, January 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19301268.

Fortino:2020:MTB

[FFM^{+}20] G. Fortino, L. Fotia, F. Messina, D. Rosaci, and G. M. L. Sarné. A meritocratic trust-based group formation in an IoT environment for smart cities. Future Generation Computer Systems, 108(??):34– 45, July 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19334375.

Forti:2021:CLS

[FGB21a] Stefano Forti, Marco Gaglianese, and Antonio Brogi. Corrigendum to “Lightweight self-organising distributed monitoring of Fog infrastructures” [Future Gener. Comput. Syst. 114 (2020) 605–618]. Future Generation Computer Systems, 118(??):495, May 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21000194.

See [FGB21b].

Forti:2021:LSO

[FGB21b] Stefano Forti, Marco Gaglianese, and Antonio Brogi. Lightweight self-organising distributed monitoring of fog infrastructures. Future Generation Computer Systems, 114(??):605–618, January 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19334582.

See corrigendum [FGB21a].

Francia:2021:MDP

[FGG^{+}21] Matteo Francia, Enrico Gallinucci, Matteo Golfarelli, Anna Giulia Leoni, Stefano Rizzi, and Nicola Santolini. Making data platforms smarter with MOSES. Future Generation Computer Systems, 125(??):299–313, December 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21002260.

Franco:2023:DMS

[FGG^{+}23] Mirko Franco, Ombretta Gaggi, Barbara Guidi, Andrea Michienzi, and Claudio E. Palazzi. A decentralised messaging system robust against the unauthorised forwarding of private content. Future Generation Computer Systems, 145(??):211–222, August 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23001024.

Fanitabasi:2020:SIT

[FGP20] Farzam Fanitabasi, Edward Gaere, and Evangelos Pournaras. A selfintegration testbed for decentralized socio-technical systems. Future Generation Computer Systems, 113(??):541–555, December 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20303435.

Francia:2023:HFH

[FGP23] Matteo Francia, Joseph Giovanelli, and Giuseppe Pisano. HAMLET: a framework for humancentered AutoML via structured argumentation. Future Generation Computer Systems, 142(??):182–194, May 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22004393.

Fuchs:2020:SFC

[FHGF20] Klaus Fuchs, Mirella Haldimann, Tobias Grundmann, and Elgar Fleisch. Supporting food choices in the Internet of People: Automatic detection of diet-related activities and display of real-time interventions via mixed reality headsets. Future Generation Computer Systems, 113(??):343–362, December 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20303411.

Fumanal-Idocin:2020:CDS

[FIABC^{+}20] J. Fumanal-Idocin, A. Alonso-Betanzos, O. Cordón, H. Bustince, and M. Minárová. Community detection and social network analysis based on the Italian wars of the 15th century. Future Generation Computer Systems, 113(??):25–40, December 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20305057.

Feng:2021:EVG

[FLF^{+}21] Wenjie Feng, Shenghua Liu, Christos Faloutsos, Bryan Hooi, Huawei Shen, and Xueqi Cheng. EagleMine: Vision-guided microclusters recognition and collective anomaly detection. Future Generation Computer Systems, 115(??):236–250, February 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20316794.

Feng:2020:MMH

[FLG^{+}20] Wendi Feng, Chuanchang Liu, Zehua Guo, Thar Baker, Gang Wang, Meng Wang, Bo Cheng, and Junliang Chen. MobiGyges: a mobile hidden volume for preventing data loss, improving storage utilization, and avoiding device reboot. Future Generation Computer Systems, 109(??): 158–171, August 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1932521X.

Feng:2024:LBL

[FLH^{+}24] Libo Feng, Yifan Liu, Kai Hu, Xue Zeng, Fake Fang, Jiale Xie, and Shaowen Yao. LPP-BPSI: a location privacy-preserving scheme using blockchain and Private Set Intersection in spatial crowdsourcing. Future Generation Computer Systems, 157(??):112–123, August 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24001092.

Fan:2020:FGA

[FLTQ20] Yongkai Fan, Shengle Liu, Gang Tan, and Fei Qiao. Fine-grained access control based on Trusted Execution Environment. Future Generation Computer Systems, 109(??):551–561, August 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17319362.

Filho:2020:FES

[FMM^{+}20] Geraldo P. Rocha Filho, Rodolfo I. Meneguette, Guilherme Maia, Gustavo Pessin, Vinícius P. Gonçalves, Li Weigang, Jó Ueyama, and Leandro A. Villas. A fogenabled smart home solution for decision-making using smart objects. Future Generation Computer Systems, 103(??):18–27, February 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18322301.

Flammini:2020:SIT

[FMN^{+}20] Francesco Flammini, Stefano Marrone, Roberto Nardone, Mauro Caporuscio, and Mirko D’Angelo. Safety integrity through self-adaptation for multisensor event detection: Methodology and casestudy. Future Generation Computer Systems, 112(??):965–981, November 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19333734.

Fazili:2020:MAG

[FNRP20] Yashar Fazili, Alireza Nafarieh, Bill Robertson, and William Phillips. MultiSLA-aware green routing mechanism for GMPLS. Future Generation Computer Systems, 107(??): 705–715, June 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1731899X.

Falah:2021:IMP

[FPH^{+}21] Ahmed Falah, Lei Pan, Shamsul Huda, Shiva Raj Pokhrel, and Adnan Anwar. Improving malicious PDF classifier with feature engineering: a datadriven approach. Future Generation Computer Systems, 115(??):314–326, February 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20305082.

Fu:2024:SOC

[FPL24] Xiankun Fu, Li Pan, and Shijun Liu. To store or not: Online cost optimization for running big data jobs on the cloud. Future Generation Computer Systems, 156(??):42– 52, July 2024. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24000803.

Ferrer:2021:ACA

[FPMJ21] Ana Juan Ferrer, Javier Panadero, Joan-Manuel Marques, and Josep Jorba. Admission control for adhoc edge cloud. Future Generation Computer Systems, 114(??):548–562, January 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19329164.

Feng:2024:CCD

[FQH^{+}24] Libo Feng, Fei Qiu, Kai Hu, Bei Yu, Junyu Lin, and Shaowen Yao. CABC: a cross-domain authentication method combining blockchain with certificateless signature for IIoT. Future Generation Computer Systems, 158(??):516–529, September 2024. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24001717.

Fontenla-Romero:2023:FFH

[FRGBHPPS23] Oscar Fontenla-Romero, Bertha Guijarro-Berdiñas, Elena Hernández-Pereira, and Beatriz Pérez-Sánchez. FedHEONN: Federated and homomorphically encrypted learning method for one-layer neural networks. Future Generation Computer Systems, 149(??):200–211, December 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23002686.

Fang:2021:MRT

[FS21] Lei Fang and Mingsi Sun. Motion recognition technology of badminton players in sports video images. Future Generation Computer Systems, 124(??):381–389, November 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21001862.

Fernando:2020:OVE

[FSBŞ^{+}20] Senaka Fernando, James Scott-Brown, Ovidiu Şerban, David Birch, David Akroyd, Miguel Molina-Solana, Thomas Heinis, and Yike Guo. Open Visualization Environment (OVE): a web framework for scalable rendering of data visualizations. Future Generation Computer Systems, 112(??):785–799, November 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1932922X.

Felzmann:2020:ADL

[FSD^{+}20] Isaías B. Felzmann, Matheus M.

Susin, Liana Duenha, Rodolfo Azevedo, and Lucas F. Wanner. ADeLe: a description language for approximate hardware. Future Generation Computer Systems, 102(??):245–258, January 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19304388.

Fireman:2024:PRE

[FSP^{+}24] Daniel Fireman, Paulo Silva, Thiago Emmanuel Pereira, Luis Mafra, and Dalton Valadares. Prebaking runtime environments to improve the FaaS cold start latency. Future Generation Computer Systems, 155(??):287–299, June 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24000190.

Fakhfakh:2020:MPD

[FTM20] Faten Fakhfakh, Mohamed Tounsi, and Mohamed Mosbah. Modeling and proving distributed algorithms for dynamic graphs. Future Generation Computer Systems, 108(??):751–761, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19312245.

Filinis:2024:IDO

[FTS^{+}24] Nikos Filinis, Ioannis Tzanettis, Dimitrios Spatharakis, Eleni Fotopoulou, Ioannis Dimolitsas, Anastasios Zafeiropoulos, Constantinos Vassilakis, and Symeon Papavassiliou. Intentdriven orchestration of serverless applications in the computing continuum. Future Generation Computer Systems, 154(??):72– 86, May 2024. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23004910.

Feng:2022:BMS

[FW22] Yanhong Feng and Gai-Ge Wang. A binary moth search algorithm based on self-learning for multidimensional knapsack problems. Future Generation Computer Systems, 126(??):48–64, January 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21002958.

Fan:2021:DBC

[FWP21] Yuqi Fan, Huanyu Wu, and Hye-Young Paik. DR– BFT: a consensus algorithm for blockchain-based multi-layer data integrity framework in dynamic edge computing system. Future Generation Computer Systems, 124(??):33–48, November 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21001412.

Fu:2023:GCS

[FWX23] Zixuan Fu, Chenghua Wang, and Jiajie Xu. Graph contextualized selfattention network for software service sequential recommendation. Future Generation Computer Systems, 149(??):509–517, December 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23002972.

Fang:2022:DWE

[FWY^{+}22] Kai Fang, Tingting Wang, Xiaochen Yuan, Chunyu Miao, Yuanyuan Pan, and Jianqing Li. Detection of weak electromagnetic interference attacks based on fingerprint in IIoT systems. Future Generation Computer Systems, 126(??):295–304, January 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21003289.

Fan:2020:HVS

[FWZ^{+}20] Xingyue Fan, Ting Wu, Qiuhua Zheng, Yuanfang Chen, Muhammad Alam, and Xiaodong Xiao. HSE-Voting: a secure highefficiency electronic voting scheme based on homomorphic signcryption. Future Generation Computer Systems, 111(??):754–762, October 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1931951X.

Fan:2024:MMF

[FXZZ24] Yuwei Fan, Wei Xi, Hengyi Zhu, and Jizhong Zhao. MiniPFL: Mini federations for hierarchical personalized federated learning. Future Generation Computer Systems, 157(??):41– 50, August 2024. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24001006.

Feng:2024:BSB

[FYHZ24] Jingyu Feng, Rui Yan, Gang Han, and Wenbo Zhang. BDPM: a secure batch dynamic password management scheme in industrial internet environments. Future Generation Computer Systems, 157(??):193–209, August 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2400102X.

Feukeu:2020:DBS

[FZ20] E. A. Feukeu and T. Zuva.

Dynamic broadcast storm mitigation approach for VANETs. Future Generation Computer Systems, 107(??):1097–1104, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17315042.

Feng:2020:TRH

[FZC^{+}20] Jingyu Feng, Xinyu Zhao, Kexuan Chen, Feng Zhao, and Guanghua Zhang. Towards random-honest miners selection and multiblocks creation: Proofof-negotiation consensus mechanism in blockchain networks. Future Generation Computer Systems, 105(??):248–258, April 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167739X19313044.

Fu:2023:BED

[FZT^{+}23] Luxia Fu, Zhuang Zhang, Liang Tan, Zhengyi Yao, Hongxin Tan, Jingxue Xie, and Kun She. Blockchainenabled device command operation security for Industrial Internet of Things. Future Generation Computer Systems, 148(??): 280–297, November 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167739X23002224.

Gajaria:2022:EPE

[GA22] Dhruv Gajaria and Tosiron Adegbija. Evaluating the performance and energy of STT-RAM caches for real-world wearable workloads. Future Generation Computer Systems, 136(??):231–240, November 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22001935.

Garcia:2021:DFD

[GAA^{+}21] Jordi Garcia, Francesc Aguiló, Adrià Asensio, Ester Simó, Marisa Zaragozá, and Xavi Masip-Bruin. Data-flow driven optimal tasks distribution for global heterogeneous systems. Future Generation Computer Systems, 125(??):792–805, December 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21002806.

Gimenez-Aguilar:2021:ACB

[GAdFGMA21] Mar Gimenez-Aguilar, Jose Maria de Fuentes, Lorena Gonzalez-Manzano, and David Arroyo. Achieving cybersecurity in blockchain-based systems: a survey. Future Generation Computer Systems, 124(??):91–118, November 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21001576.

Garcia:2024:ASS

[GAP24] Rodrigo García, Jose Aguilar, and Angel Pinto. An autonomous system for the self-supervision of animal fattening in the context of precision livestock farming. Future Generation Computer Systems, 150(??):220–231, January 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X2300331X.

Gasparetti:2022:DPR

[Gas22]  Fabio Gasparetti.  Dis-

covering prerequisite relations from educational documents through word embeddings. Future Generation Computer Systems, 127(??):31–41, February 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21003290.

Gbadouissa:2020:HHB

[GAT^{+}20] Jocelyn Edinio Zacko Gbadouissa, Ado Adamou Abba Ari, Chafiq Titouna, Abdelhak Mourad Gueroui, and Ousmane Thiare. HGC: HyperGraph based clustering scheme for power aware wireless sensor networks. Future Generation Computer Systems, 105(??): 175–183, April 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1932240X.

Gu:2020:EAW

[GB20] Yi Gu and Chandu Budati. Energy-aware workflow scheduling and optimization in clouds using bat algorithm. Future Generation Computer Systems, 113(??):106–112, December 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19317066.

Guldner:2024:DER

[GBC^{+}24] Achim Guldner, Rabea Bender, Coral Calero, Giovanni S. Fernando, Markus Funke, Jens Gröger, Lorenz M. Hilty, Julian Hörnschemeyer, Geerd-Dietger Hoffmann, Dennis Junger, Tom Kennes, Sandro Kreten, Patricia Lago, Franziska Mai, Ivano Malavolta, Julien Murach, Kira Obergöker, Benno Schmidt, Arne Tarara, Joseph P. De Veaugh-Geiss, Sebastian Weber, Max Westing, Volker Wohlgemuth, and Stefan Naumann. Development and evaluation of a reference measurement model for assessing the resource and energy efficiency of software products and components

— Green Software Measurement Model (GSMM). Future Generation Computer Systems, 155(??): 402–418, June 2024. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X24000384.

Garcia-Blas:2020:TEM

[GBdRACG20] Javier Garcia-Blas, David del Rio Astorga, Jesus Carretero, and J. Daniel Garcia. Towards enhanced MRI by using a multiple back end programming framework. Future Generation Computer Systems, 112(??):467–477, November 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19318199.

Gong:2023:DTO

[GBH^{+}23] Yanqi Gong, Kun Bian, Fei Hao, Yifei Sun, and Yulei Wu. Dependent tasks offloading in mobile edge computing: a multiobjective evolutionary optimization strategy. Future Generation Computer Systems, 148(??):314–325, November 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23002364.

Garg:2020:EDI

[GBK20] Neenu Garg, Seema Bawa, and Neeraj Kumar. An efficient data integrity auditing protocol for cloud computing. Future Generation Computer Systems, 109(??):306–316, August 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19319661.

Genez:2020:TDS

[GBM20] Thiago A. L. Genez, Luiz F.

Bittencourt, and Edmundo R. M. Madeira. Timediscretization for speedingup scheduling of deadlineconstrained workflows in clouds. Future Generation Computer Systems, 107(??):1116–1129, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17316679.

Gurgul:2023:CNA

[GBP23] Grzegorz Gurgul, Bartosz Baliś, and Maciej Paszyński. Cloud-native alternating directions solver for isogeometric analysis. Future Generation Computer Systems, 140(??): 151–172, March 2023. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22003338.

Gomez-Carmona:2020:ECC

[GCCMK^{+}20] Oihane Gómez-Carmona, Diego Casado-Mansilla, Frank Alexander Kraemer, Diego López de Ipiña, and Javier García-Zubia. Exploring the computational cost of machine learning at the edge for human-centric Internet of Things. Future Generation Computer Systems, 112(??):670–683, November 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20304106.

Gong:2022:ATP

[GCH^{+}22] Bei Gong, Chi Cui, Mingsheng Hu, Chong Guo, Xiaochong Li, and Yuheng Ren. Anonymous traceability protocol based on group signature for blockchain. Future Generation Computer Systems, 127(??): 160–167, February 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167739X21003666.

Garcia-Cobo:2021:DLQ

[GCM21] Iván García-Cobo and Héctor D. Menéndez. Designing large quantum key distribution networks via medoid-based algorithms. Future Generation Computer Systems, 115(??):814–824, February 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19332339.

Gao:2024:RRB

[GCN^{+}24] Xin Gao, Naiyuan Cui, Jiawei Nian, Zongnan Liang, Jiaxuan Gao, Hongjin Liu, and Mengfei Yang. ReBEC: a replacementbased energy-efficient faulttolerance design for associative caches. Future Generation Computer Systems, 155(??):39–52, June 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24000244.

Guarascio:2022:BCT

[GCPM22] Massimo Guarascio, Nunziato Cassavia, Francesco Sergio Pisani, and Giuseppe Manco. Boosting cyberthreat intelligence via collaborative intrusion detection. Future Generation Computer Systems, 135(??):30–43, October 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22001571.

G:2020:ODA

[GCT^{+}20] S. S. Chalapathi G., Vinay Chamola, Chen-Khong Tham, Gurunarayanan S., and Nirwan Ansari. An optimal delay aware task assignment scheme for wireless SDN networked edge cloudlets. Future Generation Computer Systems, 102(??):862–875, January 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19301232.

Garcia-Diaz:2021:DMS

[GDCGCPVG21] José Antonio García-Díaz, Mar Cánovas-García, Ricardo Colomo-Palacios, and Rafael Valencia-García. Detecting misogyny in Spanish tweets. An approach based on linguistics features and word embeddings. Future Generation Computer Systems, 114(??):506–518, January 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20301928.

Garcia-Diaz:2020:ODA

[GDCGVG20] José Antonio García-Díaz, Mar Cánovas-García, and Rafael Valencia-García. Ontology-driven aspectbased sentiment analysis classification: an infodemiological case study regarding infectious diseases in Latin America. Future Generation Computer Systems, 112(??):641–657, November 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2030892X.

Garcia-Diaz:2022:PTI

[GDCPVG22] José Antonio García-Díaz, Ricardo Colomo-Palacios, and Rafael Valencia-García. Psychographic traits identification based on political ideology: an author analysis study on Spanish politicians’ tweets posted in 2020. Future Generation Computer Systems, 130 (??):59–74, May 2022. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21004921.

Gonzalez-Dominguez:2020:CJA

[GDEBC20] Jorge González-Domínguez, Roberto R. Expósito, and Verónica Bolón-Canedo. CUDA-JMI: Acceleration of feature selection on heterogeneous systems. Future Generation Computer Systems, 102(??):426–436, January 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19312968.

Giatrakos:2020:OOD

[GDGK20] Nikos Giatrakos, Antonios Deligiannakis, Minos Garofalakis, and Yannis Kotidis. Omnibus outlier detection in sensor networks using windowed locality sensitive hashing. Future Generation Computer Systems, 110(??):587–609, September 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17319180.

Gondim:2020:MSA

[GdOAO20] João J. C. Gondim, Robson de Oliveira Albuquerque, and Ana Lucila Sandoval Orozco. Mirror saturation in amplified reflection distributed denial of service: a case of study using SNMP, SSDP, NTP and DNS protocols. Future Generation Computer Systems, 108 (??):68–81, July 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19322745.

Giannakou:2020:MLA

[GDP20] Anna Giannakou, Dipankar Dwivedi, and Sean Peisert. A machine learning approach for packet loss prediction in science flows. Future Generation Computer Systems, 102(??):190–197, January 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19305850.

Gawande:2020:SDL

[GDS^{+}20] Nitin A. Gawande, Jeff A.

Daily, Charles Siegel, Nathan R. Tallent, and Abhinav Vishnu. Scaling deep learning workloads: NVIDIA DGX-1/ Pascal and Intel Knights Landing. Future Generation Computer Systems, 108(??):1162–1172, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17318599.

Gad-Elrab:2020:TTB

[GEN20] Ahmed. A. A. Gad-Elrab and Amin. Y. Noaman. A two-tier bipartite graph task allocation approach based on fuzzy clustering in cloud-fog environment. Future Generation Computer Systems, 103(??):79– 90, February 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19305175.

Gianini:2020:MPR

[GFM^{+}20] Gabriele Gianini, Leopold Ghemmogne Fossi, Corrado Mio, Olivier Caelen, Lionel Brunie, and Ernesto Damiani. Managing a pool of rules for credit card fraud detection by a game theory based approach. Future Generation Computer Systems, 102(??):549–561, January 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18317151.

Gaglianese:2023:AEC

[GFPB23] M. Gaglianese, S. Forti, F. Paganelli, and A. Brogi. Assessing and enhancing a Cloud–IoT monitoring service over federated testbeds. Future Generation Computer Systems, 147(??):77–92, October 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23001668.

Gan:2021:SMC

[GFZ21] Chenquan Gan, Qingdong Feng, and Zufan Zhang. Scalable multi-channel dilated CNN–BiLSTM model with attention mechanism for Chinese textual sentiment analysis. Future Generation Computer Systems, 118(??):297–309, May 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21000340.

Garcia-Garcia:2020:IDJ

[GGCIV20] Francisco García-García, Antonio Corral, Luis Iribarne, and Michael Vassilakopoulos. Improving distance-join query processing with Voronoi-diagram based partitioning in SpatialHadoop. Future Generation Computer Systems, 111(??):723–740, October 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19311872.

Ghomeshi:2020:NCH

[GGK20] Hossein Ghomeshi, Mohamed Medhat Gaber, and Yevgeniya Kovalchuk. A non-canonical hybrid metaheuristic approach to adaptive data stream classification. Future Generation Computer Systems, 102(??):127–139, January 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18332497.

Guerra:2024:TDT

[GHB^{+}24] Victor Guerra, Benoit Hamon, Benoit Bataillou, Adwait Inamdar, and Willem D. van Driel. Towards a digital twin architecture for the lighting industry. Future Generation Computer Systems, 155 (??):80–95, June 2024. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24000359.

Gharbia:2018:MSP

[GHEB^{+}18] Reham Gharbia, Aboul Ella Hassanien, Ali Hassan El-Baz, Mohamed Elhoseny, and M. Gunasekaran. Multi-spectral and panchromatic image fusion approach using stationary wavelet transform and swarm flower pollination optimization for remote sensing applications. Future Generation Computer Systems, 88(??):501–511, November 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18308318.

See retraction notice [GHEB^{+}23].

Gharbia:2023:RNM

[GHEB^{+}23] Reham Gharbia, Aboul Ella Hassanien, Ali Hassan El-Baz, Mohamed Elhoseny, and M. Gunasekaran. Retraction notice to “Multispectral and panchromatic image fusion approach using stationary wavelet transform and swarm flower pollination optimization for remote sensing applications” [Future Gener. Comput. Syst. 88 (2018) 501–511]. Future Generation Computer Systems, 148(??):636, November 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23002339.

See [GHEB^{+}18].

Guo:2021:BAV

[GHG^{+}21] Yi-Ming Guo, Zhen-Ling Huang, Ji Guo, Xing-Rong Guo, Hua Li, Meng-Yu Liu, Safa Ezzeddine, and Mpeoane Judith Nkeli. A bibliometric analysis and visualization of blockchain. Future Generation Computer Systems, 116(??): 316–332, March 2021. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20330004.

Gong:2020:HPS

[GHW^{+}20] Xuri Gong, Zhou Huang, Yaoli Wang, Lun Wu, and Yu Liu. High-performance spatiotemporal trajectory matching across heterogeneous data sources. Future Generation Computer Systems, 105(??):148–161, April 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19319880.

Guzzo:2020:MED

[GIPS20] Antonella Guzzo, Michele Ianni, Andrea Pugliese, and Domenico Saccà. Modeling and efficiently detecting security-critical sequences of actions. Future Generation Computer Systems, 113(??):196–206, December 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19331528.

Guo:2020:LPP

[GJC^{+}20] Cheng Guo, Xueru Jiang, Kim-Kwang Raymond Choo, Xinyu Tang, and Jing Zhang. Lightweight privacy preserving data aggregation with batch verification for smart grid. Future Generation Computer Systems, 112(??):512–523, November 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1933225X.

Gulati:2021:AED

[GK21] Nancy Gulati and Pankaj Deep Kaur. An argumentation enabled decision making approach for fall activity recognition in social IoT based ambient assisted living systems. Future Generation Computer Systems, 122(??):82– 97, September 2021. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21001205.

Goyat:2021:SLS

[GKA^{+}21] Rekha Goyat, Gulshan Kumar, Mamoun Alazab, Rahul Saha, Reji Thomas, and Mritunjay Kumar Rai. A secure localization scheme based on trust assessment for WSNs using blockchain technology. Future Generation Computer Systems, 125(??):221–231, December 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2100234X.

Garg:2020:MSA

[GKB^{+}20] Sahil Garg, Kuljeet Kaur, Shalini Batra, Georges Kaddoum, Neeraj Kumar, and Azzedine Boukerche. A multi-stage anomaly detection scheme for augmenting the security in IoTenabled applications. Future Generation Computer Systems, 104(??):105–118, March 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19319703.

Ghasempouri:2020:MCR

[GL20] Seyed Asgary Ghasempouri and Behrouz Tork Ladani. Model checking of robustness properties in trust and reputation systems. Future Generation Computer Systems, 108(??):302–319, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19317741.

Guo:2022:GAP

[GLF^{+}22] Zehua Guo, Liangjie Liu, Mofan Feng, Kai Su, Runqiu Chi, Keyi Li, Qing Lu, Xianbin Su, Lintai Da, Song Cao, Mingxuan Zhang, Luming Meng, Dan Cao, Jiayi Wang, Guang He, and Yi Shi. 3D genome assisted protein–protein interaction prediction. Future Generation Computer Systems, 137(??):87–96, December 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X2200231X.

Garavagno:2024:COL

[GLF24] Andrea Mattia Garavagno, Daniele Leonardis, and Antonio Frisoli. Colab-NAS: Obtaining lightweight task-specific convolutional neural networks following Occam’s razor. Future Generation Computer Systems, 152(??):152–159, March 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23004028.

Guidec:2021:CCB

[GLM21] Frédéric Guidec, Pascale Launay, and Yves Mahéo. Causal and Δcausal broadcast in opportunistic networks. Future Generation Computer Systems, 118(??):142–156, May 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20330867.

Gong:2024:LTA

[GLP^{+}24] Kun Gong, Senlin Luo, Limin Pan, Linghao Zhang, Yifei Zhang, and Haomiao Yu. LogETA: Timeaware cross-system logbased anomaly detection with inter-class boundary optimization. Future Generation Computer Systems, 157(??):16–28, August 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X24000992.

Guo:2020:ZBC

[GlRpG20] Hui Guo, Lan lan Rui, and Zhi peng Gao. A zone-based content precaching strategy in vehicular edge networks. Future Generation Computer Systems, 106(??):22–33, May 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1932165X.

Guan:2020:TSE

[GLW^{+}20] Zhitao Guan, Xin Lu, Naiyu Wang, Jun Wu, Xiaojiang Du, and Mohsen Guizani. Towards secure and efficient energy trading in IIoT-enabled energy Internet: a blockchain approach. Future Generation Computer Systems, 110(??):686–695, September 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19315018.

Guo:2020:NDC

[GLWP20] Xuancheng Guo, Hui Lin, Yulei Wu, and Min Peng. A new data clustering strategy for enhancing mutual privacy in healthcare IoT systems. Future Generation Computer Systems, 113(??):407–417, December 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19335319.

Gul:2022:CMF

[GMA^{+}22] Faiza Gul, Imran Mir, Laith Abualigah, Suleman Mir, and Maryam Altalhi. Cooperative multi-function approach: a new strategy for autonomous ground robotics. Future Generation Computer Systems, 134(??):361–373, September 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22001303.

Gowda:2023:BFB

[GMAL23] Naveen Chandra Gowda, Sunilkumar S. Manvi, Bharathi Malakreddy A., and Pascal Lorenz. BSKM-FC: Blockchain-based secured key management in a fog computing environment. Future Generation Computer Systems, 142 (??):276–291, May 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0167739X22004460.

Guerra-Manzanares:2023:AAL

[GMB23] Alejandro Guerra-Manzanares and Hayretdin Bahsi. On the application of active learning for efficient and effective IoT botnet detection. Future Generation Computer Systems, 141 (??):40–53, April 2023. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22003399.

Gonzalez-Manzano:2023:CDW

[GMBdF^{+}23] Lorena González-Manzano, Marta Beltrán, José María de Fuentes, Gianluca Dini, and Cristina Alcaráz. Cybersecurity in the digital world. Future Generation Computer Systems, 144(??):242–243, July 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23000729.

Guedes:2020:CAP

[GMF^{+}20] Thaylon Guedes, Lucas Bertelli Martins, Maria Luiza Furtuozo Falci, Vitor Silva, Kary A. C. S. Ocaña, Marta Mattoso, Marcos Bedo, and Daniel de Oliveira. Capturing and analyzing provenance from Spark-based scientific workflows with SAMbA-RaP. Future Generation Computer Systems, 112(??):658–669, November 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1931742X.

Gomez:2023:HLV

[GMFC23] Constantino Gómez, Filippo Mantovani, Erich Focht, and Marc Casas. HPCG on long-vector architectures: Evaluation and optimization on NEC SX-Aurora and RISC-V. Future Generation Computer Systems, 143(??):152–162, June 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23000225.

Gonzalo:2022:CNC

[GMGV^{+}22] Sergio Gonzalo, Joan Manuel Marquès, Alberto García-Villoria, Javier Panadero, and Laura Calvet. CLARA: a novel clustering-based resource-allocation mechanism for exploiting lowavailability complementarities of voluntarily contributed nodes. Future Generation Computer Systems, 128(??):248–264, March 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21003927.

Ghoneim:2020:CCC

[GMH20] Ahmed Ghoneim, Ghulam Muhammad, and M. Shamim Hossain. Cervical cancer classification using convolutional neural networks and extreme learning machines. Future Generation Computer Systems, 102(??):643–649, January 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19306065.

Garompolo:2022:BSC

[GMI22] D. Garompolo, A. Molinaro, and A. Iera. Bridging separate communities with common interest in distributed social networks through the use of social objects. Future Generation Computer Systems, 129(??):440–452, April 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21004246.

Gari:2022:QLA

[GMM22] Yisel Garí, David A.

Monge, and Cristian Mateos. A Q-learning approach for the autoscaling of scientific workflows in the cloud. Future Generation Computer Systems, 127(??):168–180, February 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21003538.

Gennari:2024:HEY

[GMMR24] Rosella Gennari, Maristella Matera, Alessandra Melonio, and Mehdi Rizvi. How to enable young teens to design responsibly. Future Generation Computer Systems, 150(??):303–316, January 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23003321.

Gogouvitis:2020:SCI

[GMP^{+}20a] Spyridon V. Gogouvitis, Harald Mueller, Sreenath Premnadh, Andreas Seitz, and Bernd Bruegge. Seamless computing in industrial systems using container orchestration. Future Generation Computer Systems, 109(??):678–688, August 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17330236.

Gourisetti:2020:CVM

[GMP20b] Sri Nikhil Gupta Gourisetti, Michael Mylrea, and Hirak Patangia. Cybersecurity vulnerability mitigation framework through empirical paradigm: Enhanced prioritized gap analysis. Future Generation Computer Systems, 105(??):410–431, April 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19307344.

Gonthier:2023:TDL

[GMT23] Maxime Gonthier, Loris Marchal, and Samuel Thibault. Taming data locality for task scheduling under memory constraint in runtime systems. Future Generation Computer Systems, 143 (??):305–321, June 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0167739X23000328.

Gheisari:2021:OOB

[GNA^{+}21] Mehdi Gheisari, Hamid Esmaeili Najafabadi, Jafar A. Alzubi, Jiechao Gao, Guojun Wang, Aaqif Afzaal Abbasi, and Aniello Castiglione. OBPP: an ontology-based framework for privacy-preserving in IoT-based smart city. Future Generation Computer Systems, 123(??):1–13, October 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21000388.

Gulcan:2023:LBL

[GOA23] Selcuk Gulcan, Muhammet Mustafa Ozdal, and Cevdet Aykanat. Load balanced locality-aware parallel SGD on multicore architectures for latent factor based collaborative filtering. Future Generation Computer Systems, 146(??):207–221, September 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23001413.

Gupta:2021:IIE

[GPC21] Shubham Gupta, Balu L.

Parne, and Narendra S. Chaudhari. ISAG: IoTenabled and Secrecy Aware Group-based handover scheme for e-health services in M2M communication network. Future Generation Computer Systems, 125(??):168–187, December 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21002338.

Gorawski:2023:SDW

[GPGG23] M. Gorawski, K. Pasterak, A. Gorawska, and M. Gorawski. The stream data warehouse: Page replacement algorithms and quality of service metrics. Future Generation Computer Systems, 142(??):212–227, May 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23000031.

Gudkov:2023:BRB

[GPR23] Andrei Gudkov, Pavel Popov, and Stepan Romanov. BalCon — resource balancing algorithm for VM consolidation. Future Generation Computer Systems, 147(??):265–274, October 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23001760.

Gari:2024:ORB

[GPR^{+}24] Yisel Garí, Elina Pacini, Luciano Robino, Cristian Mateos, and David A. Monge. Online RL-based cloud autoscaling for scientific workflows: Evaluation of Q-learning and SARSA. Future Generation Computer Systems, 157(??): 573–586, August 2024. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24001432.

Garre:2021:NML

[GPRM21] José Tomás Martínez Garre, Manuel Gil Pérez, and Antonio Ruiz-Martínez. A novel machine learningbased approach for the detection of SSH botnet infection. Future Generation Computer Systems, 115(??):387–396, February 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20303265.

Gui:2020:MMS

[GPWL20] Zhipeng Gui, Dehua Peng, Huayi Wu, and Xi Long. MSGC: Multi-scale grid clustering by fusing analytical granularity and visual cognition for detecting hierarchical spatial patterns. Future Generation Computer Systems, 112(??):1038–1056, November 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20304830.

Gramoli:2020:BCB

[Gra20] Vincent Gramoli. From blockchain consensus back to Byzantine consensus. Future Generation Computer Systems, 107(??):760–769, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17320095.

Galland:2020:RTE

[GRG20] Stéphane Galland, Sebastian Rodriguez, and Nicolas Gaud. Run-time environment for the SARL agent-programming language: the example of the Janus platform. Future Generation Computer Systems, 107(??):1105–1115, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17313419.

Gazori:2020:STC

[GRN20] Pegah Gazori, Dadmehr Rahbari, and Mohsen Nickray. Saving time and cost on the scheduling of fog-based IoT applications using deep reinforcement learning approach. Future Generation Computer Systems, 110(??):1098–1115, September 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19308702.

G:2020:ADF

[GS20] Bhuvaneswari Amma N.

G. and Selvakumar S. Anomaly detection framework for Internet of things traffic using vector convolutional deep learning approach in fog environment. Future Generation Computer Systems, 113(??):255–265, December 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19316954.

Gaamel:2020:BLM

[GSARS20] Awadh Gaamel, Tarek Sheltami, Anas Al-Roubaiey, and Elhadi Shakshuki. Broker-less middleware for WSAN performance evaluation. Future Generation Computer Systems, 110(??):372–381, September 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1732900X.

Godinho:2022:RRM

[GSCP22] Noé Godinho, Henrique Silva, Marilia Curado, and Luís Paquete. A reconfigurable resource management framework for fog environments. Future Generation Computer Systems, 133(??):124–140, August 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22000905.

Garcia-Silva:2021:IKB

[GSDGP21] Andres Garcia-Silva, Ronald Denaux, and Jose Manuel Gomez-Perez. On the impact of knowledge-based linguistic annotations in the quality of scientific embeddings. Future Generation Computer Systems, 120 (??):26–35, July 2021. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21000716.

Guo:2023:CCS

[GSG^{+}23] Xiangyu Guo, Kai Song, Mingliang Gao, Wenzhe Zhai, Qilei Li, and Gwanggil Jeon. Crowd counting in smart city via lightweight Ghost Attention Pyramid Network. Future Generation Computer Systems, 147(??):328–338, October 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23001930.

Gundersen:2022:DML

[GSI22] Odd Erik Gundersen, Saeid Shamsaliei, and Richard Juul Isdahl. Do machine learning platforms provide outof-the-box reproducibility? Future Generation Computer Systems, 126(??):34– 47, January 2022. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21002090.

Gireesha:2020:IWI

[GSKS20] Obulaporam Gireesha, Nivethitha Somu, Kannan Krithivasan, and Shankar Sriram V. S. IIVIFS-WASPAS: an integrated multi-criteria decision-making perspective for cloud service provider selection. Future Generation Computer Systems, 103(??):91–110, February 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19307307.

Gallo:2020:PUR

[GSMF20] Fabio R. Gallo, Gerardo I.

Simari, Maria Vanina Martinez, and Marcelo A. Falappa. Predicting user reactions to Twitter feed content based on personality type and social cues. Future Generation Computer Systems, 110(??):918–930, September 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19304091.

Gonzalez-Sendino:2024:MBA

[GSSB24] Rubén González-Sendino, Emilio Serrano, and Javier Bajo. Mitigating bias in artificial intelligence: Fair data generation via causal models for transparent and explainable decision-making. Future Generation Computer Systems, 155(??):384–401, June 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24000694.

Gruhl:2021:NDC

[GST21] Christian Gruhl, Bernhard Sick, and Sven Tomforde. Novelty detection in continuously changing environments. Future Generation Computer Systems, 114(??):138–154, January 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2030323X.

Guo:2021:DGN

[GTG^{+}21] Zhiwei Guo, Lianggui Tang, Tan Guo, Keping Yu, Mamoun Alazab, and Andrii Shalaginov. Deep graph neural network-based spammer detection under the perspective of heterogeneous cyberspace. Future Generation Computer Systems, 117(??):205–218, April 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20330612.

Gulec:2022:ELW

[Gul22] Omer Gulec. Extending lifetime of wireless nanosensor networks: an energy efficient distributed routing algorithm for Internet of Nano-Things. Future Generation Computer Systems, 135(??):382–393, October 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22001741.

Guri:2021:EDA

[Gur21a] Mordechai Guri. Exfiltrating data from air-gapped computers via ViBrAtIoNs. Future Generation Computer Systems, 122(??):69– 81, September 2021. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21001151.

Guri:2021:MCC

[Gur21b] Mordechai Guri. MAG-NETO: Covert channel between air-gapped systems and nearby smartphones via CPU-generated magnetic fields. Future Generation Computer Systems, 115(??):115–125, February 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X2030916X.

Garcia-Valls:2024:CCI

[GVCC24] Marisol García-Valls and Alejandro M. Chirivella-Ciruelos. CoTwin: Collaborative improvement of digital twins enabled by blockchain. Future Generation Computer Systems, 157(??):408–421, August 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X24001171.

Garcia-Valls:2020:ASE

[GVCUGF20] Marisol García-Valls, Christian Calva-Urrego, and Ana García-Fornes. Accelerating smart eHealth services execution at the fog computing infrastructure. Future Generation Computer Systems, 108(??):882–893, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17327425.

Gjurkovic:2022:SSI

[GVŠ22] Matej Gjurković, Iva Vukojević, and Jan Šnajder. SIMPA: Statement-to-Item Matching Personality Assessment from text. Future Generation Computer Systems, 130(??):114–127, May 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21004957.

Gu:2020:INI

[GW20] Zi-Min Gu and Gai-Ge Wang. Improving NSGA-III algorithms with information feedback models for large-scale manyobjective optimization. Future Generation Computer Systems, 107(??):49–69, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19320552.

Gao:2022:SPP

[GW22] Shan Gao and Hanyi Wang.

Scenario prediction of public health emergencies using infectious disease dynamics model and dynamic Bayes. Future Generation Computer Systems, 127(??):334–346, February 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21003745.

Wu:2021:EFS

[gWLWZ21] Chu ge Wu, Wei Li, Ling Wang, and Albert Y. Zomaya. An evolutionary fuzzy scheduler for multiobjective resource allocation in fog computing. Future Generation Computer Systems, 117(??):498–509, April 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20330818.

Gong:2022:SLC

[GWW^{+}22] Bei Gong, Yong Wu, Qian Wang, Yu heng Ren, and Chong Guo. A secure and lightweight certificateless hybrid signcryption scheme for Internet of Things. Future Generation Computer Systems, 127(??):23– 30, February 2022. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21003356.

Gao:2020:DNE

[GWY^{+}20] Ying Gao, Xiaoqiang Wu, Wei Yan, Lei Zhang, and Tunhua Wu. Dynamic network embedding enhanced advisor-advisee relationship identification based on Internet of Scholars. Future Generation Computer Systems, 108(??):677–686, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1933537X.

Gan:2020:MES

[GWZ20] Chenquan Gan, Lu Wang, and Zufan Zhang. Multientity sentiment analysis using self-attention based hierarchical dilated convolutional neural network. Future Generation Computer Systems, 112(??):116–125, November 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19329383.

Gu:2022:MMS

[GXS22] Qi Gu, Zhihua Xia, and Xingming Sun. MSP-PIR: Multi-Source Privacy-Preserving Image Retrieval in cloud computing. Future Generation Computer Systems, 134(??):78–92, September 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22001212.

Guo:2022:CPD

[GYAW22] Tan Guo, Keping Yu, Moayad Aloqaily, and Shaohua Wan. Constructing a prior-dependent graph for data clustering and dimension reduction in the edge of AIoT. Future Generation Computer Systems, 128(??):381–394, March 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21003903.

Guo:2022:FOC

[GZ22] Kai Guo and Ruiling Zhang. Fairness-oriented computation offloading for cloud-assisted edge computing. Future Generation Computer Systems, 128(??):132–141, March 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21003940.

Gomez:2022:ECC

[GZB^{+}22] Elizabeth Gómez, Carlos Shui Zhang, Ludovico Boratto, Maria Salamó, and Guilherme Ramos. Enabling cross-continent provider fairness in educational recommender systems. Future Generation Computer Systems, 127(??):435–447, February 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21003332.

Gao:2020:FMM

[GZF^{+}20a] Yicong Gao, Zixian Zhang, Yixiong Feng, Maria Savchenko, Ichiro Hagiwara, and Hao Zheng. Flexible mesh morphing in sustainable design using data mining and mesh subdivision. Future Generation Computer Systems, 108(??):987–994, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17317818.

Guo:2020:ERP

[GZF^{+}20b] Zehua Guo, Shaojun Zhang, Wendi Feng, Weichao Wu, and Julong Lan. Exploring the role of paths for dynamic switch assignment in software-defined networks. Future Generation Computer Systems, 107(??): 238–246, June 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18324543.

Gagliardelli:2023:BDP

[GZF^{+}23] Luca Gagliardelli, Luca Zecchini, Luca Ferretti, Domenico Beneventano, Giovanni Simonini, Sonia Bergamaschi, Mirko Orsini, Luca Magnotta, Emma Mescoli, Andrea Livaldi, Nicola Gessa, Piero De Sabbata, Gianluca D’Agosta, Fabrizio Paolucci, and Fabio Moretti. A big data platform exploiting auditable tokenization to promote good practices inside local energy communities. Future Generation Computer Systems, 141(??): 595–610, April 2023. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22004113.

Guo:2020:AAS

[GZG20] Nan Guo, Cong Zhao, and Tianhan Gao. An anonymous authentication scheme for edge computingbased car–home connectivity services in vehicular networks. Future Generation Computer Systems, 106(??):659–671, May 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167739X19323179.

Gao:2022:DCD

[GZL^{+}22] Lina Gao, Zhongying Zhao, Chao Li, Jianli Zhao, and Qingtian Zeng. Deep cognitive diagnosis model for predicting students’ performance. Future Generation Computer Systems, 126(??):252–262, January 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21003277.

Gaaloul:2020:SIF

[GZPZ20] Walid Gaaloul, Zhangbing Zhou, Hervé Panetto, and Liang Zhang. Special issue on fog and cloud computing for cooperative information system management: Challenges and opportunities. Future Generation Computer Systems, 109(??): 704–705, August 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20306609.

Gan:2021:NFC

[GZT^{+}21] Hongxiao Gan, Jinglan Zhang, Michael Towsey, Anthony Truskinger, Debra Stark, Berndt J. van Rensburg, Yuefeng Li, and Paul Roe. A novel frog chorusing recognition method with acoustic indices and machine learning. Future Generation Computer Systems, 125(??):485–495, December 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21002144.

Gong:2024:QPS

[GZXH24] Chen Gong, Nanrun Zhou, Shuhua Xia, and Shuiyuan Huang. Quantum particle swarm optimization algorithm based on diversity migration strategy. Future Generation Computer Systems, 157(??):445–458, August 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24001389.

Guo:2024:EEE

[GZZG24] Hongtai Guo, Zhangbing Zhou, Deng Zhao, and Walid Gaaloul. EGNN: Energy-efficient anomaly detection for iot multivariate time series data using graph neural network. Future Generation Computer Systems, 151(??):45– 56, February 2024. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23003606.

Haider:2020:PDP

[HAA^{+}20] Shahab Haider, Ghulam Abbas, Ziaul Haq Abbas, Saadi Boudjit, and Zahid Halim. P-DACCA: a probabilistic direction-aware cooperative collision avoidance scheme for VANETs. Future Generation Computer Systems, 103(??):1– 17, February 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1831714X.

Hasan:2020:SDA

[HAB^{+}20] Khalid Hasan, Khandakar Ahmed, Kamanashis Biswas, Md. Saiful Islam, and Omid Ameri Sianaki. Softwaredefined application-specific traffic management for wireless body area networks. Future Generation Computer Systems, 107 (??):274–285, June 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167739X19322587.

Husnoo:2023:FDI

[HAH^{+}23] Muhammad Akbar Husnoo, Adnan Anwar, Nasser Hosseinzadeh, Shama Naz Islam, Abdun Naser Mahmood, and Robin Doss. False data injection threats in active distribution systems: a comprehensive survey. Future Generation Computer Systems, 140(??):344–364, March 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22003363.

Hakak:2021:EML

[HAK^{+}21] Saqib Hakak, Mamoun Alazab, Suleman Khan, Thippa Reddy Gadekallu, Praveen Kumar Reddy Maddikunta, and Wazir Zada Khan. An ensemble machine learning approach through effective feature extraction to classify fake news. Future Generation Computer Systems, 117 (??):47–58, April 2021. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20330466.

Helmi:2023:HAR

[HAqDE23] Ahmed M. Helmi, Mohammed A. A. Al-qaness, Abdelghani Dahou, and Mohamed Abd Elaziz. Human activity recognition using marine predators algorithm with deep learning. Future Generation Computer Systems, 142(??): 340–350, May 2023. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23000134.

Ahmad:2024:AFS

[hAS24] Hussien Al haj Ahmad and Yasser Sedaghat. An automated framework for selectively tolerating SDC errors based on rigorous instruction-level vulnerability assessment. Future Generation Computer Systems, 157(??):392–407, August 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24001365.

Harth:2022:LFL

[HAVK22] Natascha Harth, Christos Anagnostopoulos, Hans-Joerg Voegel, and Kostas Kolomvatsos. Local & federated learning at the network edge for efficient predictive analytics. Future Generation Computer Systems, 134(??):107–122, September 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22001054.

Hussain:2021:CVM

[HB21] Md. Muzakkir Hussain and M. M. Sufyan Beg. CODE-V: Multi-hop computation offloading in vehicular fog computing. Future Generation Computer Systems, 116(??):86–102, March 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20303526.

Hadjres:2020:GET

[HBEK20] Souad Hadjres, Fatna Belqasmi, May El Barachi, and Nadjia Kara. A green, energy, and trust-aware multi-objective cloud coalition formation approach. Future Generation Computer Systems, 111(??):52– 67, October 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20300844.

Henaien:2024:SSI

[HBF24] Amira Henaien, Hadda Ben Elhadj, and Lamia Chaari Fourati. A sustainable smart IoT-based solid waste management system. Future Generation Computer Systems, 157(??):587–602, August 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24001183.

Happ:2021:JJP

[HBH21] Daniel Happ, Suzan Bayhan, and Vlado Handziski. JOI: Joint placement of IoT analytics operators and pub/sub message brokers in fog-centric IoT platforms. Future Generation Computer Systems, 119(??):7– 19, June 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21000364.

Hanussek:2020:BBB

[HBK20] Maximilian Hanussek, Felix Bartusch, and Jens Krüger. BOOTABLE: Bioinformatics benchmark tool suite for applications and hardware. Future Generation Computer Systems, 102(??):1016–1026, January 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1931533X.

Husak:2021:PMC

[HBSG21] Martin Husák, Václav Bartoš, Pavol Sokol, and Andrej Gajdoš. Predictive methods in cyber defense: Current experience and research challenges. Future Generation Computer Systems, 115(??):517–530, February 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20329836.

HewaNadungodage:2020:DPS

[HCB^{+}20] Chandima HewaNadungodage, Ann Christine Catlin, Andres Bejarano, Steven Clark, Guneshi Wickramaarachchi, Sumudinie Fernando, and Parameswaran Desigavinayagam. The DEEDS platform: Support for integrated data and computing across the research lifecycle. Future Generation Computer Systems, 111(??):793–805, October 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19302742.

Hou:2023:GND

[HCG^{+}23] Jiaxin Hou, Meng Chen, Haijun Geng, Rongzhen Li, and Jianyuan Lu. GP-NFSP: Decentralized task offloading for mobile edge computing with independent reinforcement learning. Future Generation Computer Systems, 141 (??):205–217, April 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167739X22003867.

Hussain:2020:NCC

[HCK20a] Abid Hussain, Jin Chun, and Maria Khan. A novel customer-centric Methodology for Optimal Service Selection (MOSS) in a cloud environment. Future Generation Computer Systems, 105(??):562–580, April 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19300950.

Hussain:2020:NFT

[HCK20b] Abid Hussain, Jin Chun, and Maria Khan. A novel framework towards viable Cloud Service Selection as a Service (CSSaaS) under a fuzzy environment. Future Generation Computer Systems, 104(??):74– 91, March 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19302912.

Hu:2022:BBG

[HCL^{+}22] Xinxin Hu, Hongchang Chen, Shuxin Liu, Haocong Jiang, Guanghan Chu, and Ran Li. BTG: a bridge to graph machine learning in telecommunications fraud detection. Future Generation Computer Systems, 137(??):274–287, December 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22002540.

Hao:2021:EAS

[HCWD21] Yongsheng Hao, Jie Cao, Qi Wang, and Jinglin Du. Energy-aware scheduling in edge computing with a clustering method. Future Generation Computer Systems, 117(??):259–272, April 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20330624.

Hu:2020:SAK

[HDD20] Xin Hu, Jiangli Duan, and Depeng Dang. Scalable aggregate keyword query over knowledge graph. Future Generation Computer Systems, 107(??):588–600, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19300251.

Havers:2020:DFE

[HDN^{+}20] Bastian Havers, Romaric Duvignau, Hannaneh Najdataei, Vincenzo Gulisano, Marina Papatriantafilou, and Ashok Chaitanya Koppisetty. DRIVEN: a framework for efficient data retrieval and clustering in vehicular networks. Future Generation Computer Systems, 107(??):1–17, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19310507.

Heidsieck:2021:CAS

[HdOP^{+}21] Gaëtan Heidsieck, Daniel de Oliveira, Esther Pacitti, Christophe Pradal, François Tardieu, and Patrick Valduriez. Cache-aware scheduling of scientific workflows in a multisite cloud. Future Generation Computer Systems, 122(??):172–186, September 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21000923.

Hua:2020:FCS

[HGK20] Yining Hua, Lin Guan, and Konstantinos G. Kyriakopoulos. A fog caching scheme enabled by ICN for IoT environments. Future Generation Computer Systems, 111(??):82–95, October 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19326457.

Huang:2023:AAE

[HGWC23] Jiwei Huang, Han Gao, Shaohua Wan, and Ying Chen. AoI-aware energy control and computation offloading for industrial IoT. Future Generation Computer Systems, 139(??):29– 37, February 2023. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22002916.

Huang:2022:HMB

[HGY^{+}22] Xiaowen Huang, Shimin Gong, Jingmin Yang, Wenjie Zhang, Liwei Yang, and Chai Kiat Yeo. Hybrid market-based resources allocation in mobile edge computing systems under stochastic information. Future Generation Computer Systems, 127(??):80– 91, February 2022. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X2100337X.

Hasan:2022:PSU

[HH22] Raiful Hasan and Ragib Hasan. Pedestrian safety using the Internet of Things and sensors: Issues, challenges, and open problems. Future Generation Computer Systems, 134(??):187–203, September 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22001170.

Hamdi:2022:TOV

[HHH22] Aisha Muhammad A. Hamdi, Farookh Khadeer Hussain, and Omar K. Hussain. Task offloading in vehicular fog computing: State-of-the-art and open issues. Future Generation Computer Systems, 133(??):201–212, August 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22000942.

Huang:2020:PPA

[HHLZ20] Yuze Huang, Jiwei Huang, Cong Liu, and Chengning Zhang. PFPMine: a parallel approach for discovering interacting data entities in data-intensive cloud workflows. Future Generation Computer Systems, 113(??):474–487, December 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19334909.

Hong:2022:WBP

[HHW^{+}22] Zhen Hong, Miao Hong, Ning Wang, Yong Ma, Xiaolong Zhou, and Wei Wang. A wearable-based posture recognition system with AI-assisted approach for healthcare IoT. Future Generation Computer Systems, 127(??):286–296, February 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21003381.

Han:2020:CVF

[HIdAR^{+}20] Tao Han, Roberto F. Ivo, Douglas de A. Rodrigues, Solon A. Peixoto, Victor Hugo C. de Albuquerque, and Pedro P. Rebouças Filho. Cascaded volumetric fully convolutional networks for wholeheart and great vessel 3D segmentation. Future Generation Computer Systems, 108(??):198–209, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19331437.

Hameed:2020:LCE

[HIMM20] Mustafa Emad Hameed, Masrullizam Mat Ibrahim, Nurulfajar Abd Manap, and Ali A. Mohammed. A lossless compression and encryption mechanism for remote monitoring of ECG data using Huffman coding and CBC–AES. Future Generation Computer Systems, 111(??):829–840, October 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19313950.

Hassan:2022:PCC

[HIU^{+}22] Md. Rafiul Hassan, Md. Fakrul Islam, Md. Zia Uddin, Goutam Ghoshal, Mohammad Mehedi Hassan, Shamsul Huda, and Giancarlo Fortino. Prostate cancer classification from ultrasound and MRI images using deep learning based explainable artificial intelligence. Future Generation Computer Systems, 127(??):462–472, February 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21003769.

Hou:2024:EET

[HJI24] Huanhuan Hou, Siti Nuraishah Agos Jawaddi, and Azlan Ismail. Energy efficient task scheduling based on deep reinforcement learning in cloud environment: a specialized review. Future Generation Computer Systems, 151(??):214–231, February 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23003771.

Hu:2020:FMR

[HJW^{+}20] Long Hu, Yinging Jiang, Fangxin Wang, Kai Hwang, M. Shamim Hossain, and Ghulam Muhammad. Follow me Robot–Mind: Cloud brain based personalized robot service with migration. Future Generation Computer Systems, 107 (??):324–332, June 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0167739X1932953X.

Han:2023:EHI

[HJX^{+}23] Zhongming Han, Xuelian Jin, Haozhen Xing, Weijie Yang, and Haitao Xiong. An effective heterogeneous information network representation learning framework. Future Generation Computer Systems, 148(??):66–78, November 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23002091.

Hudson:2024:QAE

[HKB^{+}24] Nathaniel Hudson, Hana Khamfroush, Matt Baughman, Daniel E. Lucani, Kyle Chard, and Ian Foster. QoS-aware edge AI placement and scheduling with multiple implementations in FaaS-based edge computing. Future Generation Computer Systems, 157(??):250–263, August 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24001067.

Huber:2023:BBS

[HKS23] Brennan Huber, Farah Kandah, and Anthony Skjellum. BEAST: Behavior as a Service for Trust management in IoT devices. Future Generation Computer Systems, 144(??): 165–178, July 2023. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23000444.

Huang:2024:IFL

[HL24] Guan-Ying Huang and Ching-Hung Lee. Industrial federated learning algorithm (P-PFedSGD) for tool wear estimation. Future Generation Computer Systems, 158(??):150–157, September 2024. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X24001547.

Hu:2020:TEC

[HLH^{+}20] Yang Hu, John C. S. Lui, Wenjun Hu, Xiaobo Ma, Jianfeng Li, and Xiao Liang. Taming energy cost of disk encryption software on dataintensive mobile devices. Future Generation Computer Systems, 107(??): 681–691, June 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17320113.

Huai:2023:LCD

[HLK^{+}23] Shuo Huai, Di Liu, Hao Kong, Weichen Liu, Ravi Subramaniam, Christian Makaya, and Qian Lin. Latency-constrained DNN architecture learning for edge systems using zerorized batch normalization. Future Generation Computer Systems, 142(??): 314–327, May 2023. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22004253.

Hu:2020:DSA

[HLL^{+}20] Yan Hu, Hong Li, Tom H.

Luan, An Yang, Limin Sun, Zhiliang Wang, and Rui Wang. Detecting stealthy attacks on industrial control systems using a permutation entropybased method. Future Generation Computer Systems, 108(??):1230–1240, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17329412.

Hu:2024:CCE

[HLL^{+}24] Shijing Hu, Junxiong Lin, Zhihui Lu, Xin Du, Qiang Duan, and Shih-Chia Huang. CoLLaRS: a cloud–edge– terminal collaborative lifelong learning framework for AIoT. Future Generation Computer Systems, 158(??):447–456, September 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24001870.

Hao:2021:CAI

[HLP21] Jingwei Hao, Senlin Luo, and Limin Pan. Computeraided intelligent design using deep multi-objective cooperative optimization algorithm. Future Generation Computer Systems, 124(??):49–53, November 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21001643.

He:2021:SPO

[HLT^{+}21] Chaobo He, Hai Liu, Yong Tang, Shuangyin Liu, Xiang Fei, Qiwei Cheng, and Hanchao Li. Similarity preserving overlapping community detection in signed networks. Future Generation Computer Systems, 116(??):275–290, March 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20330181.

Han:2023:FSF

[HLW^{+}23a] Biao Han, Yahui Li, Xiaoyan Wang, Hanxun Li, and Jinsen Huang. FLoRa: Sequential fuzzy extractor based physical layer key generation for LPWAN. Future Generation Computer Systems, 140(??):253–265, March 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22003284.

He:2023:DSB

[HLW^{+}23b] Huijie He, Yingxu Lai, Yipeng Wang, Siqi Le, and Zijian Zhao. A data skew-based unknown traffic classification approach for TLS applications. Future Generation Computer Systems, 138(??):1–12, January 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22002655.

Hei:2022:PAC

[HLZ^{+}22] Yiming Hei, Dawei Li, Chi Zhang, Jianwei Liu, Yizhong Liu, and Qianhong Wu. Practical AgentChain: a compatible cross-chain exchange system. Future Generation Computer Systems, 130(??):207–218, May 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2100474X.

Hajizadeh:2021:CDP

[HMA^{+}21] Fahimeh Hajizadeh, Mohammadreza Binesh Marvasti, Seyyed Amir Asghari, Mostafa Abbas Mollaei, and Amir M. Rahmani. Configurable DSI partitioned approximate multiplier. Future Generation Computer Systems, 115(??):100–114, February 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20301345.

Haseeb:2022:ABF

[HMH^{+}22] Junaid Haseeb, Masood Mansoori, Yuichi Hirose, Harith Al-Sahaf, and Ian Welch. Autoencoder-based feature construction for IoT attacks clustering. Future Generation Computer Systems, 127(??):487–502, February 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X2100371X.

Huang:2020:MMB

[HMLS20] Liwei Huang, Yutao Ma, Yanbo Liu, and Arun Kumar Sangaiah. Multimodal Bayesian embedding for point-of-interest recommendation on locationbased cyber-physical-social networks. Future Generation Computer Systems, 108(??):1119–1128, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17310191.

Hammoud:2020:CFF

[HMO^{+}20] Ahmad Hammoud, Azzam Mourad, Hadi Otrok, Omar Abdel Wahab, and Haidar Harmanani. Cloud federation formation using genetic and evolutionary game theoretical models. Future Generation Computer Systems, 104(??):92– 104, March 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19306995.

Haghi:2023:WFE

[HMSA^{+}23] Abbas Haghi, Santiago Marco-Sola, Lluc Alvarez, Dionysios Diamantopoulos, Christoph Hagleitner, and Miquel Moreto. WFA-FPGA: an efficient accelerator of the wavefront algorithm for short and long read genomics alignment. Future Generation Computer Systems, 149(??):39– 58, December 2023. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2300256X.

Heydari:2020:KUI

[HMT^{+}20] Mohammad Heydari, Alexios Mylonas, Vahid Heydari Fami Tafreshi, Elhadj Benkhelifa, and Surjit Singh. Known unknowns: Indeterminacy in authentication in IoT. Future Generation Computer Systems, 111(??):278–287, October 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1931982X.

Hasan:2023:UNO

[HMY^{+}23] Haya R. Hasan, Mohammad Madine, Ibrar Yaqoob, Khaled Salah, Raja Jayaraman, and Dragan Boscovic. Using NFTs for ownership management of digital twins and for proof of delivery of their physical assets. Future Generation Computer Systems, 146(??):1– 17, September 2023. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23001280.

Halder:2022:EST

[HN22] Subir Halder and Thomas Newe. Enabling secure time-series data sharing via homomorphic encryption in cloud-assisted IIoT. Future Generation Computer Systems, 133(??):351–363, August 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22001078.

Halder:2023:RFA

[HN23] Subir Halder and Thomas Newe. Radio fingerprinting for anomaly detection using federated learning in LoRa-enabled Industrial Internet of Things. Future Generation Computer Systems, 143(??):322–336, June 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23000298.

Huynh:2020:EAM

[HNV^{+}20] Huy M. Huynh, Loan T. T.

Nguyen, Bay Vo, Unil Yun, Zuzana Komínková Oplatková, and Tzung-Pei Hong. Efficient algorithms for mining clickstream patterns using pseudo-IDLists. Future Generation Computer Systems, 107(??):18– 30, June 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19314475.

Hammoud:2021:SFF

[HOMD21] Ahmad Hammoud, Hadi Otrok, Azzam Mourad, and Zbigniew Dziong. Stable federated fog formation: an evolutionary game theoretical approach. Future Generation Computer Systems, 124(??):21–32, November 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21001710.

Huaman:2020:AIS

[HOV20] Carlos Quinto Huamán, Ana Lucila Sandoval Orozco, and Luis Javier García Villalba. Authentication and integrity of smartphone videos through multimedia container structure analysis. Future Generation Computer Systems, 108 (??):15–33, July 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20300078.

Herrera:2020:SQN

[HPP20] José Herrera, Denis Parra, and Barbara Poblete. Social QA in non-CQA platforms. Future Generation Computer Systems, 105 (??):631–649, April 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0167739X19308180.

Hao:2020:DTM

[HPY20] Fei Hao, Zheng Pei, and Laurence T. Yang. Diversified top-k maximal clique detection in Social Internet of Things. Future Generation Computer Systems, 107(??):408–417, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19302195.

Huang:2020:AIC

[HQLH20] Fenghua Huang, Xinjiu Qi, Chuanlin Li, and Wei Hu. Aerial image classification by learning quality-aware spatial pyramid model. Future Generation Computer Systems, 111(??):271–277, October 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20300601.

Hou:2021:RCM

[HRGL21] Rui Hou, Guowen Ren, Wei Gao, and Lijun Liu. Research on cyberspace multi-objective security algorithm and decision mechanism of energy Internet. Future Generation Computer Systems, 120(??): 119–124, July 2021. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21000595.

Hossain:2020:TEA

[HRM20] M. Shamim Hossain, Md. Abdur Rahman, and Ghulam Muhammad. Towards energy-aware cloudoriented cyber-physical therapy system. Future Generation Computer Systems, 105(??):800–813, April 2020.

CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167739X1731885X.

Hao:2021:ACF

[HRX^{+}21] Xiaohan Hao, Wei Ren, Ruoting Xiong, Tianqing Zhu, and Kim-Kwang Raymond Choo. Asymmetric cryptographic functions based on generative adversarial neural networks for Internet of Things. Future Generation Computer Systems, 124(??):243–253, November 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21001801.

Hu:2021:EAR

[HRY^{+}21] Tao Hu, Quan Ren, Peng Yi, Ziyong Li, Julong Lan, Yuxiang Hu, and Qian Li. An efficient approach to robust controller placement for link failures in softwaredefined networks. Future Generation Computer Systems, 124(??):187–205, November 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21001722.

Hong:2021:OAD

[HS21] Huihui Hong and Xiaoxiao Sun. Obstacle avoidance dynamic control of manipulator based on space operator algebra. Future Generation Computer Systems, 123(??):201–205, October 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2100159X.

Hussain:2024:RAI

[HS24] Razin Farhan Hussain and Mohsen Amini Salehi. Resource allocation of industry 4.0 micro-service applications across serverless fog federation. Future Generation Computer Systems, 154(??):479–490, May 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24000177.

He:2022:DCA

[HSGX22] Xiaoyu He, Suixiang Shi, Xiulin Geng, and Lingyu Xu. Dynamic co-attention networks for multi-horizon forecasting in multivariate time series. Future Generation Computer Systems, 135(??):72–84, October 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22001583.

He:2020:TRT

[HSGY20] Guobiao He, Wei Su, Shuai Gao, and Jiarui Yue. TD-Root: a trustworthy decentralized DNS root management architecture based on permissioned blockchain. Future Generation Computer Systems, 102(??):912–924, January 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19312762.

He:2022:OAB

[HSR^{+}22] Xingqiu He, Yuhang Shen, Jing Ren, Sheng Wang, Xiong Wang, and Shizhong Xu. An online auctionbased incentive mechanism for soft-deadline tasks in Collaborative Edge Computing. Future Generation Computer Systems, 137(??):1–13, December 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22002291.

Hassan:2020:SER

[HSS20] Hadeer A. Hassan, Sameh A. Salem, and Elsayed M. Saad. A smart energy and reliability aware scheduling algorithm for workflow execution in DVFS-enabled cloud environment. Future Generation Computer Systems, 112(??):431–448, November 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19322307.

Husch:2020:CCB

[HSvB20] Marc Hüsch, Bruno U.

Schyska, and Lueder von Bremen. CorClustSTcorrelation-based clustering of big spatio-temporal datasets. Future Generation Computer Systems, 110(??):610–619, September 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17313353.

Haque:2022:BEF

[HT22] Md Enamul Haque and Mehmet Engin Tozal. Byte embeddings for file fragment classification. Future Generation Computer Systems, 127(??):448–461, February 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21003654.

Haugerud:2021:DSP

[HTAY21] Hårek Haugerud, Huy Nhut Tran, Nadjib Aitsaadi, and Anis Yazidi. A dynamic and scalable parallel Network Intrusion Detection System using intelligent rule ordering and network function virtualization. Future Generation Computer Systems, 124(??):254–267, November 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21001874.

Huynh-The:2023:BMR

[HTGW^{+}23] Thien Huynh-The, Thippa Reddy Gadekallu, Weizheng Wang, Gokul Yenduri, Pasika Ranaweera, Quoc-Viet Pham, Daniel Benevides da Costa, and Madhusanka Liyanage. Blockchain for the metaverse: a review. Future Generation Computer Systems, 143(??):401–419, June 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23000493.

Horta:2021:EKD

[HTLM21] Vitor A. C. Horta, Ilaria Tiddi, Suzanne Little, and Alessandra Mileo. Extracting knowledge from deep neural networks through graph analysis. Future Generation Computer Systems, 120(??):109–118, July 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21000613.

He:2021:PFI

[HTXW21] Xiang He, Zhiying Tu, Xiaofei Xu, and Zhongjie Wang. Programming framework and infrastructure for self-adaptation and optimized evolution method for microservice systems in cloud–edge environments. Future Generation Computer Systems, 118(??):263–281, May 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167739X21000170.

Hu:2020:MET

[Hu20] Chia-Cheng Hu. Minimizing executing and transmitting time of task scheduling and resource allocation in C-RANs. Future Generation Computer Systems, 108(??):349–360, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19321995.

Hu:2021:MSG

[Hu21]  Su Hu.  Massive-scale

graph-clustering-based data management based on multi-metrics. Future Generation Computer Systems, 118(??):432–437, May 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0167739X20330806.

Han:2023:PSB

[HWH^{+}23a] Wei Han, Junchao Wang, Shike Hou, Tong Bai, Gwanggil Jeon, and Joel J. P. C. Rodrigues. An PPG signal and body channel based encryption method for WBANs. Future Generation Computer Systems, 141(??):704–712, April 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22003818.

Huang:2023:EDB

[HWH^{+}23b] Ruimin Huang, Huaqiang Wang, Xiaoqiao Huang, Yonghang Tai, Feiyan Cheng, and Junsheng Shi. Edge device-based realtime implementation of CycleGAN for the colorization of infrared video. Future Generation Computer Systems, 149(??):402–415, December 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23002984.

Hou:2020:QPV

[HWQ^{+}20] Aiqin Hou, Chase Q. Wu, Ruimin Qiao, Liudong Zuo, Michelle M. Zhu, Dingyi Fang, Weike Nie, and Feng Chen. QoS provisioning for various types of deadline-constrained bulk data transfers between data centers. Future Generation Computer Systems, 105 (??):162–174, April 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167739X19304856.

Hussain:2022:DCE

[HWR^{+}22] Mehboob Hussain, Lian-Fu Wei, Amir Rehman, Fakhar Abbas, Abid Hussain, and Muqadar Ali. Deadlineconstrained energy-aware workflow scheduling in geographically distributed cloud data centers. Future Generation Computer Systems, 132(??):211–222, July 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22000619.

Hu:2021:DCC

[HX21] Su Hu and Yinhao Xiao. Design of cloud computing task offloading algorithm based on dynamic multiobjective evolution. Future Generation Computer Systems, 122(??):144–148, September 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21001175.

He:2023:EEC

[HXL^{+}23] Zhenli He, Yanan Xu, Di Liu, Wei Zhou, and Keqin Li. Energy-efficient computation offloading strategy with task priority in cloud assisted multi-access edge computing. Future Generation Computer Systems, 148(??):298–313, November 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23002352.

Huang:2023:ULS

[HXWX23] Jinshu Huang, Haidong Xie, Chunlin Wu, and Xueshuang Xiang. Union label smoothing adversarial training: Recognize small perturbation attacks and reject larger perturbation attacks balanced. Future Generation Computer Systems, 148(??):600–609, November 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23002480.

Hu:2021:ROS

[HY21] Su Hu and Hua Yin. Research on the optimum synchronous network search data extraction based on swarm intelligence algorithm. Future Generation Computer Systems, 125(??):151–155, December 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2100145X.

He:2021:EPS

[HYC^{+}21] Xin He, Yapeng Yao, Zhiwen Chen, Jianhua Sun, and Hao Chen. Efficient parallel A^{⋆} search on multi-GPU system. Future Generation Computer Systems, 123(??):35–47, October 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21001321.

Huang:2023:OCB

[HYC^{+}23] Jie Huang, Fan Yang, Chinmay Chakraborty, Zhiwei Guo, Huiyan Zhang, Li Zhen, and Keping Yu. Opportunistic capacity based resource allocation for 6G wireless systems with network slicing. Future Generation Computer Systems, 140(??):390–401, March 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22003557.

Hu:2020:FEF

[HYL^{+}20] Tao Hu, Peng Yi, Julong Lan, Yuxiang Hu, and Penghao Sun. FTLink: Efficient and flexible link fault tolerance scheme for data plane in softwaredefined networking. Future Generation Computer Systems, 111(??):381–400, October 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19308441.

Hu:2020:MUC

[HYRZ20] Yujiao Hu, Yuan Yao, Qian Ren, and Xingshe Zhou. 3D multi-UAV cooperative velocity-aware motion planning. Future Generation Computer Systems, 102(??):762–774, January 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19313597.

He:2022:TTN

[HYWY22] Xin He, Haochen Yang, Guanghui Wang, and Junyang Yu. Towards trusted node selection using blockchain for crowdsourced abnormal data detection. Future Generation Computer Systems, 133(??):320–330, August 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22001017.

Huang:2020:IUP

[HZ20] Xinyi Huang and Yuexin Zhang. Indistinguishability and unextractablility of password-based authentication in blockchain. Future Generation Computer Systems, 112(??):561–566, November 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20300327.

Hu:2020:CCS

[HZdLZ20] Yang Hu, Huan Zhou, Cees de Laat, and Zhiming Zhao. Concurrent container scheduling on heterogeneous clusters with multiresource constraints. Future Generation Computer Systems, 102(??):562–573, January 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19309446.

Huang:2021:ASB

[HZL^{+}21] Chenxi Huang, Guokai Zhang, Yiwen Lu, Yisha Lan, Sirui Chen, and Siyuan Guo. Automatic segmentation of bioabsorbable vascular stents in intravascular optical coherence images using weakly supervised attention network. Future Generation Computer Systems, 114(??):427–434, January 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19330298.

Han:2022:MSR

[HZL22] Dong Han, Zhengkun Zhao, and Ke Li. MeterSSD: Scale and rotate invariance lightweight detector. Future Generation Computer Systems, 135(??):223–233, October 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22001637.

He:2021:IAM

[HZLH21] Ming He, Xudong Zhao, Yu Lu, and Yi Hu. An improved AlexNet model for automated skeletal maturity assessment using hand X-ray images. Future Generation Computer Systems, 121(??):106–113, August 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21000996.

Han:2021:IEP

[HZPS21] Jing Han, Zixing Zhang, Maja Pantic, and Björn Schuller. Internet of emotional people: Towards continual affective computing cross cultures via audiovisual signals. Future Generation Computer Systems, 114(??):294–306, January 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20302594.

Huang:2023:REC

[HZS^{+}23] Yu Huang, Heli Zhang, Xun Shao, Xi Li, and Hong Ji. RoofSplit: an edge computing framework with heterogeneous nodes collaboration considering optimal CNN model splitting. Future Generation Computer Systems, 140(??):79– 90, March 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22003181.

Hu:2022:IET

[HZT^{+}22] Xiaoyan Hu, Cheng Zhu, Zhongqi Tong, Wenjie Gao, Guang Cheng, Ruidong Li, Hua Wu, and Jian Gong. Identifying Ethereum traffic based on an active node library and DEVp2p features. Future Generation Computer Systems, 132(??):162–177, July 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22000565.

Huajun:2019:CAB

[HZX^{+}19] Lin Huajun, Liu Zhao, Ma Xuemei, Wang Tingting, Xu Bodong, Zhou Xiaona, and Zhang Zhongtao. Correlation analysis of biochemical indicators in common bile duct stone patients with negative magnetic resonance cholangiopancreatography. Future Generation Computer Systems, 98(??):530– 535, September 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18326232.

See retraction notice [HZX^{+}20].

Huajun:2020:RNC

[HZX^{+}20] Lin Huajun, Liu Zhao, Ma Xuemei, Wang Tingting, Xu Bodong, Zhou Xiaona, and Zhang Zhongtao. Retraction notice to “Correlation analysis of biochemical indicators in common bile duct stone patients with negative magnetic resonance cholangiopancreatography” [Future Gener. Comput. Syst. 98 (2019) 530–535]. Future Generation Computer Systems, 107(??): 1146, June 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20306488.

See [HZX^{+}19].

Hu:2024:DMS

[HZX^{+}24] Na Hu, Dafang Zhang, Kun Xie, Wei Liang, Kuan-Ching Li, and Albert Y. Zomaya. Dynamic multi-scale spatial– temporal graph convolutional network for traffic flow prediction. Future Generation Computer Systems, 158(??):323–332, September 2024. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24001948.

Hou:2020:FGF

[HZZ^{+}20] Jianwei Hou, Minjian Zhang, Ziqi Zhang, Wenchang Shi, Bo Qin, and Bin Liang. On the fine-grained fingerprinting threat to software-defined networks. Future Generation Computer Systems, 107(??): 485–497, June 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19305321.

Iqbal:2020:SNF

[IA20] Zafar Iqbal and Zahid Anwar. SCERM — a novel framework for automated management of cyber threat response activities. Future Generation Computer Systems, 108(??):687–708, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19311586.

Iqbal:2023:CMR

[IA23] Shahid Md. Asif Iqbal and Asaduzzaman. Cache-MAB: a reinforcement learning-based hybrid caching scheme in named data networks. Future Generation Computer Systems, 147(??):163–178, October 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23001723.

Iqbal:2024:CMH

[IA24] Shahid Md. Asif Iqbal and Asaduzzaman. Cache-MCDM : a hybrid caching scheme in Mobile Named Data Networks based on multi-criteria decision making. Future Generation Computer Systems, 154 (??):344–358, May 2024. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167739X24000165.

Inoubli:2022:DIA

[IAM^{+}22] Wissem Inoubli, Sabeur Aridhi, Haithem Mezni, Mondher Maddouri, and Engelbert Mephu Nguifo. A distributed and incremental algorithm for largescale graph clustering. Future Generation Computer Systems, 134(??):334–347, September 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22001376.

Isafiade:2020:SMP

[IB20] Omowunmi E. Isafiade and Antoine B. Bagula. Series mining for public safety advancement in emerging smart cities. Future Generation Computer Systems, 108(??):777–802, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1932432X.

Iannucci:2020:HMF

[ICBB20] Stefano Iannucci, Valeria Cardellini, Ovidiu Daniel Barba, and Ioana Banicescu. A hybrid modelfree approach for the nearoptimal intrusion response control of non-stationary systems. Future Generation Computer Systems, 109(??):111–124, August 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19320424.

Islas-Cota:2022:SRI

[ICGGAR22] Eduardo Islas-Cota, J. Octavio Gutierrez-Garcia, Christian O. Acosta, and Luis-Felipe Rodríguez. A systematic review of intelligent assistants. Future Generation Computer Systems, 128(??):45–62, March 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21003812.

Iannucci:2021:EFS

[ICW21] Stefano Iannucci, Emiliano Casalicchio, and Byron Williams. Editorial for FGCS special issue: Advances in selfprotecting systems. Future Generation Computer Systems, 123(??):178–180, October 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21001552.

Iqbal:2020:BDA

[IDM^{+}20] Rahat Iqbal, Faiyaz Doctor, Brian More, Shahid Mahmud, and Usman Yousuf. Big Data analytics and computational intelligence for cyber-physical systems: Recent trends and state of the art applications. Future Generation Computer Systems, 105(??):766–778, April 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17323282.

Islam:2020:FTB

[IHA^{+}20] Naveed Islam, Khalid Haseeb, Ahmad Almogren, Ikram Ud Din, Mohsen Guizani, and Ayman Altameem. A framework for topological based map building: a solution to autonomous robot navigation in smart cities. Future Generation Computer Systems, 111(??):644–653, October 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19324227.

Ianni:2020:FEB

[IMM^{+}20] Michele Ianni, Elio Masciari, Giuseppe M. Mazzeo, Mario Mezzanzanica, and Carlo Zaniolo. Fast and effective Big Data exploration by clustering. Future Generation Computer Systems, 102(??):84–94, January 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19303838.

Imtiaz:2021:DCD

[IMuI^{+}21] Zar Bakht Imtiaz, Awais Manzoor, Saif ul Islam, Malik Ali Judge, Kim-Kwang Raymond Choo, and Joel J. P. C. Rodrigues. Discovering communities from disjoint complex networks using Multi-Layer Ant Colony Optimization. Future Generation Computer Systems, 115(??):659–670, February 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20329812.

Isik-Polat:2023:AAR

[IPPK23] Ece Isik-Polat, Gorkem Polat, and Altan Kocyigit. ARFED: Attack-resistant federated averaging based on outlier elimination. Future Generation Computer Systems, 141(??):626–650, April 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22004083.

Ibrahim:2022:NDP

[ISD22] Omar Adel Ibrahim, Savio Sciancalepore, and Roberto Di Pietro. Noise2Weight: On detecting payload weight from drones acoustic emissions. Future Generation Computer Systems, 134(??):319–333, September 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22001224.

Irfan:2022:RCA

[ISUC22] Muhammad Irfan, Abdurrashid Ibrahim Sanka, Zahid Ullah, and Ray C. C. Cheung. Reconfigurable content-addressable memory (CAM) on FPGAs: a tutorial and survey. Future Generation Computer Systems, 128(??):451–465, March 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21003836.

Ismayilov:2020:NNB

[IT20] Goshgar Ismayilov and Haluk Rahmi Topcuoglu. Neural network based multi-objective evolutionary algorithm for dynamic workflow scheduling in cloud computing. Future Generation Computer Systems, 102(??):307–322, January 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19306983.

Imtiaz:2021:DDI

[IuRJ^{+}21] Syed Ibrahim Imtiaz, Saif ur Rehman, Abdul Rehman Javed, Zunera Jalil, Xuan Liu, and Waleed S. Alnumay. DeepAMD: Detection and identification of Android malware using high-efficient deep artificial neural network. Future Generation Computer Systems, 115(??):844–856, February 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2032985X.

Jadallah:2020:SCB

[JA20] Hiba Jadallah and Zaher Al Aghbari. SwapQt: Cloudbased in-memory indexing of dynamic spatial data. Future Generation Computer Systems, 106(??): 360–373, May 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19323283.

Jararweh:2020:EFF

[JAAAZB20] Yaser Jararweh, Mahmoud Al-Ayyoub, Du’a Al-Zoubi, and Elhadj Benkhelifa. An experimental framework for future smart cities using data fusion and software defined systems: the case of environmental monitoring for smart healthcare. Future Generation Computer Systems, 107(??):883–897, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17312013.

Jin:2023:MSF

[JAC^{+}23] Chao Jin, David Abramson, Jake Carroll, Zhengchun Liu, and Rajkumar Kettimuthu. Moving small files in a networked environment. Future Generation Computer Systems, 139(??):167–180, February 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22003004.

Jansen:2020:CCF

[JAS^{+}20] Christoph Jansen, Jonas Annuscheit, Bruno Schilling, Klaus Strohmenger, Michael Witt, Felix Bartusch, Christian Herta, Peter Hufnagl, and Dagmar Krefting. Curious containers: a framework for computational reproducibility in life sciences with support for deep learning applications. Future Generation Computer Systems, 112(??):209–227, November 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19318096.

Jurczuk:2024:AMR

[JCK24] Krzysztof Jurczuk, Marcin Czajkowski, and Marek Kretowski. Adaptive inmemory representation of decision trees for GPUaccelerated evolutionary induction. Future Generation Computer Systems, 153 (??):419–430, April 2024. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167739X23004636.

Jaksic:2020:HPF

[JCP^{+}20] Zoran Jakšić, Nicola Cadenelli, David Buchaca Prats, Jordà Polo, Josep Lluís Berral Garcia, and David Carrera Perez. A highly parameterizable framework for conditional restricted Boltzmann machine based workloads accelerated with FPGAs and OpenCL. Future Generation Computer Systems, 104(??):201–211, March 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19313676.

Jiang:2023:PIN

[JCW^{+}23] Wenbin Jiang, Yuhao Chen, Suyang Wen, Long Zheng, and Hai Jin. PDAS: Improving network pruning based on Progressive Differentiable Architecture Search for DNNs. Future Generation Computer Systems, 146(??):98–113, September 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23001462.

Jin:2021:SAD

[JCX^{+}21] Yi Jin, Jiawei Cai, Jiawei Xu, Yuxiang Huan, Yulong Yan, Bin Huang, Yongliang Guo, Lirong Zheng, and Zhuo Zou. Self-aware distributed deep learning framework for heterogeneous IoT edge devices. Future Generation Computer Systems, 125(??):908–920, December 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21002715.

Jiang:2020:SBB

[JGL^{+}20] Peng Jiang, Fuchun Guo, Kaitai Liang, Jianchang Lai, and Qiaoyan Wen. Searchain: Blockchainbased private keyword search in decentralized storage. Future Generation Computer Systems, 107 (??):781–792, June 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0167739X17318630.

Jayanetti:2022:DRL

[JHB22] Amanda Jayanetti, Saman Halgamuge, and Rajkumar Buyya. Deep reinforcement learning for energy and time optimized scheduling of precedence-constrained tasks in edge–cloud computing environments. Future Generation Computer Systems, 137(??):14–30, December 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22002230.

Jayasekara:2020:UMO

[JHK20] Sachini Jayasekara, Aaron Harwood, and Shanika Karunasekera. A utilization model for optimization of checkpoint intervals in distributed stream processing systems. Future Generation Computer Systems, 110(??):68–79, September 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19320102.

Jiang:2021:FAC

[Jia21] Lianyuan Jiang. A fast and accurate circle detection algorithm based on random sampling. Future Generation Computer Systems, 123(??):245–256, October 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21001606.

Ji:2023:BJN

[JJZ^{+}23] Hangxu Ji, Su Jiang, Yuhai Zhao, Gang Wu, Guoren Wang, and George Y. Yuan. BS-Join: a novel and efficient mixed batch-stream join method for spatiotemporal data management in flink. Future Generation Computer Systems, 141 (??):67–80, April 2023. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22003776.

Jain:2020:TRM

[JKS20a] Deepak Kumar Jain, Akshi Kumar, and Vibhuti Sharma. Tweet recommender model using adaptive neuro-fuzzy inference system. Future Generation Computer Systems, 112(??):996–1009, November 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19306521.

Jindal:2020:IEB

[JKS20b] Anish Jindal, Neeraj Kumar, and Mukesh Singh. Internet of energy-based demand response management scheme for smart homes and PHEVs using SVM. Future Generation Computer Systems, 108(??):1058–1068, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17320976.

Jindal:2020:UFB

[JKS20c] Anish Jindal, Neeraj Kumar, and Mukesh Singh. A unified framework for big data acquisition, storage, and analytics for demand response management in smart cities. Future Generation Computer Systems, 108(??):921–934, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17324780.

Jiang:2021:IVA

[JL21] Yujia Jiang and Xin Liu.

Image/video aesthetic quality management based on fusing 3D CNN and shallow features. Future Generation Computer Systems, 118(??):118–123, May 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20330892.

Ji:2020:RNI

[JLC^{+}20] Yanli Ji, Wei Lu, Guowei Che, Mei Yang, and Lunxu Liu. Retraction notice to “Inhibition of KRAS gene mutation on non-small cell lung cancer and its effect on circulating tumor cells” [future gener. comput. syst. 98 (2019) 104–108]. Future Generation Computer Systems, 111(??):937, October 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19332996.

See [YWG^{+}19].

Jiang:2022:PFR

[JLK22] Hongling Jiang, Jinzhi Lin, and Haiyan Kang. FGMD: a robust detector against adversarial attacks in the IoT network. Future Generation Computer Systems, 132(??):194–210, July 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22000620.

Jiang:2021:AIL

[JLP^{+}21] Wanchun Jiang, Haoyang Li, Lijuan Peng, Jia Wu, Chang Ruan, and Jianxin Wang. Analysis and improvement of the latencybased congestion control algorithm DX. Future Generation Computer Systems, 123(??):206–218, October 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21001485.

Jansen:2023:VAE

[JLS^{+}23] Christoph Jansen, Björn Lindequist, Klaus Strohmenger, Daniel Romberg, Tobias Küster, Nick Weiss, Michael Franz, Lars Ole Schwen, Theodore Evans, André Homeyer, and Norman Zerbe. The vendoragnostic EMPAIA platform for integrating AI applications into digital pathology infrastructures. Future Generation Computer Systems, 140(??):209–224, March 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22003405.

Jiang:2021:SSM

[JLT^{+}21] Du Jiang, Gongfa Li, Chong Tan, Li Huang, Ying Sun, and Jianyi Kong. Semantic segmentation for multiscale target based on object recognition using the improved Faster-RCNN model. Future Generation Computer Systems, 123(??):94–104, October 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21001400.

Jiang:2023:ADR

[JLW^{+}23] Tian Jiang, Yunqi Liu, Xuemeng Wu, Mohan Xu, and Xiaohui Cui. Application of deep reinforcement learning in attacking and protecting structural features-based malicious PDF detector. Future Generation Computer Systems, 141(??):325–338, April 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22003740.

Johnsson:2020:TEU

[JM20] Björn A. Johnsson and Boris Magnusson. Towards end-user development of graphical user interfaces for Internet of Things. Future Generation Computer Systems, 107(??):670–680, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17321660.

Janjua:2021:TAT

[JMA^{+}21] Faisal Janjua, Asif Masood, Haider Abbas, Imran Rashid, and Malik M. Zaki Murtaza Khan. Textual analysis of traitorbased dataset through semi supervised machine learning. Future Generation Computer Systems, 125(??):652–660, December 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21002314.

John:2024:MGW

[JMHB24] Joseph John, Josh Milthorpe, Thomas Herault, and George Bosilca. Multi-GPU work sharing in a task-based dataflow programming model. Future Generation Computer Systems, 156(??):313–324, July 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24000888.

Jurado:2022:CMG

[JPJO22] Juan M. Jurado, Emilio J.

Padrón, J. Roberto Jiménez, and Lidia Ortega. An outof-core method for GPU image mapping on large 3D scenarios of the real world. Future Generation Computer Systems, 134(??):66– 77, September 2022. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22000978.

Jan:2021:SDT

[JPMR21] Aiman Jan, Shabir A.

Parah, Bilal A. Malik, and Mamoon Rashid. Secure data transmission in IoTs based on CLoG edge detection. Future Generation Computer Systems, 121(??):59–73, August 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21000856.

Jin:2020:VTS

[JPW20] Song Jin, Songwei Pei, and Yu Wang. A variation tolerant scheme for memristor crossbar based neural network designs via two-phase weight mapping and memristor programming. Future Generation Computer Systems, 106(??):270–276, May 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19314128.

Jia:2022:FBL

[JQZ^{+}22] Bing Jia, Wenling Qiao, Zhaopeng Zong, Shuai Liu, Mohammad Hijji, Javier Del Ser, and Khan Muhammad. A fingerprintbased localization algorithm based on LSTM and data expansion method for sparse samples. Future Generation Computer Systems, 137(??):380–393, December 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22002552.

Jiang:2022:RFC

[JR22] Chunhong Jiang and Guanghao Ren. Regional flow control in peak travel period based on fuzzy algorithm. Future Generation Computer Systems, 126(??):279–283, January 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21003241.

Jahan:2020:MKM

[JRW^{+}20] Sharmin Jahan, Ian Riley, Charles Walter, Rose F. Gamble, Matt Pasco, Philip K. McKinley, and Betty H. C. Cheng. MAPE–K/MAPE– SAC: an interaction framework for adaptive systems with security assurance cases. Future Generation Computer Systems, 109(??):197–209, August 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19320527.

Junoh:2023:CLA

[JSP23] Suhardi Azliy Junoh, Santosh Subedi, and Jae-Young Pyun. Crowdsourcing landmark-assisted localization with deep learning. Future Generation Computer Systems, 144(??):256–270, July 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23000584.

Jagtap:2021:HBK

[JSV21] Sujeet S. Jagtap, Shankar Sriram V. S., and Subramaniyaswamy V. A hypergraph based Kohonen map for detecting intrusions over cyber–physical systems traffic. Future Generation Computer Systems, 119(??):84–109, June 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21000467.

Jangra:2022:EAV

[JT22] Shalini Jangra and Durga Toshniwal. Efficient algorithms for victim item selection in privacy-preserving utility mining. Future Generation Computer Systems, 128(??):219–234, March 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21003988.

Jaiswal:2021:EAC

[JTGH21] Amit Kumar Jaiswal, Prayag Tiwari, Sahil Garg, and M. Shamim Hossain. Entityaware capsule network for multi-class classification of big data: a deep learning approach. Future Generation Computer Systems, 117(??):1–11, April 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20330363.

Jiang:2020:PKB

[JVH^{+}20] Jiaojiao Jiang, Steve Versteeg, Jun Han, M. D. Arafat Hossain, and Jean-Guy Schneider. A positional keyword-based approach to inferring fine-grained message formats. Future Generation Computer Systems, 102(??):369–381, January 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18329327.

Jing:2022:EDB

[JWC22] Chao Jing, Yun Wu, and Chaoyuan Cui. Ensemble dynamic behavior detection method for adversarial malware. Future Generation Computer Systems, 130(??):193–206, May 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21004945.

Ji:2021:TSH

[JWYÍ21] Zhenyan Ji, Mengdan Wu, Hong Yang, and José Enrique Armendáriz Íñigo. Temporal sensitive heterogeneous graph neural network for news recommendation. Future Generation Computer Systems, 125(??):324–333, December 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21002028.

Jin:2022:STT

[JWZ^{+}22] Guangyin Jin, Min Wang, Jinlei Zhang, Hengyu Sha, and Jincai Huang. STGNN–TTE: Travel time estimation via spatial– temporal graph neural network. Future Generation Computer Systems, 126(??):70–81, January 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21002740.

Jiang:2024:MME

[JXYC24] Qiangqiang Jiang, Xu Xin, Libo Yao, and Bo Chen. METSM: Multiobjective energy-efficient task scheduling model for an edge heterogeneous multiprocessor system. Future Generation Computer Systems, 152(??):207–223, March 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23003990.

Jeremiah:2024:DTA

[JYP24] Sekione Reward Jeremiah, Laurence Tianruo Yang, and Jong Hyuk Park. Digital twin-assisted resource allocation framework based on edge collaboration for vehicular edge computing. Future Generation Computer Systems, 150(??):243–254, January 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23003278.

Jabeur:2020:TBI

[JYSH20] Nafaâ Jabeur, Ansar Ul-Haque Yasar, Elhadi Shakshuki, and Hedi Haddad. Toward a bioinspired adaptive spatial clustering approach for IoT applications. Future Generation Computer Systems, 107(??):736–744, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17309846.

Jiang:2023:FRM

[JYSH23] Liang Jiang, Jingjing Yao, Leilei Shi, and Zixuan Han. A fusion recommendation model based on mutual information and attention learning in heterogeneous social networks. Future Generation Computer Systems, 148(??):128–138, November 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23002133.

Jan:2021:AEL

[JZK^{+}21] Mian Ahmad Jan, Muhammad Zakarya, Muhammad Khan, Spyridon Mastorakis, Varun G. Menon, Venki Balasubramanian, and Ateeq Ur Rehman. An AI-enabled lightweight data fusion and load optimization approach for Internet of Things. Future Generation Computer Systems, 122(??):40–51, September 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21001011.

Jiang:2020:EPD

[JZL^{+}20] Wenbin Jiang, Yangsong Zhang, Pai Liu, Jing Peng, Laurence T. Yang, Geyan Ye, and Hai Jin. Exploiting potential of deep neural networks by layer-wise finegrained parallelism. Future Generation Computer Systems, 102(??):210–221, January 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1832716X.

Jin:2024:FIN

[JZL^{+}24] Zhigang Jin, Junyi Zhou, Bing Li, Xiaodong Wu, and Chenxu Duan. FL-IIDS: a novel federated learningbased incremental intrusion detection system. Future Generation Computer Systems, 151(??):57–70, February 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23003503.

Ju:2022:NNG

[JZM^{+}22] Zhen Ju, Huiling Zhang, Jintao Meng, Jingjing Zhang, Jianping Fan, Yi Pan, Weiguo Liu, Xuelei Li, and Yanjie Wei. nGIA: a novel Greedy Incremental Alignment based algorithm for gene sequence clustering. Future Generation Computer Systems, 136(??):221–230, November 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22001959.

Ji:2023:LLA

[JZZ^{+}23] Shan Ji, Jiale Zhang, Yongjing Zhang, Zhaoyang Han, and Chuan Ma. LAFED: a lightweight authentication mechanism for blockchain-enabled federated learning system. Future Generation Computer Systems, 145(??):56–67, August 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23000912.

Ji:2021:BTQ

[JZZD21] Fang Ji, Heqing Zhang, Zijiang Zhu, and Weihuang Dai. Blog text quality assessment using a 3D CNN-based statistical framework. Future Generation Computer Systems, 116(??):365–370, March 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20330028.

Kolomvatsos:2021:PUD

[KA21] Kostas Kolomvatsos and Christos Anagnostopoulos. Proactive, uncertaintydriven queries management at the edge. Future Generation Computer Systems, 118(??):75–93, May 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21000029.

Kim:2022:SES

[KA22] Jun Suk Kim and Chang Wook Ahn. Size-efficient sparse population for strictly structured quantum genetic algorithm. Future Generation Computer Systems, 135(??):159–171, October 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22001595.

Kolomvatsos:2024:APD

[KA24] Kostas Kolomvatsos and Christos Anagnostopoulos. Autonomous proactive data management in support of pervasive edge applications. Future Generation Computer Systems, 155(??): 108–120, June 2024. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X24000487.

Khan:2021:SUE

[KAA^{+}21] Abd Ullah Khan, Ghulam Abbas, Ziaul Haq Abbas, Wali Ullah Khan, and Muhammad Waqas. Spectrum utilization efficiency in CRNs with hybrid spectrum access and channel reservation: a comprehensive analysis under prioritized traffic. Future Generation Computer Systems, 125(??):726–742, December 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21002867.

Kader:2020:PDN

[Kad20] Md. Fazlul Kader. A powerdomain NOMA based overlay spectrum sharing scheme. Future Generation Computer Systems, 105 (??):222–229, April 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0167739X19306764.

Khorasani:2020:RMF

[KAF^{+}20] Neda Khorasani, Saeid Abrishami, Mehdi Feizi, Mahdi Abolfazli Esfahani, and Faeze Ramezani. Resource management in the federated cloud environment using Cournot and Bertrand competitions. Future Generation Computer Systems, 113(??):391–406, December 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19324422.

Khalid:2023:IEN

[KAF^{+}23] Shah Khalid, Aftab Alam, Muhammad Fayaz, Fakhrud Din, Sehat Ullah, and Shabir Ahmad. Investigating the effect of network latency on users’ performance in Collaborative Virtual Environments using navigation aids. Future Generation Computer Systems, 145(??):68–76, August 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23000663.

Kamal:2023:EIL

[KAH^{+}23] A. H. M. Kamal, Md. Golam Rabiul Alam, Md Rafiul Hassan, Tasnim Sakib Apon, and Mohammad Mehedi Hassan. Explainable indoor localization of BLE devices through RSSI using recursive continuous wavelet transformation and XGBoost classifier. Future Generation Computer Systems, 141(??):230–242, April 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22003624.

Kumar:2024:DTE

[KAJ^{+}24] Randhir Kumar, Ahamed Aljuhani, Danish Javeed, Prabhat Kumar, Shareeful Islam, and A. K. M. Najmul Islam. Digital twinsenabled zero touch network: a smart contract and explainable AI integrated cybersecurity framework. Future Generation Computer Systems, 156(??):191–205, July 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X24000608.

Khan:2020:IPC

[KAK20] Kashif Mehboob Khan, Junaid Arshad, and Muhammad Mubashir Khan. Investigating performance constraints for blockchain based secure e-voting system. Future Generation Computer Systems, 105 (??):13–26, April 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19310805.

Khan:2023:HHG

[KAK^{+}23] Muhammad Attique Khan, Habiba Arshad, Wazir Zada Khan, Majed Alhaisoni, Usman Tariq, Hany S. Hussein, Hammam Alshazly, Lobna Osman, and Ahmed Elashry. HGRBOL^{2}: Human gait recognition for biometric application using Bayesian optimization and extreme learning machine. Future Generation Computer Systems, 143(??): 337–348, June 2023. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23000468.

Khedo:2020:IWS

[KBG20] Kavi K. Khedo, Yasdeo Bissessur, and Datta S. Goolaub. An inland wireless sensor network system for monitoring seismic activity. Future Generation Computer Systems, 105(??):520–532, April 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19301785.

Kondoro:2021:RTP

[KBTM21] Aron Kondoro, Imed Ben Dhaou, Hannu Tenhunen, and Nerey Mvungi. Real time performance analysis of secure IoT protocols for microgrid communication. Future Generation Computer Systems, 116(??):1– 12, March 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20305707.

Kabra:2020:MBB

[KBTT20] Naman Kabra, Pronaya Bhattacharya, Sudeep Tanwar, and Sudhanshu Tyagi. MudraChain: Blockchainbased framework for automated cheque clearance in financial institutions. Future Generation Computer Systems, 102(??):574–587, January 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19311896.

Ko:2020:SSN

[KCB20] Ili Ko, Desmond Chambers, and Enda Barrett. Selfsupervised network traffic management for DDoS mitigation within the ISP domain. Future Generation Computer Systems, 112(??):524–533, November 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20302193.

Koraei:2023:NOM

[KCJ23] Mostafa Koraei, Juan M.

Cebrian, and Magnus Jahre. Near-optimal multiaccelerator architectures for predictive maintenance at the edge. Future Generation Computer Systems, 140(??):331–343, March 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22003533.

Kim:2024:DPF

[KCKK24] Hyochan Kim, Ji Sub Choi, Jungrae Kim, and Jong Hwan Ko. A DNN partitioning framework with controlled lossy mechanisms for edge-cloud collaborative intelligence. Future Generation Computer Systems, 154(??): 426–439, May 2024. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24000062.

Krzywaniak:2023:DGP

[KCP23] Adam Krzywaniak, Paweł Czarnul, and Jerzy Proficz. Dynamic GPU power capping with online performance tracing for energy efficient GPU computing using DEPO tool. Future Generation Computer Systems, 145(??):396–414, August 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23001267.

K:2020:CFA

[KCR20] Dhanalakshmi B. K., Srikantaiah K. C., and Venugopal K. R. Carry forward and access control for unused resources in multi sharing system of hybrid cloud. Future Generation Computer Systems, 110(??):282–290, September 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19318746.

Kristiani:2021:UDE

[KCY^{+}21] Endah Kristiani, Yuan-An Chen, Chao-Tung Yang, Chin-Yin Huang, Yu-Tse Tsan, and Wei-Cheng Chan. Using deep ensemble for influenzalike illness consultation rate prediction. Future Generation Computer Systems, 117(??):369–386, April 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20330661.

Kang:2024:ILE

[KDX^{+}24] KaiXuan Kang, Ding Ding, HuaMao Xie, LiHong Zhao, YiNong Li, and YiXuan Xie. Imitation learning enabled fast and adaptive task scheduling in cloud. Future Generation Computer Systems, 154(??):160–172, May 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24000037.

Kanbar:2022:RAD

[KF22] Asan Baker Kanbar and Kamaran Faraj. Region aware dynamic task scheduling and resource virtualization for load balancing in IoT–fog multicloud environment. Future Generation Computer Systems, 137(??):70–86, December 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22002163.

Kanbar:2023:CRA

[KF23] Asan Baker Kanbar and Kamaran Faraj. Corrigendum to “Region aware dynamic task scheduling and resource virtualization for load balancing in IoT–fog multi-cloud environment” [Future Gener. Comput. Syst. 137C (2022) 70–86].

Future Generation Computer Systems, 138(??):212, January 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2200276X.

Kuang:2020:ODM

[KGO^{+}20] Li Kuang, Tao Gong, Shuyin OuYang, Honghao Gao, and Shuiguang Deng. Offloading decision methods for multiple users with structured tasks in edge computing for smart cities. Future Generation Computer Systems, 105(??): 717–729, April 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19323325.

Khaleel:2024:DWA

[Kha24] Mustafa Ibrahim Khaleel. A dynamic weight–assignment load balancing approach for workflow scheduling in edge–cloud computing using ameliorated moth flame and rock hyrax optimization algorithms. Future Generation Computer Systems, 155(??):465–485, June 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24000700.

Knapen:2020:UPD

[KHB20] Luk Knapen, Irith Ben-Arroyo Hartman, and Tom Bellemans. Using path decomposition enumeration to enhance route choice models. Future Generation Computer Systems, 107(??):1077–1088, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17321866.

Konig:2023:WCH

[KHB23] Jemma L. König, Annika Hinze, and Judy Bowen. Workload categorization for hazardous industries: the semantic modelling of multi-modal physiological data. Future Generation Computer Systems, 141(??):369–381, April 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22003806.

Kim:2021:IPD

[KHES21] Sunggon Kim, Jaehyun Han, Hyeonsang Eom, and Yongseok Son. Improving I/O performance in distributed file systems for flash-based SSDs by access pattern reshaping. Future Generation Computer Systems, 115(??):365–373, February 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19331000.

Khan:2021:GEE

[KHH21] Shuraia Khan, Farookh Khadeer Hussain, and Omar K. Hussain. Guaranteeing end-to-end QoS provisioning in SOA based SDN architecture: a survey and open issues. Future Generation Computer Systems, 119(??):176–187, June 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21000637.

Karimiafshar:2021:RDM

[KHHT21] Aref Karimiafshar, Massoud Reza Hashemi, Mohammad Reza Heidarpour, and Adel N. Toosi. A request dispatching method for efficient use of renewable energy in fog computing environments. Future Generation Computer Systems, 114(??):631–646, January 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19327979.

Kabir:2021:SHB

[KHHV21] Muhammad Ashad Kabir, Jun Han, Md. Arafat Hossain, and Steve Versteeg. SpecMiner: Heuristicbased mining of service behavioral models from interaction traces. Future Generation Computer Systems, 117(??):59–71, April 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2033017X.

Kim:2020:PDA

[KHL20] Chumsu Kim, Bonghee Hong, and Jiwan Lee. The promotion/demotion algorithm of moving objects with large velocity differences in timeparameterized spatio-temporal index. Future Generation Computer Systems, 107 (??):645–658, June 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167739X19304066.

Kholidy:2021:AMC

[Kho21a] Hisham A. Kholidy. Autonomous mitigation of cyber risks in the cyber– physical systems. Future Generation Computer Systems, 115(??):171–187, February 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19320680.

Kholidy:2021:DIA

[Kho21b] Hisham A. Kholidy. Detecting impersonation attacks in cloud computing environments using a centric user profiling approach. Future Generation Computer Systems, 117(??):299–320, April 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20330715.

Khan:2024:EAD

[KIJ^{+}24] Murad Ali Khan, Naeem Iqbal, Harun Jamil, Faiza Qayyum, Jong-Hyun Jang, Salabat Khan, Jae-Chul Kim, and Do-Hyeun Kim. Enhanced abnormal data detection hybrid strategy based on heuristic and stochastic approaches for efficient patients rehabilitation. Future Generation Computer Systems, 154(??):101–122, May 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23004533.

Kechedzhi:2024:EQV

[KIM^{+}24] K. Kechedzhi, S. V. Isakov, S. Mandrà, B. Villalonga, X. Mi, S. Boixo, and V. Smelyanskiy. Effective quantum volume, fidelity and computational cost of noisy quantum processing experiments. Future Generation Computer Systems, 153(??):431–441, April 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23004569.

Kumari:2024:CEF

[KJ24] Nidhi Kumari and Prasanta K. Jana. Communication efficient federated learning with data offloading in fog-based IoT environment. Future Generation Computer Systems, 158(??):158–166, September 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24001936.

Kanchan:2023:EPP

[KJYC23] Sneha Kanchan, Jae Won Jang, Jun Yong Yoon, and Bong Jun Choi. Efficient and privacy-preserving group signature for federated learning. Future Generation Computer Systems, 147(??):93–106, October 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23001528.

Kaur:2020:PPR

[KK20] Harkeerat Kaur and Pritee Khanna. Privacy preserving remote multi-server biometric authentication using cancelable biometrics and secret sharing. Future Generation Computer Systems, 102(??):30–41, January 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18330553.

Kurz:2022:PTP

[KK22] Christoph F. Kurz and Adriana N. König. Predicting time preference from social media behavior. Future Generation Computer Systems, 130(??):155–163, May 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21004982.

Ko:2024:MAP

[KKL^{+}24] Haneul Ko, Yeunwoong Kyung, Jaewook Lee, Sangheon Pack, and Namseok Ko. Mobilityaware personalized handover function provisioning system in B5G networks. Future Generation Computer Systems, 157(??): 436–444, August 2024. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24001316.

Klinkenberg:2023:HEH

[KKT^{+}23] Jannis Klinkenberg, Anara Kozhokanova, Christian Terboven, Clément Foyer, Brice Goglin, and Emmanuel Jeannot. H2M: Exploiting heterogeneous shared memory architectures. Future Generation Computer Systems, 148(??):39–55, November 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23002029.

Kumar:2022:VAS

[KLA22] Nikhil Kumar, Divya Lohani, and Debopam Acharya. Vehicle accident sub-classification modeling using stacked generalization: a multisensor fusion approach. Future Generation Computer Systems, 133(??):39–52, August 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22000802.

Kong:2021:RDD

[KLW^{+}21] Xiangjie Kong, Jiaxing Li, Luna Wang, Guojiang Shen, Yiming Sun, and Ivan Lee. Recurrent-DC: a deep representation clustering model for university profiling based on academic graph. Future Generation Computer Systems, 116(??):156–167, March 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20329964.

Kim:2020:OJP

[KMCJ20] Taesik Kim, Hong Min, Eunsoo Choi, and Jinman Jung. Optimal job partitioning and allocation for vehicular cloud computing. Future Generation Computer Systems, 108(??):82– 96, July 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19321119.

Khan:2023:THP

[KMK^{+}23] Muhammad Attique Khan, Asif Mehmood, Seifedine Kadry, Nouf Abdullah Almujally, Majed Alhaisoni, Jamel Balili, Abdullah Al Hejaili, Abed Alanazi, Shtwai Alsubai, and Abdullah Alqatani. TS^{2}HGRNet: a paradigm of two stream best deep learning feature fusion assisted framework for human gait analysis using controlled environment in smart cities. Future Generation Computer Systems, 147(??):292–303, October 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23001929.

Khan:2022:XIE

[KMR^{+}22] Izhar Ahmed Khan, Nour Moustafa, Imran Razzak, M. Tanveer, Dechang Pi, Yue Pan, and Bakht Sher Ali. XSRU–IoMT: Explainable simple recurrent units for threat detection in Internet of Medical Things networks. Future Generation Computer Systems, 127(??):181–193, February 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21003563.

Koroniotis:2020:NNF

[KMS20] Nickolaos Koroniotis, Nour Moustafa, and Elena Sitnikova. A new network forensic framework based on deep learning for Internet of Things networks: a particle deep framework. Future Generation Computer Systems, 110(??):91– 106, September 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19325105.

Kopke:2023:DSB

[KMS23] Julius Köpke, Giovanni Meroni, and Mattia Salnitri. Designing secure business processes for blockchains with SecBPMN2BC. Future Generation Computer Systems, 141(??): 382–398, April 2023. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22003764.

Kamran:2021:HDA

[KNRI21] Iqra Kamran, Saeeda Naz, Imran Razzak, and Muhammad Imran. Handwriting dynamics assessment using deep neural network for early identification of Parkinson’s disease. Future Generation Computer Systems, 117(??):234–244, April 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20330442.

Kishor:2020:GTA

[KNV20] Avadh Kishor, Rajdeep Niyogi, and Bharadwaj Veeravalli. A gametheoretic approach for costaware load balancing in distributed systems. Future Generation Computer Systems, 109(??):29–44, August 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19322824.

Kolomvatsos:2022:PIS

[Kol22] Kostas Kolomvatsos. A proactive inference scheme for data-aware decision making in support of pervasive applications. Future Generation Computer Systems, 136(??):193–204, November 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22001893.

Kadadha:2020:SBB

[KOM^{+}20] Maha Kadadha, Hadi Otrok, Rabeb Mizouni, Shakti Singh, and Anis Ouali. SenseChain: a blockchain-based crowdsensing framework for multiple requesters and multiple workers. Future Generation Computer Systems, 105(??):650–664, April 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19312579.

Kadadha:2022:CBP

[KOM^{+}22] Maha Kadadha, Hadi Otrok, Rabeb Mizouni, Shakti Singh, and Anis Ouali. On-chain behavior prediction machine learning model for blockchainbased crowdsourcing. Future Generation Computer Systems, 136(??):170–181, November 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22001960.

Kong:2021:SAB

[Kon21] Lingqiang Kong. A study on the AI-based online triage model for hospitals in sustainable smart city. Future Generation Computer Systems, 125(??):59– 70, December 2021. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21002181.

Kim:2022:RTN

[KP22] Taehoon Kim and Wooguil Pak. Real-time network intrusion detection using deferred decision and hybrid classifier. Future Generation Computer Systems, 132(??):51–66, July 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22000541.

Koslovski:2024:DBW

[KPA24] Guilherme Piêgas Koslovski, Kleiton Pereira, and Paulo Roberto Albuquerque. DAG-based workflows scheduling using Actor–Critic Deep Reinforcement Learningo. Future Generation Computer Systems, 150(??):354–363, January 2024. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23003485.

Karabulut:2024:ODT

[KPGD24] Erkan Karabulut, Salvatore F. Pileggi, Paul Groth, and Victoria Degeler. Ontologies in digital twins: a systematic literature review. Future Generation Computer Systems, 153 (??):442–456, April 2024. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167739X23004739.

Kang:2022:JSB

[KPL22] Youyou Kang, Li Pan, and Shijun Liu. Job scheduling for big data analytical applications in clouds: a taxonomy study. Future Generation Computer Systems, 135(??):129–145, October 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22001649.

Khodabandehloo:2021:HCE

[KRA21] Elham Khodabandehloo, Daniele Riboni, and Abbas Alimohammadi. HealthXAI: Collaborative and explainable AI for supporting early diagnosis of cognitive decline. Future Generation Computer Systems, 116(??):168–189, March 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20330144.

Kayes:2020:ASS

[KRW^{+}20] A. S. M. Kayes, Wenny Rahayu, Paul Watters, Mamoun Alazab, Tharam Dillon, and Elizabeth Chang. Achieving security scalability and flexibility using fog-based contextaware access control. Future Generation Computer Systems, 107(??):307–323, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19323349.

Khansari:2024:SMD

[KS24] Mina Emami Khansari and Saeed Sharifian. A scalable modified deep reinforcement learning algorithm for serverless IoT microservice composition infrastructure in fog layer. Future Generation Computer Systems, 153(??):206–221, April 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23004284.

Khan:2020:IBC

[KSA^{+}20] Shah Rukh Khan, Misba Sikandar, Ahmad Almogren, Ikram Ud Din, Antonio Guerrieri, and Giancarlo Fortino. IoMT-based computational approach for detecting brain tumor. Future Generation Computer Systems, 109(??):360–367, August 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20307688.

Kim:2020:MML

[KSC20] Minsoo Kim, Ilhyun Suh, and Yon Dohn Chung. MARS: a multi-level array representation for simulation data. Future Generation Computer Systems, 111(??):419–434, October 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19311306.

Kumari:2021:MMA

[KSDR21] Kirti Kumari, Jyoti Prakash Singh, Yogesh K. Dwivedi, and Nripendra P. Rana. Multi-modal aggression identification using convolutional neural network and binary particle swarm optimization. Future Generation Computer Systems, 118(??):187–197, May 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167739X21000248.

Kostamis:2024:DME

[KSE24] Periklis Kostamis, Andreas Sendros, and Pavlos S. Efraimidis. Data management in Ethereum DApps: a cost and performance analysis. Future Generation Computer Systems, 153(??):193–205, April 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23004351.

Khan:2021:ECE

[KSH^{+}21] Tayyab Khan, Karan Singh, Mohd Hilmi Hasan, Khaleel Ahmad, G. Thippa Reddy, Senthilkumar Mohan, and Ali Ahmadian. ETERS: a comprehensive energy aware trust-based efficient routing scheme for adversarial WSNs. Future Generation Computer Systems, 125(??):921–943, December 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21002442.

Kumar:2021:IFI

[KSLC21] Gulshan Kumar, Rahul Saha, Chhagan Lal, and Mauro Conti. Internet-of-Forensic (IoF): a blockchain based digital forensics framework for IoT applications. Future Generation Computer Systems, 120 (??):13–25, July 2021. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21000686.

Kouloumpris:2024:OFT

[KSMT24] Andreas Kouloumpris, Georgios L. Stavrinides, Maria K. Michael, and Theocharis Theocharides. An optimization framework for task allocation in the edge/ hub/cloud paradigm. Future Generation Computer Systems, 155(??):354–366, June 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24000505.

Karim:2020:BDM

[KSS^{+}20] Ahmad Karim, Aisha Siddiqa, Zanab Safdar, Maham Razzaq, Syeda Anum Gillani, Huma Tahir, Sana Kiran, Ejaz Ahmed, and Muhammad Imran. Big data management in participatory sensing: Issues, trends and future directions. Future Generation Computer Systems, 107 (??):942–955, June 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167739X17311627.

Kumar:2021:DBN

[KSS^{+}21] Adarsh Kumar, Kriti Sharma, Harvinder Singh, Sagar Gupta Naugriya, Sukhpal Singh Gill, and Rajkumar Buyya. A drone-based networked system and methods for combating coronavirus disease (COVID-19) pandemic. Future Generation Computer Systems, 115(??):1–19, February 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20317064.

Kaur:2020:CIE

[KSSR20] Manjit Kaur, Dilbag Singh, Kehui Sun, and Umashankar Rawat. Color image encryption using non-dominated sorting genetic algorithm with local chaotic search based 5D chaotic map. Future Generation Computer Systems, 107(??):333–350, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19324707.

Kuo:2023:MID

[KTC23] Shu-Yu Kuo, Fan-Hsun Tseng, and Yao-Hsin Chou. Metaverse intrusion detection of wormhole attacks based on a novel statistical mechanism. Future Generation Computer Systems, 143(??):179–190, June 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23000249.

Khan:2022:WFE

[KTIB22] Tahseen Khan, Wenhong Tian, Shashikant Ilager, and Rajkumar Buyya. Workload forecasting and energy state estimation in cloud data centres: MLcentric approach. Future Generation Computer Systems, 128(??):320–332, March 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21004155.

Kuyu:2022:GNM

[KV22] Yiğit Çağatay Kuyu and Fahri Vatansever. GOZDE: a novel metaheuristic algorithm for global optimization. Future Generation Computer Systems, 136(??):128–152, November 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22001947.

Kumar:2020:EMC

[KVCY20] Neeraj Kumar, Athanasios V. Vasilakos, Kim-Kwang Raymond Choo, and Laurence T. Yang. Energy management for cyber-physical cloud systems. Future Generation Computer Systems, 105 (??):754–756, April 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167739X19313998.

Korch:2020:ILE

[KW20] Matthias Korch and Tim Werner. Improving locality of explicit onestep methods on GPUs by tiling across stages and time steps. Future Generation Computer Systems, 102(??):889–901, January 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19307186.

Kong:2023:DMI

[KWL^{+}23] Chang Kong, Haitao Wang, Qiuming Luo, Rui Mao, and Guoliang Chen. Deep Multi-Input Multi-Stream Ordinal Model for age estimation: Based on spatial attention learning. Future Generation Computer Systems, 140(??):173–184, March 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22003247.

Kong:2023:CFS

[KXZW23] Lulu Kong, Zichen Xu, Qiaoying Zhang, and Yuhao Wang. Carbon footprint and service coverage tradeoffs in geo-diverse sites. Future Generation Computer Systems, 143(??):1– 14, June 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23000110.

Kim:2020:ARM

[KYPJ20] Hyun-Woo Kim, Gangman Yi, Jong Hyuk Park, and Young-Sik Jeong. Adaptive resource management using many-core processing for fault tolerance based on cyberphysical cloud systems. Future Generation Computer Systems, 105(??):884–893, April 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17314747.

Kim:2020:OSB

[KYY^{+}20] Jongseong Kim, Unil Yun, Eunchul Yoon, Jerry Chun-Wei Lin, and Philippe Fournier-Viger. One scan based high average-utility pattern mining in static and dynamic databases. Future Generation Computer Systems, 111(??):143–158, October 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19331103.

Kermany:2023:IUR

[KZB^{+}23] Naime Ranjbar Kermany, Weiliang Zhao, Tseesuren Batsuuri, Jian Yang, and Jia Wu. Incorporating user rating credibility in recommender systems. Future Generation Computer Systems, 147(??):30–43, October 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2300170X.

Kuzilek:2021:SSP

[KZF21] Jakub Kuzilek, Zdenek Zdrahal, and Viktor Fuglik. Student success prediction using student exam behaviour. Future Generation Computer Systems, 125(??):661–671, December 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21002703.

Kong:2022:PSA

[KZG^{+}22] Ke Kong, Zhaoxin Zhang, Changyong Guo, Jideng Han, and Gang Long. PMMSA: Security analysis system for Android wearable applications based on permission matching and malware similarity analysis. Future Generation Computer Systems, 137(??):349–362, December 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22002631.

Lee:2022:GCG

[LAHN22] Munkyu Lee, Hyunho Ahn, Cheol-Ho Hong, and Dimitrios S. Nikolopoulos. gShare: a centralized GPU memory management framework to enable GPU memory sharing for containers. Future Generation Computer Systems, 130(??):181–192, May 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21004970.

Lemoine:2020:SAF

[LAS20] Frédéric Lemoine, Tatiana Aubonnet, and Noëmie Simoni. Self-assemble-featured Internet of Things. Future Generation Computer Systems, 112(??):41– 57, November 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20302843.

Lopes:2020:AOD

[LAT^{+}20] Fábio Lopes, João Agnelo, César A. Teixeira, Nuno Laranjeiro, and Jorge Bernardino. Automating orthogonal defect classification using machine learning algorithms. Future Generation Computer Systems, 102(??):932–947, January 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19308283.

Lira:2023:AIA

[LBDP23] Cleber Lira, Ernando Batista, Flávia C. Delicato, and Cássio Prazeres. Architecture for IoT applications based on reactive microservices: a performance evaluation. Future Generation Computer Systems, 145(??):223–238, August 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23001036.

Li:2020:HAE

[LBGL20] Chunlin Li, Jingpan Bai, Yuan Ge, and Youlong Luo. Heterogeneity-aware elastic provisioning in cloudassisted edge computing systems. Future Generation Computer Systems, 112(??):1106–1121, November 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20300339.

Leng:2018:RAS

[LBJ^{+}18] Kaijun Leng, Ya Bi, Linbo Jing, Han-Chi Fu, and Inneke Van Nieuwenhuyse. Research on agricultural supply chain system with double chain architecture based on blockchain technology. Future Generation Computer Systems, 86(??):641–649, September 2018. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18304527.

See retraction notice [LBJ^{+}24].

Leng:2024:RNR

[LBJ^{+}24] Kaijun Leng, Ya Bi, Linbo Jing, Han-Chi Fu, and Inneke Van Nieuwenhuyse. Retraction notice to “Research on agricultural supply chain system with double chain architecture based on blockchain technology” [Future Generation Computer Systems 86 (2018) 641–649]. Future Generation Computer Systems, 151(??):272, February 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23003758.

Liu:2020:ADL

[LBY^{+}20] Zhuo Liu, Changchuan Bai, Hang Yu, Ying Zhu, Taihua Wu, Fanyu Bu, and Qingchen Zhang. An adaptive deep learning model to differentiate syndromes of infectious fever in smart medicine. Future Generation Computer Systems, 111(??):853–858, October 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19321843.

Luo:2020:ETQ

[LC20] Xiaoyu Luo and Zhibin Chen. English text quality analysis based on recurrent neural network and semantic segmentation. Future Generation Computer Systems, 112(??):507–511, November 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20302764.

Lei:2020:BBC

[LCB^{+}20] Ao Lei, Yue Cao, Shihan Bao, Dasen Li, Philip Asuquo, Haitham Cruickshank, and Zhili Sun. A blockchain based certificate revocation scheme for vehicular communication systems. Future Generation Computer Systems, 110(??):892–903, September 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1831522X.

Lohman:2023:BDT

[LCB^{+}23] Walter Lohman, Hans Cornelissen, Jeroen Borst, Ralph Klerkx, Yashar Araghi, and Erwin Walraven. Building digital twins of cities using the Inter Model Broker framework. Future Generation Computer Systems, 148(??):501–513, November 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23002455.

Li:2024:EMM

[LCC^{+}24a] Ya-Lun Li, Yan-Yang Cheng, Zheng-Yi Chai, Xu Liu, Hao-Le Hou, and Guoqiang Chen. Evolutionary multitasking for multiobjective optimization based on hybrid differential evolution and multiple search strategy. Future Generation Computer Systems, 158(??):230–241, September 2024. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X24001626.

Luo:2024:CSA

[LCC^{+}24b] Jiang-Yao Luo, Liang Chen, Wei-Kun Chen, Jian-Hua Yuan, and Yu-Hong Dai. A cut-and-solve algorithm for virtual machine consolidation problem. Future Generation Computer Systems, 154(??):359–372, May 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24000104.

Lyu:2021:RPP

[LCCP21] Desheng Lyu, Ziwei Chen, Zesu Cai, and Songhao Piao. Robot path planning by leveraging the graph-encoded Floyd algorithm. Future Generation Computer Systems, 122(??):204–208, September 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2100087X.

Liu:2020:HTA

[LCFM20] Xinyao Liu, Baojiang Cui, Junsong Fu, and Jinxin Ma. HFuzz: Towards automatic fuzzing testing of NB–IoT core network protocols implementations. Future Generation Computer Systems, 108(??):390–400, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19324409.

Li:2021:OMR

[LCH^{+}21] Zhongjin Li, Victor Chang, Haiyang Hu, Maozhong Fu, Jidong Ge, and Francesco Piccialli. Optimizing makespan and resource utilization for multi-DNN training in GPU cluster. Future Generation Computer Systems, 125(??):206–220, December 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21002168.

Li:2022:CDC

[LCH^{+}22] Wentao Li, Zhiwen Chen, Xin He, Guoyun Duan, Jianhua Sun, and Hao Chen. CVFuzz: Detecting complexity vulnerabilities in OpenCL kernels via automated pathological input generation. Future Generation Computer Systems, 127(??):384–395, February 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21003526.

Li:2023:AIB

[LCH^{+}23] Beibei Li, Yujie Chang, Hanyuan Huang, Wenshan Li, Tao Li, and Wen Chen. Artificial immunity based distributed and fast anomaly detection for Industrial Internet of Things. Future Generation Computer Systems, 148(??):367–379, November 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23002327.

Luo:2024:PPC

[LCH^{+}24] Guixun Luo, Naiyue Chen, Jiahuan He, Bingwei Jin, Zhiyuan Zhang, and Yidong Li. Privacy-preserving clustering federated learning for non-IID data. Future Generation Computer Systems, 154(??):384–395, May 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24000050.

Li:2020:DAU

[LCL^{+}20] Xin Li, Shidan Cheng, Zhihan Lv, Houbing Song, Tao Jia, and Ning Lu. Data analytics of urban fabric metrics for smart cities. Future Generation Computer Systems, 107(??):871–882, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1731227X.

Li:2022:ODP

[LCL22] Chunlin Li, Qianqian Cai, and Youlong Lou. Optimal data placement strategy considering capacity limitation and load balancing in geographically distributed cloud. Future Generation Computer Systems, 127(??):142–159, February 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21003228.

Lv:2021:AIS

[LCLA21] Zhihan Lv, Dongliang Chen, Ranran Lou, and Ammar Alazab. Artificial intelligence for securing industrial-based cyber– physical systems. Future Generation Computer Systems, 117(??):291–298, April 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20330636.

Lv:2021:IEC

[LCLW21] Zhihan Lv, Dongliang Chen, Ranran Lou, and Qingjun Wang. Intelligent edge computing based on machine learning for smart city. Future Generation Computer Systems, 115(??):90–99, February 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20306889.

See retraction notice [LCLW24].

Lv:2024:RNI

[LCLW24] Zhihan Lv, Dongliang Chen, Ranran Lou, and Qingjun Wang. Retraction notice to “Intelligent Edge Computing Based on Machine Learning for Smart City” [Future Generation Computer Systems 115 (2020) 90–99]. Future Generation Computer Systems, 158(??):569, September 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24001869.

See [LCLW21].

Liu:2023:RCR

[LCO^{+}23] Hongyun Liu, Peng Chen, Xue Ouyang, Hui Gao, Bing Yan, Paola Grosso, and Zhiming Zhao. Robustness challenges in reinforcement learning based time-critical cloud resource scheduling: a meta-learning based solution. Future Generation Computer Systems, 146(??):18–33, September 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23001061.

Li:2020:SMA

[LCW^{+}20] Shimin Li, Huiling Chen, Mingjing Wang, Ali Asghar Heidari, and Seyedali Mirjalili. Slime mould algorithm: a new method for stochastic optimization. Future Generation Computer Systems, 111(??):300–323, October 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19320941.

Li:2023:SSC

[LCY^{+}23a] Shuzhe Li, Wei Chen, Bingqi Yan, Zhen Li, Shunzhi Zhu, and Yanwei Yu. Self-supervised contrastive representation learning for large-scale trajectories. Future Generation Computer Systems, 148(??):357–366, November 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23002376.

Liang:2023:DDG

[LCY^{+}23b] Yi Liang, Kaizhong Chen, Lan Yi, Xing Su, and Xiaoming Jin. DeGTeC: a deep graph-temporal clustering framework for dataparallel job characterization in data centers. Future Generation Computer Systems, 141(??):81–95, April 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22003752.

Liang:2021:TRR

[LCZB21] Haibo Liang, Haifeng Chen, Jialing Zou, and Jing Bai. Technical research on realizing remote intelligent diagnosis of petroleum drilling loss circulation under smart city strategy. Future Generation Computer Systems, 125(??):91– 99, December 2021. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21002120.

Liu:2020:PPB

[LDCZ20] Qiang Liu, XiaoShe Dong, Heng Chen, and Xingjun Zhang. H^{2} Pregel: a partition-based hybrid hierarchical graph computation approach. Future Generation Computer Systems, 104(??):15–31, March 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18306010.

Long:2022:ODP

[LDD^{+}22] Linbo Long, Jinpei Du, Xuxu Deng, Renping Liu, Yi Jiang, and Yan Wang. Optimizing data placement and size configuration for morphable NVM based SPM in embedded multicore systems. Future Generation Computer Systems, 135(??):270–282, October 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22001704.

Liang:2021:POQ

[LDDL21] Helan Liang, Bincheng Ding, Yanhua Du, and Fanzhang Li. Parallel optimization of QoS-aware big service processes with discovery of skyline services. Future Generation Computer Systems, 125(??):496–514, December 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21002430.

Liang:2020:CDS

[LDGS20] Helan Liang, Yanhua Du, Enting Gao, and Jinghan Sun. Cost-driven scheduling of service processes in hybrid cloud with VM deployment and intervalbased charging. Future Generation Computer Systems, 107(??):351–367, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19324288.

Li:2020:ICP

[LDLS20] Daming Li, Lianbing Deng, Wenjian Liu, and Qinglang Su. Improving communication precision of IoT through behaviorbased learning in smart city environment. Future Generation Computer Systems, 108(??):512–520, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19334569.

Li:2022:SDB

[LDLS22] Guozhi Li, Yifan Dong, Jirui Li, and Xuekun Song. Strategy for dynamic blockchain construction and transmission in novel edge computing networks. Future Generation Computer Systems, 130 (??):19–32, May 2022. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21004866.

Lisi:2021:RRT

[LDM^{+}21] Andrea Lisi, Andrea De Salve, Paolo Mori, Laura Ricci, and Samuel Fabrizi. Rewarding reviews with tokens: an Ethereumbased approach. Future Generation Computer Systems, 120(??):36–54, July 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21000480.

Liu:2021:IFP

[LDW^{+}21] Jian Liu, Yuanmin Duan, Yuedong Wu, Rui Chen, Liang Chen, and Geng Chen. Information flow perception modeling and optimization of Internet of Things for cloud services. Future Generation Computer Systems, 115(??):671–679, February 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20329897.

Liang:2020:MAR

[LDWZ20] Bin Liang, Xiaoshe Dong, Yufei Wang, and Xingjun Zhang. Memory-aware resource management algorithm for low-energy cloud data centers. Future Generation Computer Systems, 113(??):329–342, December 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20305835.

Liu:2023:ADI

[LDX^{+}23] Guozhi Liu, Fei Dai, Xiaolong Xu, Xiaodong Fu, Wanchun Dou, Neeraj Kumar, and Muhammad Bilal. An adaptive DNN inference acceleration framework with end–edge–cloud collaborative computing. Future Generation Computer Systems, 140(??): 422–435, March 2023. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22003570.

Lim:2024:EAI

[LEWC24] Kiho Lim, Christian Esposito, Tian Wang, and Chang Choi. Editorial: Artificial intelligence in biomedical big data and digital healthcare. Future Generation Computer Systems, 152(??):343–345, March 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23003928.

Luo:2020:SDE

[LEXH20] Jia Luo, Didier El Baz, Rui Xue, and Jinglu Hu. Solving the dynamic energy aware job shop scheduling problem with the heterogeneous parallel genetic algorithm. Future Generation Computer Systems, 108(??):119–134, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19314189.

Liu:2021:ERD

[LF21] Yishu Liu and Guifang Fu. Emotion recognition by deeply learned multichannel textual and EEG features. Future Generation Computer Systems, 119 (??):1–6, June 2021. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21000200.

Lu:2023:EET

[LFHS23] Baoshan Lu, Junli Fang, Xuemin Hong, and Jianghong Shi. Energy-efficient task scheduling for mobile edge computing with virtual machine I/O interference. Future Generation Computer Systems, 148(??):538–549, November 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23002431.

Li:2022:EMC

[LFM^{+}22] Tengyue Li, Simon Fong, Sabah Mohammed, Jinan Fiaidhi, Steven Guan, and Victor Chang. Empowering multi-class medical data classification by Groupof-Single-Class-predictors and transfer optimization: Cases of structured dataset by machine learning and radiological images by deep learning. Future Generation Computer Systems, 133(??):10–22, August 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22000711.

Liang:2022:SAO

[LFYH22] Liang Liang, Rosa Filgueira, Yan Yan, and Thomas Heinis. Scalable adaptive optimizations for stream-based workflows in multi-HPC-clusters and cloud infrastructures. Future Generation Computer Systems, 128(??):102–116, March 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21003824.

Li:2021:THR

[LFZJ21] Ying Li, Binbin Fan, Weiping Zhang, and Zhiqiang Jiang. TireNet: a high recall rate method for practical application of tire defect type classification. Future Generation Computer Systems, 125(??):1–9, December 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21002041.

Lin:2021:ASG

[LGC^{+}21] Gan Lin, Zhihua Guo, Fei Chao, Longzhi Yang, Xiang Chang, Chih-Min Lin, Changle Zhou, V. Vijayakumar, and Changjing Shang. Automatic stroke generation for style-oriented robotic Chinese calligraphy. Future Generation Computer Systems, 119(??):20– 30, June 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X2100039X.

Liao:2022:BBI

[LGCY22] Chia-Hung Liao, Xue-Qin Guan, Jen-Hao Cheng, and Shyan-Ming Yuan. Blockchain-based identity management and access control framework for open banking ecosystem. Future Generation Computer Systems, 135(??):450–466, October 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22001868.

Liu:2023:SLD

[LGJ^{+}23] Yu Liu, Yunchuan Guan, Tianming Jiang, Ke Zhou, Hua Wang, Guangxing Hu, Ji Zhang, Wei Fang, Zhuo Cheng, and Ping Huang. SPAE: Lifelong disk failure prediction via end-to-end GAN-based anomaly detection with ensemble update. Future Generation Computer Systems, 148(??):460–471, November 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23002030.

Lohachab:2021:PEH

[LGKA21] Ankur Lohachab, Saurabh Garg, Byeong Ho Kang, and Muhammad Bilal Amin. Performance evaluation of hyperledger fabricenabled framework for pervasive peer-to-peer energy trading in smart cyber– physical systems. Future Generation Computer Systems, 118(??):392–416, May 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21000339.

Liu:2020:ESG

[LGL^{+}20a] Yongnan Liu, Xin Guan, Jun Li, Di Sun, Tomoaki Ohtsuki, Mohammad Mehedi Hassan, and Abdulhameed Alelaiwi. Evaluating smart grid renewable energy accommodation capability with uncertain generation using deep reinforcement learning. Future Generation Computer Systems, 110(??):647–657, September 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19323039.

Lu:2020:OLT

[LGL^{+}20b] Haifeng Lu, Chunhua Gu, Fei Luo, Weichao Ding, and Xinping Liu. Optimization of lightweight task offloading strategy for mobile edge computing based on deep reinforcement learning. Future Generation Computer Systems, 102(??):847–861, January 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19308209.

Lin:2023:AAC

[LGL^{+}23] Rongping Lin, Xuhui Guo, Shan Luo, Yong Xiao, Bill Moran, and Moshe Zukerman. Application-aware computation offloading in edge computing networks. Future Generation Computer Systems, 146(??):86– 97, September 2023. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23001437.

Losada:2020:FTM

[LGM^{+}20] Nuria Losada, Patricia González, María J. Martín, George Bosilca, Aurélien Bouteiller, and Keita Teranishi. Fault tolerance of MPI applications in exascale systems: the ULFM solution. Future Generation Computer Systems, 106(??): 467–481, May 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1930860X.

Loff:2021:NPB

[LGM^{+}21] Júnior Löff, Dalvan Griebler, Gabriele Mencagli, Gabriell Araujo, Massimo Torquati, Marco Danelutto, and Luiz Gustavo Fernandes. The NAS Parallel Benchmarks for evaluating C++ parallel programming frameworks on shared-memory architectures. Future Generation Computer Systems, 125(??):743–757, December 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21002831.

Liu:2023:IDP

[LGS^{+}23] Zhuojin Liu, Shufeng Gong, Yuxuan Su, Changyi Wan, Yanfeng Zhang, and Ge Yu. Improving density peaks clustering through GPU acceleration. Future Generation Computer Systems, 141(??):399–413, April 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167739X22004046.

Luo:2020:HTI

[LGT^{+}20] Entao Luo, Kehua Guo, Yayuan Tang, Xiangdong Ying, and Wen Huang. Hidden the true identity and dating characteristics based on quick private matching in mobile social networks. Future Generation Computer Systems, 109(??): 633–641, August 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17320149.

Liu:2021:ESP

[LGW^{+}21] Jiayu Liu, Huaxi Gu, Wenting Wei, Ziqi Chen, and Yawen Chen. An efficient shortest path algorithm for content-based routing on 2-D mesh accelerator networks. Future Generation Computer Systems, 114(??):519–530, January 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19332169.

Liu:2022:TPM

[LGW22] Jianli Liu, Bei Gong, and Qian Wang. A trusted proof mechanism of data source for smart city. Future Generation Computer Systems, 128(??):349–364, March 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21004027.

Liu:2020:APG

[LGYC20] Wei Liu, Jingzhi Guo, Feng Yao, and Deng Chen. Adaptive protocol generation for group collaborative in smart medical waste transportation. Future Generation Computer Systems, 110(??):167–180, September 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19302031.

Li:2024:TOM

[LGZ^{+}24] Jianjiang Li, Hongyaoxing Gu, Jing Zhao, Lin Qiao, Chunye Gong, and Gang Zheng. Transplantation and optimization of molecular dynamics simulation on MT-3000. Future Generation Computer Systems, 153(??):262–275, April 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23004521.

Lu:2020:GGB

[LH20] Yi-Shu Lu and Jiun-Long Huang. GLR: a graphbased latent representation model for successive POI recommendation. Future Generation Computer Systems, 102(??):230–244, January 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19303966.

Liu:2021:FSI

[LH21] Sai Liu and Wenqi Hao. Forecasting the scheduling issues in engineering project management: Applications of deep learning models. Future Generation Computer Systems, 123(??):85–93, October 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21001345.

Levinger:2020:HSU

[LHA20] Chaya Levinger, Noam Hazon, and Amos Azaria. Human satisfaction as the ultimate goal in ridesharing. Future Generation Computer Systems, 112(??):176–184, November 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19302249.

Lai:2020:QAU

[LHC^{+}20] Phu Lai, Qiang He, Guangming Cui, Xiaoyu Xia, Mohamed Abdelrazek, Feifei Chen, John Hosking, John Grundy, and Yun Yang. QoE-aware user allocation in edge computing systems with dynamic QoS. Future Generation Computer Systems, 112(??):684–694, November 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19330961.

Lee:2021:MBM

[LHC21] Gyeong Ho Lee, Jaeseob Han, and Jun Kyun Choi. MPdist-based missing data imputation for supporting big data analyses in IoTbased applications. Future Generation Computer Systems, 125(??):421–432, December 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21002375.

Li:2020:MTR

[LHD^{+}20] Yujie Li, Gang Hu, Jinyang Du, Haider Abbas, and Yin Zhang. Multi-task reading for intelligent legal services. Future Generation Computer Systems, 113(??):218–227, December 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20307780.

Li:2020:MRT

[LHF^{+}20] Qingjuan Li, Wei Huangfu, Fadi Farha, Tao Zhu, Shunkun Yang, Liming Chen, and Huansheng Ning. Multi-resident type recognition based on ambient sensors activity. Future Generation Computer Systems, 112(??):108–115, November 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19325361.

Li:2021:SEA

[LHH^{+}21] Zhongjin Li, Haiyang Hu, Hua Hu, Binbin Huang, Jidong Ge, and Victor Chang. Security and energy-aware collaborative task offloading in D2D communication. Future Generation Computer Systems, 118(??):358–373, May 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21000315.

Li:2020:IDL

[LHL20] Defeng Li, Yuan Hu, and Mingming Lan. IoT device location information storage system based on blockchain. Future Generation Computer Systems, 109(??):95–102, August 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20300169.

Li:2023:CSA

[LHLC23] Zhen Li, Xiang Huang, Yangrui Li, and Guenevere Chen. A comparative study of adversarial training methods for neural models of source code. Future Generation Computer Systems, 142(??):165–181, May 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22004332.

Liz:2023:DLU

[LHTSM^{+}23] Helena Liz, Javier Huertas-Tato, Manuel Sánchez-Montañés, Javier Del Ser, and David Camacho. Deep learning for understanding multilabel imbalanced chest X-ray datasets. Future Generation Computer Systems, 144(??):291–306, July 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23000821.

Lu:2020:HEA

[LHW20] Zhi Lu, Jin-Kao Hao, and Qinghua Wu. A hybrid evolutionary algorithm for finding low conductance of large graphs. Future Generation Computer Systems, 106(??):105–120, May 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19320540.

Li:2023:NSM

[LHW^{+}23] Jiatao Li, Dezhi Han, Zhongdai Wu, Junxiang Wang, Kuan-Ching Li, and Arcangelo Castiglione. A novel system for medical equipment supply chain traceability based on alliance chain and attribute and role access control. Future Generation Computer Systems, 142(??):195–211, May 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22004411.

Li:2022:QBB

[LHXL22] Siyuan Li, Guangji Huang, Xing Xu, and Huimin Lu. Query-based black-box attack against medical image segmentation model. Future Generation Computer Systems, 133(??):331–337, August 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2200084X.

Li:2020:FSC

[LHY^{+}20a] Daifeng Li, Lu Huang, Biyun Ye, Fangbin Wan, Andrew Madden, and Xingjian Liang. FSRM-STS: Cross-dataset pedestrian retrieval based on a four-stage retrieval model with selection–translation– selection. Future Generation Computer Systems, 107(??):601–619, June 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167739X19300433.

Lu:2020:EET

[LHY^{+}20b] Feng Lu, Jingru Hu, Laurence Tianruo Yang, Zaiyang Tang, Peng Li, Ziqian Shi, and Hai Jin. Energy-efficient traffic offloading for mobile users in two-tier heterogeneous wireless networks. Future Generation Computer Systems, 105(??):855–863, April 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17317740.

Li:2020:ROS

[Li20] Jing Li. Resource optimization scheduling and allocation for hierarchical distributed cloud service system in smart city. Future Generation Computer Systems, 107(??):247–256, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1932028X.

Loconte:2024:ECE

[LIP^{+}24] Davide Loconte, Saverio Ieva, Agnese Pinto, Giuseppe Loseto, Floriano Scioscia, and Michele Ruta. Expanding the cloud-to-edge continuum to the IoT in serverless federated learning. Future Generation Computer Systems, 155(??):447–462, June 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24000670.

Liu:2021:ODT

[Liu21] Long Liu. Objects detection toward complicated high remote basketball sports by leveraging deep CNN architecture. Future Generation Computer Systems, 119(??):31– 36, June 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21000303.

Liu:2023:QLB

[Liu23]  Guanghua Liu.  A Q-

Learning-based distributed routing protocol for frequencyswitchable magnetic inductionbased wireless underground sensor networks. Future Generation Computer Systems, 139(??):253–266, February 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22003193.

Liu:2024:HBN

[Liu24] Changping Liu. HPCLS-BC: a novel blockchain framework using heterogeneous peer-node and cloudbased ledger storage for Internet of Things applications. Future Generation Computer Systems, 150(??):364–379, January 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23003473.

Liu:2021:AAW

[LJ21] Xin Liu and Yujia Jiang. Aesthetic assessment of website design based on multimodal fusion. Future Generation Computer Systems, 117(??):433–438, April 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20330764.

Liu:2024:AAP

[LJ24] Shuai Liu and Tao Ju. APapo: an asynchronous parallel optimization method for DNN models. Future Generation Computer Systems, 152(??):317–330, March 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23004041.

Li:2020:SSB

[LJC^{+}20] Xiaoqi Li, Peng Jiang, Ting Chen, Xiapu Luo, and Qiaoyan Wen. A survey on the security of blockchain systems. Future Generation Computer Systems, 107(??):841–853, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17318332.

Li:2021:FSA

[LJL^{+}21] Jianhua Li, Jiong Jin, Lingjuan Lyu, Dong Yuan, Yingying Yang, Longxiang Gao, and Chao Shen. A fast and scalable authentication scheme in IOT for smart living. Future Generation Computer Systems, 117(??):125–137, April 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20330302.

Liu:2020:PRN

[LJW^{+}20] Ling Liu, Qixuan Jin, Dan Wang, Hongfang Yu, Gang Sun, and Shouxi Luo. PSNet: Reconfigurable network topology design for accelerating parameter server architecture based distributed machine learning. Future Generation Computer Systems, 106(??):320–332, May 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19315614.

Lahlou:2022:DOD

[LKE22] Laaziz Lahlou, Nadjia Kara, and Claes Edstrom. DAVINCI: online and Dynamic Adaptation of eVolvable vIrtual Network services over Cloud Infrastructures. Future Generation Computer Systems, 127(??):396–408, February 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21003630.

Le-Khac:2020:SVF

[LKJN^{+}20] Nhien-An Le-Khac, Daniel Jacobs, John Nijhoff, Karsten Bertens, and Kim-Kwang Raymond Choo. Smart vehicle forensics: Challenges and case study. Future Generation Computer Systems, 109(??):500–510, August 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17322422.

Lesch:2021:CMC

[LKS^{+}21] Veronika Lesch, Christian Krupitzer, Kevin Stubenrauch, Nico Keil, Christian Becker, Samuel Kounev, and Michele Segata. A comparison of mechanisms for compensating negative impacts of system integration. Future Generation Computer Systems, 116(??): 117–131, March 2021. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20330132.

Luo:2020:PGB

[LL20] Le Luo and Yi Liu. Processing graphs with barrierless asynchronous parallel model on sharedmemory systems. Future Generation Computer Systems, 106(??):641–652, May 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19322198.

Laso:2022:CNL

[LLC^{+}22] Ruben Laso, Oscar G.

Lorenzo, José C. Cabaleiro, Tomás F. Pena, Juan Ángel Lorenzo, and Francisco F. Rivera. CIMAR, NIMAR, and LMMA: Novel algorithms for thread and memory migrations in user space on NUMA systems using hardware counters. Future Generation Computer Systems, 129(??):18– 32, April 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21004374.

Li:2023:MKR

[LLC^{+}23] YongGang Li, GuoYuan Lin, Yeh-Ching Chung, YaoWen Ma, Yi Lu, and Yu Bao. MagBox: Keep the risk functions running safely in a magic box. Future Generation Computer Systems, 140(??):282–298, March 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22003582.

Li:2021:LBB

[LLCH21] Meng Li, Chhagan Lal, Mauro Conti, and Donghui Hu. LEChain: a blockchainbased lawful evidence management scheme for digital forensics. Future Generation Computer Systems, 115(??):406–420, February 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1933167X.

Liu:2021:LNA

[LLD^{+}21] Xu Liu, Weiyou Liu, Xiaoqiang Di, Jinqing Li, Binbin Cai, Weiwu Ren, and Huamin Yang. LogNADS: Network anomaly detection scheme based on log semantics representation. Future Generation Computer Systems, 124(??):390–405, November 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21001746.

Liu:2023:MEA

[LLF^{+}23] Chuanchang Liu, Jianyun Lu, Wendi Feng, Enbo Du, Luyang Di, and Zhen Song. MobiPCR: Efficient, accurate, and strict ML-based mobile malware detection. Future Generation Computer Systems, 144(??): 140–150, July 2023. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23000559.

Li:2021:BIT

[LLFQ21] Xiaoqing Li, Yu Lu, Xianghua Fu, and Yingjian Qi. Building the Internet of Things platform for smart maternal healthcare services with wearable devices and cloud computing. Future Generation Computer Systems, 118(??):282–296, May 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21000261.

Liu:2020:CAC

[LLG^{+}20] YuXin Liu, Anfeng Liu, Shuang Guo, Zhetao Li, Young-June Choi, and Hiroo Sekiya. Contextaware collect data with energy efficient in cyberphysical cloud systems. Future Generation Computer Systems, 105(??):932–947, April 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17310518.

Liu:2024:EDO

[LLLS24] Huan Liu, Shiyong Li, Wenzhe Li, and Wei Sun. Efficient decentralized optimization for edge-enabled smart manufacturing: a federated learning-based framework. Future Generation Computer Systems, 157(??):422–435, August 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24001146.

Li:2020:MFT

[LLP^{+}20a] Hui Li, Zhe Li, Sizhe Peng, Jingjing Li, and Chia Emmanuel Tungom. Mining the frequency of timeconstrained serial episodes over massive data sequences and streams. Future Generation Computer Systems, 110(??):849–863, September 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18332138.

Li:2020:MBR

[LLP^{+}20b] Jianjiang Li, Yajun Liu, Jian Pan, Peng Zhang, Wei Chen, and Lizhe Wang. Map-Balance-Reduce: an improved parallel programming model for load balancing of MapReduce. Future Generation Computer Systems, 105(??):993–1001, April 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17303710.

Liang:2024:DAK

[LLS24] Yihuai Liang, Yan Li, and Byeong-Seok Shin. Dynamic authenticated keyword search in hybridstorage blockchain. Future Generation Computer Systems, 155(??):53–65, June 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X24000323.

Loreti:2020:PML

[LLT20] Daniela Loreti, Marco Lippi, and Paolo Torroni. Parallelizing machine learning as a service for the end-user. Future Generation Computer Systems, 105(??):275–286, April 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19316346.

Liu:2022:DAF

[LLT22] Yingbo Liu, Duc Van Le, and Rui Tan. A dataassisted first-principle approach to modeling server outlet temperature in air free-cooled data centers. Future Generation Computer Systems, 129(??): 225–235, April 2022. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21004799.

Li:2020:DSS

[LLW^{+}20] Mengshan Li, Suyun Lian, Fan Wang, Yanying Zhou, Bingsheng Chen, Lixin Guan, and Yan Wu. A decision support system using hybrid AI based on multiimage quality model and its application in color design. Future Generation Computer Systems, 113(??):70– 77, December 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20308360.

Li:2022:BDA

[LLW^{+}22a] Xiaoming Li, Hao Liu, Weixi Wang, Ye Zheng, Haibin Lv, and Zhihan Lv. Big data analysis of the Internet of Things in the digital twins of smart city based on deep learning. Future Generation Computer Systems, 128(??):167–177, March 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21003964.

Liu:2022:MIF

[LLW^{+}22b] Yongqiang Liu, Bing Li, Jian Wang, Duantengchuan Li, and Yutao Ma. Multiinformation fusion based few-shot Web service classification. Future Generation Computer Systems, 130(??):231–240, May 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0167739X2100501X.

Li:2023:UIF

[LLW^{+}23a] Dongfen Li, Jinshan Lai, Ruijin Wang, Xiong Li, Pandi Vijayakumar, Brij B. Gupta, and Wadee Alhalabi. Ubiquitous intelligent federated learning privacy-preserving scheme under edge computing. Future Generation Computer Systems, 144(??):205–218, July 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23000869.

Li:2023:RSF

[LLW^{+}23b] Hao Li, Chengcheng Li, Jian Wang, Aimin Yang, Zezhong Ma, Zunqian Zhang, and Dianbo Hua. Review on security of federated learning and its application in healthcare. Future Generation Computer Systems, 144(??):271–290, July 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23000626.

Li:2024:PSI

[LLWJ24] Weimin Li, Lu Liu, Kevin I. K. Wang, and Qun Jin. Preface of special issue on heterogeneous information network embedding and applications. Future Generation Computer Systems, 152(??):331–332, March 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2300393X.

Liu:2020:PPM

[LLY^{+}20] Meng Liu, Yun Luo, Chi Yang, Shaoning Pang, Deepak Puthal, Kaijun Ren, and Xuyun Zhang. Privacy-preserving matrix product based static mutual exclusive roles constraints violation detection in interoperable role-based access control. Future Generation Computer Systems, 109(??):457–468, August 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17320551.

Liu:2020:NAR

[LLZ20] Yan Liu, Xiner Li, and Zaimei Zhang. A new approach in reject inference of using ensemble learning based on global semisupervised framework. Future Generation Computer Systems, 109(??):382–391, August 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19323635.

Liang:2021:RCM

[LLZ^{+}21] Haibo Liang, Gang Liu, Jialing Zou, Jing Bai, and Yingjun Jiang. Research on calculation model of bottom of the well pressure based on machine learning. Future Generation Computer Systems, 124(??):80– 90, November 2021. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21001618.

Liu:2022:HQH

[LLZ^{+}22] Yang Liu, Chaoqun Li, Yao Zhang, Mengying Xu, Jing Xiao, and Jie Zhou. HPCP-QCWOA: High performance clustering protocol based on quantum clone whale optimization algorithm in integrated energy system. Future Generation Computer Systems, 135(??):315–332, October 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22001662.

Liao:2024:PRH

[LLZ^{+}24a] Zhuhua Liao, Shoubin Li, Yijiang Zhao, Yizhi Liu, Wei Liang, and Shaohua Wan. Predicting ride-hailing passenger demand: a POI-based adaptive clustering federated learning approach. Future Generation Computer Systems, 156(??):168–178, July 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24000712.

Liu:2024:BTR

[LLZ^{+}24b] Yanhua Liu, Zhihuang Liu, Qiu Zhang, Jinshu Su, Zhiping Cai, and Xiaoyan Li. Blockchain and trusted reputation assessment-based incentive mechanism for healthcare services. Future Generation Computer Systems, 154 (??):59–71, May 2024. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23004843.

Li:2021:ECC

[LLZL21] Chunlin Li, Jun Liu, Qingchuan Zhang, and Youlong Luo. Efficient cooperative cache management for latency-aware data intelligent processing in edge environment. Future Generation Computer Systems, 123(??):48–67, October 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21001333.

Lee:2020:BBM

[LM20] Han Lee and Maode Ma. Blockchain-based mobility management for 5G. Future Generation Computer Systems, 110(??):638–646, September 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19306843.

Lopez-Martin:2020:ITT

[LMCSE20] Manuel Lopez-Martin, Belen Carro, and Antonio Sanchez-Esguevillas. IoT type-of-traffic forecasting method based on gradient boosting neural networks. Future Generation Computer Systems, 105(??): 331–345, April 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19322319.

Lee:2022:DBU

[LMNC22] Yunseong Lee, Arooj Masood, Wonjong Noh, and Sungrae Cho. DQN based user association control in hierarchical mobile edge computing systems for mobile IoT services. Future Generation Computer Systems, 137(??):53–69, December 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22002333.

Lamaazi:2022:SDD

[LMO^{+}22] Hanane Lamaazi, Rabeb Mizouni, Hadi Otrok, Shakti Singh, and Ernesto Damiani. Smart-3DM: Data-driven decision making using smart edge computing in hetero-crowdsensing environment. Future Generation Computer Systems, 131(??):151–165, June 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167739X2200022X.

Liu:2024:BAV

[LMW^{+}24] Hang Liu, Yang Ming, Chenhao Wang, Yi Zhao, Songnian Zhang, and Rongxing Lu. Blockchain-assisted verifiable certificate-based searchable encryption against untrusted cloud server for Industrial Internet of Things. Future Generation Computer Systems, 153(??):97–112, April 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23004260.

Liu:2022:HRR

[LMZ^{+}22] Zhaoshan Liu, Yiming Ma, Huihua Zheng, Dege Liu, and Jing Liu. Human resource recommendation algorithm based on improved frequent itemset mining. Future Generation Computer Systems, 126(??):284–288, January 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21003253.

Liu:2024:TGO

[LMZL24] Jiagang Liu, Yun Mi, Xinyu Zhang, and Xiaocui Li. Task graph offloading via deep reinforcement learning in mobile edge computing. Future Generation Computer Systems, 158(??):545–555, September 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X24001638.

Lee:2023:PGP

[LOH^{+}23] Dongwon Lee, Yongwoo Oh, Jin B. Hong, Hyoungshick Kim, and Dan Dongseong Kim. PP–GSM: Privacypreserving graphical security model for security assessment as a service. Future Generation Computer Systems, 142(??):351–363, May 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22004459.

Leotta:2022:LEA

[LOR22] Maurizio Leotta, Dario Olianas, and Filippo Ricca. A large experimentation to analyze the effects of implementation bugs in machine learning algorithms. Future Generation Computer Systems, 133(??):184–200, August 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22000796.

Li:2021:CDR

[LP21a] Sumin Li and Xiuqin Pan.

A computational drug repositioning model based on hybrid similarity side information powered graph neural network. Future Generation Computer Systems, 125(??):24–31, December 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21002132.

Liu:2021:HCS

[LP21b] Chen Liu and Songwen Pei. Heterogeneous computation in specific domain accelerations. Future Generation Computer Systems, 124(??):10–11, November 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21001679.

Luciani:2023:SDC

[LP23] Alessio Luciani and Emanuele Panizzi. Street direction classification using implicit vehicle crowdsensing and deep learning. Future Generation Computer Systems, 149(??):59–70, December 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23002650.

Lucas:2020:TAF

[LPL^{+}20] Yvan Lucas, Pierre-Edouard Portier, Léa Laporte, Liyun He-Guelton, Olivier Caelen, Michael Granitzer, and Sylvie Calabretto. Towards automated feature engineering for credit card fraud detection using multiperspective HMMs. Future Generation Computer Systems, 102(??):393–402, January 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19300664.

Liu:2022:ECE

[LPL22] Mingyu Liu, Li Pan, and Shijun Liu. Effeclouds: a cost-effective cloud-of-clouds framework for two-tier storage. Future Generation Computer Systems, 129(??):33–49, April 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21004416.

Li:2024:PAP

[LPS^{+}24] Xiulin Li, Li Pan, Wei Song, Shijun Liu, and Xiangxu Meng. Performance analysis of parallel composite service-based applications in clouds. Future Generation Computer Systems, 153(??):27–40, April 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23004302.

Livaja:2022:DGP

[LPSV22] Ivan Livaja, Krešimir Pripužić, Siniša Sovilj, and Marin Vuković. A distributed geospatial publish/subscribe system on Apache Spark. Future Generation Computer Systems, 132(??):282–298, July 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22000577.

Lombardo:2022:CRL

[LPT22] Gianfranco Lombardo, Agostino Poggi, and Michele Tomaiuolo. Continual representation learning for node classification in power-law graphs. Future Generation Computer Systems, 128(??): 420–428, March 2022. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21004015.

Lv:2020:AHB

[LQ20] Zhihan Lv and Liang Qiao. Analysis of healthcare big data. Future Generation Computer Systems, 109(??):103–110, August 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20304829.

Li:2023:CSI

[LQG^{+}23] Meizi Li, Weiqiao Que, Ziyao Geng, Maozhen Li, Zuliang Kou, Jisheng Chen, Chang Guo, and Bo Zhang. Cold-start item recommendation for representation learning based on heterogeneous information networks with fusion side information. Future Generation Computer Systems, 149(??):227–239, December 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23002534.

Li:2022:PSL

[LQML22] Jiahui Li, Xiaogang Qi, Wenchao Ma, and Lifang Liu. Path selection for link failure protection in hybrid SDNs. Future Generation Computer Systems, 137(??):201–215, December 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22002448.

Liu:2020:ESB

[LQNW20] Hua Liu, Cunquan Qu, Yawei Niu, and Guanghui Wang. The evolution of structural balance in time-varying signed networks. Future Generation Computer Systems, 102(??):403–408, January 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18330036.

Li:2020:LAS

[LQS^{+}20] Shengnan Li, Zheng Qin, Houbing Song, Chengxiang Si, Bo Sun, Xiao Yang, and Renwei Zhang. A lightweight and aggregated system for indoor/outdoor detection using smart devices. Future Generation Computer Systems, 107(??):988–997, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1730506X.

Li:2020:DPS

[LQW^{+}20] Jingwei Li, Yong Qi, Wei Wei, Jinwei Lin, Marcin Wozniak, and Robertas Damasevicius. dCCPIpredictor: a state-aware approach for effectively predicting cross-core performance interference. Future Generation Computer Systems, 105(??):184–195, April 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19312397.

Liao:2021:IDS

[LQYL21] Yangzhe Liao, Xinhui Qiao, Quan Yu, and Quan Liu. Intelligent dynamic service pricing strategy for multiuser vehicle-aided MEC networks. Future Generation Computer Systems, 114(??):15–22, January 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20301151.

Li:2024:EPV

[LRCL24] Ning Li, Xiaojun Ren, Aniello Castiglione, and Mengyun Liu. Efficient and precise visual location estimation by effective priority matching-based pose verification in edge–cloud collaborative IoT. Future Generation Computer Systems, 155(??):96–107, June 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24000402.

Liu:2021:LBP

[LRML21] Shanshan Liu, Pedro Reviriego, Paolo Montuschi, and Fabrizio Lombardi. Less-is-better protection (LBP) for memory errors in k NNs classifiers. Future Generation Computer Systems, 117(??):401–411, April 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20330776.

Leroux:2023:SRN

[LS23a] Sam Leroux and Pieter Simoens. Sparse random neural networks for online anomaly detection on sensor nodes. Future Generation Computer Systems, 144(??):327–343, July 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22004344.

Li:2023:ESW

[LS23b] Feng Li and Fengguang Song. Efficient in-situ workflow planning for geographically distributed heterogeneous environments. Future Generation Computer Systems, 149(??):105–121, December 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23002601.

Liu:2021:DLD

[LSB21] Jun Liu, Bo Sun, and Jingpan Bai. Deeply learning a discriminative spatial– temporal feature for robot action understanding. Future Generation Computer Systems, 120(??):55–60, July 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21000698.

Lu:2020:MLA

[LSGA20] Jiayi Lu, Enmin Song, Ahmed Ghoneim, and Mubarak Alrashoud. Machine learning for assisting cervical cancer diagnosis: an ensemble approach. Future Generation Computer Systems, 106(??):199–205, May 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19330092.

Li:2020:PEB

[LSH^{+}20] Shijian Li, Minhao Shi, Runhe Huang, Xinwei Chen, and Gang Pan. Perception-enhancement based task learning and action scheduling for robotic limb in CPS environment. Future Generation Computer Systems, 108(??): 1069–1083, July 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17311676.

Liu:2020:CST

[LSL^{+}20] Sufang Liu, Lei Shi, Simian Liu, Chuan Wang, and Chulin Chen. City space type recognition by leveraging deeply-learned topological features. Future Generation Computer Systems, 112(??):624–629, November 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20302752.

Liz:2021:ECN

[LSMT^{+}21] Helena Liz, Manuel Sánchez-Montañés, Alfredo Tagarro, Sara Domínguez-Rodríguez, Ron Dagan, and David Camacho. Ensembles of convolutional neural network models for pediatric pneumonia diagnosis. Future Generation Computer Systems, 122(??):220–233, September 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2100128X.

Lei:2020:NDD

[LSN^{+}20] Zhenfeng Lei, Yuan Sun, Y. A. Nanehkaran, Shuangyuan Yang, Md. Saiful Islam, Huiqing Lei, and Defu Zhang. A novel data-driven robust framework based on machine learning and knowledge graph for disease classification. Future Generation Computer Systems, 102(??):534–548, January 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19306235.

Lin:2022:IGM

[LSS^{+}22] Jionghao Lin, Shaveen Singh, Lele Sha, Wei Tan, David Lang, Dragan Gašević, and Guanliang Chen. Is it a good move? mining effective tutoring strategies from human–human tutorial dialogues. Future Generation Computer Systems, 127(??):194–207, February 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21003393.

Liu:2022:BEF

[LTB^{+}22] Lin Liu, Wei-Tek Tsai, Md.

Zakirul Alam Bhuiyan, Hao Peng, and Mingsheng Liu. Blockchain-enabled fraud discovery through abnormal smart contract detection on Ethereum. Future Generation Computer Systems, 128(??):158–166, March 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21003319.

Li:2024:CEE

[LTX^{+}24] Huifang Li, Luzhi Tian, Guanghao Xu, Julio Ruben Cañizares Abreu, Shuangxi Huang, Senchun Chai, and Yuanqing Xia. Co-evolutionary and elite learning-based biobjective poor and rich optimization algorithm for scheduling multiple workflows in the cloud. Future Generation Computer Systems, 152(??):99–111, March 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23003898.

Liu:2022:DNN

[LTXL22] Siping Liu, Xiaohan Tu, Cheng Xu, and Renfa Li. Deep neural networks with attention mechanism for monocular depth estimation on embedded devices. Future Generation Computer Systems, 131(??): 137–150, June 2022. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22000255.

Loreti:2024:PAD

[LV24] Daniela Loreti and Giorgio Visani. Parallel approaches for a decision tree-based explainability algorithm. Future Generation Computer Systems, 158(??):308–322, September 2024. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24001857.

Lopez-Villellas:2024:GGB

[LVLBB^{+}24] Lorién López-Villellas, Rubén Langarita-Benítez, Asaf Badouh, Víctor Soria-Pardos, Quim Aguado-Puig, Guillem López-Paradís, Max Doblas, Javier Setoain, Chulho Kim, Makoto Ono, Adrià Armejach, Santiago Marco-Sola, Jesús Alastruey-Benedé, Pablo Ibáñez, and Miquel Moretó. GenArch-Bench: a genomics benchmark suite for ARM HPC processors. Future Generation Computer Systems, 157(??):313–329, August 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24001250.

Lopez-Vizcaino:2021:EDC

[LVNCC21] Manuel F. López-Vizcaíno, Francisco J. Nóvoa, Victor Carneiro, and Fidel Cacheda. Early detection of cyberbullying on social media networks. Future Generation Computer Systems, 118(??):219–229, May 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21000157.

Li:2023:RGP

[LWCC23] Ke Li, Hongyu Wang, Ziwen Chen, and Lisi Chen. Relaxed group pattern detection over massivescale trajectories. Future Generation Computer Systems, 144(??):131–139, July 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23000699.

Liu:2023:HDD

[LWF^{+}23] Ji Liu, Zhihua Wu, Danlei Feng, Minxu Zhang, Xinxuan Wu, Xuefeng Yao, Dianhai Yu, Yanjun Ma, Feng Zhao, and Dejing Dou. HeterPS: Distributed deep learning with reinforcement learning based scheduling in heterogeneous environments. Future Generation Computer Systems, 148(??):106–117, November 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23002157.

Lv:2022:MTP

[LWH^{+}22] Fang Lv, Wei Wang, Linxuan Han, Di Wang, Yulong Pei, Junheng Huang, Bailing Wang, and Mykola Pechenizkiy. Mining trading patterns of pyramid schemes from financial time series data. Future Generation Computer Systems, 134(??):388–398, September 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22000607.

Liu:2022:PSI

[LWHW22] Qin Liu, Guojun Wang, Jiankun Hu, and Jie Wu. Preface of special issue on artificial intelligence: the security & privacy opportunities and challenges for emerging applications. Future Generation Computer Systems, 133(??):169–170, August 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22000966.

Lee:2024:PID

[LWHW24] Alan Y.-P. Lee, Michael I.- C. Wang, Chi-Hsiang Hung, and Charles H.-P. Wen. PS-IPS: Deploying Intrusion Prevention System with machine learning on programmable switch. Future Generation Computer Systems, 152(??):333–342, March 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23004193.

Liu:2021:RUF

[LWJ^{+}21] Zhen Liu, Ruoyu Wang, Nathalie Japkowicz, Deyu Tang, Wenbin Zhang, and Jie Zhao. Research on unsupervised feature learning for Android malware detection based on Restricted Boltzmann Machines. Future Generation Computer Systems, 120(??):91–108, July 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21000674.

Li:2023:HAA

[LWJ^{+}23] Fei Li, Yuzhu Wang, Jinrong Jiang, He Zhang, Xiaocong Wang, and Xuebin Chi. Heterogeneous acceleration algorithms for shallow cumulus convection scheme over GPU clusters. Future Generation Computer Systems, 146(??):166–177, September 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23001553.

Li:2023:PUG

[LWL23a] Qing Li, Ziyue Wang, and Zehao Li. PAGCL: an unsupervised graph poisoned attack for graph contrastive learning model. Future Generation Computer Systems, 149(??):240–249, December 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23002558.

Li:2023:CBU

[LWL23b] Zejun Li, Hao Wu, and Yunlong Lu. Coalition based utility and efficiency optimization for multi-task federated learning in Internet of Vehicles. Future Generation Computer Systems, 140(??):196–208, March 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22003302.

Lu:2020:DDA

[LWLH20] Xiaofeng Lu, Xiaoming Wang, Pietro Lio’, and Pan Hui. DADIM: a distance adjustment dynamic influence map model. Future Generation Computer Systems, 112(??):1122–1130, November 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19332704.

Li:2021:FED

[LWLW21] Zedong Li, Cunrui Wang, Xiangdong Liu, and Yuangang Wang. Facial expression description and recognition based on fuzzy semantic concepts. Future Generation Computer Systems, 114(??):619–628, January 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20309857.

Li:2022:EBR

[LWNH22] Kexin Li, Xingwei Wang, Qiang Ni, and Min Huang. Entropy-based reinforcement learning for computation offloading service in software-defined multiaccess edge computing. Future Generation Computer Systems, 136(??):241–251, November 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22002011.

Li:2023:DBA

[LWS^{+}23a] Beibei Li, Peiran Wang, Zerui Shao, Ao Liu, Yukun Jiang, and Yizhou Li. Defending Byzantine attacks in ensemble federated learning: a reputation-based phishing approach. Future Generation Computer Systems, 147(??):136–148, October 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23001772.

Li:2023:TER

[LWS^{+}23b] Wanting Li, Yongcai Wang, Yu Shao, Gaowei Hu, and Deying Li. TrackPuzzle: Efficient registration of unlabeled PDR trajectories for learning indoor route graph. Future Generation Computer Systems, 149(??):171–183, December 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23002716.

Li:2020:ETB

[LWW^{+}20] Tianpeng Li, Wenjun Wang, Xunxun Wu, Huaming Wu, Pengfei Jiao, and Yandong Yu. Exploring the transition behavior of nodes in temporal networks based on dynamic community detection. Future Generation Computer Systems, 107(??):458–468, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19326901.

Li:2022:IID

[LWW^{+}22] Liying Li, Haizhou Wang, Youyang Wang, Mingsong Chen, and Tongquan Wei. Improving IoT data availability via feedback- and voting-based anomaly imputation. Future Generation Computer Systems, 135(??):194–204, October 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X2200156X.

Li:2024:FGA

[LWW24] Yang Li, Dongrui Wu, and Suhang Wang. Futuregeneration attack and defense in neural networks. Future Generation Computer Systems, 152(??):224, March 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23003916.

Lei:2022:PSA

[LWX22] Jian Lei, Quanwang Wu, and Jin Xu. Privacy and security-aware workflow scheduling in a hybrid cloud. Future Generation Computer Systems, 131 (??):269–278, June 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167739X22000279.

Lu:2024:ABV

[LWY^{+}24] Lingling Lu, Zhenyu Wen, Ye Yuan, Qinming He, Jianhai Chen, and Zhenguang Liu. ANNProof: Building a verifiable and efficient outsourced approximate nearest neighbor search system on blockchain. Future Generation Computer Systems, 156(??):206–220, July 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24000785.

Liu:2020:SAD

[LWZ^{+}20] Yu Liu, Yangtao Wang, Ke Zhou, Yujuan Yang, and Yifei Liu. Semanticaware data quality assessment for image big data. Future Generation Computer Systems, 102(??):53– 65, January 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19302304.

Liu:2023:PQS

[LWZ^{+}23a] Jinhui Liu, Jiaming Wen, Bowen Zhang, Shunyu Dong, Bo Tang, and Yong Yu. A post quantum secure multi-party collaborative signature with deterability in the Industrial Internet of Things. Future Generation Computer Systems, 141(??):663–676, April 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22003983.

Liu:2023:TYO

[LWZ^{+}23b] Song Liu, Xinhe Wan, Zengyuan Zhang, Bo Zhao, and Weiguo Wu. TurboStencil: You only compute once for stencil computation. Future Generation Computer Systems, 146(??):260–272, September 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2300153X.

Liang:2024:TTB

[LXC^{+}24] Wei Liang, Jiahong Xiao, Yuxiang Chen, Chaoyi Yang, Kun Xie, Kuan-Ching Li, and Beniamino Di Martino. TMHD: Twinbridge scheduling of multiheterogeneous dependent tasks for edge computing. Future Generation Computer Systems, 158(??):60– 72, September 2024. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24001560.

Li:2021:LSO

[LXH^{+}21] Zhao Li, Yuying Xing, Jiaming Huang, Haobo Wang, Jianliang Gao, and Guoxian Yu. Large-scale online multi-view graph neural network and applications. Future Generation Computer Systems, 116(??):145–155, March 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20329952.

Li:2021:HOM

[LXL^{+}21] Xiaoming Li, Guangquan Xu, Changzheng Liu, Wei Yu, Zhao Liu, Zhenhuan Wu, and Xiaoping Yang. Higher-order multiple-feature-based community evolution model with potential applications in criminal network investigation. Future Generation Computer Systems, 125(??):364–375, December 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21002235.

Liu:2023:ILS

[LXL^{+}23] Junwen Liu, Ziyun Xiao, Shiyong Lu, Dunren Che, Ming Dong, and Changxin Bai. Infrastructure-level support for GPU-enabled deep learning in DATAVIEW. Future Generation Computer Systems, 141(??): 723–737, April 2023. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22004198.

Liu:2021:MSCa

[LXY21] Yi Liu, Jie Xu, and Weijie Yi. Massive-scale carbon pollution control and biological fusion under big data context. Future Generation Computer Systems, 118(??):257–262, May 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21000042.

Lin:2020:NMU

[LXZ^{+}20] Jie Lin, Biao Xiao, Hanlin Zhang, Xinyu Yang, and Peng Zhao. A novel multitype-users welfare equilibrium based realtime pricing in smart grid. Future Generation Computer Systems, 108(??): 145–160, July 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19322861.

Li:2021:NEE

[LY21] Jianghao Li and Guo Yang. Network embedding enhanced intelligent recommendation for online social networks. Future Generation Computer Systems, 119(??):68–76, June 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21000273.

Lee:2023:OPC

[LY23] Byung Moo Lee and Hong Yang. Optimized power control strategy in massive MIMO for distributed IoT networks. Future Generation Computer Systems, 141(??):433–447, April 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22004010.

Liu:2021:RAS

[LYBS21] Jun Liu, Tianfu Yang, Jingpan Bai, and Bo Sun. Resource allocation and scheduling in the intelligent edge computing context. Future Generation Computer Systems, 121(??):48– 53, August 2021. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21000704.

Liu:2022:FRI

[LYC^{+}22] Weixing Liu, Shaohong Yan, Tailong Chen, Jiaqing Cheng, Ke Wang, Jintao Song, Aimin Yang, Jie Li, Hongwei Xing, and Yuzhu Zhang. Feature recognition of irregular pellet images by regularized extreme learning machine in combination with fractal theory. Future Generation Computer Systems, 127(??):92– 108, February 2022. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21003022.

Li:2020:SMN

[LYFZ20] Wei Li, Kun Yu, Chaolu Feng, and Dazhe Zhao. SP-MIOV: a novel framework of shadow proxy based medical image online visualization in computing and storage resource restrained environments. Future Generation Computer Systems, 105(??):318–330, April 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19310362.

Li:2024:FRU

[LYG^{+}24] Jialun Li, Diying Yang, Hairui Guo, Xuan Mo, and Weigang Wu. Forecasting resource usage pattern changes in clouds via contrast graph-evolution learning. Future Generation Computer Systems, 154 (??):373–383, May 2024. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167739X24000153.

Liu:2021:MSM

[LYGF21] Qinglin Liu, Xiongzhou Yuan, Huaguo Gao, and Xueyi Fu. Mode superposition methods based on hysteretic damping assumption used in seismic calculation of hybrid structures. Future Generation Computer Systems, 124(??):361–368, November 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21002016.

Liu:2021:DLF

[LYH^{+}21] Zhiyong Liu, Chuan Yang, Jun Huang, Shaopeng Liu, Yumin Zhuo, and Xu Lu. Deep learning framework based on integration of S-Mask R-CNN and Inception-v3 for ultrasound image-aided diagnosis of prostate cancer. Future Generation Computer Systems, 114(??):358–367, January 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20317192.

Lee:2022:PQP

[LYKK22] Jemin Lee, Misun Yu, Yongin Kwon, and Taeho Kim. Quantune: Post-training quantization of convolutional neural networks using extreme gradient boosting for fast deployment. Future Generation Computer Systems, 132(??):124–135, July 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22000498.

Li:2024:LBE

[LYP^{+}24] Juyan Li, Mingyan Yan, Jialiang Peng, Haodong Huang, and Ahmed A. Abd El-Latif. A lattice-based efficient certificateless public key encryption for big data security in clouds. Future Generation Computer Systems, 158(??):255–266, September 2024. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24001687.

Liu:2020:JSD

[LYS^{+}20] Ling Liu, Hongfang Yu, Gang Sun, Long Luo, Qixuan Jin, and Shouxi Luo. Job scheduling for distributed machine learning in optical WAN. Future Generation Computer Systems, 112(??):549–560, November 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20302612.

Lu:2023:IMP

[LYW23] Ziyi Lu, Na Yu, and Xuehe Wang. Incentive mechanism and path planning for Unmanned Aerial Vehicle (UAV) hitching over traffic networks. Future Generation Computer Systems, 145(??):521–535, August 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23001334.

Li:2020:EEQ

[LYY^{+}20a] Zhihua Li, Xinrong Yu, Lei Yu, Shujie Guo, and Victor Chang. Energy-efficient and quality-aware VM consolidation method. Future Generation Computer Systems, 102(??):789–809, January 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18324713.

Lv:2020:UPS

[LYY^{+}20b] Siyi Lv, Jinhui Ye, Sijie Yin, Xiaochun Cheng, Chen Feng, Xiaoyan Liu, Rui Li, Zhaohui Li, Zheli Liu, and Li Zhou. Unbalanced private set intersection cardinality protocol with low communication cost. Future Generation Computer Systems, 102(??):1054–1061, January 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19316413.

Liu:2022:TFV

[LYY^{+}22] Yaru Liu, Jia Yu, Ming Yang, Wenqiang Hou, and Huaqun Wang. Towards fully verifiable forward secure privacy preserving keyword search for IoT outsourced data. Future Generation Computer Systems, 128(??):178–191, March 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X2100399X.

Li:2020:EAR

[LYYG20a] Yingfang Li, Bo Yang, Li Yan, and Wei Gao. Energy-aware resource management for uplink nonorthogonal multiple access: Multi-agent deep reinforcement learning. Future Generation Computer Systems, 105(??):684–694, April 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19323295.

See retraction notice [LYYG20b].

Li:2020:RNE

[LYYG20b] Yingfang Li, Bo Yang, Li Yan, and Wei Gao. Retraction notice to “Energyaware resource management for uplink nonorthogonal multiple access: Multi-agent deep reinforcement learning” [Future Gener. Comput. Syst. 105 (2020) 684–694]. Future Generation Computer Systems, 111(??):940, October 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20316897.

See [LYYG20a].

Li:2020:PIU

[LZ20a] Wei Li and Hai Zhuge.

Probabilistic inference on uncertain semantic link network and its application in event identification.

Future Generation Computer Systems, 104(??):32– 42, March 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18322210.

Liu:2020:ISS

[LZ20b] YiHe Liu and Shuang Zhang. Information security and storage of Internet of Things based on block chains. Future Generation Computer Systems, 106(??):296–303, May 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19316784.

Li:2021:UDP

[LZ21a] Xing Li and Lei Zhang. Unbalanced data processing using deep sparse learning technique. Future Generation Computer Systems, 125(??):480–484, December 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21001849.

Li:2021:SMC

[LZ21b] Yanxia Li and Ke Zhao. Sports motional characteristics modeling by leveraging multi-modal image technique. Future Generation Computer Systems, 119(??):37–42, June 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21000418.

Liu:2022:TRT

[LZ22] Jiagang Liu and Xinyu Zhang. Truthful resource trading for dependent task offloading in heterogeneous edge computing. Future Generation Computer Systems, 133(??):228–239, August 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22000851.

Liu:2020:IDF

[LZA^{+}20] Xiaozhu Liu, Rongbo Zhu, Ashiq Anjum, Jun Wang, Hao Zhang, and Maode Ma. Intelligent data fusion algorithm based on hybrid delay-aware adaptive clustering in wireless sensor networks. Future Generation Computer Systems, 104(??):1–14, March 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19308477.

Longo:2020:ATC

[LZB20] Antonella Longo, Marco Zappatore, and Mario A. Bochicchio. Apollon: Towards a citizen science methodology for urban environmental monitoring. Future Generation Computer Systems, 112(??):899–912, November 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20303940.

Liu:2021:MSCb

[LZC21] Yishu Liu, Qi Zhang, and Weixiong Chen. Massivescale complicated human action recognition: Theory and applications. Future Generation Computer Systems, 125(??):806–811, December 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21002739.

Li:2023:HSI

[LZC^{+}23a] Pei Li, Wenqiang Zhu, Jiageng Chen, Shixiong Yao, Ching-Fang Hsu, and Guangquan Xiong. High-speed implementation of rainbow table method on heterogeneous multidevice architecture. Future Generation Computer Systems, 143(??):293–304, June 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23000286.

Liu:2023:TDR

[LZC^{+}23b] Zhizhong Liu, Hedan Zheng, Dianhui Chu, Quan Z. Sheng, Jian Yu, and Xiaofei Xu. Towards dynamic reconfiguration of composite services via failure estimation of general and domain quality of services. Future Generation Computer Systems, 143(??):76– 92, June 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23000250.

Luna:2020:ACS

[LZCGMVV20] Francisco Luna, Pablo H.

Zapata-Cano, Juan C. González-Macías, and Juan F. Valenzuela-Valdés. Approaching the cell switchoff problem in 5G ultradense networks with dynamic multi-objective optimization. Future Generation Computer Systems, 110(??):876–891, September 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19315663.

Liu:2022:PRA

[LZCH22] Linyuan Liu, Haibin Zhu, Shenglei Chen, and Zhiqiu Huang. Privacy regulation aware service selection for multi-provision cloud service composition. Future Generation Computer Systems, 126(??):263–278, January 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21003186.

Li:2024:TSM

[LZF^{+}24] Lianpeng Li, Xu Zhao, Junfang Fan, Fuchao Liu, Ning Liu, and Hui Zhao. A trustworthy security model for IIoT attacks on industrial robots. Future Generation Computer Systems, 153(??):340–349, April 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X2300434X.

Li:2020:RIV

[LZH^{+}20] Jie Li, Jianming Zhi, Wei Hu, Liya Wang, and Aimin Yang. Research on the improvement of vision target tracking algorithm for Internet of things technology and simple extended application in pellet ore phase. Future Generation Computer Systems, 110(??):233–242, September 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20300662.

Lei:2023:CEN

[LZHL23] Zhongcheng Lei, Hong Zhou, Wenshan Hu, and Guo-Ping Liu. Concurrent experimentation in NC-SLab: a scalable approach for online laboratories. Future Generation Computer Systems, 148(??):139–149, November 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2300198X.

Lu:2024:MHP

[LZHS24] Chang Lu, Jie Zhu, Haiping Huang, and Yuzhong Sun. A multi-hierarchy particle swarm optimizationbased algorithm for cloud workflow scheduling. Future Generation Computer Systems, 153(??):125–138, April 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23004387.

Lu:2020:EFH

[LZJ^{+}20] Yu Lu, Xi Zhang, Liwen Jing, Xiaoqing Li, and Xianghua Fu. Estimation of the foetal heart rate baseline based on singular spectrum analysis and empirical mode decomposition. Future Generation Computer Systems, 112(??):126–135, November 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19328298.

Li:2021:CRA

[LZK21] Xing Li, Junpei Zhong, and M. M. Kamruzzaman. Complicated robot activity recognition by quality-aware deep reinforcement learning. Future Generation Computer Systems, 117(??):480–485, April 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20330417.

Li:2020:IIP

[LZL^{+}20] Youzhu Li, Huiling Zhou, Zhonglong Lin, Yifan Wang, Shunjie Chen, Chang Liu, Zhouyang Wang, Daniela Gifu, and Jingbo Xia. Investigation in the influences of public opinion indicators on vegetable prices by corpora construction and WeChat article analysis. Future Generation Computer Systems, 102(??):876–888, January 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18327341.

Lin:2021:IUF

[LZL^{+}21] Qika Lin, Yifan Zhu, Hao Lu, Kaize Shi, and Zhendong Niu. Improving university faculty evaluations via multi-view knowledge graph. Future Generation Computer Systems, 117(??):181–192, April 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20330454.

Li:2023:AAN

[LZL^{+}23] Chuanyou Li, Kun Zhang, Yifan Li, Jiangwei Shang, Xinyue Zhang, and Lei Qian. ANNA: Accelerating Neural Network Accelerator through softwarehardware co-design for vertical applications in edge systems. Future Generation Computer Systems, 140 (??):91–103, March 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167739X22003168.

Li:2024:SBT

[LZL^{+}24a] Hongjian Li, Xue Zhang, Hua Li, Xiaolin Duan, and Chen Xu. SLA-based task offloading for energy consumption constrained workflows in fog computing. Future Generation Computer Systems, 156(??):64– 76, July 2024. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X24000852.

Liu:2024:RBC

[LZL^{+}24b] Xingting Liu, Siwang Zhou, Jie Luo, Jianping Yu, and Wei Zhang. Regionbased compressive distributed storage in mobile CrowdSensing. Future Generation Computer Systems, 158(??):200–209, September 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X2400147X.

Lu:2024:AAF

[LZL^{+}24c] Renhao Lu, Weizhe Zhang, Qiong Li, Hui He, Xiaoxiong Zhong, Hongwei Yang, Desheng Wang, Zenglin Xu, and Mamoun Alazab. Adaptive asynchronous federated learning. Future Generation Computer Systems, 152(??):193–206, March 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23004004.

Li:2020:MOO

[LZLY20] Rui Li, Qinghua Zheng, Xiuqi Li, and Zheng Yan. Multi-objective optimization for rebalancing virtual machine placement. Future Generation Computer Systems, 105(??):824–842, April 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1731840X.

Li:2023:SBF

[LZP23] Yanli Li, Zhengwei Zuo, and Julong Pan. Sensorbased fall detection using a combination model of a temporal convolutional network and a gated recurrent unit. Future Generation Computer Systems, 139(??):53–63, February 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22002953.

Li:2021:DFL

[LZS^{+}21] Xing Li, Zijiang Zhu, Nan Shen, Weihuang Dai, and Yi Hu. Deeply feature learning by CMAC network for manipulating rehabilitation robots. Future Generation Computer Systems, 121(??):19–24, August 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21000455.

Liu:2022:SDD

[LZS^{+}22] Ying Liu, Ting Zhi, Ming Shen, Lu Wang, Yikun Li, and Ming Wan. Softwaredefined DDoS detection with information entropy analysis and optimized deep learning. Future Generation Computer Systems, 129(??):99–114, April 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21004386.

Lin:2024:BBC

[LZS^{+}24] Guanjie Lin, Mingyuan Zeng, Zhiguang Shan, Kaishun Wu, Guan Wang, and Kai Lei. Blockchainbased cooperative game bilateral matching architecture for shared storage. Future Generation Computer Systems, 158(??):122–137, September 2024. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24001456.

Liang:2024:RFL

[LZTM24] Xiang-Yu Liang, Heng-Ru Zhang, Wei Tang, and Fan Min. Robust federated learning with voting and scaling. Future Generation Computer Systems, 153(??):113–124, April 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23004235.

Liu:2021:DNN

[LZW21] Xin Liu, Yanju Zhou, and Zongrun Wang. Deep neural network-based recognition of entities in Chinese online medical inquiry texts. Future Generation Computer Systems, 114(??):581–604, January 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20317453.

Li:2022:GGA

[LZW^{+}22] Jingbo Li, Xingjun Zhang, Jia Wei, Zeyu Ji, and Zheng Wei. GARLSched: Generative adversarial deep reinforcement learning task scheduling optimization for large-scale high performance computing systems. Future Generation Computer Systems, 135(??):259–269, October 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22001613.

Lu:2020:DDC

[LZZ^{+}20] Jiawei Lu, Huan Zhou, Haotian Zhu, Yuanming Zhang, Qianhui Liang, and Gang Xiao. DCEM: a data cell evolution model for service composition based on bigraph theory. Future Generation Computer Systems, 112(??):330–347, November 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19330006.

Li:2023:EAE

[LZZ^{+}23] Beibei Li, Ziqing Zhu, Linghao Zhang, Zhengwei Chang, Liang Zhao, and Arun Kumar. EPPSQ: Achieving efficient and privacy-preserving statistics queries over encrypted data in smart grids. Future Generation Computer Systems, 149(??):265–279, December 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23002765.

Liu:2020:DCS

[LZZX20] Quan Liu, Jie Zuo, Chang Zhu, and Sheng Quan Xie. Design and control of soft rehabilitation robots actuated by pneumatic muscles: State of the art. Future Generation Computer Systems, 113(??):620–634, December 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20306543.

Morell:2022:DAF

[MA22] José Ángel Morell and Enrique Alba. Dynamic and adaptive fault-tolerant asynchronous federated learning using volunteer edge devices. Future Generation Computer Systems, 133(??):53–67, August 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22000735.

Mahdizadeh:2024:AMW

[MA24] Samaneh Hajy Mahdizadeh and Saeid Abrishami. An assignment mechanism for workflow scheduling in Function as a Service edge environment. Future Generation Computer Systems, 157(??):543–557, August 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24001328.

Mostafaei:2022:NAW

[MAA22] Habib Mostafaei, Shafi Afridi, and Jemal Abawajy. Network-aware worker placement for wide-area streaming analytics. Future Generation Computer Systems, 136(??):270–281, November 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22002205.

Matri:2020:MPU

[MAB^{+}20] Pierre Matri, Yevhen Alforov, Álvaro Brandon, María S. Pérez, Alexandru Costan, Gabriel Antoniu, Michael Kuhn, Philip Carns, and Thomas Ludwig. Mission possible: Unify HPC and Big Data stacks towards applicationdefined blobs at the storage layer. Future Generation Computer Systems, 109(??):668–677, August 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17330583.

Mathew:2024:PBS

[MABK24] Anil Mathew, Vasilios Andrikopoulos, Frank J. Blaauw, and Dimka Karastoyanova. Pattern-based serverless data processing pipelines for functionas-a-service orchestration systems. Future Generation Computer Systems, 154(??):87–100, May 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23004855.

McConville:2021:VDH

[MAC^{+}21] Ryan McConville, Gareth Archer, Ian Craddock, Michał Kozłowski, Robert Piechocki, James Pope, and Raul Santos-Rodriguez. Vesta: a digital health analytics platform for a smart home in a box. Future Generation Computer Systems, 114(??):106–119, January 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20307391.

Murad:2024:SPS

[MAM^{+}24] Saydul Akbar Murad, Zafril Rizal M. Azmi, Abu Jafar Md Muzahid, MD. Khairul Bashar Bhuiyan, Md Saib, Nick Rahimi, Nusrat Jahan Prottasha, and Anupam Kumar Bairagi. SG-PBFS: Shortest gappriority based fair scheduling technique for job scheduling in cloud environment. Future Generation Computer Systems, 150(??):232–242, January 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23003266.

Mann:2020:SSP

[Man20] Zoltán Ádám Mann. Secure software placement and configuration. Future Generation Computer Systems, 110(??):243–253, September 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19311525.

Muzammal:2020:DAL

[MAQ^{+}20] Muhammad Muzammal, Faima Abbasi, Qiang Qu, Romana Talat, and Jianping Fan. A decentralised approach for link inference in large signed graphs. Future Generation Computer Systems, 102(??):827–837, January 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18324968.

M:2023:DBT

[MAS23] Revanesh M., John M.

Acken, and V. Sridhar. DAG block: Trust aware load balanced routing and lightweight authentication encryption in WSN. Future Generation Computer Systems, 140(??):402–421, March 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22003272.

Martinez-Aguero:2022:ICT

[MASRAM^{+}22] Sergio Martínez-Agüero, Cristina Soguero-Ruiz, Jose M. Alonso-Moral, Inmaculada Mora-Jiménez, Joaquín Álvarez-Rodríguez, and Antonio G. Marques. Interpretable clinical time-series modeling with intelligent feature selection for early prediction of antimicrobial multidrug resistance. Future Generation Computer Systems, 133(??):68– 83, August 2022. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22000644.

Mthunzi:2020:CCS

[MBB^{+}20] Siyakha N. Mthunzi, Elhadj Benkhelifa, Tomasz Bosakowski, Chirine Ghedira Guegan, and Mahmoud Barhamgi. Cloud computing security taxonomy: From an atomistic to a holistic view. Future Generation Computer Systems, 107(??):620–644, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19300871.

Mora:2024:EGF

[MBB24] Alessio Mora, Armir Bujari, and Paolo Bellavista. Enhancing generalization in Federated Learning with heterogeneous data: a comparative literature review. Future Generation Computer Systems, 157(??):1– 15, August 2024. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24001031.

Mai:2022:LBD

[MBC22] Tai Tan Mai, Marija Bezbradica, and Martin Crane. Learning behaviours data in programming education: Community analysis and outcome prediction with cleaned data. Future Generation Computer Systems, 127(??):42–55, February 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21003344.

Molan:2023:RUA

[MBC^{+}23] Martin Molan, Andrea Borghesi, Daniele Cesarini, Luca Benini, and Andrea Bartolini. RUAD: Unsupervised anomaly detection in HPC systems. Future Generation Computer Systems, 141(??):542–554, April 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2200406X.

Montella:2020:UFI

[MBD^{+}20] Raffaele Montella, Alison Brizius, Diana Di Luccio, Cheryl Porter, Joshua Elliot, Ravi Madduri, David Kelly, Angelo Riccio, and Ian Foster. Using the FACE–IT portal and workflow engine for operational food quality prediction and assessment: an application to mussel farms monitoring in the Bay of Napoli, Italy. Future Generation Computer Systems, 110(??):453–467, September 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16308305.

MirhoseiniNejad:2021:HTA

[MBD21] SeyedMorteza MirhoseiniNejad, Ghada Badawy, and Douglas G. Down. Holistic thermal-aware workload management and infrastructure control for heterogeneous data centers using machine learning. Future Generation Computer Systems, 118(??):208–218, May 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21000169.

Mazzei:2020:BTI

[MBF^{+}20] Daniele Mazzei, Giacomo Baldi, Gualtiero Fantoni, Gabriele Montelisciani, Antonio Pitasi, Laura Ricci, and Lorenzo Rizzello. A Blockchain Tokenizer for Industrial IOT trustless applications. Future Generation Computer Systems, 105(??):432–445, April 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19317686.

Missier:2020:APP

[MBGC20] P. Missier, J. Bryans, C. Gamble, and V. Curcin. Abstracting PROV provenance graphs: a validitypreserving approach. Future Generation Computer Systems, 111(??):352–367, October 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19331760.

Mazurczyk:2020:SIA

[MBJ^{+}20] Wojciech Mazurczyk, Pascal Bisson, Roger Piqueras Jover, Koji Nakao, and Krzysztof Cabaj. Special issue on advancements in 5G networks security. Future Generation Computer Systems, 110(??):314–316, September 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20313662.

Malki:2020:TQO

[MBM^{+}20] Abdelhamid Malki, Sidi-Mohamed Benslimane, Mimoun Malki, Mahmoud Barhamgi, and Djamal Benslimane. Top-k query optimization over data services. Future Generation Computer Systems, 113(??):1–12, December 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19333540.

Masmoudi:2021:KHB

[MBZ^{+}21] Maroua Masmoudi, Sana Ben Abdallah Ben Lamine, Hajer Baazaoui Zghal, Bernard Archimede, and Mohamed Hedi Karray. Knowledge hypergraphbased approach for data integration and querying: Application to Earth observation. Future Generation Computer Systems, 115(??):720–740, February 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20311961.

Malla:2020:ESE

[MC20] Sulav Malla and Ken Christensen. The effect of server energy proportionality on data center power oversubscription. Future Generation Computer Systems, 104(??):119–130, March 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19304145.

Martinez-Casanueva:2024:CFD

[MCBGSL24] Ignacio D. Martinez-Casanueva, Luis Bellido, Daniel González-Sánchez, and Diego Lopez. CANDIL: a federated data fabric for network analytics. Future Generation Computer Systems, 158(??):98– 109, September 2024. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24001420.

Moubarak:2020:DLS

[MCF20] Joanna Moubarak, Maroun Chamoun, and Eric Filiol. On distributed ledgers security and illegal uses. Future Generation Computer Systems, 113(??):183–195, December 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17330650.

M:2022:PIA

[MCT^{+}22] Hariharan M., Sathish Kumar C., Anshul Tanwar, Krishna Sundaresan, Prasanna Ganesan, Sriram Ravi, and R. Karthik. Proximal instance aggregator networks for explainable security vulnerability detection. Future Generation Computer Systems, 134(??):303–318, September 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22001315.

Mulfari:2023:TLA

[MCV23] Davide Mulfari, Lorenzo Carnevale, and Massimo Villari. Toward a lightweight ASR solution for atypical speech on the edge. Future Generation Computer Systems, 149(??):455–463, December 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23003023.

Morell:2024:MOA

[MDC^{+}24] José Ángel Morell, Zakaria Abdelmoiz Dahi, Francisco Chicano, Gabriel Luque, and Enrique Alba. A multi-objective approach for communication reduction in federated learning under devices heterogeneity constraints. Future Generation Computer Systems, 155(??):367–383, June 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24000682.

Martinez:2021:BDC

[MDDZ21] Patricia López Martínez, Ricardo Dintén, José María Drake, and Marta Zorrilla. A big data-centric architecture metamodel for Industry 4.0. Future Generation Computer Systems, 125(??):263–284, December 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21002156.

Mou:2022:EHC

[MDG^{+}22] Jianhui Mou, Peiyong Duan, Liang Gao, Xinhua Liu, and Junqing Li. An effective hybrid collaborative algorithm for energy-efficient distributed permutation flow-shop inverse scheduling. Future Generation Computer Systems, 128(??):521–537, March 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21003939.

Ma:2023:PSB

[MDL^{+}23] Guojun Ma, Yunlong Duan, Mingze Li, Zhibin Xie, and Jin Zhu. A probability smoothing Bi-RRT path planning algorithm for indoor robot. Future Generation Computer Systems, 143(??):349–360, June 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23000456.

Madon:2024:RFH

[MDP24] Maël Madon, Georges Da Costa, and Jean-Marc Pierson. Replay with feedback: How does the performance of HPC system impact user submission behavior? Future Generation Computer Systems, 155 (??):66–79, June 2024. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24000219.

Marinakis:2020:BDS

[MDT^{+}20] Vangelis Marinakis, Haris Doukas, John Tsapelas, Spyros Mouzakitis, Álvaro Sicilia, Leandro Madrazo, and Sgouris Sgouridis. From big data to smart energy services: an application for intelligent energy management. Future Generation Computer Systems, 110(??):572–586, September 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17318769.

Mai:2021:EES

[MDZ^{+}21] Liuyang Mai, Yi Ding, Xiaoning Zhang, Lang Fan, Shui Yu, and Zhichao Xu. Energy efficiency with service availability guarantee for network function virtualization. Future Generation Computer Systems, 119(??):140–153, June 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21000479.

Mohammed:2020:ARS

[MEC^{+}20] Ali Mohammed, Ahmed Eleliemy, Florina M. Ciorba, Franziska Kasielke, and Ioana Banicescu. An approach for realistically simulating the performance of scientific applications on high performance computing systems. Future Generation Computer Systems, 111(??):617–633, October 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19308830.

Macias-Escobar:2020:PSM

[MECRFD20] Teodoro Macias-Escobar, Laura Cruz-Reyes, Héctor Fraire, and Bernabé Dorronsoro. Plane separation: a method to solve dynamic multi-objective optimization problems with incorporated preferences. Future Generation Computer Systems, 110(??):864–875, September 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19315754.

Machado:2023:MBP

[MEL^{+}23] Rafael Ravedutti Lucio Machado, Jan Eitzinger, Jan Laukemann, Georg Hager, Harald Köstler, and Gerhard Wellein. MD-Bench: a performancefocused prototyping harness for state-of-the-art short-range molecular dynamics algorithms. Future Generation Computer Systems, 149(??):25–38, December 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23002467.

Mhaisen:2020:CCR

[MFE^{+}20] Naram Mhaisen, Noora Fetais, Aiman Erbad, Amr Mohamed, and Mohsen Guizani. To chain or not to chain: a reinforcement learning approach for blockchain-enabled IoT monitoring applications. Future Generation Computer Systems, 111(??):39– 51, October 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19334399.

Martinez:2020:TMV

[MFMSG20] Víctor Martínez, Senaka Fernando, Miguel Molina-Solana, and Yike Guo. Tuoris: a middleware for visualizing dynamic graphics in scalable resolution display environments. Future Generation Computer Systems, 106(??):559–571, May 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19321880.

Ma:2024:RAM

[MGB24] Peizhe Ma, Saurabh Garg, and Mutaz Barika. Research allocation in mobile volunteer computing system: Taxonomy, challenges and future work. Future Generation Computer Systems, 154(??):251–265, May 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24000128.

Martini:2023:IBN

[MGC23] B. Martini, M. Gharbaoui, and P. Castoldi. Intent-based network slicing for SDN vertical services with assurance: Context, design and preliminary experiments. Future Generation Computer Systems, 142(??):101–116, May 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2200437X.

Mantovani:2020:PEC

[MGGG^{+}20] Filippo Mantovani, Marta Garcia-Gasulla, José Gracia, Esteban Stafford, Fabio Banchelli, Marc Josep-Fabrego, Joel Criado-Ledesma, and Mathias Nachtmann. Performance and energy consumption of HPC workloads on a cluster based on Arm ThunderX2 CPU. Future Generation Computer Systems, 112(??):800–818, November 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19309781.

Monga:2020:SDN

[MGM^{+}20] Inder Monga, Chin Guok, John MacAuley, Alex Sim, Harvey Newman, Justas Balcas, Phil DeMar, Linda Winkler, Tom Lehman, and Xi Yang. Software-defined network for end-to-end networked science at the exascale. Future Generation Computer Systems, 110(??):181–201, September 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19305618.

Mehrabi:2021:UPC

[MGS21] Mostafa Mehrabi, Nasser Giacaman, and Oliver Sinnen. Unified programming concepts for unobtrusive integration of cloudbased and local parallel computing. Future Generation Computer Systems, 115(??):700–719, February 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19302870.

Ma:2023:VMS

[MGW23] Zhanyou Ma, Shanshan Guo, and Rong Wang. The virtual machines scheduling strategy based on M/M/c queueing model with vacation. Future Generation Computer Systems, 138(??):43–51, January 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2200262X.

Mo:2023:IID

[MGX^{+}23] Zijia Mo, Zhipeng Gao, Kaile Xiao, Chen Zhao, and Xinlei Yu. IDDANet: an Input-Driven Dynamic Adaptive Network ensemble method for edge intelligence. Future Generation Computer Systems, 148(??):436–445, November 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23002200.

Malawski:2020:SES

[MGZ^{+}20] Maciej Malawski, Adam Gajek, Adam Zima, Bartosz Balis, and Kamil Figiela. Serverless execution of scientific workflows: Experiments with HyperFlow, AWS Lambda and Google Cloud functions. Future Generation Computer Systems, 110(??):502–514, September 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1730047X.

Monem:2024:SBB

[MHA^{+}24] Maruf Monem, Md Tamjid Hossain, Md. Golam Rabiul Alam, Md. Shirajum Munir, Md. Mahbubur Rahman, Salman A. AlQahtani, Samah Almutlaq, and Mohammad Mehedi Hassan. A sustainable bitcoin blockchain network through introducing dynamic block size adjustment using predictive analytics. Future Generation Computer Systems, 153 (??):12–26, April 2024. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23004065.

Mahjoub:2020:MDC

[MhCEANSM20] Yassine Idel Mahjoub, El houcine Chakir El-Alaoui, and Ahmed Nait-Sidi-Moh. Modeling and developing a conflict-aware scheduling in urban transportation networks. Future Generation Computer Systems, 107(??):1026–1036, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1732040X.

Mafarja:2020:AWF

[MHH^{+}20] Majdi Mafarja, Ali Asghar Heidari, Maria Habib, Hossam Faris, Thaer Thaher, and Ibrahim Aljarah. Augmented whale feature selection for IoT attacks: Structure, analysis and applications. Future Generation Computer Systems, 112(??):18–40, November 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19313561.

Ma:2020:BBM

[MHL20] Haiying Ma, Elmo X.

Huang, and Kwok-Yan Lam. Blockchain-based mechanism for fine-grained authorization in data crowdsourcing. Future Generation Computer Systems, 106(??):121–134, May 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167739X19300822.

Maketo:2023:MLA

[MIIN23] Lydia Maketo, Tomayess Issa, Theodora Issa, and S. Zaung Nau. M-learning adoption in higher education towards SDG4. Future Generation Computer Systems, 147(??):304–315, October 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23001917.

Marques:2020:ADC

[MIMS20] Diogo Marques, Aleksandar Ilic, Zakhar A. Matveev, and Leonel Sousa. Application-driven cacheaware roofline model. Future Generation Computer Systems, 107(??):257–273, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19309586.

Materwala:2022:ESA

[MISB22] Huned Materwala, Leila Ismail, Raed M. Shubair, and Rajkumar Buyya. Energy-SLA-aware genetic algorithm for edge–cloud integrated computation offloading in vehicular networks. Future Generation Computer Systems, 135(??):205–222, October 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22001327.

Masuduzzaman:2022:UBM

[MISS22] Md Masuduzzaman, Anik Islam, Kazi Sadia, and Soo Young Shin. UAVbased MEC-assisted automated traffic management scheme using blockchain. Future Generation Computer Systems, 134(??): 256–270, September 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167739X22001418.

Mbarek:2022:AOS

[MJB22] Atika Mbarek, Salma Jamoussi, and Abdelmajid Ben Hamadou. An across online social networks profile building approach: Application to suicidal ideation detection. Future Generation Computer Systems, 133(??): 171–183, August 2022. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22000929.

Meng:2021:DAM

[MJSW21] Lun Meng, Feng Ji, Yao Sun, and Tao Wang. Detecting anomalies in microservices with execution trace comparison. Future Generation Computer Systems, 116(??):291–301, March 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20330247.

Mwase:2022:CED

[MJW^{+}22] Christine Mwase, Yi Jin, Tomi Westerlund, Hannu Tenhunen, and Zhuo Zou. Communication-efficient distributed AI strategies for the IoT edge. Future Generation Computer Systems, 131(??):292–308, June 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22000218.

Meng:2023:ARB

[MJW23] Lubin Meng, Xue Jiang, and Dongrui Wu. Adversarial robustness benchmark for EEG-based brain– computer interfaces. Future Generation Computer Systems, 143(??):231–247, June 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23000353.

Mwase:2024:DNT

[MJW^{+}24] Christine Mwase, Yi Jin, Tomi Westerlund, Hannu Tenhunen, and Zhuo Zou. DAI-NET: Toward communicationaware collaborative training for the industrial edge. Future Generation Computer Systems, 155(??): 193–203, June 2024. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24000335.

Mu:2021:DES

[MJZC21] Lin Mu, Peiquan Jin, Jie Zhao, and Enhong Chen. Detecting evolutionary stages of events on social media: a graph-kernel-based approach. Future Generation Computer Systems, 123(??):219–232, October 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21001564.

Makkar:2020:EDL

[MK20] Aaisha Makkar and Neeraj Kumar. An efficient deep learning-based scheme for web spam detection in IoT environment. Future Generation Computer Systems, 108(??):467–487, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19326779.

Makkar:2021:POD

[MK21] Aaisha Makkar and Neeraj Kumar. PROTECTOR: an optimized deep learningbased framework for image spam detection and prevention. Future Generation Computer Systems, 125(??):41–58, December 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21002211.

Mangalampalli:2022:WDN

[MK22] Ashish Mangalampalli and Avinash Kumar. WBATimeNet: a deep neural network approach for VM Live Migration in the cloud. Future Generation Computer Systems, 135(??):438–449, October 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2200187X.

Mampage:2023:DRL

[MKB23] Anupama Mampage, Shanika Karunasekera, and Rajkumar Buyya. Deep reinforcement learning for application scheduling in resourceconstrained, multi-tenant serverless computing environments. Future Generation Computer Systems, 143(??):277–292, June 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167739X2300047X.

Murad:2024:UAA

[MKBT24] Abdulmajid Murad, Frank Alexander Kraemer, Kerstin Bach, and Gavin Taylor. Uncertainty-aware autonomous sensing with deep reinforcement learning. Future Generation Computer Systems, 156(??):242–253, July 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24000918.

Moustafa:2021:DDA

[MKC^{+}21] Nour Moustafa, Marwa Keshk, Kim-Kwang Raymond Choo, Timothy Lynar, Seyit Camtepe, and Monica Whitty. DAD: a distributed anomaly detection system using ensemble one-class statistical learning in edge networks. Future Generation Computer Systems, 118(??):240–251, May 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21000212.

Muhammad:2020:VBP

[MKK^{+}20] Khan Muhammad, Salman Khan, Neeraj Kumar, Javier Del Ser, and Seyedali Mirjalili. Vision-based personalized Wireless Capsule Endoscopy for smart healthcare: Taxonomy, literature review, opportunities and challenges. Future Generation Computer Systems, 113(??):266–280, December 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20304131.

Mehmood:2024:NEA

[MKK^{+}24] Hassan Mehmood, Ahmed Khalid, Panos Kostakos, Ekaterina Gilman, and Susanna Pirttikangas. A novel edge architecture and solution for detecting concept drift in smart environments. Future Generation Computer Systems, 150(??):127–143, January 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23003230.

Ma:2020:LGL

[MLC^{+}20] Tinghuai Ma, Qin Liu, Jie Cao, Yuan Tian, Abdullah Al-Dhelaan, and Mznah Al-Rodhaan. LGIEM: Global and local node influence based community detection. Future Generation Computer Systems, 105(??):533–546, April 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19310210.

Ma:2024:CVD

[MLN24] Kanghua Ma, Shubing Liao, and Yunyun Niu. Connected vehicles’ dynamic route planning based on reinforcement learning. Future Generation Computer Systems, 153(??): 375–390, April 2024. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2300451X.

Mao:2021:EPB

[MLP^{+}21] Qianren Mao, Xi Li, Hao Peng, Jianxin Li, Dongxiao He, Shu Guo, Min He, and Lihong Wang. Event prediction based on evolutionary event ontology knowledge. Future Generation Computer Systems, 115(??):76–89, February 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20311778.

Meng:2020:DIA

[MLWA20] Weizhi Meng, Wenjuan Li, Yu Wang, and Man Ho Au. Detecting insider attacks in medical cyberphysical networks based on behavioral profiling. Future Generation Computer Systems, 108(??):1258–1266, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18300207.

Ma:2023:IQM

[MLX23] Huihui Ma, Xuanhao Luo, and Du Xu. Intelligent queue management of open vSwitch in multitenant data center. Future Generation Computer Systems, 144(??):50–62, July 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23000596.

Martin:2022:KMC

[MLZ^{+}22] Cristian Martín, Peter Langendoerfer, Pouya Soltani Zarrin, Manuel Díaz, and Bartolomé Rubio. Kafka-ML: Connecting the data stream with ML/AI frameworks. Future Generation Computer Systems, 126(??):15–33, January 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21002995.

Min:2023:EDC

[MLZ^{+}23a] Gongshun Min, Liang Liu, Wenbin Zhai, Zijie Wang, and Wanying Lu. An efficient data collection algorithm for partitioned wireless sensor networks. Future Generation Computer Systems, 140(??):53– 66, March 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22002898.

Mo:2023:SSC

[MLZ^{+}23b] Wen Mo, Zeyuan Li, Zhiwen Zeng, Neal N. Xiong, Shaobo Zhang, and Anfeng Liu. SCTD: a spatiotemporal correlation truth discovery scheme for security management of data platform. Future Generation Computer Systems, 139(??):109–125, February 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22003065.

Madichetty:2021:NMI

[MM21a] Sreenivasulu Madichetty and Sridevi M. A novel method for identifying the damage assessment tweets during disaster. Future Generation Computer Systems, 116(??):440–454, March 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20330211.

Mandal:2021:SIW

[MM21b] Anirban Mandal and Raffaele Montella. Special issue on workflows in support of large-scale science. Future Generation Computer Systems, 123(??):105–107, October 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21001370.

Mishra:2023:RLB

[MM23] Pankaj Mishra and Ahmed Moustafa. Reinforcement learning based monotonic policy for online resource allocation. Future Generation Computer Systems, 138(??):313–327, January 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21003691.

Magalhaes:2022:DMC

[MMAH22] Filipe Magalhães, José Monteiro, Juan A. Acebrón, and José R. Herrero. A distributed Monte Carlo based linear algebra solver applied to the analysis of large complex networks. Future Generation Computer Systems, 127(??):320–330, February 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21003605.

MirhoseiniNejad:2020:JDC

[MMBD20] SeyedMorteza MirhoseiniNejad, Hosein Moazamigoodarzi, Ghada Badawy, and Douglas G. Down. Joint data center cooling and workload management: a thermal-aware approach. Future Generation Computer Systems, 104(??): 174–186, March 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19302547.

Maniriho:2022:SMS

[MMC22] Pascal Maniriho, Abdun Naser Mahmood, and Mohammad Jabed Morshed Chowdhury. A study on malicious software behaviour analysis and detection techniques: Taxonomy, current trends and challenges. Future Generation Computer Systems, 130(??):1– 18, May 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21004751.

Merino:2023:ESO

[MMC^{+}23] Miguel Guzman Merino, Maria-Cristina Marinescu, Alberto Cascajo, Jesus Carretero, and David E. Singh. Evaluating the spread of Omicron COVID-19 variant in Spain. Future Generation Computer Systems, 149(??):547–561, December 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23002753.

Macia:2023:AST

[MMFAB23] Sandra Macià, Pedro J. Martínez-Ferrer, Eduard Ayguadé, and Vicenç Beltran. Assessing Saiph, a task-based DSL for highperformance computational fluid dynamics. Future Generation Computer Systems, 147(??):235–250, October 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23001759.

Mostefaoui:2022:BDA

[MMH^{+}22] Ahmed Mostefaoui, Mohammed Amine Merzoug, Amir Haroun, Anthony Nassar, and François Dessables. Big data architecture for connected vehicles: Feedback and application examples from an automotive group. Future Generation Computer Systems, 134(??):374–387, September 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22001492.

Maronas:2020:EOM

[MMK^{+}20] Marcos Maroñas, Sergi Mateo, Kai Keller, Leonardo Bautista-Gomez, Eduard Ayguadé, and Vicenç Beltran. Extending the OpenCHK model with advanced checkpoint features. Future Generation Computer Systems, 112(??):738–750, November 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20304908.

Makuch:2022:MSS

[MMKS22] Mieszko Makuch, Maciej Malawski, Joanna Kocot, and Tomasz Szepieniec. Model and system for scientific workflows represented in file system directory tree. Future Generation Computer Systems, 133(??): 378–394, August 2022. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X2200098X.

Ma:2020:PPP

[MMM^{+}20] Zhuoran Ma, Jianfeng Ma, Yinbin Miao, Kim-Kwang Raymond Choo, Ximeng Liu, Xiangyu Wang, and Tengfei Yang. PMKT: Privacy-preserving Multiparty Knowledge Transfer for financial market forecasting. Future Generation Computer Systems, 106(??):545–558, May 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19302596.

Madeo:2020:EGT

[MMMZ20] Dario Madeo, Somnath Mazumdar, Chiara Mocenni, and Roberto Zingone. Evolutionary game for task mapping in resource constrained heterogeneous environments. Future Generation Computer Systems, 108(??):762–776, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19314219.

Ma:2023:ADM

[MMP^{+}23] Chenbin Ma, Yulan Ma, Longsheng Pan, Xuemei Li, Chunyu Yin, Rui Zong, and Zhengbo Zhang. Automatic diagnosis of multitask in essential tremor: Dynamic handwriting analysis using multi-modal fusion neural network. Future Generation Computer Systems, 145(??):429–441, August 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23001188.

Mengoni:2020:CEC

[MMPL20] Paolo Mengoni, Alfredo Milani, Valentina Poggioni, and Yuanxi Li. Community elicitation from cooccurrence of activities. Future Generation Computer Systems, 110(??):904–917, September 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19303899.

Marchal:2022:TPM

[MMPV22] Loris Marchal, Thibault Marette, Grégoire Pichon, and Frédéric Vivien. Trading performance for memory in sparse direct solvers using lowrank compression. Future Generation Computer Systems, 130(??):307–320, May 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21004994.

Malakhov:2023:ACT

[MMR23a] Ivan Malakhov, Andrea Marin, and Sabina Rossi. Analysis of the confirmation time in proof-of-work blockchains. Future Generation Computer Systems, 147(??):275–291, October 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23001504.

Massri:2023:RRT

[MMR^{+}23b] Maria Massri, Zoltan Miklos, Philippe Raipin, Pierre Meye, Amaury Bouchra Pilet, and Thomas Hassan. RTGEN++: a relative temporal graph GENerator. Future Generation Computer Systems, 146(??):139–155, September 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23001000.

Muhammad:2021:HAR

[MMU^{+}21] Khan Muhammad, Mustaqeem, Amin Ullah, Ali Shariq Imran, Muhammad Sajjad, Mustafa Servet Kiran, Giovanna Sannino, and Victor Hugo C. de Albuquerque. Human action recognition using attention based LSTM network with dilated CNN features. Future Generation Computer Systems, 125(??):820–830, December 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21002405.

Manzoor:2022:FLE

[MMZI22] Sanaullah Manzoor, Adnan Noor Mian, Ahmed Zoha, and Muhammad Ali Imran. Federated learning empowered mobility-aware proactive content offloading framework for fog radio access networks. Future Generation Computer Systems, 133(??):307–319, August 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22001005.

Moussa:2023:RLB

[MNA^{+}23] Noureddine Moussa, Edmond Nurellari, Kebira Azbeg, Abdellah Boulouz, Karim Afdel, Lahcen Koutti, Mohamed Ben Salah, and Abdelbaki El Belrhiti El Alaoui. A reinforcement learning based routing protocol for software-defined networking enabled wireless sensor network forest fire detection. Future Generation Computer Systems, 149(??):478–493, December 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23003072.

Meneses:2024:ARM

[MNFQ24] Enzo Meneses, Cristóbal A. Navarro, Héctor Ferrada, and Felipe A. Quezada. Accelerating range minimum queries with ray tracing cores. Future Generation Computer Systems, 157(??):98–111, August 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X24001110.

Miletto:2022:PAT

[MNSL22] Marcelo Cogo Miletto, Lucas Leandro Nesi, Lucas Mello Schnorr, and Arnaud Legrand. Performance analysis of taskbased multi-frontal sparse linear solvers: Structure matters. Future Generation Computer Systems, 135(??):409–425, October 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22001789.

Metin:2022:MMF

[MÖ22] Serdar Metin and Can Özturan. Max–min fairness based faucet design for blockchains. Future Generation Computer Systems, 131(??):18–27, June 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22000164.

Metin:2024:QSM

[MÖ24a] Serdar Metin and Can Özturan. Quantised and simulated max–min fairness in blockchain ecosystems. Future Generation Computer Systems, 151(??):260–271, February 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23003242.

Mohammed:2024:LBA

[MO24b] Najib Ahmed Mohammed and Mohamed Othman. A load-balanced algorithm for Internet Gateway placement in Backbone Wireless Mesh Networks. Future Generation Computer Systems, 150(??):144–159, January 2024. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23003254.

Moghaddam:2021:MIM

[MOU^{+}21] Seyedhamid Mashhadi Moghaddam, Michael O’Sullivan, Charles Peter Unsworth, Sareh Fotuhi Piraghaj, and Cameron Walker. Metrics for improving the management of cloud environments — load balancing using measures of Quality of Service, Service Level Agreement Violations and energy consumption. Future Generation Computer Systems, 123(??):142–155, October 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X2100131X.

Moghaddam:2020:EIM

[MOW^{+}20] Seyedhamid Mashhadi Moghaddam, Michael O’Sullivan, Cameron Walker, Sareh Fotuhi Piraghaj, and Charles Peter Unsworth. Embedding individualized machine learning prediction models for energy efficient VM consolidation within cloud data centers. Future Generation Computer Systems, 106(??):221–233, May 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19308969.

Makris:2024:PAC

[MPC^{+}24] Antonios Makris, Evangelos Psomakelis, Emanuele Carlini, Matteo Mordacchini, Theodoros Theodoropoulos, Patrizio Dazzi, and Konstantinos Tserpes. Pro-active component image placement in Edge computing environments. Future Generation Computer Systems, 157(??):344–359, August 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24001353.

Mothukuri:2021:SSP

[MPP^{+}21] Viraaji Mothukuri, Reza M.

Parizi, Seyedamin Pouriyeh, Yan Huang, Ali Dehghantanha, and Gautam Srivastava. A survey on security and privacy of federated learning. Future Generation Computer Systems, 115(??):619–640, February 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20329848.

Marrone:2021:EHL

[MPS21] Stefano Marrone, Cristina Papa, and Carlo Sansone. Effects of hidden layer sizing on CNN fine-tuning. Future Generation Computer Systems, 118(??):48– 55, May 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X2033082X.

Mehrabian:2023:SFI

[MR23a] Hamidreza Mehrabian and Reza Ravanmehr. Sensor fusion for indoor positioning system through improved RSSI and PDR methods. Future Generation Computer Systems, 138(??):254–269, January 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22002874.

Mishra:2023:MAE

[MR23b] Ayaskanta Mishra and Arun Kumar Ray. Multiaccess edge computing assisted ultra-low energy scheduling and harvesting in multi-hop wireless sensor and actuator network for energy neutral self-sustainable next-gen cyber-physical system. Future Generation Computer Systems, 141(??):298–324, April 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22003855.

Montezanti:2020:SED

[MRD^{+}20] Diego Montezanti, Enzo Rucci, Armando De Giusti, Marcelo Naiouf, Dolores Rexachs, and Emilio Luque. Soft errors detection and automatic recovery based on replication combined with different levels of checkpointing. Future Generation Computer Systems, 113(??):240–254, December 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19308404.

Mendes:2024:CNM

[MRM^{+}24] Aldo H. D. Mendes, Michel J. F. Rosa, Marcelo A. Marotta, Aleteia Araujo, Alba C. M. A. Melo, and Célia Ghedini Ralha. MASloud+: a novel multiagent architecture with reasoning models for resource management in multiple providers. Future Generation Computer Systems, 154(??):16–34, May 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23004776.

Mazzocca:2024:EFL

[MRMB24] Carlo Mazzocca, Nicolò Romandini, Rebecca Montanari, and Paolo Bellavista. Enabling federated learning at the edge through the IOTA tangle. Future Generation Computer Systems, 152(??):17– 29, March 2024. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23003849.

Mateless:2020:DAB

[MRMM20] Roni Mateless, Daniel Rejabek, Oded Margalit, and Robert Moskovitch. Decompiled APK based malicious code classification. Future Generation Computer Systems, 110(??):135–147, September 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19325129.

Mohanty:2020:ELI

[MRR^{+}20] Sachi Nandan Mohanty, K. C. Ramya, S. Sheeba Rani, Deepak Gupta, K. Shankar, S. K. Lakshmanaprabu, and Ashish Khanna. An efficient Lightweight integrated Blockchain (ELIB) model for IoT security and privacy. Future Generation Computer Systems, 102(??):1027–1037, January 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19319843.

Martins:2022:HBI

[MRS^{+}22] Inês Martins, João S. Resende, Patrícia R. Sousa, Simão Silva, Luís Antunes, and João Gama. Hostbased IDS: a review and open issues of an anomaly detection system in IoT. Future Generation Computer Systems, 133(??):95– 113, August 2022. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22000760.

Mishra:2020:SDS

[MS20] Sandhya Mishra and Devpriya Soni. Smishing Detector: a security model to detect smishing through SMS content analysis and URL behavior analysis. Future Generation Computer Systems, 108(??):803–815, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19318758.

Mosleh:2024:ECI

[MS24] Mohammad Reza Babaei Mosleh and Saeed Sharifian. An efficient cloudintegrated distributed deep neural network framework for IoT malware classification. Future Generation Computer Systems, 157(??):603–617, August 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24001237.

Moosavi:2024:FSB

[MSA^{+}24] Syed Kumayl Raza Moosavi, Ahsan Saadat, Zainab Abaid, Wei Ni, Kai Li, and Mohsen Guizani. Feature selection based on dataset variance optimization using Hybrid Sine Cosine– Firehawk Algorithm (HSCFHA). Future Generation Computer Systems, 155(??): 272–286, June 2024. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24000621.

Mosteiro-Sanchez:2024:EES

[MSBAU24] Aintzane Mosteiro-Sanchez, Marc Barcelo, Jasone Astorga, and Aitor Urbieta. End to end secure data exchange in value chains with dynamic policy updates. Future Generation Computer Systems, 158(??):333–345, September 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2400195X.

Mu:2023:FPC

[MSC^{+}23] Xutong Mu, Yulong Shen, Ke Cheng, Xueli Geng, Jiaxuan Fu, Tao Zhang, and Zhiwei Zhang. Fed-Proc: Prototypical contrastive federated learning on non-IID data. Future Generation Computer Systems, 143(??):93–104, June 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23000262.

Maia:2020:ECA

[MSG^{+}20] José Maia, Carlos Alberto Severiano, Frederico Gadelha Guimarães, Cristiano Leite de Castro, André Paim Lemos, Juan Camilo Fonseca Galindo, and Miri Weiss Cohen. Evolving clustering algorithm based on mixture of typicalities for stream data mining. Future Generation Computer Systems, 106(??):672–684, May 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19312786.

Mao:2021:ELC

[MSK^{+}21] Shitong Mao, Aliaa Sabry, Yassin Khalifa, James L. Coyle, and Ervin Sejdic. Estimation of laryngeal closure duration during swallowing without invasive X-rays. Future Generation Computer Systems, 115(??):610–618, February 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20304118.

Mayya:2021:MCC

[MSKG21] Veena Mayya, Sowmya Kamath S., Gokul S. Krishnan, and Tushaar Gangavarapu. Multi-channel, convolutional attention based neural model for automated diagnostic coding of unstructured patient discharge summaries. Future Generation Computer Systems, 118(??):374–391, May 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21000236.

Ma:2020:DDC

[MSLJ20] Fuzhe Ma, Tao Sun, Lingyun Liu, and Hongyu Jing. Detection and diagnosis of chronic kidney disease using deep learning-based heterogeneous modified artificial neural network. Future Generation Computer Systems, 111(??):17–26, October 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20308128.

Ma:2024:SDD

[MSLP24] Jie Ma, Wei Su, Yikun Li, and Yihua Peng. Synchronizing DDoS detection and mitigation based graph learning with programmable data plane, SDN. Future Generation Computer Systems, 154 (??):206–218, May 2024. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167739X23004909.

Mehta:2022:FGP

[MSM^{+}22] Yash Mehta, Clemens Stachl, Konstantin Markov, Joseph T. Yun, and Björn W. Schuller. Future-generation personality prediction from digital footprints. Future Generation Computer Systems, 136(??):322–325, November 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22002229.

Maroof:2022:PIP

[MSMJ22] Uzma Maroof, Arash Shaghaghi, Regio Michelin, and Sanjay Jha. iRECOVer: Patch your IoT on-the-fly. Future Generation Computer Systems, 132(??):178–193, July 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22000589.

Masood:2020:DFO

[MSR20] Zaheer Masood, Raza Samar, and Muhammad Asif Zahoor Raja. Design of fractional order epidemic model for future generation tiny hardware implants. Future Generation Computer Systems, 106(??):43– 54, May 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19317170.

Marques:2024:PRM

[MSS^{+}24] Gonçalo Marques, Carlos Senna, Susana Sargento, Luís Carvalho, Luís Pereira, and Ricardo Matos. Proactive resource management for cloud of services environments. Future Generation Computer Systems, 150(??):90–102, January 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23003059.

Marzuni:2021:CMD

[MSTN21] Saeed Mirpour Marzuni, Abdorreza Savadi, Adel N. Toosi, and Mahmoud Naghibzadeh. Cross-MapReduce: Data transfer reduction in geodistributed MapReduce. Future Generation Computer Systems, 115(??): 188–200, February 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167739X20305847.

Mallmann:2020:PAR

[MSV^{+}20] Jackson Mallmann, Altair Olivo Santin, Eduardo Kugler Viegas, Roger Robson dos Santos, and Jhonatan Geremias. PPCensor: Architecture for real-time pornography detection in video streaming. Future Generation Computer Systems, 112(??):945–955, November 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19331073.

Malik:2020:CMA

[MSY20] Haroon Malik, Elhadi M. Shakshuki, and Wook-Sung Yoo. Comparing mobile apps by identifying ‘hot’ features. Future Generation Computer Systems, 107(??):659–669, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18302644.

Ma:2020:TDE

[MSZ^{+}20] Zhulin Ma, Edwin H.-M. Sha, Qingfeng Zhuge, Weiwen Jiang, Runyu Zhang, and Shouzhen Gu. Towards the design of efficient hash-based indexing scheme for growing databases on non-volatile memory. Future Generation Computer Systems, 105 (??):1–12, April 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18331005.

Meftah:2022:THP

[MTA^{+}22] Souhail Meftah, Benjamin Hong Meng Tan, Khin Mi Mi Aung, Lu Yuxiao, Lin Jie, and Bharadwaj Veeravalli. Towards high performance homomorphic encryption for inference tasks on CPU: an MPI approach. Future Generation Computer Systems, 134(??):13–21, September 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22001145.

Moura:2022:TIC

[MTCS22] Ivan Moura, Ariel Teles, Luciano Coutinho, and Francisco Silva. Towards identifying contextenriched multimodal behavioral patterns for digital phenotyping of human behaviors. Future Generation Computer Systems, 131(??):227–239, June 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22000310.

Merlino:2024:FIE

[MTD^{+}24] Giovanni Merlino, Giuseppe Tricomi, Luca D’Agati, Zakaria Benomar, Francesco Longo, and Antonio Puliafito. FaaS for IoT: Evolving serverless towards deviceless in I/Oclouds. Future Generation Computer Systems, 154(??):189–205, May 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23004922.

Mahdavi:2024:CIF

[MTHA24a] Mahdi Mahdavi, Mohammad Hesam Tadayon, Mohammad Sayad Haghighi, and Zahra Ahmadian. Corrigendum to “IoT-friendly, pre-computed and outsourced attribute based encryption” [Future Generation Computer Systems (FGCS) volume 150 (2024) 115–126 / FGCS-D-23-00424]. Future Generation Computer Systems, 154 (??):396, May 2024. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23004880.

See [MTHA24b].

Mahdavi:2024:IFP

[MTHA24b] Mahdi Mahdavi, Mohammad Hesam Tadayon, Mohammad Sayad Haghighi, and Zahra Ahmadian. IoTfriendly, pre-computed and outsourced attribute based encryption. Future Generation Computer Systems, 150(??):115–126, January 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23003151.

See corrigendum [MTHA24a].

Mateless:2021:PHP

[MTM21] Roni Mateless, Oren Tsur, and Robert Moskovitch. Pkg2Vec: Hierarchical package embedding for code authorship attribution. Future Generation Computer Systems, 116(??):49–60, March 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20329976.

Malik:2023:HAS

[MTT^{+}23] Haroon Malik, Shahzaib Tahir, Hasan Tahir, Musfirah Ihtasham, and Fawad Khan. A homomorphic approach for security and privacy preservation of Smart Airports. Future Generation Computer Systems, 141(??):500–513, April 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22004101.

Monge:2021:CCS

[MV21] Marco Antonio Sotelo Monge and Jorge Maestre Vidal. Conceptualization and cases of study on cyber operations against the sustainability of the tactical edge. Future Generation Computer Systems, 125(??):869–890, December 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21002788.

Makki:2021:TLR

[MVLJ21] Majid Makki, Dimitri Van Landuyt, Bert Lagaisse, and Wouter Joosen. Thread-level resource consumption control of tenant custom code in a shared JVM for multitenant SaaS. Future Generation Computer Systems, 115(??):351–364, February 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19330675.

McIntosh:2021:ESA

[MWK^{+}21] Timothy McIntosh, Paul Watters, A. S. M. Kayes, Alex Ng, and Yi-Ping Phoebe Chen. Enforcing situationaware access control to build malware-resilient file systems. Future Generation Computer Systems, 115(??):568–582, February 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20305641.

Miao:2020:ITP

[MWL^{+}20] Yiming Miao, Gaoxiang Wu, Miao Li, Ahmed Ghoneim, Mabrook Al-Rakhami, and M. Shamim Hossain. Intelligent task prediction and computation offloading based on mobileedge cloud computing. Future Generation Computer Systems, 102(??):925–931, January 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19320862.

Mirampalli:2024:ENM

[MWS24] Sreenivasu Mirampalli, Rajeev Wankar, and Satish Narayana Srirama. Evaluating NiFi and MQTT based serverless data pipelines in fog computing environments. Future Generation Computer Systems, 150(??):341–353, January 2024. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2300345X.

Ma:2020:GTI

[MXL^{+}20] Laifa Ma, Zheng Xiao, Kenli Li, Shengli Li, Jianlin Li, and Xiaoping Yi. Game theoretic interpretability for learning based preoperative gliomas grading. Future Generation Computer Systems, 112(??):1–10, November 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20307512.

Meng:2022:AAT

[MXS22] Fan Meng, Danya Xu, and Tao Song. ATDNNS: an adaptive time–frequency decomposition neural networkbased system for tropical cyclone wave height realtime forecasting. Future Generation Computer Systems, 133(??):297–306, August 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22001042.

Ma:2022:BEF

[MXW22] Xuyang Ma, Du Xu, and Katinka Wolter. Blockchainenabled feedback-based combinatorial double auction for cloud markets. Future Generation Computer Systems, 127(??):225–239, February 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21003551.

Ma:2023:EEO

[MXW^{+}23] Yinan Ma, Qi Xu, Yue Wang, Jing Wu, Chengnian Long, and Yi-Bing Lin. EOS: an efficient obstacle segmentation for blind guiding. Future Generation Computer Systems, 140(??):117–128, March 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22003016.

Ma:2023:BED

[MYL^{+}23] Zhaowei Ma, Xiaoming Yuan, Kai Liang, Jie Feng, Li Zhu, Dajun Zhang, and F. Richard Yu. Blockchainescorted distributed deep learning with collaborative model aggregation towards 6G networks. Future Generation Computer Systems, 141(??):555–566, April 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22004009.

Miao:2021:DPB

[MYM^{+}21] Zhang Miao, Peng Yong, Yang Mei, Yin Quanjun, and Xie Xu. A discrete PSO-based static load balancing algorithm for distributed simulations in a cloud environment. Future Generation Computer Systems, 115(??):497–516, February 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19328031.

Ma:2021:HCC

[MYT^{+}21] Tinghuai Ma, Huimin Yang, Qing Tian, Yuan Tian, and Najla Al-Nabhan. A hybrid Chinese conversation model based on retrieval and generation. Future Generation Computer Systems, 114(??):481–490, January 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20300133.

Machida:2023:PAA

[MZA23] Fumio Machida, Qingyang Zhang, and Ermeson Andrade. Performability analysis of adaptive drone computation offloading with fog computing. Future Generation Computer Systems, 145(??):121–135, August 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23001048.

Ma:2022:SAS

[MZL^{+}22] Xiaodong Ma, Jia Zhu, Zhihao Lin, Shanxuan Chen, and Yangjie Qin. A stateof-the-art survey on solving non-IID data in Federated Learning. Future Generation Computer Systems, 135(??):244–258, October 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22001686.

Ma:2021:DRU

[MZLT21] Yixuan Ma, Zhenji Zhang, Deming Li, and Mincong Tang. Deflated reputation using multiplicative long short-term memory neural networks. Future Generation Computer Systems, 118(??):198–207, May 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21000030.

Ma:2024:SCT

[MZX^{+}24] Ruifeng Ma, Yufeng Zhan, Yuanqing Xia, Chuge Wu, Liwen Yang, and Runze Gao. Sonnet: a controltheoretic approach for resource allocation in cluster management. Future Generation Computer Systems, 153(??):169–181, April 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23004272.

Moradi:2020:CFM

[MZZ20] Hossein Moradi, Bahman Zamani, and Kamran Zamanifar. CaaSSET: a framework for model-driven development of context as a service. Future Generation Computer Systems, 105 (??):61–95, April 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19305977.

Nandy:2022:IIA

[NAC^{+}22] Sudarshan Nandy, Mainak Adhikari, Supriya Chakraborty, Ahmed Alkhayyat, and Neeraj Kumar. IBoNN: Intelligent agent-based Internet of Medical Things framework for detecting brain response from electroencephalography signal using bag-of-neural network. Future Generation Computer Systems, 130 (??):241–252, May 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167739X21005008.

Namanya:2020:SHB

[NADY20] Anitta Patience Namanya, Irfan U. Awan, Jules Pagna Disso, and Muhammad Younas. Similarity hash based scoring of portable executable files for efficient malware detection in IoT. Future Generation Computer Systems, 110(??):824–832, September 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18325913.

Nasir:2022:SBS

[NAK^{+}22] Muhammad Hassan Nasir, Junaid Arshad, Muhammad Mubashir Khan, Mahawish Fatima, Khaled Salah, and Raja Jayaraman. Scalable blockchains

— a systematic review. Future Generation Computer Systems, 126(??):136–162, January 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21002971.

Nasiraee:2020:ADA

[NAT20] Hassan Nasiraee and Maede Ashouri-Talouki. Anonymous decentralized attributebased access control for cloud-assisted IoT. Future Generation Computer Systems, 110(??):45–56, September 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19322381.

Nozal:2020:EUP

[NBB20] Raúl Nozal, Jose Luis Bosque, and Ramon Beivide. EngineCL: Usability and performance in heterogeneous computing. Future Generation Computer Systems, 107(??):522–537, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19314323.

Najafi:2021:FMO

[NBJ21] Aniseh Najafi, Majid Bayat, and Hamid Haj Seyyed Javadi. Fair multi-owner search over encrypted data with forward and backward privacy in cloud-assisted Internet of Things. Future Generation Computer Systems, 124(??):285–294, November 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21002053.

Novaes:2021:ADL

[NCLP21] Matheus P. Novaes, Luiz F. Carvalho, Jaime Lloret, and Mario Lemes Proença.

Adversarial deep learning approach detection and defense against DDoS attacks in SDN environments. Future Generation Computer Systems, 125(??):156–167, December 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21002429.

Nurmaini:2020:RDA

[NED^{+}20] Siti Nurmaini, Alexander Edo Tondas, Annisa Darmawahyuni, Muhammad Naufal Rachmatullah, Radiyati Umi Partan, Firdaus Firdaus, Bambang Tutuko, Ferlita Pratiwi, Andre Herviant Juliano, and Rahmi Khoirani. Robust detection of atrial fibrillation from short-term electrocardiogram using convolutional neural networks. Future Generation Computer Systems, 113(??):304–317, December 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20305410.

Nicolaescu:2020:HCE

[NFK^{+}20] Sergiu Stefan Nicolaescu, Adrian Florea, Claudiu Vasile Kifor, Ugo Fiore, Nicolae Cocan, Ilie Receu, and Paolo Zanetti. Human capital evaluation in knowledgebased organizations based on big data analytics. Future Generation Computer Systems, 111(??):654–667, October 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19306351.

Ntumba:2024:ASC

[NGC24] Patient Ntumba, Nikolaos Georgantas, and Vassilis Christophides. Adaptive scheduling of continuous operators for IoT edge analytics. Future Generation Computer Systems, 158(??):277–293, September 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X24001584.

Naha:2020:DBD

[NGCB20] Ranesh Kumar Naha, Saurabh Garg, Andrew Chan, and Sudheer Kumar Battula. Deadline-based dynamic resource allocation and provisioning algorithms in Fog– Cloud environment. Future Generation Computer Systems, 104(??):131–141, March 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19319016.

Nguyen:2024:HCS

[Ngu24] Tam N. Nguyen. Holistic cold-start management in serverless computing cloud with deep learning for time series. Future Generation Computer Systems, 153(??):312–325, April 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2300465X.

Ninh:2020:ERS

[NHTH20] Duong Bao Ninh, Jing He, Vu Thanh Trung, and Dang Phuoc Huy. An effective random statistical method for indoor positioning system using WiFi fingerprinting. Future Generation Computer Systems, 109(??):238–248, August 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19324835.

Nanda:2020:STP

[NHY20] Priyadarsi Nanda, Xiangjian He, and Laurence T. Yang. Security, Trust and Privacy in Cyber (STPCyber): Future trends and challenges. Future Generation Computer Systems, 109(??):446–449, August 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20304854.

Nkenyereye:2021:SCS

[NIB^{+}21] Lewis Nkenyereye, S. M. Riazul Islam, Muhammad Bilal, M. Abdullah-Al-Wadud, Atif Alamri, and Anand Nayyar. Secure crowd-sensing protocol for fog-based vehicular cloud. Future Generation Computer Systems, 120(??):61– 75, July 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21000601.

Nikzad:2020:RBT

[NJB20] Mortaza Nikzad, Kamal Jamshidi, and Ali Bohlooli. A responsibility-based transport control for named data networking. Future Generation Computer Systems, 106(??):518–533, May 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19305230.

Nguyen:2020:GCN

[NK20] Minh Tuan Nguyen and Kiseon Kim. Genetic convolutional neural network for intrusion detection systems. Future Generation Computer Systems, 113(??):418–427, December 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1932730X.

Netti:2020:MLA

[NKB^{+}20] Alessio Netti, Zeynep Kiziltan, Ozalp Babaoglu, Alina Sîrbu, Andrea Bartolini, and Andrea Borghesi. A machine learning approach to online fault classification in HPC systems. Future Generation Computer Systems, 110(??):1009–1022, September 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1932045X.

Nguyen:2023:AID

[NKG23] Tri Nguyen, Risto Katila, and Tuan Nguyen Gia. An advanced Internetof-Drones System with Blockchain for improving quality of service of Search and Rescue: a feasibility study. Future Generation Computer Systems, 140(??):36–52, March 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X2200317X.

Nguyen:2020:USD

[NLO^{+}20] Thin Nguyen, Mark Larsen, Bridianne O’Dea, Hung Nguyen, Duc Thanh Nguyen, John Yearwood, Dinh Phung, Svetha Venkatesh, and Helen Christensen. Using spatiotemporal distribution of geocoded Twitter data to predict US countylevel health indices. Future Generation Computer Systems, 110(??):620–628, September 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17312487.

Nesi:2023:AMP

[NLS23] Lucas Leandro Nesi, Arnaud Legrand, and Lucas Mello Schnorr. Asynchronous multi-phase taskbased applications: Employing different nodes to design better distributions. Future Generation Computer Systems, 147(??):119–135, October 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23001796.

Ning:2020:HEC

[NLSY20] Huansheng Ning, Yunfei Li, Feifei Shi, and Laurence T. Yang. Heterogeneous edge computing open platforms and tools for Internet of Things. Future Generation Computer Systems, 106 (??):67–76, May 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19312841.

Nour:2021:EIC

[NMR21] Boubakr Nour, Hassine Moungla, and Ammar Rayes. Editorial: Informationcentric network enabler communication for Internet of Things. Future Generation Computer Systems, 114(??):629–630, January 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20326893.

Naeem:2021:FAT

[NMRK21] Faizana Naeem, Mujahid Mohsin, Usman Rauf, and Liaqat Ali Khan. Formal approach to thwart against drone discovery attacks: a taxonomy of novel 3D obfuscation mechanisms. Future Generation Computer Systems, 115(??):374–386, February 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X2030337X.

Ning:2021:DST

[NN21] Bai Ning and Liu Na. Deep spatial/temporallevel feature engineering for tennis-based action recognition. Future Generation Computer Systems, 125(??):188–193, December 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X2100217X.

Nanda:2020:HET

[NNH^{+}20] Ashish Nanda, Priyadarsi Nanda, Xiangjian He, Aruna Jamdagni, and Deepak Puthal. A hybrid encryption technique for Secure-GLOR: the adaptive secure routing protocol for dynamic wireless mesh networks. Future Generation Computer Systems, 109(??):521–530, August 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17322409.

Nguyen:2023:TTT

[NNPP23] Tri Nguyen, Huong Nguyen, Juha Partala, and Susanna Pirttikangas. Trusted-MaaS: Transforming trust and transparency mobilityas-a-service with blockchain. Future Generation Computer Systems, 149(??): 606–621, December 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167739X23003138.

Nicolazzo:2020:PPA

[NNUV20] Serena Nicolazzo, Antonino Nocera, Domenico Ursino, and Luca Virgili. A privacypreserving approach to prevent feature disclosure in an IoT scenario. Future Generation Computer Systems, 105(??):502–519, April 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19313974.

Nguyen:2023:JOR

[NPNC23] Duong Tuan Nguyen, Chuan Pham, Kim Khoa Nguyen, and Mohamed Cheriet. Jointly optimized resource allocation for SDN control and forwarding planes in edge-cloud SDN-based networks. Future Generation Computer Systems, 145(??):176–188, August 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23000924.

Navarro:2023:SEE

[NQB^{+}23] Cristóbal A. Navarro, Felipe A. Quezada, Benjamin Bustos, Nancy Hitschfeld, and Rolando Kindelan. A scalable and energy efficient GPU thread map for m-simplex domains. Future Generation Computer Systems, 141(??):651–662, April 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22004241.

Navarro:2020:EGT

[NQH^{+}20] Cristobál A. Navarro, Felipe A. Quezada, Nancy Hitschfeld, Raimundo Vega, and Benjamin Bustos. Efficient GPU thread mapping on embedded 2D fractals. Future Generation Computer Systems, 113(??):158–169, December 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20313418.

Nadalini:2023:RPF

[NRBC23] Davide Nadalini, Manuele Rusci, Luca Benini, and Francesco Conti. Reduced precision floating-point optimization for Deep Neural Network On-Device Learning on microcontrollers. Future Generation Computer Systems, 149(??):212–226, December 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23002728.

Naseem:2020:TBD

[NRMI20] Usman Naseem, Imran Razzak, Katarzyna Musial, and Muhammad Imran. Transformer based deep intelligent contextual embedding for Twitter sentiment analysis. Future Generation Computer Systems, 113(??):58–69, December 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2030306X.

Negueroles:2024:BBD

[NSJ^{+}24] Salvador Cuñat Negueroles, Raúl Reinosa Simón, Matilde Julián, Andreu Belsa, Ignacio Lacalle, Raúl S-Julián, and Carlos E. Palau. A Blockchainbased Digital Twin for IoT deployments in logistics and transportation. Future Generation Computer Systems, 158(??):73– 88, September 2024. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X24001407.

Nolte:2023:SHW

[NSR^{+}23] Hendrik Nolte, Nicolai Spicher, Andrew Russel, Tim Ehlers, Sebastian Krey, Dagmar Krefting, and Julian Kunkel. Secure HPC: a workflow providing a secure partition on an HPC system. Future Generation Computer Systems, 141(??):677–691, April 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X2200423X.

Najm:2022:TCA

[NT22] Moustafa Najm and Venkatesh Tamarapalli. Towards costaware VM migration to maximize the profit in federated clouds. Future Generation Computer Systems, 134(??):53–65, September 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22000954.

Njoya:2022:LOD

[NTA^{+}22] Arouna Ndam Njoya, Christopher Thron, Marah Nana Awa, Ado Adamou Abba Ari, and Abdelhak Mourad Gueroui. Lifetime optimization of dense wireless sensor networks using continuous ring-sector model. Future Generation Computer Systems, 129(??): 212–224, April 2022. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21004684.

Nabavirazavi:2024:EFL

[NTI24] Seyedsina Nabavirazavi, Rahim Taheri, and Sundararaja Sitharama Iyengar. Enhancing federated learning robustness through randomization and mixture. Future Generation Computer Systems, 158(??):28–43, September 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24001390.

Nan:2021:NSK

[NTY^{+}21] Fengtao Nan, Yahui Tang, Po Yang, Zhenli He, and Yun Yang. A novel sub-kmeans based on cotraining approach by transforming single-view into multi-view. Future Generation Computer Systems, 125(??):831–843, December 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21002818.

Nguyen:2024:QSF

[NUB24] Hoa T. Nguyen, Muhammad Usman, and Rajkumar Buyya. QFaaS: a serverless function-as-a-service framework for quantum computing. Future Generation Computer Systems, 154(??):281–300, May 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X24000189.

Nikitin:2022:AEA

[NVS^{+}22] Nikolay O. Nikitin, Pavel Vychuzhanin, Mikhail Sarafanov, Iana S. Polonskaia, Ilia Revin, Irina V. Barabanova, Gleb Maximov, Anna V. Kalyuzhnaya, and Alexander Boukhanovsky. Automated evolutionary approach for the design of composite machine learning pipelines. Future Generation Computer Systems, 127(??):109–125, February 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21003307.

Nie:2023:TSI

[NZY^{+}23] Haojie Nie, Xiangguo Zhao, Xin Yao, Qingling Jiang, Xin Bi, Yuliang Ma, and Yongjiao Sun. Temporal-structural importance weighted graph convolutional network for temporal knowledge graph completion. Future Generation Computer Systems, 143(??):30–39, June 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23000195.

Oueslati:2020:RSA

[OCBO20] Oumaima Oueslati, Erik Cambria, Moez Ben HajHmida, and Habib Ounelli. A review of sentiment analysis research in Arabic language. Future Generation Computer Systems, 112(??):408–430, November 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19311537.

Ortega-Calvo:2023:AAI

[OCMJFB^{+}23] Alberto S. Ortega-Calvo, Roberto Morcillo-Jimenez, Carlos Fernandez-Basso, Karel Gutiérrez-Batista, Maria-Amparo Vila, and Maria J. Martin-Bautista. AIMDP: an artificial intelligence modern data platform. Use case for Spanish national health service data silo. Future Generation Computer Systems, 143(??):248–264, June 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23000432.

Osorio:2022:RLS

[OCSCB22] Alfonso Osorio, Maria Calle, Jose Soto, and John E. Candelo-Becerra. Routing in LoRa for smart cities: a gossip study. Future Generation Computer Systems, 136(??):84– 92, November 2022. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22001996.

Oriol:2021:ERB

[ODET21] Xavier Oriol, Giuseppe De Giacomo, Montserrat Estañol, and Ernest Teniente. Embedding reactive behavior into artifactcentric business process models. Future Generation Computer Systems, 117(??):97–110, April 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20330429.

Otte:2020:TSR

[OdVP20] Pim Otte, Martijn de Vos, and Johan Pouwelse. TrustChain: a Sybil-resistant scalable blockchain. Future Generation Computer Systems, 107(??):770–780, June 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167739X17318988.

Ocana:2020:BSG

[OGO^{+}20] Kary A. C. S. Ocaña, Marcelo Galheigo, Carla Osthoff, Luiz M. R. Gadelha, Fabio Porto, Antônio Tadeu A. Gomes, Daniel de Oliveira, and Ana Tereza Vasconcelos. BioinfoPortal: a scientific gateway for integrating bioinformatics applications on the Brazilian national high-performance computing network. Future Generation Computer Systems, 107(??):192–214, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19318230.

Orozco:2020:MLF

[OHÁV20] Ana Lucila Sandoval Orozco, Carlos Quinto Huamán, Daniel Povedano Álvarez, and Luis Javier García Villalba. A machine learning forensics technique to detect post-processing in digital videos. Future Generation Computer Systems, 111(??):199–212, October 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20306786.

Olszewski:2024:DRD

[OIG24] Dominik Olszewski, Marcin Iwanowski, and Waldemar Graniszewski. Dimensionality reduction for detection of anomalies in the IoT traffic data. Future Generation Computer Systems, 151(??):137–151, February 2024. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23003655.

Ochoa:2024:DCA

[OLLP24] William Ochoa, Jon Legaristi, Felix Larrinaga, and Alain Pérez. Dynamic contextaware workflow management architecture for efficient manufacturing: a ROS-based case study. Future Generation Computer Systems, 153(??):505–520, April 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23004831.

Ochoa:2023:CAW

[OLP23] William Ochoa, Felix Larrinaga, and Alain Pérez. Context-aware workflow management for smart manufacturing: a literature review of Semantic Webbased approaches. Future Generation Computer Systems, 145(??):38–55, August 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23000948.

Ortiz-Martin:2020:IIE

[OMPSPL20] Lara Ortiz-Martin, Pablo Picazo-Sanchez, and Pedro Peris-Lopez. Are the interpulse intervals of an ECG signal a good source of entropy? an in-depth entropy analysis based on NIST 800-90B recommendation. Future Generation Computer Systems, 105(??): 346–360, April 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19315201.

Ojagh:2020:LBO

[OMSL20] Soroush Ojagh, Mohammad Reza Malek, Sara Saeedi, and Steve Liang. A location-based orientationaware recommender system using IoT smart devices and social networks.

Future Generation Computer Systems, 108(??):97– 118, July 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1930562X.

Ono:2020:HCP

[ONK^{+}20] Kenji Ono, Jorji Nonaka, Tomohiro Kawanabe, Masahiro Fujita, Kentaro Oku, and Kazuma Hatta. HIVE: a crossplatform, modular visualization framework for largescale data sets. Future Generation Computer Systems, 112(??):875–883, November 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19308726.

Ozsoyeller:2022:MRM

[OÖA22] Deniz Ozsoyeller, Öznur Özkasap, and Moayad Aloqaily. m-RENDEZVOUS: Multi-agent asynchronous rendezvous search technique. Future Generation Computer Systems, 126(??):185–195, January 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21003125.

Omisore:2021:ALB

[OOB^{+}21] Olatunji Mumini Omisore, Bolanle Adefowoke Ojokoh, Asegunoluwa Eunice Babalola, Tobore Igbe, Yetunde Folajimi, Zedong Nie, and Lei Wang. An affective learning-based system for diagnosis and personalized management of diabetes mellitus. Future Generation Computer Systems, 117(??):273–290, April 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20330193.

Olabode:2023:COH

[OOZ^{+}23] Shola Olabode, Rebecca Owens, Viana Nijia Zhang, Jehana Copilah-Ali, Maxim Kolomeets, Han Wu, Shrikant Malviya, Karolina Markeviciute, Tasos Spiliotopoulos, Cristina Neesham, Lei Shi, and Deborah Chambers. Complex online harms and the smart home: a scoping review. Future Generation Computer Systems, 149(??):664–678, December 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23003199.

Onile:2024:SBH

[OPLB24] Abiodun E. Onile, Eduard Petlenkov, Yoash Levron, and Juri Belikov. Smartgrid-based hybrid digital twins framework for demand side recommendation service provision in distributed power systems. Future Generation Computer Systems, 156(??):142–156, July 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24000906.

Orts:2023:QAS

[OPOG23] F. Orts, A. M. Puertas, G. Ortega, and E. M. Garzón. Quantum annealing solution for the unrelated parallel machine scheduling with priorities and delay of task switching on machines. Future Generation Computer Systems, 148(??):514–523, November 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23002583.

Oliveira:2020:ABS

[ORLV20] Rui André Oliveira, Miquel Martínez Raga, Nuno Laranjeiro, and Marco Vieira. An approach for benchmarking the security of web service frameworks. Future Generation Computer Systems, 110(??):833–848, September 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1831882X.

Ortin:2020:HTS

[ORPPG20] Francisco Ortin, Oscar Rodriguez-Prieto, Nicolas Pascual, and Miguel Garcia. Heterogeneous tree structure classification to label Java programmers according to their expertise level. Future Generation Computer Systems, 105(??):380–394, April 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1931516X.

Oriol:2023:GVT

[OTMN23] Xavier Oriol, Ernest Teniente, Marc Maynou, and Sergi Nadal. Generating valid test data through data cloning. Future Generation Computer Systems, 144(??):179–191, July 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23000614.

Orzechowski:2023:ILD

[OWK^{+}23] Michał Orzechowski, Michał Wrzeszcz, Bartosz Kryza, Łukasz Dutka, Renata G. Słota, and Jacek Kitowski. Indexing legacy data-sets for global access and processing in multi-cloud environments. Future Generation Computer Systems, 148(??):150–159, November 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23002078.

Petrik:2020:UAA

[PABBA20] Olga Petrik, Muhammad Adnan, Kakali Basak, and Moshe Ben-Akiva. Uncertainty analysis of an activity-based microsimulation model for Singapore. Future Generation Computer Systems, 110(??):350–363, September 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1830150X.

Peake:2022:PVP

[PAC^{+}22] Joshua Peake, Martyn Amos, Nicholas Costen, Giovanni Masala, and Huw Lloyd. PACO–VMP: Parallel ant colony optimization for virtual machine placement. Future Generation Computer Systems, 129(??):174–186, April 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21004568.

Perez:2021:ECO

[PAM21] Sergio Pérez, Patricia Arroba, and José M. Moya. Energy-conscious optimization of edge computing through deep reinforcement learning and two-phase immersion cooling. Future Generation Computer Systems, 125(??):891–907, December 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21002934.

Pandey:2020:SMD

[Pan20] Hari Mohan Pandey. Secure medical data transmission using a fusion of bit mask oriented genetic algorithm, encryption and steganography. Future Generation Computer Systems, 111(??):213–225, October 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20303848.

Politou:2020:DCE

[PAP^{+}20] Eugenia Politou, Efthimios Alepis, Constantinos Patsakis, Fran Casino, and Mamoun Alazab. Delegated content erasure in IPFS. Future Generation Computer Systems, 112(??):956–964, November 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19323003.

Park:2020:ABE

[Par20] Hyunhee Park. Adaptive backoff enabled WUR on non-cellular local IoT for extreme low power operation. Future Generation Computer Systems, 108 (??):62–67, July 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19324203.

Paragliola:2022:ETB

[Par22] Giovanni Paragliola. Evaluation of the trade-off between performance and communication costs in federated learning scenario. Future Generation Computer Systems, 136(??):282–293, November 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22002175.

Patrono:2020:CAR

[PAS^{+}20] Luigi Patrono, Luigi Atzori, Petar Solić, Marina Mongiello, and Aitor Almeida. Challenges to be addressed to realize Internet of Things solutions for smart environments. Future Generation Computer Systems, 111(??):873–878, October 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19324628.

Patel:2023:MDM

[PB23] Yashwant Singh Patel and Jatin Bedi. MAG-D: a multivariate attention network based approach for cloud workload forecasting. Future Generation Computer Systems, 142(??):376–392, May 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2300002X.

Pereira:2022:PIH

[PBC^{+}22] Jorge Pereira, Thais Batista, Everton Cavalcante, Arthur Souza, Frederico Lopes, and Nelio Cacho. A platform for integrating heterogeneous data and developing smart city applications. Future Generation Computer Systems, 128(??):

552–566, March 2022. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2100426X.

Peterson:2022:EML

[PBK^{+}22] J. Luc Peterson, Ben Bay, Joe Koning, Peter Robinson, Jessica Semler, Jeremy White, Rushil Anirudh, Kevin Athey, Peer-Timo Bremer, Francesco Di Natale, David Fox, Jim A. Gaffney, Sam A. Jacobs, Bhavya Kailkhura, Bogdan Kustowski, Steven Langer, Brian Spears, Jayaraman Thiagarajan, Brian Van Essen, and Jae-Seung Yeom. Enabling machine learning-ready HPC ensembles with Merlin. Future Generation Computer Systems, 131(??):255–268, June 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22000322.

Peng:2023:PPF

[PBL^{+}23] Shaoliang Peng, Wenxuan Bao, Hao Liu, Xia Xiao, Jiandong Shang, Lin Han, Shan Wang, Xiaolan Xie, and Yang Xu. A peer-to-peer file storage and sharing system based on consortium blockchain. Future Generation Computer Systems, 141(??):197–204, April 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22003715.

Pagani:2022:LBS

[PBM^{+}22] Marco Pagani, Alessandro Biondi, Mauro Marinoni, Lorenzo Molinari, Giuseppe Lipari, and Giorgio Buttazzo. A Linuxbased support for developing real-time applications on heterogeneous platforms with dynamic FPGA reconfiguration. Future Generation Computer Systems, 129(??):125–140, April 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21004362.

Panigrahi:2023:FDC

[PBS23] Monalisa Panigrahi, Sourabh Bharti, and Arun Sharma. FedDCS: a distributed client selection framework for cross device federated learning. Future Generation Computer Systems, 144 (??):24–36, July 2023. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23000420.

Priyanka:2024:YBR

[PBSS24] Priyanka, Naman Baranwal, Kedar Nath Singh, and Amit Kumar Singh. YOLO-based ROI selection for joint encryption and compression of medical images with reconstruction through superresolution network. Future Generation Computer Systems, 150(??):1–9, January 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23003205.

Perez-Chacon:2024:PSB

[PCACTMÁ24] R. Pérez-Chacón, G. Asencio-Cortés, A. Troncoso, and F. Martínez-Álvarez. Pattern sequence-based algorithm for multivariate big data time series forecasting: Application to electricity consumption. Future Generation Computer Systems, 154(??):397–412, May 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23004752.

Paganelli:2021:TDS

[PCC21] Federica Paganelli, Paola Cappanera, and Giovanni Cuffaro. Tenant-defined service function chaining in a multi-site network slice. Future Generation Computer Systems, 121(??):1– 18, August 2021. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21000820.

Peng:2021:BGT

[PCCX21] Weimin Peng, Aihong Chen, Jing Chen, and Haitao Xu. Block generation in a two-dimensional space constructed by Hellinger metric and affinity for weather data fusion and learning inputs. Future Generation Computer Systems, 114(??):383–393, January 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20310050.

Piccialli:2020:PPI

[PCG^{+}20] Francesco Piccialli, Salvatore Cuomo, Fabio Giampaolo, Giampaolo Casolla, and Vincenzo Schiano di Cola. Path prediction in IoT systems through Markov chain algorithm. Future Generation Computer Systems, 109(??): 210–217, August 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19335472.

Popa:2024:SER

[PCI^{+}24] Alin-Bogdan Popa, Bogdan-Călin Ciobanu, Voichiţa Iancu, Florin Pop, and Pantelimon George Popescu. SkySwapping: Entanglement resupply by separating quantum swapping and photon exchange. Future Generation Computer Systems, 158(??):89– 97, September 2024. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X24001602.

Pawlicki:2020:DNI

[PCK20] Marek Pawlicki, Michał Choraś, and Rafał Kozik. Defending network intrusion detection systems against adversarial evasion attacks. Future Generation Computer Systems, 110(??):148–154, September 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20303368.

Pianini:2021:PIC

[PCVN21] Danilo Pianini, Roberto Casadei, Mirko Viroli, and Antonio Natali. Partitioned integration and coordination via the selforganising coordination regions pattern. Future Generation Computer Systems, 114(??):44–68, January 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20304775.

Parnell:2020:TSC

[PDA^{+}20] Thomas Parnell, Celestine Dünner, Kubilay Atasu, Manolis Sifalakis, and Haralampos Pozidis. Terascale coordinate descent on GPUs. Future Generation Computer Systems, 108(??):1173–1191, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17318204.

Prenkaj:2021:HSD

[PDFV21] Bardh Prenkaj, Damiano Distante, Stefano Faralli, and Paola Velardi. Hidden space deep sequential risk prediction on student trajectories. Future Generation Computer Systems, 125(??):532–543, December 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21002569.

Poojara:2022:SDP

[PDJS22] Shivananda R. Poojara, Chinmaya Kumar Dehury, Pelle Jakovits, and Satish Narayana Srirama. Serverless data pipeline approaches for IoT data in fog and cloud computing. Future Generation Computer Systems, 130(??):91– 105, May 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21004933.

Phetsouvanh:2021:UDS

[PDT21] Silivanxay Phetsouvanh, Anwitaman Datta, and Alwen Tiu. On unlinkability and denial of service attacks resilience of whistleblower platforms. Future Generation Computer Systems, 118(??):438–452, May 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21000406.

Ponte-Fernandez:2022:SAD

[PFGDM22] Christian Ponte-Fernández, Jorge González-Domínguez, and María J. Martín. A SIMD algorithm for the detection of epistatic interactions of any order. Future Generation Computer Systems, 132(??):108–123, July 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22000528.

Pons:2022:EHT

[PFP^{+}22] Lucía Pons, Josué Feliu, José Puche, Chaoyi Huang, Salvador Petit, Julio Pons, María E. Gómez, and Julio Sahuquillo. Effect of hyper-threading in latencycritical multithreaded cloud applications and utilization analysis of the major system resources. Future Generation Computer Systems, 131(??):194–208, June 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22000334.

Pons:2023:CWD

[PFS^{+}23] Lucía Pons, Josué Feliu, Julio Sahuquillo, María E. Gómez, Salvador Petit, Julio Pons, and Chaoyi Huang. Cloud white: Detecting and estimating QoS degradation of latency-critical workloads in the public cloud. Future Generation Computer Systems, 138(??):13–25, January 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22002734.

Pathmabandu:2023:PII

[PGCB23] Chehara Pathmabandu, John Grundy, Mohan Baruwal Chhetri, and Zubair Baig. Privacy for IoT: Informed consent management in Smart Buildings. Future Generation Computer Systems, 145(??):367–383, August 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23001322.

Paluch:2020:OSP

[PGHS20] Robert Paluch, Łukasz G. Gajewski, Janusz A. Hołyst, and Boleslaw K. Szymanski. Optimizing sensors placement in complex networks for localization of hidden signal source: a review. Future Generation Computer Systems, 112(??):1070–1092, November 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20306294.

Pastor-Galindo:2023:GTO

[PGMP23] Javier Pastor-Galindo, Félix Gómez Mármol, and Gregorio Martínez Pérez. On the gathering of Tor onion addresses. Future Generation Computer Systems, 145(??):12–26, August 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23000651.

Pastor-Galindo:2024:BDA

[PGSM^{+}24] Javier Pastor-Galindo, Hông-Ân Sandlin, Félix Gómez Mármol, Gérôme Bovet, and Gregorio Martínez Pérez. A Big Data architecture for early identification and categorization of dark web sites. Future Generation Computer Systems, 157(??):67–81, August 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24000967.

Palanca:2020:EAA

[PJBB20] Javier Palanca, Jaume Jordán, Javier Bajo, and Vicent Botti. An energyaware algorithm for electric vehicle infrastructures in smart cities. Future Generation Computer Systems, 108(??):454–466, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1930281X.

Pan:2022:GEG

[PJJ^{+}22] Shirui Pan, Shaoxiong Ji, Di Jin, Feng Xia, and Philip S. Yu. Guest editorial: Graph-powered machine learning in futuregeneration computing systems. Future Generation Computer Systems, 126(??):88–90, January 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21003046.

Park:2024:UME

[PJL^{+}24] Seongwan Park, Woojin Jeong, Yunyoung Lee, Bumho Son, Huisu Jang, and Jaewook Lee. Unraveling the MEV enigma: ABIfree detection model using Graph Neural Networks. Future Generation Computer Systems, 153(??):70– 83, April 2024. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23004223.

Park:2023:MMI

[PJLL23] Jongmin Park, Soohwan Jeong, Byung Suk Lee, and Sungsu Lim. MIGTNet: Metapath Instance-based Graph Transformation Network for heterogeneous graph embedding. Future Generation Computer Systems, 149(??):390–401, December 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23002960.

Pirahandeh:2022:MSN

[PK22] Mehdi Pirahandeh and Deok-Hwan Kim. MS scheduler: New, scalable, and high-performance sparse AVX-2 parity encoding and decoding technique for erasure-coded cloud storage systems. Future Generation Computer Systems, 126(??):123–135, January 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21003010.

Pham:2024:EFL

[PK24] Khanh Quan Pham and Taehong Kim. Elastic Federated Learning with Kubernetes Vertical Pod Autoscaler for edge computing. Future Generation Computer Systems, 158(??):501–515, September 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24001882.

Pallewatta:2022:QAP

[PKB22] Samodha Pallewatta, Vassilis Kostakos, and Rajkumar Buyya. QoS-aware placement of microservicesbased IoT applications in fog computing environments. Future Generation Computer Systems, 131 (??):121–136, June 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0167739X22000206.

Park:2022:MLB

[PKLC22] Jaeeun Park, Jangkyum Kim, Sanghyun Lee, and Jun Kyun Choi. Machine learning based photovoltaic energy prediction scheme by augmentation of on-site IoT data. Future Generation Computer Systems, 134(??):1– 12, September 2022. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22001030.

Park:2023:MFR

[PKLC23] Hyunseo Park, Nakyoung Kim, Gyeong Ho Lee, and Jun Kyun Choi. MultiCNN-FilterLSTM: Resourceefficient sensor-based human activity recognition in IoT applications. Future Generation Computer Systems, 139(??):196–209, February 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22003089.

Paknejad:2021:CIP

[PKR21] Peyman Paknejad, Reihaneh Khorsand, and Mohammadreza Ramezanpour. Chaotic improved PICEA-g-based multi-objective optimization for workflow scheduling in cloud environment. Future Generation Computer Systems, 117 (??):12–28, April 2021. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20330260.

Placzek:2024:PBD

[Pła24] Bartłomiej Płaczek. Predictionbased data reduction with dynamic target node selection in IoT sensor networks. Future Generation Computer Systems, 152(??): 225–238, March 2024. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2300403X.

Panizo-LLedot:2020:MOG

[PLBOC20] Angel Panizo-LLedot, Gema Bello-Orgaz, and David Camacho. A multiobjective genetic algorithm for detecting dynamic communities using a local search driven immigrant’s scheme. Future Generation Computer Systems, 110(??):960–975, September 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19311847.

Park:2024:MAB

[PLHC24] Hyunseo Park, Gyeong Ho Lee, Jaeseob Han, and Jun Kyun Choi. Multiclass autoencoder-based active learning for sensor-based human activity recognition. Future Generation Computer Systems, 151(??):71– 84, February 2024. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23003618.

Pan:2024:MEA

[PLL^{+}24] Yuxuan Pan, Kaiyue Luo, Yunming Liu, Chen Xu, Yu Liu, and Lin Zhang. Mobile edge assisted multiview light field video system: Prototype design and empirical evaluation. Future Generation Computer Systems, 153(??):154–168, April 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23004314.

Potena:2023:SMD

[PLMZ23] Domenico Potena, Antonella Longo, Alex Mircoli, and Marco Zappatore. Semantic modeling and design patterns for IoT ecosystems. Future Generation Computer Systems, 142 (??):1–3, May 2023. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22004290.

Penney:2023:PLD

[PLS^{+}23] Drew Penney, Bin Li, Jaroslaw J. Sydir, Lizhong Chen, Charlie Tai, Stefan Lee, Eoin Walsh, and Thomas Long. PROMPT: Learning dynamic resource allocation policies for network applications. Future Generation Computer Systems, 145(??):164–175, August 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23000936.

Pineiro:2020:IES

[PMCP20] César Piñeiro, Rodrigo Martínez-Castaño, and Juan C. Pichel. Ignis: an efficient and scalable multi-language Big Data framework. Future Generation Computer Systems, 105(??):705–716, April 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19311021.

Platenius-Mohr:2020:FAB

[PMMG^{+}20] Marie Platenius-Mohr, Somayeh Malakuti, Sten Grüner, Johannes Schmitt, and Thomas Goldschmidt. File- and API-based interoperability of digital twins by model transformation: an IIoT case study using asset administration shell. Future Generation Computer Systems, 113(??):94– 105, December 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20302600.

Pena-Monferrer:2021:HCN

[PMMSE21]  Carlos  Peña-Monferrer,

Robert Manson-Sawko, and Vadim Elisseev. HPC-cloud native framework for concurrent simulation, analysis and visualization of CFD workflows. Future Generation Computer Systems, 123(??):14–23, October 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21001291.

Pishvaei:2022:OAM

[PMT22] Seyed Mahmoud Pishvaei, Farid Tabee Miandoab, and Behzad Mozaffari Tazehkand. Outage analysis of mmWave– NOMA transmission in the presence of LOS and NLOS paths. Future Generation Computer Systems, 128 (??):88–101, March 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167739X21003873.

Phan:2021:DFF

[PNL^{+}21] Linh-An Phan, Duc-Thang Nguyen, Meonghun Lee, Dae-Heon Park, and Taehong Kim. Dynamic fog-to-fog offloading in SDN-based fog computing systems. Future Generation Computer Systems, 117(??):486–497, April 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20330831.

Pourhabibi:2021:DCC

[POBK21] Tahereh Pourhabibi, Kok-Leong Ong, Yee Ling Boo, and Booi H. Kam. Detecting covert communities in multi-layer networks: a network embedding approach. Future Generation Computer Systems, 124(??):467–479, November 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21002223.

Pandya:2020:UDS

[POMK20] Abhinay Pandya, Mourad Oussalah, Paola Monachesi, and Panos Kostakos. On the use of distributed semantics of tweet metadata for user age prediction. Future Generation Computer Systems, 102(??):437–452, January 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19304509.

Pradhan:2020:HPS

[PP20] Tribikram Pradhan and Sukomal Pal. A hybrid personalized scholarly venue recommender system integrating social network analysis and contextual similarity. Future Generation Computer Systems, 110 (??):1139–1166, September 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19307101.

Pineiro:2022:UFI

[PP22] César Piñeiro and Juan C.

Pichel. A unified framework to improve the interoperability between HPC and Big Data languages and programming models. Future Generation Computer Systems, 134(??):123–139, September 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2200125X.

Peng:2024:ITQ

[PP24] Yaqiong Peng and Haocheng Peng. InferFair: Towards QoS-aware scheduling for performance isolation guarantee in heterogeneous model serving systems. Future Generation Computer Systems, 150(??):10– 20, January 2024. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23003187.

Pournaropoulos:2024:FPF

[PPA^{+}24]  Foivos  Pournaropoulos,

Alexandros Patras, Christos D. Antonopoulos, Nikos Bellas, and Spyros Lalis. Fluidity: Providing flexible deployment and adaptation policy experimentation for serverless and distributed applications spanning cloud–edge–mobile environments. Future Generation Computer Systems, 157(??):210–225, August 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24000980.

Persia:2020:IOB

[PPG^{+}20] Fabio Persia, Giovanni Pilato, Mouzhi Ge, Paolo Bolzoni, Daniela D’Auria, and Sven Helmer. Improving orienteering-based tourist trip planning with social sensing. Future Generation Computer Systems, 110(??):931–945, September 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19303929.

Puche:2020:ECF

[PPGS20] José Puche, Salvador Petit, María E. Gómez, and Julio Sahuquillo. An efficient cache flat storage organization for multithreaded workloads for low power processors. Future Generation Computer Systems, 110(??):1037–1054, September 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1930384X.

Prandi:2024:PSI

[PPM24] Catia Prandi, Diogo Pacheco, and Costas Mourlas. Preface of special issue on future generation ICT solutions for digital social innovation and sustainable development. Future Generation Computer Systems, 153(??):537–538, April 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23004818.

Pelle:2023:PAS

[PPSC23] István Pelle, Francesco Paolucci, Balázs Sonkoly, and Filippo Cugini. P4assisted seamless migration of serverless applications towards the edge continuum. Future Generation Computer Systems, 146(??):122–138, September 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23001450.

Premarathne:2020:TBM

[PR20] U. S. Premarathne and S. Rajasingham. Trust based multi-agent cooperative load balancing system (TCLBS). Future Generation Computer Systems, 112(??):185–192, November 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19321089.

Pierro:2022:UOM

[PRD^{+}22] Giuseppe Antonio Pierro, Henrique Rocha, Stéphane Ducasse, Michele Marchesi, and Roberto Tonelli. A user-oriented model for Oracles’ Gas price prediction. Future Generation Computer Systems, 128(??): 142–157, March 2022. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21003678.

Posner:2020:CAL

[PRF20] Jonas Posner, Lukas Reitz, and Claudia Fohry. A comparison of application-level fault tolerance schemes for task pools. Future Generation Computer Systems, 105(??):119–134, April 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19308295.

Pedratscher:2022:MTF

[PRF22] Stefan Pedratscher, Sasko Ristov, and Thomas Fahringer. M2FaaS: Transparent and fault tolerant FaaSification of Node.js monolith code blocks. Future Generation Computer Systems, 135(??):57–71, October 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22001509.

Perez-Rodriguez:2020:MSH

[PRPPFRL20] Gael Pérez-Rodríguez, Martín Pérez-Pérez, Florentino Fdez-Riverola, and Anália Lourenço. Mining the sociome for health informatics: Analysis of therapeutic lifestyle adherence of diabetic patients in Twitter. Future Generation Computer Systems, 110(??):214–232, September 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19329516.

Pradeep:2020:QNQ

[PS20] D. Pradeep and C. Sundar.

QAOC: Novel query analysis and ontology-based clustering for data management in Hadoop. Future Generation Computer Systems, 108(??):849–860, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1933050X.

Parah:2020:EHR

[PSAL20] Shabir A. Parah, Javaid A. Sheikh, Jahangir A. Akhoon, and Nazir A. Loan. Electronic health record hiding in images for smart city applications: a computationally efficient and reversible information hiding technique for secure communication. Future Generation Computer Systems, 108(??):935–949, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17324676.

Panadero:2021:TSM

[PSC^{+}21] Javier Panadero, Mennan Selimi, Laura Calvet, Joan Manuel Marquès, and Felix Freitag. A twostage multi-criteria optimization method for service placement in decentralized edge micro-clouds. Future Generation Computer Systems, 121(??):90– 105, August 2021. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21000935.

Petrovic:2020:BSH

[PSH^{+}20] Filip Petrovič, David Střelák, Jana Hozzová, Jaroslav Ol’ha, Richard Trembecký, Siegfried Benkner, and Jiří Filipovič. A benchmark set of highly-efficient CUDA and OpenCL kernels and its dynamic autotuning with Kernel Tuning Toolkit. Future Generation Computer Systems, 108(??):161–177, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19327360.

Prybila:2020:RVB

[PSHW20] Christoph Prybila, Stefan Schulte, Christoph Hochreiner, and Ingo Weber. Runtime verification for business processes utilizing the Bitcoin blockchain. Future Generation Computer Systems, 107(??):816–831, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1731837X.

Paredes:2021:DMB

[PSMF21] Jose N. Paredes, Gerardo I.

Simari, Maria Vanina Martinez, and Marcelo A. Falappa. Detecting malicious behavior in social platforms via hybrid knowledge- and data-driven systems. Future Generation Computer Systems, 125(??):232–246, December 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21002284.

Perera:2023:DAP

[PSS^{+}23] Niranda Perera, Arup Kumar Sarker, Mills Staylor, Gregor von Laszewski, Kaiying Shan, Supun Kamburugamuve, Chathura Widanage, Vibhatha Abeykoon, Thejaka Amila Kanewela, and Geoffrey Fox. In-depth analysis on parallel processing patterns for highperformance dataframes. Future Generation Computer Systems, 149(??): 250–264, December 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167739X23002595.

Pease:2020:ISS

[PSvL^{+}20] Sarogini Grace Pease, Richard Sharpe, Kate van Lopik, Eleni Tsalapati, Paul Goodall, Bob Young, Paul Conway, and Andrew West. An interoperable semantic service toolset with domain ontology for automated decision support in the end-of-life domain. Future Generation Computer Systems, 112(??):848–858, November 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19326640.

Psomakelis:2020:CAT

[PTZ^{+}20] Evangelos Psomakelis, Konstantinos Tserpes, Dimitris Zissis, Dimosthenis Anagnostopoulos, and Theodora Varvarigou. Context agnostic trajectory prediction based on λarchitecture. Future Generation Computer Systems, 110(??):531–539, September 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17320204.

Parra-Ullauri:2024:KPP

[PUMN^{+}24] Juan Marcelo Parra-Ullauri, Hari Madhukumar, Adrian-Cristian Nicolaescu, Xunzheng Zhang, Anderson Bravalheri, Rasheed Hussain, Xenofon Vasilakos, Reza Nejabati, and Dimitra Simeonidou. kubeFlower: a privacy-preserving framework for Kubernetes-based federated learning in cloud– edge environments. Future Generation Computer Systems, 157(??):558–572, August 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24001134.

Psychas:2020:CTP

[PVA^{+}20] A. Psychas, J. Violos, F. Aisopos, A. Evangelinou, G. Kousiouris, I. Bouras, T. Varvarigou, G. Xidas, D. Charilas, and Y. Stavroulas. Cloud toolkit for provider assessment, optimized application cloudification and deployment on IaaS. Future Generation Computer Systems, 109(??):657–667, August 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17329357.

Peng:2022:MMS

[PWH^{+}22] Hu Peng, Cong Wang, Yupeng Han, Wenhui Xiao, Xinyu Zhou, and Zhijian Wu. Micro multi-strategy multi-objective artificial bee colony algorithm for microgrid energy optimization. Future Generation Computer Systems, 131 (??):59–74, June 2022. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X2200019X.

Papadimitriou:2021:EEO

[PWV^{+}21] George Papadimitriou, Cong Wang, Karan Vahi, Rafael Ferreira da Silva, Anirban Mandal, Zhengchun Liu, Rajiv Mayani, Mats Rynge, Mariam Kiran, Vickie E. Lynch, Rajkumar Kettimuthu, Ewa Deelman, Jeffrey S. Vetter, and Ian Foster. End-to-end online performance data capture and analysis for scientific workflows. Future Generation Computer Systems, 117(??):387–400, April 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20330570.

Prabono:2022:MID

[PYL22] Aria Ghora Prabono, Bernardo Nugroho Yahya, and Seok-Lyong Lee. Multipleinstance domain adaptation for cost-effective sensor-based human activity recognition. Future Generation Computer Systems, 133(??):114–123, August 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22000814.

Peng:2020:GAI

[PZHD20] Shenglong Peng, Liang Zhou, Xuan He, and Junyi Du. GPS-aided inter-microcell interference avoidance for requesttransmission splitting slotted ALOHA-based scheme in smart cities with connected vehicles. Future Generation Computer Systems, 108(??):501–511, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20300224.

Pu:2021:FCC

[PZLL21] Bin Pu, Ningbo Zhu, Kenli Li, and Shengli Li. Fetal cardiac cycle detection in multi-resource echocardiograms using hybrid classification framework. Future Generation Computer Systems, 115(??):825–836, February 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20307524.

Qu:2021:IVI

[QC21] Zehui Qu and Zhongsheng Chen. An intelligent vehicle image segmentation and quality assessment model. Future Generation Computer Systems, 117(??): 426–432, April 2021. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20330648.

Qi:2024:MAT

[QCG^{+}24] Pian Qi, Diletta Chiaro, Antonella Guzzo, Michele Ianni, Giancarlo Fortino, and Francesco Piccialli. Model aggregation techniques in federated learning: a comprehensive survey. Future Generation Computer Systems, 150(??):272–293, January 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23003333.

Qin:2024:MMA

[QCW^{+}24] Wei Qin, Haiming Chen, Lei Wang, Yinshui Xia, Alfredo Nascita, and Antonio Pescapè. MCOTM: Mobility-aware computation offloading and task migration for edge computing in industrial IoT. Future Generation Computer Systems, 151(??):232–241, February 2024. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23003795.

Qin:2023:HSC

[QCWY23] Tian Qin, Guang Cheng, Yichen Wei, and Zifan Yao. Hier-SFL: Clientedge-cloud collaborative traffic classification framework based on hierarchical federated split learning. Future Generation Computer Systems, 149(??):12– 24, December 2023. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23002510.

Qu:2021:DLB

[QCY^{+}21] Chengyi Qu, Prasad Calyam, Jeromy Yu, Aditya Vandanapu, Osunkoya Opeoluwa, Ke Gao, Songjie Wang, Raymond Chastain, and Kannappan Palaniappan. DroneCOCoNet: Learning-based edge computation offloading and control networking for drone video analytics. Future Generation Computer Systems, 125(??):247–262, December 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21002351.

Qin:2020:NMN

[QG20] Peng Qin and Jingzhi Guo. A novel machine natural language mediation for semantic document exchange in smart city. Future Generation Computer Systems, 102(??):810–826, January 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19305679.

Qi:2022:SBD

[QGH^{+}22] Hui Qi, Yingjun Guo, Dinghui Hou, Ziyang Xing, Weiwu Ren, Ligang Cong, and Xiaoqiang Di. SDNbased dynamic multi-path routing strategy for satellite networks. Future Generation Computer Systems, 133(??):254–265, August 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22000875.

Qin:2024:TMR

[QHC24] Peng Qin, Quanyi Hu, and Menglin Cui. Towards machine-readable semanticbased e-business contract representations using network of timed automata (NTA). Future Generation Computer Systems, 158(??):457–471, September 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24001699.

Qiang:2024:IAD

[QHCH24] Xianke Qiang, Yun Hu, Zheng Chang, and Timo Hamalainen. Importanceaware data selection and resource allocation for hierarchical federated edge learning. Future Generation Computer Systems, 154 (??):35–44, May 2024. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23004740.

Qiao:2020:PON

[QHE^{+}20] Yuchen Qiao, Kazuma Hashimoto, Akiko Eriguchi, Haixia Wang, Dongsheng Wang, Yoshimasa Tsuruoka, and Kenjiro Taura. Parallelizing and optimizing neural encoder–decoder models without padding on multi-core architecture. Future Generation Computer Systems, 108(??): 1206–1213, July 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17318150.

Qi:2021:PPB

[QHNL21] Yuanhang Qi, M. Shamim Hossain, Jiangtian Nie, and Xuandi Li. Privacypreserving blockchain-based federated learning for traffic flow prediction. Future Generation Computer Systems, 117(??):328–337, April 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X2033065X.

Qin:2020:CDP

[QHW^{+}20] Bo Qin, Jikun Huang, Qin Wang, Xizhao Luo, Bin Liang, and Wenchang Shi. Cecoin: a decentralized PKI mitigating MitM attacks. Future Generation Computer Systems, 107(??):805–815, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17318381.

Qi:2021:SPI

[QJS^{+}21] Rongxin Qi, Sai Ji, Jian Shen, Pandi Vijayakumar, and Neeraj Kumar. Security preservation in industrial medical CPS using Chebyshev map: an AI approach. Future Generation Computer Systems, 122(??):52–62, September 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21000881.

Qamar:2020:VFU

[QJZ^{+}20] Saqib Qamar, Hai Jin, Ran Zheng, Parvez Ahmad, and Mohd Usama. A variant form of 3D-UNet for infant brain segmentation. Future Generation Computer Systems, 108(??):613–623, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18332291.

Qiu:2020:ISO

[QKG20] Meikang Qiu, Sun-Yuan Kung, and Keke Gai. Intelligent security and optimization in edge/fog computing. Future Generation Computer Systems, 107(??):1140–1142, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19314852.

Qiu:2022:AMP

[QL22] Hao Qiu and Tong Li. Auction method to prevent bidrigging strategies in mobile blockchain edge computing resource allocation. Future Generation Computer Systems, 128(??):1– 15, March 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21003770.

Quilis:2023:FCA

[QLHLB23] J. Damián Segrelles Quilis, Sergio López-Huguet, Pau Lozano, and Ignacio Blanquer. A federated cloud architecture for processing of cancer images on a distributed storage. Future Generation Computer Systems, 139(??):38–52, February 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2200303X.

Lu:2020:WRC

[qLhZ20] An qi Lu and Jing hua Zhu.

Worker recruitment with cost and time constraints in mobile crowd sensing. Future Generation Computer Systems, 112(??):819–831, November 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20303952.

Qiang:2021:DCA

[QLJ21] Weizhong Qiang, Renwan Liu, and Hai Jin. Defending CNN against privacy leakage in edge computing via binary neural networks. Future Generation Computer Systems, 125(??):460–470, December 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21002326.

Qi:2020:NBM

[QMCX20] Xiaoyu Qi, Gang Mei, Salvatore Cuomo, and Lei Xiao. A networkbased method with privacypreserving for identifying influential providers in large healthcare service systems. Future Generation Computer Systems, 109(??):293–305, August 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20303629.

Quezada:2022:SEC

[QNHB22] Felipe A. Quezada, Cristóbal A. Navarro, Nancy Hitschfeld, and Benjamin Bustos. Squeeze: Efficient compact fractals for tensor core GPUs. Future Generation Computer Systems, 135(??):10–19, October 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22001522.

Qashlan:2024:DPM

[QNM24] Amjad Qashlan, Priyadarsi Nanda, and Manoranjan Mohanty. Differential privacy model for blockchain based smart home architecture. Future Generation Computer Systems, 150(??):49–63, January 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23003102.

Quezada:2023:MGD

[QNRA23] Felipe A. Quezada, Cristóbal A.

Navarro, Miguel Romero, and Cristhian Aguilera. Modeling GPU dynamic parallelism for self similar density workloads. Future Generation Computer Systems, 145(??):239–253, August 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23001292.

Qi:2022:LOB

[QPL22a] Yuxiao Qi, Li Pan, and Shijun Liu. A Lyapunov optimization-based online scheduling algorithm for service provisioning in cloud computing. Future Generation Computer Systems, 134(??):40– 52, September 2022. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22001182.

Qiao:2022:RER

[QPL^{+}22b] Sibo Qiao, Shanchen Pang, Gang Luo, Silin Pan, Zengchen Yu, Taotao Chen, and Zhihan Lv. RLDS: an explainable residual learning diagnosis system for fetal congenital heart disease.

Future Generation Computer Systems, 128(??): 205–218, March 2022. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21003915.

Qin:2021:LUA

[QRS^{+}21] Yubo Qin, Ivan Rodero, Anthony Simonet, Charles Meertens, Daniel Reiner, James Riley, and Manish Parashar. Leveraging user access patterns and advanced cyberinfrastructure to accelerate data delivery from shared-use scientific observatories. Future Generation Computer Systems, 122(??):14–27, September 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21000844.

Qiao:2024:EPC

[QSZ^{+}24] Yufei Qiao, Shihao Shen, Cheng Zhang, Wenyu Wang, Tie Qiu, and Xiaofei Wang. EdgeOptimizer: a programmable containerized scheduler of timecritical tasks in Kubernetesbased edge-cloud clusters. Future Generation Computer Systems, 156(??): 221–230, July 2024. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24000748.

Qiao:2020:IPS

[QWR^{+}20] Xiuquan Qiao, Hongyi Wang, Pei Ren, Yukai Tu, Guoshun Nan, Junliang Chen, and M. Brian Blake. Interest packets scheduling and sizebased flow control mechanism for content-centric networking web servers. Future Generation Computer Systems, 107(??):564–577, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19325208.

Qiu:2024:IBM

[QZZ^{+}24] Shaoming Qiu, Jiancheng Zhao, Xuecui Zhang, Fen Chen, and Yahui Wang. Improved binary marine predator algorithm-based digital twin-assisted edgecomputing offloading method. Future Generation Computer Systems, 155(??): 437–446, June 2024. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24000669.

Qiu:2021:MIR

[QZZH21] Defu Qiu, Lixin Zheng, Jianqing Zhu, and Detian Huang. Multiple improved residual networks for medical image super-resolution. Future Generation Computer Systems, 116(??): 200–208, March 2021. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20330259.

Raza:2020:EEC

[RAA^{+}20] Mohsin Raza, Muhammad Awais, Kamran Ali, Nauman Aslam, Vishnu Vardhan Paranthaman, Muhammad Imran, and Farman Ali. Establishing effective communications in disaster affected areas and artificial intelligence based detection using social media platform. Future Generation Computer Systems, 112(??):1057–1069, November 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20302727.

Razaque:2021:ERF

[RAA^{+}21] Abdul Razaque, Moayad Aloqaily, Muder Almiani, Yaser Jararweh, and Gautam Srivastava. Efficient and reliable forensics using intelligent edge computing. Future Generation Computer Systems, 118(??): 230–239, May 2021. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21000224.

Randhawa:2024:DRL

[RAA^{+}24] Rizwan Hamid Randhawa, Nauman Aslam, Mohammad Alauthman, Muhammad Khalid, and Husnain Rafiq. Deep reinforcement learning based Evasion Generative Adversarial Network for botnet detection. Future Generation Computer Systems, 150(??):294–302, January 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23003369.

Roman:2023:LRA

[RAB23] Roberto Román, Rosario Arjona, and Iluminada Baturone. A lightweight remote attestation using PUFs and hash-based signatures for low-end IoT devices. Future Generation Computer Systems, 148(??):425–435, November 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23002236.

Risco:2024:RSW

[RAL^{+}24] Sebastián Risco, Caterina Alarcón, Sergio Langarita, Miguel Caballer, and Germán Moltó. Rescheduling serverless workloads across the cloud-to-edge continuum. Future Generation Computer Systems, 153(??):457–466, April 2024. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0167739X23004764.

Rahman:2020:FMT

[RAN^{+}20] Fatin Hamadah Rahman, Thien-Wan Au, S. H. Shah Newaz, Wida Susanty Suhaili, and Gyu Myoung Lee. Find my trustworthy fogs: a fuzzy-based trust evaluation framework. Future Generation Computer Systems, 109(??):562–572, August 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17322367.

Rawashdeh:2020:KDA

[RAS^{+}20] Majdi Rawashdeh, Mohammed Gh. Al Zamil, Samer Samarah, M. Shamim Hossain, and Ghulam Muhammad. A knowledgedriven approach for activity recognition in smart homes based on activity profiling.

Future Generation Computer Systems, 107(??): 924–941, June 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17311342.

Rodrigues:2022:MSA

[RAS^{+}22] Vinicius F. Rodrigues, Rodolfo S. Antunes, Lucas A. Seewald, Rodrigo Bazo, Eduardo S. dos Reis, Uelison J. L. dos Santos, Rodrigo da R. Righi, Luiz G. da S., Cristiano A. da Costa, Felipe L. Bertollo, Andreas Maier, Bjoern Eskofier, Tim Horz, Marcus Pfister, and Rebecca Fahrig. A multisensor architecture combining human pose estimation and real-time location systems for workflow monitoring on hybrid operating suites. Future Generation Computer Systems, 135(??):283–298, October 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22001716.

Rzepka:2022:SBF

[RBA^{+}22] Michał Rzepka, Piotr Boryło, Marcos D. Assunção, Artur Lasoń, and Laurent Lefèvre. SDN-based fog and cloud interplay for stream processing. Future Generation Computer Systems, 131(??):1–17, June 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22000140.

Rabhi:2021:DII

[RBLD21] Fethi A. Rabhi, Madhushi Bandara, Kun Lu, and Saif Dewan. Design of an innovative IT platform for analytics knowledge management. Future Generation Computer Systems, 116(??):209–219, March 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X2032999X.

Rodriguez-Barroso:2022:DDA

[RBMCLH22] Nuria Rodríguez-Barroso, Eugenio Martínez-Cámara, M. Victoria Luzón, and Francisco Herrera. Dynamic defense against Byzantine poisoning attacks in federated learning. Future Generation Computer Systems, 133(??):1– 9, August 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22000784.

Rajput:2023:ECV

[RBSK23] Nitin Singh Rajput, Rahul Banerjee, Dheeraj Sanghi, and Chitra Kalyansundaram. Evolution of cooperation in vehicular cloud assisted networks for ITS services: a hunt gamebased approach. Future Generation Computer Systems, 146(??):62–77, September 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23001486.

Rjoub:2020:BTA

[RBW20] Gaith Rjoub, Jamal Bentahar, and Omar Abdel Wahab. BigTrustScheduling: Trust-aware big data task scheduling approach in cloud computing environments. Future Generation Computer Systems, 110 (??):1079–1097, September 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19304054.

Rosa:2021:IAD

[RCdF^{+}21] Luis Rosa, Tiago Cruz, Miguel Borges de Freitas, Pedro Quitério, João Henriques, Filipe Caldeira, Edmundo Monteiro, and Paulo Simões. Intrusion and anomaly detection for the next-generation of industrial automation and control systems. Future Generation Computer Systems, 119(??):50–67, June 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21000431.

Ranjan:2020:NAS

[RCJZ20] Rajiv Ranjan, Lydia Y. Chen, Prem Prakash Jayaraman, and Albert Y. Zomaya. A note on advances in scheduling algorithms for cyber-physical-social workflows. Future Generation Computer Systems, 108(??):1027–1029, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19314001.

Ramon-Cortes:2020:PMH

[RCLEB20] Cristian Ramon-Cortes, Francesc Lordan, Jorge Ejarque, and Rosa M. Badia. A programming model for hybrid workflows: Combining task-based workflows and dataflows allin-one. Future Generation Computer Systems, 113(??):281–297, December 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19324823.

Raptis:2024:ETP

[RCP24] Theofanis P. Raptis, Claudio Cicconetti, and Andrea Passarella. Efficient topic partitioning of Apache Kafka for highreliability real-time data streaming applications. Future Generation Computer Systems, 154(??):173–188, May 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23004892.

Repetto:2021:AMS

[RCR21] Matteo Repetto, Alessandro Carrega, and Riccardo Rapuzzi. An architecture to manage security operations for digital service chains. Future Generation Computer Systems, 115(??):251–266, February 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20303290.

Roffarello:2023:TLD

[RD23] Alberto Monge Roffarello and Luigi De Russis. Teaching and learning “Digital wellbeing”. Future Generation Computer Systems, 149(??):494–508, December 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23003035.

Rajapackiyam:2024:ECO

[RDR^{+}24] Ezhilarasie Rajapackiyam, Anousouya Devi, Mandi Sushmanth Reddy, Umamakeswari Arumugam, Subramaniyaswamy Vairavasundaram, Indragandhi Vairavasundaram, and Vishnu Suresh. An efficient computation offloading in edge environment using genetic algorithm with directed search techniques for IoT applications. Future Generation Computer Systems, 158(??):378–390, September 2024. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24001523.

Rizzo:2020:CEI

[RFd20] Giuseppe Rizzo, Nicola Fanizzi, and Claudia d’Amato. Class expression induction as concept space exploration: From DL-Foil to DL-Focl. Future Generation Computer Systems, 108(??):256–272, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19303991.

Rodrigues-Filho:2022:PHS

[RFP22] Roberto Rodrigues-Filho and Barry Porter. Hatch: Self-distributing systems for data centers. Future Generation Computer Systems, 132(??):80–92, July 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22000516.

Russo:2024:QAO

[RFP^{+}24] Gabriele Russo Russo, Daniele Ferrarelli, Diana Pasquali, Valeria Cardellini, and Francesco Lo Presti. QoS-aware offloading policies for serverless functions in the Cloud-to-Edge continuum. Future Generation Computer Systems, 156 (??):1–15, July 2024. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24000645.

Rodriguez-Gallego:2023:CSF

[RGDMMR^{+}23] Cristina Rodríguez-Gallego, Fernando Díez-Muñoz, María-Luisa Martín-Ruiz, Ana-Marta Gabaldón, María Dolón-Poza, and Iván Pau. A collaborative semantic framework based on activities for the development of applications in Smart Home living labs. Future Generation Computer Systems, 140(??):450–465, March 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2200351X.

Rivas:2022:TAM

[RGP^{+}22] Daniel Rivas, Francesc Guim, Jordà Polo, Pubudu M. Silva, Josep Ll. Berral, and David Carrera. Towards automatic model specialization for edge video analytics. Future Generation Computer Systems, 134(??):399–413, September 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22001200.

Rodriguez-Garlito:2024:ARS

[RGPGP24] Elena C. Rodríguez-Garlito, Abel Paz-Gallardo, and Antonio Plaza. Advancements in remote sensing for invasive plant mapping along the Guadiana River: the role of CNN2D. Future Generation Computer Systems, 158(??):400–409, September 2024. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24001985.

Redondo:2020:HAL

[RGRV^{+}20] Rebeca P. Díaz Redondo, Carlos Garcia-Rubio, Ana Fernández Vilas, Celeste Campo, and Alicia Rodriguez-Carrion. A hybrid analysis of LBSN data to early detect anomalies in crowd dynamics. Future Generation Computer Systems, 109(??):83– 94, August 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19309859.

Rouland:2020:FSV

[RHJ20] Quentin Rouland, Brahim Hamid, and Jason Jaskolka. Formal specification and verification of reusable communication models for distributed systems architecture. Future Generation Computer Systems, 108(??):178–197, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19311926.

Rahman:2023:IIF

[RHK^{+}23] Anichur Rahman, Kamrul Hasan, Dipanjali Kundu, Md. Jahidul Islam, Tanoy Debnath, Shahab S. Band, and Neeraj Kumar. On the ICN-IoT with federated learning integration of communication: Concepts, security-privacy issues, applications, and future perspectives. Future Generation Computer Systems, 138(??):61–88, January 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22002667.

Ruz:2020:SAT

[RHM20] Gonzalo A. Ruz, Pablo A. Henríquez, and Aldo Mascareño. Sentiment analysis of Twitter data during critical events through Bayesian networks classifiers. Future Generation Computer Systems, 106(??):92–104, May 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19303322.

Ren:2023:BBB

[RHWY23] Yongjun Ren, Ding Huang, Wenhai Wang, and Xiaofeng Yu. BSMD: a blockchain-based secure storage mechanism for big spatio-temporal data. Future Generation Computer Systems, 138(??):328–338, January 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22002928.

Huang:2024:GDD

[rHZmH^{+}24] Yi rui Huang, Jing Zhang, Hong ming Hou, Xiu cai Ye, and Yi Chen. GeoPM-DMEIRL: a deep inverse reinforcement learning security trajectory generation framework with serverless computing. Future Generation Computer Systems, 154(??):123–139, May 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24000025.

Razaque:2022:EES

[RJA^{+}22] Abdul Razaque, Yaser Jararweh, Bandar Alotaibi, Munif Alotaibi, Salim Hariri, and Muder Almiani. Energy-efficient and secure mobile fog-based cloud for the Internet of Things. Future Generation Computer Systems, 127(??):1– 13, February 2022. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21003320.

Riley:2021:EVA

[RJM^{+}21] Ian Riley, Sharmin Jahan, Allen Marshall, Charles Walter, and Rose F. Gamble. Evaluating verification awareness as a method for assessing adaptation risk. Future Generation Computer Systems, 119(??): 110–135, June 2021. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21000443.

Ragaventhiran:2020:MOR

[RK20] J. Ragaventhiran and M. K. Kavithadevi. Map-optimize-reduce: CAN tree assisted FP-growth algorithm for clusters based FP mining on Hadoop. Future Generation Computer Systems, 103(??):111–122, February 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19315729.

Robert:2020:ELN

[RKG20] Jérémy Robert, Sylvain Kubler, and Sankalp Ghatpande. Enhanced lightning network (off-chain)- based micropayment in IoT ecosystems. Future Generation Computer Systems, 112(??):283–296, November 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19322654.

Rosciszewski:2023:OTS

[RKI^{+}23] Paweł Rościszewski, Adam Krzywaniak, Sergio Iserte, Krzysztof Rojek, and Paweł Gepner. Optimizing throughput of Seq2Seq model training on the IPU platform for AI-accelerated CFD simulations. Future Generation Computer Systems, 147(??):149–162, October 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23001784.

Reshadinezhad:2023:EAC

[RKM23] Amir Reshadinezhad, Mohammad Reza Khayyambashi, and Naser Movahedinia. An efficient adaptive cache management scheme for named data networks. Future Generation Computer Systems, 148(??):79– 92, November 2023. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23001693.

Rodrigues:2021:TCN

[RKP^{+}21] Leonardo Rosa Rodrigues, Guilherme Piêgas Koslovski, Marcelo Pasin, Maurício Aronne Pillon, Omir Correia Alves, and Charles Christian Miers. Time-constrained and network-aware containers scheduling in GPU era.

Future Generation Computer Systems, 117(??):72– 86, April 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20330387.

Roy:2022:LSI

[RLCB22] Souradip Roy, Juan Li, Bong-Jin Choi, and Yan Bai. A lightweight supervised intrusion detection mechanism for IoT networks. Future Generation Computer Systems, 127(??):276–285, February 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21003733.

Ren:2022:PPI

[RLL^{+}22] Yingying Ren, Wei Liu, Anfeng Liu, Tian Wang, and Ang Li. A privacyprotected intelligent crowdsourcing application of IoT based on the reinforcement learning. Future Generation Computer Systems, 127(??):56–69, February 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21003411.

Romano:2020:DGP

[RLML20] Diego Romano, Marco Lapegna, Valeria Mele, and Giuliano Laccetti. Designing a GPU-parallel algorithm for raw SAR data compression: a focus on parallel performance estimation. Future Generation Computer Systems, 112(??):695–708, November 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19310428.

Ren:2021:MCS

[RLQ^{+}21] Yongjun Ren, Yan Leng, Jian Qi, Pradip Kumar Sharma, Jin Wang, Zafer Almakhadmeh, and Amr Tolba. Multiple cloud storage mechanism based on blockchain in smart homes. Future Generation Computer Systems, 115(??):304–313, February 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19319041.

Ren:2021:MSG

[RLZW21] Jiajia Ren, Hongxin Li, Mengmeng Zhang, and Cong Wu. Massivescale graph mining for ecommerce cold chain analysis and optimization. Future Generation Computer Systems, 125(??):526–531, December 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21002521.

Raza:2020:ESI

[RMA^{+}20] Abdur Rehman Raza, Khawir Mahmood, Muhammad Faisal Amjad, Haider Abbas, and Mehreen Afzal. On the efficiency of software implementations of lightweight block ciphers from the perspective of programming languages. Future Generation Computer Systems, 104(??):43– 59, March 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19310519.

Rashid:2021:CFR

[RMA21] Aqsa Rashid, Asif Masood, and Haider Abbas. Cryptographic framework for role control remedy: a secure role engineering mechanism for single authority organizations. Future Generation Computer Systems, 117(??):245–258, April 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20329988.

Rejiba:2021:TUC

[RMBMT21] Zeineb Rejiba, Xavier Masip-Bruin, and Eva Marín-Tordera. Towards user-centric, switching costaware fog node selection strategies. Future Generation Computer Systems, 117(??):359–368, April 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20330685.

Rahmani:2020:MOM

[RMC20] R. Rahmani, I. Moser, and A. L. Cricenti. Modelling and optimisation of microgrid configuration for green data centres: a metaheuristic approach. Future Generation Computer Systems, 108(??):742–750, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19327359.

Rana:2022:LCI

[RMI22] Muhammad Rana, Quazi Mamun, and Rafiqul Islam. Lightweight cryptography in IoT networks: a survey. Future Generation Computer Systems, 129(??):77– 89, April 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21004404.

Rashid:2021:EME

[RNA21] Zahid Rashid, Umara Noor, and Jörn Altmann. Economic model for evaluating the value creation through information sharing within the cybersecurity information sharing ecosystem. Future Generation Computer Systems, 124(??):436–466, November 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21001837.

Rahman:2022:EET

[RNA^{+}22] Fatin Hamadah Rahman, S. H. Shah Newaz, Thien-Wan Au, Wida Susanty Suhaili, M. A. Parvez Mahmud, and Gyu Myoung Lee. EnTruVe: ENergy and TRUst-aware virtual machine allocation in VEhicle fog computing for catering applications in 5G. Future Generation Computer Systems, 126(??):196–210, January 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21002983.

Romero:2023:SHS

[RNRA23] Jose Carlos Romero, Angeles Navarro, Andrés Rodríguez, and Rafael Asenjo. SkyFlow: Heterogeneous streaming for skyline computation using FlowGraph and SYCL. Future Generation Computer Systems, 141(??):269–283, April 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2200382X.

Romero:2021:EHM

[RNV^{+}21] Jose Carlos Romero, Angeles Navarro, Antonio Vilches, Andrés Rodríguez, Francisco Corbera, and Rafael Asenjo. Efficient heterogeneous matrix profile on a CPU + high performance FPGA with integrated HBM. Future Generation Computer Systems, 125(??):10–23, December 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2100220X.

Rodriguez:2020:AUM

[RPdVR20] L. Rodríguez, J. Palanca, E. del Val, and M. Rebollo. Analyzing urban mobility paths based on users’ activity in social networks. Future Generation Computer Systems, 102(??):333–346, January 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19301864.

Ristov:2021:AAF

[RPF21] Sasko Ristov, Stefan Pedratscher, and Thomas Fahringer. AFCL: an abstract function choreography language for serverless workflow specification. Future Generation Computer Systems, 114(??):368–382, January 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20302648.

Reddy:2020:DOA

[RPP^{+}20] K. Srikanth Reddy, Lokesh Kumar Panwar, B. K. Panigrahi, Rajesh Kumar, and Yan Xu. A dual objective approach for aggregator managed demand side management (DSM) in cloud based cyber physical smart distribution system. Future Generation Computer Systems, 105(??):843–854, April 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17318320.

Rodic:2023:PLL

[RPŠŠ23] Lea Dujić Rodić, Toni Perković, Maja Škiljo, and Petar Šolić. Privacy leakage of LoRaWAN smart parking occupancy sensors. Future Generation Computer Systems, 138(??):142–159, January 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22002680.

Rao:2024:DLB

[RRAB24] A. Sai Venkateshwar Rao, Prasanta Kumar Roy, Tarachand Amgoth, and Ansuman Bhattacharya. A deep learning-based authentication protocol for IoT-enabled LTE systems. Future Generation Computer Systems, 154(??): 451–464, May 2024. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24000141.

Rausch:2021:OCS

[RRD21] Thomas Rausch, Alexander Rashed, and Schahram Dustdar. Optimized container scheduling for dataintensive serverless edge computing. Future Generation Computer Systems, 114(??):259–271, January 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X2030399X.

Rojo-Rivas:2023:KBN

[RRDSAML23] María Isabel Rojo-Rivas, Daniel Díaz-Sánchez, Florina Almenarez, and Andrés Marín-Lopez. Kriper: a blockchain network with permissioned storage. Future Generation Computer Systems, 138(??):160–171, January 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22002679.

Rosa:2021:CRC

[RRHA21] Michel J. F. Rosa, Célia Ghedini Ralha, Maristela Holanda, and Aleteia P. F. Araujo. Computational resource and cost prediction service for scientific workflows in federated clouds. Future Generation Computer Systems, 125(??):844–858, December 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21002922.

Roy:2020:DLF

[RSB20] Pradeep Kumar Roy, Jyoti Prakash Singh, and Snehasish Banerjee. Deep learning to filter SMS spam. Future Generation Computer Systems, 102(??):524–533, January 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19306879.

Raman:2020:KIE

[RSBM20] Rahul Raman, Pankaj Kumar Sa, Sambit Bakshi, and Banshidhar Majhi. Kinesiology-inspired estimation of pedestrian walk direction for smart surveillance. Future Generation Computer Systems, 108(??):1008–1026, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17312414.

Rutishauser:2023:JIE

[RSFB23] Georg Rutishauser, Moritz Scherer, Tim Fischer, and Luca Benini. 7μJ/inference end-to-end gesture recognition from dynamic vision sensor data using ternarized hybrid convolutional neural networks. Future Generation Computer Systems, 149(??):717–731, December 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23002704.

Ren:2024:HLP

[RSL24] Chenshan Ren, Wei Song, and Xinchen Lyu. Hybrid learning of predictive mobile-edge computation offloading under differentlyaged network states. Future Generation Computer Systems, 156(??):301–312, July 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X24000943.

Ramperez:2021:FCP

[RSLL21] Víctor Rampérez, Javier Soriano, David Lizcano, and Juan A. Lara. FLAS: a combination of proactive and reactive auto-scaling architecture for distributed services. Future Generation Computer Systems, 118 (??):56–72, May 2021. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20330879.

Rauf:2021:FAT

[RSQS21] Usman Rauf, Mohamed Shehab, Nafees Qamar, and Sheema Sameen. Formal approach to thwart against insider attacks: a bioinspired auto-resilient policy regulation framework.

Future Generation Computer Systems, 117(??): 412–425, April 2021. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20330338.

Rahman:2024:DTF

[RSR^{+}24] Hameedur Rahman, Uzair Muzamil Shah, Syed Morsleen Riaz, Kashif Kifayat, Syed Atif Moqurrab, and Joon Yoo. Digital twin framework for smart greenhouse management using next-gen mobile networks and machine learning. Future Generation Computer Systems, 156(??):285–300, July 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24000931.

Rico:2024:EGA

[RTD24] Noelia Rico, Luigi Troiano, and Irene Díaz. Efficient GPU-algorithms for the combination of evidence in Dempster–Shafer theory. Future Generation Computer Systems, 154(??): 465–478, May 2024. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24000098.

Ruiperez-Valiente:2021:DDD

[RVJMJ^{+}21] José A. Ruipérez-Valiente, Daniel Jaramillo-Morillo, Srećko Joksimović, Vitomir Kovanović, Pedro J. Muñoz-Merino, and Dragan Gašević. Data-driven detection and characterization of communities of accounts collaborating in MOOCs. Future Generation Computer Systems, 125(??):590–603, December 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21002570.

Ren:2021:DBS

[RWG21] Wei Ren, Xutao Wan, and Pengcheng Gan. A double-blockchain solution for agricultural sampled data security in Internet of Things network. Future Generation Computer Systems, 117(??):453–461, April 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20330697.

Ruan:2020:PPB

[RWJ^{+}20] Chang Ruan, Jianxin Wang, Wanchun Jiang, Geyong Min, and Yi Pan. PTCP: a priority-based transport control protocol for timeout mitigation in commodity data center. Future Generation Computer Systems, 102(??):619–632, January 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19302390.

Rasheed:2020:DRL

[RYL20] Faizan Rasheed, Kok-Lim Alvin Yau, and Yeh-Ching Low. Deep reinforcement learning for traffic signal control under disturbances: a case study on Sunway city, Malaysia. Future Generation Computer Systems, 109(??):431–445, August 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19328249.

Rodriguez-Zurrunero:2021:AFS

[RZA21] Roberto Rodriguez-Zurrunero and Alvaro Araujo. Adaptive frequency scaling strategy to improve energy efficiency in a tick-less operating system for resourceconstrained embedded devices. Future Generation Computer Systems, 124(??):230–242, November 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21001886.

Ren:2021:TME

[RZH21] Jianguo Ren, Chunming Zhang, and Qihong Hao. A theoretical method to evaluate honeynet potency. Future Generation Computer Systems, 116(??):76– 85, March 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20308219.

Razzak:2020:RNO

[RZIX20] Imran Razzak, Khurram Zafar, Muhammad Imran, and Guandong Xu. Randomized nonlinear one-class support vector machines with bounded loss function to detect of outliers for large scale IoT data. Future Generation Computer Systems, 112(??):715–723, November 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19313913.

Shokry:2022:SSA

[SAAEK22] Mostafa Shokry, Ali Ismail Awad, Mahmoud Khaled Abd-Ellah, and Ashraf A. M. Khalaf. Systematic survey of advanced metering infrastructure security: Vulnerabilities, attacks, countermeasures, and future vision. Future Generation Computer Systems, 136(??):358–377, November 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22002254.

Sanchez-Anguix:2021:SIA

[SACN^{+}21] Victor Sanchez-Anguix, Kuo-Ming Chao, Paulo Novais, Olivier Boissier, and Vicente Julian. Social and intelligent applications for future cities: Current advances. Future Generation Computer Systems, 114(??):181–184, January 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20324006.

Su:2023:CEC

[SACW23] Xin Su, Li An, Zhen Cheng, and Yajuan Weng. Cloud– edge collaboration-based bi-level optimal scheduling for intelligent healthcare systems. Future Generation Computer Systems, 141 (??):28–39, April 2023. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22003661.

Sahin:2024:LCH

[SAD24] Emre Sahin, Mustafa Alper Akkas, and Orhan Dagdeviren. Low-cost and highperformance channel access strategies for Internet of Nano-Things applications. Future Generation Computer Systems, 153(??): 477–491, April 2024. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X2300479X.

Subratie:2023:ESO

[SAF23] Kensworth Subratie, Saumitra Aditya, and Renato J. Figueiredo. EdgeVPN: Self-organizing layer-2 virtual edge networks. Future Generation Computer Systems, 140(??):104–116, March 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22003235.

Sakhri:2024:DTB

[SAM^{+}24] Aya Sakhri, Arsalan Ahmed, Moufida Maimour, Mehdi Kherbache, Eric Rondeau, and Noureddine Doghmane. A digital twin-based energy-efficient wireless multimedia sensor network for waterbirds monitoring. Future Generation Computer Systems, 155(??):

146–163, June 2024. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24000566.

Savva:2020:ALA

[SAT20] Fotis Savva, Christos Anagnostopoulos, and Peter Triantafillou. Adaptive learning of aggregate analytics under dynamic workloads. Future Generation Computer Systems, 109(??): 317–330, August 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19329504.

Shahraki:2024:EME

[SB24] Mahdi Shahraki and Amir Jalaly Bidgoly. Edge model: an efficient method to identify and reduce the effectiveness of malicious clients in federated learning. Future Generation Computer Systems, 157(??):459–468, August 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24001201.

Schintke:2024:VCD

[SBD^{+}24] Florian Schintke, Khalid Belhajjame, Ninon De Mecquenem, David Frantz, Vanessa Emanuela Guarino, Marcus Hilbrich, Fabian Lehmann, Paolo Missier, Rebecca Sattler, Jan Arne Sparka, Daniel T. Speckhard, Hermann Stolte, Anh Duc Vu, and Ulf Leser. Validity constraints for data analysis workflows. Future Generation Computer Systems, 157(??):82– 97, August 2024. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24001079.

Shakeel:2021:ITF

[SBF^{+}21] P. Mohamed Shakeel, S. Baskar, Hassan Fouad, Gunasekaran Manogaran, Vijayalakshmi Saravanan, and Carlos Enrique Montenegro-Marin. Internet of things forensic data analysis using machine learning to identify roots of data scavenging. Future Generation Computer Systems, 115(??):756–768, February 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20329782.

Shahzaad:2021:RCD

[SBMN21] Babar Shahzaad, Athman Bouguettaya, Sajib Mistry, and Azadeh Ghari Neiat. Resilient composition of drone services for delivery. Future Generation Computer Systems, 115(??):335–350, February 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19331735.

Saprykin:2022:AAB

[SCA22] Aleksandr Saprykin, Ndaona Chokani, and Reza S. Abhari. Accelerating agentbased demand-responsive transport simulations with GPUs. Future Generation Computer Systems, 131 (??):43–58, June 2022. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22000176.

Sancho:2020:NAT

[SCÁB20] José Carlos Sancho, Andrés Caro, Mar Ávila, and Alberto Bravo. New approach for threat classification and security risk estimations based on security event management. Future Generation Computer Systems, 113(??):488–505, December 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20301849.

Sanchez:2024:AAD

[SCBP24] Pedro Miguel Sánchez Sánchez, Alberto Huertas Celdrán, Gérôme Bovet, and Gregorio Martínez Pérez. Adversarial attacks and defenses on ML- and hardware-based IoT device fingerprinting and identification. Future Generation Computer Systems, 152(??):30–42, March 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23003874.

Steer:2020:RSA

[SCC20] Benjamin Steer, Felix Cuadrado, and Richard Clegg. Raphtory: Streaming analysis of distributed temporal graphs. Future Generation Computer Systems, 102(??):453–464, January 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19301621.

Sarfaraz:2023:AAC

[SCE23] Aaliya Sarfaraz, Ripon K. Chakrabortty, and Daryl L. Essam. AccessChain: an access control framework to protect data access in blockchain enabled supply chain. Future Generation Computer Systems, 148(??):380–394, November 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23002303.

Sacoto-Cabrera:2020:EFV

[SCGVP20] Erwin J. Sacoto-Cabrera, Luis Guijarro, Jose R. Vidal, and Vicent Pla. Economic feasibility of virtual operators in 5G via network slicing. Future Generation Computer Systems, 109(??):172–187, August 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19325579.

Sun:2022:OOB

[SCK^{+}22] Xue Sun, Ping Chou, Chorng-Shiuh Koong, Chao-Chin Wu, and Liang-Rui Chen. Optimizing 2-opt-based heuristics on GPU for solving the single-row facility layout problem. Future Generation Computer Systems, 126(??):91–109, January 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21002843.

Shah:2020:HMV

[SCL20] Faaiz Shah, Arnaud Castelltort, and Anne Laurent. Handling missing values for mining gradual patterns from NoSQL graph databases. Future Generation Computer Systems, 111(??):523–538, October 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18324002.

Silva:2021:ATD

[SCP^{+}21] João A. Silva, Filipe Cerqueira, Hervé Paulino, João M. Lourenço, João Leitão, and Nuno Preguiça. It’s about thyme: On the design and implementation of a time-aware reactive storage system for pervasive edge computing environments. Future Generation Computer Systems, 118 (??):14–36, May 2021. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20330703.

Sallam:2020:TSM

[SCR20] Karam M. Sallam, Ripon K. Chakrabortty, and Michael J. Ryan. A two-stage multi-operator differential evolution algorithm for solving Resource Constrained Project Scheduling problems. Future Generation Computer Systems, 108(??):432–444, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19317480.

Sun:2022:EER

[SCW^{+}22] Dawei Sun, Yijing Cui, Minghui Wu, Shang Gao, and Rajkumar Buyya. An energy efficient and runtime-aware framework for distributed stream computing systems. Future Generation Computer Systems, 136(??):252–269, November 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22002187.

Shu:2021:RSA

[SCX21] Wanneng Shu, Ken Cai, and Neal Naixue Xiong. Research on strong agile response task scheduling optimization enhancement with optimal resource usage in green cloud computing. Future Generation Computer Systems, 124(??):12– 20, November 2021. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2100162X.

Sanchez:2024:FST

[SCX^{+}24] Pedro Miguel Sánchez Sánchez, Alberto Huertas Celdrán, Ning Xie, Gérôme Bovet, Gregorio Martínez Pérez, and Burkhard Stiller. FederatedTrust: a solution for trustworthy federated learning. Future Generation Computer Systems, 152(??):83–98, March 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23003886.

Shao:2023:VIC

[SCXZ23] Caixing Shao, Fengxin Cheng, Jingzhong Xiao, and Ke Zhang. Vehicular intelligent collaborative intersection driving decision algorithm in Internet of Vehicles. Future Generation Computer Systems, 145(??):384–395, August 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23001231.

Shen:2020:CBM

[SCZ^{+}20] Meng Shen, Guohua Cheng, Liehuang Zhu, Xiaojiang Du, and Jiankun Hu. Content-based multi-source encrypted image retrieval in clouds with privacy preservation. Future Generation Computer Systems, 109(??):621–632, August 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17321969.

Sahu:2020:KTD

[SD20] Ashish Kumar Sahu and Pragya Dwivedi. Knowledge transfer by domainindependent user latent factor for cross-domain recommender systems. Future Generation Computer Systems, 108(??):320–333, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19316176.

Sinha:2022:RAC

[SD22] Bam Bahadur Sinha and R. Dhanalakshmi. Recent advancements and challenges of Internet of Things in smart agriculture: a survey. Future Generation Computer Systems, 126(??):169–184, January 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21003113.

S:2024:UIC

[SD24] Asha S. and Shanmugapriya D. Understanding insiders in cloud adopted organizations: a survey on taxonomies, incident analysis, defensive solutions, challenges. Future Generation Computer Systems, 158(??):427–446, September 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24001614.

Srirama:2021:AFB

[SDA21] Satish Narayana Srirama, Freddy Marcelo Surriabre Dick, and Mainak Adhikari. Akka framework based on the actor model for executing distributed fog computing applications. Future Generation Computer Systems, 117(??):439–452, April 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20330739.

Santiago-Duran:2020:GMP

[SDGCB^{+}20] Miguel Santiago-Duran, J. L. Gonzalez-Compean, André Brinkmann, Hugo G. Reyes-Anastacio, Jesus Carretero, Raffaele Montella, and Gregorio Toscano Pulido. A gearbox model for processing large volumes of data by using pipeline systems encapsulated into virtual containers. Future Generation Computer Systems, 106(??):304–319, May 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19316334.

Sotenga:2020:IIL

[SDKM20] Prosper Zanu Sotenga, Karim Djouani, Anish Matthew Kurien, and Martin Mwila. Implementation of an indoor localisation algorithm for Internet of Things. Future Generation Computer Systems, 107(??): 1037–1046, June 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17320307.

Steenwinckel:2021:FMA

[SDV^{+}21] Bram Steenwinckel, Dieter De Paepe, Sander Vanden Hautte, Pieter Heyvaert, Mohamed Bentefrit, Pieter Moens, Anastasia Dimou, Bruno Van Den Bossche, Filip De Turck, Sofie Van Hoecke, and Femke Ongenae. FLAGS: a methodology for adaptive anomaly detection and root cause analysis on sensor data streams by fusing expert knowledge with machine learning. Future Generation Computer Systems, 116(??):30–48, March 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20329927.

Sgueglia:2022:SLR

[SDVC22] Arnaldo Sgueglia, Andrea Di Sorbo, Corrado Aaron Visaggio, and Gerardo Canfora. A systematic literature review of IoT time series anomaly detection solutions. Future Generation Computer Systems, 134(??):170–186, September 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22001285.

Shi:2020:AAD

[SDZ^{+}20] Yuliang Shi, Mianxiong Dong, Wenbin Zhang, Lei Liu, Yongqing Zheng, Lizhen Cui, and Junhua Zhang. AdaptScale: an adaptive data scaling controller for improving the multiple performance requirements in clouds. Future Generation Computer Systems, 105(??):814–823, April 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17318563.

Seeling:2020:WRH

[See20] Patrick Seeling. WWW retrieval handling optimization w_{ρ}^{3}: a metric for webpage timeout setting performance evaluation and comparison. Future Generation Computer Systems, 110 (??):1055–1066, September 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19310349.

Serhani:2020:SAC

[SEKS^{+}20] M. Adel Serhani, Hadeel T.

El-Kassabi, Khaled Shuaib, Alramzana N. Navaz, Boualem Benatallah, and Amine Beheshti. Self-adapting cloud services orchestration for fulfilling intensive sensory data-driven IoT workflows. Future Generation Computer Systems, 108(??): 583–597, July 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19316231.

Santos:2022:RLB

[SEL^{+}22] Guto Leoni Santos, Patricia Takako Endo, Theo Lynn, Djamel Sadok, and Judith Kelner. A reinforcement learning-based approach for availabilityaware service function chain placement in largescale networks. Future Generation Computer Systems, 136(??):93–109, November 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22001923.

Souza:2023:EPB

[SFC23] Paulo S. Souza, Tiago Ferreto, and Rodrigo N. Calheiros. EdgeSimPy: Python-based modeling and simulation of edge computing resource management policies. Future Generation Computer Systems, 148(??):446–459, November 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23002340.

S:2020:ADL

[SG20] Ashwin T. S. and Ram Mohana Reddy Guddeti. Affective database for elearning and classroom environments using Indian students’ faces, hand gestures and body postures. Future Generation Computer Systems, 108(??): 334–348, July 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19314992.

Serrano:2020:AII

[SGBC^{+}20] Estefania Serrano, Javier Garcia-Blas, Jesus Carretero, Manuel Desco, and Monica Abella. Accelerated iterative image reconstruction for cone-beam computed tomography through Big Data frameworks. Future Generation Computer Systems, 106(??):534–544, May 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19323489.

Su:2023:UCB

[SGDG23] Yuhan Su, Zhibin Gao, Xiaojiang Du, and Mohsen Guizani. User-centric base station clustering and resource allocation for celledge users in 6G ultradense networks. Future Generation Computer Systems, 141(??):173–185, April 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22003727.

Sanchez-Gallegos:2021:EPB

[SGDK^{+}21] Dante Domizzi Sánchez-Gallegos, Diana Di Luccio, Sokol Kosta, J. L. Gonzalez-Compean, and Raffaele Montella. An efficient pattern-based approach for workflow supporting large-scale science: the DagOnStar experience. Future Generation Computer Systems, 122(??):187–203, September 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21000984.

Shi:2020:WGC

[SGL^{+}20a] Kaize Shi, Changjin Gong, Hao Lu, Yifan Zhu, and Zhendong Niu. Widegrained capsule network with sentence-level feature to detect meteorological event in social network. Future Generation Computer Systems, 102(??):323–332, January 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19310908.

Sun:2020:DRR

[SGL^{+}20b] Dawei Sun, Shang Gao, Xunyun Liu, Xindong You, and Rajkumar Buyya. Dynamic redirection of realtime data streams for elastic stream computing. Future Generation Computer Systems, 112(??):193–208, November 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19323477.

Sun:2022:MLC

[SGLB22] Dawei Sun, Shang Gao, Xunyun Liu, and Rajkumar Buyya. A multilevel collaborative framework for elastic stream computing systems. Future Generation Computer Systems, 128(??):117–131, March 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21003952.

Santipantakis:2020:SSI

[SGP^{+}20a] Georgios M. Santipantakis, Apostolos Glenis, Kostas Patroumpas, Akrivi Vlachou, Christos Doulkeridis, George A. Vouros, Nikos Pelekis, and Yannis Theodoridis. SPAR-TAN: Semantic integration of big spatio-temporal data from streaming and archival sources. Future Generation Computer Systems, 110(??):540–555, September 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17320319.

Savaglio:2020:ABI

[SGP^{+}20b] Claudio Savaglio, Maria Ganzha, Marcin Paprzycki, Costin Bădică, Mirjana Ivanović, and Giancarlo Fortino. Agentbased Internet of Things: State-of-the-art and research challenges. Future Generation Computer Systems, 102(??):1038–1053, January 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19312282.

Sierra-Garcia:2024:FDR

[SGS24] J. Enrique Sierra-Garcia and Matilde Santos. Federated discrete reinforcement learning for automatic guided vehicle control. Future Generation Computer Systems, 150(??):78–89, January 2024. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23003217.

Sanchez-Gallegos:2023:BES

[SGSGGC^{+}23] Genaro Sanchez-Gallegos, Dante D. Sanchez-Gallegos, J. L. Gonzalez-Compean, Hugo G. Reyes-Anastacio, and Jesus Carretero. On the building of efficient selfadaptable health data science services by using dynamic patterns. Future Generation Computer Systems, 145(??):478–495, August 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23001243.

Shalaginov:2020:BDA

[Sha20] Andrii Shalaginov. Big data analytics and artificial intelligence for cyber crime investigation and prevention. Future Generation Computer Systems, 109(??):702–703, August 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20311432.

Sellami:2022:DRL

[SHB22] Bassem Sellami, Akram Hakiri, and Sadok Ben Yahia. Deep Reinforcement Learning for energyaware task offloading in join SDN-Blockchain 5G massive IoT edge network. Future Generation Computer Systems, 137(??):363–379, December 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22002588.

Sestak:2021:AVC

[ŠHDT21] Martina Šestak, Marjan Heričko, Tatjana Welzer Družovec, and Muhamed Turkanović. Applying k-vertex cardinality constraints on a Neo4j graph database. Future Generation Computer Systems, 115(??):459–474, February 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19324094.

Soto:2023:JFH

[SHF23] Javier E. Soto, Cecilia Hernández, and Miguel Figueroa. JACC-FPGA: a hardware accelerator for Jaccard similarity estimation using FPGAs in the cloud. Future Generation Computer Systems, 138(??):26–42, January 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22002643.

Spisakova:2023:NBE

[SHH23] Viktória Spišaková, Lukáš Hejtmánek, and Jakub Hynšt. Nextflow in bioinformatics: Executors performance comparison using genomics data. Future Generation Computer Systems, 142(??):328–339, May 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2300016X.

Song:2023:ECA

[SHKW23] Ge Song, Seong Hyeon Hong, Tristan Kyzer, and Yi Wang. Energy consumption auditing based on a generative adversarial network for anomaly detection of robotic manipulators. Future Generation Computer Systems, 149(??):376–389, December 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2300287X.

Shahryari:2020:SBF

[SHST20] Shirzad Shahryari, Seyed-Amin Hosseini-Seno, and Farzad Tashtarian. An SDN based framework for maximizing throughput and balanced load distribution in a cloudlet network. Future Generation Computer Systems, 110(??):18– 32, September 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19322447.

Said:2021:DDG

[SHT^{+}21] Anwar Said, Saeed-Ul Hassan, Suppawong Tuarob, Raheel Nawaz, and Mudassir Shabbir. DGSD: Distributed graph representation via graph statistical properties. Future Generation Computer Systems, 119(??):166–175, June 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21000571.

Shang:2024:IDP

[SHW24] Shuo Shang, Bingsheng He, and Lizhe Wang. Introduction to distributed and parallel processing of big spatiotemporal data. Future Generation Computer Systems, 151(??):98–99, February 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23003588.

Sun:2021:LSU

[SHY^{+}21] Dawei Sun, Hanyu He, Hongbin Yan, Shang Gao, Xunyun Liu, and Xinqi Zheng. Lr-Stream: Using latency and resource aware scheduling to improve latency and throughput for streaming applications. Future Generation Computer Systems, 114(??):243–258, January 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19334338.

Salinas-Hilburg:2021:FEE

[SHZMA21] Juan Carlos Salinas-Hilburg, Marina Zapater, José M. Moya, and José L. Ayala. Fast energy estimation framework for longrunning applications. Future Generation Computer Systems, 115(??):20–33, February 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20305380.

Sellami:2024:VDI

[SIG24] Youssef Sellami, Youcef Imine, and Antoine Gallais. A verifiable data integrity scheme for distributed data sharing in fog computing architecture. Future Generation Computer Systems, 150(??):64– 77, January 2024. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23003163.

Saxena:2020:INH

[SJD^{+}20] Amit Saxena, Rajendra Joshi, Hemant Darbari, Cezary Mazurek, Amar Bhat, Anil Srivastava, Kevin Wojkovich, Ken Buetow, and Sandra Gesing.

The ICTBioMed NCIP Hub: Cancer research in a science gateway consortium. Future Generation Computer Systems, 105 (??):27–32, April 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18314687.

Sharma:2020:OCB

[SJQ20] Vishal Sharma, Dushantha Nalin K. Jayakody, and Marwa Qaraqe. Osmotic computing-based service migration and resource scheduling in Mobile Augmented Reality Networks (MARN). Future Generation Computer Systems, 102(??):723–737, January 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19310738.

Santander-Jimenez:2022:EML

[SJVRS22] Sergio Santander-Jiménez, Miguel A. Vega-Rodríguez, and Leonel Sousa. Exploiting multi-level parallel metaheuristics and heterogeneous computing to boost phylogenetics. Future Generation Computer Systems, 127(??):208–224, February 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21003575.

Shahid:2020:SDS

[SK20a] Furqan Shahid and Abid Khan. Smart Digital Signatures (SDS): a post-quantum digital signature scheme for distributed ledgers. Future Generation Computer Systems, 111(??):241–253, October 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19319892.

Shoukourian:2020:FPE

[SK20b] Hayk Shoukourian and Dieter Kranzlmüller. Forecasting power-efficiency related key performance indicators for modern data centers using LSTMs. Future Generation Computer Systems, 112(??):362–382, November 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20303964.

Skandylas:2021:DIS

[SK21a] Charilaos Skandylas and Narges Khakpour. Design and implementation of self-protecting systems: a formal approach. Future Generation Computer Systems, 115(??):421–437, February 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19318643.

Sudhakar:2021:MCM

[SK21b] Sudhakar and Sushil Kumar. MCFT–CNN: Malware classification with fine-tune convolution neural networks using traditional and transfer learning in Internet of Things. Future Generation Computer Systems, 125(??):334–351, December 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21002247.

Singh:2023:EES

[SK23] Sarbjeet Singh and Dilip Kumar. Energy-efficient secure data fusion scheme for IoT based healthcare system. Future Generation Computer Systems, 143 (??):15–29, June 2023. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22004447.

Sajid:2020:SCR

[SKA^{+}20] Adnan Sajid, Bilal Khalid, Mudassar Ali, Shahid Mumtaz, Usman Masud, and Farhan Qamar. Securing cognitive radio networks using blockchains. Future Generation Computer Systems, 108(??): 816–826, July 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19315997.

Saharan:2020:ESP

[SKB20] Sandeep Saharan, Neeraj Kumar, and Seema Bawa. An efficient smart parking pricing system for smart city environment: a machine-learning based approach. Future Generation Computer Systems, 106(??):622–640, May 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167739X19322496.

Silva:2020:IBD

[SKH20] Bhagya Nathali Silva, Murad Khan, and Kijun Han. Integration of Big Data analytics embedded smart city architecture with RESTful web of things for efficient service provision and energy management. Future Generation Computer Systems, 107(??): 975–987, June 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17305174.

Skinderowicz:2020:IGB

[Ski20] Rafał Skinderowicz. Implementing a GPU-based parallel MAX–MIN ant system. Future Generation Computer Systems, 106 (??):277–295, May 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0167739X19309550.

Saini:2022:DER

[SKS22] Kanika Saini, Sheetal Kalra, and Sandeep K. Sood. Disaster emergency response framework for smart buildings. Future Generation Computer Systems, 131(??):106–120, June 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22000231.

Singh:2024:GBD

[SKTP24] Sushil Kumar Singh, Manish Kumar, Sudeep Tanwar, and Jong Hyuk Park. GRUbased digital twin framework for data allocation and storage in IoT-enabled smart home networks. Future Generation Computer Systems, 153(??):391–402, April 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23004697.

Sun:2020:VVI

[SKX^{+}20] Yuanxin Sun, Ruoxin Kuai, Sa Xiao, Wanbin Tang, and Xiaoping Li. VI-MAC: Vehicular information medium access control protocol for high reliable and low latency transmissions for vehicular ad hoc networks in smart city. Future Generation Computer Systems, 106(??):55– 66, May 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19323647.

Szustak:2023:POP

[SLA^{+}23] Lukasz Szustak, Marcin Lawenda, Sebastian Arming, Gregor Bankhamer, Christoph Schweimer, and Robert Elsässer. Profiling and optimization of Python-based social sciences applications on HPC systems by means of task and data parallelism. Future Generation Computer Systems, 148(??):623–635, November 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23002571.

Salah:2020:ACG

[SLH^{+}20] Ahmad Salah, Kenli Li, Khalid M. Hosny, Mohamed M. Darwish, and Qi Tian. Accelerated CPU-GPUs implementations for quaternion polar harmonic transform of color images. Future Generation Computer Systems, 107(??): 368–382, June 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19302171.

Smahi:2024:VCB

[SLH^{+}24] Abla Smahi, Hui Li, Wang Han, Ahmed Ameen Fateh, and Ching Chuen Chan. VFL-Chain: Bulletproofing federated learning in the V2X environments. Future Generation Computer Systems, 155(??):419–436, June 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X24000578.

Shen:2020:ECA

[SLS^{+}20] Jian Shen, Dengzhi Liu, Xingming Sun, Fushan Wei, and Yang Xiang. Efficient cloud-aided verifiable secret sharing scheme with batch verification for smart cities. Future Generation Computer Systems, 109(??):

450–456, August 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17318629.

Sun:2024:TPP

[SLX^{+}24] Zihui Sun, Anfeng Liu, Neal N. Xiong, Qian He, and Shaobo Zhang. A trust and privacy-preserving intelligent big data collection scheme in mobile edgecloud crowdsourcing. Future Generation Computer Systems, 157(??):145–163, August 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24001043.

Silvano:2020:ITC

[SM20] Wellington Fernandes Silvano and Roderval Marcelino. Iota Tangle: a cryptocurrency to communicate Internet-of-Things data. Future Generation Computer Systems, 112(??):307–319, November 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19329048.

Si-Mohammed:2024:NSI

[SMBB^{+}24] Samir Si-Mohammed, Anthony Bardou, Thomas Begin, Isabelle Guérin Lassous, and Pascale Vicat-Blanc. NS+NDT: Smart integration of Network Simulation in Network Digital Twin, application to IoT networks. Future Generation Computer Systems, 157(??):124–144, August 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24001080.

Singh:2020:IEC

[SMC^{+}20] R. Raja Singh, Yash S. M., Shubham S. C., Indragandhi V., Vijayakumar V., Saravanan P., and Subramaniyaswamy V. IoT embedded cloud-based intelligent power quality monitoring system for industrial drive application. Future Generation Computer Systems, 112(??):884–898, November 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19333795.

Sadeghiram:2023:MOD

[SMC23] Soheila Sadeghiram, Hui Ma, and Gang Chen. Multiobjective distributed Web service composition — a link-dominance driven evolutionary approach. Future Generation Computer Systems, 143(??):163–178, June 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23000018.

Sahay:2019:CID

[SME^{+}19] Rishikesh Sahay, Weizhi Meng, D. A. Sepulveda Estay, Christian D. Jensen, and Michael Bruhn Barfod. CyberShip-IoT: a dynamic and adaptive SDN-based security policy enforcement framework for ships. Future Generation Computer Systems, 100(??):736–750, November 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1930367X.

See corrigendum [SME^{+}21].

Sahay:2021:CCI

[SME^{+}21] Rishikesh Sahay, Weizhi Meng, D. A. Sepulveda Estay, Christian D. Jensen, and Michael Bruhn Barfod. Corrigendum to “CyberShip-IoT: a Dynamic and Adaptive SDN-Based Security Policy Enforcement Framework for Ships” [Future Gener. Comput. Syst. 100 (2019) 736– 750]. Future Generation Computer Systems, 118 (??):492–494, May 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0167739X20326674. See

[SME^{+}19].

Sanam:2023:SDF

[SMKA23] Kahkishan Sanam, Saif Ur Rehman Malik, Tehsin Kanwal, and Zain Ul Islam Adil. SecurePrivChain: a decentralized framework for securing the global model using cryptography. Future Generation Computer Systems, 142(??):364–375, May 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22004368.

Saenger:2020:VNC

[SMKC20] Jens Saenger, Wojciech Mazurczyk, Jörg Keller, and Luca Caviglione. VoIP network covert channels to enhance privacy and information sharing. Future Generation Computer Systems, 111(??):96–106, October 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19333965.

Sami:2024:LTC

[SMO^{+}24] Hani Sami, Rabeb Mizouni, Hadi Otrok, Shakti Singh, Jamal Bentahar, and Azzam Mourad. LearnChain: Transparent and cooperative reinforcement learning on Blockchain. Future Generation Computer Systems, 150(??):255–271, January 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23003370.

Sharif:2022:ACC

[SMS22] Samane Sharif, Mohammad Hossein Yaghmaee Moghaddam, and Seyed Amin Hosseini Seno. Adaptive cache content placement for software-defined Internet of Things. Future Generation Computer Systems, 136(??):34–48, November 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2200190X.

Selvarajan:2024:PPU

[SMS^{+}24] Shitharth Selvarajan, Hariprasath Manoharan, Achyut Shankar, Alaa O. Khadidos, Adil O. Khadidos, and Antonino Galletta. PUDT: Plummeting uncertainties in digital twins for aerospace applications using deep learning algorithms. Future Generation Computer Systems, 153(??):575–586, April 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X2300448X.

Sciullo:2024:RDT

[SMT^{+}24] Luca Sciullo, Alberto De Marchi, Angelo Trotta, Federico Montori, Luciano Bononi, and Marco Di Felice. Relativistic digital twin: Bringing the IoT to the future. Future Generation Computer Systems, 153(??):521–536, April 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23004788.

Sadiq:2021:ADT

[SMU^{+}21] Saima Sadiq, Arif Mehmood, Saleem Ullah, Maqsood Ahmad, Gyu Sang Choi, and Byung-Won On. Aggression detection through deep neural model on Twitter. Future Generation Computer Systems, 114(??):120–129, January 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19330717.

Shakshuki:2020:ASN

[SMY20] Elhadi M. Shakshuki, Haroon Malik, and Ansar-Ul-Haque Yasar. Ambient systems, networks and technologies. Future Generation Computer Systems, 107(??):1011–1013, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19317510.

Simpson:2021:FBC

[SN21] Serin V. Simpson and G. Nagarajan. A fuzzy based Co-Operative Blackmailing Attack detection scheme for edge computing nodes in MANET–IOT environment. Future Generation Computer Systems, 125(??):544–563, December 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21002478.

Sadeghi-Niaraki:2023:ITI

[SN23] Abolghasem Sadeghi-Niaraki.

Internet of Things (IoT) review of review: Bibliometric overview since its foundation. Future Generation Computer Systems, 143(??):361–377, June 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23000237.

Sajjad:2020:RPA

[SNM^{+}20] Muhammad Sajjad, Mansoor Nasir, Khan Muhammad, Siraj Khan, Zahoor Jan, Arun Kumar Sangaiah, Mohamed Elhoseny, and Sung Wook Baik. Raspberry Pi assisted face recognition framework for enhanced law-enforcement services in smart cities. Future Generation Computer Systems, 108(??):995–1007, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17309512.

Sanchez-Navarro:2021:RTT

[SNMWC21] Ignacio Sanchez-Navarro, Ana Serrano Mamolar, Qi Wang, and Jose M. Alcaraz Calero. 5GTopoNet: Real-time topology discovery and management on 5G multi-tenant networks. Future Generation Computer Systems, 114(??):435–447, January 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1933122X.

Silva:2020:ADD

[SNS^{+}20] Vítor Silva, Leonardo Neves, Renan Souza, Alvaro L. G. A. Coutinho, Daniel de Oliveira, and Marta Mattoso. Adding domain data to code profiling tools to debug workflow parallel execution. Future Generation Computer Systems, 110(??):422–439, September 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X16308275.

St-Onge:2020:DTS

[SOKW^{+}20] Cédric St-Onge, Nadjia Kara, Omar Abdel Wahab, Claes Edstrom, and Yves Lemieux. Detection of time series patterns and periodicity of cloud computing workloads. Future Generation Computer Systems, 109(??):249–261, August 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19312440.

Sasak-Okon:2024:ISQ

[SOT24] Anna Sasak-Okoń and Marek Tudruj. Improving speculative query execution support by the use of the hypergraph representation. Future Generation Computer Systems, 150(??):186–201, January 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23002893.

Semiz:2021:IMA

[SP21] Fatih Semiz and Faruk Polat. Incremental multiagent path finding. Future Generation Computer Systems, 116(??):220–233, March 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19331784.

Shaw:2022:CSB

[SP22] Rabi Shaw and Bidyut Kr. Patra. Classifying students based on cognitive state in flipped learning pedagogy. Future Generation Computer Systems, 126(??):305–317, January 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21003265.

Sethi:2023:FFD

[SP23] Vivek Sethi and Sujata Pal. FedDOVe: a federated deep Q-learning-based offloading for vehicular fog computing. Future Generation Computer Systems, 141(??):96–105, April 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22003739.

Seo:2024:SSC

[SP24] Jungwon Seo and Sooyong Park. SBAC: Substitution cipher access control based on blockchain for protecting personal data in metaverse. Future Generation Computer Systems, 151(??):85– 97, February 2024. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23003552.

Song:2022:OAO

[SPL22] Xurui Song, Li Pan, and Shijun Liu. An online algorithm for optimally releasing multiple on-demand instances in IaaS clouds. Future Generation Computer Systems, 136(??):311–321, November 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22002242.

Song:2024:LBA

[SPL24] Shihao Song, Li Pan, and Shijun Liu. A Qlearning based auto-scaling approach for provisioning big data analysis services in cloud environments. Future Generation Computer Systems, 154(??):140–150, May 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X24000049.

Singh:2021:MPP

[SPRA21] Alok Singh, Shweta Purawat, Arvind Rao, and Ilkay Altintas. Modular performance prediction for scientific workflows using machine learning. Future Generation Computer Systems, 114(??):1–14, January 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18308355.

Shen:2023:MLM

[SPWL23] Yuanyuan Shen, Manman Peng, Qiang Wu, and Renfa Li. A machine learning method to variable classification in OpenMP. Future Generation Computer Systems, 140(??):67– 78, March 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22003259.

Shen:2021:TPD

[SPWW21] Yuanyuan Shen, Manman Peng, Shiling Wang, and Qiang Wu. Towards parallelism detection of sequential programs with graph neural network. Future Generation Computer Systems, 125(??):515–525, December 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21002557.

Sheng:2022:AND

[SQ22] Hu Sheng and Xiaodong Qi. Application of new digital signal processing technology based on distributed cloud computing in electronic information engineering. Future Generation Computer Systems, 128(??):443–450, March 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21004283.

Singh:2022:FPP

[SRA^{+}22] Saurabh Singh, Shailendra Rathore, Osama Alfarraj, Amr Tolba, and Byungun Yoon. A framework for privacy-preservation of IoT healthcare data using federated learning and blockchain technology. Future Generation Computer Systems, 129(??):380–388, April 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21004726.

Stabili:2023:MDS

[SRM^{+}23] Dario Stabili, Raffaele Romagnoli, Mirco Marchetti, Bruno Sinopoli, and Michele Colajanni. A multidisciplinary detection system for cyber attacks on powertrain cyber physical systems. Future Generation Computer Systems, 144(??):151–164, July 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23000602.

Singh:2020:BBE

[SRP20] Sushil Kumar Singh, Shailendra Rathore, and Jong Hyuk Park. BlockIoTIntelligence: a blockchain-enabled intelligent IoT architecture with artificial intelligence. Future Generation Computer Systems, 110(??):721–743, September 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19316474.

Singhal:2021:FBC

[SS21] Ritu Singhal and Archana Singhal. A feedback-based combinatorial fair economical double auction resource allocation model for cloud computing. Future Generation Computer Systems, 115(??):780–797, February 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19333539.

Singh:2022:OSA

[SS22] Garima Singh and Anil Kumar Singh. Optimization of SLA aware live migration of multiple virtual machines using Lagrange multiplier. Future Generation Computer Systems, 130(??): 279–291, May 2022. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22000073.

Singh:2023:CSC

[SSA^{+}23] Roop Singh, Mukesh Saraswat, Alaknanda Ashok, Himanshu Mittal, Ashish Tripathi, Avinash Chandra Pandey, and Raju Pal. From classical to soft computing based watermarking techniques: a comprehensive review. Future Generation Computer Systems, 141(??):738–754, April 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22004204.

Starzec:2020:DDA

[SSB^{+}20] Mateusz Starzec, Grazyna Starzec, Aleksander Byrski, Wojciech Turek, and Kamil Pietak. Desynchronization in distributed ant colony optimization in HPC environment. Future Generation Computer Systems, 109(??):125–133, August 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19327402.

Souza:2020:DRS

[SSC^{+}20] Renan Souza, Vítor Silva, Alvaro L. G. A. Coutinho, Patrick Valduriez, and Marta Mattoso. Data reduction in scientific workflows using provenance monitoring and user steering. Future Generation Computer Systems, 110(??):481–501, September 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X16308238.

Sole:2022:TDS

[SSDC22] Leonardo Solé, Matteo Sammarco, Marcin Detyniecki, and Miguel Elias M. Campista. Towards drivers’ safety with multi-criteria car navigation systems. Future Generation Computer Systems, 135(??):1–9, October 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22001480.

Stefano:2022:IQT

[SSM22] Alessandro Di Stefano, Antonella Di Stefano, and Giovanni Morana. Improving QoS through network isolation in PaaS. Future Generation Computer Systems, 131(??):91–105, June 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22000188.

Silveira:2021:PUE

[SSMdS21] Bárbara Silveira, Henrique S. Silva, Fabricio Murai, and Ana Paula C. da Silva. Predicting user emotional tone in mental disorder online communities. Future Generation Computer Systems, 125(??):641–651, December 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21002764.

Singh:2021:DTL

[SSS21] Deepak Singh, Anurag Shukla, and Mohit Sajwan.

Deep transfer learning framework for the identification of malicious activities to combat cyberattack. Future Generation Computer Systems, 125(??):687–697, December 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21002776.

Singh:2024:OFD

[SSV24] Satveer Singh, Eht E. Sham, and Deo Prakash Vidyarthi. Optimizing fog device deployment for maximal network connectivity and edge coverage using metaheuristic algorithm. Future Generation Computer Systems, 157(??): 529–542, August 2024. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24001377.

Sheganaku:2023:CEA

[SSWW23] Gerta Sheganaku, Stefan Schulte, Philipp Waibel, and Ingo Weber. Costefficient auto-scaling of container-based elastic processes. Future Generation Computer Systems, 138(??):296–312, January 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22002850.

Sahmoud:2020:GFB

[ST20a] Shaaban Sahmoud and Haluk Rahmi Topcuoglu. A general framework based on dynamic multi-objective evolutionary algorithms for handling feature drifts on data streams. Future Generation Computer Systems, 102(??):42–52, January 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19301748.

Sharma:2020:MMS

[ST20b] Shivani Sharma and Durga Toshniwal. MR-I MaxMinscalable two-phase border based knowledge hiding technique using MapReduce. Future Generation Computer Systems, 109(??):538–550, August 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17322173.

Simpkin:2020:EON

[STH^{+}20] Chris Simpkin, Ian Taylor, Daniel Harborne, Graham Bent, Alun Preece, and Raghu K. Ganti. Efficient orchestration of Node-RED IoT workflows using a Vector Symbolic Architecture. Future Generation Computer Systems, 111(??):117–131, October 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19317467.

Sajina:2024:MTP

[ŠTI24] Robert Šajina, Nikola Tanković, and Ivo Ipšić. Multi-task peer-to-peer learning using an encoderonly transformer model. Future Generation Computer Systems, 152(??): 170–178, March 2024. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23004053.

Sanchez:2020:SBP

[STK20] Odnan Ref Sanchez, Ilaria Torre, and Bart P. Knijnenburg. Semantic-based privacy settings negotiation and management. Future Generation Computer Systems, 111(??):879–898, October 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18317035.

Soveizi:2023:SPC

[STK23] Nafiseh Soveizi, Fatih Turkmen, and Dimka Karastoyanova. Security and privacy concerns in cloudbased scientific and business workflows: a systematic review. Future Generation Computer Systems, 148(??):184–200, November 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23001991.

Shafiq:2020:SEM

[STS^{+}20] Muhammad Shafiq, Zhihong Tian, Yanbin Sun, Xiaojiang Du, and Mohsen Guizani. Selection of effective machine learning algorithm and Bot– IoT attacks traffic identification for Internet of Things in smart city. Future Generation Computer Systems, 107(??):433–442, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19334880.

Shandilya:2022:AAC

[SUKN22] Shishir Kumar Shandilya, Saket Upadhyay, Ajit Kumar, and Atulya K. Nagar. AI-assisted Computer Network Operations testbed for Nature-Inspired Cyber Security based adaptive defense simulation and analysis. Future Generation Computer Systems, 127(??):297–308, February 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21003642.

Sun:2020:RID

[Sun20] Chenghao Sun. Research on investment decisionmaking model from the perspective of “Internet of Things + Big data”. Future Generation Computer Systems, 107(??):286–292, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19321387.

Sikandar:2023:CBK

[SuRMA^{+}23] Hira S. Sikandar, Saif ur Rehman Malik, Adeel Anjum, Abid Khan, and Gwanggil Jeon. Cohortbased kernel principal component analysis with multipath service routing in federated learning. Future Generation Computer Systems, 149(??):518–530, December 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23002947.

Sardianos:2020:RCR

[SVD^{+}20] Christos Sardianos, Iraklis Varlamis, George Dimitrakopoulos, Dimosthenis Anagnostopoulos, Abdullah Alsalemi, Faycal Bensaali, Yassine Himeur, and Abbes Amira. REHAB-C: Recommendations for Energy HABits Change. Future Generation Computer Systems, 112(??):394–407, November 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19317327.

Sarmento:2020:IPA

[SVFdA20] Róger M. Sarmento, Francisco F. X. Vasconcelos, Pedro P. Rebouças Filho, and Victor Hugo C. de Albuquerque. An IoT platform for the analysis of brain CT images based on Parzen analysis. Future Generation Computer Systems, 105(??):135–147, April 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1932182X.

Sengan:2020:ECP

[SVN^{+}20a] Sudhakar Sengan, Subramaniyaswamy V., Sreekumar Krishnan Nair, Indragandhi V., Manikandan J., and Logesh Ravi. Enhancing cyber–physical systems with hybrid smart city cyber security architecture for secure public data-smart network. Future Generation Computer Systems, 112(??):724–737, November 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20300765.

Sun:2020:FLB

[SVN20b] Jiawen Sun, Hans Vandierendonck, and Dimitrios S. Nikolopoulos. Fast load balance parallel graph analytics with an automatic graph data structure selection algorithm. Future Generation Computer Systems, 112(??):612–623, November 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19316590.

Shu:2020:EEM

[SW20] Tong Shu and Chase Q. Wu. Energy-efficient mapping of large-scale workflows under deadline constraints in big data computing systems. Future Generation Computer Systems, 110(??):515–530, September 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17300468.

Salman:2022:NOR

[SW22] Mohammed I. Salman and Bin Wang. Near-optimal responsive traffic engineering in software defined networks based on deep learning. Future Generation Computer Systems, 135(??):172–180, October 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22001650.

Sun:2020:MCC

[SWL^{+}20] Xiao-Li Sun, Hui Wang, Xin-Ke Li, Guo-Hong Cao, Yu Kuang, and Xiao-Chen Zhang. Monte Carlo computer simulation of a camera system for proton beam range verification in cancer treatment. Future Generation Computer Systems, 102(??):978–991, January 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19321107.

Shen:2020:MCB

[SWW^{+}20] Jian Shen, Chen Wang, Anxi Wang, Qi Liu, and Yang Xiang. Moving centroid based routing protocol for incompletely predictable cyber devices in cyber-physical-social distributed systems. Future Generation Computer Systems, 108(??):1129–1139, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17311573.

Song:2023:IPA

[SXC^{+}23] Yujia Song, Ruyue Xin, Peng Chen, Rui Zhang, Juan Chen, and Zhiming Zhao. Identifying performance anomalies in fluctuating cloud environments: a robust correlative-GNN-based explainable approach. Future Generation Computer Systems, 145(??):77–86, August 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23000973.

Song:2024:ASF

[SXC^{+}24] Yujia Song, Ruyue Xin, Peng Chen, Rui Zhang, Juan Chen, and Zhiming Zhao. Autonomous selection of the fault classification models for diagnosing microservice applications.

Future Generation Computer Systems, 153(??): 326–339, April 2024. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23004661.

Song:2022:AKS

[SXF22] Xiaojia Song, Tao Xie, and Stephen Fischer. Accelerating kNN search in high dimensional datasets on FPGA by reducing external memory access. Future Generation Computer Systems, 137(??):189–200, December 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22002370.

Song:2024:DHB

[SXHD24] Xiaoling Song, Guangxia Xu, Yongfei Huang, and Jingnan Dong. DID-HVC-based Web3 healthcare data security and privacy protection scheme. Future Generation Computer Systems, 158(??):267–276, September 2024. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24001468.

Song:2022:ODT

[SXW^{+}22] Fuhong Song, Huanlai Xing, Xinhan Wang, Shouxi Luo, Penglin Dai, and Ke Li. Offloading dependent tasks in multiaccess edge computing: a multi-objective reinforcement learning approach. Future Generation Computer Systems, 128(??): 333–348, March 2022. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21004039.

Shan:2023:KDF

[SXZZ23] Chenggang Shan, Yuanqing Xia, Yufeng Zhan, and Jinhui Zhang. KubeAdaptor: a docking framework for workflow containerization on Kubernetes. Future Generation Computer Systems, 148(??):584–599, November 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2300242X.

Samuel:2020:NTP

[SYG^{+}20] Oluwarotimi Williams Samuel, Bin Yang, Yanjuan Geng, Mojisola Grace Asogbon, Sandeep Pirbhulal, Deogratias Mzurikwao, Oluwagbenga Paul Idowu, Tunde Joseph Ogundele, Xiangxin Li, Shixiong Chen, Ganesh R. Naik, Peng Fang, Fanghai Han, and Guanglin Li.

A new technique for the prediction of heart failure risk driven by hierarchical neighborhood componentbased learning and adaptive multi-layer networks. Future Generation Computer Systems, 110(??):781–794, September 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19316279.

Singh:2023:ELI

[SYHX23] Bikash Chandra Singh, Qingqing Ye, Haibo Hu, and Bin Xiao. Efficient and lightweight indexing approach for multidimensional historical data in blockchain. Future Generation Computer Systems, 139(??):210–223, February 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22002862.

Shi:2023:VDN

[SYW^{+}23] Jianzhi Shi, Bo Yi, Xingwei Wang, Min Huang, Peichen Li, Chao Zeng, and Keqin Li. Vivace-Distributed: a novel congestion control mechanism for Joint-Cloud environments. Future Generation Computer Systems, 149(??):317–329, December 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X2300273X.

Sheng:2022:AAC

[SYXL22] Deming Sheng, Jingling Yuan, Qing Xie, and Lin Li. ACMF: an attention collaborative extended matrix factorization based model for MOOC course service via a heterogeneous view. Future Generation Computer Systems, 126(??):211–224, January 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21003009.

Song:2021:DPM

[SYXW21] Yu Song, Jin Yang, Renjun Xie, and Yi Wu. Dynamic positioning model of offshore oil drilling platform based on OIPSO algorithm. Future Generation Computer Systems, 124(??):168–173, November 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21001473.

Sarwar:2021:LDC

[SYYuR21] Kinza Sarwar, Sira Yongchareon, Jian Yu, and Saeed ur Rehman.

Lightweight, divide-and-conquer privacy-preserving data aggregation in fog computing. Future Generation Computer Systems, 119(??):188–199, June 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167739X21000650.

Sarwar:2022:EPP

[SYYuR22] Kinza Sarwar, Sira Yongchareon, Jian Yu, and Saeed ur Rehman. Efficient privacy-preserving data replication in fogenabled IoT. Future Generation Computer Systems, 128(??):538–551, March 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21004209.

Shi:2022:BGE

[SZdLZ22] Zeshun Shi, Huan Zhou, Cees de Laat, and Zhiming Zhao. A Bayesian gameenhanced auction model for federated cloud services using blockchain. Future Generation Computer Systems, 136(??):49–66, November 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22001881.

Sun:2024:ALB

[SZGB24] Dawei Sun, Chunlin Zhang, Shang Gao, and Rajkumar Buyya. An adaptive load balancing strategy for stateful join operator in skewed data stream environments. Future Generation Computer Systems, 152(??):138–151, March 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23004016.

Song:2021:IHH

[SZL^{+}21] Tao Song, Honghua Zhao, Zhi Liu, Hao Liu, Yuanyuan Hu, and Dianmin Sun. Intelligent human hand gesture recognition by local– global fusing quality-aware features. Future Generation Computer Systems, 115(??):298–303, February 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20318343.

Simonsson:2021:OCE

[SZM^{+}21] Jesper Simonsson, Long Zhang, Brice Morin, Benoit Baudry, and Martin Monperrus. Observability and chaos engineering on system calls for containerized applications in Docker. Future Generation Computer Systems, 122(??):117–129, September 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21001163.

Su:2022:BRB

[SZM22] Jian Su, Leyou Zhang,

and Yi Mu. BA-RMKABSE:

Blockchain-aided ranked multi-keyword attributebased searchable encryption with hiding policy for smart health system. Future Generation Computer Systems, 132(??):299–309, July 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22000292.

Sun:2020:RRL

[SZO^{+}20] Guolin Sun, Tong Zhan, Boateng Gordon Owusu, Ayepah-Mensah Daniel, Guisong Liu, and Wei Jiang. Revised reinforcement learning based on anchor graph hashing for autonomous cell activation in cloud-RANs. Future Generation Computer Systems, 104(??):60–73, March 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18318788.

Sun:2021:LHF

[SZS^{+}21] Dianmin Sun, Honghua Zhao, Tao Song, Aiqin Liu, Jinling Cheng, Zhi Liu, and Xin Zhao. Learning hierarchical face representation to enhance HCI among medical robots. Future Generation Computer Systems, 118(??):180–186, May 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20330314.

Sanchez-Zas:2023:OBA

[SZVVB^{+}23] Carmen Sánchez-Zas, Víctor A. Villagrá, Mario Vega-Barbas, Xavier Larriva-Novo, José Ignacio Moreno, and Julio Berrocal. Ontologybased approach to realtime risk management and cyber-situational awareness. Future Generation Computer Systems, 141 (??):462–472, April 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0167739X22004058.

Shi:2023:CGA

[SZW^{+}23] Lei Shi, Yimin Zhou, Juan Wang, Zuli Wang, Ding Chen, Haifeng Zhao, Wankou Yang, and Edward Szczerbicki. Compact global association based adaptive routing framework for personnel behavior understanding. Future Generation Computer Systems, 141(??):514–525, April 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22004071.

Song:2022:VVC

[SZZY22] Jie Song, Peimeng Zhu, Yanfeng Zhang, and Ge Yu. Versatility or validity: a comprehensive review on simulation of datacenters powered by renewable energy mix. Future Generation Computer Systems, 136(??):326–341, November 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22002199.

Toosi:2021:TFA

[TA21] Ramin Toosi and Mohammad Ali Akhaee. Time– frequency analysis of keystroke dynamics for user authentication. Future Generation Computer Systems, 115(??):438–447, February 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19319247.

Tiwari:2023:EAC

[TA23] Sadhana Tiwari and Sonali Agarwal. Empirical analysis of chronic disease dataset for multiclass classification using optimal feature selection based hybrid model with spark streaming. Future Generation Computer Systems, 139(??):87–99, February 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22002990.

Tolba:2021:PDA

[TAM21] Amr Tolba and Zafer Al-Makhadmeh. Predictive data analysis approach for securing medical data in smart grid healthcare systems. Future Generation Computer Systems, 117 (??):87–96, April 2021. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20330326.

Tao:2023:SOE

[Tao23]  Ming Tao.  Semantic

ontology enabled modeling, retrieval and inference for incomplete mobile trajectory data. Future Generation Computer Systems, 145(??):1–11, August 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23000894.

Taufer:2023:ROF

[Tau23] Michela Taufer. Recognition of outstanding Future Generation Computer Systems reviewers for 2022. Future Generation Computer Systems, 143(??):420–421, June 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23000687.

Taufer:2024:RBP

[Tau24] Michela Taufer. Recognition of best paper, outstanding editors, and outstanding reviewers for Future Generation Computer Systems in 2023. Future Generation Computer Systems, 155(??):463–464, June 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24000657.

Tsai:2023:TPA

[TBA23] Yu-Hsiang Mike Tsai, Natalie Beams, and Hartwig Anzt. Three-precision algebraic multigrid on GPUs. Future Generation Computer Systems, 149(??):280–293, December 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23002741.

Tortorella:2023:RMP

[TBB^{+}23] Yvan Tortorella, Luca Bertaccini, Luca Benini, Davide Rossi, and Francesco Conti. RedMule: a mixedprecision matrix–matrix operation engine for flexible and energy-efficient on-chip linear algebra and TinyML training acceleration. Future Generation Computer Systems, 149(??):122–135, December 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23002546.

Tuli:2020:HED

[TBG^{+}20] Shreshth Tuli, Nipam Basumatary, Sukhpal Singh Gill, Mohsen Kahani, Rajesh Chand Arya, Gurpreet Singh Wander, and Rajkumar Buyya. Health-Fog: an ensemble deep learning based smart healthcare system for automatic diagnosis of heart diseases in integrated IoT and fog computing environments. Future Generation Computer Systems, 104(??): 187–200, March 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19313391.

Thabet:2023:SOP

[TBH23] Marwa Thabet, Mouhebeddine Berrima, and Brahim Hnich. A secure optimal placement strategy based on Monte Carlo simulation and hypothesis testing. Future Generation Computer Systems, 138(??):89– 103, January 2023. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2200259X.

Taghavi:2020:TSG

[TBO20] Mona Taghavi, Jamal Bentahar, and Hadi Otrok. Two-stage game theoretical framework for IaaS market share dynamics. Future Generation Computer Systems, 102(??):173–189, January 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18324531.

Tusa:2023:EES

[TC23] Francesco Tusa and Stuart Clayman. End-to-end slices to orchestrate resources and services in the cloudto-edge continuum. Future Generation Computer Systems, 141(??):473–488, April 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22003971.

Tusa:2024:MSF

[TCBF24] Francesco Tusa, Stuart Clayman, Alina Buzachis, and Maria Fazio. Microservices and serverless functions — lifecycle, performance, and resource utilisation of edge based realtime IoT analytics. Future Generation Computer Systems, 155(??):204–218, June 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24000529.

Tanganelli:2020:MDD

[TCMV20] Giacomo Tanganelli, Luca Cassano, Antonio Miele, and Carlo Vallati. A methodology for the design and deployment of distributed cyber-physical systems for smart environments. Future Generation Computer Systems, 109(??):420–430, August 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19325245.

Tang:2022:NDM

[TCW^{+}22] Dan Tang, Jingwen Chen, Xiyin Wang, Siqi Zhang, and Yudong Yan. A new detection method for LDoS attacks based on data mining. Future Generation Computer Systems, 128(??):73–87, March 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2100385X.

Tempelmeier:2021:LOK

[TD21] Nicolas Tempelmeier and Elena Demidova. Linking OpenStreetMap with knowledge graphs — link discovery for schema-agnostic volunteered geographic information. Future Generation Computer Systems, 116(??):349–364, March 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20330272.

Tserpes:2020:SIH

[TDC^{+}20] Konstantinos Tserpes, Patrizio Dazzi, Emanuele Carlini, Massimo Coppola, and Dimitrios Zissis. Special issue on high performance services computing and Internet technologies. Future Generation Computer Systems, 109(??):642–643, August 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20311705.

Teng:2021:LCP

[TDL^{+}21] Haojun Teng, Mianxiong Dong, Yuxin Liu, Wang Tian, and Xuxun Liu. A low-cost physical location discovery scheme for largescale Internet of Things in smart city through joint use of vehicles and UAVs. Future Generation Computer Systems, 118(??):310–326, May 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2100042X.

Tang:2020:DCH

[TDLT20] Chengxiang Tang, Xuefan Dong, Ying Lian, and Daisheng Tang. Do Chinese hospital services constitute an oligopoly? Evidence of the rich-club phenomenon in a patient referral network. Future Generation Computer Systems, 105(??):492–501, April 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1931622X.

Tong:2022:RTE

[TDM^{+}22] Zhao Tong, Xiaomei Deng, Jing Mei, Bilan Liu, and Keqin Li. Response time and energy consumption co-offloading with SLRTA algorithm in cloud–edge collaborative computing. Future Generation Computer Systems, 129(??):64– 76, April 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2100443X.

Tesolin:2023:EHM

[TDMC23] Julio Cesar Cardoso Tesolin, André M. Demori, David Fernandes Cruz Moura, and Maria Cláudia Cavalcanti. Enhancing heterogeneous mobile network management based on a well-founded reference ontology. Future Generation Computer Systems, 149(??):577–593, December 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23003084.

Tong:2022:BDD

[TDS^{+}22a] Wei Tong, Xuewen Dong, Yulong Shen, Xiaohong Jiang, and Zhiwei Zhang. A blockchain-driven data exchange model in multidomain IoT with controllability and parallelity. Future Generation Computer Systems, 135(??):85– 94, October 2022. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22001558.

Tong:2022:CSP

[TDS^{+}22b] Wei Tong, Xuewen Dong, Yulong Shen, Yuanyu Zhang, Xiaohong Jiang, and Wensheng Tian. CHChain: Secure and parallel crowdsourcing driven by hybrid blockchain. Future Generation Computer Systems, 131(??):279–291, June 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22000309.

Tewari:2020:SPT

[TG20] Aakanksha Tewari and B. B. Gupta. Security, privacy and trust of different layers in Internet-of-Things (IoTs) framework. Future Generation Computer Systems, 108(??):909–920, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17321003.

Tudor:2020:BDP

[TGAP20] Valentin Tudor, Vincenzo Gulisano, Magnus Almgren, and Marina Papatriantafilou. BES: Differentially private event aggregation for large-scale IoTbased systems. Future Generation Computer Systems, 108(??):1241–1257, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17322823.

Taheri:2020:SBA

[TGJ^{+}20] Rahim Taheri, Meysam Ghahramani, Reza Javidan, Mohammad Shojafar, Zahra Pooranian, and Mauro Conti. Similaritybased Android malware detection using Hamming distance of static binary features. Future Generation Computer Systems, 105 (??):230–247, April 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167739X19315122.

Talluri:2024:EDS

[THA^{+}24] Sacheendra Talluri, Nikolas Herbst, Cristina Abad, Tiziano De Matteis, and Alexandru Iosup. ExDe: Design space exploration of scheduler architectures and mechanisms for serverless data-processing. Future Generation Computer Systems, 153(??):84–96, April 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23004211.

Thabet:2023:IQC

[THB23] Marwa Thabet, Brahim Hnich, and Mouhebeddine Berrima. Investigating, quantifying and controlling the co-location attack’s conditional value at risk of VM placement strategies. Future Generation Computer Systems, 149(??):464–477, December 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23002868.

Tsanakas:2024:LWE

[THVL24] Stylianos Tsanakas, Aroosa Hameed, John Violos, and Aris Leivadeas. A lightweight edge-enabled knowledge distillation technique for next location prediction of multitude transportation means. Future Generation Computer Systems, 154 (??):45–58, May 2024. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23004867.

Tubishat:2021:EAE

[TIA21] Mohammad Tubishat, Norisma Idris, and Mohammad Abushariah. Explicit aspects extraction in sentiment analysis using optimal rules combination. Future Generation Computer Systems, 114(??):448–480, January 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1933081X.

Tan:2020:MPS

[TJG^{+}20] Zhaonian Tan, Weixing Ji, Jianhua Gao, Yueyan Zhao, Akrem Benatia, Yizhuo Wang, and Feng Shi. MMSparse: 2D partitioning of sparse matrix based on mathematical morphology. Future Generation Computer Systems, 108(??):521–532, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19327967.

Tran:2024:ORU

[TK24a] Minh-Ngoc Tran and Young-Han Kim. Optimized resource usage with hybrid auto-scaling system for knative serverless edge computing. Future Generation Computer Systems, 152(??):304–316, March 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23004156.

Turgut:2024:EHD

[TK24b] Zeynep Turgut and Arzu Gorgulu Kakisim. An explainable hybrid deep learning architecture for WiFi-based indoor localization in Internet of Things environment. Future Generation Computer Systems, 151(??):196–213, February 2024. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2300376X.

Tao:2023:ESV

[TKS^{+}23] Yunting Tao, Fanyu Kong, Yuliang Shi, Jia Yu, Hanlin Zhang, and Xiangyi Wang. Efficient, secure and verifiable outsourcing scheme for SVD-based collaborative filtering recommender system. Future Generation Computer Systems, 149(??):445–454, December 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23002996.

Tang:2020:HLS

[TLC^{+}20] Jie Tang, Shaoshan Liu, Jie Cao, Dawei Sun, Bolin Ding, Jean-Luc Gaudiot, and Weisong Shi. π-Hub: Large-scale video learning, storage, and retrieval on heterogeneous hardware platforms. Future Generation Computer Systems, 102(??):514–523, January 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1931427X.

Tang:2022:AMO

[TLJ^{+}22] Junwei Tang, Ruixuan Li, Yu Jiang, Xiwu Gu, and Yuhua Li. Android malware obfuscation variants detection method based on multi-granularity opcode features. Future Generation Computer Systems, 129(??):141–151, April 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21004349.

Tian:2021:MRP

[TLKX21] Shasha Tian, Yuanxiang Li, Yilin Kang, and Jiening Xia. Multi-robot path planning in wireless sensor networks based on jump mechanism PSO and safety gap obstacle avoidance.

Future Generation Computer Systems, 118(??):37– 47, May 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20330740.

Tang:2024:RFL

[TLL^{+}24] Yuncan Tang, Yongquan Liang, Yang Liu, Jinquan Zhang, Lina Ni, and Liang Qi. Reliable federated learning based on dual-reputation reverse auction mechanism in Internet of Things. Future Generation Computer Systems, 156(??):269–284, July 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X2400089X.

Taguchi:2021:GCN

[TLM21] Hibiki Taguchi, Xin Liu, and Tsuyoshi Murata. Graph convolutional networks for graphs containing missing features. Future Generation Computer Systems, 117(??):155–168, April 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20330405.

Tapas:2020:ESC

[TLMP20] Nachiket Tapas, Francesco Longo, Giovanni Merlino, and Antonio Puliafito. Experimenting with smart contracts for access control and delegation in IoT. Future Generation Computer Systems, 111(??):324–338, October 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18326979.

Thai:2023:TSA

[TLN23] Huy-Tan Thai, Kim-Hung Le, and Ngan Luu-Thuy Nguyen. Towards sustainable agriculture: a lightweight hybrid model and cloud-based collection of datasets for efficient leaf disease detection. Future Generation Computer Systems, 148(??):488–500, November 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23002388.

Truong:2021:BBT

[TLS^{+}21] Nguyen Truong, Gyu Myoung Lee, Kai Sun, Florian Guitton, and YiKe Guo. A blockchain-based trust system for decentralised applications: When trustless needs trust. Future Generation Computer Systems, 124(??):68–79, November 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21001758.

Tong:2024:DAH

[TLW^{+}24] Yulai Tong, Jiazhen Liu, Hua Wang, Mingjian He, Ke Zhou, Rongfeng He, Qin Zhang, and Cheng Wang. DAG-aware harmonizing job scheduling and data caching for disaggregated analytics frameworks. Future Generation Computer Systems, 156(??):116–129, July 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24000761.

Tang:2023:BSE

[TLX^{+}23] Jine Tang, Xinming Lu, Yong Xiang, Chaochen Shi, and Junhua Gu. Blockchain search engine: Its current research status and future prospect in Internet of Things network. Future Generation Computer Systems, 138(??):120–141, January 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22002692.

Tayal:2020:NMS

[TM20] Devendra K. Tayal and Kanak Meena. A new MapReduce solution for associative classification to handle scalability and skewness in vertical data structure. Future Generation Computer Systems, 103(??):44–57, February 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17322215.

Tardio:2022:BTD

[TMT22] Roberto Tardío, Alejandro Maté, and Juan Trujillo. Beyond TPC-DS, a benchmark for big data OLAP systems (BDOLAP-Bench). Future Generation Computer Systems, 132(??): 136–151, July 2022. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22000553.

Talat:2020:DAP

[TOM^{+}20] Romana Talat, Mohammad S. Obaidat, Muhammad Muzammal, Ali Hassan Sodhro, Zongwei Luo, and Sandeep Pirbhulal. A decentralised approach to privacy preserving trajectory mining. Future Generation Computer Systems, 102(??):382–392, January 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19313901.

Thi:2020:NGJ

[TPD^{+}20] Minh-Thuyen Thi, Jean-Marc Pierson, Georges Da Costa, Patricia Stolf, Jean-Marc Nicod, Gustavo Rostirolla, and Marwa Haddad. Negotiation game for joint IT and energy management in green datacenters. Future Generation Computer Systems, 110(??):1116–1138, September 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1931235X.

Tiwari:2024:KBM

[TPD^{+}24] Vaibhav Tiwari, Chandrasen Pandey, Abisek Dahal, Diptendu Sinha Roy, and Ugo Fiore. A Knapsack-based Metaheuristic for Edge Server Placement in 5G networks with heterogeneous edge capacities. Future Generation Computer Systems, 153(??):222–233, April 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23004363.

Tesone:2020:PIS

[TPF^{+}20] Pablo Tesone, Guillermo Polito, Luc Fabresse, Noury Bouraqadi, and Stéphane Ducasse. Preserving instance state during refactorings in live environments. Future Generation Computer Systems, 110(??):1–17, September 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17306568.

Tyagi:2021:RCT

[TPN^{+}21] Sumarga Kumar Sah Tyagi, Shiva Raj Pokhrel, Mahyar Nemati, Deepak Kumar Jain, Gang Li, and Jinho Choi. Redesigning compound TCP with cognitive edge intelligence for WiFi-based IoT. Future Generation Computer Systems, 125(??):859–868, December 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21002909.

Touma:2020:CCA

[TQC20] Rizkallah Touma, Anna Queralt, and Toni Cortes. CAPre: Code-analysis based prefetching for persistent object stores. Future Generation Computer Systems, 111(??):491–506, October 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19314293.

Taneja:2023:IWS

[TRB^{+}23a] Ashu Taneja, Shalli Rani, Jose Breñosa, Amr Tolba, and Seifedine Kadry. An improved WiFi sensing based indoor navigation with reconfigurable intelligent surfaces for 6G enabled IoT network and AI explainable use case. Future Generation Computer Systems, 149(??):294–303, December 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23002698.

Tanwar:2023:FBI

[TRB^{+}23b] Sudeep Tanwar, Dakshita Ribadiya, Pronaya Bhattacharya, Anuja R. Nair, Neeraj Kumar, and Minho Jo. Fusion of blockchain and IoT in scientific publishing: Taxonomy, tools, and future directions. Future Generation Computer Systems, 142(??):248–275, May 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2200440X.

Tuli:2020:SDA

[TSB20] Shreshth Tuli, Rajinder Sandhu, and Rajkumar Buyya. Shared data-aware dynamic resource provisioning and task scheduling for data intensive applications on hybrid clouds using Aneka. Future Generation Computer Systems, 106(??):595–606, May 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19319508.

Tounsi:2023:ITD

[TSKK23] Imen Tounsi, Abdessamad Saidi, Mohamed Hadj Kacem, and Ahmed Hadj Kacem. Internet of Things design patterns modeling proven correct by construction: Application to aged care solution. Future Generation Computer Systems, 148(??):395–407, November 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23002017.

Thein:2024:PFL

[TSM24] Thin Tharaphe Thein, Yoshiaki Shiraishi, and Masakatu Morii. Personalized federated learningbased intrusion detection system: Poisoning attack and defense. Future Generation Computer Systems, 153(??):182–192, April 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23003783.

Tahir:2020:PDQ

[TSR^{+}20] Shahzaib Tahir, Liutauras Steponkus, Sushmita Ruj, Muttukrishnan Rajarajan, and Ali Sajjad. A parallelized disjunctive query based searchable encryption scheme for big data. Future Generation Computer Systems, 109(??): 583–592, August 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17321842.

Tao:2024:MEO

[TSX^{+}24] Xinkun Tao, Pengfei Sun, Zhuang Xiao, Chengcheng Fu, Xiaoyun Feng, and Qingyuan Wang. Modeling and energy-optimal control for freight trains based on data-driven approaches. Future Generation Computer Systems, 152(??): 346–360, March 2024. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2300420X.

Tang:2020:MAL

[TTD^{+}20] Dan Tang, Liu Tang, Rui Dai, Jingwen Chen, Xiong Li, and Joel J. P. C. Rodrigues. MF-Adaboost: LDoS attack detection based on multi-features and improved Adaboost. Future Generation Computer Systems, 106(??):347–359, May 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19310544.

Todeschi:2020:CZO

[TTTH20] Grégoire Todeschi, Boris Teabe, Alain Tchana, and Daniel Hagimont. Cacol: a zero overhead and nonintrusive double caching mitigation system. Future Generation Computer Systems, 106(??):14–21, May 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19309239.

Tang:2021:SUS

[TTZ^{+}21] Wenyi Tang, Ling Tian, Xu Zheng, Guangchun Luo, and Zaobo He. Susceptible user search for defending opinion manipulation. Future Generation Computer Systems, 115(??):531–541, February 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20329800.

Tariq:2020:EMA

[TWI20] Umair Ullah Tariq, Hui Wu, and Suhaimi Abd Ishak. Energy and memoryaware software pipelining streaming applications on NoC-based MPSoCs. Future Generation Computer Systems, 111(??):1–16, October 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19330493.

Tao:2023:DWE

[TWL23] Ye Tao, Can Wang, and Alan Wee-Chung Liew. Dynamic weighted ensemble learning for sequential recommendation systems: the AIRE model. Future Generation Computer Systems, 149(??):162–170, December 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23002418.

Tong:2023:MTT

[TWM^{+}23] Zhao Tong, Jiake Wang, Jing Mei, Kenli Li, Wenbin Li, and Keqin Li. Multitype task offloading for wireless Internet of Things by federated deep reinforcement learning. Future Generation Computer Systems, 145(??):536–549, August 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23001383.

Tan:2023:SFM

[TWY^{+}23] Jingweijia Tan, Qixiang Wang, Kaige Yan, Xiaohui Wei, and Xin Fu. Saca-FI: a microarchitecturelevel fault injection framework for reliability analysis of systolic array based CNN accelerator. Future Generation Computer Systems, 147(??):251–264, October 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2300184X.

Tripathi:2022:NEB

[TYR22] Shashi Prakash Tripathi, Rahul Kumar Yadav, and Abhay Kumar Rai. Network embedding based link prediction in dynamic networks. Future Generation Computer Systems, 127(??):409–420, February 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21003708.

Tian:2022:FGA

[TZW^{+}22] Teng Tian, Letian Zhao, Xiaotian Wang, Qizhe Wu, Wei Yuan, and Xi Jin. FP-GNN: Adaptive FPGA accelerator for graph neural networks. Future Generation Computer Systems, 136(??):294–310, November 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22002217.

Unal:2021:SEI

[UAACH21] Devrim Unal, Abdulla Al-Ali, Ferhat Ozgur Catak, and Mohammad Hammoudeh. A secure and efficient Internet of Things cloud encryption scheme with forensics investigation compatibility based on identity-based encryption. Future Generation Computer Systems, 125(??):433–445, December 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21002454.

Uctu:2021:STE

[UADD21] Göksel Uçtu, Mustafa Alkan, İbrahim Alper Doğru, and Murat Dörterler. A suggested testbed to evaluate multicast network and threat prevention performance of next generation firewalls. Future Generation Computer Systems, 124(??):56–67, November 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21001631.

Usama:2020:ABS

[UAS^{+}20] Mohd Usama, Belal Ahmad, Enmin Song, M. Shamim Hossain, Mubarak Alrashoud, and Ghulam Muhammad. Attentionbased sentiment analysis using convolutional and recurrent neural network. Future Generation Computer Systems, 113(??):571–578, December 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19334600.

Ullah:2020:IHS

[UCO20] Sami Ullah, Joontae Choi, and Heekuck Oh. IPsec for high speed network links: Performance analysis and enhancements. Future Generation Computer Systems, 107(??):112–125, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19323143.

Uban:2021:ECB

[UCR21] Ana-Sabina Uban, Berta Chulvi, and Paolo Rosso. An emotion and cognitive based analysis of mental health disorders from social media data. Future Generation Computer Systems, 124(??):480–494, November 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21001825.

Ehatisham-ul-Haq:2020:OSI

[uHA20]  Muhammad  Ehatisham

ul Haq and Muhammad Awais Azam. Opportunistic sensing for inferring in-the-wild human contexts based on activity pattern recognition using smart computing. Future Generation Computer Systems, 106(??):374–392, May 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19314967.

Usman:2020:QSD

[UJHN20] Muhammad Usman, Mian Ahmad Jan, Xiangjian He, and Priyadarsi Nanda. QASEC: a secured data communication scheme for mobile ad-hoc networks. Future Generation Computer Systems, 109(??): 604–610, August 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1732321X.

Usman:2020:ORF

[UKY^{+}20] Muhammad Usman, Luk Knapen, Ansar-Ul-Haque Yasar, Tom Bellemans, Davy Janssens, and Geert Wets. Optimal recharging framework and simulation for electric vehicle fleet. Future Generation Computer Systems, 107(??):745–757, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17307689.

Ujjan:2020:TSA

[UPD^{+}20] Raja Majid Ali Ujjan, Zeeshan Pervez, Keshav Dahal, Ali Kashif Bashir, Rao Mumtaz, and J. González. Towards sFlow and adaptive polling sampling for deep learning based DDoS detection in SDN. Future Generation Computer Systems, 111(??):763–779, October 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19318333.

Ullah:2023:HHI

[UPK^{+}23] Fasee Ullah, Chi-Man Pun, Omprakash Kaiwartya, Ali Safaa Sadiq, Jaime Lloret, and Mohammed Ali. HIDE — Healthcare IoT Data Trust ManagEment: Attribute centric intelligent privacy approach. Future Generation Computer Systems, 148(??):326–341, November 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23001838.

Baig:2020:ASW

[uRBIBC20] Shuja ur Rehman Baig, Waheed Iqbal, Josep Lluis Berral, and David Carrera. Adaptive sliding windows for improved estimation of data center resource utilization. Future Generation Computer Systems, 104(??):212–224, March 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19309203.

urRehman:2021:DAD

[uRKI^{+}21] Saif ur Rehman, Mubashir Khaliq, Syed Ibrahim Imtiaz, Aamir Rasool, Muhammad Shafiq, Abdul Rehman Javed, Zunera Jalil, and Ali Kashif Bashir. DIDDOS: an approach for detection and identification of distributed denial of service (DDoS) cyberattacks using gated recurrent units (GRU). Future Generation Computer Systems, 118(??):453–466, May 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21000327.

urRehman:2021:DCA

[uRLW^{+}21] Muhammad Habib ur Rehman, Chee Sun Liew, Teh Ying Wah, Muhammad Imran, Khaled Salah, Nidal Nasser, and Davor Svetinovic. Device-centric adaptive data stream management and offloading for analytics applications in future internet architectures. Future Generation Computer Systems, 114(??):155–168, January 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20303204.

Ullah:2020:IEE

[URN^{+}20] Rehmat Ullah, Muhammad Atif Ur Rehman, Muhammad Ali Naeem, Byung-Seo Kim, and Spyridon Mastorakis. ICN with edge for 5g: Exploiting in-network caching in ICN-based edge computing for 5G networks. Future Generation Computer Systems, 111(??):159–174, October 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19328900.

Ullah:2022:AIT

[UUH^{+}22] Waseem Ullah, Amin Ullah, Tanveer Hussain, Khan Muhammad, Ali Asghar Heidari, Javier Del Ser, Sung Wook Baik, and Victor Hugo C. De Albuquerque. Artificial intelligence of things-assisted two-stream neural network for anomaly detection in surveillance big video data. Future Generation Computer Systems, 129(??): 286–297, April 2022. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21004295.

Usman:2021:IDM

[UUK^{+}21] Nighat Usman, Saeeda Usman, Fazlullah Khan, Mian Ahmad Jan, Ahthasham Sajid, Mamoun Alazab, and Paul Watters. Intelligent dynamic malware detection using machine learning in IP reputation for forensics data analytics. Future Generation Computer Systems, 118(??): 124–141, May 2021. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21000066.

Ullah:2021:ITF

[UYH21] Ihsan Ullah, Hee Yong Youn, and Youn-Hee Han. Integration of type-2 fuzzy logic and Dempster–Shafer theory for accurate inference of IoT-based healthcare system. Future Generation Computer Systems, 124(??):369–380, November 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21002077.

Veiga:2023:TCR

[VAKB23] Tiago Veiga, Hafiz Areeb Asad, Frank Alexander Kraemer, and Kerstin Bach. Towards containerized, reuse-oriented AI deployment platforms for cognitive IoT applications. Future Generation Computer Systems, 142(??):4– 13, May 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22004320.

Versluis:2023:LMW

[VCG^{+}23] Laurens Versluis, Mehmet Cetin, Caspar Greeven, Kristian Laursen, Damian Podareanu, Valeriu Codreanu, Alexandru Uta, and Alexandru Iosup. Less is not more: We need rich datasets to explore. Future Generation Computer Systems, 142(??):117–130, May 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22004265.

Vuurstaek:2020:GBS

[VCK^{+}20] Jan Vuurstaek, Glenn Cich, Luk Knapen, Wim Ectors, Ansar-Ul-Haque Yasar, Tom Bellemans, and Davy Janssens. GTFS bus stop mapping to the OSM network. Future Generation Computer Systems, 110(??):393–406, September 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17324330.

Viviani:2021:AVP

[VCM^{+}21] Marco Viviani, Cristina Crocamo, Matteo Mazzola, Francesco Bartoli, Giuseppe Carrà, and Gabriella Pasi. Assessing vulnerability to psychological distress during the COVID-19 pandemic through the analysis of microblogging content. Future Generation Computer Systems, 125(??):446–459, December 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21002399.

Vandervelden:2022:SKV

[VDSB22] Thibaut Vandervelden, Ruben De Smet, Kris Steenhaut, and An Braeken. SHA 3 and Keccak variants computation speeds on constrained devices. Future Generation Computer Systems, 128(??):28–35, March 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21003885.

Volpert:2023:VSM

[VEH^{+}23] Simon Volpert, Philipp Eichhammer, Florian Held, Thomas Huffert, Hans P. Reiser, and Jörg Domaschka. The view on systems monitoring and its requirements from future Cloud-to-Thing applications and infrastructures. Future Generation Computer Systems, 141(??): 243–257, April 2023. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22003843.

Vega:2020:ITD

[VFOV20] Esteban Alejandro Armas Vega, Edgar González Fernández, Ana Lucila Sandoval Orozco, and Luis Javier García Villalba. Image tampering detection by estimating interpolation patterns. Future Generation Computer Systems, 107(??):229–237, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19325282.

Valdez:2021:CBC

[VG21] Mario García Valdez and Juan J. Merelo Guervós. A container-based cloudnative architecture for the reproducible execution of multi-population optimization algorithms. Future Generation Computer Systems, 116(??):234–252, March 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20330235.

Vila:2023:CCV

[VGL23] Sergi Vila, Fernando Guirado, and Josep L. Lérida. Cloud computing virtual machine consolidation based on stock trading forecast techniques. Future Generation Computer Systems, 145(??):321–336, August 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23000961.

Vaarandi:2024:SCG

[VGM24] Risto Vaarandi and Alejandro Guerra-Manzanares. Stream clustering guided supervised learning for classifying NIDS alerts. Future Generation Computer Systems, 155(??):231–244, June 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24000396.

Vidal:2022:EFA

[VHP^{+}22] Vinicius F. Vidal, Leonardo M.

Honório, Milena F. Pinto, Mario A. R. Dantas, Maria Júlia Aguiar, and Miriam Capretz. An edge– fog architecture for distributed 3D reconstruction. Future Generation Computer Systems, 135(??):146–158, October 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22001388.

Versluis:2021:SDW

[VI21] Laurens Versluis and Alexandru Iosup. A survey of domains in workflow scheduling in computing infrastructures: Community and keyword analysis, emerging trends, and taxonomies. Future Generation Computer Systems, 123(??):156–177, October 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21001308.

Varadarajan:2022:PSI

[VKP22] Vijayakumar Varadarajan, Piet Kommers, and Vincenzo Piuri. Preface of special issue on advanced techniques and emerging trends in smart cyber– physical systems. Future Generation Computer Systems, 135(??):299–302, October 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22001765.

Vairetti:2020:ECS

[VMCM^{+}20] Carla Vairetti, Eugenio Martínez-Cámara, Sebastián Maldonado, Victoria Luzón, and Francisco Herrera. Enhancing the classification of social media opinions by optimizing the structural information. Future Generation Computer Systems, 102(??):838–846, January 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19312695.

Verderame:2020:SCE

[VMM^{+}20a] Luca Verderame, Ivan Merelli, Lucia Morganti, Elena Corni, Daniele Cesini, Daniele D’Agostino, and Alessio Merlo. A secure cloud-edges computing architecture for metagenomics analysis. Future Generation Computer Systems, 111(??):919–930, October 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19301207.

Vidal:2020:EEP

[VMM20b] Jorge Maestre Vidal, Marco Antonio Sotelo Monge, and Sergio Mauricio Martínez Monterrubio. EsPADA: Enhanced Payload Analyzer for malware Detection robust against Adversarial threats. Future Generation Computer Systems, 104(??):159–173, March 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19310350.

Vikash:2020:PER

[VMV20] Vikash, Lalita Mishra, and Shirshu Varma. Performance evaluation of realtime stream processing systems for Internet of Things applications. Future Generation Computer Systems, 113(??):207–217, December 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20302636.

Velliangiri:2020:FTE

[VP20] S. Velliangiri and Hari Mohan Pandey. Fuzzy-Taylor-elephant herd optimization inspired deep belief network for DDoS attack detection and comparison with stateof-the-arts algorithms. Future Generation Computer Systems, 110(??):80–90, September 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19332388.

Vijayaragavan:2020:OSV

[VPA20] P. Vijayaragavan, R. Ponnusamy, and M. Aramudhan. An optimal support vector machine based classification model for sentimental analysis of online product reviews. Future Generation Computer Systems, 111(??):234–240, October 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19333138.

Vitali:2022:SIC

[VPBE22] Monica Vitali, Pierluigi Plebani, David Bermbach, and Erik Elmroth. Special issue on co-design of data and computation management in fog computing. Future Generation Computer Systems, 129(??):423–424, April 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21004301.

Valero:2023:ASD

[VPSC^{+}23] Cayetano Valero, Jaime Pérez, Sonia Solera-Cotanilla, Mario Vega-Barbas, Guillermo Suarez-Tangil, Manuel Alvarez-Campana, and Gregorio López. Analysis of security and data control in smart personal assistants from the user’s perspective.

Future Generation Computer Systems, 144(??):12– 23, July 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X2300050X.

Viejo:2020:SMI

[VS20] Alexandre Viejo and David Sánchez. Secure monitoring in IoT-based services via fog orchestration. Future Generation Computer Systems, 107(??):443–457, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1932713X.

Violettas:2021:SID

[VSPM21] George Violettas, George Simoglou, Sophia Petridou, and Lefteris Mamatas. A softwarized intrusion detection system for the RPLbased Internet of Things networks. Future Generation Computer Systems, 125(??):698–714, December 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21002752.

Vercellino:2023:MLA

[VSV^{+}23] Chiara Vercellino, Alberto Scionti, Giuseppe Varavallo, Paolo Viviani, Giacomo Vitali, and Olivier Terzo. A machine learning approach for an HPC use case: the jobs queuing time prediction. Future Generation Computer Systems, 143(??):215–230, June 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23000274.

Vodyaho:2024:CAC

[VZDS24] Alexander Vodyaho, Nataly Zhukova, Radhakrishnan Delhibabu, and Alexey Subbotin. Continuous agile cyber–physical systems architectures based on digital twins. Future Generation Computer Systems, 153(??):350–359, April 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23004326.

Wang:2020:DLA

[Wan20] Jibin Wang. A deep learning approach for atrial fibrillation signals classification based on convolutional and modified Elman neural network. Future Generation Computer Systems, 102(??):670–679, January 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1931828X.

Wang:2021:DLS

[Wan21] Hechuang Wang. Deeplylearned and spatial–temporal feature engineering for human action understanding. Future Generation Computer Systems, 123(??):257–262, October 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21001424.

Wu:2020:MCB

[WBR20] Caesar Wu, Rajkumar Buyya, and Kotagiri Ramamohanarao. Modeling cloud business customers’ utility functions. Future Generation Computer Systems, 105(??):737–753, April 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19324987.

Wang:2020:SHM

[WC20] Xiaonan Wang and Shaohao Cai. Secure healthcare monitoring framework integrating NDN-based IoT with edge cloud. Future Generation Computer Systems, 112(??):320–329, November 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20301862.

Wan:2022:ABR

[WC22a] Jie Wan and Ji Chen. AHP based relay selection strategy for energy harvesting wireless sensor networks. Future Generation Computer Systems, 128(??):36– 44, March 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21003848.

Wang:2022:ARI

[WC22b] Xiaonan Wang and Xilan Chen. Accurate road information warning framework for intelligent vehicles based on resource allocation. Future Generation Computer Systems, 135(??):95–104, October 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22001510.

Wu:2023:SET

[WC23] Yulin Wu and Lanxiang Chen. Structured encryption for triangle counting on graph data. Future Generation Computer Systems, 145(??):200–210, August 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23001073.

Wang:2022:PLM

[WCD^{+}22] Ziheng Wang, Heng Chen, Xiaoshe Dong, Weilin Cai, and Xingjun Zhang. LogSC: Model-based onesided communication performance estimation. Future Generation Computer Systems, 132(??):25–39, July 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22000486.

Wang:2020:DLB

[WCHA20] Eric Ke Wang, Chien-Ming Chen, Mohammad Mehedi Hassan, and Ahmad Almogren. A deep learning based medical image segmentation technique in Internetof-Medical-Things domain. Future Generation Computer Systems, 108(??): 135–144, July 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19323313.

Wang:2024:GAG

[WCL^{+}24] Yahui Wang, Hongchang Chen, Shuxin Liu, Kai Wang, and Yuxiang Hu. Geo-aware graph-augmented self-attention network for individual mobility prediction. Future Generation Computer Systems, 151(??):1–11, February 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23003540.

Wang:2023:MRN

[WCP23] Chen Wang, Tao Chen, and Antonio Plaza. MFE-ResNet: a new extraction framework for land cover characterization in mining areas. Future Generation Computer Systems, 145(??):550–562, August 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23001358.

Wang:2024:GGP

[WCS24] Sheng Wang, Shiping Chen, and Yumei Shi. GPARS: Graph predictive algorithm for efficient resource scheduling in heterogeneous GPU clusters. Future Generation Computer Systems, 152(??):127–137, March 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23003953.

Wang:2021:ESI

[WCW^{+}21] Meng Wang, Weitong Chen, Sen Wang, Yinlin Jiang, Lina Yao, and Guilin Qi. Efficient search over incomplete knowledge graphs in binarized embedding space. Future Generation Computer Systems, 123(??):24–34, October 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21001217.

Wen:2019:ROC

[WCWC19] Ping Wen, Shengduo Chen, Jiarui Wang, and Wei Che. Receiver Operating Characteristics (ROC) analysis for decreased disease risk and elevated treatment response to pegylatedinterferon in chronic hepatitis B patients. Future Generation Computer Systems, 98(??):372–376, September 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1832692X.

See retraction notice [WCWC20].

Wen:2020:RNR

[WCWC20] Ping Wen, Shengduo Chen, Jiarui Wang, and Wei Che. Retraction notice to “Receiver Operating Characteristics (ROC) Analysis for decreased disease risk and elevated treatment response to pegylatedinterferon in chronic hepatitis B patient” [Future Gener. Comput. Syst. 98 (2019) 372–376]. Future Generation Computer Systems, 107(??):1148, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20306518.

See [WCWC19].

Wang:2022:LBP

[WCXW22] Peng Wang, Biwen Chen, Tao Xiang, and Zhongming Wang. Lattice-based public key searchable encryption with fine-grained access control for edge computing. Future Generation Computer Systems, 127(??):373–383, February 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21003587.

Wang:2020:IEG

[WCY^{+}20] Eric Ke Wang, Chien-Ming Chen, Siu Ming Yiu, Mohammad Mehedi Hassan, Majed Alrubaian, and Giancarlo Fortino. Incentive evolutionary game model for opportunistic social networks. Future Generation Computer Systems, 102(??):14–29, January 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19309173.

Wang:2021:OCR

[WCY^{+}21] Xiaogang Wang, Jian Cao, Dingyu Yang, Zhen Qin, and Rajkumar Buyya. Online cloud resource prediction via scalable window waveform sampling on classified workloads. Future Generation Computer Systems, 117(??):338–358, April 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20330673.

Wyrzykowski:2024:PSI

[WD24] Roman Wyrzykowski and Ewa Deelman. Preface of special issue on advances in algorithms, models, hardware, and software for nextgenerations HPC systems. Future Generation Computer Systems, 153(??): 539–542, April 2024. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2300482X.

Wang:2020:ECQ

[WDG20] Xinxin Wang, Depeng Dang, and Zixian Guo. Evaluating the crowd quality for subjective questions based on a Spark computing environment. Future Generation Computer Systems, 106(??):426–437, May 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19313743.

Wang:2020:DRP

[WDHY20] Shangguang Wang, Chuntao Ding, Ching-Hsien Hsu, and Fangchun Yang. Dimensionality reduction via preserving local information. Future Generation Computer Systems, 108(??):967–975, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17319696.

Wang:2021:PCI

[WDL^{+}21] Bin Wang, Shuai Ding, Xiao Liu, X. Li, and Gang Li. Predictive classification of ICU readmission using weight decay random forest. Future Generation Computer Systems, 124(??):351–360, November 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21002065.

Wang:2023:AEP

[WDS^{+}23] Zhe Wang, Matthieu Dorier, Pradeep Subedi, Philip E. Davis, and Manish Parashar. Adaptive elasticity policies for staging-based in situ visualization. Future Generation Computer Systems, 142(??):75–89, May 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22004150.

Wrzeszcz:2021:NAG

[WDSK21] Michał Wrzeszcz, Łukasz Dutka, Renata G. Słota, and Jacek Kitowski. New approach to global data access in computational infrastructures. Future Generation Computer Systems, 125(??):575–589, December 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21002491.

Wei:2021:RAP

[Wei21]  Na Wei.  Research on

the algorithm of painting image style feature extraction based on intelligent vision. Future Generation Computer Systems, 123(??):196–200, October 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21001655.

Wu:2021:OJC

[WF21] Zhaoxi Wu and Liqun Fu.

Optimizing job completion time with fairness in largescale data centers. Future Generation Computer Systems, 114(??):563–573, January 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1931667X.

Wong:2020:DLB

[WFA20] Kelvin K. L. Wong, Giancarlo Fortino, and Derek Abbott. Deep learningbased cardiovascular image diagnosis: a promising challenge. Future Generation Computer Systems, 110(??):802–811, September 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19307435.

Wang:2020:TCC

[WFL^{+}20] Jinyan Wang, Shijian Fang, Chen Liu, Jiawen Qin, Xianxian Li, and Zhenkui Shi. Top-k closed cooccurrence patterns mining with differential privacy over multiple streams. Future Generation Computer Systems, 111(??):339–351, October 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19323684.

Wang:2021:MFF

[WFL^{+}21] Peng Wang, Huitong Fu, Xiaoyan Li, Jia Guo, Zhigang Lv, and Ruohai Di. Multi-feature fusion tracking algorithm based on generative compression network. Future Generation Computer Systems, 124(??):206–214, November 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21001813.

Wang:2022:DUB

[WFLC22] Haiyan Wang, Jiaming Feng, Ke Li, and Lisi Chen. Deep understanding of big geospatial data for selfdriving: Data, technologies, and systems. Future Generation Computer Systems, 137(??):146–163, December 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22002308.

Wang:2022:POC

[WFLL22] Yan Wang, Henian Fang, Linbo Long, and Jinhui Liu. Performance-oriented cache management scheme based on a retention state for energy-harvesting nonvolatile processors. Future Generation Computer Systems, 129(??):90–98, April 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21004398.

Wang:2021:VSC

[WG21] Ai Wang and Xuedong Gao.

A variable scale case-based reasoning method for evidence location in digital forensics. Future Generation Computer Systems, 122(??):209–219, September 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2100100X.

Wang:2020:CCS

[WGG^{+}20] Yawen Wang, Yunfei Guo, Zehua Guo, Thar Baker, and Wenyan Liu. CLO-SURE: a cloud scientific workflow scheduling algorithm based on attackdefense game model. Future Generation Computer Systems, 111(??):460–474, October 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19313603.

Wang:2024:DRL

[WGGB24] Zhiyu Wang, Mohammad Goudarzi, Mingming Gong, and Rajkumar Buyya. Deep reinforcement learning-based scheduling for optimizing system load and response time in edge and fog computing environments. Future Generation Computer Systems, 152(??):55– 69, March 2024. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23003862.

Wang:2024:FCS

[WGL^{+}24] Chenfeng Wang, Xiaoguang Gao, Xinyu Li, Bo Li, and Kaifang Wan. Finding community structure in Bayesian networks by heuristic K-standard deviation method. Future Generation Computer Systems, 158(??):556–568, September 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24001158.

Wang:2020:CDS

[WGLH20] Shangguang Wang, Yan Guo, Yan Li, and Ching-Hsien Hsu. Cultural distance for service composition in cyber-physical-social systems. Future Generation Computer Systems, 108(??):1049–1057, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17310816.

Wang:2024:STI

[WGS24] Yongzhi Wang, Pengfei Gui, and Mehdi Sookhak. Sort-then-insert: a space efficient and oblivious model aggregation algorithm for top-k sparsification in federated learning. Future Generation Computer Systems, 158(??):1– 10, September 2024. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2400150X.

Wang:2020:OAS

[WGW^{+}20] Yuan Wang, Zhipeng Gui, Huayi Wu, Dehua Peng, Jinghang Wu, and Zousen Cui. Optimizing and accelerating space-time Ripley’s K function based on Apache Spark for distributed spatiotemporal point pattern analysis. Future Generation Computer Systems, 105(??):96–118, April 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19314815.

Wang:2021:IIT

[WGW^{+}21] Yawen Wang, Yunfei Guo, Wenbo Wang, Hao Liang, and Shumin Huo. IN-HIBITOR: an intrusion tolerant scheduling algorithm in cloud-based scientific workflow system. Future Generation Computer Systems, 114(??):272–284, January 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20300534.

Wang:2020:BBE

[WHA^{+}20] Junchao Wang, Kaining Han, Anastasios Alexandridis, Zhiyu Chen, Zeljko Zilic, Yu Pang, Gwanggil Jeon, and Francesco Piccialli. A blockchainbased eHealthcare system interoperating with WBANs. Future Generation Computer Systems, 110(??):675–685, September 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19321247.

Wang:2022:PPL

[WHC^{+}22] Mei Wang, Kun He, Jing Chen, Ruiying Du, Bingsheng Zhang, and Zengpeng Li. PANDA: Lightweight non-interactive privacypreserving data aggregation for constrained devices. Future Generation Computer Systems, 131(??):28– 42, June 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22000152.

Wang:2024:ERV

[WHC^{+}24] Huiqing Wang, Yongrong Huang, Zhide Chen, Xu Yang, Xun Yi, Hai Dong, and Xuechao Yang. Estimation of realized volatility of cryptocurrencies using CEEMDAN-RF-LSTM. Future Generation Computer Systems, 158(??): 219–229, September 2024. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167739X24001729.

Wang:2020:LMB

[WHF^{+}20] Junchao Wang, Kaining Han, Shengwen Fan, Ying Zhang, Honghao Tan, Gwanggil Jeon, Yu Pang, and Jinzhao Lin. A logistic mapping-based encryption scheme for wireless body area networks. Future Generation Computer Systems, 110(??):57– 67, September 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19335617.

Wang:2023:PPS

[WHF^{+}23] Junchao Wang, Dongmin Huang, Shengwen Fan, Kaining Han, Gwanggil Jeon, and Joel J. P. C. Rodrigues. PS-DCE: Physiological signalbased double chaotic encryption for instantaneous e-healthcare services. Future Generation Computer Systems, 141(??):116–128, April 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22003569.

Wang:2020:DVP

[WHJ20] Huaqun Wang, Debiao He, and Yimu Ji. Designatedverifier proof of assets for bitcoin exchange using elliptic curve cryptography.

Future Generation Computer Systems, 107(??): 854–862, June 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1731350X.

Wei:2020:EAE

[WHW20] Chen Wei, Zhi-Hua Hu, and You-Gan Wang. Exact algorithms for energyefficient virtual machine placement in data centers. Future Generation Computer Systems, 106(??):77– 91, May 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19319594.

Wang:2020:TTC

[WHZ^{+}20] Hao Wang, Guangjie Han, Chunsheng Zhu, Sammy Chan, and Wenbo Zhang. TCSLP: a trace cost based source location privacy protection scheme in WSNs for smart cities. Future Generation Computer Systems, 107(??):965–974, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17309603.

Wang:2024:RIL

[WJC^{+}24] Qinghu Wang, Jie Jia, Jian Chen, Yansha Deng, Xingwei Wang, and Abdol Hamid Aghvami. Robust indoor localization based on multi-modal information fusion and multi-scale sequential feature extraction. Future Generation Computer Systems, 155(??):164–178, June 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24000475.

Wang:2022:ECI

[WKW^{+}22] Xin Wang, Azim Khan, Jianwu Wang, Aryya Gangopadhyay, Carl Busart, and Jade Freeman. An edge–cloud integrated framework for flexible and dynamic stream analytics. Future Generation Computer Systems, 137(??):323–335, December 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22002576.

Wang:2020:UAC

[WLAC20] G. Wang, B. Lee, J. Ahn, and G. Cho. A UAVassisted CH election framework for secure data collection in wireless sensor networks. Future Generation Computer Systems, 102(??):152–162, January 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18325056.

Wang:2020:PRI

[WLC^{+}20a] Eric Ke Wang, Zuodong Liang, Chien-Ming Chen, Saru Kumari, and Muhammad Khurram Khan. PoRX: a reputation incentive scheme for blockchain consensus of IIoT. Future Generation Computer Systems, 102(??):140–151, January 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19310581.

Win:2020:FCI

[WLC^{+}20b] Khin Nandar Win, Kenli Li, Jianguo Chen, Philippe Fournier Viger, and Keqin Li. Fingerprint classification and identification algorithms for criminal investigation: a survey. Future Generation Computer Systems, 110(??):758–771, September 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19315109.

Wang:2023:DUA

[WLC23] Chenhao Wang, Ke Li, and Lisi Chen. Deep unified attention-based sequence modeling for online anomalous trajectory detection. Future Generation Computer Systems, 144(??):1– 11, July 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23000481.

Wang:2020:FHS

[WLD^{+}20a] En Wang, Dawei Li, Boxiang Dong, Huan Zhou, and Michelle Zhu. Flat and hierarchical system deployment for edge computing systems. Future Generation Computer Systems, 105(??):308–317, April 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19304042.

Wen:2020:SWP

[WLD^{+}20b] Yiping Wen, Jianxun Liu, Wanchun Dou, Xiaolong Xu, Buqing Cao, and Jinjun Chen. Scheduling workflows with privacy protection constraints for big data applications on cloud. Future Generation Computer Systems, 108(??): 1084–1091, July 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17307379.

Wang:2021:EEV

[WLL21] Bin Wang, Fagui Liu, and Weiwei Lin. Energyefficient VM scheduling based on deep reinforcement learning. Future Generation Computer Systems, 125(??):616–628, December 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21002855.

Wang:2022:SAD

[WLL22] Jiexi Wang, Yingxu Lai, and Jing Liu. Stealthy attack detection method based on multi-feature long short-term memory prediction model. Future Generation Computer Systems, 137(??):248–259, December 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22002424.

Wang:2024:DTAa

[WLL24] Jiadai Wang, Jingyi Li, and Jiajia Liu. Digital twinassisted flexible slice admission control for 5G core network: a deep reinforcement learning approach. Future Generation Computer Systems, 153(??):467–476, April 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23004806.

Wu:2020:DBP

[WLLC20] Budan Wu, Ran Liu, Rongheng Lin, and Junliang Chen. A distributed business process fragmentation method based on community discovery. Future Generation Computer Systems, 108(??):372–389, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19325580.

Wang:2020:KBD

[WLLF20] Jilong Wang, Renfa Li, Rui Li, and Bin Fu. A knowledge-based deep learning method for ECG signal delineation. Future Generation Computer Systems, 109(??):56–66, August 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1932360X.

Wang:2020:LMI

[WLLY20] Guan Wang, Jing Liu, Wei Lo, and Chun-Ming Yang. Learning multiple instance deep quality representation for robust object tracking. Future Generation Computer Systems, 113(??):298–303, December 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20310712.

Wang:2021:BEV

[WLN^{+}21] Jian Wang, Yongxin Liu, Shuteng Niu, Houbing Song, Weipeng Jing, and Jiawei Yuan. Blockchain enabled verification for cellular-connected unmanned aircraft system networking. Future Generation Computer Systems, 123(??):233–244, October 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21001461.

Wang:2020:CMA

[WLP^{+}20] Rui Wang, Miao Li, Limei Peng, Ying Hu, Mohammad Mehedi Hassan, and Abdulhameed Alelaiwi. Cognitive multi-agent empowering mobile edge computing for resource caching and collaboration. Future Generation Computer Systems, 102(??):66–74, January 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19318783.

Wu:2021:RHC

[WLR21] Cong Wu, Hongxin Li, and Jiajia Ren. Research on hierarchical clustering method based on partiallyordered Hasse graph. Future Generation Computer Systems, 125(??):785–791, December 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21002879.

Wang:2024:LGP

[WLS^{+}24] Pengfei Wang, Shiqi Li, Geng Sun, Changjun Zhou, Chengxi Gao, Sen Qiu, Tiwei Tao, and Qiang Zhang. Labeled graph partitioning scheme for distributed edge caching. Future Generation Computer Systems, 153(??):492–504, April 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23004715.

Wang:2021:TSS

[WLW^{+}21] Shudong Wang, Gaowei Liu, Xinzeng Wang, Yuanyuan Zhang, Sicheng He, and Yulin Zhang. TagSNP-set selection for genotyping using integrated data. Future Generation Computer Systems, 115(??):327–334, February 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20301072.

Wang:2024:SSI

[WLX^{+}24] Luying Wang, Anfeng Liu, Neal N. Xiong, Shaobo Zhang, Tian Wang, and Mianxiong Dong. SD-SRF: an intelligent service deployment scheme for serverlessoperated cloud-edge computing in 6G networks. Future Generation Computer Systems, 151(??):242–259, February 2024. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2300359X.

Wang:2020:IRN

[WLY^{+}20] Hongbing Wang, Jiajie Li, Qi Yu, Tianjing Hong, Jia Yan, and Wei Zhao. Integrating recurrent neural networks and reinforcement learning for dynamic service composition. Future Generation Computer Systems, 107(??):551–563, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19311240.

Wei:2023:CES

[WLY23] Wei Wei, Haoyi Li, and Weidong Yang. Costeffective stochastic resource placement in edge clouds with horizontal and vertical sharing. Future Generation Computer Systems, 138(??):213–225, January 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22002771.

Weng:2020:CMS

[WLYL20] Tien-Hsiung Weng, Kuan-Ching Li, Zhiliu Yang, and Chen Liu. On the code modernization of shared sampling alpha matting with OpenMP. Future Generation Computer Systems, 107(??):177–191, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19314116.

Wang:2020:ASB

[WLZ^{+}20] Jingjing Wang, Meiru Liu, Chunhui Zhang, Huaqiang Xu, Liren Zhang, and Yuefeng Zhao. An adaptive sparse Bayesian model combined with probabilistic label fusion for multiple sclerosis lesion segmentation in brain MRI. Future Generation Computer Systems, 105(??):695–704, April 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19325312.

Wang:2023:RLB

[WLZ^{+}23] Kan Wang, Xuan Liu, Hongfang Zhou, Dapeng Lan, Zhen Gao, Amir Taherkordi, Yujie Ye, and Yuan Gao. Reinforcement learning-based costefficient service function chaining with CoMP zeroforcing beamforming in edge networks. Future Generation Computer Systems, 141(??):355–368, April 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22003831.

Wei:2021:HOD

[WM21] Jianliang Wei and Fei Meng. How opinion distortion appears in superinfluencer dominated social network. Future Generation Computer Systems, 115(??):542–552, February 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20320021.

Wendzel:2022:ETD

[WMCH22] Steffen Wendzel, Wojciech Mazurczyk, Luca Caviglione, and Amir Houmansadr. Emerging topics in defending networked systems. Future Generation Computer Systems, 128(??):317–319, March 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21004003.

Wang:2020:BBD

[WMD^{+}20] Hao Wang, Shenglan Ma, Hong-Ning Dai, Muhammad Imran, and Tongsen Wang. Blockchain-based data privacy management with Nudge theory in open banking. Future Generation Computer Systems, 110(??):812–823, September 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18322702.

Wang:2021:ISU

[WML^{+}21] Huan Wang, Shahid Mumtaz, Houjun Li, JingXian Liu, and Fan Yang. An identification strategy for unknown attack through the joint learning of space–time features. Future Generation Computer Systems, 117(??):145–154, April 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20330478.

Wei:2023:EEF

[WML^{+}23] Shanming Wei, Shunmei Meng, Qianmu Li, Xiaokang Zhou, Lianyong Qi, and Xiaolong Xu. Edgeenabled federated sequential recommendation with knowledge-aware transformer. Future Generation Computer Systems, 148(??):610–622, November 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23002509.

Wang:2020:DDV

[WMNV20] Nan Wang, Michail Matthaiou, Dimitrios S. Nikolopoulos, and Blesson Varghese. DY-VERSE: DYnamic VERtical Scaling in multi-tenant Edge environments. Future Generation Computer Systems, 108(??):598–612, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19312403.

Wahab:2023:MLB

[WMU^{+}23] Haroon Wahab, Irfan Mehmood, Hassan Ugail, Arun Kumar Sangaiah, and Khan Muhammad. Machine learning based small bowel video capsule endoscopy analysis: Challenges and opportunities. Future Generation Computer Systems, 143(??): 191–214, June 2023. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23000171.

Wahab:2024:FDL

[WMU^{+}24] Haroon Wahab, Irfan Mehmood, Hassan Ugail, Javier Del Ser, and Khan Muhammad. Federated deep learning for wireless capsule endoscopy analysis: Enabling collaboration across multiple data centers for robust learning of diverse pathologies. Future Generation Computer Systems, 152(??):361–371, March 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23003801.

Wei:2022:SIAb

[WPPA22a] Wei Wei, Vincenzo Piuri, Witold Pedrycz, and Syed Hassan Ahmed. Special issue on Artificial Intelligence-of-Things (AIoT): Opportunities, challenges, and solutions — Part I: Artificial intelligence applications in various fields. Future Generation Computer Systems, 137(??):216–218, December 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22002527.

Wei:2022:SIAa

[WPPA22b] Wei Wei, Vincenzo Piuri, Witold Pedrycz, and Syed Hassan Ahmed. Special issue on Artificial Intelligence-of-Things (AIoT): Opportunities, challenges, and solutions — Part II: Artificial-intelligence-powered Internet of Things. Future Generation Computer Systems, 137(??): 186–188, December 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167739X22002539.

Wu:2023:TCE

[WPX^{+}23] Chunrong Wu, Qinglan Peng, Yunni Xia, Yong Jin, and Zhentao Hu. Towards cost-effective and robust AI microservice deployment in edge computing environments. Future Generation Computer Systems, 141(??):129–142, April 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22003314.

Wang:2020:PTP

[WQHX20] Qin Wang, Bo Qin, Jiankun Hu, and Fu Xiao. Preserving transaction privacy in bitcoin. Future Generation Computer Systems, 107(??):793–804, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17318393.

Wang:2023:OSC

[WS23] Xin Wang and Hong Shen. Online scheduling of coflows by attentionempowered scalable deep reinforcement learning. Future Generation Computer Systems, 146(??):195–206, September 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23001541.

Wang:2023:HBB

[WSC^{+}23] Taochun Wang, Huimin Shen, Jian Chen, Fulong Chen, Qingshan Wu, and Dong Xie. A hybrid blockchain-based identity authentication scheme for mobile crowd sensing. Future Generation Computer Systems, 143(??):40– 50, June 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23000201.

Wu:2022:AMT

[WSD^{+}22] Xiaocan Wu, Yu-E. Sun, Yang Du, Guoju Gao, He Huang, and Xiaoshuang Xing. An anti-malicious task allocation mechanism in crowdsensing systems. Future Generation Computer Systems, 127(??):347–361, February 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21003629.

Wu:2021:HSI

[WSJ^{+}21] Jimmy Ming-Tai Wu, Gautam Srivastava, Alireza Jolfaei, Philippe Fournier-Viger, and Jerry Chun-Wei Lin. Hiding sensitive information in eHealth datasets. Future Generation Computer Systems, 117(??): 169–180, April 2021. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20330594.

Wang:2021:BSA

[WSL21] Yongqiang Wang, Mei Sun, and Long Liu. Basketball shooting angle calculation and analysis by deeplylearned vision model. Future Generation Computer Systems, 125(??):949–953, December 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2100282X.

Wang:2023:DAA

[WSL^{+}23] Ying Wang, Fengjun Shang, Jianjun Lei, Xiangwei Zhu, Haoming Qin, and Jiayu Wen. Dual-attention assisted deep reinforcement learning algorithm for energy-efficient resource allocation in Industrial Internet of Things. Future Generation Computer Systems, 142(??):150–164, May 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22004137.

Wang:2021:IMS

[WSXL21] Zhixiao Wang, Chengcheng Sun, Jingke Xi, and Xiaocui Li. Influence maximization in social graphs based on community structure and node coverage gain. Future Generation Computer Systems, 118(??):327–338, May 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21000352.

Werner:2024:RAS

[WT24] Sebastian Werner and Stefan Tai. A reference architecture for serverless big data processing. Future Generation Computer Systems, 155(??):179–192, June 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24000360.

Wu:2020:DPA

[WTL^{+}20] Xiangqiong Wu, Guanghua Tan, Kenli Li, Shengli Li, Huaxuan Wen, Xianyi Zhu, and Wenli Cai. Deep parametric active contour model for neurofibromatosis segmentation. Future Generation Computer Systems, 112(??):58– 66, November 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20306191.

Wu:2022:FIR

[Wu22] Jia Wu. A fast-iterative reconstruction algorithm for sparse angle CT based on compressed sensing. Future Generation Computer Systems, 126(??):289–294, January 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21003216.

Wang:2020:DSA

[WW20] Kun Wang and Lei Wang. Detect Slitheen by analyzing the browsing behaviors and forcing retransmission. Future Generation Computer Systems, 106(??): 333–346, May 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19319314.

Wang:2024:HDS

[WWC^{+}24] Taochun Wang, Qingshan Wu, Jian Chen, Fulong Chen, Dong Xie, and Huimin Shen. Health data security sharing method based on hybrid blockchain. Future Generation Computer Systems, 153(??): 251–261, April 2024. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23004478.

Wang:2023:RFR

[WWF^{+}23] Zehao Wang, Huifeng Wu, Jin Fan, Danfeng Sun, and Jia Wu. A robust feature reinforcement framework for heterogeneous graphs neural networks. Future Generation Computer Systems, 141(??):143–153, April 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22003703.

Wang:2021:SSC

[WWH^{+}21] Shaoqiang Wang, Zhenzhen Wu, Gewen He, Shudong Wang, Hongwei Sun, and Fangfang Fan. Semisupervised classificationaware cross-modal deep adversarial data augmentation. Future Generation Computer Systems, 125(??):194–205, December 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21001795.

Wang:2021:NBI

[WWL21] Xiaonan Wang, Xingwei Wang, and Yanli Li. NDNbased IoT with edge computing. Future Generation Computer Systems, 115(??):397–405, February 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20303903.

Wei:2019:CEP

[WWP19] Li Wei, Chongling Wu, and Xie Peng. Clinical effect of percutaneous vertebroplasty (PVP) in spinal surgery on senile osteoporotic compressible fractures of the spine. Future Generation Computer Systems, 98(??):197–200, September 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18326359.

See retraction notice [WWP20].

Wei:2020:RNC

[WWP20] Li Wei, Chongling Wu, and Xie Peng. Retraction notice to “Clinical effect of percutaneous vertebroplasty (pvp) in spinal surgery on senile osteoporotic compressible fractures of the spine” [Future Gener. Comput. Syst. 98 (2019) 197– 200]. Future Generation Computer Systems, 107 (??):1149, June 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X2030652X.

See [WWP19].

Wu:2020:SDD

[WWS20] Di Wu, Hao Wang, and Razak Seidu. Smart data driven quality prediction for urban water source management. Future Generation Computer Systems, 107(??):418–432, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19302687.

Wang:2023:COK

[WWS^{+}23a] Kun Wang, Song Wu, Kun Suo, Yijie Liu, Hang Huang, Zhuo Huang, and Hai Jin. Characterizing and optimizing kernel resource isolation for containers. Future Generation Computer Systems, 141(??):218–229, April 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2200379X.

Wozniak:2023:BDN

[WWS23b] Marcin Woźniak, Michał Wieczorek, and Jakub Siłka. BiLSTM deep neural network model for imbalanced medical data of IoT systems. Future Generation Computer Systems, 141(??):489–499, April 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22004095.

Wu:2024:MOO

[WWW^{+}24] Dongkuo Wu, Xingwei Wang, Xueyi Wang, Min Huang, Rongfei Zeng, and Kaiqi Yang. Multiobjective optimization-based workflow scheduling for applications with data locality and deadline constraints in geo-distributed clouds. Future Generation Computer Systems, 157(??):485–498, August 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X2400133X.

Wang:2021:EIA

[WWY21] Fenghe Wang, Junquan Wang, and Wenfeng Yang. Efficient incremental authentication for the updated data in fog computing. Future Generation Computer Systems, 114(??):130–137, January 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20303988.

Wei:2024:DCB

[WWY^{+}24] Xiaohui Wei, Xiaonan Wang, Hengshan Yue, Nan Jiang, Jianpeng Zhao, and Meikang Qiu. DUAL-C: Building a “soft error efficient” on-the-fly compression mechanism for raw video data at edge devices. Future Generation Computer Systems, 153(??): 276–286, April 2024. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23004557.

Wei:2020:BDB

[WWZ^{+}20] PengCheng Wei, Dahu Wang, Yu Zhao, Sumarga Kumar Sah Tyagi, and Neeraj Kumar. Blockchain data-based cloud data integrity protection mechanism. Future Generation Computer Systems, 102(??):902–911, January 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19313494.

Wang:2023:DPD

[WWZW23] Yanling Wang, Qian Wang, Lingchen Zhao, and Cong Wang. Differential privacy in deep learning: Privacy and beyond. Future Generation Computer Systems, 148(??):408–424, November 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23002315.

Wang:2023:OOI

[WX23] Zhongyang Wang and Du Xu. Online optimization of intelligent reflecting surface-aided energyefficient IoT-edge computing. Future Generation Computer Systems, 141 (??):611–625, April 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0167739X22004125.

Wei:2024:HWO

[WX24] Yan Wei and Zhang Xingjun. A highly write-optimized concurrent B^{+}-tree for persistent memory. Future Generation Computer Systems, 155(??):219–230, June 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X24000530.

Wan:2024:ACA

[WXC^{+}24] Yi Wan, Xianzhong Xie, Junfan Chen, Kunpeng Xie, Dezhi Yi, Ye Lu, and Keke Gai. ADS-CNN: Adaptive dataflow scheduling for lightweight CNN accelerator on FPGAs. Future Generation Computer Systems, 158(??):138–149, September 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24001675.

Wang:2023:SBD

[WXD^{+}23] Ran Wang, Cheng Xu, Runshi Dong, Zhenghui Luo, Rong Zheng, and Xiaotong Zhang. A secured big-data sharing platform for materials genome engineering: State-of-the-art, challenges and architecture. Future Generation Computer Systems, 142(??):59– 74, May 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22004307.

Wu:2022:SHL

[WXS^{+}22] Xingjiao Wu, Luwei Xiao, Yixuan Sun, Junhang Zhang, Tianlong Ma, and Liang He. A survey of human-in-the-loop for machine learning. Future Generation Computer Systems, 135(??):364–381, October 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22001790.

Wang:2024:FMF

[WXX^{+}24] Chao Wang, Hui Xia, Shuo Xu, Hao Chi, Rui Zhang, and Chunqiang Hu. FedBnR: Mitigating federated learning non-IID problem by breaking the skewed task and reconstructing representation. Future Generation Computer Systems, 153(??):1– 11, April 2024. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23004247.

Wu:2023:BBA

[WXZ23] Nannan Wu, Lei Xu, and Liehuang Zhu. A blockchain based access control scheme with hidden policy and attribute. Future Generation Computer Systems, 141(??):186–196, April 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22003685.

Wang:2023:DNV

[WXZX23] Jin Wang, Hui Xiao, Shuwen Zhong, and Yinhao Xiao. DeepVulSeeker: a novel vulnerability identification framework via code graph structure and pre-training mechanism. Future Generation Computer Systems, 148(??):15– 26, November 2023. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23001978.

Wang:2022:JRM

[WXZZ22] Ge Wang, Fangmin Xu, Hengsheng Zhang, and Chenglin Zhao. Joint resource management for mobility supported federated learning in Internet of Vehicles. Future Generation Computer Systems, 129(??):199–211, April 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2100457X.

Wang:2020:EMC

[WYD20] Bang Wang, Qiang Yang, and Xianjun Deng. Energy management for cost minimization in green heterogeneous networks. Future Generation Computer Systems, 105(??):973–984, April 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17305599.

Wang:2024:DDA

[WYDB24] Zhousheng Wang, Geng Yang, Hua Dai, and Yunlu Bai. DAFL: Domain adaptation-based federated learning for privacypreserving biometric recognition. Future Generation Computer Systems, 150(??):436–450, January 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23003527.

Wei:2020:GSA

[WYG^{+}20] Xiaohui Wei, Hengshan Yue, Shang Gao, Lina Li, Ruyu Zhang, and Jingweijia Tan. G-SEAP: Analyzing and characterizing soft-error aware approximation in GPGPUs. Future Generation Computer Systems, 109(??):262–274, August 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19313214.

Wen:2021:BDD

[WYGP21] Chunhui Wen, Jinhai Yang, Liu Gan, and Yang Pan. Big data driven Internet of Things for credit evaluation and early warning in finance. Future Generation Computer Systems, 124(??):295–307, November 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21001977.

Wang:2021:IMC

[WYHM21] Guan Wang, Jiali Yin, M. Shamim Hossain, and Ghulam Muhammad. Incentive mechanism for collaborative distributed learning in Artificial Intelligence of Things. Future Generation Computer Systems, 125(??):376–384, December 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21002107.

Wang:2024:DTAb

[WYJ^{+}24] Kan Wang, Peng Yuan, Mian Ahmad Jan, Fazlullah Khan, Thippa Reddy Gadekallu, Saru Kumari, Hao Pan, and Lei Liu. Digital twin-assisted service function chaining in multidomain computing power networks with multi-agent reinforcement learning. Future Generation Computer Systems, 158(??):294–307, September 2024. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24001511.

Wang:2020:SSH

[WYS20] Liangmin Wang, Zhendong Yang, and Xiangmei Song. SHAMC: a secure and highly available database system in multicloud environment. Future Generation Computer Systems, 105(??):873–883, April 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17314759.

Wang:2022:FLB

[WYWS22] TianTian Wang, HaiLong Yu, KeChao Wang, and XiaoHong Su. Fault localization based on wide & deep learning model by mining software behavior. Future Generation Computer Systems, 127(??):309–319, February 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21003721.

Wang:2023:MSC

[WYX^{+}23a] Chaowei Wang, Xiaofei Yu, Lexi Xu, Ziye Wang, and Weidong Wang. Multimodal semantic communication accelerated bidirectional caching for 6G MEC. Future Generation Computer Systems, 140(??): 225–237, March 2023. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22003594.

Wang:2023:HHE

[WYX^{+}23b] Hao Wang, Ce Yu, Jian Xiao, Shanjiang Tang, Min Long, and Ming Zhu. HEGrid: a high efficient multi-channel radio astronomical data gridding framework in heterogeneous computing environments. Future Generation Computer Systems, 138(??):243–253, January 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22002886.

Wang:2020:EWR

[WYZ^{+}20] Ziyu Wang, Hui Yu, Zongyang Zhang, Jiaming Piao, and Jianwei Liu. ECDSA weak randomness in bitcoin. Future Generation Computer Systems, 102(??):507–513, January 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17330030.

Wang:2024:OLB

[WYZ^{+}24] Yuanzhang Wang, Fengkui Yang, Ke Zhou, Chunhua Li, Chong Liu, Ji Zhang, and Zhuo Cheng. An optimized learning-based directory placement policy with two-rounds selection in distributed file systems. Future Generation Computer Systems, 154(??):235–250, May 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23004727.

Wang:2020:NTM

[WZB^{+}20] Tian Wang, Guangxue Zhang, Md Zakirul Alam Bhuiyan, Anfeng Liu, Weijia Jia, and Mande Xie. A novel trust mechanism based on fog computing in sensor-cloud system. Future Generation Computer Systems, 109(??):573–582, August 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17323658.

Wu:2022:VGA

[WZC^{+}22] Chulin Wu, Heye Zhang, Jiaqi Chen, Zhifan Gao, Pengfei Zhang, Khan Muhammad, and Javier Del Ser. Vessel-GAN: Angiographic reconstructions from myocardial CT perfusion with explainable generative adversarial networks. Future Generation Computer Systems, 130(??):128–139, May 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2100488X.

Zhong:2019:SRN

[wZcZN^{+}19] Xiao wu Zhong, Ming cai Zhao, Guo Ning, Liu Meirong, and Guo Zixuan. Synthesis and research of a novel near infrared nanoprobe targeting liver cancer. Future Generation Computer Systems, 98(??):536–540, September 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18326438.

See retraction notice [wZcZN^{+}20].

Zhong:2020:RNS

[wZcZN^{+}20] Xiao wu Zhong, Ming cai Zhao, Guo Ning, Liu Meirong, and Guo Zixuan. Retraction notice to “Synthesis and research of a novel near infrared nanoprobe targeting liver cancer” [Future Gener. Comput. Syst. 98 (2019) 536–540]. Future Generation Computer Systems, 107(??):1147, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20306506.

See [wZcZN^{+}19].

Wu:2022:NBL

[WZH^{+}22] Yalin Wu, Qianjian Zhang, Yaqin Hu, Ko Sun-Woo, Xiangyan Zhang, Hongmin Zhu, Liu jie, and ShiYong Li. Novel binary logistic regression model based on feature transformation of XGBoost for type 2 Diabetes Mellitus prediction in healthcare systems. Future Generation Computer Systems, 129(??):1– 12, April 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21004325.

Wang:2023:ARD

[WZHX23] Yongkang Wang, Di-Hua Zhai, Yongping He, and Yuanqing Xia. An adaptive robust defending algorithm against backdoor attacks in federated learning. Future Generation Computer Systems, 143(??):118–131, June 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X2300033X.

Wei:2022:PDD

[WZJ^{+}22] Jia Wei, Xingjun Zhang, Zeyu Ji, Zheng Wei, and Jingbo Li. DPLRS: Distributed population learning rate schedule. Future Generation Computer Systems, 132(??):40–50, July 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22000449.

Wang:2020:DCW

[WZL^{+}20] Tian Wang, Jiandian Zeng, Yongxuan Lai, Yiqiao Cai, Hui Tian, Yonghong Chen, and Baowei Wang. Data collection from WSNs to the cloud based on mobile fog elements. Future Generation Computer Systems, 105(??):864–872, April 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17315315.

Wu:2022:AFL

[WZS^{+}22] Xiang Wu, Yongting Zhang, Minyu Shi, Pei Li, Ruirui Li, and Neal N. Xiong. An adaptive federated learning scheme with differential privacy preserving. Future Generation Computer Systems, 127(??):362–372, February 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21003617.

Wang:2023:SSE

[WZS^{+}23] Chen Wang, Tianqi Zhou, Jian Shen, Weizheng Wang, and Xiaokang Zhou. Searchable and secure edge precache scheme for intelligent 6G wireless systems. Future Generation Computer Systems, 140(??):129–137, March 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22003296.

Wu:2020:SPS

[WZT^{+}20] Fan Wu, Qinhua Zheng, Feng Tian, Zhihai Suo, Yuan Zhou, Kuo-Ming Chao, Mo Xu, Nazaraf Shah, Jun Liu, and Fei Li. Supporting povertystricken college students in smart campus. Future Generation Computer Systems, 111(??):599–616, October 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19300949.

Wang:2020:DTR

[WZTL20] Jing Wang, Huyin Zhang, Xing Tang, and Zongpeng Li. Delay-tolerant routing and message scheduling for CR–VANETs. Future Generation Computer Systems, 110(??):291–309, September 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19320503.

Wang:2020:ESW

[WZW^{+}20] Qin Wang, Haitao Zhao, Qianqian Wang, Haotong Cao, Gagangeet Singh Aujla, and Hongbo Zhu. Enabling secure wireless multimedia resource pricing using consortium blockchains. Future Generation Computer Systems, 110(??): 696–707, September 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167739X19315791.

Wang:2023:MTB

[WZW^{+}23] Kun Wang, Chen Zheng, Min Wang, Zhigang Chen, Zhipeng Zhao, and Lizhong Zhang. MAT-transformer-based state forecasting method for information devices. Future Generation Computer Systems, 147(??):360–370, October 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23001097.

Wu:2021:DCE

[WZX^{+}21] Kewei Wu, Yuming Zhang, Zhao Xie, Dan Guo, and Xin An. DDFPN: Context enhanced network for object detection. Future Generation Computer Systems, 124(??):133–141, November 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21001680.

Wang:2021:EVD

[WZXX21] Qiang Wang, Fucai Zhou, Jian Xu, and Zifeng Xu. Efficient verifiable databases with additional insertion and deletion operations in cloud computing. Future Generation Computer Systems, 115(??):553–567, February 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19321569.

Wan:2023:SEE

[WZZD23] Guojia Wan, Zhengyun Zhou, Zhigao Zheng, and Bo Du. Sub-entity embedding for inductive spatiotemporal knowledge graph completion. Future Generation Computer Systems, 148(??):240–249, November 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23002108.

Xiong:2020:SDD

[XCB^{+}20] Jinbo Xiong, Lei Chen, Md Zakirul Alam Bhuiyan, Chunjie Cao, Minshen Wang, Entao Luo, and Ximeng Liu. A secure data deletion scheme for IoT devices through key derivation encryption and data analysis. Future Generation Computer Systems, 111(??):741–753, October 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19314499.

Xin:2024:FGR

[XCGZ24] Ruyue Xin, Peng Chen, Paola Grosso, and Zhiming Zhao. A fine-grained robust performance diagnosis framework for runtime cloud applications. Future Generation Computer Systems, 155(??):300–311, June 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24000591.

Xia:2020:GBD

[XCH^{+}20] Xiaoyu Xia, Feifei Chen, Qiang He, Guangming Cui, Phu Lai, Mohamed Abdelrazek, John Grundy, and Hai Jin. Graph-based data caching optimization for edge computing. Future Generation Computer Systems, 113(??):228–239, December 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19330973.

Xu:2020:ILS

[XCL^{+}20] Rongbin Xu, Yongliang Cheng, Zhiqiang Liu, Ying Xie, and Yun Yang. Improved long short-term memory based anomaly detection with concept drift adaptive method for supporting IoT services. Future Generation Computer Systems, 112(??):228–242, November 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20302235.

Xing:2022:DDG

[XCS^{+}22] Yunlong Xing, Jiahao Cao, Kun Sun, Fei Yan, and Shengye Wan. The devil is in the detail: Generating system call whitelist for Linux seccomp. Future Generation Computer Systems, 135(??):105–113, October 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2200139X.

Xu:2020:OPA

[XCSF20] Jie Xu, Baojiang Cui, Ruisheng Shi, and Qingling Feng. Outsourced privacy-aware task allocation with flexible expressions in crowdsourcing. Future Generation Computer Systems, 112(??):383–393, November 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20305653.

Xu:2020:DRE

[XCW20] Ye Xu, Yun Chi, and Zhuo Wang. Dynamic robustness and efficiency on link-weighted scale-free networks by gray infos. Future Generation Computer Systems, 107(??):865–870, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17305630.

Xu:2022:ZTM

[XCZ^{+}22] Shiwei Xu, Xiaowen Cai, Yizhi Zhao, Zhengwei Ren, Le Du, Qin Wang, and Jianying Zhou. zkrpChain: Towards multiparty privacy-preserving data auditing for consortium blockchains based on zero-knowledge range proofs. Future Generation Computer Systems, 128(??):490–504, March 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21003800.

Xue:2023:IBE

[XCZ^{+}23] He Xue, Dajiang Chen, Ning Zhang, Hong-Ning Dai, and Keping Yu. Integration of blockchain and edge computing in Internet of Things: a survey. Future Generation Computer Systems, 144(??):307–326, July 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22003521.

Xie:2020:NTM

[XFJ^{+}20] Xia Xie, Yu Fu, Hai Jin, Yaliang Zhao, and Wenzhi Cao. A novel text mining approach for scholar information extraction from web content in Chinese. Future Generation Computer Systems, 111(??):859–872, October 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19305084.

Xiao:2020:EAT

[XGS^{+}20] Kaile Xiao, Zhipeng Gao, Weisong Shi, Xuesong Qiu, Yang Yang, and Lanlan Rui. EdgeABC: an architecture for task offloading and resource allocation in the Internet of Things. Future Generation Computer Systems, 107(??):498–508, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19323738.

Xia:2020:ENA

[XGX20] Junxu Xia, Deke Guo, and Junjie Xie. Efficient innetwork aggregation mechanism for data block repairing in data centers. Future Generation Computer Systems, 105(??):33– 43, April 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19310374.

Xiao:2023:MRA

[XGY^{+}23] Jianmao Xiao, Qinghang Gao, Zhenyue Yang, Yuanlong Cao, Hao Wang, and Zhiyong Feng. Multi-round auction-based resource allocation for edge computing: Maximizing social welfare. Future Generation Computer Systems, 140(??): 365–375, March 2023. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22003508.

Xue:2020:PHD

[XHW20] Dongliang Xue, Linpeng Huang, and Chentao Wu. A pure hardware-driven scheduler for enhancing bank-level parallelism in a persistent memory controller. Future Generation Computer Systems, 107 (??):383–393, June 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0167739X18332382.

Xu:2024:CDA

[XJL^{+}24] Zhuo Xu, Mengqing Jin, Jian Lin, Yuelong Liu, Jianlong Xu, Zhi Xiong, and Hao Cai. Connectiondensity-aware satellite-ground federated learning via asynchronous dynamic aggregation. Future Generation Computer Systems, 155 (??):312–323, June 2024. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0167739X24000517.

Xhafa:2020:EIS

[XKK20] Fatos Xhafa, Burak Kilic, and Paul Krause. Evaluation of IoT stream processing at edge computing layer for semantic data enrichment. Future Generation Computer Systems, 105(??):730–736, April 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19321296.

Xiao:2020:DIW

[XLCB20] Wenjing Xiao, Miao Li, Min Chen, and Ahmed Barnawi. Deep interaction: Wearable robotassisted emotion communication for enhancing perception and expression ability of children with Autism Spectrum Disorders. Future Generation Computer Systems, 108(??):709–716, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19328389.

Xie:2023:UUU

[XLG^{+}23] Zichao Xie, Zeyuan Li, Jinsong Gui, Anfeng Liu, Neal N. Xiong, and Shaobo Zhang. UWPEE: Using UAV and wavelet packet energy entropy to predict trafficbased attacks under limited communication, computing and caching for 6G wireless systems. Future Generation Computer Systems, 140(??):238–252, March 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22003260.

Xiao:2024:SMP

[XLH^{+}24] Ke Xiao, Jiayang Li, Yunhua He, Xu Wang, and Chao Wang. A secure multi-party payment channel on-chain and off-chain supervisable scheme. Future Generation Computer Systems, 154(??):330–343, May 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X2400013X.

Xu:2020:SEC

[XLL20a] Guangxia Xu, Weifeng Li, and Jun Liu. A social emotion classification approach using multimodel fusion. Future Generation Computer Systems, 102(??):347–356, January 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1930888X.

Xu:2020:FDP

[XLL^{+}20b] Yingying Xu, Zhi Liu, Yujun Li, Haixia Hou, Yankun Cao, Yuefeng Zhao, Wei Guo, and Lizhen Cui. Feature data processing: Making medical data fit deep neural networks. Future Generation Computer Systems, 109(??):149–157, August 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1932686X.

Xiao:2020:SDO

[XLLL20] Surong Xiao, Chubo Liu, Kenli Li, and Keqin Li. System delay optimization for mobile edge computing. Future Generation Computer Systems, 109(??):17– 28, August 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19310787.

Xiao:2022:ABD

[XLMC22] Weichu Xiao, Hongli Liu, Ziji Ma, and Weihong Chen. Attention-based deep neural network for driver behavior recognition. Future Generation Computer Systems, 132(??): 152–161, July 2022. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22000504.

Xie:2021:MSD

[XLS^{+}21] Xiaoyun Xie, Hui Liu, Minglei Shu, Qing Zhu, Anpeng Huang, Xiangpu Kong, and Yinglong Wang. A multi-stage denoising framework for ambulatory ECG signal based on domain knowledge and motion artifact detection. Future Generation Computer Systems, 116(??):103–116, March 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20330016.

Xiao:2021:HSP

[XLX^{+}21] Zhiyong Xiao, Xu Liu, Jingheng Xu, Qingxiao Sun, and Lin Gan. Highly scalable parallel genetic algorithm on Sunway manycore processors. Future Generation Computer Systems, 114(??):679–691, January 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19333187.

Xu:2022:PEB

[XLZ^{+}22] Ruihong Xu, Sihang Liu, Qingwang Zhang, Zemeng Yang, and Jianxiao Liu. PEWOBS: an efficient Bayesian network learning approach based on permutation and extensible ordering-based search. Future Generation Computer Systems, 128(??):505–520, March 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21004222.

Xu:2022:NIV

[XPT^{+}22] Yi Xu, Changgen Peng, Weijie Tan, Youliang Tian, Minyao Ma, and Kun Niu. Non-interactive verifiable privacy-preserving federated learning. Future Generation Computer Systems, 128(??):365–380, March 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21004131.

Xu:2024:PPP

[XQW^{+}24] Haitao Xu, Saiyu Qi, Junzhe Wei, Chenyang Li, Yong Qi, Wei Wei, and Yanan Qiao. POMF: a Privacy-preserved On-chain Matching Framework. Future Generation Computer Systems, 150(??):424–435, January 2024. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23003308.

Xie:2021:CAS

[XRHS21] Mande Xie, Yingying Ruan, Haibo Hong, and Jun Shao. A CP–ABE scheme based on multi-authority in hybrid clouds for mobile devices. Future Generation Computer Systems, 121(??):114–122, August 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21001023.

Xiong:2024:CCR

[XRZ^{+}24] Ruoting Xiong, Wei Ren, Shenghui Zhao, Jie He, Yi Ren, Kim-Kwang Raymond Choo, and Geyong Min. CoPiFL: a collusionresistant and privacy-preserving federated learning crowdsourcing scheme using blockchain and homomorphic encryption. Future Generation Computer Systems, 156(??):95–104, July 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24000876.

Xu:2021:STD

[XSW^{+}21] Peng Xu, Ruijie Sun, Wei Wang, Tianyang Chen, Yubo Zheng, and Hai Jin. SDD: a trusted display of FIDO2 transaction confirmation without trusted execution environment. Future Generation Computer Systems, 125(??):32–40, December 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21002296.

Xie:2023:ALM

[XTL^{+}23] Xiurui Xie, Maojun Tian, Guangchun Luo, Guisong Liu, Yizhe Wu, and Ke Qin. Active learning in multilabel image classification with graph convolutional network embedding. Future Generation Computer Systems, 148(??):56–65, November 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2300211X.

Xu:2021:CHA

[Xu21] Jianming Xu. Complicated human action understanding by massive-scale graph discovering technique. Future Generation Computer Systems, 123(??):68–72, October 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21001357.

Xu:2021:MVC

[XW21] Yuanjin Xu and Ming Wei.

Multi-view clustering toward aerial images by combining spectral analysis and local refinement. Future Generation Computer Systems, 117(??):138–144, April 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20330296.

Xiao:2023:EDR

[XW23] Shan Xiao and Chunyi Wu. Explore deep reinforcement learning for efficient task processing based on federated optimization in big data. Future Generation Computer Systems, 149(??):150–161, December 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23002492.

Xie:2020:RMV

[XWD20] Yanghao Xie, Sheng Wang, and Yueyue Dai. Revenuemaximizing virtualized network function chain placement in dynamic environment. Future Generation Computer Systems, 108(??):650–661, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19322836.

Xu:2021:CAR

[XWG^{+}21] Yueshen Xu, Yinchen Wu, Honghao Gao, Shengli Song, Yuyu Yin, and Xichu Xiao. Collaborative APIs recommendation for artificial intelligence of things with information fusion. Future Generation Computer Systems, 125(??):471–479, December 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21002582.

Xu:2021:IIC

[XWK21] Yuanjin Xu, Ming Wei, and M. M. Kamruzzaman. Inter/intra-category discriminative features for aerial image classification: a quality-aware selection model. Future Generation Computer Systems, 119 (??):77–83, June 2021. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20330399.

Xu:2020:HQT

[XWLC20] Zhiyi Xu, Yuzhe Wu, Qing Li, and Dingpan Chen. High-quality topological structures selection for smart city land spatial understanding and governance. Future Generation Computer Systems, 112(??):709–714, November 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20306014.

Xiao:2020:EPA

[XWM20] Lei Xiao, Shuangyan Wang, and Gang Mei. Efficient parallel algorithm for detecting influential nodes in large biological networks on the Graphics Processing Unit. Future Generation Computer Systems, 106 (??):1–13, May 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19318187.

Xue:2024:CGH

[XWR24] Weicheng Xue, Hongyu Wang, and Christopher J. Roy. CPU–GPU heterogeneous code acceleration of a finite volume Computational Fluid Dynamics solver. Future Generation Computer Systems, 158(??):367–377, September 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X24001900.

Xu:2020:PPD

[XWW^{+}20] Jian Xu, Laiwen Wei, Wei Wu, Andi Wang, Yu Zhang, and Fucai Zhou.

Privacy-preserving data integrity verification by using lightweight streaming authenticated data structures for healthcare cyberphysical system. Future Generation Computer Systems, 108(??):1287–1296, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17326067.

Xu:2024:GRL

[XWZM24] Jiawei Xu, Yufeng Wang, Bo Zhang, and Jianhua Ma. A graph reinforcement learning based SDN routing path selection for optimizing long-term revenue. Future Generation Computer Systems, 150(??):412–423, January 2024. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23003497.

Xu:2024:FPM

[XXL^{+}24] Shuo Xu, Hui Xia, Peishun Liu, Rui Zhang, Hao Chi, and Wei Gao. FLPM: a property modification scheme for data protection in federated learning. Future Generation Computer Systems, 154(??):151–159, May 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23004946.

Xu:2023:TES

[XXY^{+}23] Yueshen Xu, Weihao Xiao, Xiaoxian Yang, Rui Li, Yuyu Yin, and Zhiping Jiang. Towards effective semantic annotation for mobile and edge services for Internet-of-Things ecosystems. Future Generation Computer Systems, 139(??):64–73, February 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22003053.

Xuan:2020:DDP

[XY20] Wei Xuan and Guangqiang You. Detection and diagnosis of pancreatic tumor using deep learningbased hierarchical convolutional neural network on the Internet of Medical Things platform. Future Generation Computer Systems, 111(??):132–142, October 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20307779.

Xie:2024:NVC

[XYH^{+}24] Mande Xie, Xuekang Yang, Haibo Hong, Guiyi Wei, and Zhen Zhang. A novel verifiable Chinese multikeyword fuzzy rank searchable encryption scheme in cloud environments. Future Generation Computer Systems, 153(??):287–300, April 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23004259.

Xiao:2020:CHM

[XYL^{+}20] Feiyun Xiao, Decai Yang, Zhongming Lv, Xiaohui Guo, Zhengshi Liu, and Yong Wang. Classification of hand movements using variational mode decomposition and composite permutation entropy index with surface electromyogram signals. Future Generation Computer Systems, 110(??):1023–1036, September 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19319612.

Xu:2020:ISL

[XZ20] Bei Xu and Hai Zhuge. The influence of semantic link network on the ability of question-answering system. Future Generation Computer Systems, 108(??):1– 14, July 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19316127.

Xu:2024:EAP

[XZC^{+}24] Bo Xu, Haitao Zhao, Haotong Cao, Sahil Garg, Georges Kaddoum, and Mohammad Mehedi Hassan. Edge aggregation placement for semidecentralized federated learning in Industrial Internet of Things. Future Generation Computer Systems, 150(??):160–170, January 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23002923.

Xiong:2021:SSG

[XZD^{+}21] Fei Xiong, Yu Zheng, Weiping Ding, Hao Wang, Xinyi Wang, and Hongshu Chen. Selection strategy in graphbased spreading dynamics with limited capacity. Future Generation Computer Systems, 114(??):307–317, January 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20313194.

Xie:2023:TTE

[XZH^{+}23] Hui Xie, Jun Zheng, Teng He, Shengjun Wei, and Changzhen Hu. TEBDS: a Trusted Execution Environmentand-Blockchain-supported IoT data sharing system. Future Generation Computer Systems, 140(??): 321–330, March 2023. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22003326.

Xie:2020:EMM

[XZJ^{+}20] Guoqi Xie, Gang Zeng, Junqiang Jiang, Chunnian Fan, Renfa Li, and Keqin Li. Energy management for multiple realtime workflows on cyberphysical cloud systems. Future Generation Computer Systems, 105(??):916–931, April 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1731066X.

Xu:2020:BVM

[XZK^{+}20] Xiaolong Xu, Xuyun Zhang, Maqbool Khan, Wanchun Dou, Shengjun Xue, and Shui Yu. A balanced virtual machine scheduling method for energyperformance trade-offs in cyber-physical cloud systems. Future Generation Computer Systems, 105 (??):789–799, April 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0167739X17318927.

Xiong:2022:RTP

[XZTC22] Zhi Xiong, Min Zhao, Linhui Tan, and Lingru Cai. Real-time power optimization for application server clusters based on Mixed-Integer Programming. Future Generation Computer Systems, 137(??):260–273, December 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22002436.

Xiao:2022:PPW

[XZYH22] Yao Xiao, Amelie Chi Zhou, Xuan Yang, and Bingsheng He. Privacy-preserving workflow scheduling in geodistributed data centers. Future Generation Computer Systems, 130(??):46– 58, May 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21004854.

Xiong:2020:PCF

[XZZ^{+}20a] Ping Xiong, Lefeng Zhang, Tianqing Zhu, Gang Li, and Wanlei Zhou. Private collaborative filtering under untrusted recommender server. Future Generation Computer Systems, 109(??):511–520, August 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17320654.

Xu:2020:AMW

[XZZ^{+}20b] Chang Xu, Lvhan Zhang, Liehuang Zhu, Chuan Zhang, Xiaojiang Du, Mohsen Guizani, and Kashif Sharif. Aggregate in my way: Privacy-preserving data aggregation without trusted authority in ICN. Future Generation Computer Systems, 111(??):107–116, October 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19326056.

Yang:2021:NSD

[Yan21] Shuigen Yang. A novel study on deep learning framework to predict and analyze the financial time series information. Future Generation Computer Systems, 125(??):812–819, December 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2100279X.

Yu:2020:TIT

[YAZ^{+}20] Heejung Yu, Muhammad Khalil Afzal, Yousaf Bin Zikria, Abderrezak Rachedi, and Frank H. P. Fitzek. Tactile Internet: Technologies, test platforms, trials, and applications. Future Generation Computer Systems, 106(??):685–688, May 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20304192.

Yuan:2020:SCC

[YBC^{+}20] Shuai Yuan, Steven R. Brandt, Qin Chen, Ling Zhu, Reza Salatin, and Rion Dooley. A sustainable collaboratory for coastal resilience research. Future Generation Computer Systems, 111(??):786–792, October 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1930295X.

Yang:2023:NIS

[YBX^{+}23] Aimin Yang, Yunjie Bai, Tao Xue, Yifan Li, and Jie Li. A novel image steganography algorithm based on hybrid machine leaning and its application in cyberspace security. Future Generation Computer Systems, 145(??):293–302, August 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23001206.

Yakubu:2022:BBA

[YC22] Abukari Mohammed Yakubu and Yi Ping Phoebe Chen. A blockchain-based application for genomic access and variant discovery using smart contracts and homomorphic encryption. Future Generation Computer Systems, 137(??):234–247, December 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22002400.

Yang:2020:SLA

[YCG^{+}20] Guang Yang, Jun Chen, Zhifan Gao, Shuo Li, Hao Ni, Elsa Angelini, Tom Wong, Raad Mohiaddin, Eva Nyktari, Ricardo Wage, Lei Xu, Yanping Zhang, Xiuquan Du, Heye Zhang, David Firmin, and Jennifer Keegan. Simultaneous left atrium anatomy and scar segmentations via deep learning in multiview information with attention. Future Generation Computer Systems, 107(??):215–228, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19312051.

Yue:2020:PIS

[YCS^{+}20] Songshan Yue, Min Chen, Jie Song, Wenping Yuan, Tiexi Chen, Guonian Lü, Chaoran Shen, Zaiyang Ma, Kai Xu, Yongning Wen, and Hongquan Song. Participatory intercomparison strategy for terrestrial carbon cycle models based on a service-oriented architecture. Future Generation Computer Systems, 112(??):449–466, November 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20304246.

Yao:2023:POS

[YCYO23] Xuyi Yao, Ningjiang Chen, Xuemei Yuan, and Pingjie Ou. Performance optimization of serverless edge computing function offloading based on deep reinforcement learning. Future Generation Computer Systems, 139(??):74–86, February 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X2200293X.

Yurekten:2021:SBC

[YD21] Ozgur Yurekten and Mehmet Demirci. SDN-based cyber defense: a survey. Future Generation Computer Systems, 115(??):126–149, February 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20303277.

Yeh:2020:SIF

[YDK20] Kuo-Hui Yeh, Robert H.

Deng, and Hiroaki Kikuchi. Special issue on FinTech security and privacy. Future Generation Computer Systems, 112(??):1172–1173, November 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20324158.

Yu:2020:PNB

[YDL^{+}20] Wangyang Yu, Zhijun Ding, Lu Liu, Xiaoming Wang, and Richard David Crossley. Petri netbased methods for analyzing structural security in e-commerce business processes. Future Generation Computer Systems, 109(??):611–620, August 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17323671.

Yu:2024:ESB

[YFL^{+}24] Wangyang Yu, Jie Feng, Lu Liu, Xiaojun Zhai, and Yumeng Cheng. Enhancing security in e-business processes: Utilizing dynamic slicing of Colored Petri Nets for logical vulnerability detection. Future Generation Computer Systems, 158(??):210–218, September 2024. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X2400164X.

Yan:2022:REM

[YFQ^{+}22] He Yan, Liyong Fu, Yong Qi, Dong-Jun Yu, and Qiaolin Ye. Robust ensemble method for short-term traffic flow prediction. Future Generation Computer Systems, 133(??):395–410, August 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22001157.

Yangui:2021:FGS

[YGD^{+}21] Sami Yangui, Andrzej Goscinski, Khalil Drira, Zahir Tari, and Djamal Benslimane. Future generation of service-oriented computing systems. Future Generation Computer Systems, 118(??):252–256, May 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21000297.

Yaqoob:2021:MHF

[YGE21] Mahnoor Yaqoob, Orhan Gemikonakli, and Enver Ever. Modelling heterogeneous future wireless cellular networks: an analytical study for interaction of 5G femtocells and macro-cells. Future Generation Computer Systems, 114(??):82– 95, January 2021. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20306282.

Yang:2024:GNB

[YGP^{+}24] Xiangyang Yang, Xuefeng Guan, Zhaoxing Pang, Xing Kui, and Huayi Wu. GridMesa: a NoSQLbased big spatial data management system with an adaptive grid approximation model. Future Generation Computer Systems, 155(??):324–339, June 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24000542.

Yin:2021:NMT

[YGR21] Xiaoyu Yin, Dagmar Gromann, and Sebastian Rudolph. Neural machine translating from natural language to SPARQL. Future Generation Computer Systems, 117(??):510–519, April 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20330752.

Yin:2022:MSF

[YGS^{+}22] Meng-Meng Yin, Ying-Lian Gao, Junliang Shang, Chun-Hou Zheng, and Jin-Xing Liu. Multi-similarity fusion-based label propagation for predicting microbes potentially associated with diseases. Future Generation Computer Systems, 134(??):247–255, September 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22001352.

Yang:2020:AMD

[YHC20] Ta-Wei Yang, Yu-Han Ho, and Cheng-Fu Chou. Achieving M2M-device authentication through heterogeneous information bound with USIM card. Future Generation Computer Systems, 110(??):629–637, September 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19307228.

Yang:2022:DBF

[YHC^{+}22] Bo Yang, Xuelin Huang, Weizheng Cheng, Tao Huang, and Xu Li. Discrete bacterial foraging optimization for community detection in networks. Future Generation Computer Systems, 128(??):192–204, March 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21004052.

Yang:2022:MAC

[YhSL^{+}22] Yifan Yang, Run hua Shi, Kunchang Li, Zhiwei Wu, and Shuhao Wang. Multiple access control scheme for EHRs combining edge computing with smart contracts. Future Generation Computer Systems, 129(??):453–463, April 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21004313.

Yang:2020:ORT

[YHW^{+}20] Jiachen Yang, Yurong Han, Yafang Wang, Bin Jiang, Zhihan Lv, and Houbing Song. Optimization of realtime traffic network assignment based on IoT data using DBN and clustering model in smart city. Future Generation Computer Systems, 108(??):976–986, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17310609.

Yi:2021:SHO

[YJB^{+}21] Chunzhi Yi, Feng Jiang, Md Zakirul Alam Bhuiyan, Chifu Yang, Xianzhong Gao, Hao Guo, Jiantao Ma, and Shen Su. Smart healthcare-oriented online prediction of lower-limb kinematics and kinetics based on data-driven neural signal decoding. Future Generation Computer Systems, 114(??):96–105, January 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20307627.

Yu:2020:MAM

[YJF^{+}20] Wangyang Yu, Menghan Jia, Xianwen Fang, Yao Lu, and Jianchun Xu. Modeling and analysis of medical resource allocation based on timed colored Petri net. Future Generation Computer Systems, 111(??):368–374, October 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19334028.

Yang:2020:IRL

[YJH^{+}20] Yikun Yang, Shengjie Jiao, Jinrong He, Bisheng Xia, Jiabo Li, and Ru Xiao. Image retrieval via learning content-based deep quality model towards big data. Future Generation Computer Systems, 112(??):243–249, November 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20301424.

Yang:2020:TSA

[YJLC20] Jiachen Yang, Bin Jiang, Zhihan Lv, and Kim-Kwang Raymond Choo. A task scheduling algorithm considering game theory designed for energy management in cloud computing. Future Generation Computer Systems, 105 (??):985–992, April 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167739X17304673.

Yu:2023:MSI

[YJQ^{+}23] Runde Yu, Fusheng Jin, Zhuang Qiao, Ye Yuan, and Guoren Wang. Multi-scale image–text matching network for scene and spatiotemporal images. Future Generation Computer Systems, 142(??):292–300, May 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23000122.

Yoon:2020:PAB

[YK20a] Su-Kyung Yoon and Shin-Dug Kim. Pattern analysis based data management method and memory-disk integrated system for high performance computing. Future Generation Computer Systems, 106(??): 185–198, May 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19320655.

Yu:2020:OUR

[YK20b] Heejung Yu and Taejoon Kim. Optimization of uplink rate and fronthaul compression in cloud radio access networks. Future Generation Computer Systems, 102(??):465–471, January 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18318867.

Yan:2020:RDP

[YL20a] Zhiwei Yan and Jong-Hyouk Lee. The road to DNS privacy. Future Generation Computer Systems, 112(??):604–611, November 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20307068.

Yu:2020:PLS

[YL20b] Heejung Yu and Il-Gu Lee. Physical layer security based on NOMA and AJ for MISOSE channels with an untrusted relay. Future Generation Computer Systems, 102(??):611–618, January 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19308386.

Yu:2023:EMS

[YLC23] Xiaofeng Yu, Jia Liu, and Zhanqi Cui. On efficient matching of spatiotemporal rules. Future Generation Computer Systems, 146(??):250–259, September 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23001632.

Yang:2023:PUV

[YLD^{+}23] Yuanfeng Yang, Lin Li, Husheng Dong, Gang Liu, Xun Sun, and Zhaobin Liu. Progressive unsupervised video person reidentification with accumulative motion and tracklet spatial–temporal correlation. Future Generation Computer Systems, 142(??):90–100, May 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22004277.

Ye:2023:NSC

[YLF^{+}23] Guodong Ye, Xin Liu, Siqi Fan, Yuan Tan, Qingguo Zhou, Rui Zhou, and Xiaokang Zhou. Novel supply chain vulnerability detection based on heterogeneous-graph-driven hash similarity in IoT. Future Generation Computer Systems, 148(??):201–210, November 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23002194.

Yuan:2024:CED

[YLG^{+}24] Zhu Yuan, Xueqiang Lv, Yunchao Gong, Ping Xie, Taifu Yuan, and Xindong You. Cost-effective data classification storage through text seasonal features. Future Generation Computer Systems, 158(??):472–487, September 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24002085.

Yang:2021:GCG

[YLGG21] Liang Yang, Weixun Li, Yuanfang Guo, and Junhua Gu. Graph-CAT: Graph co-attention networks via local and global attribute augmentations. Future Generation Computer Systems, 118(??): 170–179, May 2021. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20330880.

Yang:2021:MST

[YLGZ21] Jiaqi Yang, Yongjun Li, Congjie Gao, and Yinyin Zhang. Measuring the short text similarity based on semantic and syntactic information. Future Generation Computer Systems, 114(??):169–180, January 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20306671.

Yang:2023:SCF

[YLH^{+}23] Pan Yang, Hongbo Li, Jianye Huang, Hao Zhang, Man Ho Allen Au, and Qiong Huang. Secure channel free public key authenticated encryption with multi-keyword search on healthcare systems. Future Generation Computer Systems, 145(??):511–520, August 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23000791.

Yang:2020:SVA

[YLKK20] Yousung Yang, Joohyung Lee, Nakyoung Kim, and Kwihoon Kim. Socialviewport adaptive caching scheme with clustering for virtual reality streaming in an edge computing platform. Future Generation Computer Systems, 108(??):424–431, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19318242.

Yang:2022:VAB

[YLL22] Yousung Yang, Seongsoo Lee, and Joohyung Lee. Video analyticsbased real-time intelligent crossing detection system (RICDS): Killer app for edge computing. Future Generation Computer Systems, 133(??):84–94, August 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22000887.

Yang:2023:FLB

[YLM23] Yuting Yang, Yue Lu, and Gang Mei. A federated learning based approach for predicting landslide displacement considering data security. Future Generation Computer Systems, 149(??):184–199, December 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23002662.

Yang:2021:CEC

[YLS21] Kai Yang, Xiaodong Lin, and Limin Sun. CShield: Enabling code privacy for cyber–physical systems. Future Generation Computer Systems, 125(??):564–574, December 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21002466.

Yan:2022:BAF

[YLSL22a] Binghao Yan, Qinrang Liu, JianLiang Shen, and Dong Liang. BatchUp: Achieve fast TCAM update with batch processing optimization in SDN. Future Generation Computer Systems, 134(??):93–106, September 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22001236.

Yan:2022:FLM

[YLSL22b] Binghao Yan, Qinrang Liu, JianLiang Shen, and Dong Liang. Flowlet-level multipath routing based on graph neural network in OpenFlow-based SDN. Future Generation Computer Systems, 134(??):140–153, September 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22001297.

Yao:2022:EEA

[YLTH22] Chunrong Yao, Wantao Liu, Weiqing Tang, and Songlin Hu. EAIS: Energyaware adaptive scheduling for CNN inference on highperformance GPUs. Future Generation Computer Systems, 130(??):253–268, May 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22000127.

Yu:2023:RTA

[YLX^{+}23] Jiayi Yu, Zeyuan Li, Naixue Xiong, Shaobo Zhang, Anfeng Liu, and Athanasios V. Vasilakos. A reliability and truth-aware based online digital data auction mechanism for cybersecurity in MCS. Future Generation Computer Systems, 141(??):526–541, April 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22003995.

Yao:2023:PGG

[YLY^{+}23] Zhexin Yao, Qiuming Liu, Jingkang Yang, Yanan Chen, and Zhen Wu. PPUP-GAN: a GAN-based privacy-protecting method for aerial photography. Future Generation Computer Systems, 145(??):284–292, August 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23001085.

Ye:2024:EFI

[YLZ^{+}24] Zhiwei Ye, Jun Luo, Wen Zhou, Mingwei Wang, and Qiyi He. An ensemble framework with improved hybrid breeding optimization-based feature selection for intrusion detection. Future Generation Computer Systems, 151(??):124–136, February 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23003667.

Yang:2021:CBP

[YLZL21] Haomiao Yang, Shaopeng Liang, Yi Zhang, and Xiong Li. Cloud-based privacy- and integrityprotecting density peaks clustering. Future Generation Computer Systems, 125(??):758–769, December 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21002892.

Yahyaoui:2022:TBM

[YMAAH22] Hamdi Yahyaoui, Zakaria Maamar, Mohamed Alkhafajiy, and Hamid Al-Hamadi. Trust-based management in IoT federations. Future Generation Computer Systems, 136(??):182–192, November 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22002023.

Yasar:2020:SIE

[YMS20a] Ansar-Ul-Haque Yasar, Haroon Malik, and Elhadi M. Shakshuki. Special issue on emerging trends in ubiquitous systems and pervasive networks. Future Generation Computer Systems, 110(??):317–319, September 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20313674.

Yasar:2020:SIF

[YMS20b] Ansar-Ul-Haque Yasar, Haroon Malik, and Elhadi M. Shakshuki. Special issue on future networks and pervasive systems. Future Generation Computer Systems, 102(??):950–951, January 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19321466.

Yun:2021:HCI

[YMY21] Yuliang Yun, Dexin Ma, and Meihong Yang. Human– computer interaction-based decision support system with applications in data mining. Future Generation Computer Systems, 114(??):285–289, January 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20312450.

Yun:2020:ETD

[YNK^{+}20] Unil Yun, Hyoju Nam, Jongseong Kim, Heonho Kim, Yoonji Baek, Judae Lee, Eunchul Yoon, Tin Truong, Bay Vo, and Witold Pedrycz. Efficient transaction deleting approach of pre-large based high utility pattern mining in dynamic databases. Future Generation Computer Systems, 103(??):58– 78, February 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19313937.

Yaacoub:2020:SIM

[YNN^{+}20] Jean-Paul A. Yaacoub, Mohamad Noura, Hassan N. Noura, Ola Salman, Elias Yaacoub, Raphaël Couturier, and Ali Chehab. Securing Internet of Medical Things systems: Limitations, issues and recommendations. Future Generation Computer Systems, 105 (??):581–606, April 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0167739X19305680.

Yungaicela-Naula:2023:SNB

[YNVRPD23] Noe M. Yungaicela-Naula, Cesar Vargas-Rosales, and Jesús A. Pérez-Díaz. SDN/ NFV-based framework for autonomous defense against slow-rate DDoS attacks by using reinforcement learning. Future Generation Computer Systems, 149(??):637–649, December 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23003047.

Ylipulli:2023:PLP

[YPEK23] Johanna Ylipulli, Matti Pouke, Nils Ehrenberg, and Turkka Keinonen. Public libraries as a partner in digital innovation project: Designing a virtual reality experience to support digital literacy. Future Generation Computer Systems, 149(??):594–605, December 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23003011.

Yang:2024:FCQ

[YPL24] Hao Yang, Li Pan, and Shijun Liu. Faster or cheaper: a Q-learning based costeffective mixed cluster scaling method for achieving low tail latencies. Future Generation Computer Systems, 157(??):264–274, August 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2400116X.

You:2020:IAA

[YPX^{+}20] Lan You, Qingxi Peng, Zenggang Xiong, Du He, Meikang Qiu, and Xuemin Zhang. Integrating aspect analysis and local outlier factor for intelligent review spam detection. Future Generation Computer Systems, 102(??):163–172, January 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19307551.

Yao:2024:KGE

[YPZ^{+}24] Xingyi Yao, Xiaogang Pan, Tao Zhang, Wenhua Li, and Jianjiang Wang. Knowledge-guided evolutionary algorithm for multisatellite resource scheduling optimization. Future Generation Computer Systems, 156(??):130–141, July 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24000773.

Qian:2020:MOE

[yQhJL20] Si yuan Qian, Zhao hong Jia, and Kai Li. A multiobjective evolutionary algorithm based on adaptive clustering for energyaware batch scheduling problem. Future Generation Computer Systems, 113(??):441–453, December 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19323593.

Yan:2022:EML

[YRV^{+}22] Runze Yan, Whitney R.

Ringwald, Julio Vega, Madeline Kehl, Sang Won Bae, Anind K. Dey, Carissa A. Low, Aidan G. C. Wright, and Afsaneh Doryab. Exploratory machine learning modeling of adaptive and maladaptive personality traits from passively sensed behavior. Future Generation Computer Systems, 132(??):266–281, July 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2200053X.

You:2022:MES

[YSL^{+}22] Xindong You, Dawei Sun, Xueqiang Lv, Shang Gao, and Rajkumar Buyya. MQDS: an energy saving scheduling strategy with diverse QoS constraints towards reconfigurable cloud storage systems. Future Generation Computer Systems, 129(??):252–268, April 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21004696.

Yan:2024:HET

[YSL^{+}24] Kaige Yan, Yanshuang Song, Tao Liu, Jingweijia Tan, Xiaohui Wei, and Xin Fu. HSAS: Efficient task scheduling for large scale heterogeneous systolic array accelerator cluster. Future Generation Computer Systems, 154(??):440–450, May 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24000220.

Yao:2019:HRU

[YTQ19] Qingrong Yao, Wu Tian, and Li Qiu. High-resolution ultrasound images in gouty arthritis to evaluate relationship between tophi and bone erosion. Future Generation Computer Systems, 98(??):131–134, September 2019. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1831851X.

See retraction notices [YTQ20a, YWG^{+}19].

Yao:2020:RNHa

[YTQ20a] Qingrong Yao, Wu Tian, and Li Qiu. Retraction notice to “High-resolution ultrasound images in gouty arthritis to evaluate relationship between tophi and bone erosion” [future gener. comput. syst. 98 (2018) 131–134]. Future Generation Computer Systems, 106(??):689, May 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20305689.

See [YTQ19].

Yao:2020:RNHb

[YTQ20b] Qingrong Yao, Wu Tian, and Li Qiu. Retraction notice to “High-resolution ultrasound Images in gouty arthritis to Evaluate Relationship between tophi and bone erosion” [Future Gener. Comput. Syst. 98 (2019) 131–134]. Future Generation Computer Systems, 111(??):938, October 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19332984.

See [YTQ19].

Ye:2020:EDA

[YTW^{+}20] Xiaochun Ye, Xu Tan, Meng Wu, Yujing Feng, Da Wang, Hao Zhang, Songwen Pei, and Dongrui Fan. An efficient dataflow accelerator for scientific applications. Future Generation Computer Systems, 112(??):580–588, November 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19313986.

Yu:2021:SBD

[Yu21] Linchen Yu. SVC-based dynamic caching for smart media streaming over the Internet of Things. Future Generation Computer Systems, 114(??):219–228, January 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19305941.

Yuen:2020:PPA

[Yue20] Tsz Hon Yuen. PAChain: Private, authenticated & auditable consortium blockchain and its implementation. Future Generation Computer Systems, 112(??):913–929, November 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19328547.

Yang:2022:LBP

[YVSG22] Huijie Yang, Pandi Vijayakumar, Jian Shen, and Brij B. Gupta. A locationbased privacy-preserving oblivious sharing scheme for indoor navigation. Future Generation Computer Systems, 137(??):42–52, December 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22002278.

Yun:2020:SMM

[YVW^{+}20] Joseph T. Yun, Nickolas Vance, Chen Wang, Luigi Marini, Joseph Troy, Curtis Donelson, Chieh-Li Chin, and Mark D. Henderson. The Social Media Macroscope: a science gateway for research using social media data. Future Generation Computer Systems, 111(??):819–828, October 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19304443.

Yan:2021:FDC

[YW21] Shankai Yan and Ka-Chun Wong. Future DNA computing device and accompanied tool stack: Towards highthroughput computation. Future Generation Computer Systems, 117(??): 111–124, April 2021. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20330223.

Yang:2023:LTE

[YWDC23] Shangdong Yang, Huihui Wang, Shaokang Dong, and Xingguo Chen. Leveraging transition exploratory bonus for efficient exploration in hard-transiting reinforcement learning problems. Future Generation Computer Systems, 145(??):442–453, August 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X2300136X.

Yanli:2019:IKG

[YWG^{+}19] Ji Yanli, Lu Wei, Che Guowei, Yang Mei, and Liu Lunxu. Inhibition of KRAS gene mutation on non-small cell lung cancer and its effect on circulating tumor cells. Future Generation Computer Systems, 98(??):104–108, September 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18326426.

See retraction notices [YWG^{+}20b, JLC^{+}20].

Yang:2020:AEE

[YWG^{+}20a] Jun Yang, Rui Wang, Xin Guan, Mohammad Mehedi Hassan, Ahmad Almogren, and Ahmed Alsanad. AI-enabled emotionaware robot: the fusion of smart clothing, edge clouds and robotics. Future Generation Computer Systems, 102(??):701–709, January 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19305643.

Yanli:2020:RNI

[YWG^{+}20b] Ji Yanli, Lu Wei, Che Guowei, Yang Mei, and Liu Lunxu. Retraction notice to “Inhibition of KRAS gene mutation on non-small cell lung cancer and its effect on circulating tumor cells” [Future Gener. Comput. Syst. 98 (2019) 104– 108]. Future Generation Computer Systems, 107 (??):1143, June 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20306361.

See [YWG^{+}19].

Yu:2021:RIS

[YWH^{+}21] Gang Yu, Yi Wang, Min Hu, Lihua Shi, Zeyu Mao, and Vijayan Sugumaran. RIOMS: an intelligent system for operation and maintenance of urban roads using spatio-temporal data in smart cities. Future Generation Computer Systems, 115(??):583–609, February 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X2030501X.

Yao:2023:EEL

[YWH^{+}23] Wenbin Yao, Zhuqing Wang, Yingying Hou, Xikang Zhu, Xiaoyong Li, and Yamei Xia. An energyefficient load balance strategy based on virtual machine consolidation in cloud environment. Future Generation Computer Systems, 146(??):222–233, September 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23001498.

Yang:2021:VQE

[YWS21] Wei Yang, Jian Wang, and Jinlong Shi. Video quality evaluation toward complicated sport activities for clustering analysis. Future Generation Computer Systems, 119(??):43– 49, June 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21000285.

Yao:2020:JOF

[YXL^{+}20] Hong Yao, Muzhou Xiong, Hui Li, Lin Gu, and Deze Zeng. Joint optimization of function mapping and preemptive scheduling for service chains in network function virtualization. Future Generation Computer Systems, 108(??):1112–1118, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17324305.

Yang:2021:DFI

[YXL^{+}21] Jiachen Yang, Shuai Xiao, Aiyun Li, Guipeng Lan, and Huihui Wang. Detecting fake images by identifying potential texture difference. Future Generation Computer Systems, 125(??):127–135, December 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21002387.

Yang:2020:WHF

[YXLB20] Jun Yang, Wenjing Xiao, Huimin Lu, and Ahmed Barnawi. Wireless highfrequency NLOS monitoring system for heart disease combined with hospital and home. Future Generation Computer Systems, 110(??):772–780, September 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19323246.

Yang:2023:ABS

[YXS23a] Xu Yang, Hongyan Xing, and Xin Su. AI-based sound source localization system with higher accuracy. Future Generation Computer Systems, 141 (??):1–15, April 2023. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22003387.

Yuan:2023:EEI

[YXS^{+}23b] Jingling Yuan, Hua Xiao, Zhishu Shen, Tiehua Zhang, and Jiong Jin. ELECT : Energy-efficient intelligent edge–cloud collaboration for remote IoT services. Future Generation Computer Systems, 147(??):179–194, October 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23001681.

Yi:2020:TOS

[YXYH20] Na Yi, Jianjun Xu, Limei Yan, and Lin Huang. Task optimization and scheduling of distributed cyberphysical system based on improved ant colony algorithm. Future Generation Computer Systems, 109(??):134–148, August 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19327608.

Yuan:2021:FPP

[YYB^{+}21] Yachao Yuan, Yali Yuan, Thar Baker, Lutz Maria Kolbe, and Dieter Hogrefe. FedRD: Privacy-preserving adaptive Federated learning framework for intelligent hazardous Road Damage detection and warning. Future Generation Computer Systems, 125(??):385–398, December 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21002302.

Yun:2020:APB

[YYKK20] Ji-Tae Yun, Su-Kyung Yoon, Jeong-Geun Kim, and Shin-Dug Kim. Access pattern-based highperformance main memory system for graph processing on single machines. Future Generation Computer Systems, 108(??):560–573, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19323945.

Yu:2022:HPI

[YYL22] Jing Yu, Xiaojun Ye, and Hongbo Li. A high precision intrusion detection system for network security communication based on multi-scale convolutional neural network. Future Generation Computer Systems, 129(??):399–406, April 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21004143.

Yang:2020:BVP

[YYN^{+}20] Xuechao Yang, Xun Yi, Surya Nepal, Andrei Kelarev, and Fengling Han. Blockchain voting: Publicly verifiable online voting protocol without trusted tallying authorities. Future Generation Computer Systems, 112(??):859–874, November 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17327656.

Yu:2021:EDM

[YYW^{+}21] Haiyang Yu, Zhen Yang, Muhammad Waqas, Shanshan Tu, Zhu Han, Zahid Halim, Richard O. Sinnott, and Udaya Parampalli. Efficient dynamic multi-replica auditing for the cloud with geographic location. Future Generation Computer Systems, 125(??):285–298, December 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21002004.

Yang:2024:PPP

[YYW^{+}24] Mengda Yang, Wenzhe Yi, Juan Wang, Hongxin Hu, Xiaoyang Xu, and Ziang Li. Penetralium : Privacypreserving and memoryefficient neural network inference at the edge. Future Generation Computer Systems, 156(??):30–41, July 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24000797.

Yang:2023:LAW

[YYXZ23] Liwen Yang, Lingjuan Ye, Yuanqing Xia, and Yufeng Zhan. Look-ahead workflow scheduling with width changing trend in clouds. Future Generation Computer Systems, 139(??):139–150, February 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22002965.

Yuan:2023:DBA

[YYY^{+}23] Lixiang Yuan, Siyang Yu, Zhibang Yang, Mingxing Duan, and Kenli Li. A data balancing approach based on generative adversarial network. Future Generation Computer Systems, 141(??):768–776, April 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22004289.

Yu:2024:WLA

[YYZ^{+}24] Zhiwang Yu, Chaoshu Yang, Runyu Zhang, Pengpeng Tian, Xianyu He, Lening Zhou, Hui Li, and Duo Liu. Wear-leveling-aware buddy-like memory allocator for persistent memory file systems. Future Generation Computer Systems, 150(??):37– 48, January 2024. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23003114.

Yang:2020:OSM

[YZC^{+}20] Chaoshu Yang, Qingfeng Zhuge, Xianzhang Chen, Edwin H.-M. Sha, Duo Liu, and Runyu Zhang. Optimizing synchronization mechanism for blockbased file systems using persistent memory. Future Generation Computer Systems, 111(??):288–299, October 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19316504.

Yuan:2020:HPH

[YZJ^{+}20] Bin Yuan, Deqing Zou, Hai Jin, Shui Yu, and Laurence T. Yang. Host-Watcher: Protecting hosts in cloud data centers through software-defined networking. Future Generation Computer Systems, 105(??):964–972, April 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167739X17306477.

Yu:2020:SGN

[YZL^{+}20] Mei Yu, Zhuo Zhang, Xuewei Li, Jian Yu, Jie Gao, Zhiqiang Liu, Bo You, Xiaoshan Zheng, and Ruiguo Yu. Superposition graph neural network for offshore wind power prediction. Future Generation Computer Systems, 113(??):145–157, December 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19330626.

Yang:2023:JHA

[YZL^{+}23] Zhao Yang, Shengbing Zhang, Chuxi Li, Miao Wang, Jiaying Yang, and Meng Zhang. Joint heterogeneity-aware personalized federated search for energy efficient batterypowered edge computing. Future Generation Computer Systems, 146(??):178–194, September 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23001644.

Yang:2024:EKM

[YZL^{+}24] Zhao Yang, Shengbing Zhang, Chuxi Li, Miao Wang, Haoyang Wang, and Meng Zhang. Efficient knowledge management for heterogeneous federated continual learning on resource-constrained edge devices. Future Generation Computer Systems, 156 (??):16–29, July 2024. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24000633.

Yu:2023:CTS

[YZR23] Xiaofeng Yu, Shunzhi Zhu, and Yongjun Ren. Continuous trajectory similarity search with result diversification. Future Generation Computer Systems, 143(??):392–400, June 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23000523.

Yu:2021:AGR

[YZS^{+}21] Ye Yu, Zhongheng Zhang, Rongju Sun, Haiping Liu, Suwei Yuan, Ting Jiang, Meng Wu, Cheng Guo, Yuelei Guo, Jianchao Weng, Xingdong Zheng, and Feng Yuan. AI-guided resource allocation and rescue decision system for medical applications. Future Generation Computer Systems, 118(??):485–491, May 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20330727.

Yin:2022:CEC

[YZW22] Bo Yin, Weilong Zeng, and Xuetao Wei. Cost-effective crowdsourced join queries for entity resolution without prior knowledge. Future Generation Computer Systems, 127(??):240–251, February 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X2100354X.

Yang:2023:NAK

[YZW^{+}23] Qing Yang, Xiaoqian Zhu, Xiaoliang Wang, Junjie Fu, Jing Zheng, and Yuzhen Liu. A novel authentication and key agreement scheme for Internet of Vehicles. Future Generation Computer Systems, 145(??):415–428, August 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X2300122X.

Yin:2023:CEM

[YZX^{+}23] Bo Yin, Peng Zhang, Binyao Xu, Hang Chen, and Youlin Ji. Crowdenabled multiple Paretooptimal queries for multicriteria decision-making services. Future Generation Computer Systems, 148(??):342–356, November 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23002212.

Yang:2023:HHI

[YZZ^{+}23] Xiaobo Yang, Daosen Zhai, Ruonan Zhang, Haotong Cao, Sahil Garg, and Mohammad Mehedi Hassan. Human-to-human interaction behaviors sensing based on complexvalued neural network using Wi-Fi channel state information. Future Generation Computer Systems, 148(??):160–172, November 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23002145.

Zhang:2020:SIC

[ZA20] Yin Zhang and Haider Abbas. Special issue: Cognitive Internet of Things assisted by cloud computing and big data. Future Generation Computer Systems, 108(??):1214–1216, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X1932014X.

Zolfaghari:2022:MCW

[ZA22] Behrooz Zolfaghari and Saeid Abrishami. A multiclass workflow ensemble management system using on-demand and spot instances in cloud. Future Generation Computer Systems, 137(??):97–110, December 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22002369.

Zahed:2020:PCC

[ZAH^{+}20] M. Ishtiaque A. Zahed, Iftekhar Ahmad, Daryoush Habibi, Quoc Viet Phung, and Md Munjure Mowla. Proactive content caching using surplus renewable energy: a win-win solution for both network service and energy providers. Future Generation Computer Systems, 105(??):210–221, April 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19316061.

Zhang:2022:TBE

[ZBF22] Wenbo Zhang, Yayu Bai, and Jingyu Feng. TIIA: a blockchain-enabled Threat Intelligence Integrity Audit scheme for IIoT. Future Generation Computer Systems, 132(??):254–265, July 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22000723.

Zhou:2023:CBA

[ZBS23] Chen Zhou, Masoud Barati, and Omair Shafiq. A compliance-based architecture for supporting GDPR accountability in cloud computing. Future Generation Computer Systems, 145(??):104–120, August 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23000985.

Zamani:2020:EAA

[ZBTV^{+}20] Ali Reza Zamani, Daniel Balouek-Thomert, J. J. Villalobos, Ivan Rodero, and Manish Parashar. An edgeaware autonomic runtime for data streaming and intransit processing. Future Generation Computer Systems, 110(??):107–118, September 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19307265.

Zaarour:2022:PSD

[ZC22] Tarek Zaarour and Edward

Curry.  SemanticPeer:

a distributional semantic peer-to-peer lookup protocol for large content spaces at Internet-scale. Future Generation Computer Systems, 132(??):239–253, July 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22000590.

Zhang:2021:MPR

[ZCF21] Kun Zhang, Kai Chen, and Binghui Fan. Massive picture retrieval system based on big data image mining. Future Generation Computer Systems, 121(??):54– 58, August 2021. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2100090X.

Zheng:2024:PPP

[ZCL24] Jiali Zheng, Yixin Chen, and Qijia Lai. PPSFL: Privacy-Preserving Split Federated Learning for heterogeneous data in edgebased Internet of Things. Future Generation Computer Systems, 156(??): 231–241, July 2024. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2400092X.

Zhang:2022:TSC

[ZCLL22] Weiwen Zhang, Lei Chen, Jinzhou Luo, and Jianqi Liu. A two-stage container management in the cloud for optimizing the load balancing and migration cost. Future Generation Computer Systems, 135(??):303–314, October 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22001674.

Zeng:2023:MTS

[ZCQ^{+}23] Fanyu Zeng, Mengdong Chen, Cheng Qian, Yanyang Wang, Yijun Zhou, and Wenzhong Tang. Multivariate time series anomaly detection with adversarial transformer architecture in the Internet of Things. Future Generation Computer Systems, 144(??):244–255, July 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23000560.

Zong:2020:ITD

[ZCS20] Wei Zong, Yang-Wai Chow, and Willy Susilo. Interactive three-dimensional visualization of network intrusion detection data for machine learning. Future Generation Computer Systems, 102(??):292–306, January 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18331091.

Zimba:2020:MDM

[ZCWC20] Aaron Zimba, Hongsong Chen, Zhaoshun Wang, and Mumbi Chishimba. Modeling and detection of the multi-stages of Advanced Persistent Threats attacks based on semi-supervised learning and complex networks characteristics. Future Generation Computer Systems, 106(??):501–517, May 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19316267.

Zhai:2022:ARM

[ZDC22] Lijie Zhai, Haisheng Duan, and Donghui Chen. An adaptive recognition method for take-off action images of back-style high jump based on feature extraction. Future Generation Computer Systems, 126(??):65– 69, January 2022. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21002946.

Zheng:2024:BRD

[ZDLD24] Mengke Zheng, Xin Du, Zhihui Lu, and Qiang Duan. A balanced and reliable data replica placement scheme based on reinforcement learning in edge– cloud environments. Future Generation Computer Systems, 155(??):132–145, June 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24000499.

Zong:2021:HLC

[ZDZ21] Xulin Zong, Xiangqian Ding, and Zhen Zhou. Health level classification by fusing medical evaluation from multiple social networks. Future Generation Computer Systems, 114(??):574–580, January 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20317933.

Zavala:2020:HAS

[ZFMB20] Edith Zavala, Xavier Franch, Jordi Marco, and Christian Berger. HAFLoop: an architecture for supporting Highly Adaptive Feedback Loops in selfadaptive systems. Future Generation Computer Systems, 105(??):607–630, April 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19318448.

Zhang:2023:RMD

[ZFZS23] Mengjie Zhang, Renhai Feng, Hehe Zhang, and Yishan Su. A recommendation management defense mechanism based on trust model in underwater acoustic sensor networks. Future Generation Computer Systems, 145(??):466–477, August 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23001310.

Zhang:2023:SNI

[ZG23] Xiaoxia Zhang and Hao Gan. STF-Net: an improved depth network based on spatio-temporal data fusion for PM_{2.5} concentration prediction. Future Generation Computer Systems, 144(??):37–49, July 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X2300064X.

Zhong:2024:LBC

[ZG24] Yue Zhong and Jieming Gu. Lightweight block ciphers for resource-constrained environments: a comprehensive survey. Future Generation Computer Systems, 157(??):288–302, August 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24001298.

Zhang:2022:RVV

[ZGK^{+}22] Peiying Zhang, Peng Gan, Neeraj Kumar, Ching-Hsien Hsu, Shigen Shen, and Shibao Li. RKD-VNE: Virtual network embedding algorithm assisted by resource knowledge description and deep reinforcement learning in IIoT scenario. Future Generation Computer Systems, 135(??):426–437, October 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2200173X.

Zhang:2023:AAA

[ZGL^{+}23] Shaobo Zhang, Tao Guo, Qin Liu, Entao Luo, Kim-Kwang Raymond Choo, and Guojun Wang. ALPS: Achieving accuracy-aware location privacy service via assisted regions. Future Generation Computer Systems, 145(??):189–199, August 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23000997.

Zhao:2020:FCB

[ZGN^{+}20] Chen Zhao, Wu Gao, Feiping Nie, Fei Wang, and Huiyang Zhou. Fair and cache blocking aware warp scheduling for concurrent kernel execution on GPU. Future Generation Computer Systems, 112(??):1093–1105, November 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19306545.

Zhang:2023:HMC

[ZGW^{+}23a] Xi Zhang, Xiaohu Guo, Yue Weng, Xianwei Zhang, Yutong Lu, and Zhong Zhao. Hybrid MPI and CUDA paralleled finite volume unstructured CFD simulations on a multi-GPU system. Future Generation Computer Systems, 139(??):1–16, February 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22002904.

Zhao:2023:FCE

[ZGW^{+}23b] Chen Zhao, Zhipeng Gao, Qian Wang, Kaile Xiao, Zijia Mo, and M. Jamal Deen. FedSup: a communicationefficient federated learning fatigue driving behaviors supervision approach. Future Generation Computer Systems, 138(??):52– 60, January 2023. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22002709.

Zeng:2020:TEE

[ZGY20] Deze Zeng, Lin Gu, and Hong Yao. Towards energy efficient service composition in green energy powered cyber–physical fog systems. Future Generation Computer Systems, 105 (??):757–765, April 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167739X18301961.

Zhao:2021:FPP

[ZGZX21] Shajunyi Zhao, Dongyuan Ge, Jingfeng Zhao, and Wenjiang Xiang. Fingerprint pre-processing and feature engineering to enhance agricultural products categorization. Future Generation Computer Systems, 125(??):944–948, December 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21002594.

Zhu:2020:NDQ

[ZH20] Hongpeng Zhu and TongCheng Huang. A novel deep quality-aware CNN for image edge smoothening. Future Generation Computer Systems, 113(??):468–473, December 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20309961.

Zhang:2020:DAF

[Zha20]  Changhao Zhang.  De-

sign and application of fog computing and Internet of Things service platform for smart city. Future Generation Computer Systems, 112(??):630–640, November 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19331024.

Zhang:2021:CPR

[Zha21] Peng Zhang. E-commerce products recognition based on a deep learning architecture: Theory and implementation. Future Generation Computer Systems, 125(??):672–676, December 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21002533.

Zhang:2020:TFD

[ZHD^{+}20] Chaolong Zhang, Yigang He, Bolun Du, Lifen Yuan, Bing Li, and Shanhe Jiang. Transformer fault diagnosis method using IoT based monitoring system and ensemble machine learning. Future Generation Computer Systems, 108(??): 533–545, July 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19334727.

Zhang:2020:DPO

[ZHGX20] Long Zhang, Jinhua Hu, Chao Guo, and Haitao Xu. Dynamic power optimization for secondary wearable biosensors in ehealthcare leveraging cognitive WBSNs with imperfect spectrum sensing. Future Generation Computer Systems, 112(??):67– 92, November 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19320710.

Zhang:2023:DTP

[ZHH^{+}23] Jing Zhang, Qihan Huang, Yirui Huang, Qian Ding, and Pei-Wei Tsai. DP-TrajGAN: a privacy-aware trajectory generation model with differential privacy. Future Generation Computer Systems, 142(??):25– 40, May 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22004319.

Zou:2020:AHU

[ZHJW20] Shaojun Zou, Jiawei Huang, Wanchun Jiang, and Jianxin Wang. Achieving high utilization of flowlet-based load balancing in data center networks. Future Generation Computer Systems, 108(??):546–559, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19319272.

Zhang:2023:SER

[ZHL^{+}23] Qizhi Zhang, Yale He, Ruilin Lai, Zhihao Hou, and Gansen Zhao. A survey on the efficiency, reliability, and security of data query in blockchain systems. Future Generation Computer Systems, 145(??):303–320, August 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23001309.

Zhu:2024:PMF

[ZHL24] Changpeng Zhu, Bo Han, and Gang Li. PAC: a monitoring framework for performance analysis of compression algorithms in Spark. Future Generation Computer Systems, 157(??):237–249, August 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24000554.

Zhang:2024:MSV

[ZHLL24] Shiwen Zhang, Jiayi He, Wei Liang, and Keqin Li. MMDS: a secure and verifiable multimedia data search scheme for cloudassisted edge computing. Future Generation Computer Systems, 151(??):32– 44, February 2024. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23003564.

Zhu:2020:WEC

[ZHLM20] Huanzhou Zhu, Ligang He, Matthew Leeke, and Rui Mao. WolfGraph: the edge-centric graph processing on GPU. Future Generation Computer Systems, 111(??):552–569, October 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X18325251.

Zhu:2021:FRC

[ZHP^{+}21] Hongbo Zhu, Guangjie Han, Yan Peng, Wenbo Zhang, Chuan Lin, and Hai Zhao. Functional-realistic CT image super-resolution for early-stage pulmonary nodule detection. Future Generation Computer Systems, 115(??):475–485, February 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20301138.

Zhu:2020:IQA

[Zhu20]  Hongpeng Zhu.  Image

quality assessment model based on multi-feature fusion of energy Internet of Things. Future Generation Computer Systems, 112(??):501–506, November 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20300625.

Zhu:2021:CED

[Zhu21] Xu Zhu. Complex event detection for commodity distribution Internet of Things model incorporating radio frequency identification and Wireless Sensor Network. Future Generation Computer Systems, 125(??):100–111, December 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21002193.

Zhao:2020:MOP

[ZHX^{+}20] Jinyuan Zhao, Zhigang Hu, Bing Xiong, Liu Yang, and Keqin Li. Modeling and optimization of packet forwarding performance in software-defined WAN. Future Generation Computer Systems, 106(??):412–425, May 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19315377.

Zhou:2023:CMN

[ZHZS23] Yi Zhou, Lei Huang, Tao Zhou, and Hanshi Sun. Combating medical noisy labels by disentangled distribution learning and consistency regularization. Future Generation Computer Systems, 141(??):567–576, April 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22004228.

Zegarra:2020:VER

[ZIOT^{+}20] Fabian Colque Zegarra, Juan C. Carbajal Ipenza, Behrooz Omidvar-Tehrani, Viviane P. Moreira, Sihem Amer-Yahia, and João L. D. Comba. Visual exploration of rating datasets and user groups. Future Generation Computer Systems, 105(??):547–561, April 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19303309.

Zhan:2022:NQD

[ZJL^{+}22] Jinyu Zhan, Wei Jiang, Ying Li, Junting Wu, Jianping Zhu, and Jinghuan Yu. NIC-QF: a design of FPGA based Network Interface Card with Query Filter for big data systems. Future Generation Computer Systems, 136(??):153–169, November 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X2200200X.

Zhang:2020:RBS

[ZJW^{+}20] Lingling Zhang, Hong Jiang, Fang Wang, Dan Feng, and Yanwen Xie. Reservoir-based sampling over large graph streams to estimate triangle counts and node degrees. Future Generation Computer Systems, 108(??):244–255, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19331577.

Zaman:2021:RRC

[ZKD21] Jesse Zaman, Kennedy Kambona, and Wolfgang De Meuter. A reusable & reconfigurable Citizen Observatory platform. Future Generation Computer Systems, 114(??):195–208, January 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19321648.

Zhao:2020:MCL

[ZKGB20] Zhongliang Zhao, Mostafa Karimzadeh, Florian Gerber, and Torsten Braun. Mobile crowd location prediction with hybrid features using ensemble learning. Future Generation Computer Systems, 110(??):556–571, September 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17318058.

Zhang:2023:EBE

[ZKL^{+}23] Baochen Zhang, Lanju Kong, Qingzhong Li, Xinping Min, Yuan Liu, and Zhengwei Che. EB-BFT: an elastic batched BFT consensus protocol in blockchain. Future Generation Computer Systems, 139(??):267–279, February 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22002989.

Zhang:2021:ARN

[ZL21] Dehua Zhang and Sha Lou. The application research of neural network and BP algorithm in stock price pattern classification and prediction. Future Generation Computer Systems, 115(??):872–879, February 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20329861.

Zhang:2022:IID

[ZL22] Ying Zhang and Qiang Liu. On IoT intrusion detection based on data augmentation for enhancing learning on unbalanced samples. Future Generation Computer Systems, 133(??):213–227, August 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22000826.

Zhang:2023:RLB

[ZL23] Lei Zhang and Panyue Lin. Reinforcement learning based energy-neutral operation for hybrid EH powered TBAN. Future Generation Computer Systems, 140(??):311–320, March 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22003600.

Zhou:2021:DCA

[ZLC^{+}21] Huaman Zhou, Zonghang Li, Qingqing Cai, Hongfang Yu, Shouxi Luo, Long Luo, and Gang Sun. DGT: a contribution-aware differential gradient transmission mechanism for distributed machine learning. Future Generation Computer Systems, 121(??):35–47, August 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21000868.

Zhang:2023:EPP

[ZLF^{+}23a] Chuan Zhang, Xingqi Luo, Qing Fan, Tong Wu, and Liehuang Zhu. Enabling privacy-preserving multiserver collaborative search in smart healthcare. Future Generation Computer Systems, 143(??):265–276, June 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23000304.

Zhang:2023:SBB

[ZLF^{+}23b] Mengxiao Zhang, Jiamou Liu, Kaiyu Feng, Fernando Beltran, and Zi-

jian Zhang. SmartAuction:

a blockchain-based secure implementation of private data queries. Future Generation Computer Systems, 138(??):198–211, January 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22002618.

Zhu:2024:SSR

[ZLG^{+}24] Anqi Zhu, Huimin Lu, Songtao Guo, Zhiwen Zeng, Mingfang Ma, and Zongtan Zhou. SyRoC: Symbiotic robotics for QoSaware heterogeneous applications in IoT-edge-cloud computing paradigm. Future Generation Computer Systems, 150(??):202–219, January 2024. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2300328X.

Zuo:2024:FFC

[ZLH^{+}24] Xiaojiang Zuo, Yaxin Luopan, Rui Han, Qinglong Zhang, Chi Harold Liu, Guoren Wang, and Lydia Y. Chen. FedViT: Federated continual learning of vision transformer at edge. Future Generation Computer Systems, 154(??):1– 15, May 2024. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23004879.

Zheng:2023:PAS

[ZLL^{+}23] Senjiong Zheng, Bo Liu, Weiwei Lin, Xiaoying Ye, and Keqin Li. A packageaware scheduling strategy for edge serverless functions based on multistage optimization. Future Generation Computer Systems, 144(??):105–116, July 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23000547.

Zhang:2024:DRR

[ZLL^{+}24a] Huabing Zhang, Liang Li, Qiong Lu, Yi Yue, Yakun Huang, and Schahram Dustdar. Distributed realtime rendering in decentralized network for mobile web augmented reality. Future Generation Computer Systems, 158(??):530–544, September 2024. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24001924.

Zhang:2024:PTP

[ZLL^{+}24b] Yiguang Zhang, Junxiong Lin, Zhihui Lu, Qiang Duan, and Shih-Chia Huang. PBRL-TChain: a performance-enhanced permissioned blockchain for time-critical applications based on reinforcement learning. Future Generation Computer Systems, 154(??):301–313, May 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23004934.

Zhang:2021:MDF

[ZLLD21] Hao Zhang, Jie-Ling Li, Xi-Meng Liu, and Chen Dong. Multi-dimensional feature fusion and stacking ensemble mechanism for network intrusion detection. Future Generation Computer Systems, 122(??):130–143, September 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2100114X.

Zappatore:2023:SMI

[ZLM^{+}23] Marco Zappatore, Antonella Longo, Angelo Martella, Beniamino Di Martino, Antonio Esposito, and Serena Angela Gracco. Semantic models for IoT sensing to infer environment–wellness relationships. Future Generation Computer Systems, 140(??):1–17, March 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22003211.

Zhu:2020:CIC

[ZLML20] Rongbo Zhu, Lu Liu, Maode Ma, and Hongxiang Li. Cognitive-inspired computing: Advances and novel applications. Future Generation Computer Systems, 109(??):706–709, August 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20308384.

Zhao:2022:USP

[ZLP^{+}22] Lei Zhao, Kenli Li, Bin Pu, Jianguo Chen, Shengli Li, and Xiangke Liao. An ultrasound standard plane detection model of fetal head based on multitask learning and hybrid knowledge graph. Future Generation Computer Systems, 135(??):234–243, October 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22001340.

Zhang:2021:SBB

[ZLPZ21] Yifei Zhang, Senlin Luo, Limin Pan, and Hanqing Zhang. Syscall–BSEM: Behavioral semantics enhancement method of system call sequence for high accurate and robust host intrusion detection. Future Generation Computer Systems, 125(??):112–126, December 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21002259.

Zhang:2023:EEP

[ZLQ23] Peng Zhang, Brian Lee, and Yuansong Qiao. Experimental evaluation of the performance of Gpipe parallelism. Future Generation Computer Systems, 147(??):107–118, October 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23001735.

Zhu:2020:CTB

[ZLS^{+}20] Liehuang Zhu, Nassoro M. R. Lwamo, Kashif Sharif, Chang Xu, Xiaojiang Du, Mohsen Guizani, and Fan Li. T-CAM: Timebased content access control mechanism for ICN subscription systems. Future Generation Computer Systems, 106(??):607–621, May 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19307770.

Zhang:2022:TAM

[ZLS^{+}22a] Jixian Zhang, Wenlu Lou, Hao Sun, Qian Su, and Weidong Li. Truthful auction mechanisms for resource allocation in the Internet of Vehicles with public blockchain networks. Future Generation Computer Systems, 132(??):11– 24, July 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22000450.

Zhang:2022:ODT

[ZLS22b] Lanyong Zhang, Chengyu Li, and Hongfang Sun. Object detection/tracking toward underwater photographs by remotely operated vehicles (ROVs). Future Generation Computer Systems, 126(??):163–168, January 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21002727.

Zhang:2023:BDA

[ZLS23] Zicheng Zhang, Xinyue Lin, and Shaonan Shan. Big data-assisted urban governance: an intelligent real-time monitoring and early warning system for public opinion in government hotline. Future Generation Computer Systems, 144(??):90–104, July 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2300081X.

Zhao:2023:DDB

[ZLST23] Dan Zhao, Hao Li, Xiuwen Sun, and Yazhe Tang. Detecting DGA-based botnets through effective phonicsbased features. Future Generation Computer Systems, 143(??):105–117, June 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23000341.

Lu:2020:BSS

[zLsZjX20] Chao ze Lu, Guo sun Zeng, and Ying jie Xie. Bigraph specification of software architecture and evolution analysis in mobile computing environment. Future Generation Computer Systems, 108(??):662–676, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19320011.

Zhang:2022:MOP

[ZLW^{+}22] Yongde Zhang, Fagui Liu, Bin Wang, Weiwei Lin, Guoxiang Zhong, Minxian Xu, and Keqin Li. A multioutput prediction model for physical machine resource usage in cloud data centers. Future Generation Computer Systems, 130(??): 292–306, May 2022. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22000103.

Zhuang:2024:EHF

[ZLW^{+}24] Hongbin Zhuang, Xiao-Yan Li, Dajin Wang, Cheng-Kuan Lin, and Kun Zhao. Enabling high faulttolerant embedding capability of alternating group graphs. Future Generation Computer Systems, 158(??):110–121, September 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24001481.

Zuo:2023:TAP

[ZLWH23] Fang Zuo, Yajie Li, Guanghui Wang, and Xin He. Towards accurate and privacypreserving localization using anchor quality assessment in Internet of Things. Future Generation Computer Systems, 148(??):524–537, November 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23002443.

Zhang:2020:PCE

[ZLXH20] Jing Zhang, Minhao Lou, Lin Xiang, and Long Hu. Power cognition: Enabling intelligent energy harvesting and resource allocation for solar-powered UAVs. Future Generation Computer Systems, 110(??):658–664, September 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19308349.

Zhang:2020:SMP

[ZLZ^{+}20a] Yi Zhang, Yu Liu, Junlong Zhou, Jin Sun, and Keqin Li. Slow-movement particle swarm optimization algorithms for scheduling security-critical tasks in resource-limited mobile edge computing. Future Generation Computer Systems, 112(??):148–161, November 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19333552.

Zhao:2020:MVF

[ZLZ^{+}20b] Feng Zhao, Jing Li, Lu Zhang, Zhe Li, and Sang-Gyun Na. Multi-view face recognition using deep neural networks. Future Generation Computer Systems, 111(??):375–380, October 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20300972.

Zhang:2021:PFV

[ZLZ21] Xialin Zhang, Lingkun Lian, and Fukang Zhu. Parameter fitting of variogram based on hybrid algorithm of particle swarm and artificial fish swarm. Future Generation Computer Systems, 116(??):265–274, March 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20320768.

Zhang:2023:DRL

[ZLZ23a] Tinghao Zhang, Kwok-Yan Lam, and Jun Zhao. Deep reinforcement learning based scheduling strategy for federated learning in sensor-cloud systems. Future Generation Computer Systems, 144(??):219–229, July 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23000870.

Zhang:2023:MAD

[ZLZ^{+}23b] Yaqiang Zhang, Ruyang Li, Yaqian Zhao, Rengang Li, Yanwei Wang, and Zhangbing Zhou. Multiagent deep reinforcement learning for online request scheduling in edge cooperation networks. Future Generation Computer Systems, 141(??):258–268, April 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22003788.

Zhiqiang:2022:IDW

[ZMJ^{+}22] Liu Zhiqiang, Ghulam Mohiuddin, Zheng Jiangbin, Muhammad Asim, and Wang Sifei. Intrusion detection in wireless sensor network using enhanced empirical based component analysis. Future Generation Computer Systems, 135(??):181–193, October 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22001534.

Zhang:2019:SAP

[ZMZ^{+}19] Xingwei Zhang, Rui Mo, Huijun Zhao, Xi Luo, and Yunsheng Yang. Statistical analysis of photodynamic therapy and stent drainage for unresectable cholangiocarcinoma. Future Generation Computer Systems, 91(??):511–517, February 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0167739X18318296.

See retraction notice [ZMZ^{+}20].

Zhang:2020:RNS

[ZMZ^{+}20] Xingwei Zhang, Rui Mo, Huijun Zhao, Xi Luo, and Yunsheng Yang. Retraction notice to “Statistical analysis of photodynamic therapy and stent drainage for unresectable cholangiocarcinoma” [Future Gener. Comput. Syst. 91 (2019) 511–517]. Future Generation Computer Systems, 107(??):1145, June 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20306397.

See [ZMZ^{+}19].

Zhan:2021:MTG

[ZN21] Kun Zhan and Chaoxi Niu. Mutual teaching for graph convolutional networks. Future Generation Computer Systems, 115(??):837–843, February 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20329939.

Zhang:2023:TSF

[ZNX23] Jiangjiang Zhang, Zhenhu Ning, and Fei Xue. A two-stage federated optimization algorithm for privacy computing in Internet of Things. Future Generation Computer Systems, 145(??):354–366, August 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23001279.

Zhang:2023:NMV

[ZNZ^{+}23] De-Gan Zhang, Chen-Hao Ni, Jie Zhang, Ting Zhang, and Zhi-Hao Zhang. New method of vehicle cooperative communication based on fuzzy logic and signaling game strategy. Future Generation Computer Systems, 142(??):131–149, May 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22004435.

Zhang:2022:PKM

[ZP22] Haixia Zhang and Qingxiu Peng. PSO and kmeans-based semantic segmentation toward agricultural products. Future Generation Computer Systems, 126(??):82–87, January 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21002545.

Zhao:2023:MOC

[ZPK^{+}23] Yang Zhao, Yuwei Pang, Xingyu Ke, Bintao Wang, Guobin Zhu, and Mingsheng Cao. A metaverseoriented CP-ABE scheme with cryptographic reverse firewall. Future Generation Computer Systems, 147(??):195–206, October 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23001656.

Zhang:2020:BRC

[ZPLQ20] Ronghui Zhang, Tao Peng, Zhihan Lv, and Zhijun Qiu. Bifurcation and robust control analysis to tractorsemitrailer with interference on rainy slippery road. Future Generation Computer Systems, 107(??): 126–143, June 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19305497.

Zhang:2021:PPQ

[ZPQH21] Yiwen Zhang, Jie Pan, Lianyong Qi, and Qiang He. Privacy-preserving quality prediction for edgebased IoT services. Future Generation Computer Systems, 114(??):336–348, January 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20312668.

Zhou:2024:PCU

[ZPS^{+}24] Erqiang Zhou, Yejian Peng, Guanghui Shao, Fuhu Deng, Yurun Miao, and Wulong Fan. Password cracking using chunk similarity. Future Generation Computer Systems, 150(??):380–394, January 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23003382.

Zhang:2023:IRC

[ZRH^{+}23] Zhi-Yuan Zhang, Hao Ren, Zhenli He, Wei Zhou, and Di Liu. Improving robustness of convolutional neural networks using elementwise activation scaling. Future Generation Computer Systems, 149(??):136–149, December 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23002637.

Zhang:2023:HSL

[ZrHhH^{+}23] Jing Zhang, Yi rui Huang, Qi han Huang, Yan zi Li, and Xiu cai Ye. Hasse sensitivity level: a sensitivityaware trajectory privacyenhanced framework with reinforcement learning. Future Generation Computer Systems, 142(??):301–313, May 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23000158.

Zhang:2023:MSA

[ZSL^{+}23a] Tao Zhang, Jian Shen, Chin-Feng Lai, Sai Ji, and Yongjun Ren. Multiserver assisted data sharing supporting secure deduplication for metaverse healthcare systems. Future Generation Computer Systems, 140(??):299–310, March 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22003545.

Zhou:2023:PPL

[ZSL^{+}23b] Yousheng Zhou, Liyuan Song, Yuanni Liu, Pandi Vijayakumar, Brij B. Gupta, Wadee Alhalabi, and Hind Alsharif. A privacypreserving logistic regressionbased diagnosis scheme for digital healthcare. Future Generation Computer Systems, 144(??):63–73, July 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23000638.

Zhao:2020:APE

[ZST^{+}20] Linchang Zhao, Zhaowei Shang, Jin Tan, Xiaoliu Luo, Taiping Zhang, Yu Wei, and Yuan Yan Tang. Adaptive parameter estimation of GMM and its application in clustering. Future Generation Computer Systems, 106(??):

250–259, May 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19322484.

Zhang:2024:PIS

[ZSZ^{+}24] Qian Zhang, Haoyun Song, Kaiyan Zhou, Jianhao Wei, and Chuqiao Xiao. A prefetching indexing scheme for in-memory database systems. Future Generation Computer Systems, 156(??):179–190, July 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24000840.

Zhang:2024:EFI

[ZSZX24] Kunpeng Zhang, Edwin Hsing-Mean Sha, Qingfeng Zhuge, and Rui Xu. An efficient flattened index structure with lazy restructuring and hotness awareness. Future Generation Computer Systems, 153(??):139–153, April 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23004338.

Zhang:2022:HMD

[ZT22a] Yifan Zhang and Peter J.

Thorburn. Handling missing data in near realtime environmental monitoring: a system and a review of selected methods. Future Generation Computer Systems, 128(??):63– 72, March 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21003794.

Zhang:2022:PWR

[ZT22b] Yuran Zhang and Ziyan Tang. PSO-weighted random forest for attractive tourism spots recommendation. Future Generation Computer Systems, 127(??):421–425, February 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21003757.

Zhou:2023:MSR

[ZTB23] Guangyao Zhou, Wenhong Tian, and Rajkumar Buyya. Multi-search-routes-based methods for minimizing makespan of homogeneous and heterogeneous resources in cloud computing. Future Generation Computer Systems, 141(??):414–432, April 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22004022.

Zhang:2020:ECP

[ZTC20] Ying Zhang, Guohui Tian, and Huanzhao Chen. Exploring the cognitive process for service task in smart home: a robot service mechanism. Future Generation Computer Systems, 102(??):588–602, January 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19311197.

Zhao:2020:EFS

[ZTP20] Zhiming Zhao, Ian Taylor, and Radu Prodan. Editorial for FGCS special issue on “Time-critical applications on software-defined infrastructures”. Future Generation Computer Systems, 112(??):1170–1171, November 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20324146.

Zhou:2020:EIE

[ZTQ^{+}20] Yi Zhou, Shubbhi Taneja, Xiao Qin, Wei-Shinn Ku, and Jifu Zhang. EDOM: Improving energy efficiency of database operations on multicore servers. Future Generation Computer Systems, 105(??):1002–1015, April 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X17303023.

Zhai:2022:DNN

[ZWC^{+}22] Daosen Zhai, Chen Wang, Haotong Cao, Sahil Garg, Mohammad Mehedi Hassan, and Salman A. AlQahtani. Deep neural network based UAV deployment and dynamic power control for 6G-Envisioned intelligent warehouse logistics system. Future Generation Computer Systems, 137(??):164–172, December 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22002382.

Zhu:2023:MCP

[ZWCS23] Xi Zhu, Junbo Wang, Wuhui Chen, and Kento Sato. Model compression and privacy preserving framework for federated learning. Future Generation Computer Systems, 140(??):376–389, March 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22003491.

Zhang:2020:LLD

[ZWH^{+}20] Wenbo Zhang, Zonglin Wu, Guangjie Han, Yongxin Feng, and Lei Shu. LDC: a lightweight data consensus algorithm based on the blockchain for the industrial Internet of Things for smart city applications. Future Generation Computer Systems, 108(??):574–582, July 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19333503.

Zhang:2021:EFS

[ZWH21a] Yan Zhang, Kun Wang, and Lei He. Editorial for FGCS special issue: Computation intelligence for energy Internet. Future Generation Computer Systems, 115(??):295–297, February 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20327631.

Zhao:2021:SNS

[ZWH^{+}21b] Meng Zhao, Hao Wang, Ying Han, Xiaokang Wang, Hong-Ning Dai, Xuguo Sun, Jin Zhang, and Marius Pedersen. SEENS: Nuclei segmentation in Pap smear images with selective edge enhancement. Future Generation Computer Systems, 114(??):185–194, January 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20304271.

Zhang:2020:NOA

[ZWL20] Yulai Zhang, Yuchao Wang, and Guiming Luo. A new optimization algorithm for non-stationary time series prediction based on recurrent neural networks. Future Generation Computer Systems, 102(??):738–745, January 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X18332540.

Zheng:2021:AER

[ZWL21] Yi Zheng, Youqiang Wang, and Jixin Liu. Analysis and experimental research on stability characteristics of squatting posture of wearable lower limb exoskeleton robot. Future Generation Computer Systems, 125(??):352–363, December 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2100248X.

Zheng:2022:RSO

[ZWL22] Yi Zheng, Youqiang Wang, and Jixin Liu. Research on structure optimization and motion characteristics of wearable medical robotics based on Improved Particle Swarm Optimization Algorithm. Future Generation Computer Systems, 129(??):187–198, April 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21004581.

Zhang:2023:DLW

[ZWM^{+}23] Hao Zhang, Tingting Wu, Zhifeng Ma, Feng Li, and Jie Liu. Dynamic layerwise sparsification for distributed deep learning. Future Generation Computer Systems, 147(??):1–15, October 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23001620.

Zhang:2020:NGU

[ZWW^{+}20a] Jindan Zhang, Baocang Wang, Xu An Wang, Han Wang, and Shuai Xiao. New group user based privacy preserving cloud auditing protocol. Future Generation Computer Systems, 106(??):585–594, May 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19321922.

Zhou:2020:TBC

[ZWW^{+}20b] Zhangbing Zhou, Jinfeng Wen, Yasha Wang, Xiao Xue, Patrick C. K. Hung, and Long D. Nguyen. Topic-based crossing-workflow fragment discovery. Future Generation Computer Systems, 112(??):1141–1155, November 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20301084.

Zhang:2023:SGS

[ZWW^{+}23] Qinglin Zhang, Menghan Wang, Haiyan Wang, Xuan Rao, and Lisi Chen. SSAR-GNN: Self-Supervised Artist Recommendation from spatiotemporal perspectives in art history with Graph Neural Networks. Future Generation Computer Systems, 144(??):230–241, July 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23000808.

Zhu:2021:EIB

[ZWWC21] Hongfei Zhu, Ye Wang, Chonghua Wang, and Xiaochun Cheng. An efficient identity-based proxy signcryption using lattice. Future Generation Computer Systems, 117(??):321–327, April 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20330582.

Zhang:2023:RMB

[ZWX^{+}23] Bowei Zhang, Xiaoliang Wang, Ru Xie, Chuncao Li, Huazheng Zhang, and Frank Jiang. A reputation mechanism based Deep Reinforcement Learning and blockchain to suppress selfish node attack motivation in Vehicular Ad-Hoc Network. Future Generation Computer Systems, 139(??):17–28, February 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22002941.

Zhou:2021:BBD

[ZWY^{+}21] Zhili Zhou, Meimin Wang, Ching-Nung Yang, Zhangjie Fu, Xingming Sun, and Q. M. Jonathan Wu. Blockchain-based decentralized reputation system in e-commerce environment. Future Generation Computer Systems, 124(??):155–167, November 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21001850.

Zhao:2021:POL

[ZWZ^{+}21] Yong Zhao, Xingwei Wang, Chuangchuang Zhang, Qiang He, and Min Huang. Power optimization with less state transition for green software defined networking. Future Generation Computer Systems, 114(??):69– 81, January 2021. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20310578.

Zhang:2023:BBS

[ZWZ^{+}23] Feng Zhang, Hao Wang, Lu Zhou, Dequan Xu, and Liang Liu. A blockchainbased security and trust mechanism for AI-enabled IIoT systems. Future Generation Computer Systems, 146(??):78–85, September 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23000882.

Zeng:2024:JOM

[ZWZ^{+}24] Chao Zeng, Xingwei Wang, Rongfei Zeng, Ying Li, Jianzhi Shi, and Min Huang. Joint optimization of multi-dimensional resource allocation and task offloading for QoE enhancement in Cloud–Edge– End collaboration. Future Generation Computer Systems, 155(??):121–131, June 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24000311.

Zhang:2024:SCR

[ZWZB24] Hang Zhang, Jinsong Wang, Hongwei Zhang, and Chao Bu. Security computing resource allocation based on deep reinforcement learning in serverless multicloud edge computing. Future Generation Computer Systems, 151(??):152–161, February 2024. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23003461.

Zheng:2020:OSC

[ZXD^{+}20] Zibin Zheng, Shaoan Xie, Hong-Ning Dai, Weili Chen, Xiangping Chen, Jian Weng, and Muhammad Imran. An overview on smart contracts: Challenges, advances and platforms. Future Generation Computer Systems, 105 (??):475–491, April 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0167739X19316280.

Zhu:2020:DPM

[ZXL^{+}20] Tianqing Zhu, Ping Xiong, Gang Li, Wanlei Zhou, and Philip S. Yu. Differentially private model publishing in cyber physical systems. Future Generation Computer Systems, 108(??): 1297–1306, July 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17325554.

Zhao:2020:ICI

[ZXW^{+}20] Xin Zhao, Wei Xiao, Lu Wu, Zhigang Zhao, Jidong Huo, Shi Wang, Zhenhua Guo, and Dianmin Sun. Intelligent city intelligent medical sharing technology based on Internet of Things technology. Future Generation Computer Systems, 111(??):226–233, October 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20300170.

Zhang:2020:RCU

[ZXX^{+}20] Bin Zhang, Wentao Xiao, Xi Xiao, Arun Kumar Sangaiah, Weizhe Zhang, and Jiajia Zhang. Ransomware classification using patch-based CNN and self-attention network on embedded n-grams of opcodes. Future Generation Computer Systems, 110(??):708–720, September 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19315912.

Zhang:2023:LDN

[ZXX23] Ziming Zhang, Xiaolong Xu, and Fu Xiao. LGAN-DP: a novel differential private publication mechanism of trajectory data. Future Generation Computer Systems, 141(??):692–703, April 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22004162.

Zhang:2021:NWS

[ZXY^{+}21] Dehai Zhang, Xiaoqiang Xia, Yun Yang, Po Yang, Cheng Xie, Menglong Cui, and Qing Liu. A novel word similarity measure method for IoT-enabled healthcare applications. Future Generation Computer Systems, 114(??):209–218, January 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19325440.

Zhu:2023:EES

[ZXZ^{+}23] Liehuang Zhu, Yumeng Xie, Yuao Zhou, Qing Fan, Chuan Zhang, and Ximeng Liu. Enabling efficient and secure health data sharing for Healthcare IoT systems. Future Generation Computer Systems, 149(??):304–316, December 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23002881.

Zhang:2020:IIF

[ZY20] Diming Zhang and Shaodi You. iFlask: Isolate flask security system from dangerous execution environment by using ARM TrustZone. Future Generation Computer Systems, 109(??):531–537, August 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X17322239.

Zhou:2021:RPD

[ZY21] Hongzhi Zhou and Gan Yu. Research on pedestrian detection technology based on the SVM classifier trained by HOG and LTP features. Future Generation Computer Systems, 125(??):604–615, December 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21002119.

Zhang:2022:EED

[ZYF^{+}22] Yujin Zhang, Luo Yu, Zhijun Fang, Neal N. Xiong, Lijun Zhang, and Haiyue Tian. An end-to-end deep learning model for robust smooth filtering identification. Future Generation Computer Systems, 127(??):263–275, February 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21003423.

Zeng:2020:SCP

[ZYL^{+}20] Jing Zeng, Laurence T.

Yang, Man Lin, Huansheng Ning, and Jianhua Ma. A survey: Cyber– physical–social systems and their system-level design methodology. Future Generation Computer Systems, 105(??):1028–1042, April 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X1630228X.

Zhu:2022:ICF

[ZYL^{+}22] Min Zhu, Han Yu, Zhiyuan Liu, Bingqing Shen, Lihong Jiang, and Hongming Cai. An intelligent collaboration framework of IoT applications based on event logic graph. Future Generation Computer Systems, 137(??):31–41, December 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X2200228X.

Zhu:2024:LAT

[ZYW24] Xikang Zhu, Wenbin Yao, and Wenhao Wang. Loadaware task migration algorithm toward adaptive load balancing in Edge Computing. Future Generation Computer Systems, 157(??):303–312, August 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X24000864.

Zhang:2020:OAM

[ZYX^{+}20] Jixian Zhang, Xutao Yang, Ning Xie, Xuejie Zhang, Athanasios V. Vasilakos, and Weidong Li. An online auction mechanism for time-varying multidimensional resource allocation in clouds. Future Generation Computer Systems, 111(??):27–38, October 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X19331127.

Zeng:2023:IEE

[ZYX^{+}23] Xuanwei Zeng, Yong Yang, Qiaoqiao Xu, Huimiao Zhan, Haoan Lv, Zhiqiang Zhou, Xin Ma, Xiaojuan Liu, Jiaojiao Gui, Qianruo Kang, Neal Xiong, Junfeng Gao, and Hua Zheng. Intraoperative enhancement of effective connectivity in the default mode network predicts postoperative delirium following cardiovascular surgery. Future Generation Computer Systems, 145(??):27–37, August 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23000845.

Zhang:2023:SLM

[ZYY^{+}23] Weisheng Zhang, Zhibang Yang, Shenghong Yang, Mingxing Duan, and Kenli Li. SPsync: Lightweight multi-terminal big spatiotemporal data synchronization solution. Future Generation Computer Systems, 141(??):106–115, April 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X2200365X.

Zhang:2021:HQF

[ZZ21a] Long Zhang and Lin Zhao. High-quality face image generation using particle swarm optimization-based generative adversarial networks. Future Generation Computer Systems, 122(??):98–104, September 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21001126.

Zhang:2021:PVP

[ZZ21b] Ruo-Xi Zhang and Le-Min Zhang. Panoramic visual perception and identification of architectural cityscape elements in a virtual-reality environment. Future Generation Computer Systems, 118(??): 107–117, May 2021. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X20330843.

Zhang:2024:TST

[ZZ24] Xiaoxia Zhang and Pengcheng Zhou. A transferred spatio-temporal deep model based on multi-LSTM autoencoder for air pollution time series missing value imputation. Future Generation Computer Systems, 156(??):325–338, July 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24000839.

Zhang:2022:IFM

[ZZB^{+}22] Guoming Zhang, Xuyun Zhang, Muhammad Bilal, Wanchun Dou, Xiaolong Xu, and Joel J. P. C. Rodrigues. Identifying fraud in medical insurance based on blockchain and deep learning. Future Generation Computer Systems, 130(??):140–154, May 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X21004878.

Zhang:2024:CFH

[ZZB^{+}24] Zao Zhang, Yuning Zhang, Wei Bao, Changyang Li, and Dong Yuan. Coarseto-fine: a hierarchical DNN inference framework for edge computing. Future Generation Computer Systems, 157(??):180–192, August 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X24000736.

Zhao:2022:AIP

[ZZD22] Zezheng Zhao, Jianyin Zhou, and Huadong Du. Artificial intelligence powered forecast of oceanic mesoscale phenomena: a typhoon cold wake case occurring in Northwest Pacific Ocean. Future Generation Computer Systems, 129 (??):389–398, April 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0167739X21004271.

Zhang:2024:FAR

[ZZF^{+}24] Yucheng Zhang, Wenxuan Zhu, Dan Feng, Wei Huang, Nan Jiang, Meng Chen, and Renxin Xia. A fragmentationaware redundancy elimination scheme for inline backup systems. Future Generation Computer Systems, 156(??):53–63, July 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X2400075X.

Zhang:2022:SRD

[ZZG^{+}22] Zhao Zhang, Yong Zhang, Da Guo, Lei Yao, and Zhao Li. SecFedNIDS: Robust defense for poisoning attack against federated learningbased network intrusion detection system. Future Generation Computer Systems, 134(??):154–169, September 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X22001339.

Zhang:2024:BBS

[ZZG^{+}24] Yi Zhang, Peiying Zhang, Mohsen Guizani, Jianyong Zhang, Jian Wang, Hailong Zhu, Kostromitin Konstantin Igorevich, and Huiling Shi. Blockchain-based secure communication of Internet of Things in space– air–ground integrated network. Future Generation Computer Systems, 158(??):391–399, September 2024. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X24001559.

Zhang:2021:CAC

[ZZJC21] Tao Zhang, Yan Zhao, Wenjing Jia, and Mu-Yen Chen. Collaborative algorithms that combine AI with IoT towards monitoring and control system. Future Generation Computer Systems, 125(??):677–686, December 2021. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21002697.

Zhang:2022:LBT

[ZZL^{+}22] Tao Zhang, Qianqiang Zhang, Yasi Lei, Shaojun Zou, Juan Huang, and Fangmin Li. Load balancing with traffic isolation in data center networks. Future Generation Computer Systems, 127(??):126–141, February 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X2100340X.

Zhang:2021:CMA

[ZZLF21] Yongnan Zhang, Yonghua Zhou, Huapu Lu, and Hamido Fujita. Cooperative multi-agent actor– critic control of traffic network flow based on edge computing. Future Generation Computer Systems, 123(??):128–141, October 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21001394.

Zhang:2023:PPA

[ZZP^{+}23] Xiangying Zhang, Pai Zheng, Tao Peng, Dai Li, Xujun Zhang, and Renzhong Tang. Privacypreserving activity recognition using multimodal sensors in smart office. Future Generation Computer Systems, 148(??):27– 38, November 2023. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X23002042.

Zhuang:2021:UDB

[ZZPK21] Xu Zhuang, Yan Zhu, Qiang Peng, and Faisal Khurshid. Using deep belief network to demote web spam. Future Generation Computer Systems, 118(??):94–106, May 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20330855.

Zhang:2021:PHP

[ZZQ21] Zhishuo Zhang, Wei Zhang, and Zhiguang Qin. A partially hidden policy CP– ABE scheme against attribute values guessing attacks with online privacyprotective decryption testing in IoT assisted cloud computing. Future Generation Computer Systems, 123(??):181–195, October 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21001436.

Zou:2022:TCM

[ZZT^{+}22] Qiang Zou, Yifeng Zhu, Yujuan Tan, Yuhui Deng, and Wei Chen. Temporal characterization of memory access behaviors in SPEC CPU2017 workloads: Analysis and synthesis. Future Generation Computer Systems, 130 (??):33–45, May 2022. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21004908.

Zhang:2020:SCS

[ZZXH20] Quanxin Zhang, Xiaosong Zhang, Yuan Xue, and Jingjing Hu. A stealthy covert storage channel for asymmetric surveillance VoLTE endpoints. Future Generation Computer Systems, 102(??):472–480, January 2020. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X19317303.

Zhong:2021:EOS

[ZZZ^{+}21a] Hong Zhong, Yiyuan Zhou, Qingyang Zhang, Yan Xu, and Jie Cui. An efficient and outsourcingsupported attribute-based access control scheme for edge-enabled smart healthcare. Future Generation Computer Systems, 115(??):486–496, February 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X2030100X.

Zhou:2021:AMV

[ZZZ21b] Jianhang Zhou, Qi Zhang, and Bob Zhang. An automatic multi-view disease detection system via collective deep region-based feature representation. Future Generation Computer Systems, 115(??):59–75, February 2021. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167739X20303897.

Zhang:2022:SOR

[ZZZ^{+}22] Hongfei Zhang, Li Zhu, Liwen Zhang, Tao Dai, Xi Feng, Li Zhang, Kaiqi Zhang, and Yutian Yan. Smart objects recommendation based on pretraining with attention and the thing–thing relationship in social Internet of Things. Future Generation Computer Systems, 129(??):347–357, April 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X21004350.

Zhang:2023:DTL

[ZZZ^{+}23] Dong Zhang, Heye Zhang, Hongwei Zhang, Lei Xu, Jinglin Zhang, and Zhifan Gao. Distance transform learning for structural and functional analysis of coronary artery from dualview angiography. Future Generation Computer Systems, 145(??):136–149, August 2023. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X23000833.

Zhou:2022:FSB

[ZZZX22] Lu Zhou, Ye Zhu, Tianrui Zong, and Yong Xiang. A feature selectionbased method for DDoS attack flow classification. Future Generation Computer Systems, 132(??):67– 79, July 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0167739X22000474.